A Complete Bibliography of Publications in Dædalus: proceedings of the American Academy of Arts and

Sciences

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

23 May 2024 Version 1.27

Title word cross-reference

250^{∘} [3]. 355^{∘} [3].

&c [2984].

-abilitie [3243].

10 [4199]. 10th [695]. 1780 [3020]. 19 [4131]. 1940s [2561]. 1950s [2561, 1378]. 1956 [11]. 1960s [2564]. 1964 [470]. 1965 [556, 636]. 1966 [663]. 1967 [668]. 1970s [3929]. 1980s [1827, 1581, 1561, 1558, 1572, 1786, 1596, 1525]. 1990s [2182, 2247].

2000 [2875]. 21st [3627, 3568, 3608, 4128, 3606]. 22 [356].

3 [4094].

400 [3020].

’70 [1030]. 700 [2875].

9 [11].

= [360].

A-Bomb [2697]. A.D [2875, 3020]. Abandoned [610]. Abatement [678]. abilitie [3243]. Abortion [3584]. Above [3447]. Abraham [1087]. Abroad [2926, 779, 2783]. Abundance [1672]. Abuse [2634, 2922, 904, 2011, 2273]. Abuses [594]. Aby [1277]. Abyss [3111].

Academe [1181]. Academia [1157]. Academic [2611, 3951, 2555, 2598, 3082, 842, 2551, 2596, 2731, 1194, 2602, 2606, 2607, 2597, 2600, 1118, 869, 1180, 1424, 2599, 91, 2608, 1202, 930, 1129, 3363]. Academy [556, 3311, 3330, 3367, 3387, 3400, 4010, 93, 1461, 13]. Accept [1949]. Acceptable [1021, 2013]. Acceptance [3594]. Access [3909, 1894, 1967, 3901, 3922, 3907, 3685, 3915, 2055, 3680, 3911, 3905, 3914]. Access-to-Justice [3907]. Accommodation [2422, 985, 3547, 441, 2793]. Accomplishing [3871]. Accountability [2213, 3334, 3876]. Accounts [324]. Accra [4059]. Accused [1439, 2019]. Accuser [1439, 2019]. Achievement [3024, 777, 3497]. Achievements [2499, 1772]. Achieving [1712, 3945]. Acknowledgment [2164]. Acquisition [2637]. Acronyms [1824, 1837]. across [4210, 3949]. Act [3535, 4167, 321]. Action [1030, 3587, 3786, 2838, 3499, 1939, 3994, 4040]. Actionism [1205]. Active [1285, 3707]. Activism [605, 4010, 699, 3671]. Activists [367]. Activities [1929, 42]. Activity [515, 3314]. Actor [386, 4181]. Acts [2338]. Ad [157]. Adam [347]. Adams [1081, 1081]. Adaptation [4058, 3720, 2683, 3032, 3731, 619, 4069]. Adapted [2043, 1951]. Adapting [3976]. Adaptive [2238, 871]. Addams [442]. Added [1871, 3102]. Addendum [2589]. Address [3561, 3942, 1846, 4216, 470]. Addressing [4202]. Adequate [1879, 3223]. Adjudication [222]. Adjust [848]. Administrative [4048, 4044, 4045, 49, 4053, 4046, 286, 4055, 2822, 4043, 4042, 4052]. Administrator [306]. Ado [1985]. Adolescence [945, 953, 941, 943, 946, 952]. Adolescent [947, 944, 948, 949]. Adult [744, 1285, 1323, 3718, 1778]. Adulthood [1286, 1282, 1318, 1317, 1284, 1287, 1321, 2015, 1291, 1289, 1290, 1320, 1288].

Adultlike [1778]. Advancement

[792, 1906, 22]. Advancing [4076, 3884]. Adventures [983, 2509]. Adversarial [3816]. Adversaries [986]. Adversary [3896]. Advertisement [157]. Aerospace [2229]. Aesthetic [163, 680]. Aesthetics [1861, 3130, 2708]. Affairs [1508, 3512, 2049]. Affecting [519]. Affections [153]. Affects [2682]. Affinity [3247]. Affirmation [2683]. Affirmative [3499, 4040]. Afraid [2211, 2787]. Africa [1642, 2800, 1100, 1654, 2213, 2808, 1651, 2804, 623, 1655, 2791, 536, 3824, 2811, 4056, 2809, 1650, 4068, 4067, 4060, 2541, 2050, 1645, 3933, 4065, 2803, 1102, 3692, 2489, 2801, 1649, 2810, 2806, 916, 4069, 4064, 2802, 1744, 2448, 2805, 2086]. African [2575, 2421, 3508, 1098, 2419, 3016, 1652, 2450, 1653, 3498, 624, 1648, 3481, 697, 571, 1644, 1101, 2812, 3506]. Afrikaner [2802]. Afrikaners [2575]. Afro [1099, 614]. Afrofuturism [3926]. After [2377, 2386, 1209, 2388, 2819, 2118, 126, 1231, 1234, 3473, 2226, 1642, 1182, 2260, 2895, 2281, 3784, 2357]. Aftermath [4223, 4077, 365]. Aftermaths [4207]. Afterword [4109]. Against [3776, 3862, 2978, 2856, 3965, 4132, 3819, 339, 345, 995, 1059, 979, 617, 456, 3459, 888]. Agassiz [130]. Age [1760, 1456, 2530, 81, 353, 1668, 3766, 2424, 1629, 2945, 2854, 3369, 1318, 4056, 2446, 2245, 1778, 1870, 1337, 1214, 753, 1796, 2544, 3433, 3439, 3477, 3483, 3359, 3501]. Aged [942]. Agencies [804]. Agency [3835, 1257]. Agenda [1548, 2295, 3934, 2247, 2755, 2710]. Agent [4090]. Ages [1078]. Aggregate [3684]. Aging [1869, 3144, 3720, 3723, 3170, 1879, 3721, 1878, 3141, 3719, 1867, 1883, 3143, 1873, 1876, 1880, 1870, 1881, 3138, 1884, 1875, 1453, 1874, 3716, 3140, 1882, 3724]. Agonism [3615]. Agonistic [1157]. Agreed [1849]. Agreements [214]. Agriculture [677, 2664, 1113, 2148, 3741, 421, 3736]. Aha [3860]. Ahead [3071, 3965, 4008, 3134]. AI [4103, 4096, 4083, 4108, 1974, 4102, 4100]. AI-Assisted [4103]. Aid [2040, 3921, 3919, 3920, 4155]. Aiding [4009]. AIDS [2065, 2077, 2082, 1943, 2072, 2087, 2073, 2071, 2069, 2067, 2079, 2083, 2074, 2084, 2088, 2070, 2068, 2066, 2085, 2080, 2086, 2807]. Aims [3794, 897, 96]. Air [3325, 4006, 678]. Air-Pollution [678]. Akosombo [4062]. Alarmed [2486]. Alaska [3863, 3859]. Album [3643]. Aleppo [4193]. Alexander [3475]. Algeria [2447]. Algorithms [4169]. Alienated [153]. Aliens [3392, 2327]. Alive [1442, 2070]. All-Volunteer [3516]. Alley [766]. Alliance [1576, 338, 1574, 2836, 1984, 1842]. Allied [1026]. Allocated [2546]. Allocation [4066]. Aloïs [1278]. Alone [3947]. Alonso [1496]. Alphonse [1942]. Alter [4221]. Alter-Native [4221]. Altered [3975]. Altering [3138]. Alternate [2719]. Alternative [637, 3600, 767, 3558, 3588, 3430, 3043, 1638]. Alternatives [3083, 3896, 3684, 1580, 2269]. Always [3818]. Alzheimer [3140]. Am [2358, 3174]. Amar [1290]. Amartya [2061]. Amateur [818]. Amazon [2766]. Ambiguity [2489]. Ambition [3852]. Ambivalence [2339]. America [1969, 434, 672, 598, 1395, 1601, 1613, 1721, 1789, 1887, 2126, 2292, 2705, 2770, 1869, 1849, 1850, 1848, 945, 1099, 599, 2007, 3121, 3720, 3493, 3249, 3346, 1782, 2659, 3170, 1595, 998, 2601, 3721, 1895, 2300, 3376, 280, 985, 2215, 3161, 2424, 437, 1565, 130, 2870, 1120, 1743, 756, 3101, 3483, 3587, 602, 2731, 256, 995, 471, 3660, 3638, 4071, 4128, 2995, 1317, 1399, 1458, 298, 49, 1890, 994, 3610, 3035, 2493, 500, 3448, 3449, 603, 1636, 3296, 3619, 1412, 3522, 4078, 50, 1413, 3717, 2038, 628, 2341, 3065, 1637, 258, 2247, 699, 351, 3456, 3683, 1408, 3543, 3455, 2858, 3724].

American [2421, 3311, 3330, 3367, 3387, 3400, 329, 343, 524, 544, 1124, 1316, 1585, 2337, 2417, 2455, 2551, 2596, 2677, 3480, 1176, 252, 93, 48, 123, 3546, 677, 2421, 2611, 3061, 1497, 2883, 358, 359, 556, 3562, 1639, 3248, 333, 990, 3517, 777, 3244, 2555, 1545, 1099, 2552, 86, 3508, 933, 2130, 4129, 2169, 2686, 3863, 3542, 87, 2598, 3505, 1588, 3376, 2303, 442, 605, 781, 3158, 297, 2419, 2469, 767, 999, 938, 324, 1400, 2297, 536, 3379, 528, 3083, 537, 1546, 3164, 1334, 3073, 3688, 759, 1948, 1797, 3515, 778, 3611, 1172, 1731, 253, 2707, 4131, 3496, 2749, 3621, 3502]. American [331, 1373, 1786, 1059, 2457, 661, 2164, 2470, 51, 1635, 2539, 3548, 3326, 52, 3590, 1580, 2606, 3457, 2685, 1792, 769, 774, 3549, 122, 1000, 2162, 1321, 2015, 3378, 997, 1783, 2287, 3481, 3494, 1164, 22, 3372, 4212, 1180, 4219, 576, 2609, 782, 1598, 870, 1398, 3585, 1573, 2560, 3072, 3545, 296, 535, 1162, 2926, 710, 534, 354, 2267, 53, 2608, 1563, 2558, 3586, 873, 3690, 2193, 2710, 124, 332, 1407, 2562, 3500, 2475, 3366, 4217, 380, 3622, 2610, 1602, 3788, 2281, 3507, 2714, 247, 350, 1551, 1405, 4208, 3506]. Americanization [2617]. Americans [4041, 1411, 1589, 3862, 2849, 2985, 2625, 3498, 4040, 3623, 257, 1409]. Americanus [397]. Amid [3342, 3441, 3577]. Amidst [2253]. Among [2202, 4037, 1042, 2620]. Amory [21]. Analogy [1533]. Analysis [1386, 3950, 374, 1159, 138, 4054, 680]. Analyst [944]. Analytical [455]. Anarchical [561, 2444]. Anarchism [995]. Anatomy [1517]. Ancestors [4172]. Ancestry [4171, 3861]. Ancient [2956, 3760, 3761, 4018, 1221, 3764, 2970, 2653, 1219]. Andes [628]. Angelos [593]. Angels [3892]. Angle [2476]. Angles [3941]. Anglo [3546, 257]. Anglo-Americans [257]. Animadversions [355]. Animal [3390, 2881, 3278, 1426]. Animals [3393, 3287]. Anniversary [695, 691]. Annunciation [3012].

Anomalies [539]. Anon [3289]. Answer [1949, 3845, 3890]. Answers [1690]. Antecedents [3140]. Anthropocene [4013, 3734]. Anthropocentrism [3006]. Anthropological [795, 1374]. Anthropologist [479]. Anthropology [32, 1043, 1443]. Anti [1849, 1850, 1848, 1832, 3995, 4040, 368, 4132, 3371, 318, 3881, 3691]. Anti-Asian [4040]. Anti-Ballistic [1849, 1850, 1848]. Anti-Corruption [3881]. Anti-Government [4132]. Anti-Intellectualism [3371, 318]. Anti-Islamic [3995]. Anti-Realists [368]. Anti-Satellite [1832]. Anti-Science [4132]. Anticipations [2179, 1776]. Anticorruption [3875, 3871, 3884, 3876]. Anticosmopolitan [3317]. Antidiscrimination [4240]. Antipolitics [1469]. Antipoverty [2295]. Antiquity [3763]. Antiracism [3858]. Anxiety [1668, 2333]. Any [1737, 279]. Anzac [1961]. Apart [2577]. Apartheid [2791, 4067, 2810, 2802]. Ape [3006]. Apercu [1977]. Apes [1443]. Aphrodisias [3761]. Apocalypse [2960, 3359]. Apogee [365]. Apollonians [1112]. Apparatus [1858]. Appeal [2484]. Appearance [3030, 3006]. Appendix [1262, 1261, 1845]. Appetite [1494]. Application [3907]. Applications [4086]. Applied [2578, 2667]. Applying [3824]. Appointments [3100]. Appraisal [127]. Appreciation [1120]. Approach [566, 1727, 1943, 1716, 804, 800, 1907, 572, 569, 413, 3527, 3836, 2193, 768]. Approaches [449, 1158, 3877, 2689, 4073]. Appropriate [2345]. Appropriation [4065]. Appropriations [2484]. Aptitude [2460]. Arab [3844, 2331, 980, 3427]. Arabic [4025, 588]. Arabs [2699]. Archaeology [4171, 3376, 903, 1375]. Archaic [29]. Archetypal [293]. Archetypes [1540]. Archipelago [2700]. Architectural [683]. Architecture [154, 152, 2632, 1279, 422, 2717]. Archives [915]. Areas [3849, 4059, 683, 245]. Argument [2920, 3131, 3615]. Ariel [983]. Aristotle [3026]. Armaments [1236, 227]. Armed [4183, 3787, 3513, 3522, 3790, 4181]. Armonica [24]. Arms [2179, 1236, 219, 220, 213, 2182, 212, 1235, 3977, 1581, 2192, 1231, 1234, 2181, 1568, 1232, 221, 228, 2183, 214, 1229, 2178, 2184, 226, 2188, 215, 2185, 2186, 1580, 231, 1230, 2224, 249, 2187, 2191, 1241, 223, 1240, 3124, 1524, 1233, 2193, 222, 3984, 216, 2190, 224, 211, 3978]. Arms-Control [219, 214, 1240]. Arms-Limitation [224]. Army [3520]. Art [170, 2624, 1370, 154, 1674, 3653, 1910, 3161, 151, 4096, 1755, 2024, 1277, 1902, 7, 1440, 2707, 1311, 2510, 3032, 159, 2963, 1728, 2923, 674, 3759, 164, 1903, 29, 1860, 2631, 1278, 156, 1271]. Article [2060, 2061, 3336]. Artifacts [1522]. Artificial [1777, 4101, 1978, 1986, 1975, 4104, 1973, 4097, 1983, 4105, 1987, 1977, 1980, 1979, 4087, 4084, 1976, 4098, 1984]. Artificially [4106]. Artist [183, 1727, 3142, 1905, 741, 1155, 181]. Artistic [168, 1661]. Artists [1723, 156]. Arts [3311, 3330, 3367, 3387, 3400, 2682, 1722, 2691, 153, 2684, 2688, 152, 187, 461, 2782, 1721, 2677, 4121, 2679, 2687, 1724, 148, 483, 2685, 2689, 3954, 2681, 2680, 3354, 825, 155]. ASAT [1833]. Ascendant [2267]. Ashes [2260]. Asia [2050, 1563, 2647, 1007, 2748, 2651]. Asian [4040, 3623, 3885, 614, 4208]. Ask [2787]. Asking [2725]. Aspects [282, 400, 789, 2365, 788, 1773, 626, 2406, 870, 1364, 163]. Aspiration [2997, 2996]. Aspirations [1792]. Assassinating [2215]. Assault [2999, 3612]. Assembly [1210]. Assessed [1409]. Assessing [3173, 1829, 3780, 2461, 3829]. Assessment [1590, 2461, 1831]. assets [3456]. Assimilation [256, 2797, 3626, 2787]. Assistance [3912, 310]. Assisted [4103].

Association [4233, 4234]. Assumptions [2475]. Assurance [3528]. Astronomical [3699]. Astronomy [1385, 3702]. Asymmetric [3793]. Asymmetry [2258]. Atatürk [740]. Atlantic [1576, 1575, 1574]. Atlas [3809]. Atom [3699]. Atomic [383]. Atomized [3836]. Atoms [56]. Atrocity [2940, 2486]. Attempts [1418]. Attitudes [3344, 553]. Auden [1668]. Audience [7, 3641, 4117]. audiences [3445]. Audiovisual [1696]. Augmentation [4099]. Australia [1802, 1817, 1819, 1807, 1815, 1808, 1961, 1811, 1818]. Australian [1810]. Australians [1809]. Austria [724]. Austria-Hungary [724]. Autarkic [2228]. Autarky [2224]. Authenticity [15]. Author [693, 2672]. Authoritarian [2649, 3825, 3773]. Authorities [647]. Authority [1479, 874, 869, 1383, 2098]. Authorization [1847]. Autism [3392]. Autobiographic [879]. Autocrat [745, 105]. Automated [4047, 4048]. Automatic [1052]. Automation [3010, 4099, 4100]. Automatism [371]. Autonomous [3780, 1191]. Autonomy [2573, 2226, 371, 1748, 2271, 971]. Autophysiopsychic [3932]. Available [2727]. Avant [3214]. Avant-Gardes [3214]. Avoid [2806]. Awake [3062]. Awareness [925]. Away [4194, 3050, 1239]. Axel [1273]. Axes [3940]. Axiomatic [330]. Axioms [1951].

B [1262]. B.C [1213, 2875, 3020]. B.C. [1222]. Babel [1041, 1733]. Back [1134, 46, 58, 67, 78, 97, 114, 131, 144, 171, 190, 208, 233, 250, 262, 276, 288, 302, 315, 327, 340, 362, 376, 390, 405, 432, 444, 458, 472, 491, 507, 522, 541, 557, 573, 584, 595, 611, 631, 688, 670, 712, 733, 749, 771, 785, 805, 828, 837, 851, 865, 875, 890, 906, 921, 939, 954, 966, 976, 987, 1001, 1017, 1034, 1054, 1071, 1091, 1105, 1122, 1173, 1211, 1227, 1242, 1263, 1280, 1292, 1314, 1329, 1352, 1367, 1380, 1392, 1414, 1433, 1448, 1464, 1476, 1487, 1498]. Back [1512, 1529, 1542, 1556, 1569, 1583, 1599, 1611, 1624, 1640, 1656, 1669, 1684, 1705, 1719, 1734, 1749, 1762, 1774, 1787, 1799, 1820, 1834, 1851, 1864, 1885, 1900, 1911, 1923, 1933, 1944, 1955, 1970, 1988, 2000, 2025, 2044, 2063, 2075, 2089, 2108, 2124, 2139, 2158, 2175, 2194, 2206, 2220, 2233, 2248, 2263, 2277, 2289, 2305, 2319, 2335, 2348, 2372, 2384, 2398, 2414, 2428, 2440, 2453, 2479, 2496, 2511, 2526, 2548, 2567, 2579, 2594, 2612, 2626, 2640, 2654, 2668, 2692, 2703, 2721, 2737, 2675, 2751, 2768, 2784, 2798, 2813, 2823]. Back [2834, 2851, 2871, 2893, 2916, 2932, 2948, 2965, 2987, 3001, 3018, 3033, 3048, 3063, 3080, 3098, 3115, 3132, 3153, 3171, 3191, 3207, 3227, 3245, 3262, 3279, 3293, 3309, 3328, 3347, 3364, 3385, 3398, 3417, 1595, 1975, 4004, 3422, 2149, 930]. Back-End [3422]. Backbone [894]. Background [1246, 1421]. Backlash [3966, 4030]. Backward [1755, 2024]. Backwardness [2382]. Bad [3649]. Badly [2138]. Bads [1061]. Bahnhof [2358]. Baited [3661]. Bakke [1595]. Balance [3103, 2875, 3338]. Balancing [2338, 3403]. Balinese [958, 3122]. Balkan [2585]. Balkanizing [2383]. Balkans [3849]. Ball [3537]. Ballistic [1849, 1850, 1848, 1826, 1839]. Ban [3982]. Bandwidth [3714]. banker [3468]. Banking [2085, 3467]. Barack [3509, 3482, 3500, 3485, 3486, 3506]. Barbarism [1310]. Barbecue [2304]. Bargain [3025]. Bargains [1239]. Barrier [2797]. Barriers [547, 3729, 4206]. Based [3593, 3114, 4184, 3696]. Baselines [636]. Basic [213, 1033, 1890, 521, 106, 95]. Basis [2838, 484]. Bassa [4061]. Bassetto [1662]. Bastille [1918, 1919]. Battle [2118, 3863, 3791, 3791]. Battleships [2458]. Baudelaire [3653]. Bayh [3892]. Be [2764, 2585, 642, 4228, 1746, 3576, 3028, 1490, 3366, 1903, 3820, 3605, 3488]. Beard [1082]. Beast [959, 21]. Beasts [3297]. Beauty [1722, 2922, 2919, 2925, 3095, 2921, 3168, 2920, 3131]. Became [3286, 2408]. Because [3305]. Beckett [3656]. Become [1285, 4226, 810, 3749]. Becoming [3944, 4226, 4000, 3623]. Been [1951, 3498, 1899, 3305]. Before [2377, 3146, 3305, 4143, 2357, 1197]. Beginning [3627, 3282, 2531, 4084]. Beguiling [3247]. Behavior [2637, 3593, 1981, 1050, 652, 3534]. Behavioral [3953]. Behaviors [3229]. Behind [3006, 2765, 3058]. Beholder [2926]. Being [1285, 3021, 2201, 2410, 2196, 3587, 1323, 3547, 1112, 3526, 3308, 2923, 2855, 2459, 2714, 37]. Beings [795]. Belief [1623, 1159, 4243]. Beliefs [3864, 2979]. Believers [1622]. Believing [4009]. Bell [2992]. Bellagio [560, 2443]. Beloved [3660]. Below [2495, 1171]. Benefit [4054]. Benefits [2153, 4054, 3625]. Benjamin [3379, 24, 3243]. Bentham [3026]. Benthamite [1807]. Berkeley [2887]. Berlin [2358, 2369, 3930]. Berlioz [1921]. Bernard [357]. Bess [4022]. Best [3346, 3548]. Better [3863, 3944, 2688, 1332, 1206, 2897, 1340]. Between [1861, 3130, 2115, 471, 4067, 1307, 55, 1849, 1850, 1848, 3951, 2211, 3493, 3863, 1993, 2788, 1112, 4150, 909, 2413, 90, 3026, 1537, 2383, 436, 2004, 2273, 1979, 1759, 1032]. Beyoncé [3487]. Beyond [4246, 2424, 2729, 3412, 2723, 4114, 3803, 2936, 3090, 576, 2855, 4051, 4052, 549, 3984]. Bialik [577]. Bias [4246, 4229, 3862, 4237, 4239, 4240, 3995, 4238, 4242, 4241, 4236, 4230, 4233, 4243, 4245]. Biases [4235]. Bibliography [232]. Big [1029, 3919]. Bigger [297, 3465]. Bighorn [3791]. Bilingualism [3112]. Bill [2972]. Billiards [2240]. Billion [2518]. Biocapital [3284]. Biodiversity [3577]. Bioethical [2726].

Bioethicists [2734]. Bioethics

[2735, 2736, 2732, 2729, 2725, 2723, 2728]. Biographical [877, 878]. Biography [2695, 2334]. Biological [4223, 2288, 651, 2952]. Biologist [1420, 2018, 3118]. Biology [1962, 4147, 1534, 4224, 3697, 2860, 885, 1390]. Biomedical [2067, 1379]. Biophilia [4152]. Biophilic [4151, 4150]. Biosocial [3201, 1355]. Bipartisan [4029]. Bipolar [451, 450]. Bipolarity [2819]. Bird [4191]. Bird’s [1819]. Bird’s-Eye [1819]. Bismarck [742]. Bizen [731]. Black [1100, 933, 4022, 925, 4207, 3505, 1095, 938, 3162, 2420, 923, 1642, 863, 934, 3496, 2450, 936, 932, 1590, 3461, 3504, 929, 937, 1103, 3645, 1372, 1104, 3703, 3506, 1409]. Black-White [2420]. Blackface [2087]. Blackness [3508]. Blacks [1585, 1883, 1137, 1193, 2014]. Blending [3553]. Blind [3596, 3045]. blocks [3454]. Blood [2085]. Blow [2605]. Blowing [2562]. Blue [3645]. Blues [1181]. Blunt [3086]. BMD [1827, 1845, 1833, 1840]. boîtes [3656]. Boarding [1620]. Boardroom [4007]. Body [3306, 2488, 2639, 3175, 2487, 4063, 151, 3182, 959, 3141, 3179, 600, 2848]. Bogdanovka [4189]. Bohemia [2116]. Bohr [57]. Bollinger [3067]. Bomb [2697, 383]. Bombers [3217]. Bonds [2085]. Bongo [4058]. Book [2334, 344, 2877, 3660, 354, 352, 1695, 350]. Book-Reviewing [354]. Books [2674, 346, 344, 1663, 1266, 2528]. Bookshop [351]. Bookstore [1693]. Boom [3810]. Boon [3135]. Bootstrapping [3009]. Border [2609]. Borders [4180, 2347]. Borg [1283]. Boris [207]. Born [748]. Boss [3872]. Bostonians [4021]. Both [2311, 1727]. Bother [3232]. Bottleneck [306]. Boundaries [2729, 3070, 4122, 2609, 802, 3301, 2650]. Boundary [1537]. Bourgeoisie [1634]. Bouts [3866]. Boys [942]. Brain [2642, 2639, 2410, 3182, 1975, 2630, 3225, 2632, 2628, 2955, 2638, 4236, 3713, 2629, 3177, 3630, 1533, 3711, 3707, 2631]. Brains [3200, 2636, 1980]. Brainwashing [361]. Branch [3920, 3605]. Branches [2620]. Branchpoint [1975]. Brazil [2753, 2754, 2767, 2757, 996, 3882, 2760, 2755, 2759]. Brazilian [2762, 2763, 591]. Break [202]. Breakdown [2252]. Breakthrough [1216]. Breed [605]. Brexit [3814]. Bricks [2528, 2540]. Bridge [2702]. Bridging [3899]. Bridgman [126]. Brief [3922, 981, 389, 898, 3734, 3868]. Bright [3696]. Bringing [4120, 3641, 1463, 2149, 2759]. Britain [625, 998, 2071, 411, 626, 2590]. British [701, 1679]. Broadband [3526]. Broadband-Connected [3526]. Broader [2241]. Broadway [3554, 3644]. Brothers [3754]. Brown [1595]. Brubeck [3928]. Brunelleschi [1954]. Bruni [1306]. bubble [3456]. Bucha [4190]. Buddhism [1006, 581, 2219]. Buddhist [1008, 2846]. Budget [1238, 1847]. Budgetary [2822]. Building [3835, 3849, 2630, 4150, 2199, 3852, 2761, 881, 1982]. Burden [3135, 2753]. Burdens [4039, 1906]. Bureaucracy [428, 4051]. Bureaucratic [1760, 1423]. Bureaucratization [1327]. Bureaucrats [4047]. Burma [1006]. Burmese [709]. Burnt [4189]. Burnt-Out [4189]. Bus [4189]. Bush [3038]. Business [777, 2804, 2427, 775, 781, 3919, 773, 778, 4007, 776, 782, 2434, 784, 783, 1547]. Businessman [550, 774]. Bust [3810]. Butler [1090, 3379]. Buy [3027, 1067, 2531]. Buying [2924]. Buzzer [2992]. Bytes [2528, 2540]. Byzantium [1303].

C [3550]. C. [2060]. Cacophony [1791]. Cahora [4061]. Calculated [330]. California [1119]. Call [346, 2468]. Calls [3890]. Calvino [2720]. Cambridge [3016]. Campaign [3585, 3434]. Campus [3260, 3947]. Can [1949, 3883, 3979, 4226, 3935, 4066, 642, 2728, 3820, 2403, 3571, 2861, 4167, 3814, 3908, 3605, 3028, 2806, 3736, 2518]. Canada [2028, 2027, 2037, 2032, 2030, 1025, 2035, 2042]. Canadian [2040, 2036, 2274, 2034, 2031]. Canberra [1805]. Candor [1793]. Cannot [761]. Canonical [1704]. Canonizing [3691]. Can’t [2464, 3916, 3930, 3927]. Canteen [3642]. Capabilities [117]. Capability [3918]. Capacity [4045, 3714]. Capital [2210, 1805]. Capitalism [4209, 3255, 3251, 3254, 1739]. Capitalist [3247]. Capriciousness [2593]. Capture [3883]. Capturing [4052]. Car [2154]. Carbon [3593, 2519, 3407, 3696]. Carbon-Based [3593]. Carbon-Constrained [3407]. Cards [2576]. Care [4183, 1894, 1879, 1878, 1349, 2034, 2402, 2862, 2663, 2069, 1887, 4226, 2407, 1348, 4228, 2403, 2287, 1339, 2068, 4181, 4148, 1361]. Career [4242, 91]. Careers [1967]. Caregiving [4147]. Caribbean [1098, 1565, 1097]. Carnegie [3445]. Carnegie-Knight [3445]. Carpet [4021]. Carreta [1594]. Cartographical [915]. Cartoon [2095]. Carver [1615]. Case [3733, 2449, 4120, 3761, 4092, 2315, 905, 2228, 1897, 2959, 3609, 2939, 2695, 3792, 229, 1496, 4239, 3584, 3889, 995, 731, 2687, 583, 1059, 1221, 2942, 1717, 3820, 3891, 2975, 3107, 4232, 1999, 2520, 1457, 972, 706, 1839, 4184, 4245, 2789, 3552, 2974, 1271, 3440]. Cases [3681, 2590]. Cash [3640]. Caso [2829]. Cast [297]. Castle [1273]. Casualties [3792, 3524, 3799]. Cat [3688]. Catalog [2535]. Catalogs [2535]. Catalogue [3]. Catalyst [1196]. Catalysts [3945]. Catastrophe [2049]. Catastrophes [2366]. Catch [2765, 356]. Catch-22 [356]. Categories [622, 972]. Category [3313]. Catholic [1633, 2764, 975, 120].

Catholicism [1632, 2912]. Catullus [3758]. Caught [1112]. Causal [324]. Causes [2258, 2131, 2188, 1604, 4008, 2366]. Cautionary [1714]. Cautiously [4145]. Cell [3454]. Census [3068]. Center [3910, 1012, 1028, 2197, 3127, 1724]. Center-Periphery [1012]. Centers [2713, 1347, 2716]. Central [1565, 2111, 2113, 2260, 884, 628, 2121]. Centrality [3144, 2537]. Centralization [638]. Centrifugal [463]. Centripetal [463]. Centrists [2211]. Centuries [727, 1301, 725]. Century [908, 1074, 3877, 4128, 3674, 1636, 662, 2559, 1494, 3134, 3546, 260, 3627, 3568, 280, 974, 2467, 905, 3856, 2870, 2437, 3164, 728, 729, 35, 3638, 1324, 3608, 3188, 3516, 982, 2629, 3503, 2200, 784, 2710, 319, 3606, 2720, 2448, 2314, 2542, 37]. CEOs [3749]. Ceremonies [1805]. Certain [726]. Certainty [3895]. Ceylon [1006]. Chains [4156]. Chairman [640, 666]. Chalice [1855]. Challenge [827, 992, 777, 1651, 2469, 3352, 291, 2067, 3738, 2450, 3681, 2934, 3295, 3664, 1339, 3822, 3947, 3799, 2542, 3448, 3449]. Challenges [1654, 4178, 4115, 3723, 1782, 3665, 2788, 2420, 4177, 3558, 4090, 1772, 3965, 3608, 3674, 2681, 2790, 2609, 3663, 2801, 2471, 3834, 4203, 2452]. Challenging [3231, 3500, 4043]. Chance [3281, 3695]. Chances [3281]. Change [3561, 2117, 481, 1223, 1595, 2422, 813, 950, 1628, 2023, 3355, 545, 2420, 3558, 291, 575, 978, 759, 4010, 2516, 1137, 743, 4056, 618, 298, 4013, 3808, 299, 926, 3809, 2838, 2352, 1540, 4062, 1577, 2599, 1162, 534, 2285, 3806, 1453, 2928, 269, 1346, 1457, 3563, 4243, 3403, 4002, 3037, 3805, 4008, 4009, 2282, 1154, 2544, 3441]. Changed [3113]. Changes [430, 418, 3593, 2663, 2362, 2760, 1404]. Changing [1397, 1892, 3342, 1485, 3728, 1898, 437, 2438, 1951, 244, 536, 672, 1107, 2196, 1218, 3685, 520, 1482, 2597, 3549, 419, 2761, 2600, 1930, 1241, 3389, 4118, 3789, 709, 1408, 4069, 2204, 1365, 165, 2533]. Chaos [2658]. Chapters [62]. Character [1296, 2831, 1621, 1609, 1781, 3788]. Characteristics [726, 214, 943, 618]. Charge [3094]. Charisma [739]. Charismatic [1644, 738]. Charlemagne [1077]. Cheap [4183]. Check [2806]. Cheers [1060]. Cheese [2830]. Chemical [3700, 3698]. Chemistry [3698, 3697, 885, 3695]. Chessboard [3240]. Chicago [3185]. Chicano [1593]. Chicanos [1585, 1591]. Child [944, 2303, 1170, 2295, 320, 1356, 1778, 2406, 1361]. Child-Care [1361]. Childhood [128, 2294, 3014, 259, 2292, 2411, 2299]. Childlike [1778]. Children [2300, 2103, 4037, 2931, 1616, 1445, 3910, 2832, 4235, 2293, 4189, 2830, 1875, 1350, 2409, 2302]. Chilean [700]. China [2307, 2311, 361, 219, 2315, 3665, 3677, 337, 3676, 3668, 3669, 1742, 3825, 3670, 2316, 2317, 2313, 3664, 3886, 3663, 2200, 2490, 204, 1219, 2310, 2650, 3675, 2197, 2262, 2491, 3127, 2847, 3673, 1564, 3666, 2651, 3671, 2314]. Chinese [2196, 3975, 2201, 2315, 2309, 2202, 707, 3326, 3674, 2205, 972, 2204, 1938, 2203, 25]. Ch’ing [971]. Chipping [1239]. Chocolate [3397]. Choice [3562, 439, 2338, 1337, 644, 3093, 3092]. Choices [3559, 331, 2343, 3748, 3750]. Chord [3649]. Choreographies [3927]. Christendom [601]. Christian [1286]. Christianity [616, 3077, 3547, 1267, 2842, 610]. Chromatic [1858]. Chronology [1261, 2957]. Chronometry [2950]. Church [2788, 2216, 1304, 2215, 2210, 2368, 653, 1997]. Churches [2422, 605, 600]. Churchill [2334]. Cigarette [2169]. Cinematography [161]. circa [4111]. Cities [403, 759, 649, 3936, 760, 237, 248, 1664, 2720]. Citizen [305, 3533]. Citizens [3821, 3883, 2327, 2759, 3610]. Citizenship [2754, 3937, 1959, 1502, 1792, 535, 2801, 2760, 2591, 2789]. City [403, 1366, 754, 3761, 244, 751, 769, 1000, 752, 494, 155, 681, 3625, 753, 247, 1142]. Cityscape [246]. Civic [580, 1794, 3948, 3608, 2903, 1307, 3612]. Civil [1639, 599, 2007, 3121, 1300, 3906, 554, 3830, 1745, 324, 3827, 3379, 3833, 3921, 532, 3841, 3846, 3853, 3854, 3844, 3828, 3843, 3920, 1768, 3494, 3260, 3829, 3838, 3848, 3633, 4176, 3834, 3612, 3069, 3831, 1409, 2749, 2216]. Civilian [3792, 3799]. Civilians [3788]. Civility [1553]. Civilization [1310, 1316, 127, 3591]. Civilizations [2981, 1224, 1386]. Civilized [4082]. Civilizing [2791, 1148, 3084, 2913]. Claims [3795]. Clash [2981]. Class [418, 545, 2902, 1731, 3961, 1927, 419, 626, 2763, 1598, 1596, 628, 552, 2370, 420, 1080, 3455, 1647]. Classical [1967, 1950, 3757, 3759, 823]. Classics [1074, 3755, 3157, 3765, 822]. Classification [2120, 3065]. Classified [2322]. Classroom [1550, 3952, 4020]. Clearing [3456]. Clergy [398]. Clerisy [986]. Cliff [357]. Clifford [3261]. Climacteric [1177]. Climate [3561, 430, 3726, 4205, 4010, 4007, 4056, 4006, 4013, 4062, 4005, 4001, 3563, 3403, 4002, 4008, 4069, 4009]. Clinical [798, 72, 2069, 2731, 802, 2868, 1426]. Clio [919, 1142]. Clock [2952]. Clocks [2951]. Close [3950]. Closely [1711]. Closer [2759]. Closing [90, 3739, 3749]. Clouds [2690]. Clowns [3242]. CNN [3851]. Coalitions [1603]. Cockfight [958, 3122]. Coda [669]. Codex [3476]. Coding [4093]. Coeducation [1158]. Coercion [3781]. Coercive [2939]. Cognate [3695]. Cognition [3714]. Cognitive [855, 4237, 1907, 4087]. Cognitive-Scientific [1907]. Cohesion [479]. Cold [1231, 1929, 2918, 2281, 3828]. Coleridge [514]. Collaboration [483].

Collapse [3591]. Colleagueship [91].

Collective

[3776, 1546, 1277, 2648, 4172, 2651]. Collective-Identity [2651]. College [3943, 3944, 576, 873, 2682, 2686, 924, 925, 1170, 461, 1166, 2683, 934, 931, 2679, 2687, 936, 932, 2685, 937, 928]. Colleges [933, 2691, 3864, 2601, 2684, 2688, 923, 2677, 927, 2689, 926, 2681, 929, 2680, 3354]. Collegiate [3953, 3354]. Collide [2639]. Colloquy [77]. Collusion [3819]. Colombian [3837]. Colonial [3094, 1006]. Colonialism [1095, 1093]. Colonization [4207]. Colony [3515]. Color [616, 3627, 623, 3802, 613, 615, 620, 618, 624, 627, 626, 2763, 630, 628, 619, 614, 621]. Colorblindness [2797]. Colored [625]. Colorism [4033]. Colossus [3083]. Colyton [727]. Combat [4238]. Combating [3877, 3939, 3885]. Come [4238, 2756]. Comedy [36]. Comes [3798, 2906, 3554]. Comfort [4006]. Coming [677, 2330, 3576]. Commemoration [3761]. Commencement [470]. Comment [768]. Commentary [1795]. Comments [2054, 2488, 2056, 454, 178, 2055, 2058, 2053, 2050, 2060, 2049, 267, 2057, 249, 39, 123, 2061, 270, 2051]. Commercial [1723, 348]. Commission [667, 768]. Committed [4166, 4192]. Committee [93]. Committees [11]. Common [1849, 2904, 3609, 3603, 3617, 3616, 3888, 1688, 2465, 2232, 3604, 490, 3683, 3615]. Common-Law [3683]. Commonplaces [965]. Commons [3535, 3617, 3533]. Commonsense [4092]. Commonwealth [3269]. Communication [243, 2507, 4121, 246, 1389, 659]. Communications [1485, 506, 1681]. Communicative [2376]. Communism [2250, 2386, 2255, 2743, 2261, 2395]. Communist [704, 2252, 3773, 1510, 219, 204]. Communists [2695]. Communities [4229, 2537, 1779, 4076, 4217]. Community [4142, 2601, 1622, 545, 2034, 843, 1702, 931, 3913, 1896, 932, 3903, 2538, 2485, 1741, 2085, 614, 4145, 1207, 4184, 1695, 1471, 4073, 2155]. Community-Based [4184]. Compact [3954]. Companies [1255, 1254]. Company [1674, 3522]. Comparative [2611, 1536, 3205, 1224, 2644, 696, 2133, 3499, 2218, 3519, 1831, 3775, 3666, 3159, 3457]. Comparison [2449, 3393, 729, 2696, 204]. Comparisons [3940]. Compassion [2935, 2855]. Compatible [1879]. competing [3377]. Competition [3975, 3966, 4102, 2344, 1032]. Compiling [2869]. Complacency [1059, 3664]. Complaint [1171]. Complementa [1308]. Complementarity [886, 2010]. Complementary [42, 3775]. Complex [3632, 3833, 2238, 794, 3521]. Complexity [758, 2599, 534, 3569]. Complicated [3866]. Composition [2426]. Comprehending [1618]. Comprehensibility [1748]. Comprehensive [2295, 1618, 1615, 224]. Compromise [3616]. Compromised [3105, 3150]. Computation [2236]. Computational [3698]. Computer [4104, 4088, 1717]. Computers [2241, 889]. Computing [4086, 2239, 2244, 2237, 2245]. concentration [3451]. Concept [1898, 1958, 3643, 551, 3491, 3316, 2860, 1320, 165]. Conception [40, 946, 1573]. Concepts [1827, 481, 3009, 1826, 1451]. Conceptual [2276, 1709]. concerning [2878]. Concerns [1418, 3718, 687]. Conciliation [3484]. Conclusion [3985, 3855, 4186]. Condition [1111, 4082, 3315, 2742]. Conditions [580, 564, 559, 560, 2443, 4156]. Condorcet [3041, 1269, 2876]. Conduct [3793, 3799]. Conducting [1031]. Conference [556, 560, 2443, 815, 820]. Confession [1440]. Confessional [3993]. Confidence [4129, 3607, 3276]. Confidence-Man [3276]. Confidences [3397]. Confinement [3891]. Confirmation [3339]. Conflict [4183, 4178, 1889, 985, 454, 453, 3968, 4177, 446, 1009, 1324, 3787, 3847, 441, 4181, 4179, 1602, 1647]. Conflict-Prevention [3847]. Conflicting [3859]. Conflicts [2883, 3218]. Conflictual [3775]. Conform [3783]. Confrontation [1421]. Confronting [1731, 3799]. Confronts [2779]. Confucian [2748, 1364, 1288, 2847, 2651]. Confused [2138]. Confusion [1480]. Congress [1847, 1238, 2530, 285, 3259, 2967, 3340, 3958, 3338, 3771, 3339]. Connected [3526]. Connecticut [3858]. Connectionism [1979]. Connections [3358, 953]. Connectors [3903]. Conquistador [1780]. Conscience [751, 752, 2219, 3994]. Conscious [373, 2242]. Consciousness [1141, 2896, 1635, 604, 3715, 1538, 1681]. Conscription [3519]. Consensual [177]. Consensus [1545, 1548, 1745, 1544, 1324, 2352, 1554, 1602]. Consequences [2258, 3526, 3739, 2346, 1881, 3630, 269]. Conservation [673, 2766]. Conservative [1218]. Conservatives [4132]. Consideration [2476, 981, 15, 1375]. Considerations [970]. Consilience [2620]. Consistent [994]. Constitutes [1915]. Constitutional [3544, 1402, 3543, 3692]. Constrained [3407]. Constraining [4051]. Constraint [1534]. Constraints [3953, 3587, 3402]. Construction [2488, 2487, 2393, 4069, 2203]. Constructive [3043]. Constructs [1962]. Consumer [1345, 2809]. Consumption [3593]. Container [2830]. Contemplative [1285]. Contemporary [3061, 4178, 1782, 4023, 1158, 1589, 849, 1120, 38, 304, 460, 2896, 602, 65, 2687, 587, 462, 52, 419, 1442, 1164, 591, 50, 604, 2099, 2822, 1447, 588, 4176, 3624, 2717, 3926, 1547, 1364, 582]. Content [1046, 3603, 82, 50]. Contested [3303, 2333]. Context [3599, 3975, 1838, 1959, 856, 1873, 2476, 4074, 4230, 296, 1233]. Contexts [2294, 1748, 2050]. Contextual [3527]. Continent [1780]. Continental [2380]. Continents [2328]. Contingent [3744]. Continue [2900]. Continued [3393]. Continuing [3505, 3688, 2664, 2667, 467]. Continuity [40, 1005, 882]. Continuous [3625]. Contract [3810]. Contractors [3517]. contradictions [3453]. Contradictory [1503, 2376]. Contraria [1308]. Contrarian [2509]. Contrary [1996]. Contrasting [2731]. Contributions [2736, 3621]. Contributors [3416, 3431]. Control [2179, 1236, 219, 121, 220, 213, 2182, 212, 3977, 1581, 2192, 798, 4079, 1231, 1943, 2072, 2181, 1568, 3677, 1232, 221, 228, 2183, 214, 1229, 2178, 2079, 2275, 119, 447, 2184, 226, 797, 2188, 2185, 2268, 2186, 1580, 231, 1230, 2049, 3873, 1430, 249, 2187, 227, 2191, 2172, 652, 1241, 1240, 1524, 1233, 117, 2193, 222, 3984, 216, 120, 2190, 211, 3978]. Controls [2227]. Controversial [3508]. Controversies [317, 116]. Controversy [2713, 1526]. Convection [31, 30]. Conventional [217, 2191]. Conventions [1677]. Converge [4212]. Convergence [3697]. Conversation [154, 2954, 4011, 3211, 4002, 3276]. convictions [3202]. Convocation [2457]. Cooperation [3968, 2189]. Cooperative [1196]. Coping [758]. Coral [4012]. Core [2317, 3847]. Corner [1541]. Corporate [2781, 3879, 3917, 3249]. Corporation [2506, 3611, 3914, 1401]. Corporations [4242]. Correcting [3667]. Correction [3509, 3839]. Correspondence [129]. Corruption [3883, 3877, 3882, 3873, 3886, 3885, 3878, 3874, 3880, 3879, 3881]. Cosmetics [2926]. Cosmic [3700, 2848]. Cosmology [3705]. Cosmopolitan [3313, 3316, 3315]. Cosmopolitanism [3321, 3318]. Cosmopolitans [2652]. Cosmos [2845]. Cost [1894, 4054]. Costs [3559, 3952, 4054]. Could [4228]. Council [4183]. Counsel [3906]. Count [5, 31]. Counter [1140]. Counter-Elites [1140]. Counterrevolution [3512]. Countess [3397]. Countries [1066, 704, 3406, 1025, 2058, 1248, 309, 3885, 2408]. Country [2764, 3640, 4066, 2283, 3509, 3486]. Countryside [2773]. County [1722]. Coup [739]. Couple [3703]. Courageous [3792]. Course [3139]. Court [3343, 3686, 2166, 456, 3897, 3100, 3584, 3889, 607, 3498, 3692, 3383, 3067, 2967, 3340, 3338, 3606, 3339, 3881]. Courtesy [2355, 2358, 2361, 2364, 2367, 2369]. Courting [3689]. Courts [3342, 4080, 3688, 3343, 3686, 3691, 3678, 2994, 3685, 3888, 3681, 3680, 3344, 3345, 3332, 3679, 3687, 3916, 3690, 2423, 3689]. Covenant [3544]. Coverage [3668]. Covering [4022]. Covert [1939]. COVID [4131]. COVID-19 [4131]. Cracks [2273]. Craft [366]. Crafts [998]. Crane [3059]. Cranes [2830]. crash [3472]. Crawford [3639]. CREATE [3950]. Creating [4117]. Creation [514, 1521, 1540, 4099]. Creationism [1119]. Creative [183, 4058, 516, 161, 308, 1785]. Creativity [515, 509, 1907, 149, 4093, 1728, 1538, 519, 510]. Credibility [1131]. Credo [1172]. Crick [884]. Cries [1786]. Crime [4180, 1662, 3833, 2854, 3460, 2833, 1404, 3453, 3458, 2810]. Crimen [4187]. Crimes [3795]. Criminal [4079, 4080, 3683]. Criminalization [3632, 4078, 4070]. Criminalizing [4035]. Crises [1653, 3423]. Crisi [3286]. Crisis [1150, 2611, 1261, 1639, 1179, 2825, 335, 1479, 2912, 2870, 3677, 879, 4205, 1244, 3686, 2660, 4218, 471, 2375, 1575, 4006, 2362, 3832, 2761, 869, 3664, 604, 809, 1247, 2191, 4216, 1255, 3473, 3470]. Critic [984]. Critical [2734, 856, 3868, 232]. Criticism [170, 965, 1701, 348, 856, 2557, 185, 1428, 2558, 873, 1383]. Critics [1963]. Critique [872, 3797, 1270, 1422]. Croatian [2584]. Cromwell [3087]. Cross [1165, 2210, 1897, 3023, 61]. Cross-Cultural [2210, 1897, 3023]. Cross-Linguistic [3023]. Crosscurrents [1590]. Crosshairs [3752].

Crossings [2609]. Crossroads [2819, 2914, 2354]. Crow [3493]. Crucial [3101, 2302]. Crucible [3638, 552]. Cruelty [3462, 1660]. Cruiser [4238]. Cult [1631]. Cultivation [517]. Cultural [481, 3864, 2315, 430, 2210, 1897, 3039, 3755, 2275, 2616, 1814, 2484, 2205, 926, 3157, 1929, 3504, 273, 2313, 2558, 186, 1457, 2200, 873, 270, 2310, 2197, 2308, 3127, 1952, 3213, 3023, 1011]. Culturally [4209, 3055]. Culture [1593, 175, 183, 1495, 488, 2169, 1301, 3054, 2831, 842, 1587, 1932, 1594, 957, 1671, 2551, 2615, 2771, 3515, 2604, 2779, 178, 485, 4238, 3621, 862, 476, 185, 2276, 184, 174, 1813, 3408, 479, 769, 477, 626, 1768, 487, 2396, 484, 2352, 2792, 2092, 39, 2646, 1673, 2895, 2156, 2837, 2795, 2332, 176, 2796, 3690, 2900, 2794, 3615, 486]. Cultures [2702, 1206, 1010, 1785, 898, 271, 1443]. Cumulative [692]. Cunning [1793]. Cures [2131]. Curious [4092]. Current [3827, 317, 355, 116, 3718, 2554]. Currents [2684, 4212]. Curricular [1154]. Curriculum [1150, 2904, 1046, 3097, 1995, 3765]. Curse [1041]. Cut [2311]. Cyber [3980, 3781]. Cybersecurity [3530]. Cyberspace [3529]. Cycle [3862, 1283, 3422, 655, 3427]. Czechoslovakia [2259].

Dædalus [2671, 2674, 3480, 1, 3151, 3349, 3017, 2002, 2]. Dakota [2776]. Dam [4061, 4062, 4064]. Damaged [2130]. Damasio [3178]. Dammed [4063]. Dance [1727, 3642, 3927, 3181]. Dancer [3094]. Danger [2929]. Dangers [2578, 3272, 184, 3459]. Daoist [2848]. Dar [4058]. Dark [1780]. Darkened [3240]. Darkly [2374]. Darkness [3661]. Darwin [514, 3231, 2888]. Darwinian [130]. Data [4107, 4182, 719, 4232]. databases [3442]. Dating [2957]. Dave [3928]. Davis [2870]. Day [3357]. Days [2982, 868].

De-Enlightenment [1536]. Dead [4012]. Deal [3421]. Dealing [2371]. Dealt [2366]. Death [358, 3166, 3554, 383, 2493, 2910]. Debate [55, 2899, 2592, 2702, 130, 3265, 3780, 2591, 1361]. Debates [2041, 3369, 2565]. Debris [2383]. Debye [3202]. Decade [4086, 1209, 2304]. Decadent [1443]. Decades [2226, 2002, 1409]. Decent [4153, 4082]. Decentralization [638, 1523]. Decentralized [2759]. Deciding [1746]. Deciphering [960, 3140, 3575]. Decision [4103, 1347, 3814, 1713]. Decision-Makers [1713]. Decision-Making [4103, 1347]. Decline [1300, 1306, 724, 2274, 1546, 1075, 1294, 3688, 1299, 661, 2903, 3149, 1664, 1028, 2434, 1309, 2051]. Declining [4129, 3495]. Decolonial [4126]. Decolonial/Postcolonial [4126]. Deconstruction [4050]. Dedication [3165]. Deep [4086, 958, 3122]. Defeat [2575]. Defective [2433]. Defending [3522]. Defense [1238, 1842, 334, 1579, 2228, 2402, 1826, 1229, 1558, 1572, 1761, 387, 1841, 1580, 2232, 3852, 1192, 1811, 898, 1839, 1847, 1828, 1830]. Defensible [3992]. Deferential [3989]. Defiant [2227]. Deficient [3681]. Definition [3193, 2585]. Definitions [1952]. Degradation [4080]. Degrees [3745]. Degrowth [4152]. Deliberate [652]. Deliberation [3825, 3816, 3814, 3819, 3822]. Deliberative [3821, 3815, 3253, 3824, 3820, 3817, 3818, 2929, 3823]. Deliso [357]. Delivery [4182, 4185]. Delusion [4061]. Demands [4010]. Dematerialization [2522]. Demedicalization [1334]. Demise [2258]. Democracy [3226, 3248, 2449, 260, 2808, 3986, 4220, 2381, 2754, 3250, 3249, 2130, 416, 497, 998, 2684, 3815, 187, 1951, 3247, 3253, 3824, 3255, 428, 3988, 3251, 278, 2279, 3987, 3948, 2995, 4105, 3768, 3820, 3817, 2216, 3818, 2353, 2313, 2801, 2828, 3605, 4162, 3819, 2929, 3823, 3642, 4137, 3433]. Democrat [1497]. Democratic [3618, 1444, 1701, 995, 3613, 1327, 3820, 3610, 3254, 3814, 1770, 2758, 1710, 3351, 3679, 2393, 3614, 644, 2562, 3775, 2256]. Democratically [3992]. Democratization [3086]. Demographic [726, 3665, 724, 529, 722, 2008, 3717, 1872, 720]. Demography [728, 717, 719, 730]. Denunciation [1493]. Dependence [2915]. Dependency [2153]. Depolarize [4002]. Deportation [4039]. Depression [4225, 3474]. Deprogramming [4245]. Derived [795]. Descent [622]. Desegregation [547]. Deservingness [3937]. Design [845, 3915, 323, 312, 3750, 803, 271]. Designer [819]. Designing [3597, 2666, 1152]. Desirable [3250]. Desire [3233, 2081, 570]. Despair [2115]. Despondency [366]. Destabilization [3832]. Destabilizing [3502]. Destinations [3626]. Destiny [1086, 1753]. Destruction [4050]. Destructive [2491]. Determinants [1236, 2733, 2401, 3630]. Determination [1202, 2323]. Deterrence [1567, 3976, 1825]. Deterrent [3412]. deux [3656]. Devastating [1731]. Developed [2408]. Developing [3406, 2666, 2058, 2328, 245, 1979]. Development [3508, 1968, 516, 101, 2736, 2056, 2410, 2297, 3039, 2047, 731, 721, 661, 4066, 4061, 567, 676, 379, 2053, 1648, 1929, 2057, 1247, 646, 1649, 2107, 720, 2051, 942, 2610, 1744, 2667, 730, 2050]. Developmental [4075]. Developments [203]. Devices [4023, 3746]. Devon [727]. Dewey [1497, 129, 3178]. Dialectics [2436]. Dialogue [2093, 427, 2709, 1167]. Dialogues [831]. Diamonds [3548]. Diary [4193, 1290]. Diaspora [2747]. Dichotomy [3951]. Did [3935, 2553, 3123]. Diderot [2876]. Dido [3659]. Died [2592]. Diego [1496]. Difference [1452, 2301, 2793, 1964, 2310]. Differences [3053, 1325, 2859, 3587, 3949]. Different [2598, 538, 2770, 2038]. Differentials [2413]. Differentiation [1032]. Difficult [1738]. Difficulties [2131, 3939, 95]. Difficulty [3061]. Diffusion [2516]. Digital [3766, 3747, 4226, 2529, 3360, 3750, 3433]. Digital-Plus [3750]. Digitized [2545]. Dignity [3137, 1755, 2024, 4082, 29]. Dilemma [2421, 1208, 1030, 3959, 536, 867, 2417, 3613, 3961, 819, 1770, 3507, 402]. Dilemmas [1782, 2338, 2731, 2625, 1772, 1744]. Dilly [3275]. Dimension [4103]. Dimensions [3745, 855, 2839, 1526, 1237]. Diminished [2432]. Diner [3447]. Dionysians [1112]. Diplomacy [333, 3424, 2150, 2701]. Diplomatic [228]. Directive [3070]. Dis [2590]. Disabled [3687]. Disadvantage [3940, 2404]. Disamenities [1061]. Disappearance [2874]. Disarmament [2179, 55, 229, 3982, 3415]. Disassembling [1551]. Disaster [4013, 2928]. Disbound [4197]. Discipline [1027, 3382]. Disciplines [2905, 2551, 2565, 480, 826]. Disconnect [3583]. Disconnected [3526]. Discontents [1794, 3112, 4128, 2969, 2380]. Discourse [2205, 3371, 2590, 2468]. Discover [3698]. Discoveries [1391, 3699, 2993]. Discovers [31]. Discovery [3010, 948, 372, 3702]. Discretion [4239]. Discussion [2052]. Disease [1889, 1943, 2081, 3140]. Diseased [4046]. Diseases [2170]. Disenchantment [1635]. Disfarmer [3029]. Disjunction [488]. Dismay [2484]. Dismissal [3491]. Disorder [3973]. Disorders [4224, 4222, 3576]. Disparities [4238, 3938]. Displacement [274, 3688]. Displacing [2493]. Disrupting [4239, 4117]. Dissent [849, 177]. Dissenters [482]. Distant [2541, 3796]. Distinctions [2795]. Distinctively [2677, 2685]. Distinctiveness [1807]. Distinguished [3890]. Distribution [1894, 4099, 1259]. Distrust [4104, 4105]. Diverse [3073].

Diversifying [2328]. Diversity [3629, 2117, 292, 2003, 2569, 1590, 2638, 4219, 480, 3079, 3628, 2339, 3673, 2086]. Divide [3899, 3493, 4126, 3505, 4037, 3585]. Divided [3155, 4045, 2707]. Divides [3961]. Divine [3842]. Divinity [3381]. Division [655]. Divisions [3505, 3837]. DNA [884, 3058, 883]. Do [4095, 3883, 3863, 3007, 645, 3003, 1718, 3935, 3995, 3395, 1149, 3795, 2850, 3962, 3651, 3908, 3065, 3198, 4087, 2304, 1903, 1138, 1048, 1522]. Doctor [207, 112]. Doctors [1887]. Doctrine [1633, 3530, 1665, 1582]. Document [160]. Documents [3760, 17, 9, 156]. Does [3890, 3851, 3446, 2857, 3590, 3713, 2790, 3848, 3302, 4134, 3366]. Dog [3322, 1893]. Dogma [1631, 884]. Doing [4157, 1332, 2138, 3576, 1340]. Dolce [357]. Dole [3892]. Domain [2874]. Domestic [220, 1396, 2017, 562]. Dominance [1855]. Domination [3577]. Donald [3960]. Don’t [3863, 645, 3183, 3003, 2863, 3462]. Doomed [2287]. Door [3184, 930]. Dostoevsky [374, 375]. Double [2311, 373]. Double-Edged [2311]. Doubles [374]. Doubt [1213, 1337]. Doubts [1478, 2795]. down [2906, 1171]. Downs [1484]. Dr. [1283, 39]. Drain [3630]. Drama [3547]. Dramas [4023]. Dramatic [1441]. Draughtsman [15]. Drawing [389]. Dream [1436, 1665, 3485]. dreamed [3485]. Dreaming [1803]. Dreams [2502, 3482, 1588, 3481, 1085, 2756]. Dreamworld [1953]. Dred [3067]. Drivers [3402]. Drives [4243]. Drone [3779]. Drop [3066]. Drug [2634, 2274, 2275, 2276, 2268, 2272, 2269, 2267, 2270, 2271, 2080]. Drugs [2883, 798, 2266, 2265, 802]. Dryden [963]. Due [4080]. Durable [3597, 3936]. During [4156, 2377, 3379, 729, 2653]. Dutch [319]. Dyeing [1858]. Dying [2663, 3286, 2328, 3659]. Dynamic [3733, 2426]. Dynamics [1386, 2074, 2709, 2781]. Dysfunctional [3604].

Ear [3709]. Early [1905, 1359, 724, 2644, 941, 2643, 943, 325, 946, 3786, 949, 1888, 720, 2647, 1928, 2299, 2645]. Earnings [2420]. Earth [1274, 3303, 3736, 3735]. East [2748, 360, 2695, 3189, 2392, 3844, 1562, 981, 2574, 3167]. Eastern [2111, 3277, 1474, 2589, 704, 2122, 2623, 2588, 2114]. Easy [3491]. Eating [3287]. Eats [3476]. Ebb [2629]. Eckstein [1795]. Ecological [4075, 2844, 2847]. Ecology [128, 673, 2523, 1194, 2843, 2837, 880, 2846, 2836, 3442]. Economic [3226, 992, 1597, 125, 3268, 3250, 1592, 1548, 1076, 2804, 924, 2432, 798, 4159, 1246, 1878, 527, 528, 2420, 2767, 3251, 1253, 381, 910, 731, 721, 3844, 1025, 3404, 1258, 676, 379, 2053, 1648, 2152, 413, 2681, 3962, 2803, 1881, 1064, 1649, 2409, 1360, 2247, 3056, 720, 3037, 1525, 678, 563, 530, 1708, 730, 3435]. Economic-Technological [2247]. Economics [2555, 3596, 3953, 3022, 2120, 2554, 347, 137, 3963, 1781, 2842, 2873, 2553, 3123, 3273, 1387, 239, 680, 679]. Economies [1470, 3833, 1653]. Economists [4161]. Economy [4159, 4143, 4168, 2222, 4147, 4150, 1810, 2058, 2354, 4156, 2848, 3664, 2820, 124, 4164]. Ecos [2849]. Ecosystem [685]. Ecosystems [3577]. Ecumenism [2777]. Eden [1981]. Edge [3746, 3111]. Edged [2311]. Editing [188]. Editor [1691, 1225, 3017, 210]. Editors [207, 3151]. Edmund [1273]. educated [3445]. Educating [2926]. Education [827, 2421, 708, 1723, 2135, 1991, 305, 2808, 1651, 1603, 1179, 1727, 3955, 836, 86, 933, 81, 1444, 2909, 1147, 423, 1198, 3863, 311, 3864, 84, 1798, 2765, 1046, 2688, 87, 1992, 1605, 3945, 1145, 1188, 2467, 2469, 2603, 711, 848, 1726, 1344, 1891, 1144, 1201, 1178, 2244, 2725, 1124, 1990, 2455, 2656, 2466, 1126, 1797, 935, 1172, 1609, 1176, 3949, 511, 1137, 1204, 2457, 82, 2464, 307, 2055, 1714, 3608, 2470, 1193, 2014, 801, 1149, 3674, 1196, 513, 2465, 2472, 1792, 92, 309, 3993, 1164, 1185, 22, 1148, 1209, 2903, 1876, 89, 687, 1190, 1146, 1728, 85, 88, 2088]. Education [710, 2895, 2661, 871, 258, 1713, 518, 2459, 1081, 83, 1152, 2475, 2059, 1202, 901, 2500, 1207, 841, 1189, 2610, 1602, 2299, 1156, 96, 510, 1812, 2667, 4202, 1551, 1187, 467, 1197, 2302, 93, 3445]. Educational [1151, 1552, 2625, 1125, 1793, 425, 1159, 2658, 686, 650, 2657, 517]. Edward [1294, 1308, 1295, 1312, 55]. Effect [3851]. Effective [2088]. Effectiveness [2507, 2407, 3946]. Effects [1628, 2023, 4239, 1731, 676, 383, 4076, 258, 4176]. Effervescence [2262]. Efficiency [2407, 2354]. efficient [3466]. Effort [848, 2460]. Efforts [4219, 652]. Egalitarian [1503, 4149, 2858]. Ego [370]. Egypt [2211, 2447]. Egyptian [3427]. Eight [756]. Eighteenth [280, 974, 905, 729, 35, 725]. Eighteenth- [974]. Eighteenth-Century [280, 905, 35]. Eighty [2387]. Eighty-Nine [2387]. Einstein [2954, 732, 4247, 2616, 2991]. Eleanor [443]. Elected [11]. Election [3096, 3333]. Elections [2754, 3342, 3343, 4139]. Elective [3247]. Electoral [2257]. Electorate [3586]. Electric [3141]. Electrical [2521]. Electronic [1696, 1778, 1673, 1698]. Elektron [2521]. Element [754]. Elementary [2300, 1371]. Elements [3700]. Elevated [4243]. Elgin [2925]. Elite [1545, 2683, 3837]. Elites [1927, 2394, 1140]. Elitism [1143]. Ella [3930]. Ellison [3375, 3485]. Else [1814]. Elusive [1597, 3267, 2599]. Emancipation [3087]. Embattled [840]. Embodiment [3176, 2489]. Embrace [1407]. Embryo [3283]. Emergence [1890, 1677, 2051, 4124]. Emergent [1981]. Emerging [2192, 2836, 3442]. Emerson [143, 3129]. Emigration [3630]. Emile [1444]. Emilia [2830]. Emissions [3590]. Emotion [3177].

Emotional [4039]. Emphasis [744].

Empire [3753, 3087, 1076, 3083, 1294, 3515, 1113, 3084, 3086, 2388, 1300, 3762, 1295]. Empirical [1709]. Empiricism [3907]. Employment [526]. Empowering [4227]. Enabling [3680]. Enacted [3547]. Enclosure [2874]. Encounter [1986]. Encouraging [283, 1187]. Encyclopedia [2869]. End [2819, 2334, 991, 3977, 2960, 2226, 2523, 1544, 3669, 728, 4007, 2387, 3422, 2531]. Endangered [4019, 4204]. Endangers [3612]. Endeavor [4115]. Endgame [2254, 3015]. Ending [3846]. Endless [997, 1423]. Endowment [1932]. Ends [1793, 82]. Enduring [1179, 3790]. Enemy [2887]. Energetic [3709]. Energy [3562, 3600, 3599, 3173, 3597, 3593, 3598, 3566, 3405, 3726, 3558, 3588, 3592, 3406, 3560, 3559, 3594, 3595, 1549, 1257, 3564, 2519, 3206, 2520, 3403, 3565]. Enforcement [222]. Engagement [3721, 3107, 3612]. Engaging [3360]. Engineer [1269]. Engineering [311, 3097, 2436, 1528]. Engineers [3890]. Engines [3702]. England [1456, 2217, 1620, 3304, 727]. English [2556, 1699, 3841, 1000, 2558, 900]. Englishness [1699]. Enhance [3408]. Enhancing [3847]. Enigma [1752]. Enlightenment [1536, 3547, 1630]. Enough [3849, 1586, 2022]. Enrichment [3429]. Entailments [3790]. Entangled [4205]. Entanglement [4173]. Enterprise [2094, 2345, 2726, 380, 779]. entre [3656]. Entrepreneurial [2499]. Entrepreneurs [2501]. Environment [2515, 2168, 1046, 2056, 673, 345, 2402, 672, 2513, 3738, 2517, 3303, 2520, 755, 2525, 2405, 3736, 517]. Environmental [2839, 3596, 2840, 2845, 3677, 4004, 3762, 676, 2838, 687, 3739, 3577, 2986, 4064, 3868]. Environmentalism [3295]. Environments [3733, 3229, 520, 4069]. Envisioning [2848]. Envy [3051]. Epidemic [726, 2069, 2079, 3576, 2066]. Epidemics [3307]. Epidemiologic [3799]. Epidemiological [2404]. Epidemiology [2074]. Epilogue [1432, 1070]. Epistemology [1491]. Epitaph [1139]. Epoch [3042]. Eppuor [452]. Equal [1757, 3848]. Equality [1597, 3937, 1208, 3959, 2420, 3479, 1326, 1552, 549, 3496, 3961, 3768, 3969, 2860, 436, 2004, 1437, 1759, 3970]. Equitable [2468]. Equity [3303, 4219]. Era [991, 2419, 1891, 2236, 3669, 1948, 200, 3489, 2387, 4105, 1696, 3972, 3604, 2662, 2227, 2531, 3750, 917, 4038, 2143, 2148, 2155]. Erich [2367]. Erik [361]. Erikson [361]. Erosion [1129]. Eroticism [993, 2012]. Errand [603]. Erudition [1309]. Escalation [3980]. Escaping [3006]. Essay [2488, 1662, 856, 511, 2219]. Essence [2635]. Essential [3720]. Established [482]. Establishing [1701, 1548, 2972, 484]. Establishment [1637]. Eternity [3652]. Ethical [4178, 3271, 2839, 1335, 3781, 789, 788, 4090, 2728, 802, 3952, 1220, 803]. Ethicist [2727]. Ethics [121, 2840, 2845, 2730, 2729, 4182, 3861, 3283, 3055, 3742, 3780, 1856, 51, 3783, 3284, 3999, 3778, 4098, 1222, 4003, 3441]. Ethiopian [4064, 3361]. Ethnic [253, 252, 2587, 3218]. Ethnicity [2122, 1326, 3070, 2763, 2304, 1408, 1647]. Ethnography [2728]. Ethnology [1536]. Ethos [576]. Eudaimonic [4154]. Eugene [1703]. Eugenics [505, 2005]. Eupsychia [1436]. Europe [2118, 408, 1467, 2111, 2374, 2581, 409, 410, 1469, 2381, 423, 416, 1474, 2435, 2379, 414, 704, 1905, 1967, 428, 411, 2122, 431, 425, 471, 583, 2375, 2113, 2283, 2185, 426, 2260, 1482, 1475, 2582, 2328, 2396, 413, 2572, 1472, 1577, 2818, 279, 417, 1250, 2817, 1486, 421, 2121, 2391, 2593, 2376, 2590, 2114, 2382, 420, 429, 2397]. European [1478, 1786, 3849, 1470, 424, 418, 3190, 3929, 384, 415, 2392, 1473, 2184, 2450, 1025, 2387, 3326, 412, 994, 419, 389, 2224, 702, 2285, 422, 2591, 2378, 2742]. Evacuation [4195]. Evaluated [642]. Evaluating [3797, 1342, 3037, 3795]. Evaluation [1151, 3898, 1069, 1240, 3946]. Evasion [4180, 1489, 4187]. Eve [196]. Evening [1407]. Events [1136]. Ever [2787]. Every [959]. Everyday [2729, 4059, 2389, 2793, 2941]. Everyman [959]. Everyone [4112]. Everything [1754, 2787]. Everywhere [3574]. Evidence [70, 73, 72, 737, 3114, 3898, 71, 3895, 4242, 2638, 74, 69, 3963, 4233]. Evidence-Based [3114]. Evil [1132, 2853, 2890]. Evolution [1519, 265, 1951, 4044, 3700, 415, 264, 2035, 52, 2716, 1376, 2636, 267, 273, 3177, 197, 3633, 1661, 270, 268, 1156, 1681, 3704]. Evolutionary [3232, 1534, 2827]. Evolutions [272]. Evolving [4080, 3097, 1475]. Exacerbate [3949]. Example [727]. Excellence [285, 281, 1726, 278, 283, 287, 279, 3866, 284]. Exceptional [2228]. Exceptionalism [2164, 3354, 3549]. Exceptions [3549]. Excerpt [4192, 3647]. Excerpts [815, 820, 3654, 129, 3648, 3382]. Excess [4070]. Excessive [3683]. Exchanges [2944]. Executive [532, 3920]. Exemplar [1897]. Exemplary [1615, 1617, 1619]. Exemptions [3992]. Exercise [2712, 2795]. Exhibition [2708]. Exhibitions [2709]. Exile [3094, 3111, 1645]. exiles [4196]. Existence [1451]. Existentialism [455, 457, 446]. Exit [2261, 2250, 3580]. Exoplanets [3701]. Expand [3724]. Expanded [4142]. Expanding [3785]. Expansion [1185, 321]. Expansive [4115]. Expectations [1643, 2659, 3344, 257]. Experience [3561, 1151, 3943, 1306, 3180, 1095, 608, 3088, 4222, 3373, 2484, 2728, 1521, 2152, 2662, 1679, 3058, 3499, 1661, 2242]. Experiences [4059, 2411]. Experiential [4120]. Experiment [3180, 711]. Experimental [3396, 1905, 1891, 500, 681].

Experimentation [791, 796, 800, 789, 793, 788, 799]. Experimenting [790]. Experiments [803]. Expert [386, 4117]. Expert/Audience [4117]. Expertise [3896, 401, 3946, 3377]. Experts [3897]. Explained [2322]. Explaining [3787]. Explanation [972]. Explanations [4221]. Explicating [3758]. Explicit [284]. Exploding [1372]. Exploitation [1591, 915]. Exploiting [1540]. Exploration [3568, 2712]. Exploratory [2219]. Exploring [1236]. Export [2227]. Exposure [608]. Expression [150, 246, 3058]. Extending [1431]. Extensions [4170]. External [2435, 3811, 1590]. Extra [4227]. Extrasolar [3046]. Extremist [3272]. Eye [1819, 2036, 1862, 2926]. Eyes [1000]. Eyewitness [3893].

Fabulous [3696]. Face [927, 3389, 3791]. Faces [2866, 4077]. Facets [1218]. Facing [2772, 3137, 2338, 425, 3608, 2475]. Fact [945, 1676]. Factor [3955]. Factors [2210, 529, 1033, 720, 519]. Facts [3723]. Faculty [1160]. Fail [2287]. Failed [1139, 2309, 4046, 3850, 1554]. Failing [2272, 2273]. Failure [1135, 1605, 3938, 3885]. Failures [2946]. Fair [3332]. Fairest [4244]. Faith [1859, 608, 2122, 1635, 1108]. Faiths [1478]. Fall [1889, 3380, 1294, 1299, 3762, 3374, 1664]. Fallacies [379]. Fallacy [682]. Fallen [3892]. Falling [2577, 2273]. Fallout [55]. False [3951]. Familiar [3235, 2578]. Families [4041, 2297, 3910, 188]. Family [1366, 1362, 3166, 950, 1359, 1363, 1354, 1873, 1357, 3496, 1356, 2832, 2790, 526, 552, 1360, 1364, 1358]. Fancy [3643]. Fanon [1274]. Fantasies [2619]. Farewell [13]. Fascist [917]. Fashioned [822]. Fatal [2271]. Fate [4142, 583]. Fates [3160]. Father [3482, 1463]. Fatherless [1364].

Fats [3644]. Fault [1545, 1215]. Faustian [3025]. Faustus [112]. Favor [3777]. Fear [2618, 3119, 2946, 443, 1718, 2930, 3312, 2271]. Fearful [2258]. Fears [2163]. Feasibility [216]. Feasible [3250]. Features [1359, 2657]. February [1847, 663]. Fecundity [729]. Federal [2033, 804, 1201, 532, 521, 2327, 918, 647, 2346, 3335]. Federalism [3149, 2759, 1403]. Federation [2587]. Federico [357]. Feed [3742, 3736]. Feel [3640]. Feeling [4019, 1332, 2138, 2658, 3646, 1340]. Feelings [3713]. Fellini [357]. Fellows [3970]. Female [3217, 2619, 3216, 2796]. Feminism [3960, 1959, 2790]. Feminist [1963]. Feminists [3966]. Fencing [2874]. Ferl [2367]. Ferment [2777]. Fertility [724, 721, 119, 725, 723]. Few [1620, 2832]. fewer [3458]. Fiasco [3619]. Fiction [2100, 274, 586, 3446]. Fictional [3836]. Fidelity [292, 2003]. Fiedler [358]. Field [1621, 795, 21, 1384]. Fields [4116]. Fifth [1301]. Fifty [4129, 2420, 2551, 885]. Fight [4242, 3886, 3970, 888]. Fighting [3965, 3874]. Figure [1495]. Figurine [3656]. File [3428]. Filipino [589]. Fillmore [1724]. Film [1859, 32, 159]. Filming [1860, 3518]. Films [2118]. Filthy [1659]. Fin [2603]. Finance [2501, 3469]. Finances [239]. Financial [3953, 4039, 3143, 3473, 3464, 3470, 3467]. Financing [4066, 1348, 3516, 928, 3565]. Finding [4018, 3005, 3604]. Findings [3815, 2313]. Fine [20]. Fingerprint [3895]. Finite [1066]. Fire [3082, 3852]. Firewood [3539]. Firm [780]. Firms [4165, 4166, 4167]. First [2241, 62, 1170, 130, 3016, 3824, 1213, 3877, 2881, 3516, 370, 1660, 2796, 2467, 3856, 3503, 784, 2710, 1222, 2720, 2542]. Fiscal [3471, 3372]. Fish [2962]. Fisheries [4062]. Fishing [3857]. Fittest [2603]. Fitzgerald [3930]. Five [2226, 1724]. fix [3468]. Flaubert [375]. Flavors [3273]. Flawed [2657]. Flemings [3218]. Fletcher [1829].

Flexner [1087, 1087]. Floating [1468, 2020]. Flood [1653]. Floor [3642]. Flourishing [4144, 4125]. Flow [1589]. Flows [2953]. Flux [2956, 2540]. Flying [1755, 2024]. Focus [4125]. Foe [3498]. Fog [2578]. Folk [1757]. Folklorist [104]. Following [3925]. Food [3740, 3726, 3742, 3738, 448, 3739, 579, 3736, 3735]. Football [2763]. Foraging [3139]. Force [1341, 3720, 1082, 3781, 2779, 2942, 3516, 3808, 3809, 601, 2927, 4008]. Forces [86, 3835, 1951, 463, 255, 3513, 3522, 3790]. Ford [1929, 1928]. Forecasting [663, 643]. Foreign [1654, 2819, 990, 1238, 335, 2266, 1232, 329, 331, 1738, 562, 2762, 3378, 330, 1554, 332, 1843]. Forensic [2161, 3894]. Foresight [3695]. Forest [2911]. Foreword [3368, 3601, 147, 116, 2456, 525, 2678, 1823, 3510, 193, 1, 12, 19, 28, 211]. Forgotten [1875]. Form [2470, 374, 3504, 1677, 1378, 163]. Formation [1296, 3704]. Formed [2717]. Forming [2717]. Forms [2573, 1951, 240, 1013]. Formula [2394]. Fortune [1753]. fortunes [3469]. Forty [2671, 2260]. Forward [3983]. Fossils [3574]. Fostering [308]. Found [2390]. Foundation [1929, 3351, 2460, 1928]. Foundational [4078]. Foundations [4142, 1042, 1927, 2288, 3372, 465]. Founder [740]. Four [664, 3094, 2551, 379]. Fox [4055]. Fractionating [454, 453]. Fragile [3101, 1692, 4069]. Fragility [4161, 3836, 3467]. Fragmentation [3532]. Fragments [2878, 1468, 2020]. Frames [1907]. Framework [791, 2276, 641, 1713, 1240, 3876]. Frameworks [3233, 793]. France [726, 23, 430, 905, 1359, 2788, 729, 1483, 2590, 300]. Franchise [1325]. Francis [21, 884]. Frank [1088, 326, 39]. Frankfurt [2355]. Franklin [24, 23]. Frantic [462]. Frantz [1274]. Fraternity [254, 1326]. Fraught [4004]. Fraying [1816]. Frazer [1460]. Free [125, 3587, 3413, 3414, 2795, 644, 380]. Freedom [4123, 3745, 1425, 3082, 2946, 2788, 3479, 3039, 2878, 2972, 1635, 2095, 22, 1424]. Freedoms [329]. Freeing [2519]. French [2543, 3073, 1941, 2326, 915, 1413, 1298, 1309]. Freud [1907, 113, 3662, 1085]. Freudian [40]. Friend [3548, 3498]. Friendly [2727]. Front [10, 18, 27, 33, 47, 59, 68, 79, 98, 115, 132, 145, 172, 191, 209, 234, 251, 263, 277, 289, 303, 316, 328, 341, 363, 377, 391, 406, 433, 445, 459, 473, 492, 508, 523, 542, 558, 574, 585, 596, 612, 632, 689, 671, 713, 734, 750, 772, 694, 786, 806, 829, 838, 852, 866, 876, 891, 907, 922, 940, 955, 967, 977, 988, 1002, 1018, 1035, 1055, 1072, 1092, 1106, 1123, 1174, 1212, 1228, 1243, 1264, 1281, 1293, 1315, 1330, 1353, 1368, 1381, 1393, 1415, 1434, 1449]. Front [1465, 1477, 1488, 1499, 1513, 1530, 1543, 1557, 1570, 1584, 1600, 1612, 1625, 1641, 1657, 1670, 1685, 1706, 1720, 1735, 1750, 1763, 1775, 1788, 1800, 1821, 1835, 1852, 1865, 1886, 1901, 1912, 1924, 1934, 1945, 1956, 1971, 1989, 2001, 2026, 2045, 2064, 2076, 2090, 2109, 2125, 2140, 2159, 2176, 2195, 2207, 2221, 2234, 2249, 2264, 2278, 2290, 2306, 2320, 2336, 2349, 2373, 2385, 2399, 2415, 2429, 2441, 2454, 2480, 2497, 2512, 2527, 2549, 2568, 2580, 2595, 2613, 2627, 2641, 2655, 2669, 2693, 2704, 2722, 2738, 2676, 2752, 2769]. Front [2785, 2799, 2814, 2824, 2835, 2852, 2872, 2894, 2917, 2933, 2949, 2966, 2988, 3002, 3019, 3034, 3049, 3064, 3081, 3099, 3116, 3133, 3154, 3172, 3192, 3208, 3228, 3246, 3263, 3280, 3294, 3310, 3329, 3348, 3365, 3386, 3399, 3418, 4103]. Fronteras [4187]. Frontier [1475, 844, 1423]. Frontiers [2391, 1784]. Frugal [3571]. Fruitful [2882, 2818]. Fuel [3420, 3742, 3422, 3427]. Führerprinzip [3776]. Full [535, 3735]. Function [7, 1610, 3918, 1378]. Functional [2632, 569]. Functioning [4090]. Functions [3706, 846, 3847, 1779]. Fundamental [3572]. ’Fundamentalism [2970]. Fundamentalists [2211]. Funding [3686, 3363]. Further [669, 1518, 4032, 2332]. Fusions [3926]. Future [808, 4041, 2711, 636, 3600, 3572, 990, 1914, 2592, 3508, 933, 3812, 924, 4079, 3807, 2537, 1400, 3558, 3588, 3719, 236, 807, 923, 1124, 1913, 2430, 2753, 565, 2953, 1176, 3036, 3949, 3201, 3136, 4223, 4121, 1199, 927, 2980, 3718, 3964, 3404, 2185, 4053, 3445, 4236, 2606, 2762, 3748, 774, 2042, 3428, 113, 3494, 2681, 484, 4219, 3894, 929, 2868, 3497, 3206, 3591, 937, 506, 4133, 1649, 1486, 660, 1741, 4089, 3519, 644, 2382, 2190, 928, 3069, 3362, 3696, 1405, 1197, 3432, 3437, 3434, 3438, 3444]. Futures [637, 664, 4113, 3856, 767, 1946, 3860, 3430, 3858, 4042, 3446]. FY [1847].

Galaxies [3704]. Galicia [3277]. Gallery [2355, 2358, 2361, 2364, 2367, 2369]. Games [4027]. Gandhi [737, 3212]. Gap [2120, 90, 3497]. Gaps [1454, 3739, 3666]. Gardes [3214]. Gas [3590, 3563]. Gassett [1083]. Gastrosophy [993, 2012]. Gateways [3624]. Gaulle [741]. Geertz [3261]. Gender [1968, 4207, 1815, 1958, 1966, 3126, 3959, 1962, 1957, 3961, 3964, 3522, 3962, 1874, 3970, 4184, 2425]. Gender-Based [4184]. Gene [3058]. Genealogies [962]. General [2555, 1806, 518, 3379]. Generality [1260]. Generals [4140]. Generate [3055]. Generating [2687]. Generation [1642, 2694, 2660, 3934, 3936]. Generational [1459, 1452, 1451, 1457, 3505]. Generations [1456, 295, 2129, 1450, 1458, 1454, 440]. Genes [3229]. Genesis [62]. Genetic [3861, 267, 266]. Geneticist [2859]. Genetics [3232, 1716, 3201, 882]. Genital [2796]. Genre [4025, 2485]. Geoffrey [3203].

Geography [2964]. Geological [1376]. Geopolitical [4210, 3975]. Georg [2364]. George [1615, 1089, 3038]. Geri [3925]. Geriatric [1877]. German [2284, 2695, 2363, 2354, 2359, 2368]. Germans [2357, 2366]. Germany [2351, 2356, 2360, 2788, 2365, 2327, 2362, 2363, 2354, 2546, 1118, 2352, 918, 2353, 1483, 2590, 2378, 2370]. Gesta [861]. Get [2464, 2553, 3123]. Gets [2862]. Getting [4083, 2389, 3876]. Ghana [698, 4059, 4062]. Ghetto [767]. Ghosts [1359, 1133]. Giacometti [3656]. Gibbon [1294, 1300, 1313, 1301, 1296, 1306, 1304, 1297, 1310, 1299, 1308, 1311, 1295, 1302, 1312, 1305, 1307, 1303, 1298, 1309]. Gideon [3682]. Giedion [1279]. Gift [3960]. Girl [3548]. Girls [942]. Given [1746]. Glasper [3645]. Glass [2374]. Glasses [451]. Global [3217, 3723, 2839, 2959, 1232, 2222, 4114, 3370, 4056, 3738, 2934, 3832, 3409, 3504, 3739, 3206, 3421, 3031, 2941, 3043, 3984, 3577, 2223, 3734, 4009, 3831, 2742, 2086]. Global-Scale [3734]. Globalism [2452]. Globalization [2829, 2854, 3675]. Globally [3319]. Glory [2334]. Glossary [275, 1824, 1837]. Glück [3658]. Gnosis [1109]. Go [2379, 2850, 3050, 3065, 1138]. Goal [1058, 3267, 1608, 1738]. Goals [1210, 212, 756, 1125, 555]. God [1537]. Goes [3857]. Going [2457, 1462, 1899]. Gold [3291]. Golden [4086, 1796]. Good [2054, 3849, 3103, 2733, 2729, 3609, 3931, 3603, 3617, 3104, 3616, 503, 3649, 4228, 3604, 4108, 3795, 372, 3351, 3563, 3970, 3615]. Goodheart [1703]. Goodman [2355, 2358, 2361, 2364, 2367, 2369, 1050]. Google [3359]. Gorbachev [1936]. Gordian [1644]. Got [1700, 3962, 4043, 2802]. Gotha [1270]. Govern [4164]. Governance [1210, 4144, 834, 835, 3849, 4171, 3677, 3676, 3844, 2936, 846, 4067, 3847, 3604, 4108, 2758, 3224, 2941, 4172]. Governed [4138].

Governing [4182, 3270]. Government [1760, 1550, 761, 2054, 225, 1345, 776, 226, 4138, 2053, 4132, 2389, 2759, 2697]. Governmental [804]. Governments [3592, 935, 2162]. Governor [1806]. Governor-General [1806]. GPT [4094]. GPT-3 [4094]. Gr [2369]. Graduate [1201, 1161]. Grammar [3660]. Gramsci [2907]. Gran [4187]. Grand [3738, 4064]. Grandeur [741]. Grandmother [3110]. Grass [1589]. Grassroots [4118]. Gray [130]. Great [4180, 3974, 4177, 549, 2031, 202, 1079, 3585, 2285, 2590, 4102, 3089, 6, 2620, 3454, 3474]. Greatest [3410]. Greco [3766]. Greece [2956, 3751]. Greek [1221, 3316, 1451, 1271]. Greeks [1215]. Green [3291, 1145, 3044]. Greenhouse [3590]. Greenlaw [3556]. Grief [2490]. Grin [3688]. Groove [3645]. Ground [2465]. Grounded [2845]. Grounds [3937]. Group [515, 4159, 259, 4090, 253, 618, 627, 255]. Groups [3201, 252]. Grow [4039]. Growing [3938, 3663, 1322, 2304]. Grows [3142, 1589]. Growth [1068, 241, 1065, 2294, 1058, 3251, 1056, 3669, 1067, 3402, 725, 3514, 1185, 1069, 1061, 39, 1064, 2409, 1525, 942, 519, 1063]. Grudging [2268]. Grutter [3067]. Guarantees [3420]. Guarantors [3974]. Guernica [169]. Guestworkers [1483]. Guide [3308, 4196]. Gulf [1561]. Gun [4217, 4073]. Gunnar [2419]. Guns [4074].

H [361, 3]. H. [577]. Habitable [3568, 1994]. Hail [4145]. Hail-and [4145]. Half [3653, 662, 2559]. Hall [2564]. Hamburgerstrasse [2369]. Hamilton [3475]. Hand [3795]. Handled [1033]. Happen [3935]. Happened [2764, 1539]. Happening [3358, 3443]. Happens [764]. ’Happines [3023]. Happiness [3024, 3022, 3027, 4146, 3025, 3028, 3020]. Harare [4057]. Harassment [3967]. Hard [3712, 3987, 1180, 4161]. Hardship [3810].

Harm [3795]. Harry [1795]. Hart [3059]. Hate [3066, 4040, 4212]. Hawaiian [3860]. Hawks [2267]. Hazards [215, 350]. Healing [4077]. Health [1877, 4183, 4178, 4215, 1351, 2054, 4130, 2168, 1335, 1879, 2410, 2733, 1878, 4063, 2034, 2402, 4182, 2862, 2663, 2401, 1343, 1332, 2400, 4226, 3861, 1347, 4221, 4223, 2171, 2411, 3667, 3370, 3718, 1428, 2407, 3738, 2167, 1348, 4228, 3047, 2403, 2272, 4060, 2413, 2406, 4185, 2084, 2345, 4216, 4181, 4179, 4217, 2405, 1379, 3735, 3866, 4073, 1340, 2412]. Health-Care [1878]. Healthy [3720, 3078]. Hear [3870]. Hearing [2647]. Heart [182, 3060]. Hedgehog [4055]. Heidegger [3165]. Heller [356]. Héloïse [1268]. Hence [1182]. Henri [1077]. Henry [1081, 4021, 1081, 4020]. Here [3643, 1468, 2020, 2464, 2850, 3065, 1138, 3305]. Heritage [200, 2031, 2545, 266, 1102]. Hero [200]. Heroes [1804]. Heroine [442]. Hibbing [3239]. Hierarchy [4117]. High [1606, 1722, 1605, 4168, 1618, 4170, 1615, 3239, 1607, 2156, 95]. High-School [1607]. High-Tech [4168, 4170]. Higham [359]. Higher [827, 708, 1651, 1179, 836, 933, 3864, 1188, 848, 1144, 1178, 1124, 1126, 935, 1172, 1125, 1176, 1137, 1193, 2014, 3674, 1185, 1209, 1190, 1146, 710, 871, 1152, 1207, 841, 1189, 2610, 4202, 1187, 1197]. Highland [1617]. Highlands [4065]. Hill [3203]. Him [1279]. Hindu [2790, 2844, 2333]. Hip [3504]. Hip-Hop [3504]. Hirohito [2141]. Hiroshima [383]. Hispanic [3718]. Hispanics [3068, 1882]. Historian [1100, 2532, 908, 863, 1295, 4013, 1307, 912]. Historians [1304, 1094]. Historical [718, 917, 3508, 101, 775, 737, 1932, 2401, 531, 3125, 3478, 714, 893, 863, 1299, 4221, 253, 1318, 1357, 717, 3940, 719, 981, 386, 1391, 2406, 918, 1664, 197, 2760, 2128, 715, 2066, 3624, 930, 3042, 730]. Historically [4063]. Historicism [1278]. Historien [1312]. Historien-Philosophe [1312].

Historiography [1377]. History [170, 1370, 359, 70, 832, 1536, 1076, 1782, 3004, 2702, 3358, 2506, 3979, 4079, 3205, 911, 4025, 3161, 1663, 859, 1082, 1591, 1732, 2543, 903, 1826, 1277, 902, 1310, 2557, 897, 2771, 3922, 899, 2957, 3487, 910, 858, 1311, 895, 2583, 3188, 1452, 909, 716, 3040, 894, 299, 2162, 313, 495, 904, 961, 2011, 3157, 2224, 3764, 722, 2008, 1999, 914, 881, 3759, 3067, 1103, 2726, 2200, 898, 1964, 880, 3241, 3500, 900, 901, 913, 1378, 3866, 916, 3734, 1044, 1216, 1384, 2621, 3120, 1358, 3160, 361]. Hit [3931]. HIV [2074]. Hobbes [1267]. Hoffman [1829]. Hold [4004]. Hole [4021]. Holes [1372, 3703]. Holistic [1938]. Holland [1494]. Hollow [3186]. Hollywood [3641, 3642]. Holocaust [2486, 2491]. Holy [603]. Homage [3930, 2720]. Home [1724, 2282]. Homeland [2329]. Homelands [2435]. Homer [3756, 3104, 1931]. Hominem [157]. Homo [397]. Homophobia [3638]. Hooks [3661]. Hop [3504]. Hope [1568, 2115, 2928]. Horizon [1578]. Horizons [311]. Horse [3791]. Host [4066]. House [1115, 2707, 2842, 13]. Household [3593]. Housing [548, 1363, 3472, 549, 4164]. Howe [2961]. hsi [2199]. Hub [2809]. Huizinga [1078]. Human [4180, 3137, 3053, 3955, 4144, 2937, 3320, 791, 56, 36, 3054, 4101, 1716, 1532, 1737, 265, 2859, 804, 800, 1743, 2073, 789, 793, 3039, 788, 1036, 1736, 2569, 4226, 160, 620, 1742, 3995, 3395, 1748, 1738, 790, 3057, 2517, 3282, 3697, 2638, 1740, 799, 4091, 2049, 3209, 795, 1862, 4065, 3299, 794, 4125, 652, 4094, 3058, 1739, 1785, 3389, 1741, 803, 518, 2563, 3056, 3315, 2525, 1379, 3577, 1747, 2964, 3734, 4009, 1744, 3042]. Human-Like [4101]. Humanism [3162, 2847, 3000]. Humanist [43, 861, 1307, 818]. Humanistic [833, 853, 862, 814, 4098, 826, 854]. Humanists [3352, 3651, 3394]. Humanitarian [4178, 4182, 4177, 3832, 4176, 4179]. Humanitarians [1756]. Humanities [808, 4123, 3353, 3163, 4120, 4115, 855, 1143, 84, 469, 4113, 4126, 3355, 821, 3352, 807, 1721, 2896, 1020, 4114, 4121, 857, 4112, 4122, 3157, 42, 3360, 3354, 809, 864, 4125, 4118, 4116, 812, 2009, 3361, 4119, 2690, 3156, 3350, 4111, 4124, 1022, 3362, 3363]. Humanorum [861]. Humans [796, 3390, 3392, 2524, 3394, 4124]. Hume [1307]. Humor [1297]. Hundred [2525]. Hungary [724, 2436]. Hunters [2880]. Hurts [2730]. Hydropolitics [4065]. Hyper [684]. Hyper-Productive [684]. hyperincarceration [3455]. Hypocrisy [1495, 1489, 1496]. Hypocrites [1492]. Hypocritical [1490]. Hypotheses [325]. Hypothesis [3734].

Iconoclasm [1938]. Idea

[945, 634, 1306, 2947, 2142, 2504, 2875, 1452, 3157, 3296, 382, 2718, 1855]. Ideal [1898, 2925, 1327]. Idealism [2938, 3610]. Ideals [1020, 2687, 184, 1503, 380]. Ideas [2874, 1889, 3807, 381, 746, 1376, 1453, 1276]. Identification [3893, 255]. Identifying [4176]. Identities [2257, 3197, 2802, 2203]. Identity [3194, 1593, 744, 3193, 2592, 109, 3864, 3761, 4159, 2201, 1587, 551, 879, 4160, 2365, 3861, 3201, 3769, 2679, 471, 2375, 2565, 3195, 2648, 618, 1814, 3196, 2331, 2762, 2359, 627, 552, 2793, 2200, 3197, 2310, 1909, 2098, 2333, 2204, 2651]. Identity/Difference [2793]. Ideological [1746, 2491]. Ideology [3807, 1807, 437, 1084, 2051, 380, 203, 1858, 105]. Idols [1020]. If [3869, 3944, 4085, 3488]. Ignorance [1908, 3333, 3689]. Ignore [2473]. II [3, 2077, 835, 1094, 1363, 1572, 1836, 1176, 820]. III [3336]. Iliad [3756]. Ill [1144]. Ill-Tempered [1144]. Illegality [3631, 4034]. Illegitimate [2697]. Illiberal [3966]. Illicit [3833, 3880]. Illiteracies [2130]. Illiteracy [1710]. Illness [4217]. Ills [1061]. Illusion [3271, 3247, 1489, 3036, 2980, 1665, 1493]. Illusions [3078]. Illustration [2027]. Illustrations [830, 2642, 3781, 4109]. Illyrianism [2584]. I’m [1814, 2589, 2588, 3657]. Image [1485, 1853, 1858]. Images [1869, 3812, 2296, 1277, 2484, 1863, 257]. Imaginary [1921, 3321]. Imagination [947, 1441, 128, 2618, 3119, 1531, 2510, 3095, 2570, 3128]. Imaginations [2619]. Imaginative [1674]. Imagined [2773, 3337, 4122]. Immigrant [625, 3624]. Immigrants [3632, 4037, 3621, 2327, 2451, 3625]. Immigration [2356, 3627, 3959, 3632, 3620, 4030, 2031, 4031, 3631, 3681, 3619, 3633, 3628, 3622, 3507, 3625]. Immigration-Industrial [3632]. Immobile [2279]. Immodest [2286, 436, 2004]. Immovable [2779]. Immunities [1129]. Impact [2711, 2241, 140, 1363, 3341, 1203, 2411, 1298, 2810, 1525]. Impair [3734]. Impair-then-Repair [3734]. Impartial [3342]. Imperatives [1335]. Imperial [2586]. Imperialism [1095, 3090, 3085, 3089]. Imperium [20]. Implementing [3876]. Implications [241, 220, 3861, 3336, 4065, 3421, 2748, 2847, 3946, 3506, 3363, 3433]. Implicit [4246, 4229, 4237, 4239, 4240, 4241, 4236, 4230, 4235, 4233, 4234, 4243, 4245]. Importance [3579, 719, 2501]. Important [2237, 3491, 3689]. Impossible [1548, 674]. Impressionism [370]. Imprint [3504]. Imprinting [2619]. imprisonment [3452, 3457]. Improve [3685, 3047]. Improved [1517, 2462, 1187]. Improvement [1715, 3945, 1797, 2667]. Improving [2135, 3943, 3898, 3496, 3224, 3914, 530, 3445]. Improvisations [3059]. Inaccurate [3893]. Inadequacy [1547]. Inadvertent [3980]. Incarceration [4039, 3450, 3454, 3448, 3449, 3451, 3459]. Incarnate [3178]. Incarnation [3174]. Incentives [2659, 3676, 1346]. Inclusion [219, 1794, 2327, 2797]. Inclusive [4159, 3852]. Incognita [1802]. Income [1362, 3901, 3938, 526, 4099, 3666]. Incomplete [2750]. Incorporated [2434]. Increasing [4129, 3673]. Independence [1643, 3334, 3337, 1642, 3341, 3331, 3336, 2450, 3335, 1649]. Independent [3860, 3332, 1614]. Index [693, 692, 2672, 2673]. India [2091, 708, 2547, 622, 2106, 322, 1009, 2105, 2746, 2097, 2975, 1010, 2861, 3421, 2993, 2646, 1222, 2098, 2333]. Indian [3863, 2730, 2100, 1859, 2094, 2093, 2092, 1665, 2099, 2104, 2107, 2570, 3128, 3866]. Indian/Alaska [3863]. Indians [1585, 2106, 3863, 1897]. Indies [630]. Indigenous [3869, 3865, 3856, 2849, 4221, 3870, 3858, 3867, 3868, 2766]. Indigenous-Language [3870]. Indirect [3874, 3799]. Indispensable [3726]. Individual [3282, 1338, 299, 4140, 942, 3724]. Individual-Level [4140]. Individualistic [1479, 3995, 417]. Individuality [50]. Individuals [3198]. Indoctrination [259]. Industrial [305, 836, 3632, 3521, 1459, 1653, 1481, 727]. Industry [1198, 2523, 2805]. Inequalities [2733, 3370]. Inequality [2424, 3901, 2854, 2985, 3949, 3938, 3940, 4074, 1042, 2857, 3939, 2892, 3934, 3936, 3818, 3962, 2803, 3941, 2861, 3942, 3916, 3822, 2425, 4202, 2542, 3460, 3450]. Inequities [4221]. Infant [729]. Inference [70, 73, 71, 74, 69]. Inflammation [3141]. Inflation [1653]. Influence [189, 2566, 240, 863, 1114, 1854, 4136]. Influences [35, 1682, 21]. Informal [1691]. Information [80, 146, 235, 290, 342, 407, 474, 543, 597, 690, 787, 839, 892, 956, 1003, 1073, 1175, 1265, 1331, 1394, 1466, 1514, 1571, 1626, 1686, 1751, 1801, 1866, 1925, 1972, 2046, 2110, 2177, 2235, 2291, 2350, 2416, 2481, 2550, 2614, 2670, 2739, 3419, 889, 2530, 1908, 2531, 644, 1672, 1909, 2610, 3442]. Information-Processing [889]. Informationally [3687]. Informed [511, 3445]. Infrastructural [4064]. Inheritance [3482]. Initial [802]. Initiative [1841, 3445, 1828]. Initiatives [2931]. Injury [2638]. ink [3454]. Innately [3055]. Inner [2198, 435, 3005, 1009]. Innocence [991]. Innocent [1440]. Innovating [3685]. Innovation [1639, 1016, 1235, 1861, 3130, 2498, 2516, 3915, 2665, 2809, 1196, 760, 3564, 802, 3571]. Innovations [94]. Innovative [2461, 4049]. Innovator [2780]. Innovators [2716]. Innuendo [4212]. Inquiries [3907]. Inquiry [1127, 3781, 1416, 1531, 1418, 1432, 1426]. Inquisition [1496]. Insanity [387]. Insecure [2438]. Insecurely [3111]. Insecurity [4059]. Inseparable [3726]. Inside-out [1048]. Insights [3561, 4059]. Insoluble [1206]. Inspection [221]. Inspire [3864]. Inspissation [3075]. Instability [1840]. Instance [3354]. Instinct [388]. Institute [2943, 518, 470]. Institution [880]. Institutional [3596, 1951, 569, 871, 1346, 653, 1154, 1183]. Institutionalized [3814]. Institutions [260, 4129, 2243, 3598, 2715, 4131, 4150, 3731, 650, 3679, 3318, 1202, 380, 2827]. Instruction [1127]. Instrument [1727, 2605]. Integration [555, 529, 413, 2590]. Intel [2503]. Intellect [2552, 1532, 1531, 746, 979]. Intellectual [3248, 2241, 973, 974, 3205, 430, 1531, 897, 602, 2875, 1221, 2545, 982, 984, 1768, 2529, 3372, 980, 722, 2008, 975, 1217, 480, 2873, 179, 898, 614, 1938, 2262, 247, 2155]. Intellectualism [3371, 318]. Intellectuals [3807, 974, 985, 969, 968, 978, 981, 979, 122, 986, 1413, 123, 972, 970, 971]. Intelligence [1777, 2243, 4101, 4092, 1978, 1986, 1975, 4104, 1973, 4097, 2866, 1981, 1983, 4105, 1987, 1977, 1980, 1979, 4087, 4084, 1976, 4098, 1984]. Intelligent [4106, 1982]. Intelligentsia [196, 198, 194, 205, 193, 197, 195, 204]. Intensive [3139]. Intention [371]. Intentional [4243]. Interaction [2818]. Interactions [3986, 3734]. Interactive [1596]. Interbranch [3334]. Interdependence [1188, 2189]. Interdependent [1066, 3335]. Interdisciplinary [3934]. Interest [2883, 3268, 3264, 3265, 3267, 3607, 4141, 2529, 3273, 4245, 3270, 3266]. Interests [3346, 2072, 1561, 1565, 1299, 1562, 1563, 2378, 3771, 3377]. Intergenerational [3719, 4075]. Intergroup [3503]. Internal [1590]. International [2611, 4210, 754, 2072, 140, 3473, 3827, 3833, 615, 3853, 1373, 3828, 2762, 1257, 1929, 2187, 568, 2231, 662, 660, 1741, 1233, 3834, 563, 321, 1708, 3439, 1855, 3881]. Internationalism [3090, 2330]. Internationalist [1943]. Internationalization [2347]. Internet [3234, 3535, 3744, 3532, 3746, 3748, 3444, 3533, 3749, 3671]. Interpersonal [1051]. Interpretation [577, 2199, 1085, 106, 1245]. Interpretive [1702, 1779]. Interrupting [3862]. Intersections [4207]. Interview [2101]. Interwar [2572]. Intra [3505, 3503]. Intra-generational [3505]. Intra-minority [3503]. Intractability [3893]. Intractable [2899]. Intravenous [2080]. Intrinsic [2132]. Introduction [808, 1917, 1920, 3388, 4215, 3743, 4110, 81, 2418, 4127, 3767, 2499, 2816, 3525, 3800, 1958, 3887, 3541, 3650, 3634, 3826, 3840, 2644, 3813, 789, 3463, 3432, 1369, 3678, 364, 264, 476, 878, 1382, 174, 2078, 3511, 3956, 148, 2483, 2724, 1333, 3972, 3900, 4230, 3449, 393, 3693, 752, 3725, 4175, 1417, 99, 1057, 3602, 1868, 715, 3999, 736, 3751, 2786, 2820, 2048, 2836, 4042, 2308, 2514, 3868, 4200, 4015]. Introductory [4083, 3662]. Intuitions [3055]. Intuitive [3055]. Invented [2145]. Inventing [4209, 3258]. Invention [3678, 2505, 1812]. Inventions [2315, 2623]. Inventor [2500]. Investigation [798, 1426]. Investigations [3868]. investigative [3437]. Investigator [801]. Investment [2409]. Invisible [3737, 3902, 4077, 2720]. Involved [255]. Involvement [1204]. IP [3535]. Iran [2212, 3428]. Iranian [1495]. Iraq [3035]. Ironies [3990, 2304]. Irony [1700]. Irrationality [1758]. Irresistible [2779]. Islam [2211, 3205, 2745, 1634, 3077, 1013, 2843, 1287, 2699, 2973, 1011]. Islamic [2213, 2212, 2744, 3995, 1287]. Islamists [2447]. Island [1113]. isma [2154]. Israel [1630]. Israeli [590]. Israelis [3218]. Issue [808, 133, 1019, 81, 789, 278, 236, 329, 317, 304, 291, 343, 392, 378, 460, 408, 446, 434, 509, 524, 493, 559, 586, 544, 575, 672, 613, 598, 633, 751, 714, 735, 831, 788, 807, 773, 840, 877, 867, 923, 893, 908, 941, 968, 978, 957, 989, 1036, 1056, 1004, 1124, 1093, 1107, 1074, 1229, 1244, 1213, 1282, 1316, 1294, 1266, 1332, 1354, 1450, 1416, 1395, 1435, 1489, 1467, 1478, 1500, 1544, 1531, 1515, 1558, 1585, 1601, 1613, 1572, 1642, 1671, 1627, 1658, 1721, 1736, 1687, 1707, 1776, 1752, 1764, 1789, 1802, 1853, 1822, 1836, 1867, 1887, 1902, 1913, 1946, 1957, 1926, 1935, 1973, 1990, 2028, 2002, 2091, 2065, 2047, 2111, 2126, 2160, 2141, 2178, 2196, 2208, 2222, 2250, 2279, 2236, 2265, 2292, 2337, 2307, 2321, 2386, 2374, 2351, 2400]. Issue [2417, 2455, 2442, 2430, 2528, 2513, 2498, 2482, 2551, 2596, 2569, 2581, 2628, 2643, 2656, 2615, 2705, 2723, 2677, 2694, 2753, 2740, 2770, 2800, 853, 1176, 364, 252, 264, 34, 116, 476, 878, 1432, 174, 525, 2078, 148, 69, 393, 752, 99, 249, 193, 715, 48, 736, 2048, 211]. Issues [677, 2674, 1516, 1579, 2839, 1878, 2815]. Italian [2831, 2821, 917, 2825, 2261, 1305, 2833, 2827]. Italiano [2829]. Italo [2720]. Italy [2815, 2819, 2592, 2826, 560, 2443, 2818, 2817, 2828, 2822, 2820]. Itself [995, 4078]. Ivanovitch [1076]. Ivory [844]. Izium [4198].

Jacobins [2652]. Jacques [1442, 1437].

Jainism [2845]. James

[4021, 1460, 747, 748, 4020]. Jane [442]. Japan [2141, 973, 2649, 731, 2648, 2146, 2280, 2153, 2150, 299, 2493, 2406, 2149, 2225, 2431, 1251, 621, 2282]. Japanese [587, 2148, 1289, 2151, 706, 2281, 1364, 1952]. Japanese-American [2281]. Jarrett [3929]. Jay [361, 3490]. Jay-Z [3490]. Jazz [3929, 3924, 3931, 3933, 3926, 3927]. Jean [1442, 1437]. Jean-Jacques [1442, 1437]. Jerome [1088]. Jesse [2779]. Jewish [608, 3292, 609]. Jews [2977, 1937]. Jihadi [3843]. Jihadism [3830]. Jim [3493]. Jobs [4153, 4154]. Johan [1078]. John [1497, 359, 129, 1931]. Johnny [3640]. Jordan [4184]. José [1083]. Joseph [356]. Journal [3311, 3330, 3367, 3387, 3400]. Journalism [3445, 3614, 3441, 3436, 3435, 3437, 3440]. journalists [3445, 3440]. Joy [2298]. Jr. [3485]. Judaism [3996, 2841]. Judge [3898]. Judges [3921, 3336, 3333, 3335]. Judgment [1541]. Judgments [1460, 3168]. Judicial [3342, 3334, 3337, 3341, 3336, 3888, 3345, 3605, 4051, 1403, 3339]. Judiciaries [3335]. Judiciary [3333]. Julie [1268]. June [2443, 470, 560, 4199]. Jungle [3297]. Junkie [3461]. Juror [3076]. Jury [3898]. Just [3781, 3817, 3797, 2214, 4076, 1407, 4008, 3779, 3793, 3799]. Justice [3902, 2733, 2939, 4097, 381, 3922, 3907, 3904, 3915, 3684, 4077, 2934, 3795, 4005, 3997, 3692, 3918, 4206, 3914, 3318, 3484, 4082, 4203, 4064, 3796, 4200, 4136, 3460]. Justifying [2174]. juvenile [3453].

Kafka [3236]. Kamasutra [3235, 2886]. Karl [3087, 1079, 2251, 1084]. Karneval [3397]. Keep [1674]. Keeping [3342, 1859]. Keeps [2955]. Keith [3929]. Keller [360]. Kenya [2731, 1365]. Kepler [4247]. Key [3815, 1116, 1012, 3422]. Keynes [1455]. Kill [1285]. Killer [3052]. Killing [3779]. Kin [1359, 3488]. Kind [3645]. King [3485, 3094]. Kingdom [2029, 1025]. Kings [735, 736]. Kinship [3963]. Klimt [1661]. Kneel [3146]. Knew [950]. Knight [3445]. Knot [1644]. Know [3007, 2246, 645, 3183, 3003, 2477, 3395, 4007, 3274, 2787, 3689]. Knowing [3856, 814, 2536, 1535, 2021, 4082, 2137]. Knowledge [1019, 56, 1030, 1539, 1967, 1343, 857, 2055, 1521, 2545, 2662, 1140, 4117, 3351, 1109, 2608, 2347]. Korea [2651]. Korean [3423, 4013]. Kuan [2199]. Kuan-hsi [2199]. Kyiv [4189].

Labor [3720, 1400, 3961, 1770, 3612]. Labor-Force [3720]. Laboratories [1117, 1032, 887]. Laboratory [369, 4094]. Labyrinth [3968]. Lake [4062]. Lamartine [1942]. Lamoille [3447]. Land [3740, 3694, 1586, 2022, 2030, 2518, 3931]. Länder [2589, 2588]. Lands [3312]. Landscape [2775, 162, 2118, 189, 2566, 1805, 389, 1614, 4118, 504, 1803]. Landscapes [4210]. Language [4095, 1038, 2488, 1046, 4201, 2487, 1040, 4205, 1594, 322, 1036, 3258, 3395, 1037, 1051, 1941, 2889, 1042, 1049, 4211, 1050, 1045, 3818, 4091, 964, 3870, 1039, 4093, 4206, 3628, 1047, 1044, 4200]. Languages [2745, 586, 833, 1052, 4204, 3870, 4203]. Lankan [3846]. Large [4095, 1618, 466, 239]. Largest [2764]. Larkin [3237]. Last [3841, 3487, 1318, 4224, 3786, 1541, 3145]. Late [325, 2761, 2448]. Lateef [3932]. Latin [1099, 2215, 1743, 3072, 699, 3622]. Latinization [1882]. Latino [3586]. Latinos [4038]. Lavinia [3556]. Law [1088, 2311, 3163, 3899, 1595, 2736, 4079, 4048, 3890, 4104, 3523, 1323, 4240, 71, 3913, 797, 4053, 2083, 4067, 231, 3035, 3670, 3891, 3344, 4078, 568, 3908, 2758, 3067, 3806, 2873, 4055, 3790, 3683, 2986]. Law-Stability [3670]. Lawlessness [4078]. Laws [3794, 3965, 3783, 2861]. Lawyer [4004]. Lawyers [3922, 3903]. Lead [3925]. Leader [125, 3777, 1644]. Leaders [3967]. Leadership [282, 933, 3776, 285, 280, 3772, 278, 735, 759, 3769, 743, 3773, 3774, 3768, 3971, 279, 3770, 286, 736, 3867, 4243, 3771, 738]. Leaderships [3775]. Leading [1809]. League [2572]. Learn [1177, 3814]. Learned [3234, 3423]. Learning [3951, 4120, 3943, 2686, 77, 3004, 764, 3953, 1046, 4086, 2715, 872, 509, 1957, 3003, 2462, 3948, 512, 1792, 2662, 3619, 1148, 2830, 3885, 2661, 1535, 2021, 1047, 3008, 2620, 3466]. Learnings [1614]. Lease [2531]. Leave [1724]. Leaves [3179]. Leaving [4006]. Lecture [2564]. Lectures [3662]. Led [4184]. Left [1511, 2826, 3374, 3060, 710, 2698]. Legacies [1132]. Legacy [2702, 3269, 2918, 37]. Legal [4229, 798, 3912, 1838, 2210, 1024, 3921, 3901, 3919, 396, 793, 4004, 3896, 4037, 3922, 3907, 3685, 934, 2173, 3681, 3920, 138, 3670, 2999, 3908, 2172, 3911, 3914, 195, 3917, 4036]. Legalism [2940]. Legalization [2268]. Legally [4241]. Legislating [2788]. Legislation [3590]. Legislative [4045, 3340, 3771]. Legislatures [3252]. Legitimacy [2311, 3320, 2450, 2699, 1108]. Legitimate [729]. Legitimize [3938]. Legitimizing [1200]. Lehrter [2358]. Leipzig [2367]. Leisure [149, 1080]. Leninist [2254]. Lens [2476, 3964, 3631]. Leo [1291, 3185]. Leonardo [2624]. Lesbianism [3237]. Leslie [358]. Lesotho [4065]. Less [4010, 1746, 3458]. Lesson [724, 3091, 3092]. Lessons [4220, 1726, 3470, 3423, 3607, 3373, 4156, 2976, 2461, 3886, 279, 3791, 2107, 3434, 3210]. Let [4014, 1490, 1549]. Lethe [2200]. Letter [1691, 207, 3017, 2877]. Letters [3290, 3151, 999, 2879]. Level [3582, 307, 4140]. Levelers [1209]. Lever [2507]. Leveraging [4182]. Leverkusen [2364]. Lewis [127]. Liberal [2682, 2691, 2686, 2684, 339, 1269, 2688, 461, 2978, 2677, 2679, 2687, 2685, 2689, 3954, 1164, 2681, 22, 1148, 85, 2680, 3354, 768]. Liberalism [3480, 1950, 2380, 3085, 3989]. Liberals [3196]. Liberation [2515, 2513]. Liberatory [4213]. Liberties [3260]. Liberty [3990, 1326]. Libertyville [1729, 1732, 1730, 1733]. Librarianship [2532, 2533]. Libraries [2547, 2530, 2537, 2539, 2534, 2540, 2546, 2541, 3750, 2544, 2542]. Library [2530, 2532, 2543, 3359, 2529, 2538]. Libya [3784]. Life [3135, 4123, 1115, 1285, 2649, 3285, 780, 1168, 2729, 3281, 1359, 1363, 1629, 443, 1283, 1873, 252, 3139, 503, 3282, 3286, 655, 1431, 2471, 2389, 3284, 2793, 2520, 3223, 3984, 1365, 3288]. Life-Extending [1431]. Life-Styles [2520]. Lifelong [2411]. Lifton [361]. Light [608, 2658, 1754]. Lights [4238]. Like [4101, 766, 4016, 2605]. Limitation [1849, 1850, 1848, 1234, 224]. Limitations [663, 184]. Limited [3849, 2466, 217, 3848]. Limiting [1420, 2018, 3118, 3792]. Limits [3137, 3955, 1536, 1532, 3083, 1416, 1531, 2938, 1432, 4170, 1258, 1540, 3714, 1533, 972, 1372, 1538, 1426, 1784]. Line [3627]. Lines [1545, 4103]. Linguistic [4210, 4201, 1050, 3023, 4202]. Linguistics [4213, 1053, 1043, 4208]. Links [3829, 1114]. Linus [55]. List [3575]. Listening [3637, 3927]. Literacies [2129, 1779]. Literacy [1712, 2130, 1716, 1697, 2132, 1707, 2126, 1608, 2127, 1717, 1709, 2133, 2128, 2134, 2471, 2137, 1708]. Literary [965, 859, 496, 856, 2557, 858, 374, 1781, 2646, 2558]. Literature [3753, 1098, 1461, 3314, 1699, 2036, 905, 3158, 3014, 3553, 3950, 347, 1321, 2015, 2958, 1675, 1298, 1101, 3356, 3159, 1271]. Litigation [3684, 3990]. Little [4241, 3791, 3689]. Live [3353, 1168, 1700, 3580, 2446, 3466]. Lived [3313, 1872]. Lives [2687, 1189, 1952]. Living [2845, 2663, 2065, 2077, 2196, 4068, 2842, 3850, 504]. Loathing [2618, 3119]. Local [3676, 899, 647, 1693, 900, 3451]. Localizing [4185, 4184]. Log [3094]. Logic [126, 2464, 1049, 994, 1987, 2993, 1638]. Logical [2727]. Logics [4208]. Lombardi [2870]. London [728]. Long [3580, 2539, 4156, 3644, 3659, 1872, 1379]. Long-Lived [1872]. Long-Range [1379]. Longer [3135]. Longevity [3718, 1915, 1871]. Longitudes [3]. Longley [3538]. Look [2734, 1268]. Looked [2157]. Looking [1134, 1467, 3071, 3377]. Looks [944, 2532, 43]. Losing [4028, 2807, 3509, 3486]. Loss [2432, 4077, 1604]. Lost [3644, 2390, 4008, 2858, 3266]. Lot [4241]. Louise [3658]. Love [358, 1780]. Lovejoy [1940]. Loving [3930]. Lower [3393, 552]. Lower-Class [552]. Lucky [3944, 2622]. Ludwigstrasse [2361]. Lure [3272]. Lustration [3384]. Luther [361, 3485].

Maastricht [2377]. Mach [732]. Machiavelli [1307]. Machine [2504, 494, 1909]. Machines [889, 1983, 3570, 4089, 1982]. Macro [3583, 3582]. Mad [3322]. Made [2736, 1594]. Madness [3062]. Magellan [2910]. Maghrib [1016, 1014]. Magnitude [811]. Mahler [1661]. Mainline [1637]. Mainstream [3623]. Maintaining [1209]. Mainzerstrasse [2355]. Maison [1268]. maisons [3656]. Major [2537]. Majority [3068]. Make [2047, 3713]. Makers [466, 1713]. Makes [4039, 2301]. Making [4153, 1674, 4103, 3314, 225, 3851, 1975, 877, 1347, 3860, 2679, 878, 1983, 1805, 159, 2343, 2545, 3783, 2763, 870, 1010, 3177, 105]. Male [2619]. Malthusian [2806]. Man [361, 1444, 40, 1086, 3393, 3325, 14, 264, 651, 1607, 21, 314, 3045, 29, 3276, 489]. Management [3424, 3677, 3420, 1402, 1517, 3426, 420, 2170]. Managerial [2424, 760]. Managers [2501]. Managing [2498, 4068]. Mandate [2312]. Manifestation [4237]. Manners [1396, 2017]. Mannheim [1084]. Manning [3516]. Many [3199, 2712, 1153, 1974, 2742]. Maoism [2491]. Map [2027, 3873, 2802]. Mapping [2950, 3787, 3829, 3702]. Marbles [2925]. March [1846]. Marginality [1438, 1483]. Margins [2205]. Marian [2355, 2358, 2361, 2364, 2367, 2369]. Marital [723]. Mark [2379, 1279]. Market [3561, 2663, 3961, 2232, 1481]. Marketing [3733, 2883, 3912]. Markets [3733, 4151, 4163, 3935, 3685, 3904, 4162, 3102, 4164, 3466]. Marriage [2791, 723]. Martin [3165, 3485]. Marvels [3573]. Marx [2864, 3077, 2251, 1271]. Marxism [1740]. Mary [1082]. Masculinities [3506]. Masculinity [2870]. Masks [3199, 2492, 3636]. Mass [183, 178, 862, 185, 184, 174, 108, 1209, 1121, 179, 176, 181, 841, 680, 1681, 3454, 3448, 3449, 3451, 3459]. Massachusetts [470]. Masses [1083, 3684]. Massively [2237]. Mastery [3103]. Matching [2470]. Material [2730, 570, 2404]. Materialization [2522]. Materializing [3760]. Materials [3097, 3696]. Maternal [2406]. Mates [3573]. Mathematical [1987, 2993]. Mathematics [2246, 134, 2689, 3095]. Matrix [4076]. Matter [10, 18, 27, 33, 46, 58, 67, 78, 47, 59, 68, 79, 97, 114, 131, 144, 98, 115, 132, 145, 171, 190, 208, 233, 172, 191, 209, 234, 250, 262, 276, 288, 251, 263, 277, 289, 302, 315, 327, 340, 303, 316, 328, 341, 362, 376, 390, 405, 363, 377, 391, 406, 432, 444, 458, 472, 433, 445, 459, 473, 491, 507, 522, 541, 492, 508, 523, 542, 557, 573, 584, 595, 558, 574, 585, 596, 611, 631, 688, 670, 612, 632, 689, 671, 712, 733, 749, 771]. Matter [713, 734, 750, 772, 694, 785, 805, 828, 837, 786, 806, 829, 838, 851, 865, 875, 890, 852, 866, 876, 891, 906, 921, 939, 954, 907, 922, 940, 955, 966, 976, 987, 1001, 967, 977, 988, 1002, 1017, 1034, 1054, 1071, 1018, 1035, 1055, 1072, 1091, 1105, 1122, 1173, 1092, 1106, 1123, 1174, 1211, 1227, 1242, 1263, 1212, 1228, 1243, 1264, 1280, 1292, 1314, 1329, 1281, 1293, 1315, 1330, 1352, 1367, 1380, 1392, 1353, 1368, 1381, 1393, 1414, 1433, 1448, 1464, 1415, 1434, 1449, 1465, 1476, 1487, 1498]. Matter [1512, 1477, 1488, 1499, 1513, 1529, 1542, 1556, 1569, 1530, 1543, 1557, 1570, 1583, 1599, 1611, 1624, 1584, 1600, 1612, 1625, 1640, 1656, 1669, 1684, 1641, 1657, 1670, 1685, 1705, 1719, 1734, 1749, 1706, 1720, 1735, 1750, 1762, 1774, 1787, 1799, 1763, 1775, 1788, 1800, 1820, 1834, 1851, 1864, 1821, 1835, 1852, 1865, 1885, 1900, 1911, 1923, 1886, 1901, 1912, 1924, 1933, 1944, 1955, 1970, 1934, 1945, 1956, 1971, 1988, 2000, 2025, 2044, 1989, 2001, 2026, 2045, 2063, 2075, 2089, 2108, 2064, 2076, 2090, 2109, 2124, 2139, 2158]. Matter [2175, 2125, 2140, 2159, 2176, 2194, 2206, 2220, 2233, 2195, 2207, 2221, 2234, 2248, 2263, 2277, 2289, 2249, 2264, 2278, 2290, 2305, 2319, 2335, 2348, 2306, 2320, 2336, 2349, 2372, 2384, 2398, 2414, 2373, 2385, 2399, 2415, 2428, 2440, 2453, 2479, 2429, 2441, 2454, 2480, 2496, 2511, 2526, 2548, 2497, 2512, 2527, 2549, 2567, 2579, 2594, 2612, 2568, 2580, 2595, 2613, 2626, 2640, 2654, 2668, 2627, 2641, 2655, 2669, 2692, 2703, 2721, 2737, 2675, 2693, 2704, 2722, 2738, 2676, 2751, 2768, 2784, 2798, 2752, 2769, 2785, 2799, 2813]. Matter [2823, 2834, 2851, 2814, 2824, 2835, 2852, 2871, 2893, 2916, 2932, 2872, 2894, 2917, 2933, 2948, 2965, 2987, 3001, 2949, 2966, 2988, 3002, 3018, 3033, 3048, 3063, 3019, 3034, 3049, 3064, 3080, 3098, 3115, 3132, 3081, 3099, 3116, 3133, 3153, 3171, 3191, 3207, 3154, 3172, 3192, 3208, 3227, 3245, 3262, 3279, 3228, 3246, 3263, 3280, 3293, 3309, 3328, 3347, 3294, 3310, 3329, 3348, 3364, 3385, 3398, 3417, 3365, 3386, 3399, 3418, 4096, 2857, 4241, 3759, 3198, 3343]. Matters [4215, 2968, 3264, 3924, 2800, 3074, 1026, 2796]. Mature [3136]. Maturity [1287]. Mausoleum [2123]. Maximizing [765]. May [11, 556]. Maynes [2060]. McCarthyism [2887]. Mcdermott [1888]. McNamara [3219]. Me [3322, 3476, 4014, 3462]. Meal [960]. Mean [2937, 4134, 3366]. Meandering [2712]. Meaning [488, 2196, 2205, 50, 2992, 3581, 486, 854]. Meaningful [3250, 2466]. Meaninglessness [486]. Meanings [2757, 2712, 100, 2726, 2333]. Means [1058, 2466, 4087]. Measure [3257, 1667, 3847]. Measurement [490, 3581]. Measures [223, 3124, 2522]. Measuring [3143]. Mechanical [2145, 3575]. Mechanisms [265, 2410, 1347, 447, 268, 3707]. Media [3851, 4121, 3370, 174, 2167, 1813, 4112, 108, 3585, 1121, 4133, 179, 1683, 3433, 3439, 4134]. Mediation [1277]. Medical [792, 1894, 791, 796, 1895, 804, 118, 1344, 1891, 2725, 1887, 1203, 2079, 1204, 4225, 1284, 2272, 2287, 1893, 1888, 2697, 4124]. Medicalization [1334]. Medicine [2883, 1877, 4130, 1895, 1897, 1891, 2731, 2492, 1896, 2868, 2726, 1406, 1336, 2016]. Medicus [397]. Medieval [1223, 325, 980, 2993, 1421]. Meditation [3477, 814]. Medium [2819]. Meeting [3643, 2450]. Meetings [8, 16, 26]. Meets [2779, 2146]. Meiji [2225]. Melancholy [498]. Melons [2148]. Melville [3276]. Members [11, 667]. Membership [3937, 259]. Memoir [2245]. Memorandum [635, 654, 1125]. Memorials [1961].

Memoriam [1270]. Memories [2492, 3481, 2503, 2706]. Memory [3761, 1277, 1459, 2283, 3714, 3692, 2200, 3711, 2494]. Men [2983, 1961, 3138, 2304, 3970, 6, 3506]. Menaces [3573]. Mental [2633, 4215, 1351, 3393, 4226, 4221, 4223, 4224, 4228, 4216, 4217]. Meritocracy [658]. Mesopotamian [1218, 1217]. Messengers [3658]. Messiah [3276]. Metaphor [3753]. Metaphors [959]. Metaphysics [3276]. Methods [3896, 3341, 897, 3871, 854]. Metropolis [236, 243, 762, 242, 765, 239]. Metropolitan [241, 245]. Mexican [257]. Mexican-Americans [257]. Mexico [4180, 4187]. Mezzogiorno [2821]. Mice [2881, 3138]. Michael [1076, 3538]. Michelle [3487]. Micro [3583]. Micro/Macro [3583]. Microbes [3573]. Micropolitics [2793]. Microprocessor [2503]. Mid [1636, 2870]. Mid-Century [1636, 2870]. Middle [1078, 3844, 1607, 2259, 4002, 3189, 1562, 981, 2574, 3167]. Middlemarch [3655]. Might [1177]. Migración [4187]. Migrants [2805]. Migration [4079, 4035, 1483, 105]. Mike [3029]. Miles [2870]. Milestones [4044]. Militarism [1559]. Military [1236, 3517, 1235, 3523, 3515, 3607, 1560, 401, 2942, 2150, 1258, 915, 4152, 4140, 1524, 1831, 1825, 3521]. Milky [3]. Mill [3026, 747]. Millennials [3505]. Millennium [1213, 1222, 2646, 3361]. Miller [360, 1829]. Mind [1088, 1296, 780, 151, 1975, 3176, 3183, 3179, 323, 3178, 2996, 3585, 3140, 3181, 3042, 1862]. Minding [3175]. Minds [3006]. Minerals [2805]. Ming [971, 2650]. Mingled [1667]. Mini [3820]. Mini-Publics [3820]. Minimum [3412]. Mining [2805]. Minnesota [2770, 2775, 2773, 2772, 2778, 2777, 2771, 2779, 2774, 2780, 2776, 2781, 2783]. Minorities [2573, 2435, 2797, 2572, 1598, 1555]. Minority [4041, 1588, 259, 2133, 1875, 680, 4208, 3503]. Minority-Group [259]. Minority-White [4041]. Minus [360]. Mir [207]. Miracle [2575]. Mirror [2541, 1954, 4244, 37]. Misapprehension [994]. Miscegenation [3929]. Misconceptions [2619]. Misgivings [1431]. Missile [1850, 1848, 1826, 1839]. Missiles [1849]. Missing [4216]. Mission [2885, 2342]. Missionary [3038]. Missions [3084]. Mistaken [2331]. Mistakes [3667, 3619]. Misunderstandings [2383]. Mitigate [3724]. Mitigating [3952]. Mitteleuropa [2112, 2121]. Mixed [4041, 4090, 3522]. Mixing [2777]. Mobility [4180, 3755, 4036]. Mobilization [1479, 3810, 417]. Mobilizing [4141]. Model [1892, 1128, 1996, 795, 4202, 4208]. Modeling [1975]. Models [4095, 639, 4135, 2531, 3435]. Moderation [2422]. Modern [1506, 1088, 889, 126, 3321, 2315, 1905, 1359, 1224, 1861, 3130, 608, 1967, 877, 1515, 325, 1950, 1517, 34, 878, 2616, 2055, 707, 981, 3372, 980, 2224, 2970, 480, 2104, 2837, 1766, 2490, 720, 3834, 1938, 486, 1421]. Modern-Secular [608]. moderne [3653]. Modernism [4168, 4170]. Modernities [2644, 2741, 2643, 2740, 2744, 2747, 2645]. Modernity [2743, 859, 2745, 2252, 1446, 2309, 2954, 2749, 2746, 2750, 972, 2332, 2647, 2576, 2742]. Modernization [1628, 2023, 578, 3630, 2760]. Modernizing [2664, 646]. Modes [449]. Modest [1797]. Modification [2636]. Modify [652]. Mohammed [1077]. Molecular [885]. Molecules [2633]. Moment [2539]. Momentum [554]. Monarchies [578]. Monetary [1412, 3470]. Money [2881, 1114, 2151, 3276]. Monism [2978]. Monitoring [2406]. Monopolies [3879]. Monotheism [1220]. Monster [2240, 1109].

Montagne [1275]. Montaigne [1493].

Mood [4222, 3581]. Moral

[3396, 4103, 2939, 4168, 1504, 2731, 3998, 4165, 1610, 2728, 4156, 4166, 3215, 2485, 1337, 4164, 1638, 3788]. Moralistic [2779]. Morality [3780, 1554]. Morally [1879, 4167, 4243]. Morass [1423]. Morphology [4188]. Mortality [729, 727]. Moscow [2186]. Most [1738, 3689]. Mostly [1539]. Motherhood [2406]. Mothers [1952]. Motivation [2132]. Motivator [4005]. Motor [2637]. Motorizing [2154]. Mountain [3555]. Mouth [1537]. Movement [554, 3669, 863, 707, 706]. Movements [3712, 2293, 2747]. Mover [1725]. Movies [1857]. Moving [1853, 2253, 4032, 3984, 1858]. Moynihan [3480]. Mr. [3804]. Much [2725, 1718, 3559, 2403, 2832, 1746, 1985, 2518]. Muddling [2761]. Muhammad [1302]. Multi [4090, 2363]. Multi-Agent [4090]. Multi-German [2363]. Multicultural [2840]. Multidimensional [3934]. Multigenerational [1882]. Multilateral [3966, 3420, 3973]. Multilingual [1486]. Multiple [2741, 2740, 3775, 2747]. Multipolar [3974]. Multiracialism [3069, 3072]. Mumford [127]. Munich [2361]. Muove [452]. Murder [3322]. Muses [2706]. Museum [2711, 3865, 3641, 2707, 2709, 2718, 2708, 2717, 181, 2720, 2714]. Museums [2711, 2713, 2719, 2715, 2705, 2712, 2716, 2706, 4118, 2710, 3327]. Music [813, 3634, 3640, 3641, 3638, 2880, 3930, 3927]. Musical [3553, 3554, 3641, 3292, 3644, 1102]. Musings [1783]. Muslim [2211, 2788, 2976]. Muslims [2977]. Mutilation [2796]. Mutual [1484, 4155, 257]. Mwasi [1646]. My [3482, 3060, 2154, 888]. My-Car-isma [2154]. Myopia [1171]. Myrdal [2427, 2424, 2419]. Mysteries [1783]. Mysterious [2989]. Mystery [998, 2039]. Mysticism [2963, 653]. Myth [963, 109, 3505, 104, 274, 957, 2283, 100, 108, 99, 1537, 110, 301, 1759, 1672, 3072]. Mythmaking [103, 107, 99]. Mythology [101, 1766, 3727]. Mythos [3545]. Myths [1869, 103, 1919].

N [577]. Nadia [2789]. Naked [3076]. Names [645]. Narayan [2101]. Narrative [3546, 2318, 3542, 3550, 3548, 3549, 1665, 1677, 3543]. Narratives [3544, 4204]. Nation [2037, 1008, 2592, 2432, 572, 2571, 583, 2326, 1790, 1507, 2328, 3481, 2761, 2776, 1846, 3224, 1766, 2434, 2324, 2357, 2325]. Nation-Building [2761]. Nation-State [2432, 572, 2571, 583, 1507, 2761, 3224, 2434]. Nation-States [2328]. National [2179, 1397, 1511, 2117, 1579, 2435, 2831, 3566, 2543, 1025, 92, 2053, 2231, 3340, 1846, 53, 48, 186, 2273, 4102, 2378, 1683, 2323, 2333, 2223]. Nationalism [1253, 2587, 2331, 3808, 4029, 2332, 2333, 2324, 2256, 582]. Nationalisms [2322]. Nationality [2757]. Nationalizing [2435]. Nations [410, 3835, 414, 2939, 2321, 1761, 2951, 2572, 2803, 2776, 1309, 335, 3854]. Native [3376, 1589, 4221, 1462, 1931, 3863, 3862]. Natives [2791]. Natural [206, 1052, 961, 3391, 1976]. Nature [663, 3740, 3054, 40, 1086, 1621, 2947, 2859, 489, 72, 737, 3325, 1711, 65, 4150, 3057, 1510, 3296, 45, 330, 2524, 1039, 3299, 3050, 3058, 1111, 3389, 3302, 3056, 3300, 2841, 2518, 2774]. Navajo [1897]. Nazi [2492]. Necessity [3783, 110, 1426]. Need [639, 665, 648, 2477, 1204, 2464, 287, 3908, 1903]. Needs [1502, 1350, 4176, 2059, 2223, 3881]. Negotiated [3796]. Negotiating [1430]. Negotiation [2802]. Negro [556, 524, 544, 3480, 3488, 548, 924, 2427, 527, 536, 528, 537, 550, 931, 529, 927, 769, 926, 526, 535, 552, 553, 530, 928, 533, 930]. Negroes [3488, 935]. Neighborhood [4074]. Neither [2268, 3656]. Nenabozho [3857]. Neo [1422]. Neo-Romantic [1422].

Neoliberal [4161]. Neptune [963]. Nervous [3706]. Net [3534]. Netherlands [730]. Nets [1978, 3668]. Network [2199]. Networked [4133]. Networks [3746]. Neues [2589, 2588]. Neural [1978]. Neurologic [3576]. Neurology [1045]. Neurons [3007]. Neuroscience [3706, 3891, 1979]. Neurosciences [1027]. Never [950]. Newbury [13]. News [2589, 1226, 3433, 2588, 1677, 3443, 3446, 3432, 3434, 3438, 3444, 3442, 3440]. Newspapers [353]. Newton [745]. Next [3798, 2694, 3804, 641, 662, 2525]. NGO [4185]. Nice [3287]. Niebuhr [1086]. Niels [57]. Nietzsche [457, 3062, 1271]. Nigeria [3884]. Night [1539]. Nightingale [2145]. Nine [2387]. Nineteen [2387, 1230]. Nineteen-Seventies [1230]. Nineteenth [725, 3546, 974]. Nineteenth-Century [974]. Nineties [2184]. Niño [3031]. Nixon [990, 991]. Nkrumah [739]. No [1700, 1056, 1067, 1069, 3644, 2200]. No-Growth [1056]. Nomad [1931]. Non [3821, 1179, 2210, 1224, 2202, 2993, 4094, 3836, 898, 2107, 4181, 2203]. Non-Chinese [2202, 2203]. Non-Crisis [1179]. Non-Human [4094]. Non-Legal [2210]. Non-Lessons [2107]. Non-Modern [1224]. Non-State [4181]. Non-Western [2993, 3836, 898]. Nonconstitutional [1402]. None [2742]. Nonfictional [3546]. Nonimmigrant [2797]. Nonpoets [3152]. Nonpoint [3729]. Nonprofit [3613]. Nonproliferation [3408]. Nonverbal [307]. nor [2268]. Norden [1753, 1759]. Nordic [1752, 1764, 1758, 1765, 1771, 1773, 1768]. Norm [3966]. Norm-Setting [3966]. Normal [2280]. Normative [665, 3391]. Norms [3965]. North [4088, 2213, 1767, 3423, 2493, 697, 2038]. Northeast [1563]. Northern [623, 547, 1222]. Nose [3110]. Nostalgia [498]. not-so-efficient [3466]. Note [133, 1225, 345, 192, 236, 1203, 695, 691, 200, 1941, 210, 1518, 1528, 1858]. Notebook [3536]. Notebooks’ [2908]. Notes [1511, 439, 183, 488, 3193, 1579, 1621, 430, 443, 551, 879, 565, 958, 3122, 728, 2902, 312, 246, 816, 1510, 4083, 3157, 861, 823, 1503, 91, 658, 186, 1384, 1197]. Nothing [443, 1718, 1856, 3302]. Notice [3839]. Notion [4080]. Nouvelle [1268]. Novel [590, 358, 369, 4025, 589, 364, 4027, 587, 365, 592, 2006, 591, 4016, 588]. Novelist [107, 371]. Novelistic [4023]. Novelists [366]. Novels [4018, 593]. Novyi [207]. Noxious [4162]. NPT [3424, 3425]. Nuclear [3975, 3410, 2192, 1231, 3985, 2439, 73, 3429, 3405, 3974, 3406, 3420, 3404, 3782, 217, 3408, 3972, 3402, 3783, 1567, 3413, 3428, 3422, 3409, 3401, 3430, 3973, 3982, 3414, 3425, 1582, 3976, 3407, 3415, 3211, 3411, 3427, 117, 2193, 3403, 3426, 3984, 1825, 3423]. Nuclear-Weapon [3982]. Nuclear-Weapons [3413]. Numbers [134, 3095]. Numeracy [2137, 2136]. Numerical [2867]. Nurture [3050].

Obama [3509, 3482, 3477, 3483, 3587, 3487, 3501, 3500, 3485, 3486, 3506]. Object [2779, 2712, 854]. Objectives [997, 3859]. Objectivity [824]. Objects [3358, 2712]. Obligation [2213]. Obligations [4138, 3488]. Obscures [3491]. Observations [199, 998, 1910, 1960, 294, 124, 2370, 1011]. observatories [3442]. Observed [1711, 1818]. Obsolete [583]. Obstacles [2466, 743]. Obstinate [583]. Obvious [1731]. Occupation [2147]. Occupational [3951]. October [636, 3011]. Odd [3703]. OECD [1257]. off [2874]. Offer [4066]. Office [1806]. Officers [11]. Oil [1261, 1244, 1653, 1248, 2231, 1255, 1254]. Oil-Producing [1248]. Ojibwe [2776]. Okayama [731]. Old [410, 3137, 1674, 2777, 3142, 2215, 821, 2695, 1478, 1687, 2581, 3265, 1318, 2492, 3547, 2574, 1580, 2582, 1133, 1117, 755, 822]. Old-Fashioned [822]. Older [3718]. Oligarchs [3809]. Once [3748, 916, 1364]. One [3290, 1165, 3322, 2982, 1539, 3542, 1716, 2478, 3066, 1206, 1153, 1974, 2698, 4191, 4081, 2742, 814]. Ongoing [2714]. Online [3955, 3528, 3527]. Ontologies [4173]. OPEC [1256]. Open [843, 3817, 3749]. Opening [4116]. Openness [566, 216]. Opera [1914, 1918, 1913, 1915, 1919]. Ophelia [3383]. Opinion [3583, 3579, 3582, 3584, 227, 3586, 3788]. Opinions [2451]. Opportunities [4223, 3729, 350, 3724]. Opportunity [2601, 1515, 1552, 3935, 3408]. Opposed [569]. Opposition [1547]. Optimal [2294]. Optimism [3911]. Options [3307, 1744]. Opulence [2402]. Oral [3553, 916]. Orchid [3540]. Ordeal [545, 526]. Order [561, 2444, 938, 580, 3833, 564, 559, 2442, 572, 3502, 2443, 661, 567, 3828, 1006, 569, 581, 571, 195, 2434, 3213, 2651, 1410, 1592, 560]. Ordering [570]. Ordinary [4227]. Organ [2730]. Organic [2043, 4026, 163]. Organizado [4187]. Organization [554, 3992, 226, 1033, 63, 1346, 463]. Organizational [2609, 1237, 3377]. Organizations [1400, 2661, 4184]. Organize [1023]. Organized [4180, 3833]. Origin [3009, 1372]. Origins [3226, 3053, 4030, 4223, 3607, 1042, 4029, 1451, 3683, 2742, 2494, 4247]. Oriya [2790]. Ortega [1083]. Oscar [357]. Osten [2589, 2588]. Oswald [1075]. Othello [1491]. Other [2118, 2826, 4023, 1299, 2380, 2328, 3006, 1528, 2304, 1032, 3552]. Otolith [3557]. Otto [2622]. Ottoman [3088]. Ought [3651]. Our [1456, 2241, 3137, 2733, 3634, 3747, 3200, 1435, 3005, 3559, 3136, 3269, 2484, 2971, 3870, 266, 2460, 4089, 3687, 3683, 3689, 37, 3509, 3464, 3486]. Ours [2031]. Outcome [3942]. Outcome-Policy [3942]. Outcomes [3772, 4176]. Outdoor [683, 684]. Outer [435]. Outline [619]. Outlook [924, 3143, 675]. Outreach [3345]. Outside-in [1048]. Overstory [4026]. Oversupply [1161]. Overt [1939]. Overview [2827].

P. [126]. Packages [3336]. Pain [2488, 2487]. Pains [3663]. Painter [818, 25]. Painting [817, 2099, 319]. Palestinians [3218]. Pan [3501]. Pan-Racial [3501]. Pandemic [1943, 2072, 4046, 4156, 2086]. Pandemics [3831]. Pane [2490]. Panel [2052]. Panels [1829]. Paper [39]. Papers [6]. Parades [3369]. Paradigm [2578, 2119, 2035, 1664]. Paradigms [2392]. Paradise [102, 2991, 961, 3323]. Paradox [1536, 2261, 1269, 1904, 2035, 3670, 3837, 336, 3452]. Paradoxes [1441, 3801, 3252, 1130]. Parallel [2239, 2244, 2240, 2237, 180]. Parallelism [2241, 2242]. Paralysis [1183]. Parental [2619]. Parenting [1355]. Parents [1455]. Paris [3648]. Park [1617]. Parliamentary [416, 2827]. Parliaments [415]. Parochial [258]. Part [3792, 2077]. Partial [1272]. Participant [912]. Participation [3720, 850, 2828]. Participatory [3915, 3820, 3439]. Particle [1371]. Particular [898]. Parties [668, 1770]. Partisan [3333, 3771]. Partisanship [3609]. Partnership [2534]. Parts [2546, 3575]. Party [3390, 1510, 3259, 3671, 3509, 3486]. Party-State [3671]. Passion [3103]. Passive [2809]. Passport [3648]. Past [3277, 3147, 1400, 629, 3620, 1935, 2371, 2953, 3560, 3667, 2283, 4236, 2093, 2762, 3210, 774, 3494, 3619, 3209, 3757, 3497, 1157, 3084, 810, 2190, 3000, 1409, 2753, 471, 2812]. Pasternak [207]. Pastoral [141]. Pasts [3860, 4042]. Patent [2881]. Patents [2883]. Pater [1272]. Paternelle [1268]. Path [2033, 856, 3786, 4212]. Pathological [2166].

Pathology [1340]. Paths [2644, 3999, 3315, 180]. Patient [2068, 1938]. Patients [4124]. Patriarchal [3968]. Patriotism [3998, 3319]. Patriots [2652]. Patron [3641, 4008]. Patronage [1932, 1926]. Patronal [3803]. Pattern [2749, 1771, 242, 63]. Patterns [1485, 293, 2516, 1633, 3718, 534, 106, 1647]. Paul [1050, 1619]. Pauling [55]. pay [3468]. Paying [3559, 3930]. Peace [566, 3785, 2356, 3189, 3837, 3085, 3167, 1846]. Peaceable [2029]. Peaceful [3852]. Peak [1891]. Peanuts [1013]. Peasant [2315]. Pedagogies [4209, 3106]. Pedagogy [1966, 3126]. Peintre [3653]. penal [3454]. Peninsula [4013]. People [2775, 360, 1700, 2145, 1675, 3633, 2518, 2766, 2311, 1742, 3211, 2059]. Peoples [3869]. Peopling [2771]. Perceived [1404]. Perceiving [3708, 451]. Perception [166, 167, 743, 3220, 621, 1288, 2163]. Perceptions [3587, 1473, 4112]. Perfect [2854]. Perfectible [1443]. Performance [2659, 1649, 1471, 1187]. Performances [959]. Performing [3361]. Peril [4101]. Perils [3991, 3221]. Period [2156]. Peripheralization [2201]. Periphery [2571, 2205, 1012, 2586, 2197, 3127]. Permanent [2545]. Perpetual [2504, 3085]. Persian [1561]. Persistence [1008, 1793, 3934, 1007]. Person [3774, 769, 370]. Personal [183, 2782, 2715, 3769, 2205, 2245, 881, 2490, 1938]. Personality [512, 652]. Perspective [1100, 1654, 2611, 1420, 2018, 3118, 2555, 3975, 2804, 3953, 2859, 2401, 1244, 2047, 166, 2275, 2411, 609, 850, 2184, 4075, 696, 1230, 2301, 2133, 4206, 2066, 1355, 3499, 4216, 3427, 3624, 4181, 3056, 3775, 1708, 3666, 3023, 1254, 930, 1830, 3457]. Perspectives [3723, 3358, 1798, 4103, 2239, 384, 773, 1213, 1316, 1613, 2731, 364, 253, 2571, 34, 1284, 2272, 2042, 2838, 657, 2844, 1241, 2128, 2218, 1360, 1872, 2308, 1843]. Persuasion [2983].

Persuasions [959]. Pessimism [1495]. Peter [3202]. Petit [1634]. Petrified [2911]. Ph.D. [3016]. Pharaohs [2211]. Pharmacology [2265]. Phenomena [619]. Phenomenology [422]. Philanthropy [1932, 1926, 3613, 1327, 1930, 2781]. Philip [3237]. Philipp [326]. Philistines [200]. Philosophe [1312]. Philosopher [948]. Philosophers [1986, 735, 35, 736]. Philosophical [608, 790, 313, 1307, 1089, 4094, 429, 2397]. Philosophy [206, 2984, 4123, 1993, 455, 3164, 4147, 3111, 3308, 326, 2524, 3757, 2560, 2559, 3051, 1378]. Philosophy-Envy [3051]. Phoenix [64]. Phosphorus [3694]. Photograph [2355, 2361, 2364, 2367, 2369, 3029, 2358]. Photographing [2997, 2996]. Photography [158]. Photosynthesis [5, 4]. Physical [4, 618, 2288]. Physician [1898, 801, 1890, 1342]. Physician-Basic [1890]. Physician-Investigator [801]. Physics [3572, 126, 1371, 2990, 2953, 65, 2621, 3120, 135]. Physiocentric [2524]. Physiognomic [167]. Piano [3649]. Picasso [169]. Pickwick [964]. Picture [2630, 3042]. Pictures [166, 1679]. Piece [4216]. Pieces [3094]. Pied [1361]. Pioneers [3643]. Piper [1361]. Pirate [2880]. Pirates [3749]. Pirenne [1077]. Place [2686, 1158, 2036, 437, 814, 1482, 769, 85, 2608, 2998, 2796, 4043, 3451]. Places [2846]. Plague [726, 2979]. Planetary [3568, 4126, 3221, 3220, 2842]. Planets [3046]. Planners [403]. Planning [645, 565, 682, 3136, 1327, 248, 758, 3876, 563]. Plant [3575]. Plate [1376]. Play [3851, 958, 3122, 1549]. Playbook [3875]. Playground [2774]. Pleasant [1407]. Pleasure [3108, 3645, 1692]. Plebiscites [3837]. Plenary [2062]. Plenty [2589, 2588, 1130]. Plight [1104]. Plumed [54]. Pluralism [1992, 3073, 2977, 3187, 3996, 607, 4149].

Pluribus [1632, 3569]. Plus [3750]. Pluto [3220]. Pneuma [3492]. poem [3447, 3461, 4190, 3476, 3357, 3754]. Poems [3556, 2961, 3538, 2104, 3462]. Poet [189, 2566, 1942]. Poetry [3061, 577, 3152, 2947, 2889, 3489, 1053, 860, 64, 2846]. Poets [4025, 1089]. Poincaré [2954]. Point [1096, 3889, 1305, 685]. Poison [75]. Poisoned [1855]. Poland [592, 2006]. Polanyi [1079]. Polarization [4129, 4045, 4031, 3259]. Polarized [3605]. Polaroid [2502]. Police [2811]. Policies [125, 3720, 1362, 1363, 3566, 2071, 4076, 2451, 2726, 2822, 3565]. Policing [4239, 4238, 4135]. Policy [1654, 2819, 990, 1238, 2821, 3268, 335, 2778, 334, 1030, 3173, 2736, 3597, 339, 1116, 2295, 225, 3851, 2274, 1232, 1504, 329, 1229, 1558, 1572, 1948, 3560, 1517, 2275, 2931, 331, 1738, 1560, 3667, 1324, 2276, 1428, 51, 1327, 562, 2268, 1580, 2762, 1549, 2272, 3783, 2658, 3378, 3619, 330, 770, 3962, 1554, 870, 2346, 1884, 1114, 3942, 2267, 245, 466, 186, 332, 1340, 2080, 2452, 3472, 3470]. Policy-Making [3851, 870]. Polish [1937]. Politic [3196, 600]. Political [947, 992, 2334, 260, 1603, 2117, 241, 2831, 798, 3807, 4159, 2315, 1807, 911, 4143, 2419, 228, 648, 2265, 3187, 3532, 759, 3587, 4147, 2779, 3114, 3369, 1205, 1014, 3940, 741, 1560, 1960, 4031, 578, 661, 618, 2164, 2394, 3787, 3820, 138, 1006, 2561, 2257, 4156, 3816, 3344, 3469, 4212, 2352, 1010, 2092, 2489, 3533, 2114, 3442, 898, 699, 3819, 498, 2562, 1237, 3805, 2370, 4164, 4172, 429, 2397, 680, 1340, 4136]. Politician [2041]. Politicians [3821]. Politicized [3991]. Politics [703, 708, 2213, 3194, 2217, 3271, 2552, 1469, 2825, 2754, 2778, 2212, 3505, 404, 1966, 3126, 704, 140, 2360, 2695, 698, 1546, 322, 2757, 1789, 1957, 1926, 2208, 3265, 3640, 700, 3100, 3803, 701, 3108, 1633, 3220, 2665, 3631, 746, 2853, 2746, 1458, 1635, 1942, 3595, 2327, 894, 3501, 122, 419, 696, 1012, 758, 697, 89, 1190, 2149, 2708, 2187, 3652, 2084, 2068, 705, 466, 1481, 2218, 2165, 123, 1677, 709, 2310, 2339, 3500, 1644, 2333, 971, 3673, 533, 1522, 1361, 2209]. Polity [2031, 2646]. Poll [3580]. Polls [3824]. Pollution [3325, 4006, 3729, 678]. Pomp [1932]. Poor [1592, 1066, 1616, 1067, 309, 1746, 2863]. Pop [2880]. Popes [1633]. Popular [178, 485, 3837, 2701]. Population [1068, 718, 121, 3720, 3723, 1058, 446, 714, 1652, 731, 2411, 119, 447, 1059, 2517, 716, 725, 448, 1884, 715, 4216, 120, 1882]. Populations [2413]. Populism [4052]. Porgy [4022]. Portrait [1272]. Portraits [1613, 1615, 1617, 1619]. Posed [2790]. Position [2554, 1217, 3038, 120]. Positive [1508, 800, 4125, 2806]. Possibilities [3986, 4079]. Possibility [334]. Possible [2131, 139, 2718]. Post [3094, 1008, 3493, 2791, 3505, 1094, 1861, 1005, 1004, 3483, 1013, 920, 1948, 3496, 3491, 3773, 2450, 3828, 4067, 1006, 844, 3481, 2810, 2802, 753, 1015, 3501]. Post-Apartheid [2791, 4067, 2810, 2802]. Post-City [753]. Post-Cold [3828]. Post-Colonial [3094, 1006]. Post-Communist [3773]. Post-Independence [2450]. Post-Modern [1861]. Post-Racial [3505, 3483, 3496, 3491, 3481, 3501]. Post-Racialism [3493]. Post-Reagan [1948]. Post-Soviet [2450]. Post-Traditional [1008, 1005, 1004, 1013, 1015]. Post-World [1094]. Postcolonial [4126]. Postcommunism [2393]. Postcommunist [2256]. Postconflict [3849, 3847]. Postconventional [948]. Postgraduate [2603]. Postlude [3578]. Postmodern [3130, 3160, 2304, 3327]. Postmortem [2395]. Posture [1237]. Postwar [2283, 299, 2227, 702, 2593, 2431]. Potentials [4115]. Poverty [2056, 2295, 3942]. Power [2819, 561, 2444, 3909, 1892, 3745, 2432, 1857, 4045, 3974, 2071, 3352, 4177, 1957, 863, 1347, 2505, 1560, 401, 3888, 3404, 746, 4146, 3971, 4067, 1258, 624, 3402, 3818, 2848, 3401, 3838, 3340, 3645, 1453, 3407, 3211, 4102, 3427, 783, 3655, 336, 4064, 1259, 680]. Powers [3393, 218, 3770, 4026]. Practical [4103, 3603, 502]. Practicalities [3850]. Practically [4241]. Practice [3161, 1791, 3913, 1648, 795, 2461]. Practices [2711, 3720, 2790]. Practitioner [2555]. Pragmatic [965]. Pragmatism [3373]. Prague [3221]. Praise [1266]. Praxis [3870]. Pre [195, 727]. Pre-Industrial [727]. Pre-Revolutionary [195]. Precarious [1115]. Precautions [3300]. Precious [3481]. Precision [3514]. Precision-Strike [3514]. Predicament [1478, 1059]. Predicting [645]. Prediction [449, 446, 1904]. Predictions [642]. Predominantly [933]. Preface [1019, 3117, 4229, 278, 329, 317, 304, 291, 343, 392, 378, 460, 408, 446, 434, 475, 509, 524, 493, 559, 586, 544, 575, 672, 613, 598, 633, 751, 714, 735, 831, 788, 807, 773, 840, 877, 867, 923, 893, 908, 941, 968, 978, 957, 989, 1036, 1056, 1004, 1124, 1093, 1107, 1074, 1229, 1244, 1213, 1282, 1316, 1294, 1266, 1332, 1354, 1450, 1416, 1395, 1435, 1489, 1467, 1478, 1500, 1544, 1531, 1515, 1558, 1585, 1601, 1613, 1572, 1642, 1671, 1627, 1658, 1721, 1736, 1687, 1707, 1776, 1752, 1764, 1789, 1802, 1853, 1822, 1836, 1867, 1887, 1902, 1913, 1946, 1957, 1926, 1935, 1973, 1990, 2028, 2002, 2091, 2065, 2047, 2077, 2111, 2126, 2160, 2141, 2178, 2196, 2208, 2222, 2250, 2279, 2236, 2265, 2292, 2337, 2307, 2321, 2386, 2374, 2351, 2400, 2417, 2455]. Preface [2442, 2430, 2528, 2513, 2498, 2482, 2551, 2596, 2569, 2581, 2628, 2643, 2656, 2615, 2705, 2723, 2677, 2694, 2753, 2740, 2770, 2815, 2800, 853, 1176, 173, 252, 4214, 69, 48, 4174]. Prefatory [133, 192, 236, 210]. Preferences [3940]. Prejudice [3225]. Preliminary [635, 449, 1621, 970]. Prelude [3567].

Prepared [3055]. Prerequisite [92]. Prerequisites [228]. Prescriptions [1965]. Presence [1098]. Present [3277, 1099, 3147, 1144, 3620, 1935, 3104, 3752, 4236, 2093, 2762, 3109, 3497, 1157, 3084, 95, 3362, 3000]. Presentation [93]. Presenting [3896]. Presents [3860]. President [285, 1829, 1846]. Presidential [744, 3369, 3096, 3434, 3770]. Presidentialism [3772]. Press [2878, 1678]. Presumptions [1419]. Presuppositions [1531, 1748]. Pretrial [4080]. Prevailing [2425]. Preventing [3429, 3882]. Prevention [3845, 1662, 3847, 2084, 2088]. Preventive [139]. Price [1653, 2153, 971]. Prices [3253]. Primacy [1509]. Primary [1339]. Prime [1725]. Primitive [102]. Principal [1725, 1607, 2301]. Principle [3783, 3816, 216]. Principles [1504, 76, 4067, 2846]. Print [1671, 1679, 1673, 1677]. Priorities [1429, 3420]. Priority [2462]. Prison [3461, 2908]. prisoners [3458]. Prisons [2423]. Privacy [2082, 656, 3748, 3527]. Private [3535, 1198, 2096, 1601, 1192, 1189]. Privilege [2879]. Privileged [1620]. Privileges [1129]. Prize [21]. Proactive [4202]. Probable [2131]. Problem [548, 761, 295, 3835, 791, 1429, 4072, 3902, 800, 1967, 1036, 1515, 1402, 1050, 137, 2317, 448, 535, 15, 2270, 1937, 3879, 2323]. Problematic [3682]. Problems [170, 1029, 395, 1351, 1894, 416, 424, 1046, 1040, 4, 1037, 425, 119, 656, 49, 513, 1195, 1389, 338, 518, 2393, 579, 2690, 1237, 2544]. Procedures [802]. Process [757, 4080, 2664, 512, 107, 1256, 3167, 3942, 1785, 1374]. Process-Policy [3942]. Processes [3938, 3934, 3837, 2760]. Processing [889]. Producing [1248]. Production [371, 117]. Productive [684]. Productivity [3721, 2507, 308, 1708]. Profession [404, 396, 2596, 3922, 401, 2606, 2597, 2600, 1180, 3107, 2599, 2608, 3765, 2533].

Professional [439, 2727, 1479, 1148, 794, 1965, 1406, 4003, 1405, 402]. Professionalism [2071, 3108]. Professionalization [400]. Professionals [2734, 4000, 4003, 3169]. Professions [395, 2598, 3103, 392, 3101, 394, 309, 393, 1528, 3106, 3102]. Professoriate [2611, 2603, 2609]. Professors [349]. Profile [528]. Profiling [4210]. Profits [2883]. Progeny [1359]. Prognostication [2732]. Program [219, 1270, 1722, 1847]. Programs [1831]. Progress [2775, 669, 1351, 1605, 1377, 633, 2694, 264, 4224, 1590, 2930, 382, 860, 3571, 885, 338, 3729, 570, 1409, 3042]. Progressive [3959]. Prohibition [2268, 2269, 2271]. Prohibitionism [2274]. Project’ [2913, 1918, 3480]. Projections [539]. Prolegomenon [1517]. Proliferation [3401, 3377]. Prologue [3560]. Promiscuity [3230]. Promise [3955, 4101, 2753, 2995, 1289, 3796]. Promises [1430, 3682]. Promote [3566, 3496]. Proofing [3777]. Proper [86, 4163, 312]. Property [2879, 2875, 2529, 2873, 2794]. Proportion [165]. Proposal [483, 2286, 436, 2004]. Proposals [214, 1797, 3336, 1240]. Propositions [3871]. Proprietary [2882]. Prose [4025]. Prosocial [3534]. Prosopography [896]. Prospect [2521, 1379]. Prospects [3802, 2977, 778, 2079, 1195, 42, 338, 3806, 1982, 1471]. Prosperity [2153]. Protect [3784, 3535]. Protecting [1426]. Protection [3785, 4183, 2573, 3832]. Protest [2422, 3672, 4219, 1222]. Protestant [121, 602, 3547, 2368]. Protests [576]. Protocol [1850]. Provider [1345]. Province [731]. Psychiatry [4218, 139, 4219, 402]. Psycho [737]. Psycho-Historical [737]. Psychoanalysis [361, 3662, 1984]. Psychodynamic [282]. Psychological [3004, 374, 383, 495, 657, 534]. Psychologists [273]. Psychology [361, 3396, 439, 3021, 1553, 1780, 3769, 904, 2011, 2915]. Psychosis [4219]. Public [3562, 1603, 3535, 4115, 1922, 2649, 3268, 4130, 2874, 1461, 4113, 2601, 2096, 1427, 2295, 2729, 2246, 4063, 3264, 2537, 3583, 2882, 3579, 3582, 1120, 1200, 1504, 3988, 343, 1107, 1601, 3265, 2811, 3267, 3584, 2602, 1517, 3269, 2173, 3607, 3940, 1324, 1428, 3996, 3333, 51, 3594, 2539, 1604, 3590, 2472, 4112, 3993, 2529, 2538, 1195, 3344, 4117, 2084, 2653, 1191, 1884, 227, 2341, 1114, 3351, 3836, 4118, 2490, 3586, 3690, 4119, 3273, 3879, 1747, 1189, 1602, 3788, 2650, 4245, 4073, 4052, 3270, 3266, 3362, 3442, 3908]. Public-Private [3535]. Public-Service [1200]. Publics [3820]. Publishers [2534, 349]. Publishing [905, 348, 347, 6, 350]. Puerto [1585, 1592, 1313, 1594]. Puff [3537]. Puff-Ball [3537]. Pump [306]. Punishment [4080, 4078, 1404, 3453, 3456, 3451]. Punitive [4078, 2267, 4070]. Purity [3878]. Purpose [1519, 682, 1791, 1560]. Purposes [1147, 1125, 94, 502]. Pursue [4146]. Pursuit [1388, 3492, 2471, 3440]. Push [3668]. Pushkin [3238]. Putin [3803, 3804, 3806]. Putin-Style [3806]. Putinism [3801]. Putting [1660]. Puzzling [3584, 3335]. Pyramid [1868]. Pyrrhic [3459].

Qing [2650]. Qualitative [1235, 138]. Quality [133, 1208, 3945, 719, 675, 137, 1209, 3739, 1728, 1809, 135, 3436]. Quandary [2728]. Quantification [136]. Quantitative [902, 915, 138]. Quantity [133, 675, 137, 135]. Quantization [2622]. Quantum [1384, 135]. Quasars [3703]. Quebec [2039, 2038]. Quentin [3640]. Queries [551]. Quest [1597, 3712, 2635, 606, 3162, 2442, 2585, 4004, 3932, 3085, 2562, 2584, 1938]. Question [1158, 3804, 1463, 3159]. Questions [798, 3987, 3491, 1690, 1197]. qui [3656]. Quicker [4039]. Quo [41].

R. [2101]. r2p [3785]. Race

[3509, 2984, 1096, 622, 1235, 1595, 3482, 3477, 4096, 551, 4160, 3479, 546, 531, 3125, 3478, 2757, 613, 615, 3487, 3074, 3070, 4031, 3489, 215, 462, 3068, 630, 2860, 4233, 1596, 628, 3067, 3486, 2891, 4036, 2425, 3495, 4136, 3455, 3962, 2803]. Racial [4209, 3493, 4080, 3505, 2424, 3073, 3483, 3480, 3587, 3496, 4238, 3071, 3502, 3491, 4074, 4236, 3501, 3481, 2426, 4235, 4078, 534, 3065, 3646, 3484, 3500, 1681]. Racialism [3493]. Racialization [4034, 4208]. Raciolinguistic [4207, 4206]. Raciolinguistics [4209]. Racism [616, 4038, 1093, 4237, 4230, 4005, 4161, 3947, 930]. Radical [3959, 4077]. Radicalization [2314]. Radio [3645]. Rainbow [3501]. Rajput [1290]. Ralph [3375, 3485]. Random [179, 1998]. Range [1379]. Rapid [3669]. Rapid-Growth [3669]. Rates [3138]. Rational [1905, 1756]. Rationale [781]. Rationality [1758, 567, 1582, 2340, 644, 678]. Ray [1385]. Raymond [1276]. Re [1548, 4022]. Re-Establishing [1548]. Reaching [1738, 1896]. Reactions [3811, 3698]. Reactivation [1015]. Reader [1688, 352, 1698]. Readers [1694]. Reading [2135, 344, 1491, 1539, 905, 3235, 1700, 343, 1687, 1689, 3108, 1696, 3662, 1779]. Ready [2379, 4143]. Reagan [1948]. Real [2773, 2439, 3337, 3940, 1980, 1404]. Real-State [2439]. Realigning [3586]. Realism [2340, 319]. Realists [368]. Realities [545, 2767, 1658, 2731, 1580]. Reality [2318, 488, 3271, 2502, 161, 2990, 256, 732, 1537, 301, 1759, 678, 156]. Realizing [3267]. Really [3548]. Realm [1921, 825, 1494, 1747]. Reaping [3625].

Reappraisal [217]. Reason [1700, 3988, 3996, 3177, 3970]. Reasoning [4091]. Reasons [3993, 2855]. Reassembling [3747]. Reassessing [2691, 3751]. Rebalancing [3378]. Rebel [984]. Rebellion [1300, 3794]. Rebels [3842, 3843]. Rebuilding [2357]. Reception [3755, 1271]. recession [3454, 3474]. Recipe [2867]. Reciprocal [223, 3124]. Reclaiming [3862, 3903]. Reclamation [3870]. Recognition [3858, 2793, 4081, 284]. Recognizing [4229]. Recommendations [3428]. Reconceptualizing [3531]. Reconcile [4007, 2728]. Reconciliation [546]. Reconciling [3376, 2129]. Reconfigurations [2555]. Reconsideration [1593]. Reconsidered [1566]. Reconstructing [3700, 2321]. Reconstruction [359, 2318, 1005, 2843, 3844, 2493, 540, 4081]. Records [8, 16, 26]. Recreation [683, 684]. Recruitment [91]. Recurrent [103]. red [3454, 3791]. Redefining [2538]. Redefinition [487]. Redemption [61]. Rediscovery [2113]. Redistribution [3937]. Redress [2856]. Redressing [3686]. Reduce [3590, 3976]. Reducing [3410, 3593, 4160, 4217, 4073]. Reemergence [3187]. Reestablishing [3617]. Reference [1027, 984, 898, 1007]. References [2761]. Referendum [3814]. Reflecting [3350]. Reflection [2782, 2728, 1938]. Reflections [847, 1150, 1456, 333, 2508, 2041, 3493, 796, 2831, 2438, 129, 1144, 435, 849, 1283, 608, 1932, 1178, 1128, 2523, 863, 1126, 1780, 3267, 2604, 2492, 707, 790, 1193, 2014, 1458, 742, 949, 1422, 929, 2147, 1290, 1457, 2332, 3519, 111, 2431, 913, 841, 318, 1154, 1358]. Reflectons [1110]. Reform [361, 1232, 2466, 1194, 2312, 3186, 2476, 2464, 1714, 3336, 3875, 2120, 2465, 1936, 2472, 3670, 2658, 2287, 770, 3664, 2661, 2822, 2459, 1237, 3884, 3454]. Reformation [2957]. Reformations [4026]. Reformers [501]. Reforming [2905, 2666, 3875, 770, 1064]. Refracted [1291]. Refrain [3290]. Reframing [4115]. Refugee [3832]. Refuse [4066]. Refusing [4204]. Regarding [3734, 2891, 1849]. Reggio [2830]. Regime [3811, 3853, 3514, 3409, 2653, 3984]. Regimes [2252]. Region [1097, 4189]. Regional [1565, 572, 1562, 2809, 2153, 3832, 1573]. Regionalization [3427]. Regions [3]. Regulate [1418, 3685]. Regulation [4106, 804, 3961, 3473, 3469]. Regulator [1524]. Reimagining [4143]. Reinforcing [1209]. Reinhold [1086]. Reinventing [3679, 2225]. Reinvention [3732]. Rejoinder [3821]. Relation [515, 1050, 1527, 29]. Relations [1096, 2284, 3334, 551, 3719, 615, 620, 1373, 1786, 909, 1012, 1472, 3503, 1481, 2281]. Relationship [3163, 86, 1833, 2788, 1107, 1376, 647, 4148]. Relationships [998, 1116, 4182, 1051, 1979]. Relevance [811, 1484, 1581, 606, 1196, 2716, 2790, 826]. Relevant [2059]. Religio [2971]. Religio-Secular [2971]. Religion [1991, 1639, 3986, 2217, 599, 2007, 3121, 1993, 3991, 2968, 2212, 605, 537, 3988, 598, 1627, 1990, 2208, 3987, 2571, 594, 1995, 2097, 1636, 3997, 314, 1637, 2837, 2218, 385, 653, 1222, 2979, 1747, 2209, 4137]. Religionist [3842]. Religions [2210, 1999, 1638, 2836]. Religious [1397, 1141, 2777, 1992, 3992, 2210, 1628, 2023, 2788, 2972, 607, 3990, 1635, 1287, 3993, 4122, 604, 63, 3079, 2218, 60, 3673]. Reluctant [3614]. Remain [462]. Remains [4023]. Remaking [755, 4172]. Remapping [2391]. Remarks [57, 564, 487, 179]. Rembrandt [15]. Remembering [3710]. Renaissance [3429, 4064, 2957, 3292, 3408, 2973].

Renewable [3566, 3592]. Renewed [3265, 610]. Reorganization [686]. Repackaging [2812]. Repair [3734, 4081]. Repertoire [1040]. Repetition [1861, 3130]. Replacing [4047]. Reply [1703, 1050]. Report [1087, 560, 2443, 1618, 1620, 2167, 3480, 2457, 93, 1847, 768]. Reporting [3895, 3439]. Representation [65, 3681]. Representations [3862, 1658]. Representative [1760, 3817]. Representing [3771]. Reprocessing [3429]. Reproduction [1652]. Republic [2327, 2311, 3155, 2212, 2879, 1742, 918]. Republics [1849, 1850, 1848, 3073]. Republished [2674]. Reputations [261, 142]. Request [1847]. Requiem [2251]. Requirements [213]. Requisites [2313]. Rereading [1969, 3070, 3662]. Rescue [2885]. Research [1506, 1099, 1023, 2691, 1031, 1429, 4171, 2338, 3815, 804, 73, 1024, 1201, 2337, 1020, 3359, 2067, 312, 1025, 2343, 2342, 521, 74, 1026, 795, 2868, 1430, 794, 1893, 2341, 871, 1028, 466, 2347, 2339, 1202, 2344, 1032, 888, 2667]. Researchers [466]. Researches [21]. Resetting [3722]. Reshape [2241]. Residential [2677, 2683, 2687]. Resilience [4058, 3868]. Resiliency [4077]. Resilient [3720, 2815]. Resistance [3960, 1591]. Resisting [3532]. Resituation [314]. Resolution [4183]. Resolving [3583]. Resort [3786]. Resource [677, 2161, 675, 1875]. Resources [792, 1789, 1650, 686, 2846, 2223, 679]. Respect [1319]. Respecting [4217]. Respond [3938]. Responding [2072]. Response [2623, 1131, 49, 4046, 1884, 871, 1309, 1154]. Responses [1654, 4178, 4104, 3073, 2811, 2173, 1896, 4176, 4179, 4184, 4009]. Responsibilities [867, 4138, 3415]. Responsibility [2082, 2939, 1546, 932, 1338, 569, 4001, 3984, 826, 4172, 3784]. Restive [1427]. Restoring [3338, 4217, 4137]. Restraint [3792, 3425, 1494, 3819, 3984].

Restraints [329, 331, 332]. Restructuring [2474]. Resurgence [2274]. Rethinking [3598, 4218, 3769, 4108, 3942, 4206, 2270]. Retirement [1880]. Retooling [4242]. Retrieval [2843]. Retrospect [2521]. Retrospective [2179, 1136, 2193]. Return [2216, 2913]. Returns [3179]. Revanche [3811]. revanchist [3455]. Revelation [1213]. Reverie [1440]. Review [2334, 1709, 4051]. Reviewing [355, 354]. Revised [3495]. Revising [1509, 3875]. Revisionist [3797]. Revisionists [1509]. Revisited [2427, 3477, 1074, 2417, 2375, 1790, 1940, 964, 3495, 1443]. Revisiting [3644]. Revival [3975]. Revived [1631]. Reviving [4217]. Revolt [339, 1083, 1634, 617]. Revolution [847, 346, 2909, 3864, 3802, 430, 437, 196, 3077, 2312, 4223, 2749, 1941, 3950, 2380, 1328, 3215, 3156, 540, 1011]. Revolutionary [2185, 742, 195]. Revolutions [1139, 1386]. Revolves [1922]. Rhetoric [3272]. Rhetorical [744]. Rhythms [3711]. Rican [1594]. Ricans [1585, 1592]. Rice [2148]. Rich [1066, 2408, 2863]. Richard [990, 4026]. Rico [1313]. Riddle [577]. Riegl [1278]. Right [3906, 4157, 3959, 1328, 4083, 4032]. Righteous [2312]. Rights [2937, 3320, 4159, 554, 1737, 1745, 1743, 532, 867, 1736, 1742, 3995, 1748, 1738, 1740, 2162, 3494, 1739, 3633, 1741, 4217, 1747, 1744, 3069, 1409]. Rigorism [2563]. Rigorously [179]. Riot [768]. Rise [1889, 3929, 536, 3935, 2875, 3374, 4040, 3973, 3216, 1481, 2434, 2748]. Rising [3665, 3901, 3938, 3663]. Risk [1425, 2169, 2161, 2160, 2173, 1790, 2164, 2167, 2534, 2172, 3976, 2165, 3403, 4217, 2174, 2170, 2163]. Risking [3375]. Risks [3410, 1832, 2171, 2162, 1063, 3724]. Rites [1659, 1459]. Ritual [3369]. River [1165, 3447]. Rivers [1589, 4068]. Road [2255, 2884, 3873]. Robbing [3610]. Robert [361, 3219, 1275, 3645]. Robotic [3780].

Robust [3426]. Role [1238, 439, 1892, 335, 1508, 3531, 2684, 3721, 775, 2734, 2410, 218, 3851, 546, 2173, 618, 2280, 93, 1053, 3786, 375, 3903, 1768, 844, 3422, 1412, 568, 975, 480, 937, 1533, 619, 3058, 3914, 3790, 2368, 3970, 928, 779]. Roles [933, 872, 4240, 1221, 982, 655]. Roman [1076, 1294, 1632, 3753, 1300, 3766, 3762, 1295, 3316]. Romance [1895, 1779, 1276]. Romania [2585]. Romantic [1668, 1422, 3371]. Romanticism [2925]. Romantics [2652]. Rome [3256, 1664, 3751]. Romes [105]. Room [1586, 2022, 3818]. Roosevelt [443, 1678]. Roots [150, 2127, 886, 2010, 2616, 4078, 2204, 3089]. Rope [1816]. Rostovtzeff [1076]. Rotten [2583]. Rough [2983]. Rousseau [1435, 1438, 1441, 1446, 1445, 3304, 1442, 3317, 1447, 1437, 1268, 1443]. Rove [3087]. Row [2708]. Rubens [963]. Ruins [2253]. Rule [3066, 3035, 3344, 2758, 3806]. Rules [2032, 1808, 2842, 3789, 3791]. Ruling [1927]. Rumford [5, 31]. Ruminations [1730]. Run [1618]. Running [2959]. Russia [2255, 3802, 3807, 974, 3974, 3810, 3808, 1291, 3809, 2390, 2389]. Russian [193, 3811, 3429, 196, 3803, 2587, 197, 2382, 3805, 2544]. Russians [360]. Ruth [3639].

Sacred [2096, 1856]. Sad [182, 3274]. Safety [3529, 3668, 2811, 3409]. Said [2459]. Saint [4008]. Salaam [4058]. Salome [3094]. SALT [1234]. Same [538]. San [3640]. Sangha [1008]. Santayana [1089]. Sappho [3754]. Sargent [1931]. Satchmo [3647]. Satellite [1832]. Save [3736]. Says [4112]. Scale [3598, 490, 2345, 466, 4087, 4145, 3734]. Scale-Community [4145]. Scales [312]. Scandal [3030]. Scandinavia [1772]. Scandinavian [703]. Scarce [792]. Scarcity [4060, 4057, 2404]. Scenarios [2804]. Scene [87, 88, 755]. Scenes [1805]. Schemata [683]. Schemes [3668]. Scholar [1768, 930]. Scholarly [2531]. Scholars [516, 466]. Scholarship [917, 863, 312, 2396, 918, 314]. School [1606, 1722, 3863, 424, 547, 546, 2131, 2476, 1618, 1620, 1610, 1619, 949, 1607, 4189, 2134, 258, 1997, 95, 1615, 3239]. Schoolhouse [2901]. Schooling [3863, 1731, 1616, 1620, 2128]. Schools [3629, 2659, 1605, 2300, 2474, 1601, 1613, 1721, 1789, 1203, 1608, 1791, 2458, 1724, 1604, 1615, 1617, 1619, 3993, 1730, 2301, 2461, 1614, 3908, 1893, 2460]. Schwab [1276]. Science [515, 1506, 2624, 1029, 1715, 889, 1520, 1420, 2018, 3118, 3899, 1115, 400, 2106, 81, 1030, 36, 2702, 2736, 1910, 2688, 2845, 1895, 1905, 1116, 2618, 3119, 2946, 140, 3890, 489, 4104, 1027, 1501, 2882, 1120, 1200, 3097, 38, 304, 877, 1107, 1887, 1902, 2615, 1418, 325, 3114, 3889, 4240, 2994, 485, 3395, 14, 520, 4131, 511, 4238, 3013, 1373, 2889, 34, 476, 878, 1906, 2616, 3950, 2166, 1428, 1714, 3731, 3888, 76, 2995, 43, 2809, 1033, 1963, 93, 2914, 2035, 3303, 426, 2716, 909, 326, 1682, 2606, 2689, 92, 2561, 2999, 997, 745, 1422, 3894, 4233, 42, 2629]. Science [85, 2346, 3571, 39, 4132, 1121, 2992, 3576, 3031, 2993, 1114, 685, 1111, 1109, 1713, 1108, 1419, 2562, 164, 1903, 1379, 1336, 2016, 2697, 1378, 95, 94, 44, 486, 468, 1784, 203, 1421, 2452, 3446, 3438, 1427]. Sciences [206, 3600, 3869, 449, 1023, 2691, 2735, 483, 2288, 1376, 480, 1533, 2345, 2563, 3868, 510, 1390, 3311, 3330, 3367, 3387, 3400]. Scientia [478]. Scientific [1712, 4229, 3712, 481, 1716, 1416, 1707, 1907, 3896, 35, 3898, 312, 1432, 2510, 3950, 2930, 308, 513, 309, 1424, 1709, 650, 1710, 1028, 269, 2562, 519, 1708, 887, 163, 3763, 4247]. Scientist [3005, 1890, 2607, 1110]. Scientists [3869, 3890, 202, 1963, 569, 1118, 90, 4001, 3897, 3169]. Scope [3598, 4163, 3588, 3340]. Scott [3067]. Screamers [3635]. Screen [297]. Scriptum [920]. Scrutiny [2341]. SDI [1844, 1843].

Sea [2284, 3694, 2285]. Search [1019, 3568, 3046, 3807, 1349, 737, 75, 550, 1343, 1021, 2013, 2028, 2811, 2119, 1459, 732, 2280, 2394, 2716, 3972, 2316, 1994, 997, 603, 2699, 4173, 352, 1384]. Searching [4088, 2222, 1317, 2535]. Seas [4004]. Second [2146, 1679, 2330, 1766, 1047, 540]. Secondary [1608, 1616, 1620, 1604, 1615, 1617, 1619]. Secret [3241]. Secretary [1847]. Sector [3613, 3731, 1192, 2828]. Sectors [3949]. Secular [1141, 2727, 1629, 608, 1323, 3995, 609, 3993, 603, 2971, 1577, 2837]. Secularism [2969, 1011]. Securitization [3467]. Security [2179, 4183, 3835, 1844, 3728, 3170, 3531, 1561, 2438, 4063, 1565, 2222, 1771, 4056, 1562, 3739, 4065, 1573, 3722, 2231, 1846, 2445, 1563, 2225, 4102, 1683, 3043, 3143]. See [989, 4074, 1983, 1754]. Seeger [3639]. Seeing [2036, 4231, 3490, 4074, 4009]. Seek [45]. Seeking [1745]. Segregation [549]. Selected [232]. Selection [3341, 3336, 3231, 517]. Selective [1261]. Self [1674, 1319, 608, 2585, 3395, 3607, 1773, 948, 1159, 2154, 141, 824, 1202, 2323, 106, 1421]. Self-Analysis [1159]. Self-Definition [2585]. Self-Determination [1202, 2323]. Self-Exposure [608]. Self-Interest [3607]. Self-Interpretation [106]. Self-Making [1674]. Self-Respect [1319]. Self-Understanding [3395]. Selfish [2855]. Selves [3199, 3747, 3200, 3136]. Semantics [477]. Semiconductors [2230]. Sen [2061]. Sensation [136]. Sense [965, 2036, 3888, 2989, 3177, 64, 1169, 1421]. Sensibility [3929, 168]. Sensitive [3319, 680]. Sensory [2636]. Sentences [2423, 1048]. Separate [2455]. Separation [2210, 3770]. Sequence [942]. Sequences [2257]. Serbelloni [560, 2443]. Serbs [2583]. Seriously [2269]. Serpent [54]. Servant [2303, 1768]. Service [4182, 1200, 844, 763].

Services [1347, 2780, 3908, 3914, 3917]. Serving [3346, 3897, 3464]. Session [2062]. Sessions [3639]. Set [2179]. Setting [212, 3966]. Settings [4178]. Seven [3873]. Seventeenth [728, 982, 1494, 319, 37]. Seventeenth-Century [982, 319, 37]. Seventies [1481, 1469, 1230]. Several [586]. Sex [2984, 3234, 3232, 3236, 2861, 1874]. Sexes [436, 2004, 1759]. Sexual [3967, 372, 3231, 3379]. Sexuality [4207]. Sexually [2170]. Shadow [811, 1062, 373, 3647, 3467]. Shakespeare [1666, 64, 1492]. Shale [3563]. Shaped [2805]. Shapes [3144]. Shaping [2775, 1341]. Shared [484, 3415]. Shares [3397]. Shift [2821, 1892]. Shifting [2420, 1633]. Shir [357]. Shir-Cliff [357]. Shoes [4181]. Short [3094, 3004, 2506, 1700]. Shortages [3866]. Should [2473, 2246, 3919, 1067, 3196, 1149, 266, 1746, 3576, 4164]. Show [2810]. Showa [2141, 2142, 2150, 2143, 2152, 2148, 2154, 2156, 2155]. Shrinking [3068, 824]. Shrugged [3809]. Si [452]. Sick [748]. Sides [4225]. Siècle [2603]. Sigfried [1279]. Sigmund [1085]. Signature [3106]. Significance [4041, 620, 4241, 4033, 2425, 3495]. Signifying [3930]. Signor [1662]. Signs [1388, 4096, 4177]. Silence [1791]. Silences [2041, 1767, 2463]. Silent [3256]. Silicon [2700]. Siloviki [3805]. Simple [4084]. Simplified [3916]. since [2421, 917, 1363, 1636, 1649, 2225]. Sine [20]. Singer [1931]. Singh [1290]. Single [682]. Single-Purpose [682]. Sings [1181]. Situation [260, 2419, 624, 205, 918]. Sixteen [941, 2104, 942]. Sixteenth [982]. Sixteenth- [982]. Sixth [1301]. Skaftafell [3060]. Skeptic [3405]. Sketch [3041, 961, 3042]. Skill [3139]. Skill-Intensive [3139]. Skilling [1146]. Skills [3951, 3496]. Skin [620, 4033, 621, 3636]. Sky [3702]. Slavery [1094, 1093]. Sleep [2635, 3711]. Slow [4013]. Slowly [610]. Small [2684, 2598, 1761, 1724, 3570]. Smaller [218]. Smell [2583]. Smile [29]. Smith [347]. Snapshots [2744]. Soak [2863]. Soaring [3666]. Social [1506, 3600, 1438, 2356, 945, 4144, 449, 488, 1023, 622, 3937, 2691, 1076, 3170, 1595, 3321, 606, 2735, 2736, 1325, 798, 1269, 2734, 2422, 4207, 813, 950, 605, 2733, 243, 1951, 527, 938, 3355, 2402, 1166, 528, 1277, 856, 2482, 759, 602, 3143, 943, 1731, 3810, 35, 238, 2436, 1517, 3769, 1633, 895, 3013, 1373, 4225, 185, 642, 298, 2486, 74, 479, 3672, 2413, 2352, 2293, 2780, 1770, 3722, 4001, 480, 881, 534, 2828, 1453, 1871, 1111, 4206, 269, 1661, 2755, 2405, 1222, 4203, 4069, 621, 4036, 2491, 2282, 61, 2404, 2425, 1390, 4200, 255, 1647, 2544, 2080]. Social [3724, 2542, 730, 3460, 3450]. Social-Democratic [1770]. Socialism [499]. Socialist [1849, 1850, 1848]. Socialization [1778]. Socializing [4107]. Societies [1008, 836, 3849, 418, 1005, 3719, 1004, 3136, 3852, 3716, 2256]. Society [1869, 175, 305, 1008, 3147, 973, 3144, 86, 1223, 1301, 1335, 1879, 3505, 1807, 1962, 1334, 843, 2365, 304, 1056, 1867, 1883, 3143, 549, 1873, 2571, 1459, 609, 895, 1813, 1510, 4083, 1164, 1876, 1880, 294, 1870, 1217, 1881, 296, 3918, 3351, 684, 1875, 1874, 660, 3614, 176, 644, 1872, 709, 4082, 3612, 181, 1882, 763, 1405, 1015, 177]. Sociobiology [1442]. Sociocultural [2406, 268]. Socioeconomic [567]. Sociogenic [4218]. Sociological [1386, 2199, 535, 2888]. Sociology [718, 395, 3147, 905, 2419, 1275, 864, 880, 706]. Softening [1407]. Solace [2200]. Soldier [372]. Solidarities [4150]. Solidarity [3195]. Solipsistic [3571]. Solitary [200, 3891]. Soluble [1206]. Solutions [3723, 4218, 3738, 1604]. Solve [761, 3916]. Somatic [3284]. Some [1029, 1150, 4019, 183, 395, 488, 1516, 1579, 1335, 4103, 798, 1621, 2215, 1359, 324, 580, 1144, 1797, 3987, 1748, 355, 3738, 1193, 2014, 215, 1604, 676, 2205, 100, 626, 4109, 657, 21, 870, 294, 2147, 4161, 269, 258, 179, 2795, 2657, 3519, 970, 268, 4054, 106, 1364, 1011, 3506, 1197]. Somebody [2714]. Someone [1168]. Something [1814, 2714]. Sometime [4022]. Somewhere [3493, 3501]. Song [3290, 3238, 3644]. Songs [1981]. sont [3656]. Sophiengemeinde [2369]. Sorcerers [4047]. Sorrow [2490]. Sorting [449]. Soul [3306, 2316, 2116, 3926, 748, 3930]. Sound [3638, 1679, 3646]. Source [1532, 2491]. Sources [605, 4104, 602, 4140, 3425, 2841]. South [1565, 2800, 1783, 2310, 2808, 2804, 1655, 2791, 2811, 2575, 2809, 4067, 624, 2541, 3933, 4065, 2803, 3692, 2489, 2801, 2810, 2806, 2647, 2802, 2805, 2812]. Southeast [1007]. Southern [3553, 3928, 1475, 935]. Sovereign [2586]. Sovereignty [384, 3842, 3974, 3847, 3857, 1406]. Soviet [1849, 1850, 1848, 199, 2228, 1252, 1560, 1841, 2450, 202, 198, 1936, 2257, 984, 1559, 301, 1831, 336, 3377, 203]. Space [1845, 1838, 435, 2622, 1822, 1836, 2376, 1372, 1831, 1825]. Spain [982, 2216, 2257, 725, 205, 2653]. Spanish [1496]. Spans [3135]. Spare [2518]. Speak [2095]. Speakers [1042]. Speaking [3380, 2919, 3999]. Speaks [3533]. Special [1027, 799, 984, 1007]. Species [3139, 4084, 3052]. Specific [4140]. Speculation [641]. Speech [1042]. Spend [1407]. Spengler [1075]. Spent [3420]. Sphere [2653, 2650]. Spheres [3836]. Spirit [3618, 603, 3284]. Spiritual [1639, 1629, 603, 4155]. Spiritualism [1289]. Spirituality [2777, 2849]. Spirochetes [3062]. Spooked [3551]. Spot [3596]. Spread [3429, 3514]. Spring [860, 3274]. Square [4081]. Squaring [1868]. Sri [3846]. St [1619]. Stability [703, 3985, 3981, 1566, 2188, 3972, 3670, 1162, 2191, 450]. Stabilization [223, 3124]. Stabilizing [3835, 1412]. Stable [3583]. Stacks [3359]. Stage [1318]. Stake [2340]. Staking [2465]. Stalin [200, 2123]. Stalled [4223]. Stamp [1145]. Stand [3869]. Standardization [4201]. Standards [281]. Star [2622]. Stars [3694, 4088]. Stasi [2371]. State [1506, 1758, 1511, 3244, 1008, 1062, 3883, 4072, 2432, 4048, 2215, 2438, 2210, 2439, 4044, 1501, 3677, 2437, 1144, 874, 2788, 3668, 1500, 2430, 3343, 572, 3686, 1509, 1950, 2571, 464, 3844, 583, 1507, 1186, 2447, 1510, 3672, 4046, 2216, 3680, 1505, 3344, 2761, 3345, 3372, 2092, 4049, 2653, 1559, 1503, 3224, 740, 2434, 4181, 2433, 4043, 2586, 2431, 4111, 4042, 4052, 3671, 1647, 1508, 935]. State-Court [3343]. Statehood [3849, 3848, 2584]. Statelessness [2448]. Statement [345]. Statements [665, 1849, 156]. States [2633, 2435, 989, 1332, 2321, 2436, 2450, 2328, 3850, 3982, 3836, 2448, 1817, 992, 1849, 1850, 1848, 1947, 991, 335, 775, 404, 225, 1565, 1363, 527, 1344, 1144, 1591, 337, 1594, 2071, 2226, 3922, 398, 2171, 3071, 3718, 2146, 1025, 996, 627, 2577, 984, 2658, 2227, 2426, 1249, 3962, 4033, 1472, 2861, 63, 2068, 3628, 1104, 900, 3622, 1032, 2204, 1681, 4200, 300]. Station [844]. Stationary [1062]. Statistical [1262, 413]. Statistics [2412]. Status [1099, 4207, 527, 3974, 4037, 934, 122, 2133, 4032, 123, 530, 4036]. Stay [3795]. Steady [1186]. Stealing [2810]. STEM [3945]. Step [2038]. Steps [3873, 3414]. Stereotypes [2215]. Stern [2622]. Stew [3204]. Steward [3605]. Stewards [2601, 3614]. Still [2968, 1589, 1568, 3924, 3183, 2455, 3074, 3547, 2029, 894, 2562]. Stimulating [3564]. stimulus [3471]. Stop [3930]. Stories [1226, 3113, 2846]. Story [2547, 4071, 2881, 2546, 2541, 2962, 4014, 3857, 3866]. Storywork [3870]. Straddling [4126]. Strain [1447]. Strains [703]. Strange [3235, 3312]. Strangers [3312]. Strasse [2367, 2364]. Strategic [1842, 3985, 3981, 1234, 3512, 1566, 1841, 4066, 3972, 1840, 1237, 1828, 1830]. Strategies [2462, 1682, 3847, 2057, 2273, 3884, 3876, 3724]. Strategy [1701, 3792, 3496, 3928, 1567, 1649, 3874, 3790]. Stratification [4159, 4033, 2760]. Strauss [3185]. Street [2564, 13]. Strengthen [266]. Strengthening [3943]. Strengths [768]. Strike [3514]. Strong [2436, 3948]. Structural [1878, 2495, 4230, 885, 3441]. Structure [488, 418, 3598, 562, 198, 419, 3963, 964, 39, 3838, 2172, 1871]. Structured [1145]. Structures [1157]. Structuring [4169]. Struggle [345, 2807]. Struggles [4057]. Struggling [4007]. Struth [2355, 2358, 2361, 2364, 2367, 2369]. Student [703, 708, 3953, 2474, 700, 701, 707, 576, 697, 710, 699, 706, 709, 517, 1154]. Students [2682, 199, 704, 698, 1161, 696, 702, 705, 1104, 1796]. Studies [2556, 3508, 1158, 3766, 3583, 714, 735, 877, 893, 833, 853, 3755, 862, 878, 4122, 3157, 823, 715, 736, 854, 3868, 1271]. Study [361, 718, 1096, 1495, 1038, 2432, 905, 3706, 2896, 374, 3006, 706, 3763]. Style [744, 603, 1540, 2833, 3806, 53, 48]. Styles [827, 842, 3641, 1025, 1026, 313, 2520, 1365]. Subject [2673, 3158]. Subjective [3021]. Subjects [804, 789, 788, 790, 799, 794, 3695]. Submission [3209]. Subordination [4201]. Substance [2470, 2273]. Suburban [1618]. Subversive [962]. Subverting [2436]. Succeed [4085]. Success [2508, 2474, 1505, 2346, 3885, 1769]. Successful [3716]. Succession [3804, 1644]. Suffer [2696]. Sufferers [2697]. Suffering [2073, 2495, 2482, 2492, 2484, 2486, 2493, 2485, 2491, 2494]. suggest [3446]. Suggestion [3645]. Suicide [3217]. Summary [640, 666]. Summer [668]. Sun [3222]. Sunt [1308]. Supermarket [182]. Supernatural [1223]. Superpowers [1574, 1564]. Superstar [3936]. Supper [3145]. Supply [4058, 3420, 4156, 3732, 3739]. Support [1362, 1201, 3921, 3919, 3592, 2288, 1604, 521, 3590, 3920, 1191, 3917, 2667]. Supporting [1022]. Supreme [3100, 3584, 3889, 607, 3498, 3067, 3606]. Surge [4041]. Surgical [802]. Surrender [2575]. surrendering [3440]. Surrogate [2168]. Surveillance [3748]. Survey [775, 3890, 389, 2313, 284]. Surveys [3702]. Survival [4024, 2603, 4114, 2031, 92, 1965, 1410, 4196]. Survive [1285]. sus [4187]. Susan [2961]. Sustainability [3599, 4068, 3577]. Sustainable [2297, 2661, 3741]. Sustained [3945]. Sustaining [3436, 4209, 2525]. Sustenance [4155]. Swain [1722]. Sweden [2071, 2451, 1766]. Sweet [2583]. Swept [4194]. Swing [3642]. Swords [2311]. Symbol [957, 60, 1374]. Symbolic [150]. Symbolism [319, 66]. Symmetry [164]. Symptom [4046]. Synthesis [42, 685, 1374]. Synthetic [3288]. Syphilis [3062]. Syria [3851, 3784]. System [2033, 4130, 754, 1235, 424, 1878, 2228, 3827, 3901, 3706, 3896, 238, 2173, 661, 797, 936, 3774, 1510, 1936, 1412, 765, 1598, 881, 662, 1503, 871, 3911, 2657, 4082, 3834, 3426, 1374, 570, 2412, 3464]. Systemic [4237, 3882, 3874]. Systems [1850, 1848, 2521, 1833, 4086, 4090, 2240, 2238, 2666, 4242, 1159, 2636, 2057, 224, 1830].

Tables [1262]. Taborsky [3322]. Tail [1893]. Taipei [3490]. Take [3094, 2034]. Taken [4096]. Taking [3414, 4172]. Tale [1714]. Talent [517]. Tales [2102]. Talk [4183, 57, 4010, 4161, 3462]. Talking [3947]. Tanzania [2731]. Tape [3257]. Targeted [3779]. Targeting [3783, 3788]. Task [2735, 1521, 4008, 4137]. Tasks [230]. Taught [3028]. Tea [3509, 3486]. Teach [1606, 3979, 3950, 2861]. Teacher [3944, 516, 3324, 818, 399]. Teacher-Scholars [516]. Teachers [2135, 2906, 2301]. Teaching [1715, 3943, 345, 3190, 1732, 1623, 2689, 3107, 4203, 95, 94, 3946]. Team [888]. Teams [887]. Tech [4168, 4170]. Technical [1967, 4090, 310]. Technics [127]. Technique [1521]. Techniques [221]. Techno [3911, 2452]. Techno-Globalism [2452]. Techno-Optimism [3911]. Technocracy [4052]. Technological [3558, 1578, 1337, 269, 643, 2247]. Technologies [1827, 1845, 3981, 3429, 3594, 1431]. Technology [2711, 3785, 1520, 1341, 1968, 1065, 1516, 1519, 1579, 1526, 1908, 140, 4182, 240, 304, 1515, 485, 1517, 2809, 2517, 2035, 426, 1342, 3564, 506, 1675, 3778, 1524, 1672, 1909, 1525, 470, 1336, 2016, 2610, 4245, 1858, 2452, 2529]. Tectonics [1376]. Television [3244, 353, 1676, 2303, 3369, 1856, 1682, 1680, 1322, 1677, 1407, 1854, 1855]. Televisual [4023]. Tell [4014]. Teller [55]. Telling [1226, 2102]. Tempered [1144]. Tempo [942]. Temptation [2271]. Ten [693, 692, 1182, 2518]. Ten-Year [693, 692]. Tendencies [1400]. Tenor [1922]. Tense [3109]. Tension [1993, 2882, 1457]. Tensions [3986, 4115, 3676, 2651]. Tentative [204]. Tenth [691, 3042]. Tenure [3336]. Terra [1802]. Terrain [3829]. Territorial [2651]. Territoriality [2648]. Terror [2935]. Terrorism [3410, 3212, 2890]. Terzakis [593]. Test [4233, 4234]. Testing [2883, 3861]. Text [1702, 1704]. Textbooks [716]. Texts [1994]. Thailand [1007]. Thaw [201]. Theater [4080, 2782]. Theatre [2105]. Theft [3410]. Their [2300, 2410, 3341, 1203, 35, 994, 122, 4189, 123, 349, 4181, 854, 3363, 156, 3445]. Them [4222, 2728, 3713, 2792, 2366, 4244]. Thematic [4247]. Theme [383, 372, 1785, 2218]. Themes [1495, 2192, 378, 103, 1101]. Theocratic [2212]. Theodore [1678]. Theology [602, 3998]. Theoretical [3603]. Theories [3004, 1959, 104, 4017, 2163]. Theorists [3814]. Theory [3561, 2424, 2578, 514, 1386, 648, 4097, 256, 853, 910, 1502, 1053, 1050, 3797, 1648, 3870, 881, 864, 1710, 3231, 2888, 498, 321, 738, 1080, 1384, 3793]. Therapeutic [4023, 802, 1337]. Therapy [4220]. There [3542, 1737, 1568, 4163, 2464, 601, 3657, 3203, 3434]. Thermal [31, 30]. Thermodynamics [594]. Thief [2303]. Thing [4022]. Things [1722, 4241, 3298, 3746]. Think [3833, 1814, 4087, 2698]. Thinkable [3301]. Thinking [4063, 2265, 1694, 1160, 2269, 660, 1863]. Third [2991, 601, 2828, 2218]. Thirtieth [2180]. Thirty [2179, 126, 2183, 2178, 641]. Thirty-Three [641]. Thomas [2355, 2358, 2361, 2364, 2367, 2369]. Thorstein [1080]. Those [868, 1757, 4222]. Thought [1218, 35, 3316, 1451, 1309, 429, 2397, 4247, 361]. Thoughts [2054, 2296, 1125, 1199, 1786, 2326, 2205, 812, 2009, 540, 1998]. Thousand [2982]. Thread [4143, 2517]. Threat [3535, 2402, 3254, 4032, 3831, 3379]. Threatens [3901]. Threats [3959, 3811, 3337, 3344, 2680, 1430, 3829]. Three [3290, 1456, 62, 999, 2002, 1907, 2866, 35, 641, 2042, 1089, 105, 727]. ThriveNYC [4220]. Tidal [1653]. Tide [412]. Time [2956, 2243, 2950, 2303, 2959, 2960, 1435, 2953, 4018, 4028, 2955, 2516, 1357, 4131, 3948, 2146, 3939, 2517, 3644, 2958, 1372, 1523, 4008, 3642, 916]. Timeliness [3926]. Times [2484, 1180, 1930, 3771, 3788]. Tinsel [1495]. Titan [1109]. Today [2808, 3493, 280, 297, 2419, 2100, 3039, 893, 2196, 3101, 2326, 817, 148, 2094, 2396, 205, 2803, 2718, 2888, 783, 3705, 1101, 429, 2397, 1521]. Together [3947]. Tokugawa [731, 2648]. Tolerance [3079]. Toleration [2268]. Tolstoi [1291]. Tomorrow [2656, 4172, 683, 783]. Tone [4033]. Tongue [3540, 2690]. Tongues [1480]. Tons [3589].

Too [2725, 3559, 2403, 2832, 4124]. Tool [4076, 3086]. Toolbox [3785]. Tools [4004]. Top [2950]. Torchlight [3369]. Torque [3642]. Torture [3148, 3035]. Total [334]. Totalism [361]. Totalitarianism [3214]. Tottering [3111]. Tourism [2812]. Tower [844, 1733, 4064]. Track [4028]. Tracking [2558]. Trade [339, 1013, 2877, 3880, 321]. Tradition [1016, 1668, 1306, 2043, 969, 1005, 102, 550, 575, 968, 1009, 587, 1287, 975, 702, 972, 880, 970, 1007, 2858, 1015]. Traditional [1008, 2845, 2201, 1005, 1004, 1013, 3896, 578, 366, 1015, 2766]. Traditions [1008, 2777, 974, 1014, 248, 980, 2844, 2114, 2570, 3128, 970, 916]. Tragedy [1491, 3752]. Trained [1888]. Training [3946]. Trajectory [634, 3508]. Transaction [765]. Transaction-Maximizing [765]. Transactional [4160]. Transactions [2488, 2487]. Transatlantic [427]. Transcendence [1221, 1214, 1219]. Transcending [3609, 2325, 1533]. Transcript [556]. Transcripts [427]. Transformation [2556, 1877, 2825, 2307, 2551, 2811, 1079, 3717, 3043, 1007, 2714]. Transformational [4243]. Transformative [2539, 2662]. Transformed [1269, 3973]. Transforming [2687]. Transgressing [1538]. Transgression [1539, 702]. Transition [3629, 724, 3588, 378, 2351, 721, 2446, 918, 2259, 1637, 3775, 841, 2404, 3565, 2256]. Transitional [3997]. Transitions [2645]. Transitory [2545]. Translating [3756]. Translation [3753, 963, 1052, 1863]. Transmission [4075, 2074]. Transmitted [2170]. Transnational [3830, 3966, 3595, 3829, 2747]. Transplantation [2730]. Trap [4101, 3631]. Trauma [1397, 4221, 587, 4075, 61, 4039]. Traumatic [2494]. Treasured [1875]. Treating [3853]. Treatment [716, 2273, 3989]. Treaty [1849, 1850, 1848, 3982]. Tree [2196].

Trembling [3312]. Tremors [2473]. Trenches [3823]. Trend [3217, 1577, 2267]. Trends [2821, 832, 424, 265, 2274, 731, 3071, 2288, 3787, 1475, 2560, 2522, 3363]. Trial [4211, 456, 2789]. Trials [3688, 2994, 3795, 802, 1646]. Tribal [3864, 1459]. Tribes [3861]. Tribulations [2994]. Tribute [1050]. Trillion [3589]. Trip [3239]. Triumph [680]. Trollope [1454]. Tropical [1650]. Troubled [2699, 3456]. Troubles [2598, 1203]. True [1269, 3559, 2468]. Truly [1701, 1982]. Trump [4038, 3960]. Trust [4130, 4129, 3528, 4182, 4131, 3333, 4128, 4140, 4135, 4133, 4139, 4136, 4137, 4134]. Trusting [3686]. Trustworthiness [3528]. Trustworthy [4138]. Truth [4128, 3040, 3168, 3994]. Truths [2282]. Trying [2765]. Tune [2996]. Tunisian [3775]. Turf [1013]. Turing [4101]. Turkey [2436, 705, 1011, 2974]. Turkish [1012]. Turn [1594, 3278, 2847]. Turned [4132]. Turning [1096, 3291]. Turtle [64]. Tweed [3872]. Twelfth [3357]. Twelve [449, 3815, 941, 942]. Twentieth [908, 1074, 260, 2437, 3164, 3638, 1324, 3188, 2200, 2448, 2314]. Twentieth-Century [1074, 2437, 3638, 1324, 2200, 2448]. Twenty [3877, 2467, 3856, 3516, 3503, 784, 2710, 2720, 2542]. Twenty-First [3877, 2467, 3856, 3503, 784, 2710, 2720, 2542]. Twenty-First-Century [3516]. Twice [748]. Twice-Born [748]. Twilight [3923]. Two [827, 1139, 577, 855, 2702, 1158, 804, 3477, 531, 3125, 3478, 2466, 2571, 4225, 1206, 313, 4017, 2803, 1664, 440, 1060, 3484, 2137, 2525, 2366, 1409, 2531]. Types [287]. Typology [2719]. Tyranny [1555].

U [1594]. U-Turn [1594]. U.S [1558, 1572, 4216]. U.S. [1844, 3863, 3173, 1561, 3598, 3435, 1562, 3888, 4242, 675, 3783, 3513, 4140, 3522, 3421, 3499, 3958, 3790, 1831, 3771, 779, 4052, 768].

U.S.A [460]. U.S.S.R. [1033]. Uganda [2449, 3824]. Ukraine [4177, 2586]. Ultras [3215]. Unapproved [2884]. Unbound [1103]. Unceasing [4033]. Uncertain [2033, 1124, 1176, 4053, 3894, 1527, 3031]. Uncertainties [2600]. Uncertainty [3895, 2930, 4069]. Unchaining [4158]. Uncharted [1199]. Uncivil [2571]. Uncommon [3683]. Unconscious [514, 3083, 375, 373]. Uncovering [4236]. Underclass [2424]. Underdevelopment [743]. Undergraduate [1147, 3945, 1145, 3949, 3950, 1149, 3107]. Underground [1511]. Undermine [3333]. Underpinnings [3248]. Underrepresentation [3958]. Underserved [4059]. Understand [4095, 3698, 1048]. Understanding [3712, 1224, 3395, 3950, 816, 375, 4091, 809, 2796, 4009, 2209]. Understandings [1849]. Understood [3607]. Undesirables [2424]. Uneasy [3163]. Unequal [2859, 2455]. Unfaithful [2303]. Unfinished [2909]. Unfit [4024]. Unfolding [3856]. Unhealthy [2979]. Unification [384, 2362, 2368, 2357]. Unifying [1785, 4136]. Unilateral [1849, 229]. Union [1849, 1850, 1848, 1252, 1841, 2257, 984, 203, 3849]. Unions [697, 3612]. Uniquely [3057]. Uniqueness [1260]. Uniquenesses [2598]. United [2356, 2360, 989, 1332, 1817, 992, 1849, 1850, 1848, 1947, 991, 335, 775, 404, 225, 1565, 1363, 527, 1344, 1144, 1591, 337, 1594, 2071, 2226, 3922, 3854, 398, 2171, 3071, 3718, 2146, 1025, 996, 3610, 627, 2577, 984, 2658, 2227, 2426, 1249, 3962, 4033, 1472, 2861, 63, 2068, 3628, 1104, 900, 3622, 1032, 2204, 1681, 4200, 300]. Uniting [1148]. Unity [77, 2617, 2638, 45, 480, 1184, 1384]. Universal [3668, 2790, 2593, 614, 841]. Universality [3569]. Universals [3054]. Universe [2945, 3699, 1372, 3702]. Universities [834, 933, 835, 3864, 2338, 3190, 3892, 874, 2602, 464, 2288, 3326, 465, 2346, 1191, 2341, 2347, 2339, 2344]. University [2552, 199, 460, 840, 867, 2337, 1203, 700, 870, 1162, 3946, 847, 1150, 1210, 1141, 1135, 1208, 845, 1158, 938, 1166, 1200, 843, 649, 1205, 1204, 1163, 1186, 2343, 2342, 1995, 844, 1195, 310, 1398, 1133, 1028, 1155, 385, 1032, 463, 1184, 468, 467]. Unjust [3781, 2214, 3779]. Unknown [3111]. Unlikely [3809]. Unmaking [1010]. Unmasking [2492]. Unmet [3908]. Unprecedented [272]. Unpredictability [3695]. Unprepared [2383]. Unpublished [129]. Unrest [1154]. Unrevolution [2329]. Unruly [1494]. Unscientific [1730]. Unseeing [2036]. Unseen [4231]. Unselfish [2923, 2855]. Unsent [3165]. Unsettled [2918]. Unsolved [513]. Unspoken [2475]. Unsustainability [3730]. Untrue [2756]. Untruth [1493]. Unum [1632, 3569]. upon [189, 2566, 1932, 916]. Ups [1484]. Uptight [952]. Urban [1068, 2037, 757, 761, 240, 4074, 4060, 3732, 1195, 770, 4057, 2057, 2461, 646, 755, 4179, 913]. Urbanism [3374]. Urbanized [580]. Urbanizing [756]. Urgency [579]. US-German [2284]. Use [1722, 2634, 804, 161, 3782, 904, 2011, 2927, 373, 2080]. Used [2764]. Useful [2144, 1967]. User [3531]. Uses [254, 244, 3781, 594, 1163, 817, 3820, 2490, 66, 1825]. Using [3561, 1031, 4238, 3698]. USSR [206, 1473]. Usual [4007]. Utilitarians [747]. Utility [501, 3603, 1108]. Utopia [497, 493, 2936, 503, 500, 494, 504, 1084, 498, 505, 2005, 499, 502]. Utopian [248]. Utopias [501, 118, 496, 495].

v [3067]. Vadis [41]. Vain [3302]. Value [4160, 857, 3196, 3335, 3284, 4099, 3102, 1278]. Values [1397, 1166, 1967, 1789, 2683, 52, 617, 825, 2593, 156]. Valuing [3616]. Vanishes [1142]. Vanishing [3158]. Variability [3144]. Variable [3055, 1779, 2302]. Variables [645]. Variation [1040, 942]. Variations [3987, 655]. Varieties [985, 3772, 496, 2645]. Various [3641, 1025]. Veblen [1080]. Veena [2488]. Velvet [2380]. Ventura [2779]. Verbal [963]. Verification [965]. Vernacular [2646]. Versatile [1662]. versus [1208, 1975, 130, 2447, 330, 4065, 2451, 972, 4243, 247]. Very [3137, 1797, 1985]. Vicissitudes [2252]. Victim [1505]. Victims [2697]. Victorians [4023]. victories [3459]. Victory [3186, 2395]. Video [1671, 4027]. vie [3653]. viel [2589, 2588]. Vienna [1457]. Vietnam [1680, 2651]. View [2773, 1819, 2702, 3429, 2644, 2495, 3405, 38, 531, 3125, 3478, 1136, 34, 2186, 996, 479, 3513, 2901, 1305, 782, 3591, 3858, 534, 1596, 822]. Viewed [273, 2559]. Viewing [791]. Viewpoint [571]. Views [154, 1301, 1889, 994, 519]. Vignettes [2647]. Villa [560, 2443]. Vince [2870]. Vinci [2624]. Violence [1511, 4072, 2495, 3833, 1941, 3965, 4074, 4071, 4077, 4075, 3787, 3209, 3216, 2758, 4076, 111, 4217, 4184, 4073, 4070, 3213, 3379]. Virgil [963]. Virginia [3654]. Virtue [1997, 1952, 3020]. Virtues [1908, 3055, 2892]. Visibility [4232]. Visible [4077]. Vision [1090, 4088, 1829, 511, 2486, 3043, 1938, 3707]. Visions [2466, 1518, 3484]. Visit [1618, 1620]. Visits [3461]. Visual [963, 3553, 148, 483]. Vita [357]. Vivum [157]. Vocation [2040, 1169, 44]. Voice [4209, 2485, 3533]. Voices [1765, 2103, 1916, 3553, 1764, 3870, 4116, 2647]. Vol [1822, 1836]. Vollmar [2364]. Volta [4062]. Voltaire [36]. Volume [80, 146, 235, 290, 342, 407, 474, 543, 597, 690, 787, 839, 892, 956, 1003, 1073, 1175, 1265, 1331, 1394, 1466, 1514, 1571, 1626, 1686, 1751, 1801, 1866, 1925, 1972, 2046, 2110, 2177, 2235, 2291, 2350, 2416, 2481, 2550, 2614, 2670, 2739, 3419, 1124, 1572, 1176]. Voluntarism [1930]. Voluntary [3712, 1637]. Volunteer [2942, 3516]. Volunteerism [3721]. Vortex [2690]. Vote [3957]. Votes [1013]. vs [2882, 3212, 1069, 3814, 3211, 3102]. Vulgar [680]. Vulnerable [2815].

W [126, 2060, 3038, 1847]. Wages [1061, 1769]. Wagging [1893]. wake [3472]. Waldorf [3647]. Wales [2217]. Walk [4010]. Wall [4244]. Waller [3644]. Walloons [3218]. Walsh [1888]. Walter [1272]. Waning [1078, 2439, 2925, 1406]. Want [3274, 4082]. Wanted [2787]. War [3517, 2258, 4188, 1300, 334, 230, 1231, 3985, 1094, 2144, 1363, 2437, 3794, 3827, 3781, 3846, 3853, 3844, 1507, 1961, 2188, 3843, 217, 3797, 3783, 3513, 1680, 3795, 1679, 1929, 914, 3848, 3778, 3789, 3791, 4176, 2918, 2979, 3788, 2281, 3793, 3089, 3518, 3831, 3799, 324, 3379, 3841, 2749, 3828, 2216]. Warburg [1277]. Warfare [3980, 3780, 3210, 3779, 3793]. Warming [2959, 3031]. Warning [3786]. Wars [3830, 2266, 2604, 3854, 3013, 3828, 2606, 3829, 3838, 2895, 3834, 2900]. Was [3094]. Washington [1615]. Wasp [3461]. Waste [2523]. Watchfulness [1938]. Water [3733, 4058, 3728, 4063, 3726, 4059, 4056, 3731, 4067, 4060, 3732, 4057, 4065, 2844, 3730, 4069, 4064, 3734, 3727]. Water-Supply [3732]. Waters [3546]. Wave [1653]. Waves [3654]. Way [3517, 2765, 3640, 3685, 1666, 3287, 1746, 2897, 1407, 2553, 3123, 3983, 1796, 3]. Ways [2311, 1285, 3856, 2137]. Weaken [266]. Weaknesses [768]. Wealth [1592, 2400, 2951, 4099, 1812]. Weapon [3982]. Weapons [1236, 1845, 1838, 2439, 3974, 1822, 1836, 1832, 919, 3780, 3782, 3413, 3514, 3414, 3411, 117]. Weather [951]. Wedding [2865]. Weighing [1832]. Weight [629]. Welfare [1758, 1743, 3186, 1505, 2915]. We’ll [2034, 3021, 4157, 2410, 3607, 1618, 2167, 3308, 1442, 3464]. Well-Being [3021, 2410, 3308]. Well-Run [1618]. We’re [3509, 3486, 868, 3298, 2787, 1796, 3203, 3488]. Werner [360]. West [1075, 409, 1949, 360, 1655, 2392, 1786, 1667, 630, 2383]. Western [2993, 3836, 3546, 260, 1469, 1842, 3849, 2623, 1473, 583, 617, 1472, 3591, 279, 702, 898, 2570, 3128, 420, 582]. Whatever [2764]. Where [3353, 764, 2639, 2730, 2457, 2850, 3298, 3065, 1899, 1138, 1899]. Wherever [3297]. Which [2387, 3335, 1888]. while [1209]. Whimpers [1786]. White [1115, 2420, 863, 3496, 742, 3068, 553, 3636, 4041, 4209, 4030, 4029]. Whites [937]. Whither [1552, 1693, 2324]. Who [1606, 2032, 2862, 1808, 4222, 2145, 1502, 2383, 3533, 540, 2163, 4244, 3488, 2802]. Whole [14, 1858]. Whom [1652, 2383]. Whose [3520]. Wide [1165, 3641, 2476, 117]. Wide-Angle [2476]. Wider [1712, 673, 297]. Widowhood [2489]. WikiLeaks [3535]. Wild [3298, 2576]. Wilhelmian [1118]. Will [1606, 3869, 3949, 741, 3333, 1131, 2472, 1028, 810, 2491]. William [1090, 748]. Williams [860]. Wilson [1273]. Wind [1667, 2562]. Winds [2928]. Wine [2777]. Wing [4191]. Winning [3792]. Winslow [1931]. Wisdom [1213, 2844, 3440]. Wish [2478]. Withdrawal [255]. Within [3143, 2559, 2981, 2413, 2776, 871, 373]. Without [4143, 2754, 3937, 230, 437, 3688, 2575, 134, 1454, 1185, 3401, 3982, 2389, 3791, 3411, 1535, 2021, 2347, 3356]. Witness [2298, 861]. Witnesses [4013]. Witnessing [4000, 3999, 4002, 4003]. WMD [3377]. Woman [4022, 1082, 1969, 434, 3461, 1158]. Womanhood [4022, 4207, 435]. Women [439, 3737, 437, 3967, 1957, 1960, 3965, 1961, 3961, 1963, 3971, 1482, 3963, 2792, 438, 3958, 3771, 3957, 4184, 2282, 3452, 1158, 1952]. Women-Led [4184]. Womenkind [1666]. Wonders [4096]. Won’t [3050]. Wood [2844]. Woods [3184]. Woolf [3654]. Word [1485, 2891]. Words [813, 1021, 2013, 4094, 1863]. Work [2775, 669, 998, 1319, 3103, 762, 317, 633, 2694, 3105, 3104, 116, 3964, 1399, 149, 1880, 2828, 3150, 4087, 4100, 4169, 156]. Worker [762]. Workers [4158, 3669]. Working [668, 4156, 107, 3714, 438, 420, 1110]. Workplace [4242]. Workplaces [2602]. Works [3183]. World [566, 125, 1947, 2284, 2764, 3728, 230, 1066, 1508, 1411, 2950, 3314, 2639, 950, 1094, 2438, 3865, 1859, 1363, 580, 2198, 536, 608, 3974, 451, 564, 38, 559, 908, 2047, 2442, 572, 4007, 2924, 560, 2443, 2936, 320, 34, 1960, 3526, 661, 3773, 567, 49, 948, 3047, 2058, 2232, 231, 2545, 3852, 3413, 237, 569, 2971, 1679, 3973, 980, 89, 581, 2976, 571, 88, 3414, 2750, 702, 3838, 1888, 3407, 2218, 3411, 2434, 1447, 579, 3427, 117, 3327, 3736, 106, 2836, 570, 450, 2847, 2783, 3881, 3356, 4015, 2881, 1412, 2867, 582]. World-Making [3314]. World-Wide [117]. Worlds [577, 3568, 2598, 3477, 531, 3125, 3478]. Worms [3138]. Worn [1128]. Worn-out [1128]. Worry [2898]. Worse [1332, 1340]. Worth [3285, 2325]. Would [3640, 2698, 3488]. Wretched [1274]. Wright [360]. Write [4016, 4190]. Writer [1768, 1320, 3015]. Writers [201]. Writes [1962]. Writing [1782, 905, 1226]. writings [3475]. Wrong [448, 354].

X [1385]. X-Ray [1385].

Yan’an [2318]. Yang [2536]. Yanking [4143]. Year [2179, 667, 693, 692, 634, 633, 649, 656, 651, 657, 2180, 1564]. Yearning [102]. Years [677, 2671, 4129, 126, 951, 2420, 2183, 2178, 2551, 1182, 641, 2260, 885, 1871, 1888, 2525, 1928]. Yeats [1090]. Yellow [3636]. yes [3446]. Yesterday [2656, 1521]. Yield [3739]. Yin [2536]. Yoked [1903]. York [1142, 2355, 2358, 2361, 2364, 2367, 2369, 1916, 3625]. Young [4041, 2864, 4235, 947, 944, 2801, 361].

You’re [3944]. Youth

[3505, 297, 293, 292, 654, 849, 2003, 291, 298, 299, 294, 296, 755, 301, 300]. Yugoslavia [2258, 2257, 2383]. Yusef [3932].

Z [3490]. Zarathustra [457]. Zeal [1793]. Zero [1068, 360, 1065, 1058, 3253]. Zhivago [207]. Zone [3923, 4006]. Zones [4179]. Zoning [3253]. ZPG [1060].

References

Whitehill:1955:FD

[1] Walter Muir Whitehill. A foreword to “Dædalus”. Dædalus, 86(1):3–5, May 1955. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                              tronic). URL http://www.jstor.org/
                                              stable/20026384.

McCord:1955:D

[2] David McCord. Dædalus. Dædalus, 86(1):6–8, May 1955. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                              www.jstor.org/stable/20026385.

Bok:1955:CHI

[3] Bart J. Bok, Michiel J. Bester, and Campbell M. Wade. Catalogue of H(II) regions in the Milky Way, for longitudes 250^{∘}–355^{∘}. Dædalus, 86(1):9–16, May 1955. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                              tronic). URL http://www.jstor.org/
                                              stable/20026386.

Franck:1955:PPP

[4] James Franck. Physical problems of photosynthesis. Dædalus, 86(1):17–42, May 1955. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026387.

Brown:1955:CRP

[5] Sanborn C. Brown. Count Rumford on photosynthesis. Dædalus, 86(1):43–46, May 1955. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026388.

Whitehill:1955:PPG

[6] Walter Muir Whitehill, Julian P. Boyd, Leonard W. Labaree, L. H. Butterfield, Wilmarth S. Lewis, and Waldo G. Leland. Publishing the papers of great men. Dædalus, 86(1):47–79, May 1955. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026389.

Gaertner:1955:AFA

[7] Johannes A. Gaertner. Art as the function of an audience. Dædalus, 86(1):80–93, May 1955. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026390.

Anonymous:1955:RM

[8] Anonymous. Records of meetings. Dædalus, 86(1):94–99, May 1955. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026391.

Washington:1955:D

[9] G. Washington and Albert Schweitzer. Documents. Dædalus, 86(1):100–103, May 1955. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026392.

Anonymous:1955:FM

[10] Anonymous. Front matter. Dædalus, 86(1):??, May 1955. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026383.

Anonymous:1956:OCN

[11] Anonymous. Officers, committees, and new members, as elected on 9 May 1956. Dædalus, 86(2):1–8, September 1956. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026401.

Whitehill:1956:F

[12] Walter Muir Whitehill. Foreword. Dædalus, 86(2):107–109, September 1956. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026395.

Shapley:1956:FAH

[13] Harlow Shapley. A farewell to the Academy House on Newbury Street. Dædalus, 86(2):110–112, September 1956. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026396.

Haskins:1956:SWM

[14] Caryl P. Haskins. Science and the whole man. Dædalus, 86(2):113–121, September 1956. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20026397.

Rosenberg:1956:RDC

[15] Jakob Rosenberg. Rembrandt the draughtsman with consideration of the problem of authenticity. Dædalus, 86 (2):122–136, September 1956. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026398.

Anonymous:1956:RM

[16] Anonymous. Records of meetings. Dædalus, 86(2):137–166, September 1956. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026399.

Copley:1956:D

[17] John Singleton Copley and John Adams. Documents. Dædalus, 86(2): 167–168, September 1956. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

/www.jstor.org/stable/20026400.

Anonymous:1956:FM

[18] Anonymous. Front matter. Dædalus, 86(2):??, September 1956. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026394.

Whitehill:1957:Fa

[19] Walter Muir Whitehill. Foreword. Dædalus, 86(3):171–173, May 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026404.

Burchard:1957:ISF

[20] John E. Burchard. Imperium sine fine. Dædalus, 86(3):174–189, May 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026405.

Millin:1957:MBF

[21] Terence Millin. Man, beast, and field: Some influences of the Francis Amory Prize researches. Dædalus, 86(3):190–205, May 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026406.

Medina:1957:LEA

[22] Harold R. Medina. A liberal education and the advancement of American freedom. Dædalus, 86(3):206–213, May 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026407.

Bishop:1957:FF

[23] Morris Bishop. Franklin in France. Dædalus, 86(3):214–230, May 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20026408.

Biggs:1957:BFA

[24] E. Power Biggs. Benjamin Franklin and the Armonica. Dædalus, 86 (3):231–241, May 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026409.

Yee:1957:CP

[25] Chiang Yee. The Chinese painter. Dædalus, 86(3):242–252, May 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20026410.

Anonymous:1957:RM

[26] Anonymous. Records of meetings. Dædalus, 86(3):253–280, May 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026411.

Anonymous:1957:FMa

[27] Anonymous. Front matter. Dædalus, 86(3):??, May 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026403.

Whitehill:1957:Fb

[28] Walter Muir Whitehill. Foreword. Dædalus, 86(4):283–284, October 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026414.

Taylor:1957:ASR

[29] Francis Henry Taylor. The archaic smile: The relation of art and the dignity of man. Dædalus, 86(4): 285–322, October 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026415.

Chandrasekhar:1957:TC

[30] S. Chandrasekhar. Thermal convection. Dædalus, 86(4):323–339, October 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026416.

Brown:1957:CRD

[31] Sanborn C. Brown. Count Rumford discovers thermal convection. Dædalus, 86(4):340–343, October 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026417.

Gardner:1957:AF

[32] Robert Gardner. Anthropology and film. Dædalus, 86(4):344–352, October 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026418.

Anonymous:1957:FMb

[33] Anonymous. Front matter. Dædalus, 86(4):??, October 1957. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026413.

Holton:1958:PIS

[34] Gerald Holton. Perspectives on the issue “Science and the Modern World View”. Dædalus, 87(1):3–7, Winter 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026421.

Guerlac:1958:TEC

[35] Henry Guerlac. Three Eighteenth-Century social philosophers: Scientific influences on their thought. Dædalus, 87(1):8–24, Winter 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026422.

Brown:1958:SHC

[36] Harcourt Brown. Science and the human comedy: Voltaire. Dædalus, 87(1):25–34, Winter 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026423.

deSantillana:1958:SCL

[37] Giorgio de Santillana. The Seventeenth-Century legacy: Our mirror of being. Dædalus, 87(1):35–56, Winter 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026424.

Frank:1958:CSC

[38] Philipp Frank. Contemporary science and the contemporary world view. Dædalus, 87(1):57–66, Winter 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026425.

Oppenheimer:1958:GSS

[39] Robert Oppenheimer. The growth of science and the structure of culture: Comments on Dr. Frank’s paper. Dædalus, 87(1):67–76, Winter 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026426.

Bruner:1958:FCM

[40] Jerome S. Bruner. The Freudian conception of man and the continuity of nature. Dædalus, 87(1):77–84, Winter 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026427.

Bridgman:1958:QV

[41] P. W. Bridgman. Quo vadis. Dædalus, 87(1):85–93, Winter 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026428.

Morris:1958:PNS

[42] Charles Morris. Prospects for a new synthesis: Science and the humanities as complementary activities. Dædalus, 87(1):94–101, Winter 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026429.

Jones:1958:HLS

[43] Howard Mumford Jones. A humanist looks at science. Dædalus, 87 (1):102–110, Winter 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026430.

Weber:1958:SV

[44] Max Weber. Science as a vocation. Dædalus, 87(1):111–134, Winter 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026431.

Mather:1958:NUW

[45] Kirtley F. Mather. On the nature of the unity we seek. Dædalus, 87 (1):135–138, Winter 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026432.

Anonymous:1958:BMa

[46] Anonymous. Back matter. Dædalus, 87(1):??, Winter 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026433.

Anonymous:1958:FMa

[47] Anonymous. Front matter. Dædalus, 87(1):??, Winter 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026420.

Rostow:1958:PIA

[48] W. W. Rostow. Preface to the issue “The American National Style”. Dædalus, 87(2):3–4, Spring 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026436.

Kennan:1958:AAR

[49] George F. Kennan. America’s administrative response to its world problems. Dædalus, 87(2):5–24, Spring 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026437.

Murray:1958:IMC

[50] Henry A. Murray. Individuality: The meaning and content of individuality in contemporary America. Dædalus, 87(2):25–47, Spring 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026438.

Kaplan:1958:AEP

[51] Abraham Kaplan. American ethics and public policy. Dædalus, 87(2):48–77, Spring 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026439.

Kluckhohn:1958:ECA

[52] Clyde Kluckhohn. The evolution of contemporary American values. Dædalus, 87(2):78–109, Spring 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026440.

Rostow:1958:ANS

[53] W. W. Rostow. The American national style. Dædalus, 87(2):110–144, Spring 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026441.

Lawrence:1958:PS

[54] D. H. Lawrence. The plumed serpent. Dædalus, 87(2):145–146, Spring 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026442.

Anonymous:1958:FDD

[55] Anonymous. Fallout and disarmament: a debate between Linus Pauling and Edward Teller. Dædalus, 87 (2):147–163, Spring 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026443.

Bohr:1958:AHK

[56] Niels Bohr. On atoms and human knowledge. Dædalus, 87(2):164–175, Spring 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026444.

Bridgman:1958:RNB

[57] P. W. Bridgman. Remarks on Niels Bohr’s talk. Dædalus, 87(2):175–177, Spring 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026445.

Anonymous:1958:BMb

[58] Anonymous. Back matter. Dædalus, 87(2):??, Spring 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026446.

Anonymous:1958:FMb

[59] Anonymous. Front matter. Dædalus, 87(2):??, Spring 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026435.

Tillich:1958:RS

[60] Paul Tillich. The religious symbol. Dædalus, 87(3):3–21, Summer 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026449.

Wilder:1958:CST

[61] Amos N. Wilder. The cross: Social trauma or redemption. Dædalus, 87(3):22–36, Summer 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026450.

Burke:1958:FTC

[62] Kenneth Burke. On the first three chapters of Genesis. Dædalus, 87 (3):37–64, Summer 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026451.

Parsons:1958:PRO

[63] Talcott Parsons. The pattern of religious organization in the United States. Dædalus, 87(3):65–85, Summer 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026452.

Richards:1958:SPS

[64] I. A. Richards. The sense of poetry: Shakespeare’s “The Phoenix and the Turtle”. Dædalus, 87(3):86–94, Summer 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026453.

Heisenberg:1958:RNC

[65] Werner Heisenberg. The representation of nature in contemporary physics. Dædalus, 87(3):95–108, Summer 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026454.

Whitehead:1958:US

[66] A. N. Whitehead. Uses of symbolism. Dædalus, 87(3):109–123, Summer 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026455.

Anonymous:1958:BMc

[67] Anonymous. Back matter. Dædalus, 87(3):??, Summer 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026456.

Anonymous:1958:FMc

[68] Anonymous. Front matter. Dædalus, 87(3):??, Summer 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026448.

Lerner:1958:PIE

[69] Daniel Lerner. Preface to the issue “On Evidence and Inference”. Dædalus, 87(4):3–10, Fall 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026460.

Aron:1958:EIH

[70] Raymond Aron, Suzanne Keller, and Mrs. Judith K. Davison. Evidence and inference in history. Dædalus, 87(4):11–39, Fall 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026461.

Hart:1958:EIL

[71] Henry M. Hart, Jr. and John T. Mc-Naughton. Evidence and inference in the law. Dædalus, 87(4):40–64, Fall 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026462.

Erikson:1958:NCE

[72] Erik H. Erikson. The nature of clinical evidence. Dædalus, 87(4):65–87, Fall 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026463.

Deutsch:1958:EIN

[73] Martin Deutsch. Evidence and inference in nuclear research. Dædalus, 87(4):88–98, Fall 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026464.

Lazarsfeld:1958:EIS

[74] Paul F. Lazarsfeld. Evidence and inference in social research. Dædalus, 87(4):99–130, Fall 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026465.

Fine:1958:SP

[75] Jacob Fine. In search of a poison. Dædalus, 87(4):131–147, Fall 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026466.

Jevons:1958:PS

[76] W. S. Jevons. Principles of science. Dædalus, 87(4):148–154, Fall 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026467.

Bruner:1958:CUL

[77] Jerome S. Bruner, Philipp Frank, Paul Tillich, Lewis Mumford, Ernest Nagel, Walter Rosenblith, and Henry A. Murray. A colloquy on the unity of learning. Dædalus, 87(4):155–165, Fall 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026468.

Anonymous:1958:BMd

[78] Anonymous. Back matter. Dædalus, 87(4):??, Fall 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026469.

Anonymous:1958:FMd

[79] Anonymous. Front matter. Dædalus, 87(4):??, Fall 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026459.

Anonymous:1958:VI

[80] Anonymous. Volume information. Dædalus, 87(4):??, Fall 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026458.

Blanshard:1959:IIE

[81] Brand Blanshard. Introduction to the issue “Education in the Age of Science”. Dædalus, 88(1):3–6, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026472.

Hook:1959:ECE

[82] Sidney Hook. The ends and content of education. Dædalus, 88(1):7–24, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026473.

Shuster:1959:WE

[83] George N. Shuster. What is education? Dædalus, 88(1):25–39, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026474.

Bush:1959:EH

[84] Douglas Bush. Education and the humanities. Dædalus, 88(1):40–55, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026475.

Nagel:1959:PSL

[85] Ernest Nagel. The place of science in a liberal education. Dædalus, 88 (1):56–74, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026476.

Bestor:1959:EPR

[86] Arthur Bestor. Education and its proper relationship to the forces of American society. Dædalus, 88(1):75– 90, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20026477.

Childs:1959:EAS

[87] John L. Childs. Education and the American scene. Dædalus, 88(1):91– 106, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20026478.

Niebuhr:1959:EWS

[88] Reinhold Niebuhr. Education and the world scene. Dædalus, 88(1):107–120, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026479.

Morgenthau:1959:EWP

[89] Hans J. Morgenthau. Education and world politics. Dædalus, 88(1):121–138, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026480.

Mead:1959:CGB

[90] Margaret Mead. Closing the gap between the scientists and the others. Dædalus, 88(1):139–146, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026481.

Riesman:1959:ACN

[91] David Riesman. The academic career: Notes on recruitment and colleagueship. Dædalus, 88(1):147–169, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026482.

LeCorbeiller:1959:ESP

[92] Philippe LeCorbeiller. Education in science: Prerequisite for national survival. Dædalus, 88(1):170–174, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026483.

Kemble:1959:PRA

[93] Edwin C. Kemble. Presentation of the Report of the Academy Committee on the Role of Science in American Education. Dædalus, 88(1):175–181, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026484.

Weaver:1959:PIS

[94] Warren Weaver. Purposes and innovations in science teaching. Dædalus, 88(1):182–185, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026485.

Watson:1959:BDP

[95] Fletcher G. Watson. Basic difficulties in present high school science teaching. Dædalus, 88(1):186–191, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026486.

Whitehead:1959:AE

[96] A. N. Whitehead. The aims of education. Dædalus, 88(1):192–205, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026487.

Anonymous:1959:BMa

[97] Anonymous. Back matter. Dædalus, 88(1):??, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026488.

Anonymous:1959:FMa

[98] Anonymous. Front matter. Dædalus, 88(1):??, Winter 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026471.

Murray:1959:IIM

[99] Henry A. Murray. Introduction to the issue “Myth and Mythmaking”. Dædalus, 88(2):211–222, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026491.

Levin:1959:SMM

[100] Harry Levin. Some meanings of myth. Dædalus, 88(2):223–231, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026492.

Campbell:1959:HDM

[101] Joseph Campbell. The historical development of mythology. Dædalus, 88 (2):232–254, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026493.

Eliade:1959:YPP

[102] Mircea Eliade. The yearning for paradise in primitive tradition. Dædalus, 88(2):255–267, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026494.

Kluckhohn:1959:RTM

[103] Clyde Kluckhohn. Recurrent themes in myths and mythmaking. Dædalus, 88(2):268–279, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026495.

Dorson:1959:TMF

[104] Richard M. Dorson. Theories of myth and the folklorist. Dædalus, 88 (2):280–290, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026496.

Wolff:1959:TRM

[105] Robert Lee Wolff. The three Romes: The migration of an ideology and the making of an autocrat. Dædalus, 88 (2):291–311, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026497.

Topitsch:1959:WIS

[106] Ernst Topitsch. World interpretation and self-interpretation: Some basic patterns. Dædalus, 88(2):312–325, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026498.

Lytle:1959:WNM

[107] Andrew Lytle. The working novelist and the mythmaking process. Dædalus, 88(2):326–338, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026499.

McLuhan:1959:MMM

[108] Marshall McLuhan. Myth and mass media. Dædalus, 88(2):339–348, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026500.

Bruner:1959:MI

[109] Jerome S. Bruner. Myth and identity. Dædalus, 88(2):349–358, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026501.

Schorer:1959:NM

[110] Mark Schorer. The necessity of myth. Dædalus, 88(2):359–362, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026502.

Sorel:1959:RV

[111] Georges Sorel. Reflections on violence. Dædalus, 88(2):363–368, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026503.

Mann:1959:DF

[112] Thomas Mann. Doctor Faustus. Dædalus, 88(2):369–373, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026504.

Mann:1959:FF

[113] Thomas Mann. Freud and the future. Dædalus, 88(2):374–378, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026505.

Anonymous:1959:BMb

[114] Anonymous. Back matter. Dædalus, 88(2):??, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026506.

Anonymous:1959:FMb

[115] Anonymous. Front matter. Dædalus, 88(2):??, Spring 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026490.

Holton:1959:FIC

[116] Gerald Holton. Foreword to the issue “Current Work and Controversies”. Dædalus, 88(3):383–384, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026509.

Simons:1959:WWC

[117] Howard Simons. World-wide capabilities for production and control of nuclear weapons. Dædalus, 88(3): 385–409, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026510.

Dubos:1959:MU

[118] Rene J. Dubos. Medical utopias. Dædalus, 88(3):410–424, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026511.

Hoagland:1959:PPC

[119] Hudson Hoagland. Population problems and the control of fertility.

Dædalus, 88(3):425–443, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026512.

Thomas:1959:CPP

[120] John L. Thomas. The Catholic position on population control. Dædalus, 88(3):444–453, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026513.

Bennett:1959:PEP

[121] John C. Bennett. Protestant ethics and population control. Dædalus, 88 (3):454–459, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026514.

Lipset:1959:AIT

[122] Seymour Martin Lipset. American intellectuals: Their politics and status. Dædalus, 88(3):460–486, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026515.

Schlesinger:1959:CAI

[123] Arthur Schlesinger, Jr., Karl W. Deutsch, David Riesman, Talcott Parsons, and Daniel Bell. Comments on “American Intellectuals: Their Politics and Status”. Dædalus, 88(3): 487–498, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026516.

Slichter:1959:OAE

[124] Sumner H. Slichter. Observations on the American economy. Dædalus, 88 (3):499–504, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026517.

Barkin:1959:EPL

[125] Solomon Barkin. Economic policies for a leader of the free world. Dædalus, 88(3):505–517, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026518.

Bridgman:1959:PWB

[126] P. W. Bridgman. P. W. Bridgman’s “The Logic of Modern Physics” after thirty years. Dædalus, 88(3): 518–526, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026519.

Mumford:1959:ALM

[127] Lewis Mumford. An appraisal of Lewis Mumford’s “Technics and Civilization” (1934). Dædalus, 88(3):527–536, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026520.

Cobb:1959:EIC

[128] Edith Cobb. The ecology of imagination in childhood. Dædalus, 88 (3):537–548, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026521.

Dewey:1959:RJD

[129] John Dewey. Reflections of John Dewey: Excerpts from unpublished correspondence. Dædalus, 88(3): 549–559, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026522.

Dupree:1959:FDD

[130] A. Hunter Dupree. The first Darwinian debate in America: Gray versus Agassiz. Dædalus, 88(3):560–569, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026523.

Anonymous:1959:BMc

[131] Anonymous. Back matter. Dædalus, 88(3):??, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026524.

Anonymous:1959:FMc

[132] Anonymous. Front matter. Dædalus, 88(3):??, Summer 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026508.

Anonymous:1959:PNI

[133] Anonymous. Prefatory note on the issue “Quantity and Quality”. Dædalus, 88(4):575–576, Fall 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026528.

Kemeny:1959:MN

[134] John G. Kemeny. Mathematics without numbers. Dædalus, 88(4):577–591, Fall 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026529.

Weisskopf:1959:QQQ

[135] Victor F. Weisskopf. Quality and quantity in quantum physics. Dædalus, 88(4):592–605, Fall 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026530.

Stevens:1959:QS

[136] S. S. Stevens. The quantification of sensation. Dædalus, 88(4):606–621, Fall 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026531.

Leontief:1959:PQQ

[137] Wassily Leontief. The problem of quality and quantity in economics. Dædalus, 88(4):622–632, Fall 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026532.

Lasswell:1959:QQP

[138] Harold D. Lasswell. The qualitative and the quantitative in political and legal analysis. Dædalus, 88(4):633–645, Fall 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026533.

Kubie:1959:PPP

[139] Lawrence S. Kubie. Is preventive psychiatry possible? Dædalus, 88(4):646– 668, Fall 1959. CODEN DAEDAU.

ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20026534.

Deutsch:1959:IST

[140] Karl W. Deutsch. The impact of science and technology on international politics. Dædalus, 88(4):669–685, Fall 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026535.

Poggioli:1959:PS

[141] Renato Poggioli. The pastoral of the self. Dædalus, 88(4):686–699, Fall 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026536.

Morris:1959:R

[142] Clarence Morris. Reputations. Dædalus, 88(4):700–711, Fall 1959. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20026537.

Frost:1959:E

[143] Robert Frost. On Emerson. Dædalus, 88(4):712–718, Fall 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026538.

Anonymous:1959:BMd

[144] Anonymous. Back matter. Dædalus, 88(4):??, Fall 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026539.

Anonymous:1959:FMd

[145] Anonymous. Front matter. Dædalus, 88(4):??, Fall 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026527.

Anonymous:1959:VI

[146] Anonymous. Volume information. Dædalus, 88(4):??, Fall 1959. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026526.

H:1960:F

[147] G. H. Foreword. Dædalus, 89(1):1–2, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026542.

Kepes:1960:IIV

[148] Gyorgy Kepes. Introduction to the issue “The Visual Arts Today”. Dædalus, 89(1):3–12, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026543.

Mead:1960:WLC

[149] Margaret Mead. Work, leisure, and creativity. Dædalus, 89(1):13–23, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026544.

Giedion:1960:RSE

[150] Sigfried Giedion. The roots of symbolic expression. Dædalus, 89(1):24–33, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026545.

Deutsch:1960:BMA

[151] Felix Deutsch. Body, mind, and art. Dædalus, 89(1):34–45, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026546.

Corbusier:1960:AA

[152] Le Corbusier. Architecture and the arts. Dædalus, 89(1):46–51, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026547.

Burchard:1960:AAA

[153] John E. Burchard. Alienated affections in the arts. Dædalus, 89 (1):52–61, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026548.

Belluschi:1960:VAA

[154] Pietro Belluschi, Harry Bertoia, Reg. Butler, Eduardo Chillida, Jimmy Ernst, Walter Gropius, Le Corbusier, Richard Lippold, Walter Netsch, Irene Rice Pereira, and José Luis Sert. Views on art and architecture: a conversation. Dædalus, 89(1):62–73, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026549.

Sekler:1960:CA

[155] Eduard F. Sekler. The city and the arts. Dædalus, 89(1):74–78, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026550.

vanGogh:1960:SDA

[156] Vincent van Gogh, Wassily Kandinsky, Juan Gris, Paul Klee, Fernand Léger, Piet Mondrian, Joan Miró, Jean Dubuffet, José Ortega y Gasset, Albert Camus, Pierre Soulages, Jean Hélion, Josef Albers, Theodore Roszak, Marcel Duchamp, Naum Gabo, Jean Arp, Stuart Davis, Oskar Kokoschka, Ben Shahn, and Saul Steinberg. Statements and documents: Artists on art and reality, on their work, and on values. Dædalus, 89(1):79–126, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026551.

Rand:1960:AAV

[157] Paul Rand and Ann Rand. Advertisement: Ad vivum or ad hominem? Dædalus, 89(1):127–135, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026552.

Steichen:1960:P

[158] Edward Steichen. On photography. Dædalus, 89(1):136–137, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026553.

Kaufman:1960:FMA

[159] Boris Kaufman. Film making as an art. Dædalus, 89(1):138–143, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026554.

Gardner:1960:HD

[160] Robert Gardner. A human document. Dædalus, 89(1):144–149, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026555.

Deren:1960:CCU

[161] Maya Deren. Cinematography: The creative use of reality. Dædalus, 89 (1):150–167, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026556.

Anonymous:1960:NL

[162] Anonymous. The new landscape. Dædalus, 89(1):169–176, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20026557.

Weiss:1960:OFS

[163] Paul Weiss. Organic form: Scientific and aesthetic aspects. Dædalus, 89 (1):177–190, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026558.

Speiser:1960:SSA

[164] Andreas Speiser. Symmetry in science and art. Dædalus, 89(1):191–198, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026559.

Wittkower:1960:CCP

[165] Rudolf Wittkower. The changing concept of proportion. Dædalus, 89 (1):199–215, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026560.

Gibson:1960:PPP

[166] James J. Gibson. Pictures, perspective, and perception. Dædalus, 89 (1):216–227, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026561.

Gombrich:1960:PP

[167] E. H. Gombrich. On physiognomic perception. Dædalus, 89(1):228–241, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026562.

Langer:1960:AS

[168] Suzanne K. Langer. On artistic sensibility. Dædalus, 89(1):242–244, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026563.

Sandberg:1960:PG

[169] W. J. H. B. Sandberg. Picasso’s “Guernica”. Dædalus, 89(1):245–252, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026564.

Ackerman:1960:AHP

[170] James S. Ackerman. Art history and the problems of criticism. Dædalus, 89(1):253–263, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026565.

Anonymous:1960:BMa

[171] Anonymous. Back matter. Dædalus, 89(1):??, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026566.

Anonymous:1960:FMa

[172] Anonymous. Front matter. Dædalus, 89(1):??, Winter 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026541.

H:1960:P

[173] G. H. Preface. Dædalus, 89(2):271–272, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026569.

Jacobs:1960:IIM

[174] Norman Jacobs. Introduction to the issue “Mass Culture and Mass Media”. Dædalus, 89(2):273–277, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026570.

Arendt:1960:SC

[175] Hannah Arendt. Society and culture. Dædalus, 89(2):278–287, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026571.

Shils:1960:MSC

[176] Edward Shils. Mass society and its culture. Dædalus, 89(2):288–314, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026572.

vandenHaag:1960:DCS

[177] Ernest van den Haag. A dissent from the consensual society. Dædalus, 89 (2):315–324, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026573.

Handlin:1960:CMP

[178] Oscar Handlin. Comments on mass and popular culture. Dædalus, 89 (2):325–332, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026574.

Rosten:1960:IMM

[179] Leo Rosten. The intellectual and the mass media: Some rigorously random remarks. Dædalus, 89(2):333–346, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026575.

Stanton:1960:PP

[180] Frank Stanton. Parallel paths. Dædalus, 89(2):347–353, Spring 1960. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20026576.

Sweeney:1960:AMM

[181] James Johnson Sweeney. The artist and the museum in a mass society. Dædalus, 89(2):354–358, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026577.

Jarrell:1960:SHS

[182] Randall Jarrell. A sad heart at the supermarket. Dædalus, 89(2):359–372, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026578.

Baldwin:1960:MCC

[183] James Baldwin. Mass culture and the creative artist: Some personal notes. Dædalus, 89(2):373–376, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026579.

Hyman:1960:IDL

[184] Stanley Edgar Hyman. Ideals, dangers, and limitations of mass culture. Dædalus, 89(2):377–387, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026580.

Hughes:1960:MCS

[185] H. Stuart Hughes. Mass culture and social criticism. Dædalus, 89 (2):388–393, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026581.

Schlesinger:1960:NNC

[186] Arthur Schlesinger, Jr. Notes on a national cultural policy. Dædalus, 89 (2):394–400, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026582.

DeTocqueville:1960:TDA

[187] Alexis-Charles-Henri Clérel De Tocqueville. De Tocqueville on democracy and the arts. Dædalus, 89 (2):401–409, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026583.

Nichols:1960:EF

[188] William I. Nichols. Editing for 13,000,000 families. Dædalus, 89 (2):410–418, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026584.

Eliot:1960:ILU

[189] T. S. Eliot. The influence of landscape upon the poet. Dædalus, 89 (2):419–428, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026585.

Anonymous:1960:BMb

[190] Anonymous. Back matter. Dædalus, 89(2):??, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026586.

Anonymous:1960:FMb

[191] Anonymous. Front matter. Dædalus, 89(2):??, Spring 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026568.

G:1960:PN

[192] S. R. G. Prefatory note. Dædalus, 89(3):437, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026589.

Pipes:1960:FIR

[193] Richard Pipes. Foreword to the issue “The Russian Intelligentsia”. Dædalus, 89(3):438–440, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026590.

Malia:1960:WI

[194] Martin Malia. What is the intelligentsia? Dædalus, 89(3):441–458, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20026591.

Schapiro:1960:PRI

[195] Leonard Schapiro. The pre-revolutionary intelligentsia and the legal order. Dædalus, 89(3):459–471, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026592.

Elkin:1960:RIE

[196] Boris Elkin. The Russian intelligentsia on the eve of the revolution. Dædalus, 89(3):472–486, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026593.

Pipes:1960:HER

[197] Richard Pipes. The historical evolution of the Russian intelligentsia. Dædalus, 89(3):487–502, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026594.

Labedz:1960:SSI

[198] Leopold Labedz. The structure of the Soviet intelligentsia. Dædalus, 89 (3):503–519, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026595.

Burg:1960:OSU

[199] David Burg. Observations on Soviet university students. Dædalus, 89 (3):520–540, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026596.

Haimson:1960:SHP

[200] Leopold H. Haimson. The solitary hero and the philistines. a note on the heritage of the Stalin era. Dædalus, 89 (3):541–550, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026597.

Hayward:1960:TW

[201] Max Hayward. The thaw and the writers. Dædalus, 89(3):551–561, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026598.

Joravsky:1960:SSG

[202] David Joravsky. Soviet scientists and the great break. Dædalus, 89(3): 562–580, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026599.

Wetter:1960:ISS

[203] Gustav Wetter. Ideology and science in the Soviet union recent developments. Dædalus, 89(3):581–603, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026600.

Schwartz:1960:ICC

[204] Benjamin Schwartz. The intelligentsia in Communist China. a tentative comparison. Dædalus, 89(3):604–621, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026601.

Marias:1960:SIS

[205] Julián Marías. The situation of the intelligentsia in Spain today. Dædalus, 89(3):622–631, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026602.

Anonymous:1960:PNS

[206] Anonymous. Philosophy and the natural sciences in the USSR. Dædalus, 89(3):632–647, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026603.

Anonymous:1960:DZL

[207] Anonymous. “Doctor Zhivago”: Letter to Boris Pasternak from the Editors of “Novyi Mir”. Dædalus, 89 (3):648–668, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026604.

Anonymous:1960:BMc

[208] Anonymous. Back matter. Dædalus, 89(3):??, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026605.

Anonymous:1960:FMc

[209] Anonymous. Front matter. Dædalus, 89(3):??, Summer 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026588.

Holton:1960:EPN

[210] Gerald Holton. Editor’s prefatory note. Dædalus, 89(4):674–676, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026609.

Wiesner:1960:FIA

[211] Jerome B. Wiesner. Foreword to the issue “Arms Control”. Dædalus, 89(4):677–680, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026610.

Brennan:1960:SGA

[212] D. G. Brennan. Setting and goals of arms control. Dædalus, 89(4):681–707, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026611.

Bowie:1960:BRA

[213] Robert R. Bowie. Basic requirements of arms control. Dædalus, 89(4):708–722, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026612.

Frye:1960:CRA

[214] William R. Frye. Characteristics of recent arms-control proposals and agreements. Dædalus, 89(4):723–743, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026613.

Kahn:1960:ARS

[215] Herman Kahn. The arms race and some of its hazards. Dædalus, 89(4):744–780, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026614.

Teller:1960:FAC

[216] Edward Teller. The feasibility of arms control and the principle of openness. Dædalus, 89(4):781–799, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026615.

Kissinger:1960:LWC

[217] Henry A. Kissinger. Limited war: Conventional or nuclear? A reappraisal. Dædalus, 89(4):800–817, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026616.

Doty:1960:RSP

[218] Paul Doty. The role of the smaller powers. Dædalus, 89(4):818–830, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026617.

Barnett:1960:ICC

[219] A. Doak Barnett. The inclusion of Communist China in an arms-control program. Dædalus, 89(4):831–845, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026618.

Boulding:1960:DIA

[220] Kenneth E. Boulding. The domestic implications of arms control. Dædalus, 89(4):846–859, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026619.

Feld:1960:ITA

[221] Bernard T. Feld. Inspection techniques of arms control. Dædalus, 89(4):860–878, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026620.

Sohn:1960:AEA

[222] Louis B. Sohn. Adjudication and enforcement in arms control. Dædalus, 89(4):879–891, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026621.

Schelling:1960:RMA

[223] T. C. Schelling. Reciprocal measures for arms stabilization. Dædalus, 89(4):892–914, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026622.

Wiesner:1960:CAL

[224] Jerome B. Wiesner. Comprehensive arms-limitation systems. Dædalus, 89(4):915–950, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026623.

Davis:1960:RPM

[225] Saville R. Davis. Recent policy making in the United States government. Dædalus, 89(4):951–966, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026624.

Humphrey:1960:GOA

[226] Hubert H. Humphrey. Government organization for arms control. Dædalus, 89(4):967–983, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026625.

Pool:1960:POC

[227] Ithiel de Sola Pool. Public opinion and the control of armaments. Dædalus, 89(4):984–999, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026626.

Fox:1960:PDP

[228] William T. R. Fox. Political and diplomatic prerequisites of arms control. Dædalus, 89(4):1000–1014, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026627.

Fromm:1960:CUD

[229] Erich Fromm. The case for unilateral disarmament. Dædalus, 89 (4):1015–1028, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026628.

Brown:1960:TWW

[230] Harrison Brown. Tasks for a world without war. Dædalus, 89(4):1029– 1038, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20026629.

Larson:1960:ACT

[231] Arthur Larson. Arms control through world law. Dædalus, 89(4):1039–1054, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026630.

Wright:1960:SCB

[232] Christopher Wright. Selected critical bibliography. Dædalus, 89(4):1055– 1070, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20026631.

Anonymous:1960:BMd

[233] Anonymous. Back matter. Dædalus, 89(4):??, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026632.

Anonymous:1960:FMd

[234] Anonymous. Front matter. Dædalus, 89(4):??, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026608.

Anonymous:1960:VI

[235] Anonymous. Volume information. Dædalus, 89(4):??, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026607.

G:1961:PNI

[236] S. R. G. Prefatory note to the issue “The Future Metropolis”. Dædalus, 90(1):3, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026635.

Lynch:1961:WC

[237] Kevin Lynch and Lloyd Rodwin. A world of cities. Dædalus, 90(1):4–10, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026636.

Handlin:1961:SS

[238] Oscar Handlin. The social system. Dædalus, 90(1):11–30, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026637.

Vernon:1961:EFL

[239] Raymond Vernon. The economics and finances of the large Metropolis. Dædalus, 90(1):31–47, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026638.

Fleisher:1961:ITU

[240] Aaron Fleisher. The influence of technology on urban forms. Dædalus, 90(1):48–60, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026639.

Banfield:1961:PIM

[241] Edward C. Banfield. The political implications of metropolitan growth. Dædalus, 90(1):61–78, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026640.

Lynch:1961:PM

[242] Kevin Lynch. The pattern of the Metropolis. Dædalus, 90(1):79–98, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026641.

Deutsch:1961:SCM

[243] Karl W. Deutsch. On social communication and the Metropolis. Dædalus, 90(1):99–110, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026642.

Dyckman:1961:CUC

[244] John Dyckman. The changing uses of the city. Dædalus, 90(1):111–131, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026643.

Rodwin:1961:MPD

[245] Lloyd Rodwin. Metropolitan policy for developing areas. Dædalus, 90 (1):132–146, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026644.

Kepes:1961:NEC

[246] Gyorgy Kepes. Notes on expression and communication in the cityscape. Dædalus, 90(1):147–165, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026645.

White:1961:AIV

[247] Morton White and Lucia White. The American intellectual versus the American city. Dædalus, 90(1):166–179, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026646.

Meyerson:1961:UTP

[248] Martin Meyerson. Utopian traditions and the planning of cities. Dædalus, 90(1):180–193, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026647.

Noel-Baker:1961:CIA

[249] Philip Noel-Baker, John Boyd Orr, Alastair Buchan, Raymond Aron, and Jules Moch. Comments on the issue “Arms Control”. Dædalus, 90 (1):194–213, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026648.

Anonymous:1961:BMa

[250] Anonymous. Back matter. Dædalus, 90(1):??, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026649.

Anonymous:1961:FMa

[251] Anonymous. Front matter. Dædalus, 90(1):??, Winter 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026634.

Handlin:1961:PIE

[252] Oscar Handlin. Preface to the issue “Ethnic Groups in American Life”. Dædalus, 90(2):219, Spring 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026652.

Handlin:1961:HPA

[253] Oscar Handlin. Historical perspectives on the American ethnic group. Dædalus, 90(2):220–232, Spring 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026653.

Benne:1961:UF

[254] Kenneth D. Benne. The uses of fraternity. Dædalus, 90(2):233–246, Spring 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026654.

Yinger:1961:SFI

[255] J. Milton Yinger. Social forces involved in group identification or withdrawal. Dædalus, 90(2):247–262, Spring 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026655.

Gordon:1961:AAT

[256] Milton M. Gordon. Assimilation in America: Theory and reality. Dædalus, 90(2):263–285, Spring 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026656.

Simmons:1961:MIE

[257] Ozzie G. Simmons. The mutual images and expectations of Anglo-Americans and Mexican-Americans. Dædalus, 90 (2):286–299, Spring 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026657.

Rossi:1961:SEP

[258] Peter H. Rossi and Alice S. Rossi. Some effects of parochial school education in America. Dædalus, 90 (2):300–328, Spring 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026658.

Fishman:1961:CIM

[259] Joshua A. Fishman. Childhood indoctrination for minority-group membership. Dædalus, 90(2):329–349, Spring 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026659.

Aron:1961:SDW

[260] Raymond Aron. The situation of democracy: Western political institutions in the Twentieth Century. Dædalus, 90(2):350–370, Spring 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026660.

Leach:1961:R

[261] Edmund R. Leach and Herbert Weisinger. Reputations. Dædalus, 90 (2):371–399, Spring 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

                                        /www.jstor.org/stable/20026661.

Anonymous:1961:BMb

[262] Anonymous. Back matter. Dædalus, 90(2):??, Spring 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026662.

Anonymous:1961:FMb

[263] Anonymous. Front matter. Dædalus, 90(2):??, Spring 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026651.

Hoagland:1961:IIE

[264] Hudson Hoagland and Ralph W. Burhoe. Introduction to the issue “Evolution and Man’s Progress”. Dædalus, 90(3):411–415, Summer 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026665.

Crow:1961:MTH

[265] James F. Crow. Mechanisms and trends in human evolution. Dædalus, 90(3):416–431, Summer 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026666.

Muller:1961:SWW

[266] Hermann J. Muller. Should we weaken or strengthen our genetic heritage? Dædalus, 90(3):432–450, Summer 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026667.

Morison:1961:CGE

[267] Robert S. Morison, Garrett Hardin, J. Paul Scott, Lawrence K. Frank, Ernst Mayr, Theodosius Dobzhansky, James F. Crow, Hermann J. Muller, I. Michael Lerner, Ralph W. Gerard, George G. Simpson, and Donald H. Fleming. Comments on genetic evolution. Dædalus, 90(3):451–476, Summer 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026668.

Steward:1961:SMS

[268] Julian H. Steward and Demitri B. Shimkin. Some mechanisms of sociocultural evolution. Dædalus, 90 (3):477–497, Summer 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026669.

Rosenblith:1961:SSC

[269] Walter A. Rosenblith. On some social consequences of scientific and technological change. Dædalus, 90(3): 498–513, Summer 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026670.

Simpson:1961:CCE

[270] George G. Simpson, Ralph W. Gerard, Ward H. Goodenough, and Alex Inkeles. Comments on cultural evolution. Dædalus, 90(3):514–533, Summer 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026671.

Skinner:1961:DC

[271] B. F. Skinner. The design of cultures. Dædalus, 90(3):534–546, Summer 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026672.

Murray:1961:UE

[272] Henry A. Murray. Unprecedented evolutions. Dædalus, 90(3):547–569, Summer 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026673.

Murray:1961:CEV

[273] Henry A. Murray, B. F. Skinner, Abraham H. Maslow, Carl R. Rogers, Lawrence K. Frank, Anatol Rapoport, and Hallock Hoffman. Cultural evolution as viewed by psychologists. Dædalus, 90(3):570–586, Summer 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026674.

Frye:1961:MFD

[274] Northrop Frye. Myth, fiction, and displacement. Dædalus, 90(3):587–605, Summer 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20026675.

Anonymous:1961:G

[275] Anonymous. Glossary. Dædalus, 90 (3):608–609, Summer 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026676.

Anonymous:1961:BMc

[276] Anonymous. Back matter. Dædalus, 90(3):??, Summer 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026677.

Anonymous:1961:FMc

[277] Anonymous. Front matter. Dædalus, 90(3):??, Summer 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026664.

G:1961:PIE

[278] S. R. G. Preface to the issue “Excellence and Leadership in a Democracy”. Dædalus, 90(4):619–627, Fall 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026681.

Peyre:1961:ELW

[279] Henri Peyre. Excellence and leadership: Has Western Europe any lessons for us? Dædalus, 90(4):628–651, Fall 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026682.

Commager:1961:LEC

[280] Henry Steele Commager. Leadership in Eighteenth-Century America and today. Dædalus, 90(4):652–673, Fall 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026683.

Conway:1961:SE

[281] John Conway. Standards of excellence. Dædalus, 90(4):674–692, Fall 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026684.

Abse:1961:PAL

[282] D. Wilfred Abse and Lucie Jessner. The psychodynamic aspects of leadership. Dædalus, 90(4):693–710, Fall 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026685.

McClelland:1961:EE

[283] David C. McClelland. Encouraging excellence. Dædalus, 90(4):711–724, Fall 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026686.

Yarmolinsky:1961:ERE

[284] Adam Yarmolinsky. The explicit recognition of excellence: a survey. Dædalus, 90(4):725–733, Fall 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026687.

Burns:1961:ELP

[285] James MacGregor Burns. Excellence and leadership in President and Congress. Dædalus, 90(4):734–749, Fall 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026688.

Price:1961:AL

[286] Don K. Price. Administrative leadership. Dædalus, 90(4):750–763, Fall 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026689.

Morison:1961:NNT

[287] Robert S. Morison. The need for new types of excellence. Dædalus, 90(4):764–781, Fall 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026690.

Anonymous:1961:BMd

[288] Anonymous. Back matter. Dædalus, 90(4):??, Fall 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026691.

Anonymous:1961:FMd

[289] Anonymous. Front matter. Dædalus, 90(4):??, Fall 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026680.

Anonymous:1961:VI

[290] Anonymous. Volume information. Dædalus, 90(4):??, Fall 1961. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026679.

G:1962:PIY

[291] S. R. G. Preface to the issue “Youth: Change and Challenge”. Dædalus, 91(1):3–4, Winter 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026694.

Erikson:1962:YFD

[292] Erik H. Erikson. Youth: Fidelity and diversity. Dædalus, 91(1):5–27, Winter 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026695.

Eisenstadt:1962:APY

[293] S. N. Eisenstadt. Archetypal patterns of youth. Dædalus, 91(1):28–46, Winter 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026696.

Naegele:1962:YSS

[294] Kaspar D. Naegele. Youth and society: Some observations. Dædalus, 91(1):47–67, Winter 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026697.

Bettelheim:1962:PG

[295] Bruno Bettelheim. The problem of generations. Dædalus, 91(1):68–96, Winter 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026698.

Parsons:1962:YCA

[296] Talcott Parsons. Youth in the context of American society. Dædalus, 91(1):97–123, Winter 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026699.

Denney:1962:AYT

[297] Reuel Denney. American youth today: a bigger cast, a wider screen. Dædalus, 91(1):124–144, Winter 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026700.

Keniston:1962:SCY

[298] Kenneth Keniston. Social change and youth in America. Dædalus, 91 (1):145–171, Winter 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026701.

Lifton:1962:YHI

[299] Robert Jay Lifton. Youth and history: Individual change in postwar Japan. Dædalus, 91(1):172–197, Winter 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026702.

Wylie:1962:YFU

[300] Laurence Wylie. Youth in France and the United States. Dædalus, 91 (1):198–215, Winter 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026703.

Sherman:1962:SYM

[301] George Sherman. Soviet youth: Myth and reality. Dædalus, 91(1):216–237, Winter 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026704.

Anonymous:1962:BMa

[302] Anonymous. Back matter. Dædalus, 91(1):??, Winter 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026705.

Anonymous:1962:FMa

[303] Anonymous. Front matter. Dædalus, 91(1):??, Winter 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026693.

G:1962:PIS

[304] S. R. G. Preface to the issue “Science and Technology in Contemporary Society”. Dædalus, 91(2):247–248, Spring 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026708.

Aron:1962:ECI

[305] Raymond Aron. The education of the citizen in industrial society. Dædalus, 91(2):249–263, Spring 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026709.

Ashby:1962:ABP

[306] Eric Ashby. The administrator: Bottleneck or pump? Dædalus, 91 (2):264–278, Spring 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026710.

Huxley:1962:ENL

[307] Aldous Huxley. Education on the nonverbal level. Dædalus, 91(2):279–293, Spring 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026711.

Kubie:1962:FCS

[308] Lawrence S. Kubie. The fostering of creative scientific productivity. Dædalus, 91(2):294–309, Spring 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026712.

Lewis:1962:ESP

[309] W. Arthur Lewis. Education for scientific professions in the poor countries. Dædalus, 91(2):310–318, Spring 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026713.

Morison:1962:UTA

[310] Robert S. Morison. The University and technical assistance. Dædalus, 91 (2):319–340, Spring 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026714.

Brown:1962:NHE

[311] Gordon S. Brown. New horizons in engineering education. Dædalus, 91 (2):341–361, Spring 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026715.

Holton:1962:SRS

[312] Gerald Holton. Scientific research and scholarship notes toward the design of proper scales. Dædalus, 91 (2):362–399, Spring 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026716.

Manuel:1962:TSP

[313] Frank E. Manuel. Two styles of philosophical history. Dædalus, 91 (2):400–417, Spring 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026717.

Ong:1962:RSR

[314] Walter J. Ong. Religion, scholarship, and the resituation of man. Dædalus, 91(2):418–436, Spring 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026718.

Anonymous:1962:BMb

[315] Anonymous. Back matter. Dædalus, 91(2):??, Spring 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026719.

Anonymous:1962:FMb

[316] Anonymous. Front matter. Dædalus, 91(2):??, Spring 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026707.

G:1962:PIC

[317] S. R. G. Preface to the issue “Current Work and Controversies”—2. Dædalus, 91(3):455–456, Summer 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026722.

White:1962:RAI

[318] Morton White. Reflections on antiintellectualism. Dædalus, 91(3): 457–468, Summer 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

/www.jstor.org/stable/20026723.

Slive:1962:RSS

[319] Seymour Slive. Realism and symbolism in Seventeenth-Century Dutch painting. Dædalus, 91(3):469–500, Summer 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026724.

Hofstadter:1962:CW

[320] Richard Hofstadter. The child and the world. Dædalus, 91(3):501–526, Summer 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026725.

Travis:1962:ITT

[321] William P. Travis. International trade theory and the trade expansion act of 1962. Dædalus, 91(3):527–542, Summer 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026726.

Friedrich:1962:LPI

[322] Paul Friedrich. Language and politics in India. Dædalus, 91(3):543–559, Summer 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026727.

Hawkins:1962:DM

[323] David Hawkins. Design for a mind. Dædalus, 91(3):560–577, Summer 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026728.

Dray:1962:SCA

[324] William Dray and Newton Garver. Some causal accounts of the American Civil War. Dædalus, 91(3): 578–598, Summer 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026729.

Grant:1962:HLM

[325] Edward Grant and Benjamin Nelson. Hypotheses in late medieval and early modern science. Dædalus, 91 (3):599–616, Summer 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026730.

Kursanov:1962:PFH

[326] G. A. Kursanov and Alexander G. Korol. Philipp Frank and his philosophy of science. Dædalus, 91(3): 617–641, Summer 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026731.

Anonymous:1962:BMc

[327] Anonymous. Back matter. Dædalus, 91(3):??, Summer 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026732.

Anonymous:1962:FMc

[328] Anonymous. Front matter. Dædalus, 91(3):??, Summer 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026721.

G:1962:PIA

[329] S. R. G. Preface to the issue “American Foreign Policy: Freedoms and Restraints”. Dædalus, 91(4):651–652, Fall 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026736.

May:1962:NFP

[330] Ernest R. May. The nature of foreign policy: The calculated versus the axiomatic. Dædalus, 91(4):653–667, Fall 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026737.

Hoffmann:1962:RCA

[331] Stanley Hoffmann. Restraints and choices in American foreign policy. Dædalus, 91(4):668–704, Fall 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026738.

Smith:1962:RAF

[332] Roger Smith. Restraints on American foreign policy. Dædalus, 91(4):705–716, Fall 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026739.

Aron:1962:RAD

[333] Raymond Aron. Reflections on American diplomacy. Dædalus, 91(4):717– 732, Fall 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20026740.

Brodie:1962:DPP

[334] Bernard Brodie. Defense policy and the possibility of total war. Dædalus, 91(4):733–748, Fall 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026741.

Bloomfield:1962:UNC

[335] Lincoln P. Bloomfield. The United Nations in crisis: The role of the United Nations in United States foreign policy. Dædalus, 91(4):749–765, Fall 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026742.

Stender:1962:PSP

[336] Peter Paul Stender. The paradox of Soviet power. Dædalus, 91(4):766–782, Fall 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026743.

Feis:1962:USC

[337] Herbert Feis. The United States and China. Dædalus, 91(4):783–799, Fall 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026744.

Plank:1962:APP

[338] John N. Plank. The alliance for progress: Problems and prospects. Dædalus, 91(4):800–811, Fall 1962. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20026745.

Caves:1962:LRA

[339] Richard E. Caves. The liberal revolt against liberal trade policy. Dædalus, 91(4):812–819, Fall 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026746.

Anonymous:1962:BMd

[340] Anonymous. Back matter. Dædalus, 91(4):??, Fall 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026747.

Anonymous:1962:FMd

[341] Anonymous. Front matter. Dædalus, 91(4):??, Fall 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026735.

Anonymous:1962:VI

[342] Anonymous. Volume information. Dædalus, 91(4):??, Fall 1962. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026734.

G:1963:PIA

[343] S. R. G. Preface to the issue “The American Reading Public”. Dædalus, 92(1):3–4, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026750.

Brower:1963:BRR

[344] Reuben A. Brower. Book reading and the reading of books. Dædalus, 92(1):5–20, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026751.

DeMott:1963:SSN

[345] Benjamin De Mott. Statement and struggle: a note on teaching against the environment. Dædalus, 92(1):21–29, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026752.

Best:1963:BTC

[346] Marshall A. Best. In books, they call it revolution. Dædalus, 92(1):30–41, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026753.

Lacy:1963:EPA

[347] Dan Lacy. The economics of publishing, or Adam Smith and literature. Dædalus, 92(1):42–62, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026754.

Epstein:1963:CCP

[348] Jason Epstein. A criticism of commercial publishing. Dædalus, 92(1):63–67, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026755.

Shugg:1963:PTP

[349] Roger W. Shugg. The professors and their publishers. Dædalus, 92 (1):68–77, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026756.

Wilson:1963:ABP

[350] Thomas J. Wilson. American book publishing: Hazards and opportunities. Dædalus, 92(1):78–91, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026757.

Shils:1963:BA

[351] Edward Shils. The bookshop in America. Dædalus, 92(1):92–104, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026758.

Shatzkin:1963:BSR

[352] Leonard Shatzkin. The book in search of a reader. Dædalus, 92(1):105–115, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026759.

Bogart:1963:NAT

[353] Leo Bogart. Newspapers in the age of television. Dædalus, 92(1):116–127, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026760.

Peyre:1963:WWA

[354] Henri Peyre. What is wrong with American book-reviewing? Dædalus, 92(1):128–144, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026761.

Hollander:1963:SAC

[355] John Hollander. Some animadversions on current reviewing. Dædalus, 92 (1):145–154, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026762.

Anonymous:1963:CJH

[356] Anonymous. Catch-22 by Joseph Heller. Dædalus, 92(1):155–165, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026763.

Anonymous:1963:DVF

[357] Anonymous. La Dolce Vita by Federico Fellini; Oscar Deliso; Bernard Shir-Cliff. Dædalus, 92(1):165–167, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026764.

Anonymous:1963:LDA

[358] Anonymous. Love and death in the American novel by Leslie Fiedler. Dædalus, 92(1):167–172, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026765.

Anonymous:1963:RAH

[359] Anonymous. The reconstruction of American history by John Higham. Dædalus, 92(1):172–174, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20026766.

Anonymous:1963:RPW

[360] Anonymous. Russians as people by Wright Miller; East minus West = zero by Werner Keller. Dædalus, 92 (1):174–178, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026767.

Anonymous:1963:BML

[361] Anonymous. Young Man Luther. A Study in Psychoanalysis and History by Erik H. Erikson; Thought Reform and the Psychology of Totalism. A Study of “Brainwashing” in China by Robert Jay Lifton. Dædalus, 92 (1):178–187, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026768.

Anonymous:1963:BMa

[362] Anonymous. Back matter. Dædalus, 92(1):??, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026769.

Anonymous:1963:FMa

[363] Anonymous. Front matter. Dædalus, 92(1):??, Winter 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026749.

Guerard:1963:IIP

[364] Albert J. Guerard. Introduction to the issue “Perspectives on the Novel”. Dædalus, 92(2):197–205, Spring 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026772.

Levin:1963:AAN

[365] Harry Levin. Apogee and aftermath of the novel. Dædalus, 92(2):206–219, Spring 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026773.

Kiely:1963:CDT

[366] Robert Kiely. The craft of despondency: The traditional novelists. Dædalus, 92(2):220–237, Spring 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026774.

Stevenson:1963:A

[367] David L. Stevenson. The activists. Dædalus, 92(2):238–249, Spring 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026775.

Littlejohn:1963:AR

[368] David Littlejohn. The anti-realists. Dædalus, 92(2):250–264, Spring 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20026776.

Brooks:1963:LN

[369] Peter Brooks. In the laboratory of the novel. Dædalus, 92(2):265–280, Spring 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026777.

Lytle:1963:IEF

[370] Andrew Lytle. Impressionism, the ego, and the first person. Dædalus, 92 (2):281–296, Spring 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026778.

Ghiselin:1963:AIA

[371] Brewster Ghiselin. Automatism, intention, and autonomy in the Novelist’s production. Dædalus, 92(2):297–311, Spring 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026779.

Moser:1963:TGS

[372] Thomas Moser. Towards “The Good Soldier”: Discovery of a sexual theme. Dædalus, 92(2):312–325, Spring 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026780.

Rosenfield:1963:SWC

[373] Claire Rosenfield. The shadow within: The conscious and unconscious use of the double. Dædalus, 92(2):326–344, Spring 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026781.

Kohlberg:1963:PAL

[374] Lawrence Kohlberg. Psychological analysis and literary form: a study of the doubles in Dostoevsky. Dædalus, 92(2):345–362, Spring 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026782.

Lesser:1963:RUU

[375] Simon O. Lesser. The role of unconscious understanding in Flaubert and Dostoevsky. Dædalus, 92(2):363–382, Spring 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026783.

Anonymous:1963:BMb

[376] Anonymous. Back matter. Dædalus, 92(2):??, Spring 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026784.

Anonymous:1963:FMb

[377] Anonymous. Front matter. Dædalus, 92(2):??, Spring 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026771.

G:1963:PIT

[378] S. R. G. Preface to the issue “Themes in Transition”. Dædalus, 92 (3):393–395, Summer 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026787.

Letwin:1963:FFA

[379] William Letwin. Four fallacies about economic development. Dædalus, 92 (3):396–414, Summer 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026788.

Taylor:1963:FEI

[380] Overton Hume Taylor. The “Free Enterprise” ideology and American ideals and institutions. Dædalus, 92 (3):415–432, Summer 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026789.

Gordon:1963:IEJ

[381] H. Scott Gordon. Ideas of economic justice. Dædalus, 92(3):433–446, Summer 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026790.

Mazlish:1963:IP

[382] Bruce Mazlish. The idea of progress. Dædalus, 92(3):447–461, Summer 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026791.

Lifton:1963:PEA

[383] Robert Jay Lifton. Psychological effects of the atomic Bomb in Hiroshima: The theme of death. Dædalus, 92 (3):462–497, Summer 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026792.

Etzioni:1963:EUP

[384] Amitai Etzioni. European unification and perspectives on sovereignty. Dædalus, 92(3):498–520, Summer 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026793.

Stendahl:1963:RU

[385] Krister Stendahl. Religion in the University. Dædalus, 92(3):521–528, Summer 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026794.

Kelly:1963:EHA

[386] George A. Kelly. The expert as historical actor. Dædalus, 92(3):529–548, Summer 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20026795.

Goldstein:1963:WID

[387] Joseph Goldstein and Jay Katz. Why an “Insanity Defense”. Dædalus, 92 (3):549–563, Summer 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026796.

Barnett:1963:I

[388] S. A. Barnett. Instinct. Dædalus, 92 (3):564–580, Summer 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026797.

Mongan:1963:ELD

[389] Agnes Mongan. European landscape drawing 1400–1900: a brief survey. Dædalus, 92(3):581–635, Summer 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026798.

Anonymous:1963:BMc

[390] Anonymous. Back matter. Dædalus, 92(3):??, Summer 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026799.

Anonymous:1963:FMc

[391] Anonymous. Front matter. Dædalus, 92(3):??, Summer 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026786.

G:1963:PIP

[392] S. R. G. Preface to the issue “The Professions”. Dædalus, 92(4):647–648, Fall 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026803.

Lynn:1963:IIP

[393] Kenneth S. Lynn. Introduction to the issue “The Professions”. Dædalus, 92(4):649–654, Fall 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026804.

Hughes:1963:P

[394] Everett C. Hughes. Professions. Dædalus, 92(4):655–668, Fall 1963. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20026805.

Barber:1963:SPS

[395] Bernard Barber. Some problems in the sociology of the professions. Dædalus, 92(4):669–688, Fall 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026806.

Freund:1963:LP

[396] Paul A. Freund. The legal profession. Dædalus, 92(4):689–700, Fall 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026807.

Means:1963:HMA

[397] James Howard Means. Homo Medicus Americanus. Dædalus, 92(4):701–723, Fall 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026808.

Gustafson:1963:CUS

[398] James M. Gustafson. The clergy in the United States. Dædalus, 92(4):724– 744, Fall 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20026809.

Wittlin:1963:T

[399] Alma S. Wittlin. The teacher. Dædalus, 92(4):745–763, Fall 1963. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20026810.

Beer:1963:APS

[400] John J. Beer and W. David Lewis. Aspects of the professionalization of science. Dædalus, 92(4):764–784, Fall 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026811.

Huntington:1963:PEM

[401] Samuel P. Huntington. Power, expertise and the military profession. Dædalus, 92(4):785–807, Fall 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026812.

Zinberg:1963:PPD

[402] Norman E. Zinberg. Psychiatry: a professional dilemma. Dædalus, 92 (4):808–823, Fall 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026813.

Alonso:1963:CCP

[403] William Alonso. Cities and city planners. Dædalus, 92(4):824–839, Fall 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026814.

Conway:1963:PPU

[404] John Conway. Politics as a profession in the United States. Dædalus, 92(4):840–856, Fall 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026815.

Anonymous:1963:BMd

[405] Anonymous. Back matter. Dædalus, 92(4):??, Fall 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026816.

Anonymous:1963:FMd

[406] Anonymous. Front matter. Dædalus, 92(4):??, Fall 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026802.

Anonymous:1963:VI

[407] Anonymous. Volume information. Dædalus, 92(4):??, Fall 1963. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026801.

G:1964:PIN

[408] S. R. G. Preface to the issue “A New Europe?”. Dædalus, 93(1):3–5, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026819.

Albonetti:1964:NEW

[409] Achille Albonetti. The new Europe and the West. Dædalus, 93(1):6–42, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026820.

Aron:1964:ONN

[410] Raymond Aron. Old nations, new Europe. Dædalus, 93(1):43–66, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026821.

Franks:1964:BE

[411] Oliver Franks. Britain and Europe. Dædalus, 93(1):67–82, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026822.

Kohnstamm:1964:ET

[412] Max Kohnstamm. The European tide. Dædalus, 93(1):83–108, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026823.

Mayne:1964:EIN

[413] Richard Mayne. Economic integration in the new Europe: A statistical approach. Dædalus, 93(1):109–133, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026824.

Clark:1964:NEN

[414] William Clark. New Europe and the new nations. Dædalus, 93(1):134–152, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026825.

Grosser:1964:EEP

[415] Alfred Grosser. The evolution of European parliaments. Dædalus, 93 (1):153–178, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026826.

Bracher:1964:PPD

[416] Karl Dietrich Bracher. Problems of parliamentary democracy in Europe. Dædalus, 93(1):179–198, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026827.

Pizzorno:1964:IME

[417] Alessandro Pizzorno. The individualistic mobilization of Europe. Dædalus, 93(1):199–224, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026828.

Dahrendorf:1964:RCC

[418] Ralf Dahrendorf. Recent changes in the class structure of European societies. Dædalus, 93(1):225–270, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026829.

Lipset:1964:CCS

[419] Seymour Martin Lipset. The changing class structure and contemporary European politics. Dædalus, 93 (1):271–303, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026830.

Touraine:1964:MWC

[420] Alain Touraine. Management and the working class in Western Europe. Dædalus, 93(1):304–334, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026831.

Rossi-Doria:1964:AE

[421] Manlio Rossi-Doria. Agriculture and Europe. Dædalus, 93(1):335–357, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026832.

Rogers:1964:PEA

[422] Ernesto N. Rogers. The phenomenology of European architecture. Dædalus, 93(1):358–372, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026833.

Bowles:1964:ENE

[423] Frank Bowles. Education in the new Europe. Dædalus, 93(1):373–393, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026834.

Cavazza:1964:ESS

[424] Fabio Luca Cavazza. The European school system: Problems and trends. Dædalus, 93(1):394–415, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026835.

Hessel:1964:EFE

[425] Stephane Hessel. Is Europe facing its educational problems? Dædalus, 93 (1):416–433, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026836.

King:1964:STN

[426] Alexander King. Science and technology in the new Europe. Dædalus, 93 (1):434–458, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026837.

Larrabee:1964:TTD

[427] Eric Larrabee. Transcripts of a transatlantic dialogue. Dædalus, 93 (1):459–470, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026838.

Frankel:1964:BDN

[428] Charles Frankel. Bureaucracy and democracy in the new Europe. Dædalus, 93(1):471–492, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20026839.

Weil:1964:PPT

[429] Eric Weil. Philosophical and political thought in Europe today. Dædalus, 93(1):493–513, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026840.

Crozier:1964:CRN

[430] Michel Crozier. The cultural revolution: Notes on the changes in the intellectual climate of France. Dædalus, 93(1):514–542, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026841.

Graubard:1964:NE

[431] Stephen R. Graubard. A new Europe? Dædalus, 93(1):543–566, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026842.

Anonymous:1964:BMa

[432] Anonymous. Back matter. Dædalus, 93(1):??, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026843.

Anonymous:1964:FMa

[433] Anonymous. Front matter. Dædalus, 93(1):??, Winter 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026818.

G:1964:PIW

[434] S. R. G. Preface to the issue “The Woman in America”. Dædalus, 93 (2):579–581, Spring 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026846.

Erikson:1964:IOS

[435] Erik H. Erikson. Inner and outer space: Reflections on womanhood. Dædalus, 93(2):582–606, Spring 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026847.

Rossi:1964:EBS

[436] Alice S. Rossi. Equality between the sexes: an immodest proposal. Dædalus, 93(2):607–652, Spring 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026848.

Degler:1964:RIC

[437] Carl N. Degler. Revolution without ideology: The changing place of women in America. Dædalus, 93 (2):653–670, Spring 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026849.

Peterson:1964:WW

[438] Esther Peterson. Working women. Dædalus, 93(2):671–699, Spring 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20026850.

Bailyn:1964:NRC

[439] Lotte Bailyn. Notes on the role of choice in the psychology of professional women. Dædalus, 93(2):700–710, Spring 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026851.

Riesman:1964:TG

[440] David Riesman. Two generations. Dædalus, 93(2):711–735, Spring 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026852.

Rostow:1964:CA

[441] Edna G. Rostow. Conflict and accommodation. Dædalus, 93(2):736–760, Spring 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026853.

Conway:1964:JAA

[442] Jill Conway. Jane Addams: an American heroine. Dædalus, 93(2):761–780, Spring 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026854.

Erikson:1964:NFN

[443] Joan M. Erikson. Nothing to fear: Notes on the life of Eleanor Roosevelt. Dædalus, 93(2):781–801, Spring 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026855.

Anonymous:1964:BMb

[444] Anonymous. Back matter. Dædalus, 93(2):??, Spring 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026856.

Anonymous:1964:FMb

[445] Anonymous. Front matter. Dædalus, 93(2):??, Spring 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026845.

G:1964:PIP

[446] S. R. G. Preface to the issue “Population, Prediction, Conflict, Existentialism”. Dædalus, 93(3):811, Summer 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026859.

Hoagland:1964:MPC

[447] Hudson Hoagland. Mechanisms of population control. Dædalus, 93(3): 812–829, Summer 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026860.

Mayer:1964:FPW

[448] Jean Mayer. Food and population: The wrong problem? Dædalus, 93 (3):830–844, Summer 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026861.

Bell:1964:TMP

[449] Daniel Bell. Twelve modes of prediction: a preliminary sorting of approaches in the social sciences. Dædalus, 93(3):845–880, Summer 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026862.

Waltz:1964:SBW

[450] Kenneth N. Waltz. The stability of a bipolar world. Dædalus, 93 (3):881–909, Summer 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026863.

Fisher:1964:PWT

[451] Roger Fisher. Perceiving the world through bipolar glasses. Dædalus, 93 (3):910–915, Summer 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026864.

Griffith:1964:ESM

[452] William E. Griffith. Eppuor Si Muove. Dædalus, 93(3):916–919, Summer 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026865.

Fisher:1964:FC

[453] Roger Fisher. Fractionating conflict. Dædalus, 93(3):920–941, Summer 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026866.

Finkelstein:1964:CFC

[454] Lawrence S. Finkelstein. Comments on “Fractionating Conflict”. Dædalus, 93 (3):942–945, Summer 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026867.

Cavell:1964:EAP

[455] Stanley Cavell. Existentialism and analytical philosophy. Dædalus, 93 (3):946–974, Summer 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026868.

Politzer:1964:TAC

[456] Heinz Politzer. The trial against the court. Dædalus, 93(3):975–997, Summer 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026869.

Fischer:1964:ENZ

[457] Kurt Rudolf Fischer. The existentialism of Nietzsche’s Zarathustra. Dædalus, 93(3):998–1016, Summer 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026870.

Anonymous:1964:BMc

[458] Anonymous. Back matter. Dædalus, 93(3):??, Summer 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026871.

Anonymous:1964:FMc

[459] Anonymous. Front matter. Dædalus, 93(3):??, Summer 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026858.

G:1964:PIC

[460] S. R. G. Preface to the issue “The Contemporary University:

U.S.A.”. Dædalus, 93(4):1027–1032, Fall 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026875.

Devane:1964:CLA

[461] William C. Devane. The College of liberal arts. Dædalus, 93(4):1033–1050, Fall 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026876.

Kerr:1964:FRR

[462] Clark Kerr. The frantic race to remain contemporary. Dædalus, 93 (4):1051–1070, Fall 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026877.

Wallis:1964:CCF

[463] W. Allen Wallis. Centripetal and centrifugal forces in University organization. Dædalus, 93(4):1071–1082, Fall 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026878.

Heimberger:1964:SU

[464] Frederic Heimberger. The state universities. Dædalus, 93(4):1083–1108, Fall 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026879.

Morison:1964:FU

[465] Robert S. Morison. Foundations and universities. Dædalus, 93(4):1109– 1141, Fall 1964. CODEN DAEDAU.

ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20026880.

Rossi:1964:RSP

[466] Peter H. Rossi. Researchers, scholars and policy makers: The politics of large scale research. Dædalus, 93 (4):1142–1161, Fall 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026881.

Ziegler:1964:CEU

[467] Jerome M. Ziegler. Continuing education in the University. Dædalus, 93(4):1162–1183, Fall 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026882.

Weiss:1964:SU

[468] Paul Weiss. Science in the University. Dædalus, 93(4):1184–1218, Fall 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026883.

Bush:1964:H

[469] Douglas Bush. The humanities. Dædalus, 93(4):1219–1237, Fall 1964. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20026884.

Stratton:1964:CAM

[470] Julius A. Stratton. Commencement address, Massachusetts Institute of Technology, June 12, 1964. Dædalus, 93 (4):1238–1243, Fall 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026885.

Hoffmann:1964:EIC

[471] Stanley Hoffmann. Europe’s identity crisis: Between the past and America. Dædalus, 93(4):1244–1297, Fall 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026886.

Anonymous:1964:BMd

[472] Anonymous. Back matter. Dædalus, 93(4):??, Fall 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026887.

Anonymous:1964:FMd

[473] Anonymous. Front matter. Dædalus, 93(4):??, Fall 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026874.

Anonymous:1964:VI

[474] Anonymous. Volume information. Dædalus, 93(4):??, Fall 1964. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026873.

G:1965:P

[475] S. R. G. Preface. Dædalus, 94(1):iii–iv, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026890.

Holton:1965:IIS

[476] Gerald Holton. Introduction to the issue “Science and Culture”. Dædalus, 94(1):v–xxix, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026891.

Levin:1965:SC

[477] Harry Levin. Semantics of culture. Dædalus, 94(1):1–13, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026892.

Ackerman:1965:S

[478] James S. Ackerman. On scientia. Dædalus, 94(1):14–23, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026893.

Leach:1965:CSC

[479] Edmund R. Leach. Culture and social cohesion: an Anthropologist’s view. Dædalus, 94(1):24–38, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026894.

Parsons:1965:UDM

[480] Talcott Parsons. Unity and diversity in the modern intellectual disciplines: The role of the social sciences. Dædalus, 94(1):39–65, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026895.

Brooks:1965:SCC

[481] Harvey Brooks. Scientific concepts and cultural change. Dædalus, 94 (1):66–83, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026896.

Price:1965:ED

[482] Don K. Price. The established dissenters. Dædalus, 94(1):84–116, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026897.

Kepes:1965:VAS

[483] Gyorgy Kepes. The visual arts and sciences: a proposal for collaboration. Dædalus, 94(1):117–134, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026898.

Mead:1965:FBE

[484] Margaret Mead. The future as the basis for establishing a shared culture. Dædalus, 94(1):135–155, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026899.

Handlin:1965:STP

[485] Oscar Handlin. Science and technology in popular culture. Dædalus, 94 (1):156–170, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026900.

Weil:1965:SMC

[486] Eric Weil. Science in modern culture, or the meaning of meaninglessness. Dædalus, 94(1):171–189, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026901.

Marcuse:1965:RRC

[487] Herbert Marcuse. Remarks on a redefinition of culture. Dædalus, 94 (1):190–207, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026902.

Bell:1965:DCS

[488] Daniel Bell. The disjunction of culture and social structure: Some notes on the meaning of social reality. Dædalus, 94(1):208–222, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026903.

Dubos:1965:SMN

[489] René Dubos. Science and Man’s nature. Dædalus, 94(1):223–244, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026904.

Morison:1965:TCS

[490] Robert S. Morison. Toward a common scale of measurement. Dædalus, 94(1):245–262, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026905.

Anonymous:1965:BMa

[491] Anonymous. Back matter. Dædalus, 94(1):??, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026906.

Anonymous:1965:FMa

[492] Anonymous. Front matter. Dædalus, 94(1):??, Winter 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026889.

G:1965:PIU

[493] S. R. G. Preface to the issue “Utopia”. Dædalus, 94(2):iii–vi, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026909.

Mumford:1965:UCM

[494] Lewis Mumford. Utopia, the city and the machine. Dædalus, 94(2):271–292, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026910.

Manuel:1965:TPH

[495] Frank E. Manuel. Toward a psychological history of utopias. Dædalus, 94(2):293–322, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026911.

Frye:1965:VLU

[496] Northrop Frye. Varieties of literary utopias. Dædalus, 94(2):323–347, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026912.

Brinton:1965:UD

[497] Crane Brinton. Utopia and democracy. Dædalus, 94(2):348–366, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026913.

Shklar:1965:PTU

[498] Judith Shklar. The political theory of utopia: From melancholy to Nostalgia. Dædalus, 94(2):367–381, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026914.

Ulam:1965:SU

[499] Adam Ulam. Socialism and utopia. Dædalus, 94(2):382–400, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20026915.

Lockwood:1965:EUA

[500] Maren Lockwood. The experimental utopia in America. Dædalus, 94 (2):401–418, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026916.

Bloch-Laine:1965:UUR

[501] François Bloch-Lainé. The utility of utopias for reformers. Dædalus, 94 (2):419–436, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026917.

deJouvenel:1965:UPP

[502] Bertrand de Jouvenel. Utopia for practical purposes. Dædalus, 94(2):437– 453, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20026918.

Kateb:1965:UGL

[503] George Kateb. Utopia and the good life. Dædalus, 94(2):454–473, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026919.

Sears:1965:ULL

[504] Paul B. Sears. Utopia and the living landscape. Dædalus, 94(2):474–486, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026920.

Smith:1965:EU

[505] John Maynard Smith. Eugenics and utopia. Dædalus, 94(2):487–505, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026921.

Pierce:1965:CTF

[506] John R. Pierce. Communications technology and the future. Dædalus, 94 (2):506–517, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026922.

Anonymous:1965:BMb

[507] Anonymous. Back matter. Dædalus, 94(2):??, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026923.

Anonymous:1965:FMb

[508] Anonymous. Front matter. Dædalus, 94(2):??, Spring 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026908.

G:1965:PIC

[509] S. R. G. Preface to the issue “Creativity and Learning”. Dædalus, 94 (3):iii–iv, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026926.

Wiesner:1965:ECS

[510] Jerome B. Wiesner. Education for creativity in the sciences. Dædalus, 94 (3):527–537, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026927.

Hawkins:1965:IVE

[511] David Hawkins. The informed vision: an essay on science education. Dædalus, 94(3):538–552, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026928.

Kagan:1965:PLP

[512] Jerome Kagan. Personality and the learning process. Dædalus, 94(3): 553–563, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026929.

Kubie:1965:UPS

[513] Lawrence S. Kubie. Unsolved problems of scientific education. Dædalus, 94(3):564–587, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026930.

Eiseley:1965:DCT

[514] Loren Eiseley. Darwin, Coleridge, and the theory of unconscious creation. Dædalus, 94(3):588–602, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026931.

Abelson:1965:RGA

[515] Philip H. Abelson. Relation of group activity to creativity in science. Dædalus, 94(3):603–614, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026932.

Brown:1965:DCT

[516] J. Douglas Brown. The development of creative teacher-scholars. Dædalus, 94(3):615–631, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026933.

Wing:1965:SSE

[517] Cliff W. Wing, Jr. Student selection, the educational environment, and the cultivation of talent. Dædalus, 94 (3):632–641, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026934.

Sanford:1965:HPI

[518] Nevitt Sanford. The human problems institute and general education. Dædalus, 94(3):642–662, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026935.

Torrance:1965:SVC

[519] E. Paul Torrance. Scientific views of creativity and factors affecting its growth. Dædalus, 94(3):663–681, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026936.

Haskins:1965:CES

[520] Caryl P. Haskins. The changing environments of science. Dædalus, 94 (3):682–712, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026937.

Kistiakowsky:1965:FSB

[521] George B. Kistiakowsky. On Federal support of basic research. Dædalus, 94 (3):713–734, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026938.

Anonymous:1965:BMc

[522] Anonymous. Back matter. Dædalus, 94(3):??, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026939.

Anonymous:1965:FMc

[523] Anonymous. Front matter. Dædalus, 94(3):??, Summer 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026925.

G:1965:PIN

[524] S. R. G. Preface to the issue “The Negro American”. Dædalus, 94(4):iii–iv, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026943.

Johnson:1965:FI

[525] Lyndon B. Johnson. Foreword to the issue. Dædalus, 94(4):743–744, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026944.

Moynihan:1965:EIO

[526] Daniel Patrick Moynihan. Employment, income, and the ordeal of the Negro family. Dædalus, 94(4):745–770, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026945.

Drake:1965:SES

[527] Clair Drake. The social and economic status of the Negro in the United States. Dædalus, 94(4):771–814, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026946.

Fein:1965:ESP

[528] Rashi Fein. An economic and social profile of the Negro American. Dædalus, 94(4):815–846, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026947.

Hauser:1965:DFI

[529] Philip M. Hauser. Demographic factors in the integration of the Negro. Dædalus, 94(4):847–877, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026948.

Tobin:1965:IES

[530] James Tobin. On improving the economic status of the Negro. Dædalus, 94(4):878–898, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026949.

Franklin:1965:TWR

[531] John Hope Franklin. The two worlds of race: a historical view. Dædalus, 94(4):899–920, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026950.

Fleming:1965:FEC

[532] Harold C. Fleming. The Federal executive and civil rights: 1961– 1965. Dædalus, 94(4):921–948, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026951.

Wilson:1965:NP

[533] James Q. Wilson. The Negro in politics. Dædalus, 94(4):949–973, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026952.

Pettigrew:1965:CCA

[534] Thomas F. Pettigrew. Complexity and change in American racial patterns: A social psychological view. Dædalus, 94(4):974–1008, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026953.

Parsons:1965:FCN

[535] Talcott Parsons. Full citizenship for the Negro American? A sociological problem. Dædalus, 94(4):1009–1054, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026954.

Emerson:1965:ADC

[536] Rupert Emerson and Martin Kilson. The American dilemma in a changing world: The rise of Africa and the Negro American. Dædalus, 94 (4):1055–1084, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026955.

Fichter:1965:ARN

[537] Joseph H. Fichter. American religion and the Negro. Dædalus, 94 (4):1085–1106, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026956.

Coles:1965:SD

[538] Robert Coles. It’s the same, but It’s different. Dædalus, 94(4):1107–1132, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026957.

Hughes:1965:AP

[539] Everett C. Hughes. Anomalies and projections. Dædalus, 94(4):1133–1147, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026958.

Turner:1965:WRT

[540] John B. Turner and Whitney M. Young, Jr. Who has the revolution or thoughts on the second reconstruction. Dædalus, 94(4):1148–1163, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20026959.

Anonymous:1965:BMd

[541] Anonymous. Back matter. Dædalus, 94(4):??, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026960.

Anonymous:1965:FMd

[542] Anonymous. Front matter. Dædalus, 94(4):??, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026942.

Anonymous:1965:VI

[543] Anonymous. Volume information. Dædalus, 94(4):??, Fall 1965. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20026941.

G:1966:PIN

[544] S. R. G. Preface to the issue “The Negro American—2”. Dædalus, 95 (1):iii–vi, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026963.

Edwards:1966:CCR

[545] G. Franklin Edwards. Community and class realities: The ordeal of change. Dædalus, 95(1):1–23, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026964.

Fischer:1966:RRR

[546] John H. Fischer. Race and reconciliation: The role of the school. Dædalus, 95(1):24–44, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026965.

Dentler:1966:BNS

[547] Robert A. Dentler. Barriers to Northern school desegregation. Dædalus, 95(1):45–63, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026966.

Abrams:1966:HPN

[548] Charles Abrams. The housing problem and the Negro. Dædalus, 95 (1):64–76, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026967.

Grier:1966:EBH

[549] Eunice Grier and George Grier. Equality and beyond: Housing segregation in the great society. Dædalus, 95 (1):77–106, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026968.

Foley:1966:NBS

[550] Eugene P. Foley. The Negro businessman: In search of a tradition. Dædalus, 95(1):107–144, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026969.

Erikson:1966:CIR

[551] Erik H. Erikson. The concept of identity in race relations: Notes and queries. Dædalus, 95(1):145–171, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026970.

Rainwater:1966:CIN

[552] Lee Rainwater. Crucible of identity: The Negro lower-class family. Dædalus, 95(1):172–216, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026971.

Sheatsley:1966:WAT

[553] Paul B. Sheatsley. White attitudes toward the Negro. Dædalus, 95 (1):217–238, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026972.

Clark:1966:CRM

[554] Kenneth B. Clark. The civil rights movement: Momentum and organization. Dædalus, 95(1):239–267, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026973.

Handlin:1966:GI

[555] Oscar Handlin. The goals of integration. Dædalus, 95(1):268–286, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026974.

Anonymous:1966:TAA

[556] Anonymous. Transcript of the American Academy Conference on the Negro American: May 14–15, 1965. Dædalus, 95(1):287–441, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026975.

Anonymous:1966:BMa

[557] Anonymous. Back matter. Dædalus, 95(1):??, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026976.

Anonymous:1966:FMa

[558] Anonymous. Front matter. Dædalus, 95(1):??, Winter 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026962.

G:1966:PIC

[559] S. R. G. Preface to the issue “Conditions of World Order”. Dædalus, 95(2):iii–iv, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026979.

Hoffmann:1966:RCC

[560] Stanley Hoffmann. Report of the Conference on Conditions of World Order: June 12–19, 1965, Villa Serbelloni, Bellagio, Italy. Dædalus, 95 (2):455–478, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026980.

Aron:1966:AOP

[561] Raymond Aron. The anarchical order of power. Dædalus, 95(2):479–502, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026981.

Kissinger:1966:DSF

[562] Henry A. Kissinger. Domestic structure and foreign policy. Dædalus, 95 (2):503–529, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026982.

Tinbergen:1966:IEP

[563] Jan Tinbergen. International economic planning. Dædalus, 95(2):530–557, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026983.

Fourastie:1966:RCW

[564] Jean Fourastié. Remarks on conditions of world order. Dædalus, 95 (2):558–571, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026984.

Gadamer:1966:NPF

[565] Hans-Georg Gadamer. Notes on planning for the future. Dædalus, 95 (2):572–589, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026985.

Aranguren:1966:OWA

[566] José Luis L. Aranguren. Openness to the world: an approach to world peace. Dædalus, 95(2):590–606, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026986.

Jaguaribe:1966:WOR

[567] Helio Jaguaribe. World order, rationality, and socioeconomic development. Dædalus, 95(2):607–626, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026987.

OConnell:1966:RIL

[568] D. P. O’Connell. The role of international law. Dædalus, 95(2):627–643, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026988.

Malek:1966:WOR

[569] Ivan Malek. World order and the responsibility of scientists: A functional as opposed to an institutional approach. Dædalus, 95(2):644–665, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026989.

Waddington:1966:DMP

[570] C. H. Waddington. The desire for material progress as a world ordering system. Dædalus, 95(2):666–673, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026990.

Nicol:1966:TWO

[571] Davidson Nicol. Toward a world order: an African viewpoint. Dædalus, 95(2):674–693, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026991.

Gladwyn:1966:WON

[572] Lord Gladwyn. World order and the nation-state: a regional approach. Dædalus, 95(2):694–703, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026992.

Anonymous:1966:BMb

[573] Anonymous. Back matter. Dædalus, 95(2):??, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026993.

Anonymous:1966:FMb

[574] Anonymous. Front matter. Dædalus, 95(2):??, Spring 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026978.

G:1966:PIT

[575] S. R. G. Preface to the issue “Tradition and Change”. Dædalus, 95 (3):iii–iv, Summer 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026996.

Meyerson:1966:EAC

[576] Martin Meyerson. The ethos of the American college student: Beyond the protests. Dædalus, 95(3): 713–739, Summer 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026997.

Bateson:1966:RTW

[577] Mary Catherine Bateson. “A Riddle of Two Worlds”: an interpretation of the poetry of H. N. Bialik. Dædalus, 95 (3):740–762, Summer 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026998.

Huntington:1966:PMT

[578] Samuel P. Huntington. The political modernization of traditional monarchies. Dædalus, 95(3):763–788, Summer 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20026999.

Scrimshaw:1966:UWF

[579] Nevin S. Scrimshaw. The urgency of world food problems. Dædalus, 95 (3):789–796, Summer 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027000.

Dyckman:1966:SCC

[580] John W. Dyckman. Some conditions of civic order in an urbanized world. Dædalus, 95(3):797–812, Summer 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027001.

Mus:1966:BWO

[581] Paul Mus. Buddhism and world order. Dædalus, 95(3):813–827, Summer 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027002.

vandenBergh:1966:CNW

[582] Godfried van Benthem van den Bergh. Contemporary nationalism in the Western World. Dædalus, 95(3): 828–861, Summer 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

                                        /www.jstor.org/stable/20027003.

Hoffmann:1966:OOF

[583] Stanley Hoffmann. Obstinate or obsolete? The fate of the nation-state and the case of Western Europe. Dædalus, 95(3):862–915, Summer 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027004.

Anonymous:1966:BMc

[584] Anonymous. Back matter. Dædalus, 95(3):??, Summer 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027005.

Anonymous:1966:FMc

[585] Anonymous. Front matter. Dædalus, 95(3):??, Summer 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20026995.

G:1966:PIF

[586] S. R. G. Preface to the issue “Fiction in Several Languages”. Dædalus, 95(4):iii–v, Fall 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027009.

Hibbett:1966:TTC

[587] Howard Hibbett. Tradition and trauma in the contemporary Japanese novel. Dædalus, 95(4):925–940, Fall 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027010.

Sfeir:1966:CAN

[588] George N. Sfeir. The contemporary Arabic novel. Dædalus, 95(4):941–960, Fall 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027011.

Gonzalez:1966:FN

[589] N. V. M. Gonzalez. The Filipino and the novel. Dædalus, 95(4):961–971, Fall 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027012.

Alter:1966:IN

[590] Robert Alter. The Israeli novel. Dædalus, 95(4):972–985, Fall 1966. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20027013.

Monegal:1966:CBN

[591] E. R. Monegal. The contemporary Brazilian novel. Dædalus, 95(4):986– 1003, Fall 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027014.

Milosz:1966:NP

[592] Czeslaw Milosz. The novel in Poland. Dædalus, 95(4):1004–1020, Fall 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027015.

Proussis:1966:NAT

[593] Costas M. Proussis. The novels of Angelos Terzakis. Dædalus, 95 (4):1021–1045, Fall 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027016.

Hiebert:1966:UAT

[594] Erwin N. Hiebert. The uses and abuses of thermodynamics in religion. Dædalus, 95(4):1046–1080, Fall 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027017.

Anonymous:1966:BMd

[595] Anonymous. Back matter. Dædalus, 95(4):??, Fall 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027018.

Anonymous:1966:FMd

[596] Anonymous. Front matter. Dædalus, 95(4):??, Fall 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027008.

Anonymous:1966:VI

[597] Anonymous. Volume information. Dædalus, 95(4):??, Fall 1966. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027007.

G:1967:PIR

[598] S. R. G. Preface to the issue “Religion in America”. Dædalus, 96 (1):iii–vii, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027021.

Bellah:1967:CRA

[599] Robert N. Bellah. Civil religion in America. Dædalus, 96(1):1–21, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027022.

Littell:1967:CBP

[600] Franklin H. Littell. The churches and the body politic. Dædalus, 96 (1):22–42, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027023.

McLoughlin:1967:TTF

[601] William G. McLoughlin. Is there a third force in Christendom? Dædalus, 96(1):43–68, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027024.

Gilkey:1967:SIS

[602] Langdon Gilkey. Social and intellectual sources of contemporary Protestant theology in America. Dædalus, 96(1):69–98, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027025.

Marty:1967:SHE

[603] Martin E. Marty. The spirit’s holy errand: The search for a spiritual style in secular America. Dædalus, 96(1):99–115, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027026.

ODea:1967:CCR

[604] Thomas F. O’Dea. The crisis of the contemporary religious consciousness. Dædalus, 96(1):116–134, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027027.

Cox:1967:NBA

[605] Harvey G. Cox. The “New Breed” in American churches: Sources of social activism in American religion. Dædalus, 96(1):135–150, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027028.

Callahan:1967:QSR

[606] Daniel Callahan. The quest for social relevance. Dædalus, 96(1):151–179, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027029.

Katz:1967:RPS

[607] Wilber G. Katz and Harold P. Southerland. Religious pluralism and the Supreme Court. Dædalus, 96(1):180– 192, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027030.

Fackenheim:1967:SEF

[608] Emil L. Fackenheim. On the selfexposure of faith to the modern-secular world: Philosophical reflections in the light of Jewish experience. Dædalus, 96(1):193–219, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027031.

Himmelfarb:1967:SSJ

[609] Milton Himmelfarb. Secular society? A Jewish perspective. Dædalus, 96 (1):220–236, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027032.

Novak:1967:CRS

[610] Michael Novak. Christianity: Renewed or slowly abandoned? Dædalus, 96 (1):237–266, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027033.

Anonymous:1967:BMa

[611] Anonymous. Back matter. Dædalus, 96(1):??, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027034.

Anonymous:1967:FMa

[612] Anonymous. Front matter. Dædalus, 96(1):??, Winter 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027020.

G:1967:PIC

[613] S. R. G. Preface to the issue “Color and Race”. Dædalus, 96(2):iii–x, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027037.

Shils:1967:CUI

[614] Edward Shils. Color, the universal intellectual community, and the Afro–Asian intellectual. Dædalus, 96 (2):279–295, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027038.

Gardiner:1967:RCI

[615] Robert K. A. Gardiner. Race and color in international relations. Dædalus, 96(2):296–311, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027039.

Bastide:1967:CRC

[616] Roger Bastide. Color, racism, and Christianity. Dædalus, 96(2):312–327, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027040.

Mason:1967:RAW

[617] Philip Mason. The revolt against Western values. Dædalus, 96(2):328–352, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027041.

Isaacs:1967:GIP

[618] Harold R. Isaacs. Group identity and political change: The role of color and physical characteristics. Dædalus, 96 (2):353–375, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027042.

Raveau:1967:ORC

[619] François Raveau. An outline of the role of color in adaptation phenomena. Dædalus, 96(2):376–389, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027043.

Gergen:1967:SSC

[620] Kenneth J. Gergen. The significance of skin color in human relations. Dædalus, 96(2):390–406, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027044.

Wagatsuma:1967:SPS

[621] Hiroshi Wagatsuma. The social perception of skin color in Japan. Dædalus, 96(2):407–443, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027045.

Beteille:1967:RDS

[622] André Béteille. Race and descent as social categories in India. Dædalus, 96(2):444–463, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027046.

Brown:1967:CNA

[623] Leon Carl Brown. Color in Northern Africa. Dædalus, 96(2):464–482, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027047.

Legum:1967:CPS

[624] Colin Legum. Color and power in the South African situation. Dædalus, 96(2):483–495, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027048.

Braithwaite:1967:CIB

[625] E. R. Braithwaite. The “Colored Immigrant” in Britain. Dædalus, 96 (2):496–511, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027049.

Little:1967:SAC

[626] Kenneth Little. Some aspects of color, class, and culture in Britain. Dædalus, 96(2):512–526, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027050.

Lincoln:1967:CGI

[627] C. Eric Lincoln. Color and group identity in the United States. Dædalus, 96(2):527–541, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027051.

Pitt-Rivers:1967:RCC

[628] Julian Pitt-Rivers. Race, color, and class in Central America and the Andes. Dædalus, 96(2):542–559, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027052.

Fernandes:1967:WP

[629] Florestan Fernandes. The weight of the past. Dædalus, 96(2):560–579, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027053.

Lowenthal:1967:RCW

[630] David Lowenthal. Race and color in the West Indies. Dædalus, 96 (2):580–626, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027054.

Anonymous:1967:BMb

[631] Anonymous. Back matter. Dædalus, 96(2):??, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027055.

Anonymous:1967:FMb

[632] Anonymous. Front matter. Dædalus, 96(2):??, Spring 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027036.

G:1967:PIT

[633] S. R. G. Preface to the issue “Toward the Year 2000: Work in Progress”. Dædalus, 96(3):v–vi, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027058.

Bell:1967:YTI

[634] Daniel Bell. The year 2000: The trajectory of an idea. Dædalus, 96 (3):639–651, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027059.

Anonymous:1967:PM

[635] Anonymous. Preliminary memorandum. Dædalus, 96(3):652–655, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027060.

Anonymous:1967:BFO

[636] Anonymous. Baselines for the future: October 22–24, 1965. Dædalus, 96 (3):656–665, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027061.

Anonymous:1967:AF

[637] Anonymous. Alternative futures. Dædalus, 96(3):666–679, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027062.

Anonymous:1967:CD

[638] Anonymous. Centralization and decentralization. Dædalus, 96(3):680–690, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027063.

Anonymous:1967:NM

[639] Anonymous. The need for models. Dædalus, 96(3):691–697, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027064.

Bell:1967:SCa

[640] Daniel Bell. A summary by the chairman. Dædalus, 96(3):698–704, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027065.

Kahn:1967:NTT

[641] Herman Kahn and Anthony J. Wiener. The next thirty-three years: a framework for speculation. Dædalus, 96 (3):705–732, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027066.

Ikle:1967:CSP

[642] Fred Charles Iklé. Can social predictions be evaluated? Dædalus, 96 (3):733–758, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027067.

Schon:1967:FTF

[643] Donald A. Schon. Forecasting and technological forecasting. Dædalus, 96 (3):759–770, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027068.

Shubik:1967:IRF

[644] Martin Shubik. Information, rationality, and free choice in a future democratic society. Dædalus, 96(3): 771–778, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027069.

Duhl:1967:PPW

[645] Leonard J. Duhl. Planning and predicting: Or what to do when you don’t know the names of the variables. Dædalus, 96(3):779–788, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027070.

Perloff:1967:MUD

[646] Harvey S. Perloff. Modernizing urban development. Dædalus, 96(3): 789–800, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027071.

Moynihan:1967:RFL

[647] Daniel P. Moynihan. The relationship of Federal to local authorities. Dædalus, 96(3):801–808, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027072.

Frank:1967:NNP

[648] Lawrence K. Frank. The need for a new political theory. Dædalus, 96 (3):809–816, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027073.

Graubard:1967:UCY

[649] Stephen R. Graubard. University cities in the year 2000. Dædalus, 96 (3):817–822, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027074.

Orlans:1967:ESI

[650] Harold Orlans. Educational and scientific institutions. Dædalus, 96(3): 823–831, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027075.

Mayr:1967:BMY

[651] Ernst Mayr. Biological man and the year 2000. Dædalus, 96(3):832–836, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027076.

Quarton:1967:DEC

[652] Gardner C. Quarton. Deliberate efforts to control human behavior and modify personality. Dædalus, 96(3): 837–853, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027077.

Stendahl:1967:RMI

[653] Krister Stendahl. Religion, mysticism, and the institutional Church. Dædalus, 96(3):854–859, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027078.

Erikson:1967:MY

[654] Erik H. Erikson. Memorandum on youth. Dædalus, 96(3):860–870, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027079.

Mead:1967:LCV

[655] Margaret Mead. The life cycle and its variations: The division of roles. Dædalus, 96(3):871–875, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027080.

Kalven:1967:PPY

[656] Harry Kalven, Jr. The problems of privacy in the year 2000. Dædalus, 96 (3):876–882, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027081.

Miller:1967:SPP

[657] George A. Miller. Some psychological perspectives on the year 2000. Dædalus, 96(3):883–896, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027082.

Riesman:1967:NM

[658] David Riesman. Notes on meritocracy. Dædalus, 96(3):897–908, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027083.

Pierce:1967:C

[659] John R. Pierce. Communication. Dædalus, 96(3):909–921, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027084.

Rostow:1967:TAF

[660] Eugene V. Rostow. Thinking about the future of international society. Dædalus, 96(3):922–926, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027085.

Huntington:1967:PDD

[661] Samuel P. Huntington. Political development and the decline of the American system of world order. Dædalus, 96(3):927–929, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027086.

Pool:1967:ISN

[662] Ithiel de Sola Pool. The international system in the next half century. Dædalus, 96(3):930–935, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027087.

Anonymous:1967:NLF

[663] Anonymous. The nature and limitations of forecasting: February 11–12, 1966. Dædalus, 96(3):936–947, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027088.

Anonymous:1967:FF

[664] Anonymous. Four futures. Dædalus, 96(3):948–963, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027089.

Anonymous:1967:NNS

[665] Anonymous. The need for normative statements. Dædalus, 96(3): 964–974, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027090.

Bell:1967:SCb

[666] Daniel Bell. A summary by the chairman. Dædalus, 96(3):975–977, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027091.

Anonymous:1967:MCY

[667] Anonymous. Members of the commission on the year 2000. Dædalus, 96(3):978, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027092.

Anonymous:1967:WPS

[668] Anonymous. The working parties: Summer 1967. Dædalus, 96(3):979–984, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027093.

Bell:1967:CWF

[669] Daniel Bell. Coda: Work in further progress. Dædalus, 96(3):985–988, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027094.

Anonymous:1967:BMe

[670] Anonymous. Back matter. Dædalus, 96(3):??, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027095.

Anonymous:1967:FMe

[671] Anonymous. Front matter. Dædalus, 96(3):??, Summer 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027057.

G:1967:PIA

[672] S. R. G. Preface to the issue “America’s Changing Environment”. Dædalus, 96(4):iii–viii, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027099.

Darling:1967:WEE

[673] F. Fraser Darling. A wider environment of ecology and conservation. Dædalus, 96(4):1003–1019, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027100.

Patterson:1967:AI

[674] Robert W. Patterson. The art of the impossible. Dædalus, 96(4):1020–1033, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027101.

Landsberg:1967:URO

[675] Hans H. Landsberg. The U.S. resource outlook: Quantity and quality. Dædalus, 96(4):1034–1057, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027102.

Krutilla:1967:SEE

[676] John V. Krutilla. Some environmental effects of economic development. Dædalus, 96(4):1058–1070, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027103.

Allee:1967:AAR

[677] David Allee. American agriculture: Its resource issues for the coming years. Dædalus, 96(4):1071–1081, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027104.

Teller:1967:APA

[678] Azriel Teller. Air-pollution abatement: Economic rationality and reality. Dædalus, 96(4):1082–1098, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027105.

Wollman:1967:NER

[679] Nathaniel Wollman. The new economics of resources. Dædalus, 96 (4):1099–1114, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027106.

Wildavsky:1967:APT

[680] Aaron Wildavsky. Aesthetic power or the triumph of the sensitive minority over the vulgar mass: a political analysis of the new economics. Dædalus, 96(4):1115–1128, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027107.

Spilhaus:1967:EC

[681] Athelstan Spilhaus. The experimental city. Dædalus, 96(4):1129–1141, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027108.

Gilliam:1967:FSP

[682] Harold Gilliam. The fallacy of single-purpose planning. Dædalus, 96 (4):1142–1157, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027109.

Papageorgiou:1967:ASO

[683] Alexis Papageorgiou. Architectural schemata for outdoor recreation areas of Tomorrow. Dædalus, 96(4):1158– 1171, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027110.

Revelle:1967:ORH

[684] Roger Revelle. Outdoor recreation in a hyper-productive society. Dædalus, 96(4):1172–1191, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027111.

Ripley:1967:ESP

[685] S. Dillon Ripley and Helmut K. Buechner. Ecosystem science as a point of synthesis. Dædalus, 96(4):1192–1199, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027112.

Maslach:1967:RER

[686] George J. Maslach. The reorganization of educational resources. Dædalus, 96(4):1200–1209, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027113.

Morison:1967:EEC

[687] Robert S. Morison. Education for environmental concerns. Dædalus, 96 (4):1210–1223, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027114.

Anonymous:1967:BMc

[688] Anonymous. Back matter. Dædalus, 96(4):??, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027115.

Anonymous:1967:FMc

[689] Anonymous. Front matter. Dædalus, 96(4):??, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027098.

Anonymous:1967:VI

[690] Anonymous. Volume information. Dædalus, 96(4):??, Fall 1967. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027097.

Graubard:1968:NTA

[691] Stephen R. Graubard. Note on the tenth anniversary. Dædalus, 97(??):1– 6, ???? 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/

stable/20023854. Special issue with the Tenth Anniversary Index: 1958– 1968.

Anonymous:1968:TYC

[692] Anonymous. Ten-year cumulative index 1958–1968. Dædalus, 97(??):7–27, ???? 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/

stable/20023855. Special issue with the Tenth Anniversary Index: 1958– 1968.

Anonymous:1968:TYA

[693] Anonymous. Ten-year author index 1958–1968. Dædalus, 97(??):29–45, ???? 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/

stable/20023856. Special issue with the Tenth Anniversary Index: 1958– 1968.

Anonymous:1968:FMe

[694] Anonymous. Front matter. Dædalus, 97(??):??, ???? 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.jstor.org/stable/20023853.

Special issue with the Tenth Anniversary Index: 1958–1968.

Graubard:1968:NA

[695] Stephen R. Graubard. Note on the 10th anniversary. Dædalus, 97(1):v–xii, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023795.

Lipset:1968:SPC

[696] Seymour Martin Lipset. Students and politics in comparative perspective. Dædalus, 97(1):1–20, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023796.

Moore:1968:SUN

[697] Clement H. Moore and Arlie R. Hochschild. Student unions in North African politics. Dædalus, 97(1):21–50, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023797.

Finlay:1968:SPG

[698] David J. Finlay. Students and politics in Ghana. Dædalus, 97(1):51–69, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023798.

Scott:1968:SPA

[699] Robert E. Scott. Student political activism in Latin America. Dædalus, 97(1):70–98, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023799.

Glazer:1968:SPC

[700] Myron Glazer. Student politics in a Chilean university. Dædalus, 97 (1):99–115, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023800.

Halsey:1968:BSP

[701] A. H. Halsey and Stephen Marks. British student politics. Dædalus, 97 (1):116–136, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023801.

Pinner:1968:TTW

[702] Frank A. Pinner. Tradition and transgression: Western European students in the postwar world. Dædalus, 97 (1):137–155, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023802.

Allardt:1968:SSS

[703] Erik Allardt and Richard F. Tomasson. Stability and strains in Scandinavian student politics. Dædalus, 97 (1):156–165, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023803.

Cornell:1968:SPC

[704] Richard Cornell. Students and politics in the communist countries of Eastern Europe. Dædalus, 97(1):166–183, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023804.

Roos:1968:SPT

[705] Leslie L. Roos, Jr., Noralou P. Roos, and Gary R. Field. Students and politics in Turkey. Dædalus, 97 (1):184–203, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023805.

Shimbori:1968:SSM

[706] Michiya Shimbori. The sociology of a student movement: a Japanese case study. Dædalus, 97(1):204–228, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023806.

Israel:1968:RMC

[707] John Israel. Reflections on the modern Chinese student movement. Dædalus, 97(1):229–253, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023807.

Altbach:1968:SPH

[708] Philip G. Altbach. Student politics and higher education in India. Dædalus, 97(1):254–273, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023808.

Silverstein:1968:BSP

[709] Josef Silverstein. Burmese student politics in a changing society. Dædalus, 97(1):274–292, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023809.

Peterson:1968:SLA

[710] Richard E. Peterson. The student left in American higher education. Dædalus, 97(1):293–317, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023810.

Duberman:1968:EE

[711] Martin Duberman. An experiment in education. Dædalus, 97(1):318–341, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023811.

Anonymous:1968:BMa

[712] Anonymous. Back matter. Dædalus, 97(1):??, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023812.

Anonymous:1968:FMa

[713] Anonymous. Front matter. Dædalus, 97(1):??, Winter 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023794.

G:1968:PIH

[714] S. R. G. Preface to the issue “Historical Population Studies”. Dædalus, 97(2):v–vi, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023815.

Revelle:1968:IIH

[715] Roger Revelle. Introduction to the issue “Historical Population Studies”. Dædalus, 97(2):353–362, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023816.

Landes:1968:TPH

[716] David Landes. The treatment of population in history textbooks. Dædalus, 97(2):363–384, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023817.

Henry:1968:HD

[717] Louis Henry. Historical demography. Dædalus, 97(2):385–396, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023818.

Banks:1968:HSS

[718] J. A. Banks. Historical sociology and the study of population. Dædalus, 97 (2):397–414, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023819.

Hollingsworth:1968:IQD

[719] T. H. Hollingsworth. The importance of the quality of the data in historical demography. Dædalus, 97 (2):415–432, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023820.

Spengler:1968:DFE

[720] Joseph J. Spengler. Demographic factors and early modern economic development. Dædalus, 97(2):433–446, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023821.

Heer:1968:EDF

[721] David M. Heer. Economic development and the fertility transition. Dædalus, 97(2):447–462, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023822.

Noonan:1968:IDH

[722] John T. Noonan, Jr. Intellectual and demographic history. Dædalus, 97 (2):463–485, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023823.

vandeWalle:1968:MMF

[723] Etienne van de Walle. Marriage and marital fertility. Dædalus, 97 (2):486–501, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023824.

Demeny:1968:EFD

[724] Paul Demeny. Early fertility decline in Austria-hungary: a lesson in demographic transition. Dædalus, 97 (2):502–522, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023825.

Livi-Bacci:1968:FPG

[725] Massimo Livi-Bacci. Fertility and population growth in Spain in the Eighteenth and Nineteenth Centuries. Dædalus, 97(2):523–535, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023826.

Biraben:1968:CDC

[726] Jean-Noël Biraben. Certain demographic characteristics of the plague epidemic in France, 1720–22. Dædalus, 97(2):536–545, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023827.

Wrigley:1968:MPI

[727] E. A. Wrigley. Mortality in preindustrial England: The example of Colyton, Devon, over three centuries. Dædalus, 97(2):546–580, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023828.

Glass:1968:NDL

[728] D. V. Glass. Notes on the demography of London at the end of the Seventeenth Century. Dædalus, 97 (2):581–592, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023829.

Goubert:1968:LFI

[729] Pierre Goubert. Legitimate fecundity and infant mortality in France during the Eighteenth Century: A comparison. Dædalus, 97(2):593–603, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023830.

vanBath:1968:HDS

[730] Bernard H. Slicher van Bath. Historical demography and the social and economic development of the Netherlands. Dædalus, 97(2):604–621, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023831.

Hanley:1968:PTE

[731] Susan B. Hanley. Population trends and economic development in Tokugawa Japan: The case of Bizen Province in Okayama. Dædalus, 97 (2):622–635, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023832.

Holton:1968:MES

[732] Gerald Holton. Mach, Einstein, and the search for reality. Dædalus, 97(2):636– 673, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20023833. Reprinted in

[4247, pp. 219–259].

Anonymous:1968:BMb

[733] Anonymous. Back matter. Dædalus, 97(2):??, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023834.

Anonymous:1968:FMb

[734] Anonymous. Front matter. Dædalus, 97(2):??, Spring 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023814.

G:1968:PIP

[735] S. R. G. Preface to the issue “Philosophers and Kings: Studies in Leadership”. Dædalus, 97(3):v–vi, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023837.

Rustow:1968:IIP

[736] Dankwart A. Rustow. Introduction to the issue “Philosophers and Kings: Studies in Leadership”. Dædalus, 97 (3):683–694, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023838.

Erikson:1968:NPH

[737] Erik H. Erikson. On the nature of psycho-historical evidence: In search of Gandhi. Dædalus, 97(3):695–730, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023839.

Tucker:1968:TCL

[738] Robert C. Tucker. The theory of charismatic leadership. Dædalus, 97 (3):731–756, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023840.

Apter:1968:NCC

[739] David E. Apter. Nkrumah, charisma, and the coup. Dædalus, 97(3):757–792, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20023841.

Rustow:1968:AFS

[740] Dankwart A. Rustow. Atatürk as founder of a state. Dædalus, 97 (3):793–828, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023842.

Hoffmann:1968:WGG

[741] Stanley Hoffmann and Inge Hoffmann. The will to grandeur: De Gaulle as political artist. Dædalus, 97(3): 829–887, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023843.

Kissinger:1968:WRR

[742] Henry A. Kissinger. The white revolutionary: Reflections on Bismarck. Dædalus, 97(3):888–924, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023844.

Hirschman:1968:UOP

[743] Albert O. Hirschman. Underdevelopment, obstacles to the perception of change, and leadership. Dædalus, 97 (3):925–937, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023845.

Barber:1968:AIP

[744] James D. Barber. Adult identity and Presidential style: The rhetorical emphasis. Dædalus, 97(3):938–968, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023846.

Manuel:1968:NAS

[745] Frank E. Manuel. Newton as autocrat of science. Dædalus, 97(3): 969–1001, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023847.

Karl:1968:PIP

[746] Barry D. Karl. The power of intellect and the politics of ideas. Dædalus, 97 (3):1002–1035, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023848.

Mazlish:1968:JMU

[747] Bruce Mazlish. James Mill and the utilitarians. Dædalus, 97(3):1036–1061, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20023849.

Strout:1968:WJT

[748] Cushing Strout. William James and the twice-born sick soul. Dædalus, 97 (3):1062–1082, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023850.

Anonymous:1968:BMc

[749] Anonymous. Back matter. Dædalus, 97(3):??, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023851.

Anonymous:1968:FMc

[750] Anonymous. Front matter. Dædalus, 97(3):??, Summer 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023836.

G:1968:PIC

[751] S. R. G. Preface to the issue “The Conscience of the City”. Dædalus, 97(4):v– vi, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013411.

Meyerson:1968:IIC

[752] Martin Meyerson. Introduction to the issue “The Conscience of the City”. Dædalus, 97(4):vii–xi, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013412.

Webber:1968:PCA

[753] Melvin M. Webber. The post-city age. Dædalus, 97(4):1091–1110, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013413.

Boulding:1968:CEI

[754] Kenneth E. Boulding. The city as an element in the international system. Dædalus, 97(4):1111–1123, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013414.

Seeley:1968:RUS

[755] John R. Seeley. Remaking the urban scene: New youth in an old environment. Dædalus, 97(4):1124–1139, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013415.

Fitch:1968:EGU

[756] Lyle C. Fitch. Eight goals for an urbanizing America. Dædalus, 97 (4):1141–1164, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013416.

Bacon:1968:UP

[757] Edmund N. Bacon. Urban process. Dædalus, 97(4):1165–1178, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013417.

Michael:1968:CCP

[758] Donald N. Michael. On coping with complexity: Planning and politics. Dædalus, 97(4):1179–1193, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013418.

George:1968:PLS

[759] Alexander L. George. Political leadership and social change in American cities. Dædalus, 97(4):1194–1217, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013419.

Lindsay:1968:MIC

[760] Franklin A. Lindsay. Managerial innovation and the cities. Dædalus, 97 (4):1218–1230, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013420.

Banfield:1968:WGC

[761] Edward C. Banfield. Why government cannot solve the urban problem. Dædalus, 97(4):1231–1241, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013421.

Drucker:1968:WWM

[762] Peter F. Drucker. Worker and work in the Metropolis. Dædalus, 97 (4):1243–1262, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013422.

Yarmolinsky:1968:SS

[763] Adam Yarmolinsky. The service society. Dædalus, 97(4):1263–1276, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013423.

Carr:1968:WLH

[764] Stephen Carr and Kevin Lynch. Where learning happens. Dædalus, 97(4): 1277–1291, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

                                        /www.jstor.org/stable/20013424.

Meier:1968:MTM

[765] Richard L. Meier. The Metropolis as a transaction-maximizing system. Dædalus, 97(4):1292–1313, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20013425.

Coles:1968:LIA

[766] Robert Coles. Like it is in the alley. Dædalus, 97(4):1315–1330, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20013426.

Downs:1968:AFA

[767] Anthony Downs. Alternative futures for the American ghetto. Dædalus, 97(4):1331–1378, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20013427.

Young:1968:LAW

[768] Michael Young. The liberal approach: Its weaknesses and its strengths: a comment on the U.S. Riot Commission Report. Dædalus, 97(4):1379–1389, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20013428.

Lerner:1968:NAH

[769] Max Lerner. The Negro American and his city: Person in place in culture. Dædalus, 97(4):1390–1408, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013429.

Meyerson:1968:UPR

[770] Martin Meyerson. Urban policy: Reforming reform. Dædalus, 97(4):1410– 1430, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20013430.

Anonymous:1968:BMd

[771] Anonymous. Back matter. Dædalus, 97(4):??, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013431.

Anonymous:1968:FMd

[772] Anonymous. Front matter. Dædalus, 97(4):??, Fall 1968. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013410.

G:1969:PIP

[773] S. R. G. Preface to the issue “Perspectives on Business”. Dædalus, 98(1):v–xx, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023860.

Letwin:1969:PFA

[774] William Letwin. The past and future of the American businessman. Dædalus, 98(1):1–22, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023861.

Chandler:1969:RBU

[775] Alfred D. Chandler, Jr. The role of business in the United States: A historical survey. Dædalus, 98(1):23–40, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023862.

Holton:1969:BG

[776] Richard H. Holton. Business and government. Dædalus, 98(1):41–59, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023863.

Averitt:1969:ABA

[777] Robert T. Averitt. American business: Achievement and challenge. Dædalus, 98(1):60–77, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023864.

Goldston:1969:NPA

[778] Eli Goldston. New prospects for American business. Dædalus, 98(1):78–112, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023865.

Vernon:1969:RUE

[779] Raymond Vernon. The role of U.S. enterprise abroad. Dædalus, 98(1):113– 133, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20023866.

Chamberlain:1969:LMF

[780] Neil W. Chamberlain. The life of the mind in the firm. Dædalus, 98 (1):134–146, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023867.

Crozier:1969:NRA

[781] Michel Crozier. A new rationale for American business. Dædalus, 98 (1):147–158, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023868.

Monsen:1969:ABV

[782] R. Joseph Monsen. The American business view. Dædalus, 98(1):159–173, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023869.

Silk:1969:BPT

[783] Leonard S. Silk. Business power, today and Tomorrow. Dædalus, 98 (1):174–190, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023870.

Shonfield:1969:BTF

[784] Andrew Shonfield. Business in the Twenty-First Century. Dædalus, 98 (1):191–207, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023871.

Anonymous:1969:BMa

[785] Anonymous. Back matter. Dædalus, 98(1):??, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023872.

Anonymous:1969:FMa

[786] Anonymous. Front matter. Dædalus, 98(1):??, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023859.

Anonymous:1969:VIa

[787] Anonymous. Volume information. Dædalus, 98(1):??, Winter 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023858.

G:1969:PIE

[788] S. R. G. Preface to the issue “Ethical Aspects of Experimentation with Human Subjects”. Dædalus, 98(2):v–vii, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023875.

Freund:1969:IIE

[789] Paul A. Freund. Introduction to the issue “Ethical Aspects of Experimentation with Human Subjects”. Dædalus, 98(2):viii–xiv, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023876.

Jonas:1969:PRE

[790] Hans Jonas. Philosophical reflections on experimenting with human subjects. Dædalus, 98(2):219–247, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023877.

Blumgart:1969:MFV

[791] Herrman L. Blumgart. The medical framework for viewing the problem of human experimentation. Dædalus, 98 (2):248–274, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023878.

Beecher:1969:SRM

[792] Henry K. Beecher. Scarce resources and medical advancement. Dædalus, 98(2):275–313, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023879.

Freund:1969:LFH

[793] Paul A. Freund. Legal frameworks for human experimentation. Dædalus, 98(2):314–324, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023880.

Parsons:1969:RHS

[794] Talcott Parsons. Research with human subjects and the “Professional Complex”. Dædalus, 98(2):325–360, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023881.

Mead:1969:RHB

[795] Margaret Mead. Research with human beings: a model derived from anthropological field practice. Dædalus, 98(2):361–386, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023882.

Calabresi:1969:RME

[796] Guido Calabresi. Reflections on medical experimentation in humans. Dædalus, 98(2):387–405, Spring 1969. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20023883.

Jaffe:1969:LSC

[797] Louis L. Jaffe. Law as a system of control. Dædalus, 98(2):406–426, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023884.

Cavers:1969:LCC

[798] David F. Cavers. The legal control of the clinical investigation of drugs: Some political, economic, and social questions. Dædalus, 98(2):427–448, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023885.

Lasagna:1969:SSH

[799] Louis Lasagna. Special subjects in human experimentation. Dædalus, 98 (2):449–462, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023886.

Edsall:1969:PAP

[800] Geoffrey Edsall. A positive approach to the problem of human experimentation. Dædalus, 98(2):463–479, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023887.

Katz:1969:EPI

[801] Jay Katz. The education of the physician-investigator. Dædalus, 98 (2):480–501, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023888.

Moore:1969:TIE

[802] Francis D. Moore. Therapeutic innovation: Ethical boundaries in the initial clinical trials of new drugs and surgical procedures. Dædalus, 98 (2):502–522, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023889.

Rutstein:1969:EDH

[803] David D. Rutstein. The ethical design of human experiments. Dædalus, 98(2):523–541, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023890.

Curran:1969:GRU

[804] William J. Curran. Governmental regulation of the use of human subjects in medical research: The approach of two Federal agencies. Dædalus, 98 (2):542–594, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023891.

Anonymous:1969:BMb

[805] Anonymous. Back matter. Dædalus, 98(2):??, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023892.

Anonymous:1969:FMb

[806] Anonymous. Front matter. Dædalus, 98(2):??, Spring 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023874.

G:1969:PIF

[807] S. R. G. Preface to the issue “The Future of the Humanities”. Dædalus, 98(3):v–vi, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023895.

Ackerman:1969:IIF

[808] James S. Ackerman. Introduction to the issue “The Future of the Humanities”. Dædalus, 98(3):605–614, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023896.

Ong:1969:CUH

[809] Walter J. Ong. Crisis and understanding in the humanities. Dædalus, 98 (3):617–640, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023897.

Roush:1969:WWB

[810] G. Jon Roush. What will become of the past? Dædalus, 98(3):641–653, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023898.

Blau:1969:RSM

[811] Herbert Blau. Relevance: The shadow of a magnitude. Dædalus, 98(3): 654–676, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023899.

Shattuck:1969:TH

[812] Roger W. Shattuck. Thoughts on the humanities. Dædalus, 98(3):677–683, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20023900.

Coles:1969:WMS

[813] Robert Coles. The words and music of social change. Dædalus, 98 (3):684–698, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023901.

Kaufmann:1969:KOP

[814] R. J. Kaufmann. On knowing One’s place: a humanistic meditation. Dædalus, 98(3):699–713, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023902.

Anonymous:1969:EBI

[815] Anonymous. Excerpts from Conference I. Dædalus, 98(3):715–735, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023903.

Kirchner:1969:NU

[816] Leon Kirchner. Notes on understanding. Dædalus, 98(3):739–746, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023904.

Kahn:1969:UPT

[817] Wolf Kahn. Uses of painting today. Dædalus, 98(3):747–754, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023905.

Reichek:1969:PTA

[818] Jesse Reichek. A painter and teacher as amateur humanist. Dædalus, 98 (3):755–764, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023906.

OHare:1969:DD

[819] Michael O’Hare. Designers’ dilemma. Dædalus, 98(3):765–777, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023908.

Anonymous:1969:EBII

[820] Anonymous. Excerpts from Conference II. Dædalus, 98(3):779–800, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023909.

Else:1969:ONH

[821] Gerald F. Else. The old and the new humanities. Dædalus, 98(3):

803–808, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023910.

Whitman:1969:WCO

[822] Cedric H. Whitman. Why not the classics? An old-fashioned view. Dædalus, 98(3):809–814, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023911.

Putnam:1969:NCS

[823] Michael C. J. Putnam. Notes on classical studies. Dædalus, 98(3): 815–823, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023912.

Steinberg:1969:OSS

[824] Leo Steinberg. Objectivity and the shrinking self. Dædalus, 98(3):824–836, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20023913.

Rosenfield:1969:ARV

[825] John M. Rosenfield. The arts in the realm of values. Dædalus, 98 (3):837–843, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023914.

Treitler:1969:RRH

[826] Leo Treitler. On responsibility and relevance in humanistic disciplines. Dædalus, 98(3):844–852, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023915.

Ackerman:1969:TSC

[827] James S. Ackerman. Two styles: a challenge to higher education. Dædalus, 98 (3):855–869, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023916.

Anonymous:1969:BMc

[828] Anonymous. Back matter. Dædalus, 98(3):??, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023917.

Anonymous:1969:FMc

[829] Anonymous. Front matter. Dædalus, 98(3):??, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023894.

Anonymous:1969:I

[830] Anonymous. [illustrations]. Dædalus, 98(3):??, Summer 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023907.

G:1969:PID

[831] S. R. G. Preface to the issue “Dialogues”. Dædalus, 98(4):v–viii, Fall 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023921.

Beer:1969:NTH

[832] Samuel H. Beer, Lee Benson, Felix Gilbert, Stephen R. Graubard, David J. Herlihy, Stanley Hoffmann, Carl Kaysen, Leonard Krieger, Thomas S. Kuhn, David Landes, Joseph Levenson, Frank E. Manuel, J. G. A. Pocock, David J. Rothman, Carl E. Schorske, Lawrence Stone, and Charles Tilly. New trends in history. Dædalus, 98(4):888–976, Fall 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023922.

Geertz:1969:LHS

[833] Clifford Geertz, Martin Malia, Talcott Parsons, Roger Revelle, and Eric Weil. The languages of the humanistic studies. Dædalus, 98(4):978–1028, Fall 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023923.

Bell:1969:GU

[834] Daniel Bell, John Brademas, Jill Conway, Martin Duberman, Robert J. Glaser, Stephen R. Graubard, Hanna H. Gray, Carl Kaysen, Edward H. Levi, Martin Meyerson, Robert S. Morison, Talcott Parsons, Bruce L. Payne, David Riesman, Neil R. Rudenstine, Preble Stolz, and David M. Wax. The governance of the universities I. Dædalus, 98(4):1030–1091, Fall 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

/www.jstor.org/stable/20023924.

Bolling:1969:GUI

[835] Landrum R. Bolling, Peter J. Caws, Sarah E. Diamant, C. M. Dick, Jr., Seymour Eskow, Edgar Z. Friedenberg, Stephen R. Graubard, Andrew M. Greeley, Jeff Greenfield, Eugene E. Grollmes, Gerald Holton, Willard Hurst, Dexter M. Keezer, Clark Kerr, S. E. Luria, Jean Mayer, Walter P. Metzger, Martin Meyerson, Robert S. Morison, Henry Norr, Talcott Parsons, Bruce L. Payne, Roger Revelle, Philip C. Ritterbush, Neil R. Rudenstine, Edward Joseph Shoben, Jr., John R. Silber, Charles E. Silberman, R. L. Sproull, Kenneth S. Tollett, and George R. Waggoner. The governance of the universities II. Dædalus, 98(4):1092–1154, Fall 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023925.

Ben-David:1969:HEI

[836] Joseph Ben-David, François Bourricaud, Asa Briggs, Mauro Cappelletti, Michel Crozier, A. D. Dunton, Baron John Scott Fulton, Pierre Grappin, Stephen R. Graubard, Walther Killy, Yoichi Maeda, Martin Meyerson, Elting E. Morison, Victor G. Onushkin, Talcott Parsons, Alessandro Pizzorno, Marshall Robinson, Edward Shils, and Solly Zuckerman. Higher education in industrial societies. Dædalus, 98 (4):1157–1223, Fall 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023926.

Anonymous:1969:BMd

[837] Anonymous. Back matter. Dædalus, 98(4):??, Fall 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023927.

Anonymous:1969:FMd

[838] Anonymous. Front matter. Dædalus, 98(4):??, Fall 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023920.

Anonymous:1969:VIb

[839] Anonymous. Volume information. Dædalus, 98(4):??, Fall 1969. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023919.

G:1970:PIE

[840] S. R. G. Preface to the issue “The Embattled University”. Dædalus, 99 (1):v–xvi, Winter 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023930.

Trow:1970:RTM

[841] Martin Trow. Reflections on the transition from mass to universal higher education. Dædalus, 99(1):1–42, Winter 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023931.

Conway:1970:SAC

[842] Jill Conway. Styles of academic culture. Dædalus, 99(1):43–55, Winter 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023932.

Friedenberg:1970:UCO

[843] Edgar Z. Friedenberg. The University community in an open society. Dædalus, 99(1):56–74, Winter 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023933.

Luria:1970:RUI

[844] S. E. Luria and Zella Luria. The role of the University: Ivory tower, service station, or frontier post? Dædalus, 99(1):75–83, Winter 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023934.

Caws:1970:DU

[845] Peter J. Caws. Design for a University. Dædalus, 99(1):84–107, Winter 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023935.

Kerr:1970:GF

[846] Clark Kerr. Governance and functions. Dædalus, 99(1):108–121, Winter 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023936.

Abram:1970:RUN

[847] Morris B. Abram. Reflections on the University in the new revolution. Dædalus, 99(1):122–140, Winter 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023937.

Dungan:1970:HEE

[848] Ralph A. Dungan. Higher education: The effort to adjust. Dædalus, 99 (1):141–153, Winter 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023938.

Erikson:1970:RDC

[849] Erik H. Erikson. Reflections on the dissent of contemporary youth. Dædalus, 99(1):154–176, Winter 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023939.

Hoffmann:1970:PP

[850] Stanley Hoffmann. Participation in perspective? Dædalus, 99(1):177–221, Winter 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023940.

Anonymous:1970:BMa

[851] Anonymous. Back matter. Dædalus, 99(1):??, Winter 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023941.

Anonymous:1970:FMa

[852] Anonymous. Front matter. Dædalus, 99(1):??, Winter 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023929.

Graubard:1970:PIT

[853] Stephen R. Graubard. Preface to the issue “Theory in Humanistic Studies”. Dædalus, 99(2):v–vii, Spring 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023944.

Weil:1970:HST

[854] Eric Weil. Humanistic studies: Their object, methods, and meaning. Dædalus, 99(2):237–255, Spring 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023945.

Bloomfield:1970:TCD

[855] Morton W. Bloomfield. The two cognitive dimensions of the humanities. Dædalus, 99(2):256–267, Spring 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023946.

Frye:1970:CPE

[856] Northrop Frye. The critical path: an essay on the social context of literary criticism. Dædalus, 99(2):268–342, Spring 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023947.

Hirsch:1970:VKH

[857] E. D. Hirsch, Jr. Value and knowledge in the humanities. Dædalus, 99 (2):343–354, Spring 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023948.

Hartman:1970:TLH

[858] Geoffrey Hartman. Toward literary history. Dædalus, 99(2):355–383, Spring 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023949.

DeMan:1970:LHL

[859] Paul De Man. Literary history and literary modernity. Dædalus, 99(2):384– 404, Spring 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20023950.

Miller:1970:WSA

[860] J. Hillis Miller. Williams’ “Spring and All” and the progress of poetry. Dædalus, 99(2):405–434, Spring 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023951.

Pearce:1970:GHN

[861] Roy Harvey Pearce. Gesta humanorum: Notes on the humanist as witness. Dædalus, 99(2):435–450, Spring 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023952.

Hoggart:1970:HSM

[862] Richard Hoggart. Humanistic studies and mass culture. Dædalus, 99 (2):451–472, Spring 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023953.

Genovese:1970:IBP

[863] Eugene D. Genovese. The influence of the black power movement on historical scholarship: Reflections of a white historian. Dædalus, 99 (2):473–494, Spring 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023954.

Parsons:1970:THS

[864] Talcott Parsons. Theory in the humanities and sociology. Dædalus, 99 (2):495–523, Spring 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023955.

Anonymous:1970:BMb

[865] Anonymous. Back matter. Dædalus, 99(2):??, Spring 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023956.

Anonymous:1970:FMb

[866] Anonymous. Front matter. Dædalus, 99(2):??, Spring 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023943.

G:1970:PIR

[867] S. R. G. Preface to the issue “Rights and Responsibilities: The University’s Dilemma”. Dædalus, 99(3):v–xiv, Summer 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023959.

Bundy:1970:WD

[868] McGeorge Bundy. Were those the days? Dædalus, 99(3):531–567, Summer 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023960.

Metzger:1970:CAA

[869] Walter P. Metzger. The crisis of academic authority. Dædalus, 99(3): 568–608, Summer 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023961.

Morison:1970:SAP

[870] Robert S. Morison. Some aspects of policy-making in the American university. Dædalus, 99(3):609–644, Summer 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023962.

Ritterbush:1970:ARW

[871] Philip C. Ritterbush. Adaptive response within the institutional system of higher education and research. Dædalus, 99(3):645–660, Summer 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023963.

Ford:1970:RCL

[872] Franklin L. Ford. Roles and the critique of learning. Dædalus, 99 (3):661–675, Summer 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023964.

Shoben:1970:CCA

[873] Edward Joseph Shoben, Jr. Cultural criticism and the American college. Dædalus, 99(3):676–699, Summer 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023965.

Epstein:1970:SAS

[874] Leon D. Epstein. State authority and state universities. Dædalus, 99 (3):700–712, Summer 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023966.

Anonymous:1970:BMc

[875] Anonymous. Back matter. Dædalus, 99(3):??, Summer 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023967.

Anonymous:1970:FMc

[876] Anonymous. Front matter. Dædalus, 99(3):??, Summer 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023958.

G:1970:PIM

[877] S. R. G. Preface to the issue “The Making of Modern Science: Biographical Studies”. Dædalus, 99(4):v–vi, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023971.

Holton:1970:IIM

[878] Gerald Holton. Introduction to the issue “The Making of Modern Science: Biographical Studies”. Dædalus, 99(4):723–729, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023972.

Erikson:1970:ANI

[879] Erik H. Erikson. Autobiographic notes on the identity crisis. Dædalus, 99(4):730–759, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023973.

Shils:1970:TEI

[880] Edward Shils. Tradition, ecology, and institution in the history of sociology. Dædalus, 99(4):760–825, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023974.

Parsons:1970:BSS

[881] Talcott Parsons. On building social system theory: a personal history. Dædalus, 99(4):826–881, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023975.

Stern:1970:CG

[882] Curt Stern. The continuity of genetics. Dædalus, 99(4):882–908, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023976.

Stent:1970:D

[883] Gunther S. Stent. DNA. Dædalus, 99(4):909–937, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023977.

Olby:1970:FCD

[884] Robert Olby. Francis Crick, DNA, and the Central Dogma. Dædalus, 99(4):938–987, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023978.

Pauling:1970:FYP

[885] Linus Pauling. Fifty years of progress in structural chemistry and molecular biology. Dædalus, 99(4):988–1014, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023979.

Holton:1970:RC

[886] Gerald Holton. The roots of complementarity. Dædalus, 99(4):1015–1055, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023980.

Weinberg:1970:STS

[887] Alvin M. Weinberg. Scientific teams and scientific laboratories. Dædalus, 99(4):1056–1075, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023981.

Wilson:1970:MFA

[888] R. R. Wilson. My fight against team research. Dædalus, 99(4):1076–1087, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20023982.

Baker:1970:CIP

[889] W. O. Baker. Computers as information-processing machines in modern science. Dædalus, 99(4):1088– 1120, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20023983.

Anonymous:1970:BMd

[890] Anonymous. Back matter. Dædalus, 99(4):??, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023984.

Anonymous:1970:FMd

[891] Anonymous. Front matter. Dædalus, 99(4):??, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023970.

Anonymous:1970:VI

[892] Anonymous. Volume information. Dædalus, 99(4):??, Fall 1970. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023969.

G:1971:PIHa

[893] S. R. G. Preface to the issue “Historical Studies Today”. Dædalus, 100(1):v–xiv, Winter 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023987.

LeGoff:1971:PSB

[894] Jacques Le Goff. Is politics still the backbone of history? Dædalus, 100(1):1–19, Winter 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023988.

Hobsbawm:1971:SHH

[895] E. J. Hobsbawm. From social history to the history of society. Dædalus, 100(1):20–45, Winter 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023989.

Stone:1971:P

[896] Lawrence Stone. Prosopography. Dædalus, 100(1):46–79, Winter 1971. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20023990.

Gilbert:1971:IHA

[897] Felix Gilbert. Intellectual history: Its aims and methods. Dædalus, 100 (1):80–97, Winter 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023991.

Schwartz:1971:BDP

[898] Benjamin I. Schwartz. A brief defense of political and intellectual history … with particular reference to non-Western cultures. Dædalus, 100 (1):98–112, Winter 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023992.

Goubert:1971:LH

[899] Pierre Goubert. Local history. Dædalus, 100(1):113–127, Winter 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023993.

Stone:1971:EUS

[900] Lawrence Stone. English and United States local history. Dædalus, 100 (1):128–132, Winter 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023994.

Talbott:1971:HE

[901] John E. Talbott. The history of education. Dædalus, 100(1):133–150, Winter 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023995.

Furet:1971:QH

[902] François Furet. Quantitative history. Dædalus, 100(1):151–167, Winter 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023996.

Finley:1971:AH

[903] M. I. Finley. Archaeology and history. Dædalus, 100(1):168–186, Winter 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20023997.

Manuel:1971:UAP

[904] Frank E. Manuel. The use and abuse of psychology in history. Dædalus, 100 (1):187–213, Winter 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20023998.

Darnton:1971:RWP

[905] Robert Darnton. Reading, writing, and publishing in Eighteenth-Century France: a case study in the sociology of literature. Dædalus, 100 (1):214–256, Winter 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023999.

Anonymous:1971:BMa

[906] Anonymous. Back matter. Dædalus, 100(1):??, Winter 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024000.

Anonymous:1971:FMa

[907] Anonymous. Front matter. Dædalus, 100(1):??, Winter 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20023986.

G:1971:PIHb

[908] S. R. G. Preface to the issue “The Historian and the World of the Twentieth Century”. Dædalus, 100(2):v–xii, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024003.

Kuhn:1971:RBH

[909] Thomas S. Kuhn. The relations between history and history of science. Dædalus, 100(2):271–304, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024004.

Habakkuk:1971:EHE

[910] John Habakkuk. Economic history and economic theory. Dædalus, 100 (2):305–322, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024005.

Craig:1971:PH

[911] Gordon A. Craig. Political history. Dædalus, 100(2):323–338, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024006.

Schlesinger:1971:HP

[912] Arthur Schlesinger, Jr. The historian as participant. Dædalus, 100 (2):339–358, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024007.

Thernstrom:1971:RNU

[913] Stephan Thernstrom. Reflections on the new urban history. Dædalus, 100 (2):359–375, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024008.

Paret:1971:HW

[914] Peter Paret. The history of war. Dædalus, 100(2):376–396, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024009.

Ladurie:1971:QCE

[915] Emmanuel Le Roy Ladurie and Paul Dumont. Quantitative and cartographical exploitation of French military archives, 1819–1826. Dædalus, 100 (2):397–441, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024010.

Vansina:1971:OUT

[916] Jan Vansina. Once upon a time: Oral traditions as history in Africa. Dædalus, 100(2):442–468, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024011.

Berengo:1971:IHS

[917] Marino Berengo. Italian historical scholarship since the Fascist Era. Dædalus, 100(2):469–484, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024012.

Mommsen:1971:HST

[918] Hans Mommsen. Historical scholarship in transition: The situation in the Federal Republic of Germany. Dædalus, 100(2):485–508, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024013.

Hanham:1971:CW

[919] H. J. Hanham. Clio’s weapons. Dædalus, 100(2):509–519, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024014.

Gilbert:1971:PS

[920] Felix Gilbert. Post scriptum. Dædalus, 100(2):520–530, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024015.

Anonymous:1971:BMb

[921] Anonymous. Back matter. Dædalus, 100(2):??, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024016.

Anonymous:1971:FMb

[922] Anonymous. Front matter. Dædalus, 100(2):??, Spring 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024002.

G:1971:PIF

[923] S. R. G. Preface to the issue “The Future of the Black Colleges”. Dædalus, 100(3):v–viii, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024019.

Brimmer:1971:EOF

[924] Andrew F. Brimmer. The economic outlook and the future of the Negro College. Dædalus, 100(3):539–572, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024020.

Bullock:1971:BCN

[925] Henry Allen Bullock. The black College and the new black awareness. Dædalus, 100(3):573–602, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024021.

Lincoln:1971:NCC

[926] C. Eric Lincoln. The Negro colleges and cultural change. Dædalus, 100 (3):603–629, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024022.

Henderson:1971:NCF

[927] Vivian W. Henderson. Negro colleges face the future. Dædalus, 100 (3):630–646, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024023.

Trent:1971:FRN

[928] William J. Trent, Jr. The future role of the Negro College and its financing. Dædalus, 100(3):647–659, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024024.

Nabrit:1971:RFB

[929] S. M. Nabrit. Reflections on the future of black colleges. Dædalus, 100 (3):660–677, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024025.

Winston:1971:TBD

[930] Michael R. Winston. Through the back door: Academic racism and the Negro scholar in historical perspective. Dædalus, 100(3):678–719, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024026.

Harris:1971:NCC

[931] Patricia Roberts Harris. The Negro College and its community. Dædalus, 100(3):720–731, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024027.

Jones:1971:RBC

[932] Mack H. Jones. The responsibility of the black College to the black community: Then and now. Dædalus, 100 (3):732–744, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024028.

Blake:1971:FLR

[933] Elias Blake, Jr. Future leadership roles for predominantly black colleges and universities in American higher education. Dædalus, 100(3):745–771, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024029.

Harper:1971:LSB

[934] Conrad K. Harper. The legal status of the black College. Dædalus, 100 (3):772–782, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024030.

Godwin:1971:SSG

[935] Winfred L. Godwin. Southern State governments and higher education for Negroes. Dædalus, 100(3):783–797, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20024031.

Johnson:1971:BCS

[936] Tobe Johnson. The black College as system. Dædalus, 100(3):798–812, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20024032.

Pettigrew:1971:RWB

[937] Thomas F. Pettigrew. The role of whites in the black College of the future. Dædalus, 100(3):813–832, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024033.

Drake:1971:BUA

[938] Clair Drake. The black University in the American social order. Dædalus, 100(3):833–897, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024034.

Anonymous:1971:BMc

[939] Anonymous. Back matter. Dædalus, 100(3):??, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024035.

Anonymous:1971:FMc

[940] Anonymous. Front matter. Dædalus, 100(3):??, Summer 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024018.

G:1971:PIT

[941] S. R. G. Preface to the issue “Twelve to Sixteen: Early Adolescence”. Dædalus, 100(4):v–vii, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024039.

Tanner:1971:STI

[942] J. M. Tanner. Sequence, tempo, and individual variation in the growth and development of Boys and girls aged twelve to sixteen. Dædalus, 100(4):907–930, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024040.

Gordon:1971:SCE

[943] Chad Gordon. Social characteristics of early adolescence. Dædalus, 100(4):931–960, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024041.

Blos:1971:CAL

[944] Peter Blos. The child analyst looks at the Young adolescent. Dædalus, 100(4):961–978, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024042.

Bakan:1971:AAI

[945] David Bakan. Adolescence in America: From idea to social fact. Dædalus, 100(4):979–995, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024043.

Kagan:1971:CEA

[946] Jerome Kagan. A conception of early adolescence. Dædalus, 100(4):997– 1012, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20024044.

Adelson:1971:PIY

[947] Joseph Adelson. The political imagination of the Young adolescent. Dædalus, 100(4):1013–1050, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024045.

Kohlberg:1971:APD

[948] Lawrence Kohlberg and Carol Gilligan. The adolescent as a philosopher: The discovery of the self in a postconventional world. Dædalus, 100(4):1051–1086, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024046.

Martin:1971:REA

[949] Edward C. Martin. Reflections on the early adolescent in school. Dædalus, 100(4):1087–1103, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024047.

Conger:1971:WTN

[950] John Janeway Conger. A world they never knew: The family and social change. Dædalus, 100(4):1105–1138, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024048.

Coles:1971:WY

[951] Robert Coles. The weather of the years. Dædalus, 100(4):1139–1157, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024049.

LaFarge:1971:UA

[952] Phyllis La Farge. An uptight adolescence. Dædalus, 100(4):1159–1175, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024050.

Cottle:1971:CA

[953] Thomas J. Cottle. The connections of adolescence. Dædalus, 100 (4):1177–1219, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024051.

Anonymous:1971:BMd

[954] Anonymous. Back matter. Dædalus, 100(4):??, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024052.

Anonymous:1971:FMd

[955] Anonymous. Front matter. Dædalus, 100(4):??, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024038.

Anonymous:1971:VI

[956] Anonymous. Volume information. Dædalus, 100(4):??, Fall 1971. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024037.

G:1972:PIM

[957] S. R. G. Preface to the issue “Myth, Symbol, and Culture”. Dædalus, 101 (1):v–viii, Winter 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024055.

Geertz:1972:DPN

[958] Clifford Geertz. Deep play: Notes on the Balinese cockfight. Dædalus, 101(1):1–37, Winter 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024056.

Fernandez:1972:PPB

[959] James W. Fernandez. Persuasions and performances: Of the beast in every body … and the metaphors of everyman. Dædalus, 101(1):39–60, Winter 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024057.

Douglas:1972:DM

[960] Mary Douglas. Deciphering a meal. Dædalus, 101(1):61–81, Winter 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024058.

Manuel:1972:SNH

[961] Frank E. Manuel and Fritzie P. Manuel. Sketch for a natural history of paradise. Dædalus, 101(1):83–128, Winter 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024059.

Shklar:1972:SG

[962] Judith N. Shklar. Subversive genealogies. Dædalus, 101(1):129–154, Winter 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024060.

Brower:1972:VVT

[963] Reuben A. Brower. Visual and verbal translation of myth: Neptune in Virgil, Rubens, Dryden. Dædalus, 101 (1):155–182, Winter 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024061.

Marcus:1972:LSP

[964] Steven Marcus. Language into structure: Pickwick revisited. Dædalus, 101 (1):183–202, Winter 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024062.

Adams:1972:SVP

[965] Robert M. Adams. The sense of verification: Pragmatic commonplaces about literary criticism. Dædalus, 101 (1):203–214, Winter 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024063.

Anonymous:1972:BMa

[966] Anonymous. Back matter. Dædalus, 101(1):??, Winter 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024064.

Anonymous:1972:FMa

[967] Anonymous. Front matter. Dædalus, 101(1):??, Winter 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024054.

G:1972:PIIa

[968] S. R. G. Preface to the issue “Intellectuals and Tradition”. Dædalus, 101(2):v–xv, Spring 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024067.

Eisenstadt:1972:IT

[969] S. N. Eisenstadt. Intellectuals and tradition. Dædalus, 101(2):1–19, Spring 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024068.

Shils:1972:ITT

[970] Edward Shils. Intellectuals, tradition, and the traditions of intellectuals: Some preliminary considerations. Dædalus, 101(2):21–34, Spring 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024069.

Wakeman:1972:PAI

[971] Frederic Wakeman, Jr. The price of autonomy: Intellectuals in ming and ch’ing politics. Dædalus, 101(2):35–70, Spring 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024070.

Schwartz:1972:LTV

[972] Benjamin I. Schwartz. The limits of “Tradition versus Modernity” as categories of explanation: The case of the Chinese intellectuals. Dædalus, 101(2):71–88, Spring 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024071.

Bellah:1972:ISJ

[973] Robert N. Bellah. Intellectual and society in Japan. Dædalus, 101(2):89–115, Spring 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024072.

Confino:1972:IIT

[974] Michael Confino. On intellectuals and intellectual traditions in Eighteenth- and Nineteenth-Century Russia. Dædalus, 101(2):117–149, Spring 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024073.

ODea:1972:RIC

[975] Thomas F. O’Dea. The role of the intellectual in the Catholic tradition. Dædalus, 101(2):151–189, Spring 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024074.

Anonymous:1972:BMb

[976] Anonymous. Back matter. Dædalus, 101(2):??, Spring 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024075.

Anonymous:1972:FMb

[977] Anonymous. Front matter. Dædalus, 101(2):??, Spring 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024066.

G:1972:PIIb

[978] S. R. G. Preface to the issue “Intellectuals and Change”. Dædalus, 101(3):v–xi, Summer 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024078.

Kolakowski:1972:IAI

[979] Leszek Kolakowski. Intellectuals against intellect. Dædalus, 101(3):1– 15, Summer 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20024079.

Milson:1972:MMI

[980] Menahem Milson. Medieval and modern intellectual traditions in the Arab world. Dædalus, 101(3):17–37, Summer 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024080.

Keddie:1972:IMM

[981] Nikki R. Keddie. Intellectuals in the modern Middle East: a brief historical consideration. Dædalus, 101 (3):39–57, Summer 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024081.

Linz:1972:IRS

[982] Juan J. Linz. Intellectual roles in Sixteenth- and Seventeenth-Century Spain. Dædalus, 101(3):59–108, Summer 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024082.

Bourricaud:1972:AA

[983] François Bourricaud. The adventures of Ariel. Dædalus, 101(3):109–136, Summer 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024083.

Lipset:1972:ICR

[984] Seymour Martin Lipset and Richard B. Dobson. The intellectual as critic and rebel: With special reference to the United States and the Soviet union. Dædalus, 101(3):137–198, Summer 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024084.

Conway:1972:IAV

[985] Jill Conway. Intellectuals in America: Varieties of accommodation and conflict. Dædalus, 101(3):199–205, Summer 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024085.

Malia:1972:IAC

[986] Martin E. Malia. The intellectuals: Adversaries or clerisy? Dædalus, 101 (3):206–216, Summer 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024086.

Anonymous:1972:BMc

[987] Anonymous. Back matter. Dædalus, 101(3):??, Summer 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024087.

Anonymous:1972:FMc

[988] Anonymous. Front matter. Dædalus, 101(3):??, Summer 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024077.

G:1972:PIS

[989] S. R. G. Preface to the issue “As Others See the United States”. Dædalus, 101(4):v–ix, Fall 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024091.

Aron:1972:RNF

[990] Raymond Aron. Richard Nixon and the future of American foreign policy. Dædalus, 101(4):1–24, Fall 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024092.

Bhagwati:1972:USN

[991] Jagdish Bhagwati. The United States in the Nixon era: The end of innocence. Dædalus, 101(4):25–47, Fall 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024093.

Albonetti:1972:USE

[992] Achille Albonetti. The United States: Economic or political challenge? Dædalus, 101(4):49–66, Fall 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024094.

Paz:1972:EG

[993] Octavio Paz. Eroticism and gastrosophy. Dædalus, 101(4):67–85, Fall 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024095.

Kriegel:1972:CME

[994] Annie Kriegel. Consistent misapprehension: European views of America and their logic. Dædalus, 101(4):87– 102, Fall 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024096.

Grine:1972:AAI

[995] Omar Grine. America against itself: a case of democratic anarchism? Dædalus, 101(4):103–117, Fall 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024097.

Lafer:1972:USV

[996] Celso Lafer. The United States: a view from Brazil. Dædalus, 101(4):119–128, Fall 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024098.

Maddox:1972:ASE

[997] John Maddox. American science: Endless search for objectives. Dædalus, 101(4):129–140, Fall 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024099.

Burrage:1972:DMC

[998] Michael Burrage. Democracy and the mystery of the crafts: Observations on work relationships in America and Britain. Dædalus, 101(4):141–162, Fall 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024100.

Doxiadis:1972:TLA

[999] C. A. Doxiadis. Three letters to an American. Dædalus, 101(4):163–183, Fall 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024101.

Llewelyn-Davies:1972:ACT

[1000] Richard Llewelyn-Davies. The American city through English eyes. Dædalus, 101(4):185–193, Fall 1972. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20024102.

Anonymous:1972:BMd

[1001] Anonymous. Back matter. Dædalus, 101(4):??, Fall 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024103.

Anonymous:1972:FMd

[1002] Anonymous. Front matter. Dædalus, 101(4):??, Fall 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024090.

Anonymous:1972:VI

[1003] Anonymous. Volume information. Dædalus, 101(4):??, Fall 1972. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024089.

G:1973:PIP

[1004] S. R. G. Preface to the issue “Post-Traditional Societies”. Dædalus, 102 (1):v–vii, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024106.

Eisenstadt:1973:PTS

[1005] S. N. Eisenstadt. Post-traditional societies and the continuity and reconstruction of tradition. Dædalus, 102(1):1–27, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024107.

Leach:1973:BPC

[1006] Edmund Leach. Buddhism in the postcolonial political order in Burma and Ceylon. Dædalus, 102(1):29–54, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024108.

Tambiah:1973:PTT

[1007] S. J. Tambiah. The persistence and transformation of tradition in Southeast Asia, with special reference to Thailand. Dædalus, 102(1):55–84, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024109.

Bechert:1973:SSS

[1008] Heinz Bechert. Sangha, state, society, “Nation”: Persistence of traditions in “Post-Traditional” Buddhist societies. Dædalus, 102(1):85–95, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024110.

Heesterman:1973:IIC

[1009] J. C. Heesterman. India and the inner conflict of tradition. Dædalus, 102(1):97–113, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024111.

Nandy:1973:MUP

[1010] Ashis Nandy. The making and unmaking of political cultures in India. Dædalus, 102(1):115–137, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024112.

Yalman:1973:SOS

[1011] Nur Yalman. Some observations on secularism in Islam: The cultural revolution in Turkey. Dædalus, 102 (1):139–168, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024113.

Mardin:1973:CPR

[1012] Serif Mardin. Center-periphery relations: a key to Turkish politics? Dædalus, 102(1):169–190, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024114.

Gellner:1973:PTF

[1013] Ernest Gellner. Post-traditional forms in Islam: The turf and trade, and votes and peanuts. Dædalus, 102 (1):191–206, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024115.

Hermassi:1973:PTM

[1014] Elbaki Hermassi. Political traditions of the Maghrib. Dædalus, 102(1):207–224, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024116.

Zghal:1973:RTP

[1015] Abdelkader Zghal. The reactivation of tradition in a post-traditional society. Dædalus, 102(1):225–237, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024117.

Berque:1973:TIM

[1016] Jacques Berque. Tradition and innovation in the Maghrib. Dædalus, 102 (1):239–250, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024118.

Anonymous:1973:BMa

[1017] Anonymous. Back matter. Dædalus, 102(1):??, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024119.

Anonymous:1973:FMa

[1018] Anonymous. Front matter. Dædalus, 102(1):??, Winter 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024105.

Anonymous:1973:PIS

[1019] Anonymous. Preface to the issue “The Search for Knowledge”. Dædalus, 102(2):v–xi, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024122.

Gombrich:1973:RHI

[1020] E. H. Gombrich. Research in the humanities: Ideals and idols. Dædalus, 102(2):1–10, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024123.

Frye:1973:SAW

[1021] Northrop Frye. The search for acceptable words. Dædalus, 102(2):11–26, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024124.

Weil:1973:SH

[1022] Eric Weil. Supporting the humanities. Dædalus, 102(2):27–38, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024125.

Ben-David:1973:HOR

[1023] Joseph Ben-David. How to organize research in the social sciences. Dædalus, 102(2):39–51, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024126.

Dworkin:1973:LR

[1024] Ronald Dworkin. Legal research. Dædalus, 102(2):53–64, Spring 1973. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20024127.

Johnson:1973:NSE

[1025] Harry G. Johnson. National styles in economic research: The United States, the United Kingdom, Canada, and various European countries. Dædalus, 102(2):65–74, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024128.

Luria:1973:RSA

[1026] S. E. Luria. On research styles and Allied matters. Dædalus, 102(2):75–84, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024129.

Eccles:1973:DSS

[1027] John C. Eccles. The discipline of science with special reference to the neurosciences. Dædalus, 102(2):85–99, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024130.

Robinson:1973:WUD

[1028] David Z. Robinson. Will the University decline as the center for scientific research? Dædalus, 102 (2):101–110, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024131.

Adams:1973:SPB

[1029] J. B. Adams. Some problems of a big science. Dædalus, 102(2):111–124, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024132.

Brooks:1973:KAD

[1030] Harvey Brooks. Knowledge and action: The dilemma of science policy in the ’70’s. Dædalus, 102(2):125–143, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024133.

Branscomb:1973:CUR

[1031] Lewis M. Branscomb. Conducting and using research. Dædalus, 102 (2):145–152, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024134.

Townes:1973:DCB

[1032] C. H. Townes. Differentiation and competition between University and other research laboratories in the United States. Dædalus, 102(2):153–165, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024135.

Kapitza:1973:BFO

[1033] Peter Kapitza. Basic factors in the organization of science and how they are handled in the U.S.S.R. Dædalus, 102(2):167–176, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024136.

Anonymous:1973:BMb

[1034] Anonymous. Back matter. Dædalus, 102(2):??, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024137.

Anonymous:1973:FMb

[1035] Anonymous. Front matter. Dædalus, 102(2):??, Spring 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024121.

G:1973:PIL

[1036] S. R. G. Preface to the issue “Language as a Human Problem”. Dædalus, 102(3):v, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024140.

Haugen:1973:PL

[1037] Einar Haugen. The problems of language. Dædalus, 102(3):1–4, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024141.

Bloomfield:1973:SL

[1038] Morton Bloomfield. The study of language. Dædalus, 102(3):5–13, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024142.

Moulton:1973:NL

[1039] William G. Moulton. The nature of language. Dædalus, 102(3):17–35, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024143.

Ferguson:1973:LPV

[1040] Charles Ferguson. Language problems of variation and repertoire. Dædalus, 102(3):37–46, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024144.

Haugen:1973:CB

[1041] Einar Haugen. The curse of Babel. Dædalus, 102(3):47–57, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024145.

Hymes:1973:SLO

[1042] Dell Hymes. Speech and language: On the origins and foundations of inequality among speakers. Dædalus, 102(3):59–85, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024146.

Teeter:1973:LA

[1043] Karl V. Teeter. Linguistics and anthropology. Dædalus, 102(3):87–98, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024147.

Watkins:1973:LH

[1044] Calvert Watkins. Language and its history. Dædalus, 102(3):99–111, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024148.

Lenneberg:1973:NL

[1045] Eric H. Lenneberg. The neurology of language. Dædalus, 102(3): 115–133, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024149.

Cazden:1973:PEL

[1046] Courtney B. Cazden. Problems for education: Language as curriculum content and learning environment. Dædalus, 102(3):135–148, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024150.

Strevens:1973:SLL

[1047] Peter Strevens. Second language learning. Dædalus, 102(3):149–160, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024151.

Wanner:1973:DWU

[1048] Eric Wanner. Do we understand sentences from the outside-in or from the inside-out? Dædalus, 102(3): 163–183, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024152.

Keenan:1973:LL

[1049] Edward L. Keenan. Logic and language. Dædalus, 102(3):185–194, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024153.

Langendoen:1973:PLT

[1050] D. Terence Langendoen. The problem of linguistic theory in relation to language behavior: a tribute and reply to Paul Goodman. Dædalus, 102 (3):195–201, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024154.

Hays:1973:LIR

[1051] David G. Hays. Language and interpersonal relationships. Dædalus, 102 (3):203–216, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024155.

Kay:1973:ATN

[1052] Martin Kay. Automatic translation of natural languages. Dædalus, 102 (3):217–230, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024156.

Kiparsky:1973:RLT

[1053] Paul Kiparsky. The role of linguistics in a theory of poetry. Dædalus, 102 (3):231–244, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024157.

Anonymous:1973:BMc

[1054] Anonymous. Back matter. Dædalus, 102(3):??, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024158.

Anonymous:1973:FMc

[1055] Anonymous. Front matter. Dædalus, 102(3):??, Summer 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024139.

G:1973:PIN

[1056] S. R. G. Preface to the issue “The No-Growth Society”. Dædalus, 102(4):v– vi, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024162.

Olson:1973:I

[1057] Mancur Olson. Introduction. Dædalus, 102(4):1–13, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024163.

Davis:1973:ZPG

[1058] Kingsley Davis. Zero population growth: The goal and the means. Dædalus, 102(4):15–30, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024164.

Holdren:1973:PAP

[1059] John P. Holdren. Population and the American predicament: The case against complacency. Dædalus, 102(4): 31–43, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024165.

Ryder:1973:TCZ

[1060] Norman B. Ryder. Two cheers for ZPG. Dædalus, 102(4):45–62, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024166.

Mishan:1973:IBD

[1061] E. J. Mishan. Ills, bads, and disamenities: The wages of growth. Dædalus, 102(4):63–87, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024167.

Boulding:1973:SSS

[1062] Kenneth E. Boulding. The shadow of the stationary state. Dædalus, 102(4):89–101, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024168.

Zeckhauser:1973:RG

[1063] Richard Zeckhauser. The risks of growth. Dædalus, 102(4):103–118, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024169.

Roberts:1973:REG

[1064] Marc J. Roberts. On reforming economic growth. Dædalus, 102(4):119– 137, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024170.

Brooks:1973:TZG

[1065] Harvey Brooks. The technology of zero growth. Dædalus, 102(4):139–152, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024171.

Brown:1973:RCP

[1066] Lester Brown. Rich countries and poor in a finite, interdependent world. Dædalus, 102(4):153–164, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024172.

Johnson:1973:SPB

[1067] Willard R. Johnson. Should the poor buy no growth? Dædalus, 102(4):165–189, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024173.

Alonso:1973:UZP

[1068] William Alonso. Urban zero population growth. Dædalus, 102(4):191–206, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024174.

McKean:1973:GVN

[1069] Roland N. McKean. Growth vs. no growth: an evaluation. Dædalus, 102(4):207–227, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024175.

Olson:1973:E

[1070] Mancur Olson, Hans H. Landsberg, and Joseph L. Fisher. Epilogue. Dædalus, 102(4):229–241, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024176.

Anonymous:1973:BMd

[1071] Anonymous. Back matter. Dædalus, 102(4):??, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024177.

Anonymous:1973:FMd

[1072] Anonymous. Front matter. Dædalus, 102(4):??, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024161.

Anonymous:1973:VI

[1073] Anonymous. Volume information. Dædalus, 102(4):??, Fall 1973. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024160.

G:1974:PIT

[1074] S. R. G. Preface to the issue “Twentieth-Century Classics Revisited”. Dædalus, 103(1):v–xvii, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024180.

Frye:1974:DWO

[1075] Northrop Frye. “The Decline of the West” by Oswald Spengler. Dædalus, 103(1):1–13, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024181.

Bowersock:1974:SEH

[1076] G. W. Bowersock. “The Social and Economic History of the Roman Empire” by Michael Ivanovitch Rostovtzeff. Dædalus, 103(1):15–23, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024182.

Brown:1974:MCH

[1077] Peter Brown. “Mohammed and Charlemagne” by Henri Pirenne. Dædalus, 103(1):25–33, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024183.

Bouwsma:1974:WMA

[1078] William J. Bouwsma. “The Waning of the Middle Ages” by Johan Huizinga. Dædalus, 103(1):35–43, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024184.

Kindleberger:1974:GTK

[1079] Charles P. Kindleberger. “The Great Transformation” by Karl Polanyi. Dædalus, 103(1):45–52, Winter 1974. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20024185.

Vernon:1974:TLC

[1080] Raymond Vernon. “The Theory of the Leisure Class” by Thorstein Veblen. Dædalus, 103(1):53–57, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024186.

Shklar:1974:EHA

[1081] Judith N. Shklar. “The Education of Henry Adams” by Henry Adams. Dædalus, 103(1):59–66, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024187.

Degler:1974:WFH

[1082] Carl N. Degler. “Woman as Force in History” by Mary Beard. Dædalus, 103(1):67–73, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024188.

Clive:1974:RMJ

[1083] Geoffrey Clive. “Revolt of the Masses” by José Ortega y Gassett. Dædalus, 103(1):75–82, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024189.

Shils:1974:IUK

[1084] Edward Shils. “Ideology and Utopia” by Karl Mannheim. Dædalus, 103 (1):83–89, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024190.

Parsons:1974:IDS

[1085] Talcott Parsons. “The Interpretation of Dreams” by Sigmund Freud. Dædalus, 103(1):91–96, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024191.

Coles:1974:NDM

[1086] Robert Coles. “The Nature and Destiny of Man” by Reinhold Niebuhr. Dædalus, 103(1):97–104, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20024192.

Chapman:1974:FRA

[1087] Carleton B. Chapman. “The Flexner Report” by Abraham Flexner. Dædalus, 103(1):105–117, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024193.

Ackerman:1974:LMM

[1088] Bruce A. Ackerman. “Law and the Modern Mind” by Jerome Frank. Dædalus, 103(1):119–130, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024194.

Putnam:1974:TPP

[1089] Michael C. J. Putnam. “Three Philosophical Poets” by George Santayana. Dædalus, 103(1):131–140, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024195.

Carberg:1974:VWB

[1090] Joan S. Carberg. “A Vision” by William Butler Yeats. Dædalus, 103 (1):141–156, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024196.

Anonymous:1974:BMa

[1091] Anonymous. Back matter. Dædalus, 103(1):??, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024197.

Anonymous:1974:FMa

[1092] Anonymous. Front matter. Dædalus, 103(1):??, Winter 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024179.

G:1974:PISa

[1093] S. R. G. Preface to the issue “Slavery, Colonialism, and Racism”. Dædalus, 103(2):v–vi, Spring 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024200.

Davis:1974:SPW

[1094] David Brion Davis. Slavery and the post-world war II historians. Dædalus, 103(2):1–16, Spring 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024201.

Curtin:1974:BEC

[1095] Philip D. Curtin. The black experience of colonialism and imperialism. Dædalus, 103(2):17–29, Spring 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024202.

Banton:1974:TPS

[1096] Michael Banton. 1960: a turning point in the study of race relations. Dædalus, 103(2):31–44, Spring 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024203.

Mintz:1974:CR

[1097] Sidney W. Mintz. The Caribbean region. Dædalus, 103(2):45–71, Spring 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024204.

Brathwaite:1974:APC

[1098] Edward Kamau Brathwaite. The African presence in Caribbean literature. Dædalus, 103(2):73–109, Spring 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024205.

Bastide:1974:PSA

[1099] Roger Bastide. The present status of Afro–American research in Latin America. Dædalus, 103(2):111–123, Spring 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024206.

Ajayi:1974:BAH

[1100] J. F. Ade. Ajayi and E. J. Alagoa. Black Africa: The historians’ perspective. Dædalus, 103(2):125–134, Spring 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024207.

Wastberg:1974:TAL

[1101] Per Wästberg. Themes in African literature today. Dædalus, 103(2):135–150, Spring 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024208.

Nketia:1974:MHA

[1102] J. H. Kwabena Nketia. The musical heritage of Africa. Dædalus, 103 (2):151–161, Spring 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024209.

Quarles:1974:BHU

[1103] Benjamin Quarles. Black history unbound. Dædalus, 103(2):163–178, Spring 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024210.

Sowell:1974:PBS

[1104] Thomas Sowell. The plight of black students in the United States. Dædalus, 103(2):179–196, Spring 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024211.

Anonymous:1974:BMb

[1105] Anonymous. Back matter. Dædalus, 103(2):??, Spring 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024212.

Anonymous:1974:FMb

[1106] Anonymous. Front matter. Dædalus, 103(2):??, Spring 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024199.

G:1974:PISb

[1107] S. R. G. Preface to the issue “Science and Its Public: The Changing Relationship”. Dædalus, 103(3):v–vii, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024215.

Shils:1974:FUL

[1108] Edward Shils. Faith, utility, and the legitimacy of science. Dædalus, 103(3):1–15, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024216.

Roszak:1974:MTS

[1109] Theodore Roszak. The monster and the titan: Science, knowledge, and gnosis. Dædalus, 103(3):17–32, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024217.

Weinberg:1974:RWS

[1110] Steven Weinberg. Reflectons of a working scientist. Dædalus, 103(3):33–45, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024218.

Roberts:1974:NCS

[1111] Marc J. Roberts. On the nature and condition of social science. Dædalus, 103(3):47–64, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024219.

Holton:1974:BCB

[1112] Gerald Holton. On being caught between Dionysians and Apollonians. Dædalus, 103(3):65–81, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024220.

Mayer:1974:AIE

[1113] André Mayer and Jean Mayer. Agriculture, the Island empire. Dædalus, 103(3):83–95, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024221.

Price:1974:MIL

[1114] Don K. Price. Money and influence: The links of science to public policy. Dædalus, 103(3):97–113, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024222.

Beckler:1974:PLS

[1115] David Z. Beckler. The precarious life of science in the white house. Dædalus, 103(3):115–134, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024223.

Daddario:1974:SPR

[1116] Emilio Q. Daddario. Science policy:

Relationships are the key. Dædalus, 103 (3):135–142, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024224.

Rose:1974:NLO

[1117] David J. Rose. New laboratories for old. Dædalus, 103(3):143–155, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024225.

McCormmach:1974:ASW

[1118] Russell McCormmach. On academic scientists in Wilhelmian Germany. Dædalus, 103(3):157–171, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024226.

Moore:1974:CC

[1119] John A. Moore. Creationism in California. Dædalus, 103(3):173–189, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024227.

Etzioni:1974:PAS

[1120] Amitai Etzioni and Clyde Nunn. The public appreciation of science in contemporary America. Dædalus, 103 (3):191–205, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024228.

Perlman:1974:SMM

[1121] David Perlman. Science and the mass media. Dædalus, 103(3):207–222, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024229.

Anonymous:1974:BMc

[1122] Anonymous. Back matter. Dædalus, 103(3):??, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024230.

Anonymous:1974:FMc

[1123] Anonymous. Front matter. Dædalus, 103(3):??, Summer 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024214.

G:1974:PIA

[1124] S. R. G. Preface to the issue “American Higher Education: Toward an Uncertain Future”. Volume I. Dædalus, 103(4):v–x, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024234.

Graubard:1974:THE

[1125] Stephen R. Graubard. Thoughts on higher educational purposes and goals: A memorandum. Dædalus, 103(4):1– 11, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024235.

Gilbert:1974:RHE

[1126] Felix Gilbert. Reflections on “Higher Education”. Dædalus, 103(4):12–17, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024236.

Caws:1974:II

[1127] Peter Caws. Instruction and inquiry. Dædalus, 103(4):18–24, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024237.

Frankel:1974:RWM

[1128] Charles Frankel. Reflections on a wornout model. Dædalus, 103(4):25–32, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024238.

Woodward:1974:EAP

[1129] C. Vann Woodward. The erosion of academic privileges and immunities. Dædalus, 103(4):33–37, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024239.

Quine:1974:PP

[1130] Willard Quine. Paradoxes of plenty. Dædalus, 103(4):38–40, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024240.

Keeney:1974:CWR

[1131] Barnaby C. Keeney. Credibility, will, and response. Dædalus, 103(4):41–43, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024241.

Gerschenkron:1974:LE

[1132] Alexander Gerschenkron. The legacies of evil. Dædalus, 103(4):44–49, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024242.

Perkins:1974:UOG

[1133] James A. Perkins. The University: Old ghosts and new. Dædalus, 103(4):50–53, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024243.

Adelson:1974:LB

[1134] Joseph Adelson. Looking back. Dædalus, 103(4):54–57, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024244.

Bloom:1974:FU

[1135] Allan Bloom. The failure of the University. Dædalus, 103(4):58–66, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024245.

Hesburgh:1974:ERV

[1136] Theodore M. Hesburgh. The “Events”: a retrospective view. Dædalus, 103(4):

67–71, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024246.

Henderson:1974:BCH

[1137] Vivian W. Henderson. Blacks and change in higher education. Dædalus, 103(4):72–79, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024247.

Ulam:1974:WDW

[1138] Adam B. Ulam. Where do we go from here? Dædalus, 103(4):80–84, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024248.

Apter:1974:ETR

[1139] David E. Apter. An epitaph for two revolutions that failed. Dædalus, 103(4):85–103, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024249.

Marty:1974:KEC

[1140] Martin E. Marty. Knowledge elites and counter-elites. Dædalus, 103(4):104– 109, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024250.

Bellah:1974:NRC

[1141] Robert N. Bellah. The new religious consciousness and the secular University. Dædalus, 103(4):110–115, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024251.

Manuel:1974:CVN

[1142] Frank E. Manuel. Clio vanishes from New York City. Dædalus, 103(4):116– 127, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024252.

Bloomfield:1974:EH

[1143] Morton W. Bloomfield. Elitism in the humanities. Dædalus, 103(4):128–137, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024253.

Else:1974:SIT

[1144] Gerald F. Else. Some ill-tempered reflections on the present state of higher education in the United States. Dædalus, 103(4):138–142, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024254.

Craig:1974:GSS

[1145] Gordon A. Craig. Green stamp or structured undergraduate education? Dædalus, 103(4):143–147, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024255.

Muller:1974:HEH

[1146] Steven Muller. Higher education or higher skilling? Dædalus, 103(4):148– 158, Fall 1974. CODEN DAEDAU.

ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024256.

Bok:1974:PUE

[1147] Derek Bok. On the purposes of undergraduate education. Dædalus, 103(4):159–172, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024257.

Meyerson:1974:CEU

[1148] Martin Meyerson. Civilizing education: Uniting liberal and professional learning. Dædalus, 103(4):173–179, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024258.

Kaysen:1974:WSU

[1149] Carl Kaysen. What should undergraduate education do? Dædalus, 103(4):180–185, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024259.

Almond:1974:SRC

[1150] Gabriel A. Almond. Some reflections on the crisis of the University curriculum. Dædalus, 103(4):186–187, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024260.

Allison:1974:EEE

[1151] Elisabeth K. Allison. The evaluation of educational experience. Dædalus, 103(4):188–195, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024261.

Skinner:1974:DHE

[1152] B. F. Skinner. Designing higher education. Dædalus, 103(4):196–202, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024262.

Novak:1974:OM

[1153] Michael Novak. The one and the many. Dædalus, 103(4):203–211, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024263.

Winston:1974:RSU

[1154] Michael R. Winston. Reflections on student unrest, institutional response, and curricular change. Dædalus, 103 (4):212–216, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024264.

Sennett:1974:AU

[1155] Richard Sennett. The artist and the University. Dædalus, 103(4):217–220, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024265.

Washburn:1974:EE

[1156] S. L. Washburn. Evolution and education. Dædalus, 103(4):221–228, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024266.

Ong:1974:ASA

[1157] Walter J. Ong. Agonistic structures in academia: Past to present. Dædalus, 103(4):229–238, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024267.

Conway:1974:CWS

[1158] Jill K. Conway. Coeducation and Women’s studies: Two approaches to the question of Woman’s place in the contemporary University. Dædalus, 103(4):239–249, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024268.

Lane:1974:SAE

[1159] Robert E. Lane. The self-analysis of educational belief systems. Dædalus, 103(4):250–257, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024269.

Light:1974:TAF

[1160] Donald Light, Jr. Thinking about faculty. Dædalus, 103(4):258–264, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024270.

Franklin:1974:OGS

[1161] John Hope Franklin. On the “Oversupply” of graduate students. Dædalus, 103(4):265–268, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024271.

Parsons:1974:SCA

[1162] Talcott Parsons. Stability and change in the American university. Dædalus, 103(4):269–277, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024272.

Kagan:1974:UU

[1163] Jerome Kagan. On the uses of the University. Dædalus, 103(4):278–281, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024273.

Mattfeld:1974:LEC

[1164] Jacquelyn Anderson Mattfeld. Liberal education in contemporary American society. Dædalus, 103(4):282–287, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024274.

Adams:1974:OWR

[1165] Thomas Boylston Adams. One wide river to cross. Dædalus, 103(4):288– 296, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20024275.

Farnsworth:1974:SVC

[1166] Dana L. Farnsworth. Social values in College and University. Dædalus, 103(4):297–301, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024276.

Rhodes:1974:D

[1167] Joseph Rhodes, Jr. A dialogue. Dædalus, 103(4):302–310, Fall 1974. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20024277.

Coles:1974:HSL

[1168] Robert Coles. But how is someone to live a life? Dædalus, 103(4):311–315, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024278.

Schudson:1974:SV

[1169] Michael Schudson. On the sense of vocation. Dædalus, 103(4):316–325, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024279.

Cottle:1974:FCC

[1170] Thomas J. Cottle. The first child at College. Dædalus, 103(4):326–331, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024280.

Sizer:1974:MCB

[1171] Theodore R. Sizer. On myopia: a complaint from down below. Dædalus, 103(4):332–340, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024281.

Gradd:1974:ACL

[1172] Edward Gradd. The American credo … of higher education. Dædalus, 103(4):341–345, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024282.

Anonymous:1974:BMd

[1173] Anonymous. Back matter. Dædalus, 103(4):??, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024283.

Anonymous:1974:FMd

[1174] Anonymous. Front matter. Dædalus, 103(4):??, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024233.

Anonymous:1974:VI

[1175] Anonymous. Volume information. Dædalus, 103(4):??, Fall 1974. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024232.

Graubard:1975:PIA

[1176] Stephen R. Graubard. Preface to the issue “American Higher Education: Toward an Uncertain Future”. volume II. Dædalus, 104(1):v–viii, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024286.

Kerr:1975:WWM

[1177] Clark Kerr. What we might learn from the climacteric. Dædalus, 104(1):1–7, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024287.

Fleming:1975:RHE

[1178] R. W. Fleming. Reflections on higher education. Dædalus, 104(1):8–15, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024288.

Bartell:1975:ENC

[1179] Ernest Bartell. The enduring noncrisis in higher education. Dædalus, 104(1):16–24, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024289.

Metzger:1975:AAP

[1180] Walter P. Metzger. The American academic profession in “Hard Times”. Dædalus, 104(1):25–44, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024290.

Wood:1975:ASB

[1181] Robert Wood. Academe sings the blues. Dædalus, 104(1):45–55, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024291.

Hitch:1975:TYA

[1182] Charles J. Hitch. Ten years after, ten years hence. Dædalus, 104(1):56–60, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024292.

Yarmolinsky:1975:IP

[1183] Adam Yarmolinsky. Institutional paralysis. Dædalus, 104(1):61–67, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024293.

Wallis:1975:UU

[1184] W. Allen Wallis. Unity in the University. Dædalus, 104(1):68–77, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024294.

McCarthy:1975:HEE

[1185] Patrick E. McCarthy. Higher education: Expansion without growth. Dædalus, 104(1):78–86, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20024295.

Kemeny:1975:USS

[1186] John G. Kemeny. The University in steady state. Dædalus, 104(1):87–96, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024296.

Zeckhauser:1975:EIP

[1187] Sally Zeckhauser and Richard Zeckhauser. Encouraging improved performance in higher education. Dædalus, 104(1):97–107, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024297.

Cronin:1975:IHE

[1188] Joseph M. Cronin. The interdependence of higher education. Dædalus, 104(1):108–112, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024298.

Trow:1975:PPL

[1189] Martin Trow. The public and private lives of higher education. Dædalus, 104(1):113–127, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024299.

Moynihan:1975:PHE

[1190] Daniel P. Moynihan. The politics of higher education. Dædalus, 104 (1):128–147, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024300.

Piel:1975:PSA

[1191] Gerard Piel. Public support for autonomous universities. Dædalus, 104 (1):148–155, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024301.

Lyman:1975:DPS

[1192] Richard W. Lyman. In defense of the private sector. Dædalus, 104 (1):156–159, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024302.

Jordan:1975:BHE

[1193] Vernon E. Jordan, Jr. Blacks and higher education: Some reflections. Dædalus, 104(1):160–165, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024303.

Grant:1975:EAR

[1194] Gerald Grant and David Riesman. An ecology of academic reform. Dædalus, 104(1):166–191, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024304.

Marshak:1975:PPU

[1195] Robert E. Marshak. Problems and prospects of an urban public University. Dædalus, 104(1):192–201, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024305.

Knowles:1975:CEC

[1196] Asa S. Knowles. Cooperative education: The catalyst for innovation and relevance. Dædalus, 104(1):202–209, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024306.

Ziegler:1975:SQB

[1197] Jerome M. Ziegler. Some questions before us: Notes toward the future in higher education. Dædalus, 104 (1):210–221, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024307.

Branscomb:1975:EPI

[1198] Lewis M. Branscomb and Paul C. Gilmore. Education in private industry. Dædalus, 104(1):222–233, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024308.

Haskins:1975:TUF

[1199] Caryl P. Haskins. Thoughts on an uncharted future. Dædalus, 104 (1):234–243, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024309.

Feld:1975:LPS

[1200] B. T. Feld. On legitimizing publicservice science in the University. Dædalus, 104(1):244–247, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024310.

Feshbach:1975:GEF

[1201] Herman Feshbach. Graduate education and Federal support of research. Dædalus, 104(1):248–250, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024311.

Stever:1975:ERS

[1202] H. Guyford Stever. Education, research, and the self-determination of academic institutions. Dædalus, 104 (1):251–253, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024312.

Glaser:1975:NUT

[1203] Robert J. Glaser. A note on the “University Troubles” and their impact on the medical schools. Dædalus, 104 (1):254–258, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024313.

Hiatt:1975:NUI

[1204] Howard H. Hiatt. The need for University involvement in medical education. Dædalus, 104(1):259–264, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024314.

Hauser:1975:PAU

[1205] Philip M. Hauser. Political actionism in the University. Dædalus, 104 (1):265–272, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024315.

Luria:1975:SIT

[1206] S. E. Luria and Zella Luria. The soluble and the insoluble or are two cultures better than one? Dædalus, 104 (1):273–277, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024316.

Tollett:1975:CHE

[1207] Kenneth S. Tollett. Community and higher education. Dædalus, 104(1): 278–297, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024317.

Boulding:1975:QVE

[1208] Kenneth E. Boulding. Quality versus equality: The dilemma of the University. Dædalus, 104(1):298–303, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024318.

Meyerson:1975:ADL

[1209] Martin Meyerson. After a decade of the levelers in higher education: Reinforcing quality while maintaining mass education. Dædalus, 104 (1):304–321, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024319.

Anonymous:1975:AUG

[1210] Anonymous. The assembly on University goals and governance. Dædalus, 104(1):322–346, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024320.

Anonymous:1975:BMa

[1211] Anonymous. Back matter. Dædalus, 104(1):??, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024321.

Anonymous:1975:FMa

[1212] Anonymous. Front matter. Dædalus, 104(1):??, Winter 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024285.

G:1975:PIW

[1213] S. R. G. Preface to the issue “Wisdom, Revelation, and Doubt: Perspectives on the First Millennium B.C.”. Dædalus, 104(2):v–vi, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024324.

Schwartz:1975:AT

[1214] Benjamin I. Schwartz. The age of transcendence. Dædalus, 104(2):1–7, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024325.

Momigliano:1975:FG

[1215] Arnaldo Momigliano. The fault of the Greeks. Dædalus, 104(2):9–19, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024326.

Weil:1975:WBH

[1216] Eric Weil. What is a breakthrough in history? Dædalus, 104(2):21–36, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024327.

Oppenheim:1975:PIM

[1217] A. Leo Oppenheim. The position of the intellectual in Mesopotamian society. Dædalus, 104(2):37–46, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024328.

Garelli:1975:CFC

[1218] Paul Garelli. The changing facets of conservative Mesopotamian thought. Dædalus, 104(2):47–56, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20024329.

Schwartz:1975:TAC

[1219] Benjamin I. Schwartz. Transcendence in ancient China. Dædalus, 104(2):57–68, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024330.

Nikiprowetzky:1975:EM

[1220] V. Nikiprowetzky. Ethical monotheism. Dædalus, 104(2):69–89, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024331.

Humphreys:1975:TIR

[1221] S. C. Humphreys. “Transcendence”

and intellectual roles: The ancient Greek case. Dædalus, 104(2):91–118, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024332.

Thapar:1975:ERS

[1222] Romila Thapar. Ethics, religion, and social protest in the First Millennium B.C. in Northern India. Dædalus, 104 (2):119–132, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024333.

Brown:1975:SSM

[1223] Peter Brown. Society and the supernatural: a medieval change. Dædalus, 104(2):133–151, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024334.

Dumont:1975:CUN

[1224] Louis Dumont. On the comparative understanding of non-modern civilizations. Dædalus, 104(2):153–172, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024335.

Anonymous:1975:NE

[1225] Anonymous. A note from the Editor. Dædalus, 104(2):173–174, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024336.

Darnton:1975:WNT

[1226] Robert Darnton. Writing news and telling stories. Dædalus, 104(2):175– 194, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20024337.

Anonymous:1975:BMb

[1227] Anonymous. Back matter. Dædalus, 104(2):??, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024338.

Anonymous:1975:FMb

[1228] Anonymous. Front matter. Dædalus, 104(2):??, Spring 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024323.

G:1975:PIA

[1229] S. R. G. Preface to the issue “Arms, Defense Policy, and Arms Control”. Dædalus, 104(3):v–vi, Summer 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024341.

Long:1975:ACP

[1230] F. A. Long. Arms control from the perspective of the Nineteen-Seventies. Dædalus, 104(3):1–13, Summer 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024342.

Chayes:1975:NAC

[1231] Abram Chayes. Nuclear arms control after the cold war. Dædalus, 104 (3):15–33, Summer 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024343.

Falk:1975:ACF

[1232] Richard A. Falk. Arms control, foreign policy, and global reform. Dædalus, 104(3):35–52, Summer 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024344.

Shulman:1975:ACI

[1233] Marshall D. Shulman. Arms control in an international context. Dædalus, 104(3):53–61, Summer 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024345.

Doty:1975:SAL

[1234] Paul Doty. Strategic arms limitation after SALT I. Dædalus, 104 (3):63–74, Summer 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024346.

Brooks:1975:MIS

[1235] Harvey Brooks. The military innovation system and the qualitative arms race. Dædalus, 104(3):75–97, Summer 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024347.

Allison:1975:AAC

[1236] Graham T. Allison and Frederic A. Morris. Armaments and arms control: Exploring the determinants of military weapons. Dædalus, 104 (3):99–129, Summer 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024348.

Steinbruner:1975:OPD

[1237] John Steinbruner and Barry Carter. Organizational and political dimensions of the strategic posture: The problems of reform. Dædalus, 104 (3):131–154, Summer 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024349.

Aspin:1975:DBF

[1238] Les Aspin. The defense budget and foreign policy: The role of Congress. Dædalus, 104(3):155–174, Summer 1975. CODEN DAEDAU.

ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024350.

Woolsey:1975:CB

[1239] R. James Woolsey. Chipping away at the bargains. Dædalus, 104(3): 175–185, Summer 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024351.

Schelling:1975:FEA

[1240] Thomas C. Schelling. A framework for the evaluation of arms-control proposals. Dædalus, 104(3):187–200, Summer 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024352.

Rathjens:1975:CPA

[1241] G. W. Rathjens. Changing perspectives on arms control. Dædalus, 104 (3):201–214, Summer 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024353.

Anonymous:1975:BMc

[1242] Anonymous. Back matter. Dædalus, 104(3):??, Summer 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024354.

Anonymous:1975:FMc

[1243] Anonymous. Front matter. Dædalus, 104(3):??, Summer 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024340.

G:1975:PIO

[1244] S. R. G. Preface to the issue “The Oil Crisis: In Perspective”. Dædalus, 104(4):v–vii, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024358.

Vernon:1975:I

[1245] Raymond Vernon. An interpretation. Dædalus, 104(4):1–14, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024359.

Darmstadter:1975:EB

[1246] Joel Darmstadter and Hans H. Landsberg. The economic background. Dædalus, 104(4):15–37, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024360.

Penrose:1975:DC

[1247] Edith Penrose. The development of crisis. Dædalus, 104(4):39–57, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024361.

Lenczowski:1975:OPC

[1248] George Lenczowski. The oil-producing countries. Dædalus, 104(4):59–72, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024362.

McKie:1975:US

[1249] James W. McKie. The United States. Dædalus, 104(4):73–90, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024363.

Prodi:1975:E

[1250] Romano Prodi and Alberto Clô. Europe. Dædalus, 104(4):91–112, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024364.

Tsurumi:1975:J

[1251] Yoshi Tsurumi. Japan. Dædalus, 104(4):113–127, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024365.

Goldman:1975:SU

[1252] Marshall I. Goldman. The Soviet union. Dædalus, 104(4):129–143, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024366.

Girvan:1975:EN

[1253] Norman Girvan. Economic nationalism. Dædalus, 104(4):145–158, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024367.

Wilkins:1975:OCP

[1254] Mira Wilkins. The oil companies in perspective. Dædalus, 104(4):159–178, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024368.

Stobaugh:1975:OCC

[1255] Robert B. Stobaugh. The oil companies in the crisis. Dædalus, 104(4):179–202, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024369.

Mikdashi:1975:OP

[1256] Zuhayr Mikdashi. The OPEC process. Dædalus, 104(4):203–215, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024370.

Lantzke:1975:OIE

[1257] Ulf Lantzke. The OECD and its international energy agency. Dædalus, 104(4):217–227, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024371.

Knorr:1975:LEM

[1258] Klaus Knorr. The limits of economic and military power. Dædalus, 104(4):229–243, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024372.

Vernon:1975:DP

[1259] Raymond Vernon. The distribution of power. Dædalus, 104(4):245–257, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024373.

Smart:1975:UG

[1260] Ian Smart. Uniqueness and generality. Dædalus, 104(4):259–281, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024374.

Anonymous:1975:ASC

[1261] Anonymous. Appendix a: Selective chronology of the oil crisis. Dædalus, 104(4):283–284, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024375.

Anonymous:1975:ABS

[1262] Anonymous. Appendix b: Statistical tables. Dædalus, 104(4):285–291, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024376.

Anonymous:1975:BMd

[1263] Anonymous. Back matter. Dædalus, 104(4):??, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024377.

Anonymous:1975:FMd

[1264] Anonymous. Front matter. Dædalus, 104(4):??, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024357.

Anonymous:1975:VI

[1265] Anonymous. Volume information. Dædalus, 104(4):??, Fall 1975. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024356.

G:1976:PIP

[1266] S. R. G. Preface to the issue “In Praise of Books”. Dædalus, 105(1):v–vii, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024380.

Letwin:1976:HC

[1267] Shirley Robin Letwin. Hobbes and Christianity. Dædalus, 105(1):1–21, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024381.

Tanner:1976:JMP

[1268] Tony Tanner. Julie and “La Maison Paternelle”: Another look at Rousseau’s “La Nouvelle Héloïse”. Dædalus, 105 (1):23–45, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024382.

Cazes:1976:CTP

[1269] Bernard Cazes. Condorcet’s true paradox, or, the liberal transformed into social engineer. Dædalus, 105(1):47–58, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024383.

Manuel:1976:MCG

[1270] Frank E. Manuel. In memoriam: “Critique of the Gotha Program”, 1875– 1975. Dædalus, 105(1):59–77, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024384.

vonStaden:1976:NMG

[1271] Heinrich von Staden. Nietzsche and Marx on Greek art and literature: Case studies in reception. Dædalus, 105(1):79–96, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024385.

Miller:1976:WPP

[1272] J. Hillis Miller. Walter pater: a partial portrait. Dædalus, 105(1):97–113, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024386.

Bell:1976:EWA

[1273] Pearl K. Bell. Edmund Wilson’s “Axel’s Castle”. Dædalus, 105(1):115– 125, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024387.

Burke:1976:FFW

[1274] Edmund Burke III. Frantz Fanon’s “The Wretched of the Earth”. Dædalus, 105(1):127–135, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024388.

Gellner:1976:SRM

[1275] Ernest Gellner. The sociology of Robert Montagne (1893–1954). Dædalus, 105(1):137–150, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024389.

Said:1976:RSR

[1276] Edward W. Said. Raymond Schwab and the romance of ideas. Dædalus, 105(1):151–167, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024390.

Forster:1976:AWH

[1277] Kurt W. Forster. Aby Warburg’s history of art: Collective memory and the social mediation of images. Dædalus, 105(1):169–176, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024391.

Zerner:1976:ARA

[1278] Henri Zerner. Aloïs Riegl: Art, value, and historicism. Dædalus, 105 (1):177–188, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024392.

Kostof:1976:AYH

[1279] Spiro Kostof. Architecture, you and him: The mark of Sigfried Giedion. Dædalus, 105(1):189–204, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024393.

Anonymous:1976:BMa

[1280] Anonymous. Back matter. Dædalus, 105(1):??, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024394.

Anonymous:1976:FMa

[1281] Anonymous. Front matter. Dædalus, 105(1):??, Winter 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024379.

G:1976:PIAa

[1282] S. R. G. Preface to the issue “Adulthood”. Dædalus, 105(2):v–viii, Spring 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024397.

Erikson:1976:RDB

[1283] Erik H. Erikson. Reflections on Dr.

Borg’s life cycle. Dædalus, 105(2):1–28, Spring 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024398.

Katchadourian:1976:MPA

[1284] Herant A. Katchadourian. Medical perspectives on adulthood. Dædalus, 105(2):29–56, Spring 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024399.

Bellah:1976:KSB

[1285] Robert N. Bellah. To kill and survive or to die and become: The active life and the contemplative life as ways of being adult. Dædalus, 105(2):57–76, Spring 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024400.

Bouwsma:1976:CA

[1286] William J. Bouwsma. Christian adulthood. Dædalus, 105(2):77–92, Spring 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024401.

Lapidus:1976:AIR

[1287] Ira M. Lapidus. Adulthood in Islam: Religious maturity in the Islamic tradition. Dædalus, 105(2):93–108, Spring 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024402.

Wei-Ming:1976:CPA

[1288] Tu Wei-Ming. The Confucian perception of adulthood. Dædalus, 105 (2):109–123, Spring 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024403.

Rohlen:1976:PAJ

[1289] Thomas P. Rohlen. The promise of adulthood in Japanese spiritualism. Dædalus, 105(2):125–143, Spring 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024404.

Rudolph:1976:RAR

[1290] Susanne Hoeber Rudolph and Lloyd I. Rudolph. Rajput adulthood: Reflections on the amar singh diary. Dædalus, 105(2):145–167, Spring 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024405.

Malia:1976:ARR

[1291] Martin E. Malia. Adulthood refracted:

Russia and Leo Tolstoi. Dædalus, 105 (2):169–183, Spring 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024406.

Anonymous:1976:BMb

[1292] Anonymous. Back matter. Dædalus, 105(2):??, Spring 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024407.

Anonymous:1976:FMb

[1293] Anonymous. Front matter. Dædalus, 105(2):??, Spring 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024396.

G:1976:PIE

[1294] S. R. G. Preface to the issue “Edward Gibbon and the Decline and Fall of the Roman Empire”. Dædalus, 105(3):v–vi, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024410.

Jordan:1976:EGH

[1295] David P. Jordan. Edward Gibbon: The historian of the Roman Empire. Dædalus, 105(3):1–12, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024411.

Brownley:1976:GFM

[1296] Martine Watson Brownley. Gibbon:

The formation of mind and character. Dædalus, 105(3):13–25, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024412.

Clive:1976:GH

[1297] John Clive. Gibbon’s humor. Dædalus, 105(3):27–35, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024413.

Shackleton:1976:IFL

[1298] Robert Shackleton. The impact of French literature on Gibbon. Dædalus, 105(3):37–48, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024414.

Giarrizzo:1976:TDF

[1299] Giuseppe Giarrizzo. Toward the “Decline and Fall”: Gibbon’s other historical interests. Dædalus, 105(3):49–62, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024415.

Bowersock:1976:GCW

[1300] G. W. Bowersock. Gibbon on civil war and rebellion in the decline of the Roman Empire. Dædalus, 105(3):63–71, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024416.

Brown:1976:GVC

[1301] Peter Brown. Gibbon’s views on culture and society in the Fifth and Sixth Centuries. Dædalus, 105(3):73–88, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024417.

Lewis:1976:GM

[1302] Bernard Lewis. Gibbon on Muhammad. Dædalus, 105(3):89–101, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024418.

Runciman:1976:GB

[1303] Steven Runciman. Gibbon and Byzantium. Dædalus, 105(3):103–110, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024419.

Chadwick:1976:GCH

[1304] Owen Chadwick. Gibbon and the Church historians. Dædalus, 105 (3):111–123, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024420.

Momigliano:1976:GIP

[1305] Arnaldo Momigliano. Gibbon from an Italian point of view. Dædalus, 105(3):125–135, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024421.

Burke:1976:TEI

[1306] Peter Burke. Tradition and experience: The idea of decline from Bruni to Gibbon. Dædalus, 105(3):137–152, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024422.

Pocock:1976:BMH

[1307] J. G. A. Pocock. Between Machiavelli and Hume: Gibbon as civic humanist and philosophical historian. Dædalus, 105(3):153–169, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024423.

Graubard:1976:EGC

[1308] Stephen R. Graubard. Edward Gibbon: Contraria sunt complementa. Dædalus, 105(3):171–187, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024424.

Starobinski:1976:DED

[1309] Jean Starobinski. From the decline of erudition to the decline of nations: Gibbon’s response to French thought. Dædalus, 105(3):189–207, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024425.

Furet:1976:CBG

[1310] François Furet. Civilization and barbarism in Gibbon’s history. Dædalus, 105(3):209–216, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024426.

Haskell:1976:GHA

[1311] Francis Haskell. Gibbon and the history of art. Dædalus, 105(3):217–229, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024427.

Manuel:1976:EGH

[1312] Frank E. Manuel. Edward Gibbon: Historien-philosophe. Dædalus, 105(3):231–245, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024428.

Brower:1976:GPR

[1313] Reuben A. Brower. With Gibbon in Puerto Rico. Dædalus, 105(3):247–249, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024429.

Anonymous:1976:BMc

[1314] Anonymous. Back matter. Dædalus, 105(3):??, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024430.

Anonymous:1976:FMc

[1315] Anonymous. Front matter. Dædalus, 105(3):??, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024409.

G:1976:PIAb

[1316] S. R. G. Preface to the issue “American Civilization: New Perspectives”. Dædalus, 105(4):v–viii, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024434.

Jordan:1976:SAA

[1317] Winthrop D. Jordan. Searching for adulthood in America. Dædalus, 105(4):1–11, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024435.

Hareven:1976:LSH

[1318] Tamara K. Hareven. The last stage: Historical adulthood and old age. Dædalus, 105(4):13–27, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024436.

Coles:1976:WSR

[1319] Robert Coles. Work and selfrespect. Dædalus, 105(4):29–38, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024437.

Stegner:1976:WCA

[1320] Wallace Stegner. The writer and the concept of adulthood. Dædalus, 105(4):39–48, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024438.

Lynn:1976:AAL

[1321] Kenneth S. Lynn. Adulthood in American literature. Dædalus, 105(4):49–59, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024439.

Rosenblatt:1976:GT

[1322] Roger Rosenblatt. Growing up on television. Dædalus, 105(4):61–68, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024440.

Goldstein:1976:BAB

[1323] Joseph Goldstein. On being adult and being an adult in secular law. Dædalus, 105(4):69–87, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024441.

Hurst:1976:CCT

[1324] Willard Hurst. Consensus and conflict in Twentieth-Century public policy. Dædalus, 105(4):89–101, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024442.

Casper:1976:SDF

[1325] Gerhard Casper. Social differences and the franchise. Dædalus, 105(4):103– 114, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024443.

Glazer:1976:LEF

[1326] Nathan Glazer. Liberty, equality, fraternity—and ethnicity. Dædalus, 105(4):115–127, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024444.

Karl:1976:PPP

[1327] Barry D. Karl. Philanthropy, policy planning, and the bureaucratization of the democratic ideal. Dædalus, 105(4):129–149, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024445.

Mansfield:1976:RR

[1328] Harvey C. Mansfield, Jr. The right of revolution. Dædalus, 105(4):151–162, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024446.

Anonymous:1976:BMd

[1329] Anonymous. Back matter. Dædalus, 105(4):??, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024447.

Anonymous:1976:FMd

[1330] Anonymous. Front matter. Dædalus, 105(4):??, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024433.

Anonymous:1976:VI

[1331] Anonymous. Volume information. Dædalus, 105(4):??, Fall 1976. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024432.

G:1977:PID

[1332] S. R. G. Preface to the issue “Doing Better and Feeling Worse: Health in the United States”. Dædalus, 106(1):v–vi, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024450.

Knowles:1977:I

[1333] John H. Knowles. Introduction. Dædalus, 106(1):1–7, Winter 1977. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20024451.

Fox:1977:MDA

[1334] Renée C. Fox. The medicalization and demedicalization of American society. Dædalus, 106(1):9–22, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024452.

Callahan:1977:HSS

[1335] Daniel Callahan. Health and society:

Some ethical imperatives. Dædalus, 106(1):23–33, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024453.

Thomas:1977:STM

[1336] Lewis Thomas. On the science and technology of medicine. Dædalus, 106(1):35–46, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024454.

Reiser:1977:TCM

[1337] Stanley Joel Reiser. Therapeutic choice and moral doubt in a technological age. Dædalus, 106(1):47–56, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024455.

Knowles:1977:RI

[1338] John H. Knowles. The responsibility of the individual. Dædalus, 106 (1):57–80, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024456.

Rogers:1977:CPC

[1339] David E. Rogers. The challenge of primary care. Dædalus, 106(1):81–103, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024457.

Wildavsky:1977:DBF

[1340] Aaron Wildavsky. Doing better and feeling worse: The political pathology of health policy. Dædalus, 106 (1):105–123, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024458.

Bennett:1977:TSF

[1341] Ivan L. Bennett, Jr. Technology as a shaping force. Dædalus, 106 (1):125–133, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024459.

McDermott:1977:EPH

[1342] Walsh McDermott. Evaluating the physician and his technology. Dædalus, 106(1):135–157, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024460.

Fredrickson:1977:HSN

[1343] Donald S. Fredrickson. Health and the search for new knowledge. Dædalus, 106(1):159–170, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024461.

Ebert:1977:MEU

[1344] Robert H. Ebert. Medical education in the United States. Dædalus, 106 (1):171–184, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024462.

Duval:1977:PGC

[1345] Merlin K. Duval. The provider, the government, and the consumer. Dædalus, 106(1):185–192, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024463.

Saward:1977:IOI

[1346] Ernest W. Saward. Institutional organization, incentives, and change. Dædalus, 106(1):193–202, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024464.

Ginzberg:1977:HSP

[1347] Eli Ginzberg. Health services, power centers, and decision-making mechanisms. Dædalus, 106(1):203–213, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024465.

Klarman:1977:FHC

[1348] Herbert E. Klarman. The financing of health care. Dædalus, 106 (1):215–234, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024466.

Eisenberg:1977:SC

[1349] Leon Eisenberg. The search for care. Dædalus, 106(1):235–246, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024467.

Richmond:1977:NC

[1350] Julius B. Richmond. The needs of children. Dædalus, 106(1):247–259, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024468.

Berger:1977:MHP

[1351] Philip Berger, Beatrix Hamburg, and David Hamburg. Mental health: Progress and problems. Dædalus, 106 (1):261–276, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024469.

Anonymous:1977:BMa

[1352] Anonymous. Back matter. Dædalus, 106(1):??, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024470.

Anonymous:1977:FMa

[1353] Anonymous. Front matter. Dædalus, 106(1):??, Winter 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024449.

G:1977:PIF

[1354] S. R. G. Preface to the issue “The Family”. Dædalus, 106(2):v–xiv, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024473.

Rossi:1977:BPP

[1355] Alice S. Rossi. A biosocial perspective on parenting. Dædalus, 106(2):1–31, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024474.

Kagan:1977:CF

[1356] Jerome Kagan. The child in the family. Dædalus, 106(2):33–56, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024475.

Hareven:1977:FTH

[1357] Tamara K. Hareven. Family time and historical time. Dædalus, 106(2):57–70, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024476.

Wrigley:1977:RHF

[1358] E. Anthony Wrigley. Reflections on the history of the family. Dædalus, 106(2):71–85, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024477.

Davis:1977:GKP

[1359] Natalie Zemon Davis. Ghosts, kin, and progeny: Some features of family life in early modern France. Dædalus, 106(2):87–114, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024478.

Sawhill:1977:EPF

[1360] Isabel V. Sawhill. Economic perspectives on the family. Dædalus, 106 (2):115–125, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024479.

Woolsey:1977:PPP

[1361] Suzanne H. Woolsey. Pied piper politics and the child-care debate. Dædalus, 106(2):127–145, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024480.

Blaydon:1977:ISP

[1362] Colin C. Blaydon and Carol B. Stack. Income support policies and the family. Dædalus, 106(2):147–161, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024481.

Downs:1977:IHP

[1363] Anthony Downs. The impact of housing policies on family life in the United States since world war II. Dædalus, 106(2):163–180, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024482.

Wagatsuma:1977:SAC

[1364] Hiroshi Wagatsuma. Some aspects of the contemporary Japanese family: Once Confucian, now fatherless? Dædalus, 106(2):181–210, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024483.

Whiting:1977:CLS

[1365] Beatrice B. Whiting. Changing life styles in Kenya. Dædalus, 106(2):211– 225, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024484.

Aries:1977:FC

[1366] Philippe Ariès. The family and the city. Dædalus, 106(2):227–235, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024485.

Anonymous:1977:BMb

[1367] Anonymous. Back matter. Dædalus, 106(2):??, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024486.

Anonymous:1977:FMb

[1368] Anonymous. Front matter. Dædalus, 106(2):??, Spring 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024472.

Graubard:1977:I

[1369] Stephen R. Graubard. Introduction. Dædalus, 106(3):v–xix, Summer 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024489.

Alpers:1977:AH

[1370] Svetlana Alpers. Is art history? Dædalus, 106(3):1–13, Summer 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024490.

Drell:1977:EPP

[1371] Sidney D. Drell. Elementary particle physics. Dædalus, 106(3):15–31, Summer 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024491.

Sciama:1977:LST

[1372] D. W. Sciama. The limits of space and time: Exploding black holes and the origin of the universe. Dædalus, 106(3):33–40, Summer 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024492.

Hoffmann:1977:ASS

[1373] Stanley Hoffmann. An American social science: International relations. Dædalus, 106(3):41–60, Summer 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024493.

Turner:1977:PSS

[1374] Victor Turner. Process, system, and symbol: a new anthropological synthesis. Dædalus, 106(3):61–80, Summer 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024494.

Willey:1977:CA

[1375] Gordon R. Willey. A consideration of archaeology. Dædalus, 106 (3):81–95, Summer 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024495.

McKenzie:1977:PTR

[1376] D. P. McKenzie. Plate tectonics and its relationship to the evolution of ideas in the geological sciences. Dædalus, 106(3):97–124, Summer 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024496.

Finley:1977:PH

[1377] M. I. Finley. “Progress” in historiography. Dædalus, 106(3):125–142, Summer 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024497.

Toulmin:1977:FFP

[1378] Stephen Toulmin. From form to function: Philosophy and history of science in the 1950s and now. Dædalus, 106 (3):143–162, Summer 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024498.

Thomas:1977:BSH

[1379] Lewis Thomas. Biomedical science and human health: The long-range prospect. Dædalus, 106(3):163–171, Summer 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024499.

Anonymous:1977:BMc

[1380] Anonymous. Back matter. Dædalus, 106(3):??, Summer 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024500.

Anonymous:1977:FMc

[1381] Anonymous. Front matter. Dædalus, 106(3):??, Summer 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024488.

Holton:1977:I

[1382] Gerald Holton. Introduction. Dædalus, 106(4):v–xiv, Fall 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024504.

Starobinski:1977:CA

[1383] Jean Starobinski, A. Cancogni, and R. Sieburth. Criticism and authority. Dædalus, 106(4):1–16, Fall 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024505.

Weinberg:1977:SUN

[1384] Steven Weinberg. The search for unity: Notes for a history of quantum field theory. Dædalus, 106(4):17–35, Fall 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024506.

Rossi:1977:XRA

[1385] Bruno Rossi. X-ray astronomy. Dædalus, 106(4):37–58, Fall 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024507.

Eisenstadt:1977:STA

[1386] Shmuel N. Eisenstadt. Sociological theory and an analysis of the dynamics of civilizations and of revolutions. Dædalus, 106(4):59–78, Fall 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024508.

Thurow:1977:E

[1387] Lester C. Thurow. Economics 1977. Dædalus, 106(4):79–94, Fall 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024509.

Culler:1977:PS

[1388] Jonathan Culler. In pursuit of signs. Dædalus, 106(4):95–111, Fall 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024510.

Miller:1977:PC

[1389] George A. Miller. Problems of communication. Dædalus, 106(4):113–125, Fall 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024511.

Wilson:1977:BSS

[1390] Edward O. Wilson. Biology and the social sciences. Dædalus, 106(4):127–140, Fall 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024512.

Ladurie:1977:RHD

[1391] Emmanuel B. LeRoy Ladurie and R. Sieburth. Recent historical “Discoveries”. Dædalus, 106(4):141–155, Fall 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024513.

Anonymous:1977:BMd

[1392] Anonymous. Back matter. Dædalus, 106(4):??, Fall 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024514.

Anonymous:1977:FMd

[1393] Anonymous. Front matter. Dædalus, 106(4):??, Fall 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024503.

Anonymous:1977:VI

[1394] Anonymous. Volume information. Dædalus, 106(4):??, Fall 1977. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024502.

G:1978:PIN

[1395] S. R. G. Preface to the issue “A New America?”. Dædalus, 107(1):v–vii, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024517.

Hardwick:1978:DM

[1396] Elizabeth Hardwick. Domestic manners. Dædalus, 107(1):1–11, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024518.

Ahlstrom:1978:NTC

[1397] Sydney E. Ahlstrom. National trauma and changing religious values. Dædalus, 107(1):13–29, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024519.

Muller:1978:NAU

[1398] Steven Muller. A new American University? Dædalus, 107(1):31–45, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024520.

Kanter:1978:WNA

[1399] Rosabeth Moss Kanter. Work in a new America. Dædalus, 107(1):47–78, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024521.

Dunlop:1978:PFT

[1400] John T. Dunlop. Past and future tendencies in American labor organizations. Dædalus, 107(1):79–96, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024522.

Wardell:1978:C

[1401] Nancy Needham Wardell. The corporation. Dædalus, 107(1):97–110, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024523.

Hamilton:1978:NMC

[1402] Edward K. Hamilton. On nonconstitutional management of a constitutional problem. Dædalus, 107(1):111–128, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024524.

Weinberg:1978:NJF

[1403] Louise Weinberg. A new judicial federalism? Dædalus, 107(1):129–141, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024525.

Wolfgang:1978:RPC

[1404] Marvin E. Wolfgang. Real and perceived changes of crime and punishment. Dædalus, 107(1):143–157, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024526.

Yarmolinsky:1978:WFP

[1405] Adam Yarmolinsky. What future for the professional in American society? Dædalus, 107(1):159–174, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024527.

Starr:1978:MWP

[1406] Paul Starr. Medicine and the waning of professional sovereignty. Dædalus, 107(1):175–193, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024528.

Smith:1978:JPW

[1407] Anthony Smith. “Just a Pleasant Way to Spend an Evening”: The softening embrace of American television. Dædalus, 107(1):195–212, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024529.

Sowell:1978:ECA

[1408] Thomas Sowell. Ethnicity in a changing America. Dædalus, 107(1):213–237, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024530.

Zashin:1978:PBA

[1409] Elliot Zashin. The progress of black Americans in civil rights: The past two decades assessed. Dædalus, 107 (1):239–262, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024531.

Yergin:1978:OS

[1410] Daniel Yergin. Order and survival. Dædalus, 107(1):263–287, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024532.

Bundy:1978:AW

[1411] McGeorge Bundy. The Americans and the world. Dædalus, 107(1):289–303, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024533.

McKinnon:1978:ARS

[1412] Ronald I. McKinnon. America’s role in stabilizing the World’s monetary system. Dædalus, 107(1):305–324, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024534.

Nora:1978:AFI

[1413] Pierre Nora and Michael Taylor. America and the French intellectuals. Dædalus, 107(1):325–337, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024535.

Anonymous:1978:BMa

[1414] Anonymous. Back matter. Dædalus, 107(1):??, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024536.

Anonymous:1978:FMa

[1415] Anonymous. Front matter. Dædalus, 107(1):??, Winter 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024516.

G:1978:PIL

[1416] S. R. G. Preface to the issue, “Limits of Scientific Inquiry”. Dædalus, 107(2):v–vi, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024539.

Morison:1978:I

[1417] Robert S. Morison. Introduction. Dædalus, 107(2):vii–xvi, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024540.

Graham:1978:CAS

[1418] Loren R. Graham. Concerns about science and attempts to regulate inquiry. Dædalus, 107(2):1–21, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024541.

Sinsheimer:1978:PS

[1419] Robert L. Sinsheimer. The presumptions of science. Dædalus, 107(2):23– 35, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024542.

Baltimore:1978:LSB

[1420] David Baltimore. Limiting science:

a Biologist’s perspective. Dædalus, 107(2):37–45, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024543.

White:1978:SSS

[1421] Lynn White, Jr. Science and the sense of self: The medieval background of a modern confrontation. Dædalus, 107(2):47–59, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024544.

Marx:1978:RNR

[1422] Leo Marx. Reflections on the neoromantic critique of science. Dædalus, 107(2):61–74, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024545.

Price:1978:EFB

[1423] Don K. Price. Endless frontier or bureaucratic morass? Dædalus, 107 (2):75–92, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024546.

Metzger:1978:AFS

[1424] Walter P. Metzger. Academic freedom and scientific freedom. Dædalus, 107(2):93–114, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024547.

Bok:1978:FR

[1425] Sissela Bok. Freedom and risk. Dædalus, 107(2):115–127, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024548.

Swazey:1978:PAN

[1426] Judith P. Swazey. Protecting the “Animal of Necessity”: Limits to inquiry in clinical investigation. Dædalus, 107 (2):129–145, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024549.

Culliton:1978:SRP

[1427] Barbara J. Culliton. Science’s restive public. Dædalus, 107(2):147–156, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024550.

Hutt:1978:PCH

[1428] Peter Barton Hutt. Public criticism of health science policy. Dædalus, 107 (2):157–169, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024551.

Brooks:1978:PRP

[1429] Harvey Brooks. The problem of research priorities. Dædalus, 107 (2):171–190, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024552.

Nelkin:1978:TPN

[1430] Dorothy Nelkin. Threats and promises: Negotiating the control of research. Dædalus, 107(2):191–209, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024553.

Morison:1978:MAL

[1431] Robert S. Morison. Misgivings about life-extending technologies. Dædalus, 107(2):211–226, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024554.

Holton:1978:EIL

[1432] Gerald Holton. Epilogue to the issue, “Limits of Scientific Inquiry”. Dædalus, 107(2):227–234, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024555.

Anonymous:1978:BMb

[1433] Anonymous. Back matter. Dædalus, 107(2):??, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024556.

Anonymous:1978:FMb

[1434] Anonymous. Front matter. Dædalus, 107(2):??, Spring 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024538.

G:1978:PIR

[1435] S. R. G. Preface to the issue, “Rousseau for Our Time”. Dædalus, 107(3):v–vii, Summer 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024559.

Manuel:1978:DE

[1436] Frank E. Manuel. A dream of eupsychia. Dædalus, 107(3):1–12, Summer 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024560.

Shklar:1978:JJR

[1437] Judith N. Shklar. Jean-Jacques Rousseau and equality. Dædalus, 107 (3):13–25, Summer 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024561.

Baczko:1978:RSM

[1438] Bronislaw Baczko and Steven Raichlen. Rousseau and social marginality. Dædalus, 107(3):27–40, Summer 1978.

CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20024562.

Starobinski:1978:AA

[1439] Jean Starobinski and Jacqueline Brovender. The accuser and the accused. Dædalus, 107(3):41–58, Summer 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024563.

Gossman:1978:IAC

[1440] Lionel Gossman. The innocent art of confession and reverie. Dædalus, 107(3):59–77, Summer 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024564.

Barber:1978:RPD

[1441] Benjamin R. Barber. Rousseau and the paradoxes of the dramatic imagination. Dædalus, 107(3):79–92, Summer 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024565.

Masters:1978:JJA

[1442] Roger D. Masters. Jean-Jacques is alive and well: Rousseau and contemporary sociobiology. Dædalus, 107 (3):93–105, Summer 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024566.

Wokler:1978:PAD

[1443] Robert Wokler. Perfectible apes in decadent cultures: Rousseau’s anthropology revisited. Dædalus, 107(3):

107–134, Summer 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024567.

Bloom:1978:EDM

[1444] Allan Bloom. The education of democratic man: Emile. Dædalus, 107 (3):135–153, Summer 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024568.

Kessen:1978:RC

[1445] William Kessen. Rousseau’s children. Dædalus, 107(3):155–166, Summer 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024569.

Featherstone:1978:RM

[1446] Joseph Featherstone. Rousseau and modernity. Dædalus, 107(3):167–192, Summer 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20024570.

Schwartz:1978:RSC

[1447] Benjamin I. Schwartz. The Rousseau strain in the contemporary world. Dædalus, 107(3):193–206, Summer 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024571.

Anonymous:1978:BMc

[1448] Anonymous. Back matter. Dædalus, 107(3):??, Summer 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024572.

Anonymous:1978:FMc

[1449] Anonymous. Front matter. Dædalus, 107(3):??, Summer 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024558.

G:1978:PIG

[1450] S. R. G. Preface to the issue, “Generations”. Dædalus, 107(4):v–vii, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024576.

Nash:1978:CEG

[1451] Laura L. Nash. Concepts of existence: Greek origins of generational thought. Dædalus, 107(4):1–21, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024577.

Kriegel:1978:GDH

[1452] Annie Kriegel and Elisabeth Hirsch.

Generational difference: The history of an idea. Dædalus, 107(4):23–38, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024578.

Riley:1978:ASC

[1453] Matilda White Riley. Aging, social change, and the power of ideas. Dædalus, 107(4):39–52, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024579.

Letwin:1978:TGG

[1454] Shirley Robin Letwin. Trollope on generations without gaps. Dædalus, 107(4):53–70, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024580.

Skidelsky:1978:KHP

[1455] Robert Skidelsky. Keynes and his parents. Dædalus, 107(4):71–79, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024581.

Annan:1978:OAR

[1456] Noel Annan. “Our Age”: Reflections on three generations in England. Dædalus, 107(4):81–109, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024582.

Schorske:1978:GTC

[1457] Carl E. Schorske. Generational tension and cultural change: Reflections on the case of Vienna. Dædalus, 107(4):111–122, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024583.

Keller:1978:RPG

[1458] Morton Keller. Reflections on politics and generations in America. Dædalus, 107(4):123–135, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024584.

Hareven:1978:SGM

[1459] Tamara K. Hareven. The search for generational memory: Tribal rites in industrial society. Dædalus, 107 (4):137–149, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024585.

Douglas:1978:JJF

[1460] Mary Douglas. Judgments on James frazer. Dædalus, 107(4):151–164, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024586.

Bush:1978:LAP

[1461] Douglas Bush. Literature, the Academy, and the public. Dædalus, 107(4):165–174, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

/www.jstor.org/stable/20024587.

Lear:1978:GN

[1462] Jonathan Lear. Going native. Dædalus, 107(4):175–188, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024588.

Isaacs:1978:BFQ

[1463] Harold R. Isaacs. Bringing up the father question. Dædalus, 107(4):189– 203, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024589.

Anonymous:1978:BMd

[1464] Anonymous. Back matter. Dædalus, 107(4):??, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024590.

Anonymous:1978:FMd

[1465] Anonymous. Front matter. Dædalus, 107(4):??, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024575.

Anonymous:1978:VI

[1466] Anonymous. Volume information. Dædalus, 107(4):??, Fall 1978. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024574.

G:1979:PIL

[1467] S. R. G. Preface to the issue “Looking for Europe”. Dædalus, 108(1):v–xi, Winter 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024593.

Hoffmann:1979:FFH

[1468] Stanley Hoffmann. Fragments floating in the here and now. Dædalus, 108(1):1–26, Winter 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024594.

Berger:1979:PAW

[1469] Suzanne Berger. Politics and antipolitics in Western Europe in the Seventies. Dædalus, 108(1):27–50, Winter 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024595.

Bressand:1979:NEE

[1470] Albert Bressand. The new European economies. Dædalus, 108(1):51–73, Winter 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024596.

Watt:1979:CPP

[1471] David Watt. The community: Performance and prospects. Dædalus, 108(1):75–85, Winter 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024597.

Nerlich:1979:WER

[1472] Uwe Nerlich. Western Europe’s relations with the United States. Dædalus, 108(1):87–111, Winter 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024598.

Hassner:1979:WEP

[1473] Pierre Hassner. Western European perceptions of the USSR. Dædalus, 108 (1):113–150, Winter 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024599.

Brown:1979:EE

[1474] Archie Brown. Eastern Europe: 1968, 1978, 1998. Dædalus, 108(1):151–174, Winter 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024600.

Linz:1979:ESF

[1475] Juan Linz. Europe’s southern frontier: Evolving trends toward what?

Dædalus, 108(1):175–209, Winter 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024601.

Anonymous:1979:BMa

[1476] Anonymous. Back matter. Dædalus, 108(1):??, Winter 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024602.

Anonymous:1979:FMa

[1477] Anonymous. Front matter. Dædalus, 108(1):??, Winter 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024592.

G:1979:PIO

[1478] S. R. G. Preface to the issue “Old Faiths and New Doubts: The European Predicament”. Dædalus, 108(2):v–xi, Spring 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024605.

Bourricaud:1979:IMC

[1479] François Bourricaud and John Muresianu. Individualistic mobilization and the crisis of professional authority. Dædalus, 108(2):1–20, Spring 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024606.

Besancon:1979:CT

[1480] Alain Besançon. The confusion of tongues. Dædalus, 108(2):21–42, Spring 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024607.

Salvati:1979:RMP

[1481] Michele Salvati and Giorgio Brosio.

The rise of market politics: Industrial relations in the seventies. Dædalus, 108(2):43–71, Spring 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024608.

Land:1979:CPW

[1482] Hilary Land. The changing place of women in Europe. Dædalus, 108 (2):73–94, Spring 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024609.

Rist:1979:MMG

[1483] Ray C. Rist. Migration and marginality: Guestworkers in Germany and France. Dædalus, 108(2):95–108, Spring 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024610.

Bressand:1979:UDM

[1484] Albert Bressand and Thierry De Montbrial. The ups and downs of mutual relevance. Dædalus, 108(2):109–132, Spring 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024611.

Briggs:1979:WIC

[1485] Asa Briggs. Word and image:

Changing patterns of communications. Dædalus, 108(2):133–149, Spring 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024612.

Ross:1979:EMF

[1486] Werner Ross. Europe’s multilingual future. Dædalus, 108(2):151–162, Spring 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024613.

Anonymous:1979:BMb

[1487] Anonymous. Back matter. Dædalus, 108(2):??, Spring 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024614.

Anonymous:1979:FMb

[1488] Anonymous. Front matter. Dædalus, 108(2):??, Spring 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024604.

G:1979:PIH

[1489] S. R. G. Preface to the issue “Hypocrisy, Illusion, and Evasion”. Dædalus, 108(3):v–viii, Summer 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024617.

Shklar:1979:LUH

[1490] Judith Shklar. Let us not be hypocritical. Dædalus, 108(3):1–25, Summer 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024618.

Cavell:1979:ETR

[1491] Stanley Cavell. Epistemology and tragedy: a reading of Othello. Dædalus, 108(3):27–43, Summer 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024619.

Vickers:1979:SH

[1492] Brian Vickers. Shakespeare’s hypocrites. Dædalus, 108(3):45–83, Summer 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024620.

Starobinski:1979:MID

[1493] Jean Starobinski and John Muresianu.

Montaigne on illusion: The denunciation of untruth. Dædalus, 108 (3):85–101, Summer 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024621.

Schama:1979:URA

[1494] Simon Schama. The unruly realm:

Appetite and restraint in Seventeenth Century Holland. Dædalus, 108(3): 103–123, Summer 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024622.

Bateson:1979:FTS

[1495] Mary Catherine Bateson. “This Figure of Tinsel”: a study of themes of hypocrisy and pessimism in Iranian culture. Dædalus, 108(3):125–134, Summer 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024623.

Gilman:1979:CDA

[1496] Stephen Gilman. The case of Diego Alonso: Hypocrisy and the Spanish Inquisition. Dædalus, 108(3):135–144, Summer 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20024624.

Anderson:1979:JDA

[1497] Quentin Anderson. John Dewey’s American Democrat. Dædalus, 108 (3):145–159, Summer 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024625.

Anonymous:1979:BMc

[1498] Anonymous. Back matter. Dædalus, 108(3):??, Summer 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024626.

Anonymous:1979:FMc

[1499] Anonymous. Front matter. Dædalus, 108(3):??, Summer 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024616.

G:1979:PIS

[1500] S. R. G. Preface to the issue, “The State”. Dædalus, 108(4):v–xix, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024630.

Eckstein:1979:SS

[1501] Harry Eckstein. On the “Science” of the state. Dædalus, 108(4):1–20, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024631.

Kelly:1979:WNT

[1502] George Armstrong Kelly. Who needs a theory of citizenship? Dædalus, 108(4):21–36, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024632.

Rae:1979:ESN

[1503] Douglas Rae. The egalitarian state:

Notes on a system of contradictory ideals. Dædalus, 108(4):37–54, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024633.

Fishkin:1979:MPP

[1504] James Fishkin. Moral principles and public policy. Dædalus, 108(4):55–67, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024634.

Logue:1979:WSV

[1505] John Logue. The welfare state: Victim of its success. Dædalus, 108(4):69–87, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024635.

Abt:1979:SSR

[1506] Clark C. Abt. Social science research and the modern state. Dædalus, 108(4):89–100, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024636.

Howard:1979:WNS

[1507] Michael Howard. War and the nationstate. Dædalus, 108(4):101–110, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024637.

Bull:1979:SPR

[1508] Hedley Bull. The State’s positive role in world affairs. Dædalus, 108 (4):111–123, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024638.

Haass:1979:PSL

[1509] Richard Haass. The primacy of the state … or revising the revisionists. Dædalus, 108(4):125–138, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024639.

Kriegel:1979:NCS

[1510] Annie Kriegel, J. Braun, and John Muresianu. The nature of the communist system: Notes on state, party, and society. Dædalus, 108(4):139–153, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024640.

Apter:1979:NUL

[1511] David E. Apter. Notes on the underground: Left violence and the national state. Dædalus, 108(4):155–172, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024641.

Anonymous:1979:BMd

[1512] Anonymous. Back matter. Dædalus, 108(4):??, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024642.

Anonymous:1979:FMd

[1513] Anonymous. Front matter. Dædalus, 108(4):??, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024629.

Anonymous:1979:VI

[1514] Anonymous. Volume information. Dædalus, 108(4):??, Fall 1979. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024628.

G:1980:PIM

[1515] S. R. G. Preface to the issue, “Modern Technology: Problem or Opportunity?”. Dædalus, 109(1):v–vi, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024645.

Brooks:1980:SIT

[1516] Harvey Brooks, Alfred Chandler, Norman Dahl, Edward E. David, Jr., Stephen R. Graubard, Albert Hirschman, Gerald Holton, David S. Landes, Michael Maccoby, Elting Morison, Robert S. Morison, Walter Rosenblith, Merritt Roe Smith, Raymond Vernon, Joseph Weizenbaum, and Langdon Winner. Some issues of technology. Dædalus, 109(1):1 3–24, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024646.

Hannay:1980:AMT

[1517] N. Bruce Hannay and Robert E. McGinn. The anatomy of modern technology: Prolegomenon to an improved public policy for the social management of technology. Dædalus, 109(1):25–53, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024647.

Morison:1980:FNV

[1518] Robert S. Morison. A further note on visions. Dædalus, 109(1):55–64, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024648.

Brooks:1980:TEP

[1519] Harvey Brooks. Technology, evolution, and purpose. Dædalus, 109 (1):65–81, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024649.

Baker:1980:ST

[1520] W. O. Baker. Science and technology. Dædalus, 109(1):83–109, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024650.

Landes:1980:CKT

[1521] David Landes. The creation of knowledge and technique: Today’s task and Yesterday’s experience. Dædalus, 109 (1):111–120, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024651.

Winner:1980:DAP

[1522] Langdon Winner. Do artifacts have politics? Dædalus, 109(1):121–136, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024652.

Spreng:1980:TD

[1523] Daniel T. Spreng and Alvin M.

Weinberg. Time and decentralization. Dædalus, 109(1):137–143, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024653.

Shapley:1980:ACR

[1524] Deborah Shapley. Arms control as a regulator of military technology. Dædalus, 109(1):145–157, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024654.

Stout:1980:ITE

[1525] David K. Stout. The impact of technology on economic growth in the 1980s. Dædalus, 109(1):159–167, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024655.

David:1980:DTC

[1526] Edward E. David, Jr. On the dimensions of the technology controversy. Dædalus, 109(1):169–177, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024656.

Morison:1980:UR

[1527] Elting Morison. The uncertain relation. Dædalus, 109(1):179–184, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024657.

Rosenblith:1980:NEO

[1528] Walter Rosenblith. A note on engineering and the other professions. Dædalus, 109(1):185–187, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024658.

Anonymous:1980:BMa

[1529] Anonymous. Back matter. Dædalus, 109(1):??, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024659.

Anonymous:1980:FMa

[1530] Anonymous. Front matter. Dædalus, 109(1):??, Winter 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024644.

G:1980:PII

[1531] S. R. G. Preface to the issue: Intellect and imagination: The limits and presuppositions of intellectual inquiry. Dædalus, 109(2):v–viii, Spring 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024662.

Cooper:1980:SLH

[1532] Leon N. Cooper. Source and limits of human intellect. Dædalus, 109(2):1–17, Spring 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024663.

Pribram:1980:RAT

[1533] Karl H. Pribram. The role of analogy in transcending limits in the brain sciences. Dædalus, 109(2):19–38, Spring 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024664.

Gould:1980:EBC

[1534] Stephen Jay Gould. The evolutionary biology of constraint. Dædalus, 109(2):39–52, Spring 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024665.

Shklar:1980:LK

[1535] Judith N. Shklar. Learning without knowing. Dædalus, 109(2):53–72, Spring 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024666.

Boon:1980:CEP

[1536] James A. Boon. Comparative Deenlightenment: Paradox and limits in the history of ethnology. Dædalus, 109(2):73–91, Spring 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024667.

OFlaherty:1980:IOM

[1537] Wendy Doniger O’Flaherty. Inside and outside the mouth of God: The boundary between myth and reality. Dædalus, 109(2):93–125, Spring 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024668.

Skura:1980:CTL

[1538] Meredith Skura. Creativity: Transgressing the limits of consciousness. Dædalus, 109(2):127–146, Spring 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024669.

Cavell:1980:KTM

[1539] Stanley Cavell. Knowledge as transgression: Mostly a reading of “It Happened One Night”. Dædalus, 109 (2):147–175, Spring 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024670.

Meyer:1980:ELC

[1540] Leonard B. Meyer. Exploiting limits: Creation, archetypes, and style change. Dædalus, 109(2):177–205, Spring 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024671.

Steinberg:1980:CLJ

[1541] Leo Steinberg. A corner of the “Last Judgment”. Dædalus, 109(2):207– 217 221–273, Spring 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024672.

Anonymous:1980:BMb

[1542] Anonymous. Back matter. Dædalus, 109(2):??, Spring 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024673.

Anonymous:1980:FMb

[1543] Anonymous. Front matter. Dædalus, 109(2):??, Spring 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024661.

G:1980:PIE

[1544] S. R. G. Preface to the issue: “The End of Consensus?”. Dædalus, 109 (3):v–xv, Summer 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024676.

Barton:1980:FLA

[1545] Allen H. Barton. Fault lines in American elite consensus. Dædalus, 109 (3):1–24, Summer 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024677.

Fiorina:1980:DCR

[1546] Morris P. Fiorina. The decline of collective responsibility in American politics. Dædalus, 109(3):25–45, Summer 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024678.

Vogel:1980:ICO

[1547] David Vogel. The inadequacy of contemporary opposition to business. Dædalus, 109(3):47–58, Summer 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024679.

Bosworth:1980:REE

[1548] Barry Bosworth. Re-establishing an economic consensus: an impossible agenda? Dædalus, 109(3):59–70, Summer 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024680.

Landsberg:1980:LAP

[1549] Hans H. Landsberg. Let’s all play energy policy! Dædalus, 109(3):71– 84, Summer 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024681.

Atkin:1980:GC

[1550] J. Myron Atkin. The government in the classroom. Dædalus, 109 (3):85–97, Summer 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024682.

Wood:1980:DAE

[1551] Robert Wood. The disassembling of American education. Dædalus, 109 (3):99–113, Summer 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024683.

Graham:1980:WEE

[1552] Patricia Albjerg Graham. Whither equality of educational opportunity? Dædalus, 109(3):115–132, Summer 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024684.

Coles:1980:CP

[1553] Robert Coles. Civility and psychology. Dædalus, 109(3):133–141, Summer 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024685.

Miller:1980:MFP

[1554] Linda B. Miller. Morality in foreign policy: a failed consensus? Dædalus, 109(3):143–158, Summer 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024686.

Staub:1980:TM

[1555] Hans O. Staub and Harry Zohn. The tyranny of minorities. Dædalus, 109 (3):159–168, Summer 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024687.

Anonymous:1980:BMc

[1556] Anonymous. Back matter. Dædalus, 109(3):??, Summer 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024688.

Anonymous:1980:FMc

[1557] Anonymous. Front matter. Dædalus, 109(3):??, Summer 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024675.

G:1980:PIU

[1558] S. R. G. Preface to the issue: “U.S. Defense Policy in the 1980s”. Dædalus, 109(4):v–xv, Fall 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024692.

Pipes:1980:MSS

[1559] Richard Pipes. Militarism and the Soviet state. Dædalus, 109(4):1–12, Fall 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024693.

Holloway:1980:MPP

[1560] David Holloway. Military power and political purpose in Soviet policy. Dædalus, 109(4):13–30, Fall 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024694.

Chubin:1980:USI

[1561] Shahram Chubin. U.S. security interests in the Persian Gulf in the 1980s. Dædalus, 109(4):31–65, Fall 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024695.

Jabber:1980:UIR

[1562] Paul Jabber. U.S. interests and regional security in the Middle East. Dædalus, 109(4):67–80, Fall 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024696.

Rowen:1980:ASI

[1563] Henry Rowen. American security interests in Northeast Asia. Dædalus, 109(4):81–96, Fall 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024697.

Whiting:1980:CST

[1564] Allen S. Whiting. China and the superpowers: Toward the year 2000. Dædalus, 109(4):97–113, Fall 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024698.

Dominguez:1980:USR

[1565] Jorge I. Domínguez. The United States and its regional security interests: The Caribbean, central, and south America. Dædalus, 109(4):115–133, Fall 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024699.

Gray:1980:SSR

[1566] Colin S. Gray. Strategic stability reconsidered. Dædalus, 109(4):135–154, Fall 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024700.

Lodal:1980:DNS

[1567] Jan M. Lodal. Deterrence and nuclear strategy. Dædalus, 109(4):155–175, Fall 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024701.

Drell:1980:ACT

[1568] Sidney D. Drell. Arms control: Is there still hope? Dædalus, 109(4):177–188, Fall 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024702.

Anonymous:1980:BMd

[1569] Anonymous. Back matter. Dædalus, 109(4):??, Fall 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024703.

Anonymous:1980:FMd

[1570] Anonymous. Front matter. Dædalus, 109(4):??, Fall 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024691.

Anonymous:1980:VI

[1571] Anonymous. Volume information. Dædalus, 109(4):??, Fall 1980. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024690.

G:1981:PIU

[1572] S. R. G. Preface to the issue, “U.S. Defense Policy in the 1980s”: Volume II. Dædalus, 110(1):v–xi, Winter 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024706.

Nacht:1981:TAC

[1573] Michael Nacht. Toward an American conception of regional security. Dædalus, 110(1):1–22, Winter 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024707.

Quester:1981:SAA

[1574] George H. Quester. The superpowers and the Atlantic alliance. Dædalus, 110(1):23–40, Winter 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024708.

Hyland:1981:AC

[1575] William G. Hyland. The Atlantic crisis. Dædalus, 110(1):41–51, Winter 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024709.

Bowie:1981:AA

[1576] Robert R. Bowie. The Atlantic alliance. Dædalus, 110(1):53–69, Winter 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024710.

Nerlich:1981:CES

[1577] Uwe Nerlich. Change in Europe:

a secular trend? Dædalus, 110 (1):71–103, Winter 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024711.

Kincade:1981:TH

[1578] William H. Kincade. Over the technological horizon. Dædalus, 110(1):105– 127, Winter 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024712.

Brooks:1981:NSI

[1579] Harvey Brooks. Notes on some issues on technology and national defense. Dædalus, 110(1):129–136, Winter 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024713.

Kruzel:1981:ACA

[1580] Joseph J. Kruzel. Arms control and American defense policy: New alternatives and old realities. Dædalus, 110 (1):137–157, Winter 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024714.

Burt:1981:RAC

[1581] Richard Burt. The relevance of arms control in the 1980s. Dædalus, 110 (1):159–177, Winter 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024715.

Rathjens:1981:NDR

[1582] George Rathjens and Jack Ruina. Nuclear doctrine and rationality. Dædalus, 110(1):179–187, Winter 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024716.

Anonymous:1981:BMa

[1583] Anonymous. Back matter. Dædalus, 110(1):??, Winter 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024717.

Anonymous:1981:FMa

[1584] Anonymous. Front matter. Dædalus, 110(1):??, Winter 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024705.

G:1981:PIAa

[1585] S. R. G. Preface to the issue, “American Indians, Blacks, Chicanos, and Puerto Ricans”. Dædalus, 110(2):v– xiii, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024720.

Franklin:1981:LRE

[1586] John Hope Franklin. The land of room enough. Dædalus, 110(2):1–12, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024721.

Deloria:1981:IC

[1587] Vine Deloria, Jr. Identity and culture. Dædalus, 110(2):13–27, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024722.

Coles:1981:MDA

[1588] Robert Coles. Minority dreams, American dreams. Dædalus, 110(2):29–41, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024723.

Dorris:1981:GSG

[1589] Michael A. Dorris. The grass still grows, the rivers still flow: Contemporary native Americans. Dædalus, 110(2):43–69, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024724.

Jones:1981:ECI

[1590] Faustine C. Jones. External crosscurrents and internal diversity: an assessment of black progress, 1960– 1980. Dædalus, 110(2):71–101, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024725.

Estrada:1981:CUS

[1591] Leobardo F. Estrada, F. Chris García, Reynaldo Flores Macías, and Lionel Maldonado. Chicanos in the United States: a history of exploitation and resistance. Dædalus, 110(2):103–131, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024726.

Bonilla:1981:WPP

[1592] Frank Bonilla and Ricardo Campos. A wealth of poor: Puerto Ricans in the New Economic Order. Dædalus, 110(2):133–176, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024727.

Arce:1981:RCC

[1593] Carlos H. Arce. A reconsideration of Chicano culture and identity. Dædalus, 110(2):177–191, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024728.

Flores:1981:CMU

[1594] Juan Flores, John Attinasi, and Pedro Pedraza, Jr. “La Carreta Made a U-Turn”: Puerto Rican language and culture in the United States. Dædalus, 110(2):193–217, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024729.

Burns:1981:BBB

[1595] Haywood Burns. From Brown to Bakke and back: Race, law, and social change in America. Dædalus, 110 (2):219–231, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024730.

Pettigrew:1981:RCI

[1596] Thomas F. Pettigrew. Race and class in the 1980s: an interactive view. Dædalus, 110(2):233–255, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024731.

Anderson:1981:EQE

[1597] Bernard E. Anderson and Phoebe H.

Cottingham. The elusive quest for economic equality. Dædalus, 110 (2):257–274, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024732.

Moore:1981:MAC

[1598] Joan W. Moore. Minorities in the American class system. Dædalus, 110 (2):275–299, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024733.

Anonymous:1981:BMb

[1599] Anonymous. Back matter. Dædalus, 110(2):??, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024734.

Anonymous:1981:FMb

[1600] Anonymous. Front matter. Dædalus, 110(2):??, Spring 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024719.

G:1981:PIAb

[1601] S. R. G. Preface to the issue, “America’s Schools: Public and Private”. Dædalus, 110(3):v–xxiv, Summer 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024737.

Tyack:1981:CCA

[1602] David Tyack and Elisabeth Hansot. Conflict and consensus in American public education. Dædalus, 110(3):1– 25, Summer 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20024738.

Bailey:1981:PCP

[1603] Stephen K. Bailey and Steven K. Bailey. Political coalitions for public education. Dædalus, 110(3):27–43, Summer 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024739.

Kirst:1981:LSP

[1604] Michael W. Kirst. Loss of support for public secondary schools: Some causes and solutions. Dædalus, 110 (3):45–68, Summer 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024740.

Cohen:1981:FHS

[1605] David K. Cohen and Barbara Neufeld.

The failure of high schools and the progress of education. Dædalus, 110 (3):69–89, Summer 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024741.

Atkin:1981:WWT

[1606] J. Myron Atkin. Who will teach in high school? Dædalus, 110 (3):91–103, Summer 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024742.

McAndrew:1981:HSP

[1607] Gordon L. McAndrew. The highschool principal: Man in the middle. Dædalus, 110(3):105–118, Summer 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024743.

Graham:1981:LGS

[1608] Patricia Albjerg Graham. Literacy: a goal for secondary schools. Dædalus, 110(3):119–134, Summer 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024744.

Grant:1981:CEE

[1609] Gerald Grant. The character of education and the education of character. Dædalus, 110(3):135–149, Summer 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024745.

Kagan:1981:MFS

[1610] Jerome Kagan. The moral function of the school. Dædalus, 110 (3):151–165, Summer 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024746.

Anonymous:1981:BMc

[1611] Anonymous. Back matter. Dædalus, 110(3):??, Summer 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024747.

Anonymous:1981:FMc

[1612] Anonymous. Front matter. Dædalus, 110(3):??, Summer 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024736.

G:1981:PIAc

[1613] S. R. G. Preface to the issue, “America’s Schools: Portraits and Perspectives”. Dædalus, 110(4):v–xvi, Fall 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024751.

Oates:1981:ISL

[1614] William A. Oates. Independent schools: Landscape and learnings. Dædalus, 110(4):1–16, Fall 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024752.

Lightfoot:1981:PESa

[1615] Sara Lawrence Lightfoot. Portraits of exemplary secondary schools: George Washington Carver Comprehensive High School. Dædalus, 110(4):17–37, Fall 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024753.

Jackson:1981:SSC

[1616] Philip W. Jackson. Secondary schooling for children of the poor. Dædalus, 110(4):39–57, Fall 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024754.

Lightfoot:1981:PESb

[1617] Sara Lawrence Lightfoot. Portraits of exemplary secondary schools: Highland Park. Dædalus, 110(4):59–80, Fall 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024755.

Jackson:1981:CWR

[1618] Philip W. Jackson. Comprehending a well-run comprehensive: a report on a visit to a large suburban high school. Dædalus, 110(4):81–95, Fall 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024756.

Lightfoot:1981:PESc

[1619] Sara Lawrence Lightfoot. Portraits of exemplary secondary schools: St. Paul’s school. Dædalus, 110(4):97–116, Fall 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024757.

Jackson:1981:SSP

[1620] Philip W. Jackson. Secondary schooling for the privileged few: a report on a visit to a new England boarding school. Dædalus, 110(4):117–130, Fall 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024758.

Coles:1981:NCS

[1621] Robert Coles. On the nature of character: Some preliminary field notes. Dædalus, 110(4):131–143, Fall 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024759.

Crim:1981:CB

[1622] Alonzo A. Crim. A community of believers. Dædalus, 110(4):145–162, Fall 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024760.

Healy:1981:BT

[1623] Timothy S. Healy. Belief and teaching. Dædalus, 110(4):163–175, Fall 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024761.

Anonymous:1981:BMd

[1624] Anonymous. Back matter. Dædalus, 110(4):??, Fall 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024762.

Anonymous:1981:FMd

[1625] Anonymous. Front matter. Dædalus, 110(4):??, Fall 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024750.

Anonymous:1981:VI

[1626] Anonymous. Volume information. Dædalus, 110(4):??, Fall 1981. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024749.

G:1982:PIRa

[1627] S. R. G. Preface to the issue, “Religion”. Dædalus, 111(1):v–xi, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024765.

Douglas:1982:EMR

[1628] Mary Douglas. The effects of modernization on religious change. Dædalus, 111(1):1–19, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024766.

Dupre:1982:SLS

[1629] Louis Dupré. Spiritual life in a secular age. Dædalus, 111(1):21–31, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024767.

Manuel:1982:IE

[1630] Frank E. Manuel. Israel and the enlightenment. Dædalus, 111(1):33–52, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024768.

Martin:1982:RDN

[1631] David Martin. Revived dogma and new cult. Dædalus, 111(1):53–71, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024769.

McBrien:1982:RCP

[1632] Richard P. McBrien. “Roman Catholicism”: E pluribus unum. Dædalus, 111(1):73–83, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024770.

Hebblethwaite:1982:PPS

[1633] Peter Hebblethwaite. The popes and politics: Shifting patterns in “Catholic Social Doctrine”. Dædalus, 111(1):85– 99, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024771.

Fischer:1982:IRP

[1634] Michael M. J. Fischer. Islam and the revolt of the petit bourgeoisie. Dædalus, 111(1):101–125, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024772.

Kelly:1982:FFD

[1635] George Armstrong Kelly. Faith, freedom, and disenchantment: Politics and the American religious consciousness. Dædalus, 111(1):127–148, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024773.

Marty:1982:RAS

[1636] Martin E. Marty. Religion in America since mid-century. Dædalus, 111 (1):149–163, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024774.

Roof:1982:AVE

[1637] Wade Clark Roof. America’s voluntary establishment: Mainline religion in transition. Dædalus, 111 (1):165–184, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024775.

Tipton:1982:MLA

[1638] Steven M. Tipton. The moral logic of alternative religions. Dædalus, 111 (1):185–213, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024776.

Anthony:1982:SIC

[1639] Dick Anthony and Thomas Robbins. Spiritual innovation and the crisis of American civil religion. Dædalus, 111 (1):215–234, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024777.

Anonymous:1982:BMa

[1640] Anonymous. Back matter. Dædalus, 111(1):??, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024778.

Anonymous:1982:FMa

[1641] Anonymous. Front matter. Dædalus, 111(1):??, Winter 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024764.

G:1982:PIB

[1642] S. R. G. Preface to the issue, “Black Africa: a Generation after Independence”. Dædalus, 111(2):v–xvi, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024781.

Ajayi:1982:EI

[1643] J. F. Ade Ajayi. Expectations of independence. Dædalus, 111(2):1–9, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024782.

Sylla:1982:SCL

[1644] Lanciné Sylla and Arthur Goldhammer. Succession of the charismatic leader: The Gordian knot of African politics. Dædalus, 111(2):11–28, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024783.

Mphahlele:1982:AE

[1645] Es’Kia Mphahlele. Africa in exile. Dædalus, 111(2):29–48, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024784.

Vansina:1982:MT

[1646] Jan Vansina. Mwasi’s trials. Dædalus, 111(2):49–70, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024785.

Young:1982:PSC

[1647] Crawford Young. Patterns of social conflict: State, class, and ethnicity. Dædalus, 111(2):71–98, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024786.

Leys:1982:AED

[1648] Colin Leys. African economic development in theory and practice. Dædalus, 111(2):99–124, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024787.

Roemer:1982:EDA

[1649] Michael Roemer. Economic development in Africa: Performance since independence, and a strategy for the future. Dædalus, 111(2):125–148, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024788.

Kamarck:1982:RTA

[1650] Andrew M. Kamarck. The resources of tropical Africa. Dædalus, 111 (2):149–163, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024789.

Ayandele:1982:ACH

[1651] Emmanuel A. Ayandele. Africa: The challenge of higher education. Dædalus, 111(2):165–177, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024790.

Gregory:1982:APR

[1652] Joel W. Gregory and Victor Piché. African population: Reproduction for whom? Dædalus, 111(2):179–209, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024791.

Johnson:1982:OCA

[1653] Willard R. Johnson and Ernest J. Wilson III. The “Oil Crises” and African economies: Oil wave on a tidal flood of industrial price inflation. Dædalus, 111(2):211–241, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024792.

Akinyemi:1982:ACR

[1654] A. Bolaji Akinyemi. Africa: Challenges and responses: a foreign policy perspective. Dædalus, 111(2):243–254, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024793.

Bull:1982:WSA

[1655] Hedley Bull. The West and South Africa. Dædalus, 111(2):255–270, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024794.

Anonymous:1982:BMb

[1656] Anonymous. Back matter. Dædalus, 111(2):??, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024795.

Anonymous:1982:FMb

[1657] Anonymous. Front matter. Dædalus, 111(2):??, Spring 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024780.

G:1982:PIRb

[1658] S. R. G. Preface to the issue, “Representations and Realities”. Dædalus, 111(3):v–vi, Summer 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024798.

Greenblatt:1982:FR

[1659] Stephen Greenblatt. Filthy rites. Dædalus, 111(3):1–16, Summer 1982. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20024799.

Shklar:1982:PCF

[1660] Judith N. Shklar. Putting cruelty first. Dædalus, 111(3):17–27, Summer 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024800.

Schorske:1982:MKS

[1661] Carl E. Schorske. Mahler and Klimt: Social experience and artistic evolution. Dædalus, 111(3):29–50, Summer 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024801.

Clive:1982:VSB

[1662] John Clive. The versatile Signor Di Bassetto: an essay in crime prevention. Dædalus, 111(3):51–63, Summer 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024802.

Darnton:1982:WHB

[1663] Robert Darnton. What is the history of books? Dædalus, 111 (3):65–83, Summer 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024803.

Pelikan:1982:TCD

[1664] Jaroslav Pelikan. The two cities: The decline and fall of Rome as historical paradigm. Dædalus, 111 (3):85–91, Summer 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024804.

OFlaherty:1982:DNI

[1665] Wendy Doniger O’Flaherty. The dream narrative and the Indian doctrine of illusion. Dædalus, 111(3):93–113, Summer 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024805.

Kastan:1982:SWW

[1666] David Scott Kastan. Shakespeare and “The Way of Womenkind”. Dædalus, 111(3):115–130, Summer 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024806.

Hollander:1982:WWM

[1667] John Hollander. The West wind and the mingled measure. Dædalus, 111 (3):131–148, Summer 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024807.

Boly:1982:ART

[1668] John R. Boly. Auden and the romantic tradition in “The Age of Anxiety”. Dædalus, 111(3):149–171, Summer 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024808.

Anonymous:1982:BMc

[1669] Anonymous. Back matter. Dædalus, 111(3):??, Summer 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024809.

Anonymous:1982:FMc

[1670] Anonymous. Front matter. Dædalus, 111(3):??, Summer 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024797.

G:1982:PIP

[1671] S. R. G. Preface to the issue, “Print Culture and Video Culture”. Dædalus, 111(4):v–vii, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024813.

Smith:1982:ITM

[1672] Anthony Smith. Information technology and the myth of abundance. Dædalus, 111(4):1–16, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024814.

Pool:1982:CEP

[1673] Ithiel de Sola Pool. The culture of electronic print. Dædalus, 111(4):17–31, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024815.

Booth:1982:CWK

[1674] Wayne C. Booth. The company we keep: Self-making in imaginative art, old and new. Dædalus, 111(4):33–59, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024816.

Poirier:1982:LTP

[1675] Richard Poirier. Literature, technology, people. Dædalus, 111(4):61–74, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024817.

Cavell:1982:FT

[1676] Stanley Cavell. The fact of television. Dædalus, 111(4):75–96, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024818.

Schudson:1982:PNF

[1677] Michael Schudson. The politics of narrative form: The emergence of news conventions in print and television. Dædalus, 111(4):97–112, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024819.

Juergens:1982:TRP

[1678] George Juergens. Theodore Roosevelt and the Press. Dædalus, 111(4):113– 133, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024820.

Marwick:1982:PPS

[1679] Arthur Marwick. Print, pictures, and sound: The second world war and the British experience. Dædalus, 111(4):135–155, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024821.

Mandelbaum:1982:VTW

[1680] Michael Mandelbaum. Vietnam: The television war. Dædalus, 111(4):157– 169, Fall 1982. CODEN DAEDAU.

ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20024822.

Winston:1982:RCE

[1681] Michael R. Winston. Racial consciousness and the evolution of mass communications in the United States. Dædalus, 111(4):171–182, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024823.

LaFollette:1982:STI

[1682] Marcel C. LaFollette. Science on television: Influences and strategies. Dædalus, 111(4):183–197, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024824.

Shapley:1982:MNS

[1683] Deborah Shapley. The media and national security. Dædalus, 111(4):199– 209, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20024825.

Anonymous:1982:BMd

[1684] Anonymous. Back matter. Dædalus, 111(4):??, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024826.

Anonymous:1982:FMd

[1685] Anonymous. Front matter. Dædalus, 111(4):??, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024812.

Anonymous:1982:VI

[1686] Anonymous. Volume information. Dædalus, 111(4):??, Fall 1982. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024811.

G:1983:PIR

[1687] S. R. G. Preface to the issue:

“Reading: Old and New”. Dædalus, 112(1):v–xii, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024829.

Kermode:1983:CR

[1688] Frank Kermode. The common reader. Dædalus, 112(1):1–11, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024830.

Hardwick:1983:R

[1689] Elizabeth Hardwick. Reading. Dædalus, 112(1):13–18, Winter 1983. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20024831.

Szladits:1983:AQ

[1690] Lola L. Szladits. Answers and questions. Dædalus, 112(1):19–26, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024832.

Aaron:1983:ILE

[1691] Daniel Aaron. An informal letter to the Editor. Dædalus, 112(1):27–33, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024833.

Silverman:1983:FP

[1692] Al Silverman. The fragile pleasure. Dædalus, 112(1):35–49, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024834.

Powell:1983:WLB

[1693] Walter W. Powell. Whither the local bookstore? Dædalus, 112(1):51–64, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024835.

Goodman:1983:TAR

[1694] William B. Goodman. Thinking about readers. Dædalus, 112(1):65–84, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024836.

Vaughan:1983:CB

[1695] Samuel S. Vaughan. The community of the book. Dædalus, 112 (1):85–115, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024837.

Lacy:1983:RAE

[1696] Dan Lacy. Reading in an audiovisual and electronic era. Dædalus, 112 (1):117–127, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024838.

Compaine:1983:NL

[1697] Benjamin M. Compaine. The new literacy. Dædalus, 112(1):129–142, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024839.

Starr:1983:ER

[1698] Paul Starr. The electronic reader.

Dædalus, 112(1):143–156, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024840.

Conrad:1983:EEL

[1699] Peter Conrad. The Englishness of English literature. Dædalus, 112 (1):157–173, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024841.

Fish:1983:SPG

[1700] Stanley Fish. Short people got no reason to live: Reading irony. Dædalus, 112(1):175–191, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024842.

Booth:1983:NSE

[1701] Wayne C. Booth. A new strategy for establishing a truly democratic criticism. Dædalus, 112(1):193–214, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024843.

Goodheart:1983:TIC

[1702] Eugene Goodheart. The text and the interpretive community. Dædalus, 112 (1):215–231, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024844.

Fish:1983:REG

[1703] Stanley Fish. A reply to Eugene Goodheart. Dædalus, 112(1):233–237, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024845.

Kibel:1983:CT

[1704] Alvin C. Kibel. The canonical text. Dædalus, 112(1):239–254, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024846.

Anonymous:1983:BMa

[1705] Anonymous. Back matter. Dædalus, 112(1):??, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024847.

Anonymous:1983:FMa

[1706] Anonymous. Front matter. Dædalus, 112(1):??, Winter 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024828.

G:1983:PIS

[1707] S. R. G. Preface to the issue: “Scientific Literacy”. Dædalus, 112(2):v–vii, Spring 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024850.

Walberg:1983:SLE

[1708] Herbert J. Walberg. Scientific literacy and economic productivity in international perspective. Dædalus, 112(2):1–28, Spring 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024851.

Miller:1983:SLC

[1709] Jon D. Miller. Scientific literacy: a conceptual and empirical review. Dædalus, 112(2):29–48, Spring 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024852.

Prewitt:1983:SID

[1710] Kenneth Prewitt. Scientific illiteracy and democratic theory. Dædalus, 112(2):49–64, Spring 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024853.

Hawkins:1983:NCO

[1711] David Hawkins. Nature closely observed. Dædalus, 112(2):65–89, Spring 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024854.

Arons:1983:AWS

[1712] A. B. Arons. Achieving wider scientific literacy. Dædalus, 112(2):91–122, Spring 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024855.

Rowe:1983:SEF

[1713] Mary Budd Rowe. Science education: a framework for decisionmakers. Dædalus, 112(2):123–142, Spring 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024856.

Jackson:1983:RSE

[1714] Philip W. Jackson. The reform of science education: a cautionary tale. Dædalus, 112(2):143–166, Spring 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024857.

Atkin:1983:IST

[1715] J. Myron Atkin. The improvement of science teaching. Dædalus, 112 (2):167–187, Spring 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024858.

Childs:1983:HGO

[1716] Barton Childs and Faith Hickman. Human genetics: One approach to scientific literacy. Dædalus, 112(2):189–209, Spring 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024859.

Kemeny:1983:CCL

[1717] John G. Kemeny. The case for computer literacy. Dædalus, 112(2):211– 230, Spring 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024860.

Graubard:1983:NFM

[1718] Stephen R. Graubard. Nothing to fear, much to do. Dædalus, 112 (2):231–248, Spring 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024861.

Anonymous:1983:BMb

[1719] Anonymous. Back matter. Dædalus, 112(2):??, Spring 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024862.

Anonymous:1983:FMb

[1720] Anonymous. Front matter. Dædalus, 112(2):??, Spring 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024849.

G:1983:PIA

[1721] S. R. G. Preface to the issue, “The Arts and Humanities in America’s Schools”. Dædalus, 112(3):v–xiv, Summer 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024865.

Barone:1983:TUT

[1722] Thomas Barone. Things of use and things of beauty: The Swain County High School Arts Program. Dædalus, 112(3):1–28, Summer 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024866.

Amarel:1983:ECA

[1723] Marianne Amarel. An education for commercial artists. Dædalus, 112 (3):29–58, Summer 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024867.

Jones:1983:FSS

[1724] Lonna Bush Jones. Five small schools leave home for Fillmore Arts Center. Dædalus, 112(3):59–79, Summer 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024868.

Donmoyer:1983:PPM

[1725] Robert Donmoyer. The principal as prime mover. Dædalus, 112(3):81–94, Summer 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024869.

Durden:1983:LEE

[1726] William G. Durden. Lessons for excellence in education. Dædalus, 112 (3):95–111, Summer 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024870.

Barylick:1983:BAI

[1727] Martha Barylick. Both artist and instrument: an approach to dance education. Dædalus, 112(3):113–127, Summer 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024871.

Murray:1983:ACQ

[1728] Jon J. Murray. Art, creativity, and the quality of education. Dædalus, 112 (3):129–147, Summer 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024872.

Esielonis:1983:L

[1729] Karyn Esielonis. Libertyville. Dædalus, 112(3):149–160, Summer 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20024873.

Marius:1983:URS

[1730] Richard Marius. Unscientific ruminations on schools in libertyville. Dædalus, 112(3):161–189, Summer 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024874.

Graubard:1983:COS

[1731] Stephen R. Graubard. Confronting the obvious: Social class and its devastating effects on American schooling. Dædalus, 112(3):191–209, Summer 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024875.

Evans:1983:THL

[1732] Paula M. Evans. Teaching history in libertyville. Dædalus, 112(3):211–228, Summer 1983. CODEN DAEDAU.

ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20024876.

Mueller:1983:TBL

[1733] Marlies Mueller. The tower of Babel in libertyville. Dædalus, 112 (3):229–247, Summer 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024877.

Anonymous:1983:BMc

[1734] Anonymous. Back matter. Dædalus, 112(3):??, Summer 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024878.

Anonymous:1983:FMc

[1735] Anonymous. Front matter. Dædalus, 112(3):??, Summer 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024864.

G:1983:PIH

[1736] S. R. G. Preface to the issue: “Human Rights”. Dædalus, 112(4):v–x, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024882.

Cranston:1983:THR

[1737] Maurice Cranston. Are there any human rights? Dædalus, 112(4):1–17, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024883.

Hoffmann:1983:RMD

[1738] Stanley Hoffmann. Reaching for the most difficult: Human rights as a foreign policy goal. Dædalus, 112(4):19– 49, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024884.

Rimlinger:1983:CHR

[1739] Gaston V. Rimlinger. Capitalism and human rights. Dædalus, 112(4):51–79, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024885.

Kolakowski:1983:MHR

[1740] Leszek Kolakowski. Marxism and human rights. Dædalus, 112(4):81–92, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024886.

Ruggie:1983:HRF

[1741] John Gerard Ruggie. Human rights and the future international community. Dædalus, 112(4):93–110, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024887.

Goldman:1983:HRP

[1742] Merle Goldman. Human rights in the People’s Republic of China. Dædalus, 112(4):111–138, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024888.

Farer:1983:HRH

[1743] Tom J. Farer. Human rights and human welfare in Latin America. Dædalus, 112(4):139–170, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024889.

Weinstein:1983:HRD

[1744] Warren Weinstein. Human rights and development in Africa: Dilemmas and options. Dædalus, 112(4):171–196, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024890.

Days:1983:SNC

[1745] Drew S. Days III. Seeking a new civil rights consensus. Dædalus, 112 (4):197–215, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024891.

Olson:1983:LIW

[1746] Mancur Olson. A less ideological way of deciding how much should be given to the poor. Dædalus, 112(4):217–236, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024892.

Tracy:1983:RHR

[1747] David Tracy. Religion and human rights in the public realm. Dædalus, 112(4):237–254, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024893.

Henrich:1983:CAS

[1748] Dieter Henrich and David S. Pacini.

The contexts of autonomy: Some presuppositions of the comprehensibility of human rights. Dædalus, 112(4):255–277, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024894.

Anonymous:1983:BMd

[1749] Anonymous. Back matter. Dædalus, 112(4):??, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024895.

Anonymous:1983:FMd

[1750] Anonymous. Front matter. Dædalus, 112(4):??, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024881.

Anonymous:1983:VI

[1751] Anonymous. Volume information. Dædalus, 112(4):??, Fall 1983. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024880.

G:1984:PINa

[1752] S. R. G. Preface to the issue, “The Nordic Enigma”. Dædalus, 113 (1):v–xiii, Winter 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024898.

Mead:1984:NDF

[1753] W. R. Mead. Norden: Destiny and fortune. Dædalus, 113(1):1–27, Winter 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024899.

McFate:1984:SEA

[1754] Patricia Bliss McFate. To see everything in another light. Dædalus, 113(1):29–59, Winter 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024900.

Enquist:1984:AFB

[1755] Per Olov Enquist and Verne Moberg. On the art of flying backward with dignity. Dædalus, 113(1):61–73, Winter 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024901.

Zetterberg:1984:RH

[1756] Hans L. Zetterberg. The rational humanitarians. Dædalus, 113(1):75–92, Winter 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024902.

Dahl:1984:EF

[1757] Hans F. Dahl. Those equal folk. Dædalus, 113(1):93–107, Winter 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024903.

Andersen:1984:RIN

[1758] Bent Rold Andersen. Rationality and irrationality of the Nordic welfare state. Dædalus, 113(1):109–139, Winter 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024904.

Skard:1984:EBS

[1759] Torild Skard and Elina Haavio-Mannila. Equality between the sexes: Myth or reality in norden? Dædalus, 113(1):141–167, Winter 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

/www.jstor.org/stable/20024905.

Allardt:1984:RGB

[1760] Erik Allardt. Representative government in a bureaucratic age. Dædalus, 113(1):169–197, Winter 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024906.

Gislason:1984:DSN

[1761] Gylfi Gislason. In defense of small nations. Dædalus, 113(1):199–211, Winter 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024907.

Anonymous:1984:BMa

[1762] Anonymous. Back matter. Dædalus, 113(1):??, Winter 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024908.

Anonymous:1984:FMa

[1763] Anonymous. Front matter. Dædalus, 113(1):??, Winter 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024897.

G:1984:PINb

[1764] S. R. G. Preface to the issue, “Nordic Voices”. Dædalus, 113(2):v–vii, Spring 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024911.

Anonymous:1984:NV

[1765] Anonymous. Nordic voices. Dædalus, 113(2):1–51, Spring 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024912.

Ruth:1984:SNN

[1766] Arne Ruth. The second new nation: The mythology of modern Sweden. Dædalus, 113(2):53–96, Spring 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024913.

Gustafsson:1984:SN

[1767] Madeleine Gustafsson and Verne Moberg. The silences of the North. Dædalus, 113(2):97–105, Spring 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024914.

Lonnroth:1984:ICS

[1768] Lars Lönnroth. The intellectual civil servant: The role of the writer and the scholar in Nordic culture. Dædalus, 113(2):107–136, Spring 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024915.

Rehn:1984:WS

[1769] Gösta Rehn. The wages of success. Dædalus, 113(2):137–168, Spring 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024916.

Olsen:1984:DSD

[1770] Erling Olsen. The dilemma of the social-democratic labor parties. Dædalus, 113(2):169–194, Spring 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024917.

Holst:1984:PNS

[1771] Johan Jørgen Holst. The pattern of Nordic security. Dædalus, 113 (2):195–225, Spring 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024918.

Hagtvet:1984:SAD

[1772] Bernt Hagtvet and Erik Rudeng. Scandinavia: Achievements, dilemmas, challenges. Dædalus, 113(2):227–256, Spring 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024919.

Klinge:1984:ANS

[1773] Matti Klinge. Aspects of the Nordic self. Dædalus, 113(2):257–277, Spring 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024920.

Anonymous:1984:BMb

[1774] Anonymous. Back matter. Dædalus, 113(2):??, Spring 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024921.

Anonymous:1984:FMb

[1775] Anonymous. Front matter. Dædalus, 113(2):??, Spring 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024910.

G:1984:PIA

[1776] S. R. G. Preface to the issue, “Anticipations”. Dædalus, 113(3):v–xii, Summer 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024924.

Bolter:1984:AI

[1777] J. David Bolter. Artificial intelligence. Dædalus, 113(3):1–18, Summer 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024925.

Meyrowitz:1984:ACC

[1778] Joshua Meyrowitz. The adultlike child and the childlike adult: Socialization in an electronic age. Dædalus, 113 (3):19–48, Summer 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024926.

Radway:1984:ICV

[1779] Janice Radway. Interpretive communities and variable literacies: The functions of romance reading. Dædalus, 113(3):49–73, Summer 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024927.

Gilligan:1984:CDC

[1780] Carol Gilligan. The conquistador and the dark continent: Reflections on the psychology of love. Dædalus, 113 (3):75–95, Summer 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024928.

McCloskey:1984:LCE

[1781] Donald N. McCloskey. The literary character of economics. Dædalus, 113 (3):97–119, Summer 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024929.

Brinkley:1984:WHC

[1782] Alan Brinkley. Writing the history of contemporary America: Dilemmas and challenges. Dædalus, 113(3): 121–141, Summer 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024930.

Marius:1984:MMA

[1783] Richard Marius. Musings on the mysteries of the American south. Dædalus, 113(3):143–176, Summer 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024931.

Weisskopf:1984:FLS

[1784] Victor F. Weisskopf. The frontiers and limits of science. Dædalus, 113 (3):177–195, Summer 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024932.

Root-Bernstein:1984:CPU

[1785] Robert Scott Root-Bernstein. Creative process as a unifying theme of human cultures. Dædalus, 113(3): 197–219, Summer 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024933.

Hoffmann:1984:CWT

[1786] Stanley Hoffmann. Cries and whimpers: Thoughts on West European– American relations in the 1980s. Dædalus, 113(3):221–252, Summer 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20024934.

Anonymous:1984:BMc

[1787] Anonymous. Back matter. Dædalus, 113(3):??, Summer 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024935.

Anonymous:1984:FMc

[1788] Anonymous. Front matter. Dædalus, 113(3):??, Summer 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024923.

G:1984:PIV

[1789] S. R. G. Preface to the issue, “Values, Resources, and Politics in America’s Schools”. Dædalus, 113(4):v–xi, Fall 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024939.

Holton:1984:NRR

[1790] Gerald Holton. A nation at risk revisited. Dædalus, 113(4):1–27, Fall 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024940.

Graham:1984:SCA

[1791] Patricia Albjerg Graham. Schools: Cacophony about practice, silence about purpose. Dædalus, 113(4):29–57, Fall 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024941.

Lazerson:1984:LCA

[1792] Marvin Lazerson, Judith Block McLaughlin, and Bruce McPherson. Learning and citizenship: Aspirations for American education. Dædalus, 113(4):59–74, Fall 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024942.

Graubard:1984:ZCC

[1793] Stephen R. Graubard. Zeal, cunning, candor, and persistence: To what educational ends? Dædalus, 113(4):75–106, Fall 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024943.

Eckstein:1984:CID

[1794] Harry Eckstein. Civic inclusion and its discontents. Dædalus, 113(4):107–145, Fall 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024944.

Burnham:1984:CHE

[1795] Walter Dean Burnham. A commentary on Harry Eckstein. Dædalus, 113(4):147–160, Fall 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024945.

Wolfthal:1984:WTW

[1796] Maurice Wolfthal. The way they were: Students in the Golden age. Dædalus, 113(4):161–168, Fall 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024946.

Glazer:1984:SVM

[1797] Nathan Glazer. Some very Modest proposals for the improvement of American education. Dædalus, 113(4):169– 176, Fall 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20024947.

Califano:1984:PE

[1798] Joseph A. Califano, Jr., Terrel H. Bell, Albert Shanker, Shirley Williams, Joseph S. Murphy, Charlotte Frank, John Kominski, John Brademas, Kenneth B. Clark, Roderick V. O. Boggs, Eugene Kinlow, Timothy S. Healy, James F. Hawker, John R. Silber, Charles Chatfield, Robert Baker, Derek Roberts, Ann Dominic Roach, and Eugene Sullivan. Perspectives on education. Dædalus, 113(4):177–287, Fall 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024948.

Anonymous:1984:BMd

[1799] Anonymous. Back matter. Dædalus, 113(4):??, Fall 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024949.

Anonymous:1984:FMd

[1800] Anonymous. Front matter. Dædalus, 113(4):??, Fall 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024938.

Anonymous:1984:VI

[1801] Anonymous. Volume information. Dædalus, 113(4):??, Fall 1984. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024937.

G:1985:PIA

[1802] S. R. G. Preface to the issue: “Australia: Terra Incognita?”. Dædalus, 114(1):v–xii, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024952.

Wright:1985:LD

[1803] Judith Wright. Landscape and dreaming. Dædalus, 114(1):29–56, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024954.

Clark:1985:H

[1804] Manning Clark. Heroes. Dædalus, 114(1):57–84, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024955.

Inglis:1985:CCL

[1805] K. S. Inglis. Ceremonies in a capital landscape: Scenes in the making of Canberra. Dædalus, 114(1):85–126, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024956.

Cowen:1985:OGG

[1806] Zelman Cowen. The Office of governorgeneral. Dædalus, 114(1):127–146, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024957.

Collins:1985:PIA

[1807] Hugh Collins. Political ideology in Australia: The distinctiveness of a Benthamite society. Dædalus, 114 (1):147–169, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024958.

Horne:1985:WRA

[1808] Donald Horne. Who rules Australia? Dædalus, 114(1):171–196, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024959.

Stretton:1985:QLA

[1809] Hugh Stretton. The quality of leading Australians. Dædalus, 114(1):197–230, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024960.

Jackson:1985:AE

[1810] Gordon Jackson. The Australian economy. Dædalus, 114(1):231–257, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024961.

Millar:1985:DA

[1811] T. B. Millar. The defense of Australia. Dædalus, 114(1):259–279, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024962.

Williams:1985:WIE

[1812] Bruce Williams. Wealth, invention, and education. Dædalus, 114(1):281–291, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024963.

Kramer:1985:MSC

[1813] Leonie Kramer. The media, society, and culture. Dædalus, 114(1):293–310, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024964.

Jose:1985:CIT

[1814] Nicholas Jose. Cultural identity: “I Think I’m Something Else”. Dædalus, 114(1):311–342, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024965.

Conway:1985:GA

[1815] Jill Conway. Gender in Australia. Dædalus, 114(1):343–368, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024966.

Davie:1985:FR

[1816] Michael Davie. The fraying of the rope. Dædalus, 114(1):369–393, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024967.

Albinski:1985:AUS

[1817] Henry S. Albinski. Australia and the United States. Dædalus, 114 (1):395–420, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024968.

Walsh:1985:AO

[1818] Richard Walsh. Australia observed. Dædalus, 114(1):421–438, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024969.

Blainey:1985:ABE

[1819] Geoffrey Blainey. Australia: a bird’seye view. Dædalus, 114(1):XIV 1–27, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024953.

Anonymous:1985:BMa

[1820] Anonymous. Back matter. Dædalus, 114(1):??, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024970.

Anonymous:1985:FMa

[1821] Anonymous. Front matter. Dædalus, 114(1):??, Winter 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024951.

G:1985:PIWa

[1822] S. R. G. Preface to the issue, “Weapons in Space,” Vol. I. Dædalus, 114(2):v–vii, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024973.

Long:1985:F

[1823] F. A. Long. Foreword. Dædalus, 114(2):9–14, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024974.

Anonymous:1985:GAa

[1824] Anonymous. Glossary of acronyms.

Dædalus, 114(2):15–16, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20024975.

York:1985:NDM

[1825] Herbert F. York. Nuclear deterrence and the military uses of space. Dædalus, 114(2):17–32, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024976.

Flax:1985:BMD

[1826] Alexander Flax. Ballistic missile defense: Concepts and history. Dædalus, 114(2):33–52, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024977.

Bethe:1985:BTC

[1827] Hans A. Bethe, Jeffrey Boutwell, and Richard L. Garwin. BMD technologies and concepts in the 1980s. Dædalus, 114(2):53–71, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.jstor.org/stable/20024978.

Special issue on the Strategic Defense Initiative. Volume 1. Concepts and technologies. Volume 2. Implications for security.

Yonas:1985:SDI

[1828] Gerold Yonas. The strategic defense initiative. Dædalus, 114(2):73–90, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024979.

Hafner:1985:APV

[1829] Donald L. Hafner. Assessing the President’s vision: The Fletcher, Miller, and Hoffman panels. Dædalus, 114 (2):91–107, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024980.

Zraket:1985:SDS

[1830] Charles A. Zraket. Strategic defense:

a systems perspective. Dædalus, 114 (2):109–126, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024981.

Stares:1985:USM

[1831] Paul Stares. U.S. and Soviet military space programs: A comparative assessment. Dædalus, 114(2):127–145, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024982.

Gottfried:1985:ASW

[1832] Kurt Gottfried and Richard Ned Lebow. Anti-satellite weapons: Weighing the risks. Dædalus, 114(2):147–170, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024983.

Carter:1985:RAB

[1833] Ashton B. Carter. The relationship of ASAT and BMD systems. Dædalus, 114(2):171–189, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024984.

Anonymous:1985:BMb

[1834] Anonymous. Back matter. Dædalus, 114(2):??, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024985.

Anonymous:1985:FMb

[1835] Anonymous. Front matter. Dædalus, 114(2):??, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024972.

G:1985:PIWb

[1836] S. R. G. Preface to the issue, “Weapons in Space,” vol. II. Dædalus, 114 (3):v–vii, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024988.

Anonymous:1985:GAb

[1837] Anonymous. Glossary of acronyms. Dædalus, 114(3):ix–x, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20024989.

Chayes:1985:SWL

[1838] Abram Chayes, Antonia Handler Chayes, and Eliot Spitzer. Space weapons: The legal context. Dædalus, 114(3):193–218, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20024990.

Toomay:1985:CBM

[1839] John C. Toomay. The case for ballistic missile defense. Dædalus, 114 (3):219–237, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20024991.

Rathjens:1985:BSI

[1840] George Rathjens and Jack Ruina. BMD and strategic instability. Dædalus, 114(3):239–255, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024992.

Holloway:1985:SDI

[1841] David Holloway. The strategic defense initiative and the Soviet union. Dædalus, 114(3):257–278, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024993.

Bertram:1985:SDW

[1842] Christoph Bertram. Strategic defense and the Western Alliance. Dædalus, 114(3):279–296, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024994.

Wilkinson:1985:FPS

[1843] John Wilkinson, T. B. Millar, and Marie-France Garaud. Foreign perspectives on the SDI. Dædalus, 114 (3):297–313, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024995.

Boutwell:1985:SUS

[1844] Jeffrey Boutwell and F. A. Long. The SDI and U.S. security. Dædalus, 114 (3):315–329, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024996.

Bethe:1985:WSA

[1845] Hans A. Bethe and Richard L. Garwin. Weapons in space: Appendix A:

New BMD technologies. Dædalus, 114 (3):331–368, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.jstor.org/stable/20024997.

Special issue on the Strategic Defense Initiative. Volume 1. Concepts and technologies. Volume 2. Implications for security.

Reagan:1985:ANP

[1846] Ronald Reagan. From the address to the nation by President Ronald Reagan March 23, 1983 “Peace and National Security”. Dædalus, 114 (3):369–371, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024998.

Weinberger:1985:RSD

[1847] Caspar W. Weinberger. From the Report of Secretary of Defense Caspar W. Weinberger to the Congress, February 4, 1985, on the FY 1986 Budget, FY 1987 Authorization Request, and FY 1986–90 Defense Program. Dædalus, 114(3):373–377, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20024999.

Anonymous:1985:TBU

[1848] Anonymous. Treaty between the United States of America and the union of Soviet socialist republics on the limitation of anti-ballistic missile systems. Dædalus, 114(3):379–384, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025000.

Anonymous:1985:ASC

[1849] Anonymous. Agreed statements, common understandings, and unilateral statements regarding the treaty between the United States of America and the union of Soviet socialist republics on the limitation of antiballistic missiles. Dædalus, 114(3): 385–391, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025001.

Anonymous:1985:PTB

[1850] Anonymous. Protocol to the treaty between the United States of America and the union of Soviet socialist republics on the limitation of antiballistic missile systems. Dædalus, 114 (3):392–394, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025002.

Anonymous:1985:BMc

[1851] Anonymous. Back matter. Dædalus, 114(3):??, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025003.

Anonymous:1985:FMc

[1852] Anonymous. Front matter. Dædalus, 114(3):??, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20024987.

G:1985:PIM

[1853] S. R. G. Preface to the issue, “The Moving Image”. Dædalus, 114(4):v– xiii, Fall 1985. CODEN DAEDAU.

ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20025007.

Smith:1985:IT

[1854] Anthony Smith. The influence of television. Dædalus, 114(4):1–15, Fall 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025008.

Tracey:1985:PCI

[1855] Michael Tracey. The poisoned chalice? International television and the idea of dominance. Dædalus, 114(4):17–56, Fall 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025009.

Ignatieff:1985:NSE

[1856] Michael Ignatieff. Is nothing sacred? The ethics of television. Dædalus, 114(4):57–78, Fall 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025010.

Carroll:1985:PM

[1857] Noël Carroll. The power of movies. Dædalus, 114(4):79–103, Fall 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025011.

Winston:1985:WTD

[1858] Brian Winston. A whole technology of dyeing: a note on ideology and the apparatus of the chromatic moving image. Dædalus, 114(4):105–123, Fall 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025012.

Dhondy:1985:KFI

[1859] Farrukh Dhondy. Keeping faith: Indian film and its world. Dædalus, 114(4):125–140, Fall 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025013.

Wechsler:1985:FA

[1860] Judith Wechsler. The filming of art. Dædalus, 114(4):141–159, Fall 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025014.

Eco:1985:IRB

[1861] Umberto Eco. Innovation and repetition: Between modern and postmodern aesthetics. Dædalus, 114 (4):161–184, Fall 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025015.

Miller:1985:MEH

[1862] Jonathan Miller. The Mind’s eye and the human eye. Dædalus, 114(4):185– 199, Fall 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20025016.

Shattuck:1985:WIT

[1863] Roger Shattuck. Words and images:

Thinking and translation. Dædalus, 114(4):201–214, Fall 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025017.

Anonymous:1985:BMd

[1864] Anonymous. Back matter. Dædalus, 114(4):??, Fall 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025018.

Anonymous:1985:FMd

[1865] Anonymous. Front matter. Dædalus, 114(4):??, Fall 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025006.

Anonymous:1985:VI

[1866] Anonymous. Volume information. Dædalus, 114(4):??, Fall 1985. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025005.

G:1986:PIAa

[1867] S. R. G. Preface to the issue, “The Aging Society”. Dædalus, 115(1):v–vii, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025021.

Pifer:1986:ISP

[1868] Alan Pifer and D. Lydia Bronte. Introduction: Squaring the pyramid. Dædalus, 115(1):1–11, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025022.

Achenbaum:1986:AAS

[1869] W. Andrew Achenbaum. America as an aging society: Myths and images. Dædalus, 115(1):13–30, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025023.

Neugarten:1986:AAS

[1870] Bernice L. Neugarten and Dail A.

Neugarten. Age in the aging society. Dædalus, 115(1):31–49, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025024.

Riley:1986:LSS

[1871] Matilda White Riley and John W. Riley, Jr. Longevity and social structure: The added years. Dædalus, 115(1):51–75, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025025.

Siegel:1986:DPL

[1872] Jacob S. Siegel and Cynthia M. Taeuber. Demographic perspectives on the long-lived society. Dædalus, 115 (1):77–117, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025026.

Hagestad:1986:ASC

[1873] Gunhild O. Hagestad. The aging society as a context for family life. Dædalus, 115(1):119–139, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025027.

Rossi:1986:SGA

[1874] Alice S. Rossi. Sex and gender in an aging society. Dædalus, 115 (1):141–169, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025028.

Richman:1986:CAS

[1875] Harold A. Richman and Matthew W. Stagner. Children in an aging society: Treasured resource or forgotten minority? Dædalus, 115(1):171–189, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025029.

Moody:1986:EAS

[1876] Harry R. Moody. Education in an aging society. Dædalus, 115(1):191–210, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025030.

Avorn:1986:MHG

[1877] Jerome L. Avorn. Medicine, health, and the geriatric transformation. Dædalus, 115(1):211–225, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025031.

Davis:1986:AHC

[1878] Karen Davis. Aging and the healthcare system: Economic and structural issues. Dædalus, 115(1):227–246, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025032.

Callahan:1986:AHC

[1879] Daniel Callahan. Adequate health care and an aging society: Are they morally compatible? Dædalus, 115 (1):247–267, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025033.

Morrison:1986:WRA

[1880] Malcolm H. Morrison. Work and retirement in an aging society. Dædalus, 115(1):269–293, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025034.

Palmer:1986:ECA

[1881] John L. Palmer and Stephanie G.

Gould. The economic consequences of an aging society. Dædalus, 115 (1):295–323, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025035.

Torres-Gil:1986:LMP

[1882] Fernando Torres-Gil. The latinization of a multigenerational population: Hispanics in an aging society. Dædalus, 115(1):325–348, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025036.

Gibson:1986:BAS

[1883] Rose C. Gibson. Blacks in an aging society. Dædalus, 115(1):349–371, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025037.

Pifer:1986:PPR

[1884] Alan Pifer. The public policy response to population aging. Dædalus, 115 (1):373–395, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025038.

Anonymous:1986:BMa

[1885] Anonymous. Back matter. Dædalus, 115(1):??, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025039.

Anonymous:1986:FMa

[1886] Anonymous. Front matter. Dædalus, 115(1):??, Winter 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025020.

G:1986:PIAb

[1887] S. R. G. Preface to the issue, “America’s Doctors, Medical Science, Medical Care”. Dædalus, 115(2):v–x, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025042.

Rogers:1986:EYM

[1888] David E. Rogers. The early years: The medical world in which Walsh Mcdermott trained. Dædalus, 115(2):1–18, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025043.

Cassell:1986:ICR

[1889] Eric J. Cassell. Ideas in conflict:

The rise and fall (and rise and fall) of new views of disease. Dædalus, 115(2):19–41, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025044.

Kilbourne:1986:EPB

[1890] Edwin D. Kilbourne. The emergence of the physician-basic scientist in America. Dædalus, 115(2):43–54, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025045.

Ebert:1986:MEP

[1891] Robert H. Ebert. Medical education at the peak of the era of experimental medicine. Dædalus, 115(2):55–81, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025046.

Beeson:1986:CRM

[1892] Paul B. Beeson. The changing role model, and the shift in power. Dædalus, 115(2):83–97, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025047.

Petersdorf:1986:MSR

[1893] Robert G. Petersdorf. Medical schools and research: Is the tail wagging the Dog? Dædalus, 115(2):99–118, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025048.

Blendon:1986:PCA

[1894] Robert J. Blendon. The problems of cost, access, and distribution of medical care. Dædalus, 115(2):119–135, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025049.

Cluff:1986:ARM

[1895] Leighton E. Cluff. America’s romance with medicine and medical science. Dædalus, 115(2):137–159, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025050.

Johnson:1986:RCR

[1896] Kenneth G. Johnson. Reaching out to the community: Responses by medicine. Dædalus, 115(2):161–173, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025051.

Deuschle:1986:CCM

[1897] Kurt W. Deuschle. Cross-cultural medicine: The Navajo Indians as case exemplar. Dædalus, 115(2):175–184, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025052.

Cassell:1986:CCI

[1898] Eric J. Cassell. The changing concept of the ideal physician. Dædalus, 115(2):185–208, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025053.

Rogers:1986:WWB

[1899] David E. Rogers. Where have we been?

Where are we going? Dædalus, 115 (2):209–229, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025054.

Anonymous:1986:BMb

[1900] Anonymous. Back matter. Dædalus, 115(2):??, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025055.

Anonymous:1986:FMb

[1901] Anonymous. Front matter. Dædalus, 115(2):??, Spring 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025041.

G:1986:PIAc

[1902] S. R. G. Preface to the issue, “Art and Science”. Dædalus, 115 (3):v–xvii, Summer 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025058.

Steinberg:1986:ASD

[1903] Leo Steinberg. Art and science: Do they need to be yoked? Dædalus, 115(3):1–16, Summer 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025059.

Gill:1986:PP

[1904] Susan P. Gill. The paradox of prediction. Dædalus, 115(3):17–48, Summer 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025060.

Crombie:1986:ESR

[1905] Alistair C. Crombie. Experimental science and the rational artist in early modern Europe. Dædalus, 115 (3):49–74, Summer 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025061.

Holton:1986:ASB

[1906] Gerald Holton. The advancement of science, and its burdens. Dædalus, 115 (3):75 77–104, Summer 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025062.

Gardner:1986:FTF

[1907] Howard Gardner. Freud in three frames: a cognitive-scientific approach to creativity. Dædalus, 115(3):105–134, Summer 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20025063.

Dennett:1986:ITV

[1908] Daniel C. Dennett. Information, technology, and the virtues of ignorance. Dædalus, 115(3):135–153, Summer 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025064.

Smith:1986:TII

[1909] Anthony Smith. Technology, identity, and the information machine. Dædalus, 115(3):155–169, Summer 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025065.

Cavell:1986:OAS

[1910] Stanley Cavell, Leon Cooper, Samuel Y.

Edgerton, Jr., and Victor F. Weisskopf. Observations on art and science. Dædalus, 115(3):171–189, Summer 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025066.

Anonymous:1986:BMc

[1911] Anonymous. Back matter. Dædalus, 115(3):??, Summer 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025067.

Anonymous:1986:FMc

[1912] Anonymous. Front matter. Dædalus, 115(3):??, Summer 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025057.

G:1986:PIF

[1913] S. R. G. Preface to the issue, “The Future of Opera”. Dædalus, 115(4):v–xv, Fall 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025071.

Askonas-Shepherd:1986:FO

[1914] Lies Askonas-Shepherd, Raffaello de Banfield, Fedora Barbieri, Humphrey Burton, John Cox, Bruce Crawford, Peter Diamand, August Everding, Hughes Gall, Ann Getty, Italo Gomez, Grey Gowrie, Ronald Grierson, George Harewood, Terence McEwen, Gian Carlo Menotti, Jonathan Miller, Claus Moser, Harold Prince, Harold Rosenthal, Julius Rudel, Aulis Sallinen, Beverly Sills, Patrick Smith, Elisabeth Söderström-Olow, John Tooley, Eva Wagner-Pasquier, George Weidenfeld, and Huw Wheldon. The future of opera. Dædalus, 115(4):1–92, Fall 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025072.

Leinsdorf:1986:OWC

[1915] Erich Leinsdorf. Opera: What constitutes longevity. Dædalus, 115(4):93– 118, Fall 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20025073.

Epstein:1986:VNY

[1916] Matthew Epstein, Kitty Carlisle Hart, and Harvey Lichtenstein. Voices from new York. Dædalus, 115(4):119–150, Fall 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025074.

Anonymous:1986:Ia

[1917] Anonymous. [introduction]. Dædalus, 115(4):152, Fall 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025075.

Chevrillon:1986:BOP

[1918] Olivier Chevrillon. The Bastille Opera Project. Dædalus, 115(4):153–164, Fall 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025076.

deSaintPulgent:1986:MBO

[1919] Maryvonne de Saint Pulgent. The myths of the Bastille Opera. Dædalus, 115(4):165–172, Fall 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025077.

Anonymous:1986:Ib

[1920] Anonymous. [introduction]. Dædalus, 115(4):174, Fall 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025078.

Boulez:1986:BRI

[1921] Pierre Boulez. Berlioz and the realm of the imaginary. Dædalus, 115(4):175– 184, Fall 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20025079.

Berlioz:1986:HTR

[1922] Hector Berlioz. How a tenor revolves around the public. Dædalus, 115(4):185–194, Fall 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025080.

Anonymous:1986:BMd

[1923] Anonymous. Back matter. Dædalus, 115(4):??, Fall 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025081.

Anonymous:1986:FMd

[1924] Anonymous. Front matter. Dædalus, 115(4):??, Fall 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025070.

Anonymous:1986:VI

[1925] Anonymous. Volume information. Dædalus, 115(4):??, Fall 1986. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025069.

G:1987:PIPa

[1926] S. R. G. Preface to the issue, “Philanthropy, Patronage, Politics”. Dædalus, 116(1):v–xx, Winter 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025084.

Karl:1987:FRC

[1927] Barry D. Karl and Stanley N. Katz. Foundations and ruling class elites. Dædalus, 116(1):1–40, Winter 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025085.

Sutton:1987:FFE

[1928] Francis X. Sutton. The Ford foundation: The early years. Dædalus, 116(1):41–91, Winter 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025086.

McCarthy:1987:CWC

[1929] Kathleen D. McCarthy. From cold war to cultural development: The international cultural activities of the Ford foundation, 1950–1980. Dædalus, 116(1):93–117, Winter 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025087.

Pifer:1987:PVC

[1930] Alan Pifer. Philanthropy, voluntarism, and changing times. Dædalus, 116 (1):119–131, Winter 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025088.

McCaughey:1987:NNW

[1931] Patrick McCaughey. Native and nomad: Winslow Homer and John Singer Sargent. Dædalus, 116(1):133–153, Winter 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025089.

Feingold:1987:PPP

[1932] Mordechai Feingold. Philanthropy, pomp, and patronage: Historical reflections upon the endowment of culture. Dædalus, 116(1):155–178, Winter 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025090.

Anonymous:1987:BMa

[1933] Anonymous. Back matter. Dædalus, 116(1):??, Winter 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025091.

Anonymous:1987:FMa

[1934] Anonymous. Front matter. Dædalus, 116(1):??, Winter 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025083.

G:1987:PIPb

[1935] S. R. G. Preface to the issue: “Past and Present”. Dædalus, 116(2):v–xxiii, Spring 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025094.

Lapidus:1987:GRS

[1936] Gail W. Lapidus. Gorbachev and the reform of the Soviet system. Dædalus, 116(2):1–30, Spring 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025095.

Smolar:1987:JPP

[1937] Aleksander Smolar. Jews as a Polish problem. Dædalus, 116(2):31–73, Spring 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025096.

Wei-ming:1987:IHV

[1938] Tu Wei-ming. Iconoclasm, holistic vision, and patient watchfulness: A personal reflection on the modern Chinese intellectual quest. Dædalus, 116 (2):75–94, Spring 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025097.

Treverton:1987:CAC

[1939] Gregory F. Treverton. Covert action: From “Covert” to overt. Dædalus, 116(2):95–123, Spring 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025098.

Manuel:1987:LR

[1940] Frank E. Manuel. Lovejoy revisited. Dædalus, 116(2):125–147, Spring 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025099.

Hoffmann:1987:NFR

[1941] Stanley Hoffmann. A note on the French revolution and the language of violence. Dædalus, 116(2):149–156, Spring 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025100.

Kelly:1987:ALP

[1942] George Armstrong Kelly. Alphonse De Lamartine: The poet in politics. Dædalus, 116(2):157–180, Spring 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025101.

Chen:1987:API

[1943] Lincoln C. Chen. The AIDS pandemic: an internationalist approach to disease control. Dædalus, 116 (2):181–195, Spring 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025102.

Anonymous:1987:BMb

[1944] Anonymous. Back matter. Dædalus, 116(2):??, Spring 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025103.

Anonymous:1987:FMb

[1945] Anonymous. Front matter. Dædalus, 116(2):??, Spring 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025093.

G:1987:PIF

[1946] S. R. G. Preface to the issue, “Futures”. Dædalus, 116(3):v–xix, Summer 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025106.

Bell:1987:WUS

[1947] Daniel Bell. The world and the United States in 2013. Dædalus, 116 (3):1–31, Summer 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025107.

Gilpin:1987:APP

[1948] Robert Gilpin. American policy in the post-Reagan era. Dædalus, 116 (3):33–67, Summer 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025108.

Allison:1987:CWA

[1949] Graham Allison and Albert Carnesale. Can the West accept da for an answer? Dædalus, 116(3):69–93, Summer 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025109.

Hall:1987:CLM

[1950] John A. Hall. Classical liberalism and the modern state. Dædalus, 116 (3):95–118, Summer 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025110.

Downs:1987:EDH

[1951] Anthony Downs. The evolution of democracy: How its axioms and institutional forms have been adapted to changing social forces. Dædalus, 116 (3):119–148, Summer 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025111.

White:1987:VJM

[1952] Merry White. The virtue of Japanese mothers: Cultural definitions of Women’s lives. Dædalus, 116(3): 149–163, Summer 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

/www.jstor.org/stable/20025112.

Schwartz:1987:D

[1953] Jacob T. Schwartz. Dreamworld.

Dædalus, 116(3):165–185, Summer 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025113.

Winston:1987:MB

[1954] Brian Winston. A mirror for Brunelleschi. Dædalus, 116(3):187–201, Summer 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20025114.

Anonymous:1987:BMc

[1955] Anonymous. Back matter. Dædalus, 116(3):??, Summer 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025115.

Anonymous:1987:FMc

[1956] Anonymous. Front matter. Dædalus, 116(3):??, Summer 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025105.

G:1987:PIL

[1957] S. R. G. Preface to the issue, “Learning about Women: Gender, Politics, and Power”. Dædalus, 116(4):v–xx, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025119.

Conway:1987:ICG

[1958] Jill K. Conway, Susan C. Bourque, and Joan W. Scott. Introduction: The concept of gender. Dædalus, 116(4):xxi–xxx, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025120.

Dietz:1987:CAF

[1959] Mary G. Dietz. Context is all: Feminism and theories of citizenship. Dædalus, 116(4):1–24, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025121.

Holtzman:1987:WPW

[1960] Elizabeth Holtzman and Shirley Williams. Women in the political world: Observations. Dædalus, 116(4): 25–33, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20025122.

Inglis:1987:MWW

[1961] Ken Inglis. Men, women, and war memorials: Anzac Australia. Dædalus, 116(4):35–59, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025123.

Fausto-Sterling:1987:SWB

[1962] Anne Fausto-Sterling. Society writes biology / biology constructs gender. Dædalus, 116(4):61–76, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025124.

Keller:1987:WSF

[1963] Evelyn Fox Keller. Women scientists and feminist critics of science. Dædalus, 116(4):77–91, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025125.

Scott:1987:HD

[1964] Joan W. Scott. History and difference. Dædalus, 116(4):93–118, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025126.

Slater:1987:PPS

[1965] Miriam Slater and Penina Migdal Glazer. Prescriptions for professional survival. Dædalus, 116(4):119–135, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025127.

Conway:1987:PPG

[1966] Jill K. Conway. Politics, pedagogy, and gender. Dædalus, 116(4):137–152, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025128.

Fox:1987:CVU

[1967] Robert Fox and Anna Guagnini. Classical values and useful knowledge: The problem of access to technical careers in modern Europe. Dædalus, 116(4):153–171, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025129.

Bourque:1987:TGD

[1968] Susan C. Bourque and Kay B. Warren. Technology, gender, and development. Dædalus, 116(4):173–197, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025130.

Degler:1987:RWA

[1969] Carl N. Degler. On rereading “The Woman in America”. Dædalus, 116 (4):199–210, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025131.

Anonymous:1987:BMd

[1970] Anonymous. Back matter. Dædalus, 116(4):??, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025132.

Anonymous:1987:FMd

[1971] Anonymous. Front matter. Dædalus, 116(4):??, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025118.

Anonymous:1987:VI

[1972] Anonymous. Volume information. Dædalus, 116(4):??, Fall 1987. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025117.

G:1988:PIA

[1973] S. R. G. Preface to the issue, “Artificial Intelligence”. Dædalus, 117 (1):v–viii, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025135.

Papert:1988:OAM

[1974] Seymour Papert. One AI or many? Dædalus, 117(1):1–14, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025136.

Dreyfus:1988:MMV

[1975] Hubert L. Dreyfus and Stuart E. Dreyfus. Making a mind versus modeling the brain: Artificial intelligence back at a branchpoint. Dædalus, 117(1):15–43, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025137.

Sokolowski:1988:NAI

[1976] Robert Sokolowski. Natural and artificial intelligence. Dædalus, 117 (1):45–64, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025138.

McCorduck:1988:AIA

[1977] Pamela McCorduck. Artificial intelligence: an Aperçu. Dædalus, 117 (1):65–83, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025139.

Cowan:1988:NNA

[1978] Jack D. Cowan and David H. Sharp. Neural nets and artificial intelligence. Dædalus, 117(1):85–121, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025140.

Schwartz:1988:NCD

[1979] Jacob T. Schwartz. The new connectionism: Developing relationships between neuroscience and artificial intelligence. Dædalus, 117(1):123–141, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025141.

Reeke:1988:RBA

[1980] George N. Reeke, Jr. and Gerald M. Edelman. Real brains and artificial intelligence. Dædalus, 117(1):143–173, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025142.

Hillis:1988:IEB

[1981] W. Daniel Hillis. Intelligence as an emergent behavior; or, the songs of Eden. Dædalus, 117(1):175–189, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025143.

Waltz:1988:PBT

[1982] David L. Waltz. The prospects for building truly intelligent machines. Dædalus, 117(1):191–212, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025144.

Hurlbert:1988:MMA

[1983] Anya Hurlbert and Tomaso Poggio.

Making machines (and artificial intelligence) see. Dædalus, 117(1):213–239, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025145.

Turkle:1988:AIP

[1984] Sherry Turkle. Artificial intelligence and psychoanalysis: a new alliance. Dædalus, 117(1):241–268, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025146.

Putnam:1988:MAA

[1985] Hilary Putnam. Much ado about not very much. Dædalus, 117(1):269–281, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025147.

Dennett:1988:WPE

[1986] Daniel C. Dennett. When philosophers encounter artificial intelligence.

Dædalus, 117(1):283–295, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025148.

McCarthy:1988:MLA

[1987] John McCarthy. Mathematical logic in artificial intelligence. Dædalus, 117 (1):297–311, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025149.

Anonymous:1988:BMa

[1988] Anonymous. Back matter. Dædalus, 117(1):??, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025150.

Anonymous:1988:FMa

[1989] Anonymous. Front matter. Dædalus, 117(1):??, Winter 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025134.

G:1988:PIR

[1990] S. R. G. Preface to the issue “Religion and Education”. Dædalus, 117(2):v–xiii, Spring 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025153.

Anonymous:1988:RE

[1991] Anonymous. Religion and education. Dædalus, 117(2):1–146, Spring 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025154.

Cobb:1988:ERP

[1992] John B. Cobb, Jr. Education and religious pluralism. Dædalus, 117 (2):147–149, Spring 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025155.

Callahan:1988:TBR

[1993] Daniel Callahan. Tension between religion and philosophy. Dædalus, 117 (2):151–152, Spring 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025156.

Lindbeck:1988:SHT

[1994] George Lindbeck. The search for habitable texts. Dædalus, 117(2):153–156, Spring 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025157.

Kiely:1988:RUC

[1995] Robert Kiely. Religion in (and out of)

the University curriculum. Dædalus, 117(2):157–159, Spring 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025158.

Healy:1988:CM

[1996] Timothy S. Healy. The contrary model. Dædalus, 117(2):161–162, Spring 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025159.

Tipton:1988:CSV

[1997] Steven M. Tipton. The Church as a school for virtue. Dædalus, 117 (2):163–175, Spring 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025160.

Twersky:1988:RT

[1998] Isadore Twersky. Random thoughts. Dædalus, 117(2):177–179, Spring 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025161.

OFlaherty:1988:CHR

[1999] Wendy Doniger O’Flaherty. The case for the history of religions. Dædalus, 117(2):181–186, Spring 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025162.

Anonymous:1988:BMb

[2000] Anonymous. Back matter. Dædalus, 117(2):??, Spring 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025163.

Anonymous:1988:FMb

[2001] Anonymous. Front matter. Dædalus, 117(2):??, Spring 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025152.

G:1988:PIT

[2002] S. R. G. Preface to the issue “Three Decades of Dædalus”. Dædalus, 117 (3):v–xii, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025166.

Erikson:1988:YFD

[2003] Erik H. Erikson. Youth: Fidelity and diversity. Dædalus, 117(3):1–24, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025167.

Rossi:1988:EBS

[2004] Alice S. Rossi. Equality between the sexes: an immodest proposal. Dædalus, 117(3):25–71, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025168.

Smith:1988:EU

[2005] John Maynard Smith. Eugenics and utopia. Dædalus, 117(3):73–92, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025169.

Milosz:1988:NP

[2006] Czeslaw Milosz. The novel in Poland. Dædalus, 117(3):93–96, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025170.

Bellah:1988:CRA

[2007] Robert N. Bellah. Civil religion in America. Dædalus, 117(3):97–118, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20025171.

Noonan:1988:IDH

[2008] John T. Noonan, Jr. Intellectual and demographic history. Dædalus, 117 (3):119–141, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025172.

Shattuck:1988:TH

[2009] Roger W. Shattuck. Thoughts on the humanities. Dædalus, 117(3): 143–149, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025173.

Holton:1988:RC

[2010] Gerald Holton. The roots of complementarity. Dædalus, 117(3):151–197, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20025174.

Manuel:1988:UAP

[2011] Frank E. Manuel. The use and abuse of psychology in history. Dædalus, 117 (3):199–225, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025175.

Paz:1988:EG

[2012] Octavio Paz. Eroticism and gastrosophy. Dædalus, 117(3):227–249, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025176.

Frye:1988:SAW

[2013] Northrop Frye. The search for acceptable words. Dædalus, 117(3): 251–272, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025177.

Jordan:1988:BHE

[2014] Vernon E. Jordan, Jr. Blacks and higher education: Some reflections. Dædalus, 117(3):273–281, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025178.

Lynn:1988:AAL

[2015] Kenneth S. Lynn. Adulthood in American literature. Dædalus, 117 (3):283–298, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025179.

Thomas:1988:STM

[2016] Lewis Thomas. On the science and technology of medicine. Dædalus, 117 (3):299–316, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025180.

Hardwick:1988:DM

[2017] Elizabeth Hardwick. Domestic manners. Dædalus, 117(3):317–332, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025181.

Baltimore:1988:LSB

[2018] David Baltimore. Limiting science: a Biologist’s perspective. Dædalus, 117 (3):333–344, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025182.

Starobinski:1988:AA

[2019] Jean Starobinski. The accuser and the accused. Dædalus, 117(3):345–370, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20025183.

Hoffmann:1988:FFH

[2020] Stanley Hoffmann. Fragments floating in the here and now. Dædalus, 117 (3):371–408, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025184.

Shklar:1988:LK

[2021] Judith N. Shklar. Learning without knowing. Dædalus, 117(3):409–437, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20025185.

Franklin:1988:LRE

[2022] John Hope Franklin. The land of room enough. Dædalus, 117(3):439–456, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20025186.

Douglas:1988:EMR

[2023] Mary Douglas. The effects of modernization on religious change. Dædalus, 117(3):457–484, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025187.

Enquist:1988:AFB

[2024] Per Olov Enquist and Verne Moberg.

On the art of flying backward with dignity. Dædalus, 117(3):485–498, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025188.

Anonymous:1988:BMc

[2025] Anonymous. Back matter. Dædalus, 117(3):??, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025189.

Anonymous:1988:FMc

[2026] Anonymous. Front matter. Dædalus, 117(3):??, Summer 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025165.

Anonymous:1988:IMC

[2027] Anonymous. [illustration]: Map of Canada. Dædalus, 117(4):vi–vii, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025193.

G:1988:PIS

[2028] S. R. G. Preface to the issue “In Search of Canada”. Dædalus, 117(4):ix–xii, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025194.

Kilbourn:1988:PKS

[2029] William Kilbourn. The peaceable kingdom still. Dædalus, 117(4):1–29, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025195.

Hare:1988:CL

[2030] F. Kenneth Hare. Canada: The land. Dædalus, 117(4):31–50, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025196.

Harney:1988:GHO

[2031] Robert F. Harney. “So Great a Heritage as Ours” immigration and the survival of the Canadian polity. Dædalus, 117(4):51–97, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025197.

Flaherty:1988:WRC

[2032] David H. Flaherty. Who rules Canada? Dædalus, 117(4):99–128, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025198.

Axworthy:1988:FSU

[2033] Lloyd Axworthy. The Federal system: an uncertain path. Dædalus, 117(4):129–153, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025199.

Evans:1988:WTC

[2034] Robert G. Evans. “We’ll Take Care of It for You” health care in the Canadian community. Dædalus, 117(4):155–189, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025200.

Kenney-Wallace:1988:PPE

[2035] Geraldine A. Kenney-Wallace and J. Fraser Mustard. From paradox to paradigm: The evolution of science and technology in Canada. Dædalus, 117(4):191–214, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

/www.jstor.org/stable/20025201.

Cook:1988:SUE

[2036] Eleanor Cook. “A Seeing and Unseeing in the Eye” Canadian literature and the sense of place. Dædalus, 117(4):215–235, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025202.

Artibise:1988:CUN

[2037] Alan F. J. Artibise. Canada as an urban nation. Dædalus, 117(4):237–264, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025203.

Pelletier:1988:QDS

[2038] Gérard Pelletier. Quebec: Different but in step with North America. Dædalus, 117(4):265–282, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025204.

Dion:1988:MQ

[2039] Léon Dion. The mystery of Quebec. Dædalus, 117(4):283–317, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025205.

Catley-Carlson:1988:ACV

[2040] Margaret Catley-Carlson. Aid: a Canadian vocation. Dædalus, 117(4):319– 333, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20025206.

Begin:1988:DSR

[2041] Monique Bégin. Debates and silences: Reflections of a politician. Dædalus, 117(4):335–362, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025207.

Macdonald:1988:TPC

[2042] Donald S. Macdonald. Three perspectives on Canada’s future. Dædalus, 117(4):363–380, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025208.

Conway:1988:AOT

[2043] John Conway. An “Adapted Organic Tradition”. Dædalus, 117(4):381–396, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025209.

Anonymous:1988:BMd

[2044] Anonymous. Back matter. Dædalus, 117(4):??, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025210.

Anonymous:1988:FMd

[2045] Anonymous. Front matter. Dædalus, 117(4):??, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025192.

Anonymous:1988:VI

[2046] Anonymous. Volume information. Dædalus, 117(4):??, Fall 1988. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025191.

G:1989:PIW

[2047] S. R. G. Preface to the issue “A World to Make: Development in Perspective”. Dædalus, 118(1):ix–x, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025213.

Sutton:1989:II

[2048] Francis X. Sutton. Introduction to the issue. Dædalus, 118(1):xi–xviii, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025214.

McNeill:1989:CCH

[2049] William H. McNeill and Charles P.

Kindleberger. Control and catastrophe in human affairs [with comments]. Dædalus, 118(1):1–15, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025215.

Low:1989:DCA

[2050] D. Anthony Low, Peter D. Bell, Michael Roemer, and Charles William Maynes. Development’s contexts: Asia, Africa [with comments]. Dædalus, 118(1):17–33, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025216.

Sutton:1989:DIE

[2051] Francis X. Sutton, Tom G. Kessinger, James P. Grant, and George Zeidenstein. Development ideology: Its emergence and decline [with comments]. Dædalus, 118(1):35–60, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025217.

Chenery:1989:PD

[2052] Hollis B. Chenery, James P. Grant, Susan P. Joekes, and Miguel Urrutia. Panel discussion. Dædalus, 118(1):61–67, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025218.

Lewis:1989:GNE

[2053] John P. Lewis, Cranford Pratt, Abiola Irele, Francis X. Sutton, Charles William Maynes, and Michael Roemer. Government and national economic development [with comments]. Dædalus, 118(1):69–88, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025219.

Birdsall:1989:TGH

[2054] Nancy Birdsall, James P. Grant, Arthur Kleinman, Lincoln C. Chen, and Kamla Chowdhry. Thoughts on good health and good government [with comments]. Dædalus, 118 (1):89–123, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025220.

Irele:1989:EAM

[2055] Abiola Irele, F. Champion Ward, and Robert A. LeVine. Education and access to modern knowledge [with comments]. Dædalus, 118(1):125–140, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025221.

Chowdhry:1989:PED

[2056] Kamla Chowdhry, Lincoln C. Chen, and Judith Tendler. Poverty, environment, development [with comments]. Dædalus, 118(1):141–158, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025222.

Naciri:1989:USD

[2057] Mohamed Naciri, Francis X. Sutton, Hollis B. Chenery, Clifford Geertz, and Nancy Birdsall. Urban systems and development strategies [with comments]. Dædalus, 118(1):159–181, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025223.

Krugman:1989:DCW

[2058] Paul R. Krugman, Edmar L. Bacha, Jeffrey D. Sachs, David E. Bell, and Nancy Birdsall. Developing countries in the world economy [with comments]. Dædalus, 118(1):183–210, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025224.

Soedjatmoko:1989:ERP

[2059] Soedjatmoko. Education relevant to People’s needs. Dædalus, 118(1):211– 218, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20025225.

Maynes:1989:CAC

[2060] Charles William Maynes, Alfred Stepan, Jeffrey D. Sachs, Stanley Heginbotham, Peter D. Bell, Cranford Pratt, and Judith Tendler. Comments on article by C. W. Maynes. Dædalus, 118(1):219–225, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

                                        /www.jstor.org/stable/20025226.

Sen:1989:CAA

[2061] Amartya Sen, Gelia T. Castillo, Lincoln C. Chen, Nancy Birdsall, and Arthur Kleinman. Comments on article by Amartya Sen. Dædalus, 118 (1):225–230, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025227.

Bell:1989:PS

[2062] David E. Bell, Edmar L. Bacha, Mc-George Bundy, Gelia T. Castillo, Clifford Geertz, George Zeidenstein, Cranford Pratt, Abiola Irele, Robert A. LeVine, Dwight H. Perkins, Judith Tendler, Peter D. Bell, Kenneth Prewitt, Alfred Stepan, John D. Gerhart, and Lincoln C. Chen. Plenary session. Dædalus, 118(1):231–249, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025228.

Anonymous:1989:BMa

[2063] Anonymous. Back matter. Dædalus, 118(1):??, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025229.

Anonymous:1989:FMa

[2064] Anonymous. Front matter. Dædalus, 118(1):??, Winter 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025212.

G:1989:PIL

[2065] S. R. G. Preface to the issue “Living with AIDS”. Dædalus, 118(2):v–xvi, Spring 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025232.

Rosenberg:1989:WEA

[2066] Charles E. Rosenberg. What is an epidemic? AIDS in historical perspective. Dædalus, 118(2):1–17, Spring 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025233.

Hamburg:1989:ACB

[2067] Margaret A. Hamburg and Anthony S. Fauci. AIDS: The challenge to biomedical research. Dædalus, 118(2):19–39, Spring 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025234.

Ron:1989:AUS

[2068] Aran Ron and David E. Rogers. AIDS in the United States: Patient care and politics. Dædalus, 118(2):41–58, Spring 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025235.

Friedland:1989:CCA

[2069] Gerald H. Friedland. Clinical care in the AIDS epidemic. Dædalus, 118(2):59–83, Spring 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025236.

Phoenix:1989:AA

[2070] Peter Phoenix. Alive with AIDS. Dædalus, 118(2):85–92, Spring 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20025237.

Fox:1989:PPP

[2071] Daniel M. Fox, Patricia Day, and Rudolf Klein. The power of professionalism: Policies for AIDS in Britain, Sweden, and the United States. Dædalus, 118(2):93–112, Spring 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025238.

Christakis:1989:RPI

[2072] Nicholas A. Christakis. Responding to a pandemic: International interests in AIDS control. Dædalus, 118 (2):113–134, Spring 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025239.

Farmer:1989:AHS

[2073] Paul Farmer and Arthur Kleinman.

AIDS as human suffering. Dædalus, 118(2):135–160, Spring 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025240.

May:1989:ETD

[2074] Robert M. May, Roy M. Anderson, and Sally M. Blower. The epidemiology and transmission dynamics of HIV–AIDS. Dædalus, 118(2):162–201, Spring 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025241.

Anonymous:1989:BMb

[2075] Anonymous. Back matter. Dædalus, 118(2):??, Spring 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025242.

Anonymous:1989:FMb

[2076] Anonymous. Front matter. Dædalus, 118(2):??, Spring 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025231.

G:1989:PLA

[2077] S. R. G. Preface to “Living with AIDS: Part II”. Dædalus, 118(3):v–viii, Summer 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025245.

Keniston:1989:II

[2078] Kenneth Keniston. Introduction to the issue. Dædalus, 118(3):ix–xxxii, Summer 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025246.

Haseltine:1989:PMC

[2079] William A. Haseltine. Prospects for the medical control of the AIDS epidemic. Dædalus, 118(3):1–21, Summer 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025247.

Zinberg:1989:SPA

[2080] Norman E. Zinberg. Social policy: AIDS and intravenous drug use. Dædalus, 118(3):23–46, Summer 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025248.

Gagnon:1989:DD

[2081] John H. Gagnon. Disease and desire. Dædalus, 118(3):47–77, Summer 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025249.

Bayer:1989:APR

[2082] Ronald Bayer. AIDS, privacy, and responsibility. Dædalus, 118(3):79–99, Summer 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20025250.

Kirby:1989:AL

[2083] Michael Kirby. AIDS and law.

Dædalus, 118(3):101–121, Summer 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025251.

Osborn:1989:PHP

[2084] June E. Osborn. Public health and the politics of AIDS prevention. Dædalus, 118(3):123–144, Summer 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025252.

Sapolsky:1989:ABB

[2085] Harvey M. Sapolsky. AIDS, blood banking, and the bonds of community. Dædalus, 118(3):145–163, Summer 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025253.

deZalduondo:1989:AAD

[2086] Barbara O. de Zalduondo, Gernard I.

Msamanga, and Lincoln C. Chen. AIDS in Africa: Diversity in the global pandemic. Dædalus, 118(3): 165–204, Summer 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025254.

Dalton:1989:AB

[2087] Harlon L. Dalton. AIDS in blackface. Dædalus, 118(3):205–227, Summer 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025255.

Ostrow:1989:APT

[2088] David G. Ostrow. AIDS prevention through effective education. Dædalus, 118(3):229–254, Summer 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025256.

Anonymous:1989:BMc

[2089] Anonymous. Back matter. Dædalus, 118(3):??, Summer 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025257.

Anonymous:1989:FMc

[2090] Anonymous. Front matter. Dædalus, 118(3):??, Summer 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025244.

G:1989:PIA

[2091] S. R. G. Preface to the issue “Another India”. Dædalus, 118(4):v–viii, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025261.

Nandy:1989:PCI

[2092] Ashis Nandy. The political culture of the Indian state. Dædalus, 118(4):1–26, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025262.

Kumar:1989:PPI

[2093] Ravinder Kumar. The past and the present: an Indian dialogue. Dædalus, 118(4):27–49, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025263.

Kothari:1989:IET

[2094] Rajni Kothari. The Indian enterprise today. Dædalus, 118(4):50–67, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025264.

Laxman:1989:FCF

[2095] R. K. Laxman. Freedom to cartoon, freedom to speak. Dædalus, 118(4):68–91, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025265.

Correa:1989:PPS

[2096] Charles Correa. The public, the private, and the sacred. Dædalus, 118(4):92–113, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025266.

Madan:1989:RI

[2097] T. N. Madan. Religion in India. Dædalus, 118(4):114–146, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025267.

Vaidyanathan:1989:AII

[2098] T. G. Vaidyanathan. Authority and identity in India. Dædalus, 118(4):147– 169, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20025268.

Patel:1989:CIP

[2099] Gieve Patel. Contemporary Indian painting. Dædalus, 118(4):170–205, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025269.

Desai:1989:IFT

[2100] Anita Desai. Indian fiction today. Dædalus, 118(4):206–231, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025270.

Graubard:1989:IRK

[2101] Stephen R. Graubard and R. K. Narayan. An interview with R. K. Narayan. Dædalus, 118(4):232–237, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025271.

Ramanujan:1989:TT

[2102] A. K. Ramanujan. Telling tales. Dædalus, 118(4):238–261, Fall 1989. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20025272.

Das:1989:VC

[2103] Veena Das. Voices of children. Dædalus, 118(4):262–294, Fall 1989. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20025273.

Ramanujan:1989:SMI

[2104] A. K. Ramanujan, Vinay Dharwadker, Nissim Ezekiel, Jibanananda Das, Clinton B. Seely, Nabaneeta Dev Sen, Sitanshu Yashashchandra, Saleem Peeradina, Jayant Parekh, Rasik Shah, Gulam Mohammed Sheikh, Raghuvir Sahay, Sarveshwar Dayal Saxena, Kedarnath Singh, M. Gopalakrishna Adiga, Chandrashekhara Kambara, G. Shankara Kurup, Indira Sant, Arun Kolatkar, Bhanuji Rao, Jayanta Mahapatra, Amrita Pritam, Subramania Bharati, Aziz Qaisi, Baidar Bakht, and Leslie Lavigne. Sixteen modern Indian poems. Dædalus, 118(4):295–329, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025274.

Karnad:1989:TI

[2105] Girish Karnad. Theatre in India. Dædalus, 118(4):330–352, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025275.

Bhargava:1989:IIS

[2106] Pushpa M. Bhargava and Chandana Chakrabarti. Of India, Indians, and science. Dædalus, 118(4):353–368, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025276.

Sen:1989:IDL

[2107] Amartya Sen. Indian development: Lessons and non-lessons. Dædalus, 118(4):369–392, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025277.

Anonymous:1989:BMd

[2108] Anonymous. Back matter. Dædalus, 118(4):??, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025278.

Anonymous:1989:FMd

[2109] Anonymous. Front matter. Dædalus, 118(4):??, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025260.

Anonymous:1989:VI

[2110] Anonymous. Volume information. Dædalus, 118(4):??, Fall 1989. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025259.

G:1990:PIE

[2111] S. R. G. Preface to the issue “Eastern Europe … Central Europe … Europe”. Dædalus, 119(1):i–ix, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025281.

Ash:1990:M

[2112] Timothy Garton Ash. Mitteleuropa? Dædalus, 119(1):1–21, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025282.

Judt:1990:RCE

[2113] Tony Judt. The rediscovery of Central Europe. Dædalus, 119(1):23–54, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025283.

Schopflin:1990:PTE

[2114] George Schöpflin. The political traditions of Eastern Europe. Dædalus, 119(1):55–90, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025284.

Geremek:1990:BHD

[2115] Bronislaw Geremek. Between hope and despair. Dædalus, 119(1):91–109, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025285.

Skvorecky:1990:BS

[2116] Josef Skvorecký. Bohemia of the soul. Dædalus, 119(1):111–139, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025286.

Banac:1990:PCN

[2117] Ivo Banac. Political change and national diversity. Dædalus, 119(1):141– 159, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20025287.

Biro:1990:LAB

[2118] Yvette Biró. Landscape after battle: Films from “The Other Europe”. Dædalus, 119(1):161–182, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025288.

Hankiss:1990:SP

[2119] Elemér Hankiss. In search of a paradigm. Dædalus, 119(1):183–214, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025289.

Kovacs:1990:REC

[2120] János Mátyás Kovács. Reform economics: The classification gap. Dædalus, 119(1):215–248, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025290.

Rupnik:1990:CEM

[2121] Jacques Rupnik. Central Europe or Mitteleuropa? Dædalus, 119 (1):249–278, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025291.

Gellner:1990:EFE

[2122] Ernest Gellner. Ethnicity and faith in Eastern Europe. Dædalus, 119 (1):279–294, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025292.

Z:1990:SM

[2123] Z. To the Stalin mausoleum. Dædalus, 119(1):295–344, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025293.

Anonymous:1990:BMa

[2124] Anonymous. Back matter. Dædalus, 119(1):??, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025294.

Anonymous:1990:FMa

[2125] Anonymous. Front matter. Dædalus, 119(1):??, Winter 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025280.

G:1990:PIL

[2126] S. R. G. Preface to the issue “Literacy in America”. Dædalus, 119(2):v–viii, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025297.

Hawkins:1990:RL

[2127] David Hawkins. The roots of literacy. Dædalus, 119(2):1–14, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025298.

Resnick:1990:HPL

[2128] Daniel P. Resnick. Historical perspectives on literacy and schooling. Dædalus, 119(2):15–32, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025299.

Damon:1990:RLG

[2129] William Damon. Reconciling the literacies of generations. Dædalus, 119 (2):33–53, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025300.

Botstein:1990:DLI

[2130] Leon Botstein. Damaged literacy: Illiteracies and American democracy. Dædalus, 119(2):55–84, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025301.

Gardner:1990:DSP

[2131] Howard Gardner. The difficulties of school: Probable causes, possible cures. Dædalus, 119(2):85–113, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025302.

Csikszentmihalyi:1990:LIM

[2132] Mihaly Csikszentmihalyi. Literacy and intrinsic motivation. Dædalus, 119 (2):115–140, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025303.

Ogbu:1990:MSL

[2133] John U. Ogbu. Minority status and literacy in comparative perspective. Dædalus, 119(2):141–168, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025304.

Resnick:1990:LS

[2134] Lauren B. Resnick. Literacy in school and out. Dædalus, 119(2):169–185, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025305.

Anderson:1990:IER

[2135] Richard C. Anderson, Bonnie B. Armbruster, and Mary Roe. Improving the education of reading teachers. Dædalus, 119(2):187–209, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025306.

Steen:1990:N

[2136] Lynn Arthur Steen. Numeracy.

Dædalus, 119(2):211–231, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025307.

Snyder:1990:LNT

[2137] Benson R. Snyder. Literacy and numeracy: Two ways of knowing. Dædalus, 119(2):233–256, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025308.

Graubard:1990:DBF

[2138] Stephen R. Graubard. Doing badly and feeling confused. Dædalus, 119 (2):257–279, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025309.

Anonymous:1990:BMb

[2139] Anonymous. Back matter. Dædalus, 119(2):??, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025310.

Anonymous:1990:FMb

[2140] Anonymous. Front matter. Dædalus, 119(2):??, Spring 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025296.

G:1990:PIS

[2141] S. R. G. Preface to the issue “Showa: The Japan of Hirohito”. Dædalus, 119(3):v–vi, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025313.

Gluck:1990:IS

[2142] Carol Gluck. The idea of Showa. Dædalus, 119(3):1–26, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20025314.

Kosaka:1990:SE

[2143] Masataka Kosaka. The Showa Era (1926–1989). Dædalus, 119(3):27–47, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20025315.

Dower:1990:UW

[2144] John W. Dower. The useful war. Dædalus, 119(3):49–70, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025316.

Johnson:1990:PWI

[2145] Chalmers Johnson. The people who invented the mechanical nightingale. Dædalus, 119(3):71–90, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025317.

Iokibe:1990:JMU

[2146] Makoto Iokibe. Japan meets the United States for the second time. Dædalus, 119(3):91–106, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025318.

Passin:1990:OSR

[2147] Herbert Passin. The occupation:

Some reflections. Dædalus, 119(3): 107–129, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025319.

Moriguchi:1990:RMJ

[2148] Chikashi Moriguchi. Rice and melons: Japanese agriculture in the Showa Era. Dædalus, 119(3):131–140, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025320.

Muramatsu:1990:BPB

[2149] Michio Muramatsu. Bringing politics back into Japan. Dædalus, 119 (3):141–154, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025321.

Kitaoka:1990:DMS

[2150] Shinichi Kitaoka. Diplomacy and the military in Showa Japan. Dædalus, 119 (3):155–176, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025322.

Royama:1990:MJ

[2151] Shoichi Royama. Money and the Japanese. Dædalus, 119(3):177–190, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20025323.

Lincoln:1990:SEE

[2152] Edward J. Lincoln. The Showa economic experience. Dædalus, 119(3): 191–208, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025324.

Kelly:1990:RJP

[2153] William W. Kelly. Regional Japan: The price of prosperity and the benefits of dependency. Dædalus, 119 (3):209–227, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025325.

Plath:1990:MCI

[2154] David W. Plath. My-car-isma: Motorizing the Showa self. Dædalus, 119 (3):229–244, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025326.

Yamazaki:1990:ICS

[2155] Masakazu Yamazaki. The intellectual community of the Showa Era. Dædalus, 119(3):245–264, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025327.

Rimer:1990:HCS

[2156] J. Thomas Rimer. High culture in the Showa period. Dædalus, 119 (3):265–278, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025328.

Seidensticker:1990:HTL

[2157] Edward Seidensticker. How they have looked to us. Dædalus, 119 (3):279–298, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025329.

Anonymous:1990:BMc

[2158] Anonymous. Back matter. Dædalus, 119(3):??, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025330.

Anonymous:1990:FMc

[2159] Anonymous. Front matter. Dædalus, 119(3):??, Summer 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025312.

G:1990:PIR

[2160] S. R. G. Preface to the issue “Risk”. Dædalus, 119(4):v–vi, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025334.

Douglas:1990:RFR

[2161] Mary Douglas. Risk as a forensic resource. Dædalus, 119(4):1–16, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025335.

Lowi:1990:RRH

[2162] Theodore J. Lowi. Risks and rights in the history of American governments. Dædalus, 119(4):17–40, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025336.

Wildavsky:1990:TRP

[2163] Aaron Wildavsky and Karl Dake. Theories of risk perception: Who fears what and why? Dædalus, 119(4): 41–60, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20025337.

Jasanoff:1990:AEP

[2164] Sheila Jasanoff. American exceptionalism and the political acknowledgment of risk. Dædalus, 119(4):61–81, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025338.

Sapolsky:1990:PR

[2165] Harvey M. Sapolsky. The politics of risk. Dædalus, 119(4):83–96, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025339.

Huber:1990:PSC

[2166] Peter W. Huber. Pathological science in court. Dædalus, 119(4):97–118, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025340.

Klaidman:1990:HWM

[2167] Stephen Klaidman. How well the media report health risk. Dædalus, 119(4):119–132, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025341.

Burger:1990:HSE

[2168] Edward J. Burger, Jr. Health as a surrogate for the environment. Dædalus, 119(4):133–153, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025342.

Brandt:1990:CRA

[2169] Allan M. Brandt. The cigarette, risk, and American culture. Dædalus, 119(4):155–176, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025343.

Widdus:1990:MRS

[2170] Roy Widdus, Andre Meheus, and Roger Short. The management of risk in sexually transmitted diseases. Dædalus, 119(4):177–191, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025344.

Helzlsouer:1990:RHU

[2171] Kathy J. Helzlsouer and Leon Gordis. Risks to health in the United States. Dædalus, 119(4):193–206, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025345.

Priest:1990:NLS

[2172] George L. Priest. The new legal structure of risk control. Dædalus, 119(4):207–227, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025346.

Hazard:1990:RLS

[2173] Geoffrey C. Hazard, Jr. The role of the legal system in responses to public risk. Dædalus, 119(4):229–234, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025347.

Teuber:1990:JR

[2174] Andreas Teuber. Justifying risk. Dædalus, 119(4):235–254, Fall 1990. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20025348.

Anonymous:1990:BMd

[2175] Anonymous. Back matter. Dædalus, 119(4):??, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025349.

Anonymous:1990:FMd

[2176] Anonymous. Front matter. Dædalus, 119(4):??, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025333.

Anonymous:1990:VI

[2177] Anonymous. Volume information. Dædalus, 119(4):??, Fall 1990. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025332.

G:1991:PIA

[2178] S. R. G. Preface to the issue “Arms Control: Thirty Years on”. Dædalus, 120(1):v–viii, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025352.

Adler:1991:ACD

[2179] Emanuel Adler. Arms control, disarmament, and national security: A thirty year retrospective and a new set of anticipations. Dædalus, 120(1):1–20, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025353.

Schelling:1991:TY

[2180] Thomas C. Schelling. The thirtieth year. Dædalus, 120(1):21–31, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025354.

Doty:1991:AC

[2181] Paul Doty. Arms control: 1960, 1990, 2020. Dædalus, 120(1):33–52, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025355.

Bowie:1991:AC

[2182] Robert R. Bowie. Arms control in the 1990s. Dædalus, 120(1):53–68, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025356.

Freedman:1991:ACT

[2183] Lawrence Freedman. Arms control: Thirty years on. Dædalus, 120(1):69– 82, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20025357.

Holst:1991:ACN

[2184] Johan Jørgen Holst. Arms control in the Nineties: a European perspective. Dædalus, 120(1):83–110, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025358.

Kelleher:1991:ACR

[2185] Catherine M. Kelleher. Arms control in a revolutionary future: Europe. Dædalus, 120(1):111–131, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025359.

Kokoshin:1991:ACV

[2186] A. A. Kokoshin. Arms control: a view from Moscow. Dædalus, 120 (1):133–143, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025360.

Nye:1991:ACI

[2187] Joseph S. Nye, Jr. Arms control and international politics. Dædalus, 120 (1):145–165, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025361.

Jervis:1991:ACS

[2188] Robert Jervis. Arms control, stability, and causes of war. Dædalus, 120 (1):167–181, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025362.

Weber:1991:CI

[2189] Steve Weber. Cooperation and interdependence. Dædalus, 120(1):183–201, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025363.

Trachtenberg:1991:PFA

[2190] Marc Trachtenberg. The past and future of arms control. Dædalus, 120 (1):203–216, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025364.

Posen:1991:CSC

[2191] Barry R. Posen. Crisis stability and conventional arms control. Dædalus, 120(1):217–232, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025365.

Carter:1991:ETN

[2192] Ashton B. Carter. Emerging themes in nuclear arms control. Dædalus, 120 (1):233–249, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025366.

Sims:1991:AAN

[2193] Jennifer E. Sims. The American approach to nuclear arms control: A retrospective. Dædalus, 120(1):251–272, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025367.

Anonymous:1991:BMa

[2194] Anonymous. Back matter. Dædalus, 120(1):??, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025368.

Anonymous:1991:FMa

[2195] Anonymous. Front matter. Dædalus, 120(1):??, Winter 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025351.

G:1991:PIL

[2196] S. R. G. Preface to the issue “The Living Tree: The Changing Meaning of Being Chinese Today”. Dædalus, 120(2):v–viii, Spring 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025371.

Wei-ming:1991:CCP

[2197] Tu Wei-ming. Cultural China: The periphery as the center. Dædalus, 120(2):1–32, Spring 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025372.

Elvin:1991:IW

[2198] Mark Elvin. The inner world of 1830. Dædalus, 120(2):33–61, Spring 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025373.

King:1991:KHN

[2199] Ambrose Yeo chi King. Kuan-hsi and network building: a sociological interpretation. Dædalus, 120(2):63–84, Spring 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025374.

Schwarcz:1991:NSL

[2200] Vera Schwarcz. No solace from Lethe:

History, memory, and cultural identity in Twentieth-Century China. Dædalus, 120(2):85–112, Spring 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025375.

Cohen:1991:BCP

[2201] Myron L. Cohen. Being Chinese: The peripheralization of traditional identity. Dædalus, 120(2):113–134, Spring 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025376.

Gungwu:1991:ANC

[2202] Wang Gungwu. Among non-Chinese. Dædalus, 120(2):135–157, Spring 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025377.

Wu:1991:CCN

[2203] David Yen ho Wu. The construction of Chinese and non-Chinese identities. Dædalus, 120(2):159–179, Spring 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025378.

Wang:1991:RCI

[2204] L. Ling chi Wang. Roots and changing identity of the Chinese in the United States. Dædalus, 120(2):181– 206, Spring 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20025379.

Lee:1991:MCD

[2205] Leo Ou fan Lee. On the margins of the Chinese discourse: Some personal thoughts on the cultural meaning of the periphery. Dædalus, 120 (2):207–226, Spring 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025380.

Anonymous:1991:BMb

[2206] Anonymous. Back matter. Dædalus, 120(2):??, Spring 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025381.

Anonymous:1991:FMb

[2207] Anonymous. Front matter. Dædalus, 120(2):??, Spring 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025370.

G:1991:PIR

[2208] S. R. G. Preface to the issue “Religion and Politics”. Dædalus, 120 (3):v–viii, Summer 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025384.

Wuthnow:1991:URP

[2209] Robert Wuthnow. Understanding religion and politics. Dædalus, 120 (3):1–20, Summer 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025385.

Demerath:1991:RCC

[2210] N. J. Demerath III. Religious capital and capital religions: Cross-cultural and non-legal factors in the separation of Church and state. Dædalus, 120(3):21–40, Summer 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025386.

Baker:1991:AIE

[2211] Raymond William Baker. Afraid for Islam: Egypt’s Muslim centrists between pharaohs and fundamentalists. Dædalus, 120(3):41–68, Summer 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025387.

Chehabi:1991:RPI

[2212] H. E. Chehabi. Religion and politics in Iran: How theocratic is the Islamic Republic? Dædalus, 120 (3):69–91, Summer 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025388.

Anderson:1991:OAI

[2213] Lisa Anderson. Obligation and accountability: Islamic politics in North Africa. Dædalus, 120(3):93–112, Summer 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025389.

Mardin:1991:JU

[2214] Serif Mardin. The just and the unjust. Dædalus, 120(3):113–129, Summer 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025390.

Crahan:1991:CSL

[2215] Margaret E. Crahan. Church and state in Latin America: Assassinating some old and new stereotypes. Dædalus, 120 (3):131–158, Summer 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025391.

Linz:1991:CSS

[2216] Juan J. Linz. Church and state in Spain from the Civil War to the return of democracy. Dædalus, 120 (3):159–178, Summer 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025392.

Beckford:1991:PRE

[2217] James A. Beckford. Politics and religion in England and Wales. Dædalus, 120(3):179–201, Summer 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025393.

Sanneh:1991:RPT

[2218] Lamin Sanneh. Religion and politics:

Third world perspectives on a comparative religious theme. Dædalus, 120 (3):203–218, Summer 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025394.

Obeyesekere:1991:BCE

[2219] Gananath Obeyesekere. Buddhism and conscience: an exploratory essay. Dædalus, 120(3):219–239, Summer 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025395.

Anonymous:1991:BMc

[2220] Anonymous. Back matter. Dædalus, 120(3):??, Summer 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025396.

Anonymous:1991:FMc

[2221] Anonymous. Front matter. Dædalus, 120(3):??, Summer 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025383.

G:1991:PIS

[2222] S. R. G. Preface to the issue “Searching for Security in a Global Economy”. Dædalus, 120(4):v–x, Fall 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025400.

Vernon:1991:NNG

[2223] Raymond Vernon and Ethan B. Kapstein. National needs, global resources. Dædalus, 120(4):1–22, Fall 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20025401.

Moravcsik:1991:AAM

[2224] Andrew Moravcsik. Arms and autarky in modern European history. Dædalus, 120(4):23–45, Fall 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025402.

Samuels:1991:RSJ

[2225] Richard J. Samuels. Reinventing security: Japan since Meiji. Dædalus, 120(4):47–68, Fall 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025403.

Friedberg:1991:EAU

[2226] Aaron L. Friedberg. The end of autonomy: The United States after five decades. Dædalus, 120(4):69–90, Fall 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025404.

Mastanduno:1991:USD

[2227] Michael Mastanduno. The United States defiant: Export controls in the postwar era. Dædalus, 120(4):91–112, Fall 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025405.

Davis:1991:ESC

[2228] Christopher Mark Davis. The exceptional Soviet case: Defense in an autarkic system. Dædalus, 120(4):113–134, Fall 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025406.

Moran:1991:A

[2229] Theodore H. Moran and David C.

Mowery. Aerospace. Dædalus, 120 (4):135–154, Fall 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025407.

Ziegler:1991:S

[2230] J. Nicholas Ziegler. Semiconductors. Dædalus, 120(4):155–182, Fall 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025408.

Painter:1991:ION

[2231] David S. Painter. International oil and national security. Dædalus, 120 (4):183–206, Fall 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025409.

Kurth:1991:CDW

[2232] James Kurth. The common defense and the world market. Dædalus, 120(4):207–228, Fall 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025410.

Anonymous:1991:BMd

[2233] Anonymous. Back matter. Dædalus, 120(4):??, Fall 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025411.

Anonymous:1991:FMd

[2234] Anonymous. Front matter. Dædalus, 120(4):??, Fall 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025399.

Anonymous:1991:VI

[2235] Anonymous. Volume information. Dædalus, 120(4):??, Fall 1991. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025398.

G:1992:PIN

[2236] S. R. G. Preface to the issue “A New Era in Computation”. Dædalus, 121(1):v–xvi, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025414.

Hillis:1992:WMP

[2237] W. Daniel Hillis. What is massively parallel computing, and why is it important? Dædalus, 121(1):1–15, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025415.

Holland:1992:CAS

[2238] John H. Holland. Complex adaptive systems. Dædalus, 121(1):17–30, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025416.

Deng:1992:PPC

[2239] Yuefan Deng, James Glimm, and David H. Sharp. Perspectives on parallel computing. Dædalus, 121(1):31–52, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025417.

Hasslacher:1992:PBM

[2240] Brosl Hasslacher. Parallel billiards and monster systems. Dædalus, 121 (1):53–65, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025418.

Bailey:1992:FWR

[2241] James Bailey. First we reshape our computers, Then our computers reshape us: The broader intellectual impact of parallelism. Dædalus, 121 (1):67–86, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025419.

Sokolowski:1992:PCE

[2242] Robert Sokolowski. Parallelism in conscious experience. Dædalus, 121 (1):87–103, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025420.

Browder:1992:TII

[2243] Felix E. Browder. Of time, intelligence, and institutions. Dædalus, 121 (1):105–110, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025421.

Fox:1992:PCE

[2244] Geoffrey C. Fox. Parallel computing and education. Dædalus, 121 (1):111–118, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025422.

Metropolis:1992:ACP

[2245] N. (Nicholas) Metropolis. The age of computing: a personal memoir. Dædalus, 121(1):119–130, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.

                                        bio.net/hypermail/bioforum/1992-
                                        September/000238.html;   http://
                                        www.jstor.org/stable/20025423.

Davis:1992:WSP

[2246] Philip J. Davis. What should the public know about mathematics? Dædalus, 121(1):131–138, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025424.

Schwartz:1992:AET

[2247] Jacob T. Schwartz. America’s economic-technological agenda for the 1990s. Dædalus, 121(1):139–165, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025425.

Anonymous:1992:BMa

[2248] Anonymous. Back matter. Dædalus, 121(1):??, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025426.

Anonymous:1992:FMa

[2249] Anonymous. Front matter. Dædalus, 121(1):??, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025413.

G:1992:PIE

[2250] S. R. G. Preface to the issue “The Exit from Communism”. Dædalus, 121(2):v–x, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025429.

Manuel:1992:RKM

[2251] Frank E. Manuel. A requiem for Karl Marx. Dædalus, 121(2):1–19, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025430.

Eisenstadt:1992:BCR

[2252] S. N. Eisenstadt. The breakdown of communist regimes and the vicissitudes of modernity. Dædalus, 121(2):21–41, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025431.

Kolakowski:1992:AMR

[2253] Leszek Kolakowski. Amidst moving ruins. Dædalus, 121(2):43–56, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025432.

Malia:1992:LE

[2254] Martin Malia. Leninist endgame. Dædalus, 121(2):57–75, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20025433.

Aaslund:1992:RRC

[2255] Anders Åslund. Russia’s road from Communism. Dædalus, 121(2):77–95, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025434.

Zaslavsky:1992:NDT

[2256] Victor Zaslavsky. Nationalism and democratic transition in postcommunist societies. Dædalus, 121(2):97–121, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025435.

Linz:1992:PIE

[2257] Juan J. Linz and Alfred Stepan. Political identities and electoral sequences: Spain, the Soviet union, and Yugoslavia. Dædalus, 121(2):123–139, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025436.

Banac:1992:FAW

[2258] Ivo Banac. The fearful asymmetry of war: The causes and consequences of Yugoslavia’s demise. Dædalus, 121 (2):141–174, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025437.

Musil:1992:CMT

[2259] Jirí Musil. Czechoslovakia in the middle of transition. Dædalus, 121 (2):175–195, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025438.

Kohak:1992:AAL

[2260] Erazim Kohák. Ashes, ashes … Central Europe after forty years. Dædalus, 121(2):197–215, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20025439.

Cavazza:1992:IPE

[2261] Fabio Luca Cavazza and Elaine Maclachlan. The Italian paradox: an exit from Communism. Dædalus, 121 (2):217–249, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

/www.jstor.org/stable/20025440.

Wei-ming:1992:IEC

[2262] Tu Wei-ming. Intellectual effervescence in China. Dædalus, 121(2):251–292, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20025441.

Anonymous:1992:BMb

[2263] Anonymous. Back matter. Dædalus, 121(2):??, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025442.

Anonymous:1992:FMb

[2264] Anonymous. Front matter. Dædalus, 121(2):??, Spring 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20025428.

G:1992:PIP

[2265] S. R. G. Preface to the issue “Political Pharmacology: Thinking about Drugs”. Dædalus, 121(3):v–x, Summer 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027118.

Falco:1992:FDF

[2266] Mathea Falco. Foreign drugs, foreign wars. Dædalus, 121(3):1–14, Summer 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027119.

Reuter:1992:HAP

[2267] Peter Reuter. Hawks ascendant: The punitive trend of American drug policy. Dædalus, 121(3):15–52, Summer 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027120.

Kleiman:1992:NPN

[2268] Mark A. R. Kleiman. Neither prohibition nor legalization: Grudging toleration in drug control policy. Dædalus, 121(3):53–83, Summer 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027121.

Nadelmann:1992:TSA

[2269] Ethan A. Nadelmann. Thinking seriously about alternatives to drug prohibition. Dædalus, 121(3):85–132, Summer 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027122.

Skolnick:1992:RDP

[2270] Jerome H. Skolnick. Rethinking the drug problem. Dædalus, 121(3): 133–159, Summer 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027123.

Szasz:1992:FTD

[2271] Thomas Szasz. The fatal temptation:

Drug prohibition and the fear of autonomy. Dædalus, 121(3):161–164, Summer 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027124.

Lewis:1992:MHP

[2272] David C. Lewis. Medical and health perspectives on a failing US drug policy. Dædalus, 121(3):165–194, Summer 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027125.

Schlesinger:1992:FBC

[2273] Mark Schlesinger and Robert A. Dorwart. Falling between the cracks: Failing national strategies for the treatment of substance abuse. Dædalus, 121 (3):195–237, Summer 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027126.

Erickson:1992:RTC

[2274] Patricia G. Erickson. Recent trends in Canadian drug policy: The decline and resurgence of prohibitionism. Dædalus, 121(3):239–267, Summer 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027127.

Heath:1992:UDC

[2275] Dwight B. Heath. US drug control policy: a cultural perspective. Dædalus, 121(3):269–291, Summer 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027128.

Husch:1992:CUD

[2276] Jerri A. Husch. Culture and US drug policy: Toward a new conceptual framework. Dædalus, 121(3): 293–304, Summer 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027129.

Anonymous:1992:BMc

[2277] Anonymous. Back matter. Dædalus, 121(3):??, Summer 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027130.

Anonymous:1992:FMc

[2278] Anonymous. Front matter. Dædalus, 121(3):??, Summer 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027117.

G:1992:PII

[2279] S. R. G. Preface to the issue “Immobile Democracy?”. Dædalus, 121(4):v–xiv, Fall 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027134.

Johnson:1992:JSN

[2280] Chalmers Johnson. Japan in search of a “Normal” role. Dædalus, 121(4):1–33, Fall 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027135.

Vogel:1992:JAR

[2281] Ezra F. Vogel. Japanese-American relations after the cold war. Dædalus, 121(4):35–60, Fall 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027136.

White:1992:HTW

[2282] Merry White. Home truths: Women and social change in Japan. Dædalus, 121(4):61–82, Fall 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027137.

Judt:1992:PAC

[2283] Tony Judt. The past is another country: Myth and memory in postwar Europe. Dædalus, 121(4):83–118, Fall 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027138.

Bertram:1992:UGR

[2284] Christoph Bertram. US-German relations in a world at sea. Dædalus, 121(4):119–128, Fall 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027139.

Pinto:1992:GES

[2285] Diana Pinto. The great European sea change. Dædalus, 121(4):129–150, Fall 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027140.

Peterson:1992:IP

[2286] Paul E. Peterson. An immodest proposal. Dædalus, 121(4):151–174, Fall 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027141.

Marmor:1992:AMC

[2287] Theodore R. Marmor and David Boyum. American medical care reform: Are we doomed to fail? Dædalus, 121(4):175–194, Fall 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

/www.jstor.org/stable/20027142.

Kevles:1992:FUT

[2288] Daniel J. Kevles. Foundations, universities, and trends in support for the physical and biological sciences, 1900–1992. Dædalus, 121(4):195–235, Fall 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027143.

Anonymous:1992:BMd

[2289] Anonymous. Back matter. Dædalus, 121(4):??, Fall 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027144.

Anonymous:1992:FMd

[2290] Anonymous. Front matter. Dædalus, 121(4):??, Fall 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027133.

Anonymous:1992:VI

[2291] Anonymous. Volume information. Dædalus, 121(4):??, Fall 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027132.

G:1993:PIAa

[2292] S. R. G. Preface to the issue “America’s Childhood”. Dædalus, 122(1):v–xii, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027147.

Minow:1993:SMC

[2293] Martha Minow and Richard Weissbourd. Social movements for children. Dædalus, 122(1):1–29, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027148.

Csikszentmihalyi:1993:COG

[2294] Mihaly Csikszentmihalyi. Contexts of optimal growth in childhood. Dædalus, 122(1):31–56, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027149.

Danziger:1993:CPP

[2295] Sandra K. Danziger and Sheldon Danziger. Child poverty and public policy: Toward a comprehensive antipoverty agenda. Dædalus, 122 (1):57–84, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027150.

Finn:1993:TI

[2296] David Finn. Thoughts and images. Dædalus, 122(1):85–92, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027151.

Earls:1993:TSD

[2297] Felton Earls and Mary Carlson. Towards sustainable development for American families. Dædalus, 122 (1):93–121, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027152.

Cottle:1993:WJ

[2298] Thomas J. Cottle. Witness of joy. Dædalus, 122(1):123–150, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027153.

Vinovskis:1993:ECE

[2299] Maris A. Vinovskis. Early childhood education: Then and now. Dædalus, 122(1):151–176, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027154.

Cohen:1993:ACT

[2300] David K. Cohen and S. G. Grant.

America’s children and their elementary schools. Dædalus, 122(1):177–207, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027155.

Marshall:1993:TSW

[2301] Kim Marshall. Teachers and schools:

What makes a difference: A Principal’s perspective. Dædalus, 122 (1):209–242, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027156.

Ziegler:1993:ECV

[2302] Jerome M. Ziegler. Education: The crucial variable for children? Dædalus, 122(1):243–258, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027157.

Condry:1993:TTU

[2303] John Condry. Thief of time, unfaithful servant: Television and the American child. Dædalus, 122(1):259–278, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027158.

Shweder:1993:WDM

[2304] Richard A. Shweder. “Why Do Men Barbecue?” and other postmodern ironies of growing up in the decade of ethnicity. Dædalus, 122(1):279–308, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027159.

Anonymous:1993:BMa

[2305] Anonymous. Back matter. Dædalus, 122(1):??, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027160.

Anonymous:1993:FMa

[2306] Anonymous. Front matter. Dædalus, 122(1):??, Winter 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027146.

G:1993:PIC

[2307] S. R. G. Preface to the issue “China in Transformation”. Dædalus, 122(2):v–vi, Spring 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027163.

Wei-ming:1993:ICP

[2308] Tu Wei-ming. Introduction: Cultural perspectives. Dædalus, 122(2):vii–xxiv, Spring 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027164.

Friedman:1993:FCM

[2309] Edward Friedman. A failed Chinese modernity. Dædalus, 122(2):1–17, Spring 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027165.

Siu:1993:CIP

[2310] Helen F. Siu. Cultural identity and the politics of difference in south China. Dædalus, 122(2):19–43, Spring 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027166.

Alford:1993:DES

[2311] William P. Alford. Double-edged swords cut both ways: Law and legitimacy in the People’s Republic of China. Dædalus, 122(2):45–69, Spring 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027167.

Gungwu:1993:RRU

[2312] Wang Gungwu. To reform a revolution: Under the righteous mandate. Dædalus, 122(2):71–94, Spring 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027168.

Nathan:1993:CRD

[2313] Andrew J. Nathan and Tianjian Shi. Cultural requisites for democracy in China: Findings from a survey. Dædalus, 122(2):95–123, Spring 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027169.

Yu:1993:RCT

[2314] Ying shih Yü. The radicalization of China in the Twentieth Century. Dædalus, 122(2):125–150, Spring 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027170.

Cohen:1993:CPI

[2315] Myron L. Cohen. Cultural and political inventions in modern China: The case of the Chinese “Peasant”. Dædalus, 122(2):151–170, Spring 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027171.

Lin:1993:SCS

[2316] Tongqi Lin. A search for China’s soul. Dædalus, 122(2):171–188, Spring 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027172.

Link:1993:CCP

[2317] Perry Link. China’s “Core” problem. Dædalus, 122(2):189–205, Spring 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027173.

Apter:1993:YNR

[2318] David E. Apter. Yan’an and the narrative reconstruction of reality. Dædalus, 122(2):207–232, Spring 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027174.

Anonymous:1993:BMb

[2319] Anonymous. Back matter. Dædalus, 122(2):??, Spring 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027175.

Anonymous:1993:FMb

[2320] Anonymous. Front matter. Dædalus, 122(2):??, Spring 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027162.

G:1993:PIR

[2321] S. R. G. Preface to the issue “Reconstructing Nations and States”. Dædalus, 122(3):v–viii, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027178.

Hall:1993:NCE

[2322] John A. Hall. Nationalisms: Classified and explained. Dædalus, 122 (3):1–28, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027179.

Tilly:1993:NSD

[2323] Charles Tilly. National self-determination as a problem for all of us. Dædalus, 122(3):29–36, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

/www.jstor.org/stable/20027180.

Verdery:1993:WNN

[2324] Katherine Verdery. Whither “Nation”

and “Nationalism”? Dædalus, 122 (3):37–46, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027181.

Greenfeld:1993:TNW

[2325] Liah Greenfeld. Transcending the Nation’s worth. Dædalus, 122(3):47–62, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027182.

Hoffmann:1993:TFN

[2326] Stanley Hoffmann. Thoughts on the French nation today. Dædalus, 122 (3):63–79, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027183.

Klusmeyer:1993:AIC

[2327] Douglas B. Klusmeyer. Aliens, immigrants, and citizens: The politics of inclusion in the Federal Republic of Germany. Dædalus, 122 (3):81–114, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027184.

Mann:1993:NSE

[2328] Michael Mann. Nation-states in Europe and other continents: Diversifying, developing, not dying. Dædalus, 122 (3):115–140, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027185.

Gellner:1993:HU

[2329] Ernest Gellner. Homeland of the unrevolution. Dædalus, 122(3):141–153, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027186.

Nairn:1993:ISC

[2330] Tom Nairn. Internationalism and the second coming. Dædalus, 122 (3):155–170, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027187.

Kramer:1993:ANM

[2331] Martin Kramer. Arab nationalism:

Mistaken identity. Dædalus, 122(3): 171–206, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027188.

Schwartz:1993:CMN

[2332] Benjamin I. Schwartz. Culture, modernity, and nationalism: Further reflections. Dædalus, 122(3):207–226, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027189.

Varshney:1993:CMI

[2333] Ashutosh Varshney. Contested meanings: India’s national identity, Hindu nationalism, and the politics of anxiety. Dædalus, 122(3):227–261, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027190.

Annan:1993:BRB

[2334] Noel Annan. Book review: Churchill, The End of Glory: A Political Biography. Dædalus, 122(3):263–272, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027191.

Anonymous:1993:BMc

[2335] Anonymous. Back matter. Dædalus, 122(3):??, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027192.

Anonymous:1993:FMc

[2336] Anonymous. Front matter. Dædalus, 122(3):??, Summer 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027177.

G:1993:PIAb

[2337] S. R. G. Preface to the issue “The American Research University”. Dædalus, 122(4):v–viii, Fall 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027196.

Cole:1993:BAD

[2338] Jonathan R. Cole. Balancing acts: Dilemmas of choice facing research universities. Dædalus, 122(4):1–36, Fall 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027197.

Smelser:1993:PAD

[2339] Neil J. Smelser. The politics of ambivalence: Diversity in the research universities. Dædalus, 122(4):37–53, Fall 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027198.

Searle:1993:RRW

[2340] John R. Searle. Rationality and realism, what is at stake? Dædalus, 122(4):55–83, Fall 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027199.

Prewitt:1993:ARU

[2341] Kenneth Prewitt. America’s research universities under public scrutiny. Dædalus, 122(4):85–99, Fall 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027200.

Keohane:1993:MRU

[2342] Nannerl O. Keohane. The mission of the research University. Dædalus, 122(4):101–125, Fall 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027201.

Kennedy:1993:MCR

[2343] Donald Kennedy. Making choices in the research University. Dædalus, 122(4):127–156, Fall 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027202.

Stigler:1993:CRU

[2344] Stephen M. Stigler. Competition and the research universities. Dædalus, 122(4):157–177, Fall 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027203.

Richardson:1993:ASH

[2345] William C. Richardson. The appropriate scale of the health sciences enterprise. Dædalus, 122(4):179–195, Fall 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027204.

Nichols:1993:FSP

[2346] Rodney W. Nichols. Federal science policy and universities: Consequences of success. Dædalus, 122(4):197–224, Fall 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027205.

Skolnikoff:1993:KBI

[2347] Eugene B. Skolnikoff. Knowledge without borders? Internationalization of the research universities. Dædalus, 122(4):225–252, Fall 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027206.

Anonymous:1993:BMd

[2348] Anonymous. Back matter. Dædalus, 122(4):??, Fall 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027207.

Anonymous:1993:FMd

[2349] Anonymous. Front matter. Dædalus, 122(4):??, Fall 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027195.

Anonymous:1993:VI

[2350] Anonymous. Volume information. Dædalus, 122(4):??, Fall 1993. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027194.

G:1994:PIG

[2351] S. R. G. Preface to the issue “Germany in Transition”. Dædalus, 123(1):v–xii, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027210.

Mertes:1994:GSP

[2352] Michael Mertes. Germany’s social and political culture: Change through consensus? Dædalus, 123(1):1–32, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027211.

Muller:1994:DG

[2353] Steven Muller. Democracy in Germany. Dædalus, 123(1):33–56, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027212.

Lauk:1994:GCE

[2354] Kurt J. Lauk. Germany at the crossroads: On the efficiency of the German economy. Dædalus, 123(1):57–83, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027213.

Anonymous:1994:PTSa

[2355] Anonymous. [photograph]: Thomas Struth, Mainzerstrasse, Frankfurt, 1987. Courtesy Marian Goodman Gallery, New York. Dædalus, 123(1): 84, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027214.

Bade:1994:ISP

[2356] Klaus J. Bade and Lieselotte Anderson. Immigration and social peace in united Germany. Dædalus, 123 (1):85–106, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027215.

Winkler:1994:RNG

[2357] Heinrich August Winkler, C. Michelle Murphy, Cornelius Partsch, and Susan List. Rebuilding of a nation: The Germans before and after unification. Dædalus, 123(1):107–127, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027216.

Anonymous:1994:PTSb

[2358] Anonymous. [Photograph]: Thomas Struth, Am Lehrter Bahnhof, Berlin, 1992. Courtesy Marian Goodman Gallery, New York. Dædalus, 123(1): 128, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027217.

LeGloannec:1994:GI

[2359] Anne-Marie Le Gloannec. On German identity. Dædalus, 123(1):129–148, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027218.

Eisel:1994:PUG

[2360] Stephan Eisel and Deborah Lucas Schneider. The politics of a united Germany. Dædalus, 123(1):149–171, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027219.

Anonymous:1994:PTSc

[2361] Anonymous. [photograph]: Thomas Struth, Ludwigstrasse, Munich, 1980. Courtesy Marian Goodman Gallery, New York. Dædalus, 123(1):172, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027220.

Kocka:1994:CUH

[2362] Jürgen Kocka. Crisis of unification:

How Germany changes. Dædalus, 123 (1):173–192, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027221.

Kuhnhardt:1994:MGG

[2363] Ludger Kühnhardt and Lieselotte Anderson. Multi-German Germany. Dædalus, 123(1):193–209, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027222.

Anonymous:1994:PTSd

[2364] Anonymous. [photograph]: Thomas Struth, Georg Von Vollmar Straße, Leverkusen, 1980. Courtesy Marian Goodman Gallery, New York. Dædalus, 123(1):210, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027223.

Fulbrook:1994:ASI

[2365] Mary Fulbrook. Aspects of society and identity in the new Germany. Dædalus, 123(1):211–234, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027224.

Witte:1994:TCT

[2366] Barthold C. Witte. Two catastrophes, two causes, and how the Germans dealt with them. Dædalus, 123 (1):235–249, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027225.

Anonymous:1994:PTSe

[2367] Anonymous. [photograph]: Thomas Struth, Erich Ferl Strasse, Leipzig, 1991. Courtesy Marian Goodman Gallery, New York. Dædalus, 123(1): 250, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027226.

Schroder:1994:RPC

[2368] Richard Schröder and Deborah Lucas Schneider. The role of the Protestant Church in German unification. Dædalus, 123(1):251–261, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027227.

Anonymous:1994:PTSf

[2369] Anonymous. [photograph]: Thomas Struth, Sophiengemeinde, Gr Hamburgerstrasse, Berlin, 1992. Courtesy Marian Goodman Gallery, New York. Dædalus, 123(1):262, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027228.

Thies:1994:OPC

[2370] Jochen Thies and Deborah Lucas Schneider. Observations on the political class in Germany. Dædalus, 123 (1):263–276, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027229.

Gauck:1994:DSP

[2371] Joachim Gauck and Martin Fry. Dealing with a Stasi past. Dædalus, 123 (1):277–284, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027230.

Anonymous:1994:BMa

[2372] Anonymous. Back matter. Dædalus, 123(1):??, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027231.

Anonymous:1994:FMa

[2373] Anonymous. Front matter. Dædalus, 123(1):??, Winter 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027209.

G:1994:PIE

[2374] S. R. G. Preface to the issue “Europe through a Glass Darkly”. Dædalus, 123(2):v–xvi, Spring 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027234.

Hoffmann:1994:EIC

[2375] Stanley Hoffmann. Europe’s identity crisis revisited. Dædalus, 123(2):1–23, Spring 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027235.

Schlesinger:1994:ECC

[2376] Philip R. Schlesinger. Europe’s contradictory communicative space. Dædalus, 123(2):25–52, Spring 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027236.

Cavazza:1994:MBD

[2377] Fabio Luca Cavazza, Carlo Pelanda, Anthony Molho, and Alan Ginet. Maastricht: Before, during, after. Dædalus, 123(2):53–80, Spring 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027237.

Schwarz:1994:GNE

[2378] Hans-Peter Schwarz and Deborah Lucas Schneider. Germany’s national and European interests. Dædalus, 123 (2):81–105, Spring 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027238.

Campiglio:1994:EMR

[2379] Luigi Campiglio. Europe on the mark: Ready to go? Dædalus, 123 (2):107–127, Spring 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027239.

Lilla:1994:OVR

[2380] Mark Lilla. The other velvet revolution: Continental liberalism and its discontents. Dædalus, 123(2):129–157, Spring 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027240.

Bermeo:1994:DE

[2381] Nancy Bermeo. Democracy in Europe. Dædalus, 123(2):159–178, Spring 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027241.

Snyder:1994:RBF

[2382] Jack Snyder. Russian backwardness and the future of Europe. Dædalus, 123(2):179–201, Spring 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027242.

Pavlowitch:1994:WBW

[2383] Stevan K. Pavlowitch. Who is “Balkanizing” whom? The misunderstandings between the debris of Yugoslavia and an unprepared West. Dædalus, 123(2):203–223, Spring 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027243.

Anonymous:1994:BMb

[2384] Anonymous. Back matter. Dædalus, 123(2):??, Spring 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027244.

Anonymous:1994:FMb

[2385] Anonymous. Front matter. Dædalus, 123(2):??, Spring 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027233.

G:1994:PIA

[2386] S. R. G. Preface to the issue “After Communism: What?”. Dædalus, 123(3):v–xiv, Summer 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027247.

Judt:1994:NEN

[2387] Tony Judt. Nineteen eighty-nine: The end of which European era? Dædalus, 123(3):1–19, Summer 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027248.

Szporluk:1994:AEW

[2388] Roman Szporluk. After empire: What? Dædalus, 123(3):21–39, Summer 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027249.

Rose:1994:GGE

[2389] Richard Rose. Getting by without government: Everyday life in Russia. Dædalus, 123(3):41–62, Summer 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027250.

Medish:1994:RLF

[2390] Mark Medish. Russia: Lost and found. Dædalus, 123(3):63–89, Summer 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027251.

Rupnik:1994:ENF

[2391] Jacques Rupnik. Europe’s new frontiers: Remapping Europe. Dædalus, 123(3):91–114, Summer 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027252.

Hankiss:1994:EPE

[2392] Elemér Hankiss. European paradigms:

East and West, 1945–1994. Dædalus, 123(3):115–126, Summer 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20027253.

Schopflin:1994:PPD

[2393] George Schöpflin. Postcommunism:

The problems of democratic construction. Dædalus, 123(3):127–141, Summer 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027254.

Kolankiewicz:1994:ESP

[2394] George Kolankiewicz. Elites in search of a political formula. Dædalus, 123 (3):143–157, Summer 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027255.

Rev:1994:PVC

[2395] István Rév. The postmortem victory of Communism. Dædalus, 123 (3):159–170, Summer 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027256.

Marga:1994:CSE

[2396] Andrei Marga. The culture of scholarship in Europe today. Dædalus, 123 (3):171–184, Summer 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027257.

Weil:1994:PPT

[2397] Eric Weil. Philosophical and political thought in Europe today. Dædalus, 123 (3):185–194, Summer 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027258.

Anonymous:1994:BMc

[2398] Anonymous. Back matter. Dædalus, 123(3):??, Summer 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027259.

Anonymous:1994:FMc

[2399] Anonymous. Front matter. Dædalus, 123(3):??, Summer 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027246.

G:1994:PIH

[2400] S. R. G. Preface to the issue “Health and Wealth”. Dædalus, 123(4):v–xiv, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027263.

Frank:1994:DHH

[2401] John W. Frank and J. Fraser Mustard. The determinants of health from a historical perspective. Dædalus, 123(4):1–19, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027264.

Evans:1994:HCT

[2402] Robert G. Evans. Health care as a threat to health: Defense, opulence, and the social environment. Dædalus, 123(4):21–42, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027265.

Lavis:1994:CWT

[2403] John N. Lavis and Gregory L. Stoddart. Can we have too much health care? Dædalus, 123(4):43–60, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027266.

Wilkinson:1994:ETM

[2404] Richard G. Wilkinson. The epidemiological transition: From material scarcity to social disadvantage? Dædalus, 123(4):61–77, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027267.

Syme:1994:SEH

[2405] S. Leonard Syme. The social environment and health. Dædalus, 123(4): 79–86, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20027268.

Miyaji:1994:MMS

[2406] Naoko T. Miyaji and Margaret Lock. Monitoring motherhood: Sociocultural and historical aspects of maternal and child health in Japan. Dædalus, 123(4):87–112, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027269.

Ikegami:1994:EEH

[2407] Naoki Ikegami. Efficiency and effectiveness in health care. Dædalus, 123(4):113–125, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027270.

Rosenberg:1994:HDC

[2408] Nathan Rosenberg. How the developed countries became Rich. Dædalus, 123(4):127–140, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027271.

Romer:1994:EGI

[2409] Paul M. Romer. Economic growth and investment in children. Dædalus, 123(4):141–154, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027272.

Cynader:1994:MBD

[2410] Max S. Cynader. Mechanisms of brain development and their role in health and well-being. Dædalus, 123(4):155– 165, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027273.

Hertzman:1994:LIC

[2411] Clyde Hertzman. The lifelong impact of childhood experiences: A population health perspective. Dædalus, 123(4):167–180, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027274.

Wolfson:1994:TSH

[2412] Michael C. Wolfson. Toward a system of health statistics. Dædalus, 123(4):181–195, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027275.

Marmot:1994:SDH

[2413] Michael G. Marmot. Social differentials in health within and between populations. Dædalus, 123(4):197–216, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027276.

Anonymous:1994:BMd

[2414] Anonymous. Back matter. Dædalus, 123(4):??, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027277.

Anonymous:1994:FMd

[2415] Anonymous. Front matter. Dædalus, 123(4):??, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027262.

Anonymous:1994:VI

[2416] Anonymous. Volume information. Dædalus, 123(4):??, Fall 1994. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027261.

G:1995:PIAa

[2417] S. R. G. Preface to the issue “An American Dilemma Revisited”. Dædalus, 124(1):v–xxxiv, Winter 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027280.

Bok:1995:I

[2418] Sissela Bok. Introduction. Dædalus, 124(1):1–13, Winter 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027281.

Dentler:1995:PSA

[2419] Robert A. Dentler. The political sociology of the African American situation: Gunnar Myrdal’s era and today. Dædalus, 124(1):15–36, Winter 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027282.

Ferguson:1995:SCF

[2420] Ronald F. Ferguson. Shifting challenges: Fifty years of economic change toward black-white earnings equality. Dædalus, 124(1):37–76, Winter 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027283.

Allen:1995:AAE

[2421] Walter R. Allen and Joseph O. Jewell. African American education since “An American Dilemma”. Dædalus, 124(1):77–100, Winter 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027284.

Clayton:1995:CSC

[2422] Obie Clayton, Jr. The churches and social change: Accommodation, moderation, or protest. Dædalus, 124 (1):101–117, Winter 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027285.

Spohn:1995:CSP

[2423] Cassia C. Spohn. Courts, sentences, and prisons. Dædalus, 124(1):119–143, Winter 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027286.

Darity:1995:UAU

[2424] William Darity, Jr. The undesirables, America’s underclass in the managerial age: Beyond the Myrdal theory of racial inequality. Dædalus, 124 (1):145–165, Winter 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027287.

Wilkinson:1995:GSI

[2425] Doris Y. Wilkinson. Gender and social inequality: The prevailing significance of race. Dædalus, 124(1):167–178, Winter 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027288.

McDaniel:1995:DRC

[2426] Antonio McDaniel. The dynamic racial composition of the United States. Dædalus, 124(1):179–198, Winter 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027289.

Butler:1995:MRN

[2427] John Sibley Butler. Myrdal revisited: The Negro in business. Dædalus, 124 (1):199–221, Winter 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027290.

Anonymous:1995:BMa

[2428] Anonymous. Back matter. Dædalus, 124(1):??, Winter 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027291.

Anonymous:1995:FMa

[2429] Anonymous. Front matter. Dædalus, 124(1):??, Winter 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027279.

G:1995:PIW

[2430] S. R. G. Preface to the issue “What Future for the State?”. Dædalus, 124(2):v–xxiv, Spring 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027294.

Tamamoto:1995:RJP

[2431] Masaru Tamamoto. Reflections on Japan’s postwar state. Dædalus, 124(2):1–22, Spring 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027295.

Cable:1995:DNS

[2432] Vincent Cable. The diminished nationstate: a study in the loss of economic power. Dædalus, 124(2):23–53, Spring 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027296.

Strange:1995:DS

[2433] Susan Strange. The defective state. Dædalus, 124(2):55–74, Spring 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027297.

Schmidt:1995:NWO

[2434] Vivien A. Schmidt. The new world order, incorporated: The rise of business and the decline of the nationstate. Dædalus, 124(2):75–106, Spring 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027298.

Brubaker:1995:NMN

[2435] Rogers Brubaker. National minorities, nationalizing states, and external national homelands in the new Europe. Dædalus, 124(2):107–132, Spring 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027299.

Hann:1995:SSS

[2436] Chris Hann. Subverting strong states: The dialectics of social engineering in Hungary and Turkey. Dædalus, 124 (2):133–153, Spring 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027300.

Eley:1995:WTC

[2437] Geoff Eley. War and the Twentieth-Century state. Dædalus, 124(2):155– 174, Spring 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027301.

DelRosso:1995:ISR

[2438] Stephen J. Del Rosso, Jr. The insecure state: Reflections on “The State” and “Security” in a changing world. Dædalus, 124(2):175–207, Spring 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027302.

Deudney:1995:NWW

[2439] Daniel Deudney. Nuclear weapons and the waning of the real-state. Dædalus, 124(2):209–231, Spring 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027303.

Anonymous:1995:BMb

[2440] Anonymous. Back matter. Dædalus, 124(2):??, Spring 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027304.

Anonymous:1995:FMb

[2441] Anonymous. Front matter. Dædalus, 124(2):??, Spring 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027293.

G:1995:PIQ

[2442] S. R. G. Preface to the issue “The Quest for World Order”. Dædalus, 124(3):v–xx, Summer 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027307.

Hoffmann:1995:RCC

[2443] Stanley Hoffmann. Report of the Conference on Conditions of World Order: June 12–19, 1965, Villa Serbelloni, Bellagio, Italy. Dædalus, 124 (3):1–26, Summer 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027308.

Aron:1995:AOP

[2444] Raymond Aron. The anarchical order of power. Dædalus, 124(3):27–52, Summer 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027309.

Rothschild:1995:WS

[2445] Emma Rothschild. What is security? Dædalus, 124(3):53–98, Summer 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027310.

Kissinger:1995:WLA

[2446] Henry A. Kissinger. We live in an age of transition. Dædalus, 124 (3):99–107, Summer 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027311.

Kepel:1995:IVS

[2447] Gilles Kepel. Islamists versus the state in Egypt and Algeria. Dædalus, 124 (3):109–127, Summer 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027312.

Widner:1995:SSL

[2448] Jennifer A. Widner. States and statelessness in late Twentieth-Century Africa. Dædalus, 124(3):129–153, Summer 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027313.

Apter:1995:DUC

[2449] David E. Apter. Democracy for Uganda: a case for comparison. Dædalus, 124(3):155–190, Summer 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027314.

Jewsiewicki:1995:MCL

[2450] Bogumil Jewsiewicki, V. Y. Mudimbe, Kathryn Barrett-Gaines, and Nadine LeMeur. Meeting the challenge of legitimacy: Post-independence black African and post-Soviet European states. Dædalus, 124(3):191–207, Summer 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027315.

Rosenberg:1995:SIP

[2451] Göran Rosenberg. Sweden and its immigrants: Policies versus opinions. Dædalus, 124(3):209–218, Summer 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027316.

delaMothe:1995:TGC

[2452] John de la Mothe and Paul R. Dufour. Techno-globalism and the challenges to science and technology policy. Dædalus, 124(3):219–235, Summer 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027317.

Anonymous:1995:BMc

[2453] Anonymous. Back matter. Dædalus, 124(3):??, Summer 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027318.

Anonymous:1995:FMc

[2454] Anonymous. Front matter. Dædalus, 124(3):??, Summer 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027306.

G:1995:PIAb

[2455] S. R. G. Preface to the issue “American Education: Still Separate, Still Unequal”. Dædalus, 124(4):v–xxx, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027322.

Holton:1995:F

[2456] Gerald Holton. Foreword. Dædalus, 124(4):xxxi–xxxvii, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027323.

Holton:1995:WAE

[2457] Gerald Holton and Daniel Goroff. Where is American education going? Report on a convocation. Dædalus, 124(4):1–42, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027324.

Graham:1995:BS

[2458] Patricia Albjerg Graham. Battleships and schools. Dædalus, 124(4):43–46, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027325.

Shanker:1995:ERW

[2459] Albert Shanker. Education reform: What’s not being said. Dædalus, 124(4):47–54, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027326.

Resnick:1995:AEN

[2460] Lauren B. Resnick. From aptitude to effort: a new foundation for our schools. Dædalus, 124(4):55–62, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027327.

Nathan:1995:AAL

[2461] Linda Nathan. Assessing assessment:

Lessons of innovative practice in urban schools. Dædalus, 124(4):63–68, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027328.

Howe:1995:PSI

[2462] Harold Howe II. Priority strategies for improved learning. Dædalus, 124(4):69–76, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027329.

Sizer:1995:S

[2463] Theodore R. Sizer. Silences. Dædalus, 124(4):77–83, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027330.

Howard:1995:YCG

[2464] Jeff Howard. You can’t get there from here: The need for a new logic in education reform. Dædalus, 124(4):85–92, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027331.

Kunin:1995:ERS

[2465] Madeleine M. Kunin. Education reform: Staking out common ground. Dædalus, 124(4):93–99, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027332.

Gardner:1995:LVL

[2466] Howard Gardner. Limited visions, limited means: Two obstacles to meaningful education reform. Dædalus, 124(4):101–105, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027333.

Csikszentmihalyi:1995:ETF

[2467] Mihaly Csikszentmihalyi. Education for the Twenty-First Century. Dædalus, 124(4):107–114, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027334.

Schwarz:1995:CTE

[2468] Paul Schwarz. A call for true and equitable discourse. Dædalus, 124(4):115– 117, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027335.

Dimond:1995:NCA

[2469] Paul R. Dimond. A new challenge for American education. Dædalus, 124(4):119–129, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027336.

Jennings:1995:MFS

[2470] John F. Jennings. Matching form and substance in American education. Dædalus, 124(4):131–133, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027337.

Robinson:1995:LLP

[2471] Sharon P. Robinson. Life, literacy, and the pursuit of challenges. Dædalus, 124(4):135–142, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027338.

Latanision:1995:ERP

[2472] R. M. Latanision. Education reform and the public will. Dædalus, 124(4):143–148, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027339.

DAuria:1995:TWS

[2473] John D’Auria. Tremors we should not ignore. Dædalus, 124(4):149–152, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027340.

Darling-Hammond:1995:RSS

[2474] Linda Darling-Hammond. Restructuring schools for student success. Dædalus, 124(4):153–162, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027341.

Soder:1995:AEF

[2475] Roger Soder. American education: Facing up to unspoken assumptions. Dædalus, 124(4):163–167, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027342.

Houston:1995:SRT

[2476] Paul D. Houston. School reform through a wide-angle lens: The consideration of context. Dædalus, 124 (4):169–172, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027343.

Graubard:1995:WNK

[2477] Stephen R. Graubard. We need to know more. Dædalus, 124(4):173–178, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027344.

Gluck:1995:OW

[2478] Carol Gluck. Had I one wish. Dædalus, 124(4):179–182, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027345.

Anonymous:1995:BMd

[2479] Anonymous. Back matter. Dædalus, 124(4):??, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027346.

Anonymous:1995:FMd

[2480] Anonymous. Front matter. Dædalus, 124(4):??, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027321.

Anonymous:1995:VI

[2481] Anonymous. Volume information. Dædalus, 124(4):??, Fall 1995. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027320.

G:1996:PIS

[2482] S. R. G. Preface to the issue “Social Suffering”. Dædalus, 125(1):v–x, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027349.

Kleinman:1996:I

[2483] Arthur Kleinman, Veena Das, and Margaret Lock. Introduction. Dædalus, 125(1):xi–xx, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027350.

Kleinman:1996:AED

[2484] Arthur Kleinman and Joan Kleinman.

The appeal of experience; the dismay of images: Cultural appropriations of suffering in our times. Dædalus, 125(1):1–23, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027351.

Morris:1996:ASV

[2485] David B. Morris. About suffering: Voice, genre, and moral community. Dædalus, 125(1):25–45, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027352.

Langer:1996:AVS

[2486] Lawrence L. Langer. The alarmed vision: Social suffering and Holocaust atrocity. Dædalus, 125(1):47–65, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027353.

Das:1996:LBT

[2487] Veena Das. Language and body: Transactions in the construction of pain. Dædalus, 125(1):67–91, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027354.

Cavell:1996:CVD

[2488] Stanley Cavell. Comments on veena Das’s essay “Language and Body: Transactions in the Construction of Pain”. Dædalus, 125(1):93–98, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027355.

Ramphele:1996:PWS

[2489] Mamphela Ramphele. Political widowhood in South Africa: The embodiment of ambiguity. Dædalus, 125 (1):99–117, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027356.

Schwarcz:1996:PSP

[2490] Vera Schwarcz. The pane of sorrow: Public uses of personal grief in modern China. Dædalus, 125 (1):119–148, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027357.

Wei-ming:1996:DWI

[2491] Tu Wei-ming. Destructive will and ideological Holocaust: Maoism as a source of social suffering in China. Dædalus, 125(1):149–179, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027358.

Harrington:1996:USM

[2492] Anne Harrington. Unmasking suffering’s masks: Reflections on old and new memories of Nazi medicine. Dædalus, 125(1):181–205, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027359.

Lock:1996:DSR

[2493] Margaret Lock. Displacing suffering:

The reconstruction of death in North America and Japan. Dædalus, 125 (1):207–244, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027360.

Young:1996:SOT

[2494] Allan Young. Suffering and the origins of traumatic memory. Dædalus, 125(1):245–260, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027361.

Farmer:1996:SSV

[2495] Paul Farmer. On suffering and structural violence: a view from below. Dædalus, 125(1):261–283, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027362.

Anonymous:1996:BMa

[2496] Anonymous. Back matter. Dædalus, 125(1):??, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027363.

Anonymous:1996:FMa

[2497] Anonymous. Front matter. Dædalus, 125(1):??, Winter 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027348.

G:1996:PIM

[2498] S. R. G. Preface to the issue “Managing Innovation”. Dædalus, 125(2):v–viii, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20013434.

Chandler:1996:IEA

[2499] Alfred D. Chandler, Jr. Introduction: Entrepreneurial achievements. Dædalus, 125(2):ix–xxix, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20013435.

Taplin:1996:EI

[2500] John F. Taplin. The education of an inventor. Dædalus, 125(2):1–17, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20013436.

Landau:1996:EMI

[2501] Ralph Landau. Entrepreneurs, managers, and the importance of finance. Dædalus, 125(2):19–37, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013437.

Blout:1996:PDR

[2502] Elkan Blout. Polaroid: Dreams to reality. Dædalus, 125(2):39–53, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013438.

Moore:1996:IMM

[2503] Gordon E. Moore. Intel: Memories and the microprocessor. Dædalus, 125(2):55–80, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013439.

Hatsopoulos:1996:PIM

[2504] George N. Hatsopoulos. A perpetual idea machine. Dædalus, 125(2):81–94, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013440.

Haney:1996:PI

[2505] William M. Haney III. The power of invention. Dædalus, 125(2):95–112, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013441.

Cabot:1996:SHC

[2506] Thomas D. Cabot. A short history of cabot corporation. Dædalus, 125 (2):113–136, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20013442.

Galvin:1996:CLE

[2507] Robert Galvin. Communication: The lever of effectiveness and productivity. Dædalus, 125(2):137–146, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20013443.

Bechtel:1996:RS

[2508] Stephen D. Bechtel, Jr. Reflections on success. Dædalus, 125(2):147–166, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20013444.

Johnson:1996:AC

[2509] Edward C. Johnson 3d. Adventures of a contrarian. Dædalus, 125 (2):167–182, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20013445.

Holton:1996:ASI

[2510] Gerald Holton. On the art of scientific imagination. Dædalus, 125 (2):183–208, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20013446.

Anonymous:1996:BMb

[2511] Anonymous. Back matter. Dædalus, 125(2):??, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013447.

Anonymous:1996:FMb

[2512] Anonymous. Front matter. Dædalus, 125(2):??, Spring 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013433.

G:1996:PIL

[2513] S. R. G. Preface to the issue “The Liberation of the Environment”. Dædalus, 125(3):v–viii, Summer 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027366.

White:1996:I

[2514] Robert M. White. Introduction. Dædalus, 125(3):ix–x, Summer 1996. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20027367.

Ausubel:1996:LE

[2515] Jesse H. Ausubel. The liberation of the environment. Dædalus, 125 (3):1–17, Summer 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027368.

Grubler:1996:TCP

[2516] Arnulf Grübler. Time for a change:

On the patterns of diffusion of innovation. Dædalus, 125(3):19–42, Summer 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027369.

Kates:1996:PTH

[2517] Robert W. Kates. Population, technology, and the human environment: A thread through time. Dædalus, 125 (3):43–71, Summer 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027370.

Waggoner:1996:HML

[2518] Paul E. Waggoner. How much land can ten billion people spare for nature? Dædalus, 125(3):73–93, Summer 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027371.

Nakicenovic:1996:FEC

[2519] Nebojsa Nakićenović. Freeing energy from carbon. Dædalus, 125 (3):95–112, Summer 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027372.

Schipper:1996:LSE

[2520] Lee Schipper. Life-styles and the environment: The case of energy. Dædalus, 125(3):113–138, Summer 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027373.

Ausubel:1996:EES

[2521] Jesse H. Ausubel and Cesare Marchetti. Elektron: Electrical systems in retrospect and prospect. Dædalus, 125 (3):139–169, Summer 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

/www.jstor.org/stable/20027374.

Wernick:1996:MDM

[2522] Iddo K. Wernick, Robert Herman, Shekhar Govind, and Jesse H. Ausubel. Materialization and dematerialization: Measures and trends. Dædalus, 125 (3):171–198, Summer 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027375.

Frosch:1996:TEW

[2523] Robert A. Frosch. Toward the end of waste: Reflections on a new ecology of industry. Dædalus, 125(3): 199–212, Summer 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027376.

Meyer-Abich:1996:HNT

[2524] Klaus Michael Meyer-Abich. Humans in nature: Toward a physiocentric philosophy. Dædalus, 125(3): 213–234, Summer 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027377.

Starr:1996:SHE

[2525] Chauncey Starr. Sustaining the human environment: The next two hundred years. Dædalus, 125(3):235–253, Summer 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027378.

Anonymous:1996:BMc

[2526] Anonymous. Back matter. Dædalus, 125(3):??, Summer 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027379.

Anonymous:1996:FMc

[2527] Anonymous. Front matter. Dædalus, 125(3):??, Summer 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027365.

G:1996:PIB

[2528] S. R. G. Preface to the issue “Books, Bricks, and Bytes”. Dædalus, 125(4):v–viii, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027383.

Lyman:1996:WDL

[2529] Peter Lyman. What is a digital library? Technology, intellectual property, and the public interest. Dædalus, 125(4):1–33, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027384.

Billington:1996:LLC

[2530] James H. Billington. Libraries, the library of Congress, and the information age. Dædalus, 125(4):35–54, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027385.

Okerson:1996:BLT

[2531] Ann Shumelda Okerson. Buy or lease? Two models for scholarly information at the end (or the beginning) of an era. Dædalus, 125(4):55–76, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027386.

Carpenter:1996:LHL

[2532] Kenneth E. Carpenter. A library historian looks at librarianship. Dædalus, 125(4):77–102, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027387.

Young:1996:LCP

[2533] Peter R. Young. Librarianship: a changing profession. Dædalus, 125 (4):103–125, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027388.

Lamm:1996:LPP

[2534] Donald S. Lamm. Libraries and publishers: a partnership at risk. Dædalus, 125(4):127–146, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027389.

Metzl:1996:SCC

[2535] Jamie Frederic Metzl. Searching for the catalog of catalogs. Dædalus, 125(4):147–160, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027390.

Mason:1996:YYK

[2536] Marilyn Gell Mason. The Yin and Yang of knowing. Dædalus, 125(4):161–171, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027391.

Dowlin:1996:CCF

[2537] Kenneth E. Dowlin and Eleanor Shapiro. The centrality of communities to the future of major public libraries. Dædalus, 125(4):173–190, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027392.

Marcum:1996:RCT

[2538] Deanna B. Marcum. Redefining community through the public library. Dædalus, 125(4):191–205, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027393.

Kent:1996:APL

[2539] Susan Goldberg Kent. American public libraries: a long transformative moment. Dædalus, 125(4):207–220, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027394.

Lang:1996:BBL

[2540] Brian Lang. Bricks and bytes: Libraries in flux. Dædalus, 125(4):221– 234, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20027395.

Lor:1996:DMS

[2541] Peter Johan Lor. A distant mirror: The story of libraries in South Africa. Dædalus, 125(4):235–265, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027396.

deSantAnna:1996:LSI

[2542] Affonso Romano de Sant’Anna. Libraries, social inequality, and the challenge of the Twenty-First Century. Dædalus, 125(4):267–281, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027397.

Favier:1996:HFN

[2543] Jean Favier and Mireille M. Dedios.

The history of the French national library. Dædalus, 125(4):283–291, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027398.

Zaitsev:1996:PRL

[2544] Vladimir Zaitsev. Problems of Russian libraries in an age of social change. Dædalus, 125(4):293–306, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027399.

Lehmann:1996:MTP

[2545] Klaus-Dieter Lehmann. Making the transitory permanent: The intellectual heritage in a digitized world of knowledge. Dædalus, 125(4):307–329, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027400.

Leskien:1996:APS

[2546] Hermann Leskien. Allocated parts: The story of libraries in Germany. Dædalus, 125(4):331–352, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027401.

Banerjee:1996:SLI

[2547] Dwarika N. Banerjee. The story of libraries in India. Dædalus, 125(4):353– 361, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20027402.

Anonymous:1996:BMd

[2548] Anonymous. Back matter. Dædalus, 125(4):??, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027403.

Anonymous:1996:FMd

[2549] Anonymous. Front matter. Dædalus, 125(4):??, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027382.

Anonymous:1996:VI

[2550] Anonymous. Volume information. Dædalus, 125(4):??, Fall 1996. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027381.

G:1997:PIAa

[2551] S. R. G. Preface to the issue “American Academic Culture in Transformation: Fifty Years, Four Disciplines”. Dædalus, 126(1):v–x, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027406.

Bender:1997:PIA

[2552] Thomas Bender. Politics, intellect, and the American university, 1945– 1995. Dædalus, 126(1):1–38, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027407.

Solow:1997:HDE

[2553] Robert M. Solow. How did economics get that way and what way did it get? Dædalus, 126(1):39–58, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027408.

Kreps:1997:ECP

[2554] David M. Kreps. Economics: The current position. Dædalus, 126(1):59–85, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027409.

Barber:1997:RAA

[2555] William J. Barber. Reconfigurations in American academic economics: A general Practitioner’s perspective. Dædalus, 126(1):87–103, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027410.

Abrams:1997:TES

[2556] M. H. Abrams. The transformation of English studies: 1930–1995. Dædalus, 126(1):105–131, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027411.

Gallagher:1997:HLC

[2557] Catherine Gallagher. The history of literary criticism. Dædalus, 126 (1):133–153, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027412.

Saldivar:1997:TEA

[2558] José David Saldívar. Tracking English and American literary and cultural criticism. Dædalus, 126(1):155–174, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027413.

Putnam:1997:HCP

[2559] Hilary Putnam. A half century of philosophy, viewed from within. Dædalus, 126(1):175–208, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027414.

Nehamas:1997:TRA

[2560] Alexander Nehamas. Trends in recent American philosophy. Dædalus, 126 (1):209–223, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027415.

Lindblom:1997:PS

[2561] Charles E. Lindblom. Political science in the 1940s and 1950s. Dædalus, 126 (1):225–252, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027416.

Smith:1997:SBW

[2562] Rogers M. Smith. Still blowing in the wind: The American quest for a democratic, scientific political science. Dædalus, 126(1):253–287, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027417.

Schorske:1997:NRH

[2563] Carl E. Schorske. The new rigorism in the human sciences, 1940– 1960. Dædalus, 126(1):289–309, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027418.

Katznelson:1997:SLH

[2564] Ira Katznelson. From the street to the lecture Hall: The 1960s. Dædalus, 126(1):311–332, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027419.

Hollinger:1997:DID

[2565] David A. Hollinger. The disciplines and the identity debates, 1970– 1995. Dædalus, 126(1):333–351, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027420.

Eliot:1997:ILU

[2566] T. S. Eliot. The influence of landscape upon the poet. Dædalus, 126(1):352, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027421.

Anonymous:1997:BMa

[2567] Anonymous. Back matter. Dædalus, 126(1):??, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027422.

Anonymous:1997:FMa

[2568] Anonymous. Front matter. Dædalus, 126(1):??, Winter 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027405.

G:1997:PIH

[2569] S. R. G. Preface to the issue “Human Diversity”. Dædalus, 126(2):v–xxii, Spring 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027425.

Sen:1997:ITW

[2570] Amartya Sen. Indian traditions and the Western imagination. Dædalus, 126(2):1–26, Spring 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027426.

Hann:1997:NSR

[2571] Chris Hann. The nation-state, religion, and uncivil society: Two perspectives from the periphery. Dædalus, 126(2):27–45, Spring 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027427.

Mazower:1997:MLN

[2572] Mark Mazower. Minorities and the league of nations in interwar Europe. Dædalus, 126(2):47–63, Spring 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027428.

Bogdanor:1997:FAP

[2573] Vernon Bogdanor. Forms of autonomy and the protection of minorities. Dædalus, 126(2):65–87, Spring 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027429.

Kramer:1997:MEO

[2574] Martin Kramer. The Middle East, old and new. Dædalus, 126(2):89–112, Spring 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027430.

Giliomee:1997:SDA

[2575] Hermann Giliomee. Surrender without defeat: Afrikaners and the South African “Miracle”. Dædalus, 126 (2):113–146, Spring 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027431.

Tiryakian:1997:WCM

[2576] Edward A. Tiryakian. The wild cards of modernity. Dædalus, 126 (2):147–181, Spring 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027432.

Lindholm:1997:USF

[2577] Charles Lindholm and John A. Hall. Is the United States falling apart? Dædalus, 126(2):183–209, Spring 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027433.

Edington:1997:FFP

[2578] Mark D. W. Edington. The fog of the familiar paradigm: On the dangers of applied theory. Dædalus, 126 (2):211–231, Spring 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027434.

Anonymous:1997:BMb

[2579] Anonymous. Back matter. Dædalus, 126(2):??, Spring 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027435.

Anonymous:1997:FMb

[2580] Anonymous. Front matter. Dædalus, 126(2):??, Spring 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027424.

G:1997:PIN

[2581] S. R. G. Preface to the issue “A New Europe for the Old?”. Dædalus, 126(3):v–xii, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027438.

Malia:1997:NEO

[2582] Martin Malia. A new Europe for the old? Dædalus, 126(3):1–22, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027439.

Judah:1997:SSR

[2583] Tim Judah. The Serbs: The sweet and rotten smell of history. Dædalus, 126(3):23–45, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027440.

Tanner:1997:ICQ

[2584] Marcus Tanner. Illyrianism and the Croatian quest for statehood. Dædalus, 126(3):47–62, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027441.

Gallagher:1997:BRQ

[2585] Tom Gallagher. To be or not to be Balkan: Romania’s quest for selfdefinition. Dædalus, 126(3):63–83, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027442.

Szporluk:1997:UIP

[2586] Roman Szporluk. Ukraine: From an Imperial periphery to a sovereign state. Dædalus, 126(3):85–119, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027443.

Khazanov:1997:ENR

[2587] Anatoly M. Khazanov. Ethnic nationalism in the Russian federation. Dædalus, 126(3):121–142, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027444.

Ischinger:1997:IOV

[2588] Barbara Ischinger. “Im Osten viel Neues”: Plenty of news from the Eastern Länder. Dædalus, 126(3): 143–157, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027445.

Clauss:1997:AIO

[2589] Jan U. Clauss, Daniel Herz, Maike Dury, and Martin Berger. Addendum to “Im Osten viel Neues: Plenty of News from the Eastern Länder”. Dædalus, 126(3):159–166, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027446.

Schmidt:1997:DDI

[2590] Vivien A. Schmidt. Discourse and (dis)integration in Europe: The cases of France, Germany, and great Britain. Dædalus, 126(3):167–197, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027447.

Schnapper:1997:EDC

[2591] Dominique Schnapper and Mireille M.

Dedios. The European debate on citizenship. Dædalus, 126(3):199–222, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027448.

Biocca:1997:NDD

[2592] Dario Biocca. Has the nation died?

The debate over Italy’s identity (and future). Dædalus, 126(3):223–239, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027449.

Ruth:1997:PEC

[2593] Arne Ruth. Postwar Europe: The capriciousness of universal values. Dædalus, 126(3):241–276, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027450.

Anonymous:1997:BMc

[2594] Anonymous. Back matter. Dædalus, 126(3):??, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027451.

Anonymous:1997:FMc

[2595] Anonymous. Front matter. Dædalus, 126(3):??, Summer 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027437.

G:1997:PIAb

[2596] S. R. G. Preface to the issue “The American Academic Profession”. Dædalus, 126(4):v–x, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027455.

Levine:1997:HAP

[2597] Arthur Levine. How the academic profession is changing. Dædalus, 126(4): 1–20, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027456.

Clark:1997:SWD

[2598] Burton R. Clark. Small worlds, different worlds: The uniquenesses and troubles of American academic professions. Dædalus, 126(4):21–42, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027457.

Oakley:1997:EAP

[2599] Francis Oakley. The elusive academic profession: Complexity and change. Dædalus, 126(4):43–66, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027458.

Massey:1997:UCA

[2600] Walter E. Massey. Uncertainties in the changing academic profession. Dædalus, 126(4):67–94, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027459.

Callan:1997:SOA

[2601] Patrick M. Callan. Stewards of opportunity: America’s public community colleges. Dædalus, 126(4):95–112, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027460.

Gumport:1997:PUA

[2602] Patricia J. Gumport. Public universities as academic workplaces. Dædalus, 126(4):113–136, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027461.

Douglas:1997:SFP

[2603] R. M. Douglas. Survival of the fittest?

Postgraduate education and the professoriate at the fin de siècle. Dædalus, 126(4):137–151, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027462.

Goodheart:1997:RCW

[2604] Eugene Goodheart. Reflections on the culture wars. Dædalus, 126(4):153–175, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027463.

Bernstein:1997:BLI

[2605] Charles Bernstein. A blow is like an instrument. Dædalus, 126(4):177–200, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027464.

Labinger:1997:SWF

[2606] Jay A. Labinger. The science wars and the future of the American academic profession. Dædalus, 126(4):201–220, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027465.

Leggon:1997:SA

[2607] Cheryl B. Leggon. The scientist as academic. Dædalus, 126(4):221–244, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027466.

Rothblatt:1997:PKA

[2608] Sheldon Rothblatt. The “Place” of knowledge in the American academic profession. Dædalus, 126(4):245–264, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027467.

Mitchell:1997:BCO

[2609] Theodore R. Mitchell. Border crossings: Organizational boundaries and challenges to the American professoriate. Dædalus, 126(4):265–292, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027468.

Trow:1997:DIT

[2610] Martin Trow. The development of information technology in American higher education. Dædalus, 126 (4):293–314, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027469.

Altbach:1997:IAC

[2611] Philip G. Altbach. An international academic crisis? The American professoriate in comparative perspective. Dædalus, 126(4):315–338, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027470.

Anonymous:1997:BMd

[2612] Anonymous. Back matter. Dædalus, 126(4):??, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027471.

Anonymous:1997:FMd

[2613] Anonymous. Front matter. Dædalus, 126(4):??, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027454.

Anonymous:1997:VI

[2614] Anonymous. Volume information. Dædalus, 126(4):??, Fall 1997. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027453.

G:1998:PIS

[2615] S. R. G. Preface to the issue “Science in Culture”. Dædalus, 127(1):v–x, Winter 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027474.

Holton:1998:ECR

[2616] Gerald Holton. Einstein and the cultural roots of modern science. Dædalus, 127(1):1–44, Winter 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027475.

Galison:1998:AU

[2617] Peter Galison. The Americanization of unity. Dædalus, 127(1):45–71, Winter 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027476.

Daston:1998:FLI

[2618] Lorraine Daston. Fear and loathing of the imagination in science. Dædalus, 127(1):73–95, Winter 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027477.

Doniger:1998:MFI

[2619] Wendy Doniger and Gregory Spinner. Misconceptions: Female imaginations and male fantasies in parental imprinting. Dædalus, 127(1):97–129, Winter 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027478.

Wilson:1998:CAG

[2620] Edward O. Wilson. Consilience among the great branches of learning. Dædalus, 127(1):131–149, Winter 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027479.

Weinberg:1998:PH

[2621] Steven Weinberg. Physics and history. Dædalus, 127(1):151–164, Winter 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027480.

Friedrich:1998:SQO

[2622] Bretislav Friedrich and Dudley Herschbach. Space quantization: Otto Stern’s lucky star. Dædalus, 127 (1):165–191, Winter 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027481.

Gombrich:1998:EIW

[2623] E. H. Gombrich. Eastern inventions and Western response. Dædalus, 127 (1):193–205, Winter 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027482.

Ackerman:1998:LVA

[2624] James S. Ackerman. Leonardo Da Vinci: Art in science. Dædalus, 127 (1):207–224, Winter 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027483.

Graham:1998:EDA

[2625] Patricia Albjerg Graham. Educational dilemmas for Americans. Dædalus, 127 (1):225–236, Winter 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027484.

Anonymous:1998:BMa

[2626] Anonymous. Back matter. Dædalus, 127(1):??, Winter 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027485.

Anonymous:1998:FMa

[2627] Anonymous. Front matter. Dædalus, 127(1):??, Winter 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027473.

G:1998:PIB

[2628] S. R. G. Preface to the issue “The Brain”. Dædalus, 127(2):v–viii, Spring 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027488.

Mountcastle:1998:BSC

[2629] Vernon B. Mountcastle. Brain science at the Century’s ebb. Dædalus, 127(2):1–36, Spring 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027489.

Edelman:1998:BPB

[2630] Gerald M. Edelman. Building a picture of the brain. Dædalus, 127(2):37–69, Spring 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027490.

Zeki:1998:AB

[2631] Semir Zeki. Art and the brain. Dædalus, 127(2):71–103, Spring 1998. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20027491.

Frackowiak:1998:FAB

[2632] Richard S. J. Frackowiak. The functional architecture of the brain. Dædalus, 127(2):105–130, Spring 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027492.

Bear:1998:MMS

[2633] Mark F. Bear and Leon N. Cooper.

From molecules to mental states. Dædalus, 127(2):131–144, Spring 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027493.

Changeux:1998:DUA

[2634] Jean-Pierre Changeux. Drug use and abuse. Dædalus, 127(2):145–165, Spring 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027494.

Borbely:1998:QES

[2635] Alexander A. Borbély and Giulio Tononi. The quest for the essence of sleep. Dædalus, 127(2):167–196, Spring 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027495.

Miklos:1998:EMB

[2636] George L. Gabor Miklos. The evolution and modification of brains and sensory systems. Dædalus, 127(2):197–216, Spring 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027496.

Bizzi:1998:AMB

[2637] Emilio Bizzi and Ferdinando A. Mussa-Ivaldi. The acquisition of motor behavior. Dædalus, 127(2):217–232, Spring 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027497.

Kinsbourne:1998:UDH

[2638] Marcel Kinsbourne. Unity and diversity in the human brain: Evidence from injury. Dædalus, 127 (2):233–256, Spring 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027498.

Clark:1998:WBB

[2639] Andy Clark. Where brain, body, and world collide. Dædalus, 127 (2):257–280, Spring 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027499.

Anonymous:1998:BMb

[2640] Anonymous. Back matter. Dædalus, 127(2):??, Spring 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027501.

Anonymous:1998:FMb

[2641] Anonymous. Front matter. Dædalus, 127(2):??, Spring 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027487.

Anonymous:1998:IB

[2642] Anonymous. Illustrations: The brain. Dædalus, 127(2):??, Spring 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027500.

G:1998:PIEa

[2643] S. R. G. Preface to the issue “Early Modernities”. Dædalus, 127 (3):v–viii, Summer 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027504.

Eisenstadt:1998:IPE

[2644] Shmuel N. Eisenstadt and Wolfgang Schluchter. Introduction: Paths to early modernities: A comparative view. Dædalus, 127(3):1–18, Summer 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027505.

Wittrock:1998:EMV

[2645] Björn Wittrock. Early modernities:

Varieties and transitions. Dædalus, 127(3):19–40, Summer 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027506.

Pollock:1998:IVM

[2646] Sheldon Pollock. India in the vernacular Millennium: Literary culture and polity, 1000–1500. Dædalus, 127 (3):41–74, Summer 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027507.

Subrahmanyam:1998:HVV

[2647] Sanjay Subrahmanyam. Hearing voices: Vignettes of early modernity in South Asia, 1400–1750. Dædalus, 127(3):75–104, Summer 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

                                        /www.jstor.org/stable/20027508.

Howell:1998:TCI

[2648] David L. Howell. Territoriality and collective identity in Tokugawa Japan. Dædalus, 127(3):105–132, Summer 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027509.

Berry:1998:PLA

[2649] Mary Elizabeth Berry. Public life in authoritarian Japan. Dædalus, 127 (3):133–165, Summer 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027510.

Wakeman:1998:BPS

[2650] Frederic Wakeman, Jr. Boundaries of the public sphere in ming and qing China. Dædalus, 127(3):167–189, Summer 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027511.

Woodside:1998:TOC

[2651] Alexander Woodside. Territorial order and collective-identity tensions in Confucian Asia: China, Vietnam, Korea. Dædalus, 127(3):191–220, Summer 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027512.

Giesen:1998:CPJ

[2652] Bernhard Giesen. Cosmopolitans, patriots, Jacobins, and romantics. Dædalus, 127(3):221–250, Summer 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027513.

Perez-Diaz:1998:SPS

[2653] Víctor Pérez-Díaz. State and public sphere in Spain during the ancient regime. Dædalus, 127(3):251–279, Summer 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027514.

Anonymous:1998:BMc

[2654] Anonymous. Back matter. Dædalus, 127(3):??, Summer 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027515.

Anonymous:1998:FMc

[2655] Anonymous. Front matter. Dædalus, 127(3):??, Summer 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027503.

G:1998:PIEb

[2656] S. R. G. Preface to the issue “Education Yesterday, Education Tomorrow”. Dædalus, 127(4):v–xxiv, Fall 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027519.

Sarason:1998:SFF

[2657] Seymour B. Sarason. Some features of a flawed educational system. Dædalus, 127(4):1–12, Fall 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027520.

Louis:1998:LFC

[2658] Karen Seashore Louis. “A Light Feeling of Chaos”: Educational reform and policy in the United States. Dædalus, 127(4):13–40, Fall 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027521.

Bromley:1998:EIP

[2659] Daniel W. Bromley. Expectations, incentives, and performance in America’s schools. Dædalus, 127(4):41–66, Fall 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027522.

Hampel:1998:GC

[2660] Robert L. Hampel. A generation in crisis? Dædalus, 127(4):67–88, Fall 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027523.

Resnick:1998:LOS

[2661] Lauren B. Resnick and Megan Williams Hall. Learning organizations for sustainable education reform. Dædalus, 127(4):89–118, Fall 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027524.

Marsick:1998:TLE

[2662] Victoria J. Marsick. Transformative learning from experience in the knowledge era. Dædalus, 127(4):119–136, Fall 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027525.

Fox:1998:LDM

[2663] Daniel M. Fox and John M. Ludden. Living but not dying by the market: Recent changes in health care. Dædalus, 127(4):137–158, Fall 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027526.

Huffman:1998:MAC

[2664] Wallace E. Huffman. Modernizing agriculture: a continuing process. Dædalus, 127(4):159–186, Fall 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027527.

John:1998:PI

[2665] Richard R. John. The politics of innovation. Dædalus, 127(4):187–214, Fall 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027528.

Hughes:1998:DDR

[2666] Thomas P. Hughes. Designing, developing, and reforming systems. Dædalus, 127(4):215–232, Fall 1998. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20027529.

Wilson:1998:ARD

[2667] Kenneth G. Wilson and Constance K. Barsky. Applied research and development: Support for continuing improvement in education. Dædalus, 127(4):233–258, Fall 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027530.

Anonymous:1998:BMd

[2668] Anonymous. Back matter. Dædalus, 127(4):??, Fall 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027531.

Anonymous:1998:FMd

[2669] Anonymous. Front matter. Dædalus, 127(4):??, Fall 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027518.

Anonymous:1998:VI

[2670] Anonymous. Volume information. Dædalus, 127(4):??, Fall 1998. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027517.

Anonymous:1999:FYD

[2671] Anonymous. Forty years of “Dædalus”. Dædalus, 128(??):1–4, ???? 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/

stable/20027605. Special issue with the Forty-Year Index 1958–1998.

Anonymous:1999:AI

[2672] Anonymous. Author index. Dædalus, 128(??):5–84, ???? 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.jstor.org/stable/20027606.

Special issue with the Forty-Year Index 1958–1998.

Anonymous:1999:SI

[2673] Anonymous. Subject index. Dædalus, 128(??):85–148, ???? 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.jstor.org/stable/20027607.

Special issue with the Forty-Year Index 1958–1998.

Anonymous:1999:IDR

[2674] Anonymous. Issues of “Dædalus” republished as books. Dædalus, 128 (??):149–155, ???? 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.jstor.org/stable/20027608.

Special issue with the Forty-Year Index 1958–1998.

Anonymous:1999:BMe

[2675] Anonymous. Back matter. Dædalus, 128(??):??, ???? 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.jstor.org/stable/20027609.

Special issue with the Forty-Year Index 1958–1998.

Anonymous:1999:FMe

[2676] Anonymous. Front matter. Dædalus, 128(??):??, ???? 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.jstor.org/stable/20027604.

Special issue with the Forty-Year Index 1958–1998.

G:1999:PID

[2677] S. R. G. Preface to the issue “Distinctively American: The Residential Liberal Arts Colleges”. Dædalus, 128(1):v–xii, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027534.

Koblik:1999:F

[2678] Steven Koblik. Foreword. Dædalus, 128(1):xiii–xiv, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027535.

Hawkins:1999:MLA

[2679] Hugh Hawkins. The making of the liberal arts College identity. Dædalus, 128(1):1–25, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027536.

Neely:1999:TLA

[2680] Paul Neely. The threats to liberal arts colleges. Dædalus, 128(1):27–45, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027537.

McPherson:1999:FEC

[2681] Michael S. McPherson and Morton Owen Schapiro. The future economic challenges for the liberal arts colleges. Dædalus, 128(1):47–75, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027538.

Astin:1999:HLA

[2682] Alexander W. Astin. How the liberal arts College affects students. Dædalus, 128(1):77–100, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027539.

Gomes:1999:AAV

[2683] Peter J. Gomes. Affirmation and adaptation: Values and the elite residential College. Dædalus, 128(1):101–119, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027540.

Canada:1999:CDR

[2684] Geoffrey Canada. The currents of democracy: The role of small liberal arts colleges. Dædalus, 128 (1):121–132, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027541.

Lang:1999:DAL

[2685] Eugene M. Lang. Distinctively American: The liberal arts College. Dædalus, 128(1):133–150, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027542.

Brann:1999:ACP

[2686] Eva T. H. Brann. The American College as the place for liberal learning. Dædalus, 128(1):151–171, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027543.

Hersh:1999:GIT

[2687] Richard H. Hersh. Generating ideals and transforming lives: A contemporary case for the residential liberal arts College. Dædalus, 128(1):173–194, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027544.

Cech:1999:SLA

[2688] Thomas R. Cech. Science at liberal arts colleges: a better education? Dædalus, 128(1):195–216, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027545.

Laws:1999:NAS

[2689] Priscilla W. Laws. New approaches to science and mathematics teaching at liberal arts colleges. Dædalus, 128 (1):217–240, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027546.

Sorum:1999:VCT

[2690] Christina Elliott Sorum. “Vortex, Clouds, and Tongue”: New problems in the humanities? Dædalus, 128 (1):241–264, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027547.

Bourque:1999:RRL

[2691] Susan C. Bourque. Reassessing research: Liberal arts colleges and the social sciences. Dædalus, 128(1):265–272, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027548.

Anonymous:1999:BMa

[2692] Anonymous. Back matter. Dædalus, 128(1):??, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027549.

Anonymous:1999:FMa

[2693] Anonymous. Front matter. Dædalus, 128(1):??, Winter 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027533.

G:1999:PIN

[2694] S. R. G. Preface to the issue “The Next Generation: Work in Progress”. Dædalus, 128(2):v–vi, Spring 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027552.

Epstein:1999:PBC

[2695] Catherine Epstein. The politics of biography: The case of East German old communists. Dædalus, 128(2):1–30, Spring 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027553.

Power:1999:SC

[2696] Samantha Power. To suffer by comparison? Dædalus, 128(2):31–66, Spring 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027554.

Todeschini:1999:ISB

[2697] Maya Todeschini. Illegitimate sufferers: A-bomb victims, medical science, and the government. Dædalus, 128(2):67–100, Spring 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027555.

Sferza:1999:WLL

[2698] Serenella Sferza. What is left of the left? More than one would think. Dædalus, 128(2):101–126, Spring 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027556.

Masoud:1999:AIT

[2699] Tarek E. Masoud. The Arabs and Islam: The troubled search for legitimacy. Dædalus, 128(2):127–145, Spring 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027557.

McGray:1999:SA

[2700] Douglas McGray. The Silicon Archipelago. Dædalus, 128(2):147–176, Spring 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027558.

Metzl:1999:PD

[2701] Jamie Frederic Metzl. Popular diplomacy. Dædalus, 128(2):177–192, Spring 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027559.

Burnett:1999:VBT

[2702] D. Graham Burnett. A view from the bridge: The two cultures debate, its legacy, and the history of science. Dædalus, 128(2):193–218, Spring 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027560.

Anonymous:1999:BMb

[2703] Anonymous. Back matter. Dædalus, 128(2):??, Spring 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027561.

Anonymous:1999:FMb

[2704] Anonymous. Front matter. Dædalus, 128(2):??, Spring 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027551.

G:1999:PIA

[2705] S. R. G. Preface to the issue “America’s Museums”. Dædalus, 128 (3):v–xiv, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027564.

Pitman:1999:MMM

[2706] Bonnie Pitman. Muses, museums, and memories. Dædalus, 128(3):1–31, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027565.

Harris:1999:DHA

[2707] Neil Harris. The divided house of the American art museum. Dædalus, 128(3):33–56, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027566.

Noriega:1999:MRA

[2708] Chon A. Noriega. On museum row: Aesthetics and the politics of exhibition. Dædalus, 128(3):57–81, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027567.

McLean:1999:MED

[2709] Kathleen McLean. Museum exhibitions and the dynamics of dialogue. Dædalus, 128(3):83–107, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027568.

Skramstad:1999:AAM

[2710] Harold Skramstad. An agenda for American museums in the Twenty-First Century. Dædalus, 128(3): 109–128, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027569.

Anderson:1999:MFI

[2711] Maxwell L. Anderson. Museums of the future: The impact of technology on museum practices. Dædalus, 128 (3):129–162, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027570.

Gurian:1999:WOE

[2712] Elaine Heumann Gurian. What is the object of this exercise? A meandering exploration of the many meanings of objects in museums. Dædalus, 128 (3):163–183, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027571.

Boyd:1999:MCC

[2713] Willard L. Boyd. Museums as centers of controversy. Dædalus, 128 (3):185–228, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027572.

Weil:1999:BAS

[2714] Stephen E. Weil. From being about something to being for somebody: The ongoing transformation of the American museum. Dædalus, 128 (3):229–258, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027573.

Falk:1999:MIP

[2715] John H. Falk. Museums as institutions for personal learning. Dædalus, 128 (3):259–275, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027574.

Koster:1999:SRS

[2716] Emlyn H. Koster. In search of relevance: Science centers as innovators in the evolution of museums. Dædalus, 128(3):277–296, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027575.

Sirefman:1999:FFC

[2717] Susanna Sirefman. Formed and forming: Contemporary museum architecture. Dædalus, 128(3):297–320, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027576.

Newhouse:1999:IMP

[2718] Victoria Newhouse. Is “The Idea of a Museum” possible today? Dædalus, 128(3):321–326, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027577.

Correa:1999:MAT

[2719] Charles Correa. Museums: an alternate typology. Dædalus, 128(3): 327–332, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027578.

Tschumi:1999:MTF

[2720] Bernard Tschumi. On the museum of the Twenty-First Century: an homage to Italo Calvino’s “Invisible Cities”. Dædalus, 128(3):333–337, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20027579.

Anonymous:1999:BMc

[2721] Anonymous. Back matter. Dædalus, 128(3):??, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027580.

Anonymous:1999:FMc

[2722] Anonymous. Front matter. Dædalus, 128(3):??, Summer 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027563.

G:1999:PIB

[2723] S. R. G. Preface to the issue “Bioethics and Beyond”. Dædalus, 128(4):v–vi, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027584.

Kleinman:1999:I

[2724] Arthur Kleinman, Renée C. Fox, and Allan M. Brandt. Introduction. Dædalus, 128(4):vii–x, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027585.

Fox:1999:MEA

[2725] Renée C. Fox. Is medical education asking too much of bioethics? Dædalus, 128(4):1–25, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027586.

Rosenberg:1999:MPM

[2726] Charles E. Rosenberg. Meanings, policies, and medicine: On the bioethical enterprise and history. Dædalus, 128(4):27–46, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027587.

Bosk:1999:PEA

[2727] Charles L. Bosk. Professional ethicist available: Logical, secular, friendly. Dædalus, 128(4):47–68, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027588.

Kleinman:1999:MEE

[2728] Arthur Kleinman. Moral experience and ethical reflection: Can ethnography reconcile them? A quandary for “The New Bioethics”. Dædalus, 128(4):69–97, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027589.

Das:1999:PGE

[2729] Veena Das. Public good, ethics, and everyday life: Beyond the boundaries of bioethics. Dædalus, 128(4):99–133, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027590.

Cohen:1999:WIH

[2730] Lawrence Cohen. Where it hurts: Indian material for an ethics of organ transplantation. Dædalus, 128 (4):135–165, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027591.

Good:1999:CRM

[2731] Mary-Jo DelVecchio Good, Esther Mwaikambo, Erastus Amayo, and James M’Imunya Machoki. Clinical realities and moral dilemmas: Contrasting perspectives from academic medicine in Kenya, Tanzania, and America. Dædalus, 128(4):167–196, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027592.

Christakis:1999:PB

[2732] Nicholas A. Christakis. Prognostication and bioethics. Dædalus, 128 (4):197–214, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027593.

Daniels:1999:WJG

[2733] Norman Daniels, Bruce P. Kennedy, and Ichiro Kawachi. Why justice is good for our health: The social determinants of health inequalities. Dædalus, 128(4):215–251, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027594.

Churchill:1999:WPC

[2734] Larry R. Churchill. Are we professionals? A critical look at the social role of bioethicists. Dædalus, 128(4):253–274, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027595.

Callahan:1999:SST

[2735] Daniel Callahan. The social sciences and the task of bioethics. Dædalus, 128(4):275–294, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027596.

Capron:1999:WCS

[2736] Alexander Morgan Capron. What contributions have social science and the law made to the development of policy on bioethics? Dædalus, 128(4):295–325, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027597.

Anonymous:1999:BMd

[2737] Anonymous. Back matter. Dædalus, 128(4):??, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027598.

Anonymous:1999:FMd

[2738] Anonymous. Front matter. Dædalus, 128(4):??, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027583.

Anonymous:1999:VI

[2739] Anonymous. Volume information. Dædalus, 128(4):??, Fall 1999. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027582.

G:2000:PIMa

[2740] S. R. G. Preface to the issue “Multiple Modernities”. Dædalus, 129(1):v–xii, Winter 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027612.

Eisenstadt:2000:MM

[2741] S. N. Eisenstadt. Multiple modernities. Dædalus, 129(1):1–29, Winter 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027613.

Wittrock:2000:MON

[2742] Björn Wittrock. Modernity: One, none, or many? European origins and modernity as a global condition. Dædalus, 129(1):31–60, Winter 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027614.

Arnason:2000:CM

[2743] Johann P. Arnason. Communism and modernity. Dædalus, 129(1):61–90, Winter 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027615.

Gole:2000:SIM

[2744] Nilüfer Göle. Snapshots of Islamic modernities. Dædalus, 129(1):91–117, Winter 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027616.

Eickelman:2000:ILM

[2745] Dale F. Eickelman. Islam and the languages of modernity. Dædalus, 129 (1):119–135, Winter 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027617.

Kaviraj:2000:MPI

[2746] Sudipta Kaviraj. Modernity and politics in India. Dædalus, 129(1):137–162, Winter 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027618.

Tambiah:2000:TMD

[2747] Stanley J. Tambiah. Transnational movements, diaspora, and multiple modernities. Dædalus, 129(1):163–194, Winter 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027619.

Weiming:2000:IRC

[2748] Tu Weiming. Implications of the rise of “Confucian” East Asia. Dædalus, 129(1):195–218, Winter 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027620.

Heideking:2000:PAM

[2749] Jürgen Heideking. The pattern of American modernity from the revolution to the Civil War. Dædalus, 129 (1):219–247, Winter 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027621.

Ortiz:2000:IMW

[2750] Renato Ortiz. From incomplete modernity to world modernity. Dædalus, 129 (1):249–260, Winter 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027622.

Anonymous:2000:BMa

[2751] Anonymous. Back matter. Dædalus, 129(1):??, Winter 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027623.

Anonymous:2000:FMa

[2752] Anonymous. Front matter. Dædalus, 129(1):??, Winter 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027611.

G:2000:PIB

[2753] S. R. G. Preface to the issue “Brazil:

The Burden of the past; The Promise of the Future”. Dædalus, 129(2):v–x, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027626.

Bethell:2000:PBE

[2754] Leslie Bethell. Politics in Brazil: From elections without democracy to democracy without citizenship. Dædalus, 129(2):1–27, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027627.

Schwartzman:2000:BSA

[2755] Simon Schwartzman. Brazil: The social agenda. Dædalus, 129(2):29–56, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027628.

deCarvalho:2000:DCU

[2756] José Murilo de Carvalho. Dreams come untrue. Dædalus, 129(2):57–82, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027629.

Fry:2000:PNM

[2757] Peter Fry. Politics, nationality, and the meanings of “Race” in Brazil. Dædalus, 129(2):83–118, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027630.

Pinheiro:2000:DGV

[2758] Paulo Sérgio Pinheiro. Democratic governance, violence, and the (un)rule of law. Dædalus, 129(2):119–143, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027631.

Stepan:2000:BDF

[2759] Alfred Stepan. Brazil’s decentralized federalism: Bringing government closer to the citizens? Dædalus, 129 (2):145–169, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027632.

Reis:2000:MCS

[2760] Elisa P. Reis. Modernization, citizenship, and stratification: Historical processes and recent changes in Brazil. Dædalus, 129(2):171–194, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027633.

Martins:2000:MTC

[2761] Luciano Martins. Muddling through changing references: From late nationbuilding to the crisis of the nationstate. Dædalus, 129(2):195–206, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027634.

Lafer:2000:BII

[2762] Celso Lafer. Brazilian international identity and foreign policy: Past, present, and future. Dædalus, 129 (2):207–238, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027635.

Lopes:2000:CEC

[2763] José Sergio Leite Lopes. Class, ethnicity, and color in the making of Brazilian football. Dædalus, 129(2):239–270, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027636.

Birman:2000:WHW

[2764] Patrícia Birman and Márcia Pereira Leite. Whatever happened to what used to be the largest Catholic country in the world? Dædalus, 129 (2):271–290, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027637.

Castro:2000:EWB

[2765] Claudio de Moura Castro. Education: Way behind but trying to catch up. Dædalus, 129(2):291–314, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027638.

daCunha:2000:IPT

[2766] Manuela Carneiro da Cunha and Mauro W. B. de Almeida. Indigenous people, traditional people, and conservation in the Amazon. Dædalus, 129(2):315–338, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027639.

Fishlow:2000:BER

[2767] Albert Fishlow. Brazil and economic realities. Dædalus, 129(2):339–357, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027640.

Anonymous:2000:BMb

[2768] Anonymous. Back matter. Dædalus, 129(2):??, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027641.

Anonymous:2000:FMb

[2769] Anonymous. Front matter. Dædalus, 129(2):??, Spring 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027625.

G:2000:PIMb

[2770] S. R. G. Preface to the issue “Minnesota: a Different America?”. Dædalus, 129(3):v–x, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027644.

Gilman:2000:HPM

[2771] Rhoda R. Gilman. The history and peopling of Minnesota: Its culture. Dædalus, 129(3):1–29, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027645.

Atkins:2000:FM

[2772] Annette Atkins. Facing Minnesota. Dædalus, 129(3):31–53, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027646.

Amato:2000:MRI

[2773] Joseph A. Amato and Anthony Amato. Minnesota, real and imagined: a view from the countryside. Dædalus, 129(3):55–80, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027647.

Lanegran:2000:MNP

[2774] David A. Lanegran. Minnesota: Nature’s playground. Dædalus, 129 (3):81–100, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027648.

Adams:2000:MWP

[2775] John S. Adams. Minnesota: a work in progress shaping the landscape and the people. Dædalus, 129(3):101–136, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027649.

Peacock:2000:NWN

[2776] Thomas D. Peacock and Donald R.

Day. Nations within a nation: The Dakota and Ojibwe of Minnesota. Dædalus, 129(3):137–159, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027650.

Chapman:2000:MNO

[2777] Richard M. Chapman. Mixing new and old wine in Minnesota: Spirituality, ecumenism, and religious traditions in ferment. Dædalus, 129 (3):161–190, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027651.

Brandl:2000:PPM

[2778] John E. Brandl. Policy and politics in Minnesota. Dædalus, 129 (3):191–220, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027652.

Gray:2000:IFM

[2779] Virginia Gray and Wyman Spano. The irresistible force meets the immovable object: Minnesota’s moralistic political culture confronts Jesse Ventura. Dædalus, 129(3):221–245, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027653.

OKeefe:2000:SSM

[2780] Michael O’Keefe. Social services: Minnesota as innovator. Dædalus, 129 (3):247–267, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027654.

Pratt:2000:DCP

[2781] Jon Pratt and Edson W. Spencer. Dynamics of corporate philanthropy in Minnesota. Dædalus, 129(3):269–292, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027655.

Dowling:2000:TAP

[2782] Joe Dowling. Theater and the arts:

a personal reflection. Dædalus, 129 (3):293–305, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027656.

White:2000:MWA

[2783] Robert J. White. Minnesota and the world abroad. Dædalus, 129 (3):307–334, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027657.

Anonymous:2000:BMc

[2784] Anonymous. Back matter. Dædalus, 129(3):??, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027658.

Anonymous:2000:FMc

[2785] Anonymous. Front matter. Dædalus, 129(3):??, Summer 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027643.

Shweder:2000:I

[2786] Richard A. Shweder, Martha Minow, and Hazel Rose Markus. Introduction. Dædalus, 129(4):v–ix, Fall 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027661.

Suarez-Orozco:2000:EYE

[2787] Marcelo M. Suárez-Orozco. Everything you ever wanted to know about assimilation but were afraid to ask. Dædalus, 129(4):1–30, Fall 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027662.

Ewing:2000:LRF

[2788] Katherine Pratt Ewing. Legislating religious freedom: Muslim challenges to the relationship between “Church” and “State” in Germany and France. Dædalus, 129(4):31–54, Fall 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027663.

Wikan:2000:CTN

[2789] Unni Wikan. Citizenship on trial: Nadia’s case. Dædalus, 129(4):55–76, Fall 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027664.

Menon:2000:DFU

[2790] Usha Menon. Does feminism have universal relevance? The challenges posed by Oriya Hindu family practices. Dædalus, 129(4):77–99, Fall 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027665.

Chambers:2000:CNM

[2791] David L. Chambers. Civilizing the natives: Marriage in post-apartheid South Africa. Dædalus, 129(4):101– 124, Fall 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027666.

Minow:2000:AWA

[2792] Martha Minow. About women, about culture: About them, about us. Dædalus, 129(4):125–145, Fall 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027667.

Sarat:2000:MID

[2793] Austin Sarat. The micropolitics of identity/difference: Recognition and accommodation in everyday life. Dædalus, 129(4):147–168, Fall 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027668.

Stolzenberg:2000:CP

[2794] Nomi Maya Stolzenberg. The culture of property. Dædalus, 129(4):169–192, Fall 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027669.

Sager:2000:FEC

[2795] Lawrence G. Sager. The free exercise of culture: Some doubts and distinctions. Dædalus, 129(4):193–208, Fall 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027670.

Shweder:2000:WAF

[2796] Richard A. Shweder. What about “Female Genital Mutilation”? And why understanding culture matters in the first place. Dædalus, 129(4):209–232, Fall 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027671.

Markus:2000:CBI

[2797] Hazel Rose Markus, Claude M. Steele, and Dorothy M. Steele. Colorblindness as a barrier to inclusion: Assimilation and nonimmigrant minorities. Dædalus, 129(4):233–259, Fall 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027672.

Anonymous:2000:BMd

[2798] Anonymous. Back matter. Dædalus, 129(4):??, Fall 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027673.

Anonymous:2000:FMd

[2799] Anonymous. Front matter. Dædalus, 129(4):??, Fall 2000. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027660.

G:2001:PIW

[2800] S. R. G. Preface to the issue “Why South Africa Matters”. Dædalus, 130(1):v–viii, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027676.

Ramphele:2001:CCS

[2801] Mamphela Ramphele. Citizenship challenges for South Africa’s Young democracy. Dædalus, 130(1):1–17, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027677.

Vestergaard:2001:WGM

[2802] Mads Vestergaard. Who’s got the map? The negotiation of Afrikaner identities in post-apartheid South Africa. Dædalus, 130(1):19–44, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027678.

Nattrass:2001:TNR

[2803] Nicoli Nattrass and Jeremy Seekings.

“Two Nations”? Race and economic inequality in South Africa today. Dædalus, 130(1):45–70, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027679.

Boyd:2001:ESS

[2804] Leslie Boyd, Michael Spicer, and Gavin Keeton. Economic scenarios for South Africa: a business perspective. Dædalus, 130(1):71–98, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027680.

Wilson:2001:MMH

[2805] Francis Wilson. Minerals and migrants: How the mining industry has shaped South Africa. Dædalus, 130 (1):99–121, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027681.

Simkins:2001:CSA

[2806] Charles Simkins. Can South Africa avoid a Malthusian positive check? Dædalus, 130(1):123–150, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027682.

vanderVliet:2001:ALN

[2807] Virginia van der Vliet. AIDS: Losing “The New Struggle”? Dædalus, 130(1):151–184, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027683.

Asmal:2001:EDS

[2808] Kader Asmal and Wilmot James. Education and democracy in South Africa today. Dædalus, 130(1):185–204, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027684.

Kahn:2001:STS

[2809] Michael J. Kahn and B. Daya Reddy. Science and technology in South Africa: Regional innovation hub or passive consumer? Dædalus, 130(1):205–234, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027685.

Shaw:2001:SSC

[2810] Mark Shaw and Peter Gastrow. Stealing the show? Crime and its impact in post-apartheid South Africa. Dædalus, 130(1):235–258, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027686.

Gastrow:2001:SSP

[2811] Peter Gastrow and Mark Shaw. In search of safety: Police transformation and public responses in South Africa. Dædalus, 130(1):259–275, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027687.

Witz:2001:RPS

[2812] Leslie Witz, Ciraj Rassool, and Gary Minkley. Repackaging the past for South African tourism. Dædalus, 130 (1):277–296, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027688.

Anonymous:2001:BMa

[2813] Anonymous. Back matter. Dædalus, 130(1):??, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027689.

Anonymous:2001:FMa

[2814] Anonymous. Front matter. Dædalus, 130(1):??, Winter 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027675.

G:2001:PII

[2815] S. R. G. Preface to the issues “Italy:

Resilient and Vulnerable”. Dædalus, 130(2):v–xxii, Spring 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027692.

Ciampi:2001:I

[2816] Carlo Azeglio Ciampi. Introduction. Dædalus, 130(2):1–6, Spring 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027693.

Prodi:2001:IE

[2817] Romano Prodi. Italy, Europe. Dædalus, 130(2):7–12, Spring 2001. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20027694.

Padoa-Schioppa:2001:IEF

[2818] Tommaso Padoa-Schioppa. Italy and Europe: a fruitful interaction. Dædalus, 130(2):13–44, Spring 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027695.

Andreatta:2001:ICF

[2819] Filippo Andreatta. Italy at a crossroads: The foreign policy of a medium power after the end of bipolarity. Dædalus, 130(2):45–65, Spring 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027696.

Signorini:2001:IEI

[2820] Luigi Federico Signorini. Italy’s economy: an introduction. Dædalus, 130(2):67–92, Spring 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027697.

Barca:2001:NTP

[2821] Fabrizio Barca. New trends and the policy shift in the Italian mezzogiorno. Dædalus, 130(2):93–113, Spring 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027698.

Schioppa:2001:BPA

[2822] Fiorella Kostoris Padoa Schioppa. Budgetary policies and administrative reform in contemporary Italy. Dædalus, 130(2):115–138, Spring 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027699.

Anonymous:2001:BMb

[2823] Anonymous. Back matter. Dædalus, 130(2):??, Spring 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027700.

Anonymous:2001:FMb

[2824] Anonymous. Front matter. Dædalus, 130(2):??, Spring 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027691.

Berselli:2001:CTI

[2825] Edmondo Berselli. The crisis and transformation of Italian politics. Dædalus, 130(3):1–24, Summer 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027703.

Brunetta:2001:IOL

[2826] Renato Brunetta. Italy’s other left. Dædalus, 130(3):25–45, Summer 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027704.

Violante:2001:IPI

[2827] Luciano Violante. Italian parliamentary institutions: an evolutionary overview. Dædalus, 130(3):47–71, Summer 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027705.

Ranci:2001:DWS

[2828] Costanzo Ranci. Democracy at work: Social participation and the “Third Sector” in Italy. Dædalus, 130(3):73– 84, Summer 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027706.

Berger:2001:CIG

[2829] Suzanne Berger and Richard M. Locke. Il Caso Italiano and globalization. Dædalus, 130(3):85–104, Summer 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027707.

Ovi:2001:CCC

[2830] Alessandro Ovi. Cheese, children, and container cranes: Learning from Reggio Emilia. Dædalus, 130(3):105–118, Summer 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027708.

Cavalli:2001:RPC

[2831] Alessandro Cavalli. Reflections on political culture and the “Italian National Character”. Dædalus, 130(3): 119–137, Summer 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027709.

Livi-Bacci:2001:TFC

[2832] Massimo Livi-Bacci. Too few children and too much family. Dædalus, 130 (3):139–155, Summer 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027710.

Paoli:2001:CIS

[2833] Letizia Paoli. Crime, Italian style.

Dædalus, 130(3):157–185, Summer 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027711.

Anonymous:2001:BMc

[2834] Anonymous. Back matter. Dædalus, 130(3):??, Summer 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027712.

Anonymous:2001:FMc

[2835] Anonymous. Front matter. Dædalus, 130(3):??, Summer 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027702.

Tucker:2001:IEA

[2836] Mary Evelyn Tucker and John A. Grim.

Introduction: The emerging alliance of world religions and ecology. Dædalus, 130(4):1–22, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027715.

Rupp:2001:RMS

[2837] George Rupp. Religion, modern secular culture, and ecology. Dædalus, 130(4):23–30, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027716.

McElroy:2001:PEC

[2838] Michael B. McElroy. Perspectives on environmental change: a basis for action. Dædalus, 130(4):31–57, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027717.

Brown:2001:EDG

[2839] Donald A. Brown. The ethical dimensions of global environmental issues. Dædalus, 130(4):59–76, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027718.

Callicott:2001:MEE

[2840] J. Baird Callicott. Multicultural environmental ethics. Dædalus, 130(4): 77–97, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027719.

Tirosh-Samuelson:2001:NSJ

[2841] Hava Tirosh-Samuelson. Nature in the sources of Judaism. Dædalus, 130(4):99–124, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027720.

McFague:2001:NHR

[2842] Sallie McFague. New house rules:

Christianity, economics, and planetary living. Dædalus, 130(4):125–140, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027721.

Haq:2001:IET

[2843] S. Nomanul Haq. Islam and ecology: Toward retrieval and reconstruction. Dædalus, 130(4):141–177, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027722.

Narayanan:2001:WWW

[2844] Vasudha Narayanan. Water, wood, and wisdom: Ecological perspectives from the Hindu traditions. Dædalus, 130(4):179–206, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027723.

Chapple:2001:LCJ

[2845] Christopher Key Chapple. The living cosmos of Jainism: a traditional science grounded in environmental ethics. Dædalus, 130(4):207–224, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027724.

Swearer:2001:PPP

[2846] Donald K. Swearer. Principles and poetry, places and stories: The resources of Buddhist ecology. Dædalus, 130(4):225–241, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027725.

Weiming:2001:ETN

[2847] Tu Weiming. The ecological turn in new Confucian humanism: Implications for China and the world. Dædalus, 130(4):243–264, Fall 2001. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20027726.

Miller:2001:EDB

[2848] James Miller. Envisioning the daoist body in the economy of cosmic power. Dædalus, 130(4):265–282, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027727.

Forbes:2001:IAS

[2849] Jack D. Forbes. Indigenous Americans: Spirituality and ecos. Dædalus, 130(4):283–300, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027728.

McKibben:2001:WDW

[2850] Bill McKibben. Where do we go from here? Dædalus, 130(4):301–306, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027729.

Anonymous:2001:BMd

[2851] Anonymous. Back matter. Dædalus, 130(4):??, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027730.

Anonymous:2001:FMd

[2852] Anonymous. Front matter. Dædalus, 130(4):??, Fall 2001. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027714.

Katznelson:2002:EP

[2853] Ira Katznelson. Evil & politics. Dædalus, 131(1):7–10, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027733.

Galbraith:2002:PCI

[2854] James K. Galbraith. A perfect crime:

Inequality in the age of globalization. Dædalus, 131(1):11–25, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027734.

Patterson:2002:BCS

[2855] Orlando Patterson. Beyond compassion: Selfish reasons for being unselfish. Dædalus, 131(1):26–38, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027735.

Epstein:2002:AR

[2856] Richard A. Epstein. Against redress. Dædalus, 131(1):39–48, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027736.

Jencks:2002:DIM

[2857] Christopher Jencks. Does inequality matter? Dædalus, 131(1):49–65, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027737.

Wilentz:2002:ALE

[2858] Sean Wilentz. America’s lost egalitarian tradition. Dædalus, 131(1):66–80, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027738.

Crow:2002:UNG

[2859] James F. Crow. Unequal by nature: a Geneticist’s perspective on human differences. Dædalus, 131(1):81–88, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027739.

Mayr:2002:BRC

[2860] Ernst Mayr. The biology of race and the concept of equality. Dædalus, 131(1):89–94, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027740.

Nussbaum:2002:SLI

[2861] Martha C. Nussbaum. Sex, laws, and inequality: What India can teach the United States. Dædalus, 131 (1):95–106, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027741.

Fogel:2002:WGH

[2862] Robert W. Fogel and Chulhee Lee.

Who gets health care? Dædalus, 131 (1):107–117, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027742.

Shapiro:2002:WPD

[2863] Ian Shapiro. Why the poor don’t soak the Rich. Dædalus, 131(1):118–128, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027743.

Bidart:2002:YM

[2864] Frank Bidart. Young Marx. Dædalus, 131(1):129, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027744.

Mukherjee:2002:W

[2865] Bharati Mukherjee. A wedding.

Dædalus, 131(1):131–138, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027745.

Gardner:2002:TFI

[2866] Howard Gardner. On the three faces of intelligence. Dædalus, 131(1):139–142, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027746.

Wilczek:2002:WNR

[2867] Frank Wilczek. On the World’s numerical recipe. Dædalus, 131(1):142–147, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027747.

Nathan:2002:CRF

[2868] David G. Nathan. On clinical research & the future of medicine. Dædalus, 131(1):147–150, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027748.

Smelser:2002:CNE

[2869] Neil J. Smelser. On compiling a new encyclopedia. Dædalus, 131(1):151–154, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027749.

Early:2002:MDV

[2870] Gerald Early. On Miles Davis, Vince Lombardi, & the crisis of masculinity in Mid-Century America. Dædalus, 131(1):154–159, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027750.

Anonymous:2002:BMa

[2871] Anonymous. Back matter. Dædalus, 131(1):??, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027751.

Anonymous:2002:FMa

[2872] Anonymous. Front matter. Dædalus, 131(1):??, Winter 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027732.

Posner:2002:LEI

[2873] Richard A. Posner. The law & economics of intellectual property. Dædalus, 131(2):5–12, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027754.

Boyle:2002:FIE

[2874] James Boyle. Fencing off ideas: Enclosure & the disappearance of the public domain. Dædalus, 131(2):13–25, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027755.

Hesse:2002:RIP

[2875] Carla Hesse. The rise of intellectual property, 700 B.C.–A.D. 2000: an idea in the balance. Dædalus, 131(2):26–45, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027756.

Goldhammer:2002:DC

[2876] Arthur Goldhammer. On Diderot & Condorcet. Dædalus, 131(2):46–47, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027757.

Goldhammer:2002:LBT

[2877] Arthur Goldhammer and Denis Diderot.

Letter on the book trade. Dædalus, 131(2):48–56, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

                                        /www.jstor.org/stable/20027758.

Goldhammer:2002:FCF

[2878] Arthur Goldhammer and The Marquis de Condorcet. Fragments concerning freedom of the Press. Dædalus, 131(2):57–59, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027759.

Goldhammer:2002:PPR

[2879] Arthur Goldhammer and Roger Chartier. Property & privilege in the Republic of letters. Dædalus, 131(2):60–66, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:/

                                        /www.jstor.org/stable/20027760.

Johns:2002:PMP

[2880] Adrian Johns. Pop music pirate hunters. Dædalus, 131(2):67–77, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027761.

Kevles:2002:MMS

[2881] Daniel J. Kevles. Of mice & money: The story of the World’s first animal patent. Dædalus, 131(2):78–88, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027762.

Eisenberg:2002:PVP

[2882] Rebecca S. Eisenberg and Richard R. Nelson. Public vs. proprietary science: a fruitful tension? Dædalus, 131(2):89–101, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027763.

Angell:2002:PPA

[2883] Marcia Angell and Arnold S. Relman.

Patents, profits & American medicine: Conflicts of interest in the testing & marketing of new drugs. Dædalus, 131 (2):102–111, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027764.

Muldoon:2002:UR

[2884] Paul Muldoon. Unapproved road.

Dædalus, 131(2):112–114, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027765.

Busch:2002:RM

[2885] Frederick Busch. The rescue mission. Dædalus, 131(2):115–125, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027766.

Doniger:2002:K

[2886] Wendy Doniger. On the Kamasutra. Dædalus, 131(2):126–129, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027767.

Breiman:2002:MBE

[2887] Leo Breiman. On McCarthyism at Berkeley: The enemy is us. Dædalus, 131(2):130–132, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027768.

Runciman:2002:DST

[2888] W. G. Runciman. On Darwin & sociological theory today. Dædalus, 131 (2):133–137, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027769.

Hoffmann:2002:PLS

[2889] Roald Hoffmann. On poetry & the language of science. Dædalus, 131 (2):137–140, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027770.

Zagorin:2002:TE

[2890] Perez Zagorin. Terrorism & evil. Dædalus, 131(2):141–144, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027771.

Wallace:2002:RWR

[2891] Bruce Wallace. Regarding the word “Race”. Dædalus, 131(2):144–146, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027772.

Johnson:2002:VI

[2892] D. Gale Johnson. The virtues of inequality. Dædalus, 131(2):146–147, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027773.

Anonymous:2002:BMb

[2893] Anonymous. Back matter. Dædalus, 131(2):??, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027774.

Anonymous:2002:FMb

[2894] Anonymous. Front matter. Dædalus, 131(2):??, Spring 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027753.

Ravitch:2002:EAC

[2895] Diane Ravitch. Education after the culture wars. Dædalus, 131(3):5–21, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027777.

Gardner:2002:CCS

[2896] Howard Gardner. Contemporary consciousness & the study of the humanities. Dædalus, 131(3):22–25, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027778.

Sizer:2002:BW

[2897] Theodore R. Sizer. A better way. Dædalus, 131(3):26–29, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027779.

Hirsch:2002:W

[2898] E. D. Hirsch, Jr. Not to worry?

Dædalus, 131(3):30–32, Summer 2002.

CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20027780.

Appleby:2002:ID

[2899] Joyce Appleby. An intractable debate. Dædalus, 131(3):33–35, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027781.

Stimpson:2002:CWC

[2900] Catharine R. Stimpson. The culture wars continue. Dædalus, 131 (3):36–40, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027782.

Meier:2002:VS

[2901] Deborah Meier. A view from the schoolhouse. Dædalus, 131(3):41–44, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027783.

Graham:2002:CN

[2902] Patricia Albjerg Graham. Class notes. Dædalus, 131(3):45–48, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027784.

Mirel:2002:DCE

[2903] Jeffrey Mirel. The decline of civic education. Dædalus, 131(3):49–55, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027785.

Boyers:2002:WCC

[2904] Robert Boyers. Why a common curriculum? Dædalus, 131(3):56–58, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027786.

Bender:2002:RD

[2905] Thomas Bender. Reforming the disciplines. Dædalus, 131(3):59–62, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027787.

Delbanco:2002:IAC

[2906] Andrew Delbanco. It all comes down to the teachers. Dædalus, 131 (3):63–66, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027788.

Buttigieg:2002:G

[2907] Joseph A. Buttigieg. On Gramsci. Dædalus, 131(3):67–70, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027789.

Gramsci:2002:PN

[2908] Antonio Gramsci and Joseph A.

Buttigieg. From the ‘Prison Notebooks’. Dædalus, 131(3):71–83, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027790.

Bloom:2002:EAU

[2909] David E. Bloom and Joel E. Cohen. Education for all: an unfinished revolution. Dædalus, 131(3):84–95, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027791.

Szporluk:2002:DM

[2910] Larissa Szporluk. Death of Magellan. Dædalus, 131(3):97, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027792.

Bell:2002:PF

[2911] Madison Smartt Bell. Petrified forest. Dædalus, 131(3):98–113, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027793.

Carroll:2002:CC

[2912] James Carroll. On the crisis in Catholicism. Dædalus, 131(3):114–116, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027794.

Shweder:2002:RCP

[2913] Richard A. Shweder. On the return of the ‘Civilizing Project’. Dædalus, 131 (3):117–121, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027795.

Kennedy:2002:SC

[2914] Donald Kennedy. On science at a crossroads. Dædalus, 131(3):122–125, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027796.

Sennett:2002:WPD

[2915] Richard Sennett. On welfare & the psychology of dependence. Dædalus, 131 (3):126–131, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027797.

Anonymous:2002:BMc

[2916] Anonymous. Back matter. Dædalus, 131(3):??, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027798.

Anonymous:2002:FMc

[2917] Anonymous. Front matter. Dædalus, 131(3):??, Summer 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027776.

Steinbruner:2002:ULC

[2918] John Steinbruner and Jeffrey Lewis.

The unsettled legacy of the cold war. Dædalus, 131(4):5–10, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027801.

Donoghue:2002:SB

[2919] Denis Donoghue. Speaking of beauty. Dædalus, 131(4):11–20, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027802.

Sontag:2002:AAB

[2920] Susan Sontag. An argument about beauty. Dædalus, 131(4):21–26, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027803.

Plotinus:2002:B

[2921] Plotinus and Stephen MacKenna. On beauty. Dædalus, 131(4):27–34, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027804.

Danto:2002:AB

[2922] Arthur C. Danto. The abuse of beauty. Dædalus, 131(4):35–56, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027805.

Nehamas:2002:ABU

[2923] Alexander Nehamas. The art of being unselfish. Dædalus, 131(4):57–68, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027806.

Hickey:2002:BW

[2924] Dave Hickey. Buying the world. Dædalus, 131(4):69–87, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027807.

Gurstein:2002:EMR

[2925] Rochelle Gurstein. The Elgin Marbles, romanticism & the waning of ‘ideal beauty’. Dædalus, 131(4):88–100, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027808.

Peiss:2002:EEB

[2926] Kathy Peiss. Educating the eye of the beholder: American cosmetics abroad. Dædalus, 131(4):101–109, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027809.

Phillips:2002:UF

[2927] Carl Phillips. The use of force. Dædalus, 131(4):110–111, Fall 2002. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20027810.

Rooke:2002:WCW

[2928] Leon Rooke. The winds of change, the winds of hope, the winds of disaster. Dædalus, 131(4):112–119, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027811.

Sunstein:2002:DDD

[2929] Cass R. Sunstein. On a danger of deliberative democracy. Dædalus, 131(4):120–124, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027812.

Kantrowitz:2002:FUS

[2930] Arthur Kantrowitz. On fear, uncertainty & scientific progress. Dædalus, 131(4):124–128, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027813.

Hiatt:2002:PIC

[2931] Howard Hiatt. On policy initiatives for children. Dædalus, 131(4):128–131, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027814.

Anonymous:2002:BMd

[2932] Anonymous. Back matter. Dædalus, 131(4):??, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027815.

Anonymous:2002:FMd

[2933] Anonymous. Front matter. Dædalus, 131(4):??, Fall 2002. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027800.

Lewis:2003:CGJ

[2934] Anthony Lewis. The challenge of global justice now. Dædalus, 132(1):5–9, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027818.

Nussbaum:2003:CT

[2935] Martha C. Nussbaum. Compassion & terror. Dædalus, 132(1):10–26, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027819.

Hoffmann:2003:WGB

[2936] Stanley Hoffmann. World governance:

Beyond utopia. Dædalus, 132(1):27–35, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027820.

Beitz:2003:WHR

[2937] Charles Beitz. What human rights mean. Dædalus, 132(1):36–46, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027821.

Goldsmith:2003:LI

[2938] Jack Goldsmith and Stephen D. Krasner. The limits of idealism. Dædalus, 132(1):47–63, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027822.

Elshtain:2003:RNM

[2939] Jean Bethke Elshtain. The responsibility of nations: a moral case for coercive justice. Dædalus, 132(1):64–72, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027823.

Bass:2003:AL

[2940] Gary J. Bass. Atrocity & legalism. Dædalus, 132(1):73–82, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027824.

Slaughter:2003:EGG

[2941] Anne-Marie Slaughter. Everyday global governance. Dædalus, 132 (1):83–90, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027825.

Kaysen:2003:CVM

[2942] Carl Kaysen and George Rathjens. The case for a volunteer UN military force. Dædalus, 132(1):91–103, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027826.

Hollander:2003:I

[2943] John Hollander. The institute. Dædalus, 132(1):105, Winter 2003. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20027827.

Morris:2003:E

[2944] Mary Morris. Exchanges. Dædalus, 132(1):106–120, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027828.

Freedman:2003:AU

[2945] Wendy L. Freedman. On the age of the universe. Dædalus, 132(1):122–126, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027829.

Dennett:2003:FFF

[2946] Daniel C. Dennett. On failures of freedom & the fear of science. Dædalus, 132(1):126–130, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027830.

Costello:2003:PIN

[2947] Bonnie Costello. On poetry & the idea of nature. Dædalus, 132(1):131–135, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027831.

Anonymous:2003:BMa

[2948] Anonymous. Back matter. Dædalus, 132(1):??, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027832.

Anonymous:2003:FMa

[2949] Anonymous. Front matter. Dædalus, 132(1):??, Winter 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027817.

Burnett:2003:MTC

[2950] D. Graham Burnett. Mapping time: Chronometry on top of the world. Dædalus, 132(2):5–19, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027835.

Landes:2003:CWN

[2951] David S. Landes. Clocks & the wealth of nations. Dædalus, 132(2):20–26, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027836.

Rosbash:2003:BC

[2952] Michael Rosbash. A biological clock. Dædalus, 132(2):27–36, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027837.

Gold:2003:WTF

[2953] Thomas Gold. Why time flows: The physics of past & future. Dædalus, 132(2):37–40, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027838.

Galison:2003:EPM

[2954] Peter L. Galison and D. Graham Burnett. Einstein, Poincaré & modernity: A conversation. Dædalus, 132 (2):41–55, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027839.

Groh:2003:HBK

[2955] Jennifer M. Groh and Michael S. Gazzaniga. How the brain keeps time. Dædalus, 132(2):56–61, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027840.

Allen:2003:FTA

[2956] Danielle S. Allen. The flux of time in ancient Greece. Dædalus, 132(2):62–73, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027841.

Grafton:2003:DHR

[2957] Anthony Grafton. Dating history:

The Renaissance & the reformation of chronology. Dædalus, 132(2):74–85, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027842.

Miller:2003:TL

[2958] J. Hillis Miller. Time in literature. Dædalus, 132(2):86–97, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027843.

Douglas:2003:TRC

[2959] Mary Douglas, Michael Thompson, and Marco Verweij. Is time running out? The case of global warming. Dædalus, 132(2):98–107, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027844.

Fenn:2003:AET

[2960] Richard Fenn. Apocalypse & the end of time. Dædalus, 132(2):108–112, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027845.

Howe:2003:PSH

[2961] Susan Howe. Poems by Susan Howe. Dædalus, 132(2):113–117, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027846.

Moody:2003:FS

[2962] Rick Moody. Fish story. Dædalus, 132(2):118–130, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027847.

McGinn:2003:MA

[2963] Bernard McGinn. On mysticism & art. Dædalus, 132(2):131–134, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027848.

Tuan:2003:HG

[2964] Yi-Fu Tuan. On human geography. Dædalus, 132(2):134–137, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027849.

Anonymous:2003:BMb

[2965] Anonymous. Back matter. Dædalus, 132(2):??, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027850.

Anonymous:2003:FMb

[2966] Anonymous. Front matter. Dædalus, 132(2):??, Spring 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027834.

Post:2003:CC

[2967] Robert Post. Congress & the Court. Dædalus, 132(3):5–8, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027853.

Carroll:2003:WRS

[2968] James Carroll. Why religion still matters. Dædalus, 132(3):9–13, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027854.

Keddie:2003:SD

[2969] Nikki R. Keddie. Secularism & its discontents. Dædalus, 132(3):14–30, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027855.

Munson:2003:FAM

[2970] Henry Munson. ’fundamentalism’ ancient & modern. Dædalus, 132(3):31– 41, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

(electronic). URL http://www.jstor.
org/stable/20027856.

Marty:2003:ORS

[2971] Martin E. Marty. Our religio-secular world. Dædalus, 132(3):42–48, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027857.

Jefferson:2003:BER

[2972] Thomas Jefferson. A bill for establishing religious freedom. Dædalus, 132(3):49–50, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027858.

Tamimi:2003:RI

[2973] Azzam S. Tamimi. The Renaissance of Islam. Dædalus, 132(3):51–58, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027859.

Yavuz:2003:CT

[2974] M. Hakan Yavuz. The case of Turkey. Dædalus, 132(3):59–61, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027860.

Madan:2003:CI

[2975] T. N. Madan. The case of India. Dædalus, 132(3):62–66, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027861.

Nasr:2003:LMW

[2976] Vali Nasr. Lessons from the Muslim world. Dædalus, 132(3):67–72, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027862.

Galston:2003:JMP

[2977] William Galston. Jews, Muslims & the prospects for pluralism. Dædalus, 132(3):73–77, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027863.

Elshtain:2003:ALM

[2978] Jean Bethke Elshtain. Against liberal monism. Dædalus, 132(3):78–79, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027864.

Tosteson:2003:UBR

[2979] Daniel C. Tosteson. Unhealthy beliefs:

Religion & the plague of war. Dædalus, 132(3):80–82, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027865.

Hitchens:2003:FI

[2980] Christopher Hitchens. The future of an illusion. Dædalus, 132(3):83–87, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027866.

Bilgrami:2003:CWC

[2981] Akeel Bilgrami. The clash within civilizations. Dædalus, 132(3):88–93, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027867.

Brock-Broido:2003:OTD

[2982] Lucie Brock-Broido. The one thousand days. Dædalus, 132(3):95–96, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027868.

Abbott:2003:MRP

[2983] Lee K. Abbott. Men of rough persuasion. Dædalus, 132(3):97–103, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027869.

Appiah:2003:PRS

[2984] Kwame Anthony Appiah and Robert S. Boynton. On philosophy, race, sex &c. Dædalus, 132(3):104–110, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027870.

Glazer:2003:AI

[2985] Nathan Glazer. On Americans & inequality. Dædalus, 132(3):111–115, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192

                                        (electronic). URL http://www.jstor.
                                        org/stable/20027871.

Traynor:2003:EL

[2986] Michael Traynor. On environmental law. Dædalus, 132(3):116–119, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027872.

Anonymous:2003:BMc

[2987] Anonymous. Back matter. Dædalus, 132(3):??, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027873.

Anonymous:2003:FMc

[2988] Anonymous. Front matter. Dædalus, 132(3):??, Summer 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027852.

Lightman:2003:SM

[2989] Alan Lightman. A sense of the mysterious. Dædalus, 132(4):5–21, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027876;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/001152603771338733.

Einstein:2003:PR

[2990] Albert Einstein. Physics & reality. Dædalus, 132(4):22–25, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027877;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152603771338742.

Holton:2003:ETP

[2991] Gerald Holton. Einstein’s third paradise. Dædalus, 132(4):26–34, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.

aip.org/history/einstein/essay-
Einsteins-Third-Paradise.pdf;
http://www.jstor.org/stable/20027878;
 http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152603771338751.

Pesic:2003:BBM

[2992] Peter Pesic. The bell & the buzzer:

On the meaning of science. Dædalus, 132(4):35–44, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027879;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152603771338760.

Pingree:2003:LNW

[2993] David Pingree. The logic of nonwestern science: Mathematical discoveries in medieval India. Dædalus, 132(4):45–53, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027880;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/001152603771338779.

Haack:2003:TTS

[2994] Susan Haack. Trials & tribulations: Science in the courts. Dædalus, 132(4):54–63, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027881;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/001152603771338788.

Jewett:2003:SPD

[2995] Andrew Jewett. Science & the promise of democracy in America. Dædalus, 132(4):64–70, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027882;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/001152603771338797.

Murray:2003:TYM

[2996] Les Murray. The tune on your mind and photographing aspiration. Dædalus, 132(4):71–72, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027883;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/001152603771338805.

Murray:2003:PA

[2997] Les Murray. Photographing aspiration. Dædalus, 132(4):72, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027884.

Scott:2003:P

[2998] Joanna Scott. That place. Dædalus, 132(4):73–83, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027885;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152603771338814.

Loftus:2003:SUL

[2999] Elizabeth F. Loftus. On science under legal assault. Dædalus, 132 (4):84–86, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027886;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152603771338823.

Zagorin:2003:HPP

[3000] Perez Zagorin. On humanism past & present. Dædalus, 132 (4):87–92, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027887;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152603771338832.

Anonymous:2003:BMd

[3001] Anonymous. Back matter. Dædalus, 132(4):??, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027888.

Anonymous:2003:FMd

[3002] Anonymous. Front matter. Dædalus, 132(4):??, Fall 2003. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027875.

Gardner:2004:WWD

[3003] Howard Gardner. What we do & don’t know about learning. Dædalus, 133(1):5–12, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027891;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604772746648.

Bruner:2004:SHP

[3004] Jerome Bruner. A short history of psychological theories of learning. Dædalus, 133(1):13–20, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027892;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/001152604772746657.

Gopnik:2004:FOI

[3005] Alison Gopnik. Finding our inner scientist. Dædalus, 133(1): 21–28, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027893;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604772746666.

Povinelli:2004:BAA

[3006] Daniel John Povinelli. Behind the Ape’s appearance: Escaping anthropocentrism in the study of other minds. Dædalus, 133(1): 29–41, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027894;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604772746675.

Churchland:2004:HDN

[3007] Patricia Smith Churchland. How do neurons know? Dædalus, 133 (1):42–50, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027895;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604772746684.

Tomasello:2004:LT

[3008] Michael Tomasello. Learning through others. Dædalus, 133(1):51–58, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027896;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/001152604772746693.

Carey:2004:BOC

[3009] Susan Carey. Bootstrapping & the origin of concepts. Dædalus, 133 (1):59–68, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027897;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604772746701.

Glymour:2004:AD

[3010] Clark Glymour. The automation of discovery. Dædalus, 133 (1):69–77, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027898;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604772746710.

Ferry:2004:O

[3011] David Ferry. October. Dædalus, 133(1):78, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027899;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604772746729.

Wachtel:2004:A

[3012] Chuck Wachtel. The annunciation. Dædalus, 133(1):79–90, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027900;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/001152604772746738.

Hochschild:2004:SSW

[3013] Jennifer Hochschild. On the social science wars. Dædalus, 133 (1):91–94, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027901;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604772746747.

Early:2004:LC

[3014] Gerald Early. On literature & childhood. Dædalus, 133(1):95–98, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027902;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/001152604772746756.

Stern:2004:WE

[3015] Richard Stern. On a Writer’s endgame. Dædalus, 133(1):98–102, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027903;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/001152604772746765.

Edwards:2004:CFA

[3016] A. W. F. Edwards. Cambridge’s first African Ph.D.? Dædalus, 133(1):103, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027904.

Edwards:2004:LED

[3017] Professor A. W. F. Edwards. Letter to the Editor of Dædalus. Dædalus, 133(1):103, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/001152604772746774.

Anonymous:2004:BMa

[3018] Anonymous. Back matter. Dædalus, 133(1):??, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027905.

Anonymous:2004:FMa

[3019] Anonymous. Front matter. Dædalus, 133(1):??, Winter 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027890.

McMahon:2004:HVV

[3020] Darrin M. McMahon. From the happiness of virtue to the virtue of happiness: 400 B.C.–A.D. 1780. Dædalus, 133(2):5–17, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027908;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604323049343.

Biswas-Diener:2004:PSW

[3021] Robert Biswas-Diener, Ed Diener, and Maya Tamir. The psychology of subjective well-being. Dædalus, 133(2):18–25, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027909;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604323049352.

Easterlin:2004:EH

[3022] Richard A. Easterlin. The economics of happiness. Dædalus, 133 (2):26–33, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027910;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604323049361.

Wierzbicka:2004:HCL

[3023] Anna Wierzbicka. ’happiness’ in crosslinguistic & cross-cultural perspective. Dædalus, 133(2):34–43, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027911;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/001152604323049370.

Annas:2004:HA

[3024] Julia Annas. Happiness as achievement. Dædalus, 133(2):44–51, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027912;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/001152604323049389.

Reginster:2004:HFB

[3025] Bernard Reginster. Happiness as a Faustian bargain. Dædalus, 133 (2):52–59, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027913;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604323049398.

Nussbaum:2004:MBA

[3026] Martha C. Nussbaum. Mill between Aristotle & Bentham. Dædalus, 133 (2):60–68, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027914;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604323049406.

Frank:2004:HBH

[3027] Robert H. Frank. How not to buy happiness. Dædalus, 133 (2):69–79, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027915;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604323049415.

Seligman:2004:CHT

[3028] Martin E. P. Seligman. Can happiness be taught? Dædalus, 133 (2):80–87, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027916;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152604323049424.

Howard:2004:PMD

[3029] Richard Howard. [on a photograph by Mike Disfarmer]. Dædalus, 133 (2):88–93, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027917.

McGraw:2004:AS

[3030] Erin McGraw. Appearance of scandal. Dædalus, 133(2):94–104, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027918;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/001152604323049442.

Philander:2004:NUS

[3031] S. George H. Philander. On El Niño & the uncertain science of global warming. Dædalus, 133(2): 105–108, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027919;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/001152604323049451.

Hutcheon:2004:AA

[3032] Linda Hutcheon. On the art of adaptation. Dædalus, 133(2): 108–111, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027920;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/001152604323049460.

Anonymous:2004:BMb

[3033] Anonymous. Back matter. Dædalus, 133(2):??, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027921.

Anonymous:2004:FMb

[3034] Anonymous. Front matter. Dædalus, 133(2):??, Spring 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027907.

Levinson:2004:TIR

[3035] Sanford Levinson. Torture in Iraq & the rule of law in America. Dædalus, 133(3):5–9, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027924;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526041504579.

Gray:2004:IF

[3036] John Gray. An illusion with a future. Dædalus, 133(3):10–17, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027925;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/0011526041504542.

Stiglitz:2004:EEC

[3037] Joseph E. Stiglitz. Evaluating economic change. Dædalus, 133(3): 18–25, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027926;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526041504551.

Shweder:2004:GWB

[3038] Richard A. Shweder. George W. Bush & the missionary position. Dædalus, 133(3):26–36, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027927;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526041504515.

Fukuda-Parr:2004:CFH

[3039] Sakiko Fukuda-Parr. Cultural freedom & human development today. Dædalus, 133(3):37–45, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027928;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526041504470.

Larmore:2004:HT

[3040] Charles Larmore. History & truth. Dædalus, 133(3):46–55, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20027929;
 http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526041504498.

Baker:2004:CS

[3041] Keith Michael Baker. On Condorcet’s “Sketch”. Dædalus, 133 (3):56–64, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027930;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526041504506.

deCondorcet:2004:SHP

[3042] Nicolas de Condorcet. Sketch for a historical picture of the progress of the human mind: Tenth epoch. Dædalus, 133(3):65–82, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027931;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526041504588.

Translation of Esquisse d’un tableau historique des progrès de l’esprit humain to English by Keith Michael Baker.

Steinbruner:2004:CTA

[3043] John Steinbruner and Nancy Gallagher. Constructive transformation: an alternative vision of global security. Dædalus, 133(3): 83–103, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027932;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526041504533.

Wilbur:2004:G

[3044] Richard Wilbur. Green. Dædalus, 133(3):104, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027933;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526041504524.

Robinson:2004:BM

[3045] Roxana Robinson. Blind man.

Dædalus, 133(3):105–115, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027934;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/0011526041504489.

Boss:2004:SEP

[3046] Alan P. Boss. On the search for extrasolar planets. Dædalus, 133(3): 116–119, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027935;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526041504597.

Kremer:2004:HIW

[3047] Michael Kremer. On how to improve world health. Dædalus, 133 (3):120–123, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027936;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/0011526041504560.

Anonymous:2004:BMc

[3048] Anonymous. Back matter. Dædalus, 133(3):??, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027937.

Anonymous:2004:FMc

[3049] Anonymous. Front matter. Dædalus, 133(3):??, Summer 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027923.

Pinker:2004:WNN

[3050] Steven Pinker. Why nature & nurture won’t go away. Dædalus, 133(4):5–17, Fall 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027940;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/0011526042365591.

Rorty:2004:PE

[3051] Richard Rorty. Philosophy-envy. Dædalus, 133(4):18–24, Fall 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027941;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/0011526042365537.

Wrangham:2004:KS

[3052] Richard Wrangham. Killer species. Dædalus, 133(4):25–35, Fall 2004. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20027942;
 http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526042365573.

Bateson:2004:OHD

[3053] Patrick Bateson. The origins of human differences. Dædalus, 133 (4):36–46, Fall 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027943;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526042365636.

Brown:2004:HUH

[3054] Donald E. Brown. Human universals, human nature & human culture. Dædalus, 133(4):47–54, Fall 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027944;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526042365645.

Haidt:2004:IEH

[3055] Jonathan Haidt and Craig Joseph.

Intuitive ethics: How innately prepared intuitions generate culturally variable virtues. Dædalus, 133 (4):55–66, Fall 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027945;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526042365555.

Smith:2004:HNE

[3056] Vernon L. Smith. Human nature:

an economic perspective. Dædalus, 133(4):67–76, Fall 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027946;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/0011526042365528.

Kagan:2004:UHH

[3057] Jerome Kagan. The uniquely human in human nature. Dædalus, 133(4):77–88, Fall 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027947;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/0011526042365609.

Ridley:2004:DBH

[3058] Matt Ridley. The DNA behind human nature: Gene expression and the role of experience. Dædalus, 133(4):89–98, Fall 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027948;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/0011526042365546.

Hill:2004:IHC

[3059] Geoffrey Hill. Improvisations for Hart Crane. Dædalus, 133(4):99–101, Fall 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027949;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/0011526042365564.

LaValle:2004:LMH

[3060] Victor La Valle. I left my heart in Skaftafell. Dædalus, 133(4): 102–112, Fall 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027950;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526042365582.

Altieri:2004:DCA

[3061] Charles Altieri. On difficulty in contemporary American poetry. Dædalus, 133(4):113–118, Fall 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027951;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526042365618.

Margulis:2004:SNM

[3062] Lynn Margulis. On syphilis & Nietzsche’s madness: Spirochetes awake! Dædalus, 133(4):118–125, Fall 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027952;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/0011526042365627.

Anonymous:2004:BMd

[3063] Anonymous. Back matter. Dædalus, 133(4):??, Fall 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027953.

Anonymous:2004:FMd

[3064] Anonymous. Front matter. Dædalus, 133(4):??, Fall 2004. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027939.

Prewitt:2005:RCA

[3065] Kenneth Prewitt. Racial classification in America: Where do we go from here? Dædalus, 134 (1):5–17, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027956;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053124370.

Hollinger:2005:ODR

[3066] David A. Hollinger. The one drop rule & the one hate rule. Dædalus, 134(1):18–28, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027957;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053124424.

Pollak:2005:RLH

[3067] Louis H. Pollak. Race, law & history: The Supreme Court from “Dred Scott” to “Grutter v. Bollinger”. Dædalus, 134(1):29–41, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027958;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053124488.

Lopez:2005:RCH

[3068] Ian Haney López. Race on the 2010 census: Hispanics & the shrinking white majority. Dædalus, 134 (1):42–52, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027959;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053124479.

Williams:2005:MCR

[3069] Kim M. Williams. Multiracialism & the civil rights future. Dædalus, 134(1):53–60, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027960;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053124406.

Hattam:2005:EBR

[3070] Victoria Hattam. Ethnicity & the boundaries of race: Rereading directive 15. Dædalus, 134 (1):61–69, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027961;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053124352.

Hochschild:2005:LAR

[3071] Jennifer L. Hochschild. Looking ahead: Racial trends in the United States. Dædalus, 134(1): 70–81, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027962;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053124343.

Nobles:2005:MLA

[3072] Melissa Nobles. The myth of Latin American multiracialism. Dædalus, 134(1):82–87, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027963;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053124398.

Fredrickson:2005:DRF

[3073] George M. Fredrickson. Diverse republics: French & American responses to racial pluralism. Dædalus, 134 (1):88–101, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027964;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/0011526053124451.

Hacking:2005:WRS

[3074] Ian Hacking. Why race still matters. Dædalus, 134(1):102–116, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027965;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/0011526053124460.

Hadas:2005:I

[3075] Rachel Hadas. Inspissation. Dædalus, 134(1):117, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027966;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/0011526053124415.

Nunez:2005:NJ

[3076] Sigrid Nunez. The naked juror. Dædalus, 134(1):118–125, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027967;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/0011526053124361.

Foucault:2005:MIC

[3077] Michel Foucault and Baqir Parham. On Marx, Islam, Christianity & revolution. Dædalus, 134(1):126–132, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20027968;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/0011526053124433.

Taylor:2005:HI

[3078] Shelley E. Taylor. On healthy illusions. Dædalus, 134(1):133–135, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027969;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/0011526053124442.

Quinn:2005:RDT

[3079] Philip L. Quinn. On religious diversity & tolerance. Dædalus, 134 (1):136–139, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027970;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053124389.

Anonymous:2005:BMa

[3080] Anonymous. Back matter. Dædalus, 134(1):??, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027971.

Anonymous:2005:FMa

[3081] Anonymous. Front matter. Dædalus, 134(1):??, Winter 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027955.

Cole:2005:AFU

[3082] Jonathan R. Cole. Academic freedom under fire. Dædalus, 134 (2):5–17, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027974;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053887310.

Ferguson:2005:UCL

[3083] Niall Ferguson. The unconscious Colossus: Limits of (& alternatives to) American empire. Dædalus, 134 (2):18–33, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027975;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053887419.

Pomeranz:2005:ECM

[3084] Kenneth Pomeranz. Empire & ‘civilizing’ missions, past & present. Dædalus, 134(2):34–45, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027976;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053887428.

Pagden:2005:ILQ

[3085] Anthony Pagden. Imperialism, liberalism & the quest for perpetual peace. Dædalus, 134(2): 46–57, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027977;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053887301.

Snyder:2005:EBT

[3086] Jack Snyder. Empire: a blunt tool for democratization. Dædalus, 134 (2):58–71, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027978;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053887356.

Blackburn:2005:EEC

[3087] Robin Blackburn. Emancipation & empire, from Cromwell to Karl Rove. Dædalus, 134(2):72–87, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027979;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/0011526053887383.

Greene:2005:OE

[3088] Molly Greene. The Ottoman experience. Dædalus, 134(2):88–99, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027980;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/0011526053887293.

Wesseling:2005:IRG

[3089] Henk Wesseling. Imperialism & the roots of the great war. Dædalus, 134 (2):100–107, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027981;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053887338.

Iriye:2005:BIN

[3090] Akira Iriye. Beyond imperialism: The new internationalism. Dædalus, 134 (2):108–116, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027982;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053887392.

Wright:2005:L

[3091] Franz Wright. Lesson. Dædalus, 134(2):117, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027983.

Wright:2005:LC

[3092] Franz Wright. Lesson and the choice. Dædalus, 134(2):117–118, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/0011526053887347.

Wright:2005:C

[3093] Franz Wright. The choice. Dædalus, 134(2):118, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027984.

Atwood:2005:FSP

[3094] Margaret Atwood. Four short pieces: King Log in exile, post-colonial, Salome was a dancer & take charge. Dædalus, 134(2):119–123, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027985;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/0011526053887329.

Mazur:2005:MIB

[3095] Barry Mazur and Peter Pesic. On mathematics, imagination & the beauty of numbers. Dædalus, 134 (2):124–130, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20027986;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/0011526053887365.

Hechter:2005:PE

[3096] Michael Hechter. On the 2004 Presidential election. Dædalus, 134(2): 131–133, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027987;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053887374.

Fine:2005:ECM

[3097] Morris E. Fine and Peter W. Voorhees.

On the evolving curriculum in materials science & engineering. Dædalus, 134(2):134–136, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027988;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526053887400.

Anonymous:2005:BMb

[3098] Anonymous. Back matter. Dædalus, 134(2):??, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027989.

Anonymous:2005:FMb

[3099] Anonymous. Front matter. Dædalus, 134(2):??, Spring 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027973.

Greenberg:2005:NPS

[3100] David Greenberg. The new politics of Supreme Court appointments. Dædalus, 134(3):5–12, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/

stable/20027992;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/0011526054622042.

Gardner:2005:PAT

[3101] Howard Gardner and Lee S. Shulman. The professions in America today: Crucial but fragile. Dædalus, 134(3):13–18, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027993;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526054622132.

Sullivan:2005:MVP

[3102] William M. Sullivan. Markets vs. professions: Value added? Dædalus, 134(3):19–26, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027994;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526054622150.

Damon:2005:PMB

[3103] William Damon, Anne Colby, Kendall Bronk, and Thomas Ehrlich. Passion & mastery in balance: Toward good work in the professions. Dædalus, 134 (3):27–35, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027995;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526054622105.

Goldman:2005:GWH

[3104] Harvey Goldman. Good work, from Homer to the present. Dædalus, 134 (3):36–41, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027996;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526054622033.

Gardner:2005:CW

[3105] Howard Gardner. Compromised work. Dædalus, 134(3):42–51, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20027997;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/0011526054622097.

Shulman:2005:SPP

[3106] Lee S. Shulman. Signature pedagogies in the professions. Dædalus, 134 (3):52–59, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027998;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526054622015.

Nakamura:2005:EPC

[3107] Jeanne Nakamura and Mihaly Csikszentmihalyi. Engagement in a profession: The case of undergraduate teaching. Dædalus, 134(3): 60–67, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027999;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526054622051.

Harpham:2005:PPP

[3108] Geoffrey Galt Harpham. Politics, professionalism & the pleasure of reading. Dædalus, 134 (3):68–75, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028000;

                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/0011526054622088.

Merwin:2005:PT

[3109] W. S. Merwin. To the present tense. Dædalus, 134(3):76, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20028001;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/0011526054622141.

Coover:2005:GN

[3110] Robert Coover. Grandmother’s nose. Dædalus, 134(3):77–81, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20028002;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/0011526054622024.

Kolakowski:2005:EPT

[3111] Leszek Kolakowski and Danny Postel. On exile, philosophy & tottering insecurely on the edge of an unknown abyss. Dædalus, 134 (3):82–88, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028003;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/0011526054622079.

Firmat:2005:BD

[3112] Gustavo Pérez Firmat. On bilingualism & its discontents. Dædalus, 134 (3):89–92, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028004;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/0011526054622114.

Stern:2005:HSC

[3113] Richard Stern. On how the stories changed. Dædalus, 134(3): 93–95, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028005;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526054622123.

Green:2005:EBP

[3114] Donald P. Green. On evidencebased political science. Dædalus, 134 (3):96–100, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028006;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/0011526054622060.

Anonymous:2005:BMc

[3115] Anonymous. Back matter. Dædalus, 134(3):??, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028007.

Anonymous:2005:FMc

[3116] Anonymous. Front matter. Dædalus, 134(3):??, Summer 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20027991.

Anonymous:2005:P

[3117] Anonymous.  Preface.  Dædalus,

134(4):5, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/001152605774431437.

Baltimore:2005:LSB

[3118] David Baltimore. Limiting science: a Biologist’s perspective. Dædalus, 134(4):7–15, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028010;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152605774431446.

Daston:2005:FLI

[3119] Lorraine Daston. Fear & loathing of the imagination in science. Dædalus, 134(4):16–30, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028011;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152605774431473.

Weinberg:2005:PH

[3120] Steven Weinberg. Physics & history. Dædalus, 134(4):31–39, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028012;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/001152605774431482.

Bellah:2005:CRA

[3121] Robert N. Bellah. Civil religion in America. Dædalus, 134 (4):40–55, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028013;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152605774431464.

Geertz:2005:DPN

[3122] Clifford Geertz. Deep play: Notes on the Balinese cockfight. Dædalus, 134(4):56–86, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028014;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152605774431563.

Solow:2005:HDE

[3123] Robert M. Solow. How did economics get that way & what way did it get? Dædalus, 134 (4):87–100, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028015;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152605774431518.

Schelling:2005:RMA

[3124] Thomas C. Schelling. Reciprocal measures for arms stabilization. Dædalus, 134(4):101–117, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028016;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152605774431455.

Franklin:2005:TWR

[3125] John Hope Franklin. The two worlds of race: a historical view. Dædalus, 134(4):118–133, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028017;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152605774431554.

Conway:2005:PPG

[3126] Jill K. Conway. Politics, pedagogy & gender. Dædalus, 134 (4):134–144, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028018;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/001152605774431509.

Wei-ming:2005:CCP

[3127] Tu Wei-ming. Cultural China: The periphery as the center. Dædalus, 134(4):145–167, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028019;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/001152605774431545.

Sen:2005:ITW

[3128] Amartya Sen. Indian traditions & the Western imagination. Dædalus, 134(4):168–185, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028020;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/001152605774431428.

Frost:2005:E

[3129] Robert Frost. On Emerson. Dædalus, 134(4):186–190, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028021;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/001152605774431536.

Eco:2005:IRB

[3130] Umberto Eco. Innovation & repetition: Between modern & postmodern aesthetics. Dædalus, 134 (4):191–207, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028022;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152605774431527.

Sontag:2005:AAB

[3131] Susan Sontag. An argument about beauty. Dædalus, 134(4):208–213, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028023;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/001152605774431491.

Anonymous:2005:BMd

[3132] Anonymous. Back matter. Dædalus, 134(4):??, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028024.

Anonymous:2005:FMd

[3133] Anonymous. Front matter. Dædalus, 134(4):??, Fall 2005. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028009.

Wilson:2006:CA

[3134] Chris Wilson. The century ahead.

Dædalus, 135(1):5–8, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20013450;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152606775320979.

Aaron:2006:LLS

[3135] Henry J. Aaron. Longer life spans:

Boon or burden? Dædalus, 135 (1):9–19, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013451;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152606775321022.

Harper:2006:MSP

[3136] Sarah Harper. Mature societies: Planning for our future selves. Dædalus, 135(1):20–31, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013452;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152606775321040.

Baltes:2006:FOL

[3137] Paul B. Baltes. Facing our limits: Human dignity in the very old. Dædalus, 135(1):32–39, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013453;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152606775321086.

Partridge:2006:WMM

[3138] Linda Partridge. Of worms, mice & men: Altering rates of aging. Dædalus, 135(1):40–47, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013454;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152606775321013.

Kaplan:2006:LCS

[3139] Hillard Kaplan. The life course of a skill-intensive foraging species. Dædalus, 135(1):48–57, Winter 2006. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20013455;
 http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152606775320960.

Selkoe:2006:AMD

[3140] Dennis J. Selkoe. The aging mind:

Deciphering Alzheimer’s disease & its antecedents. Dædalus, 135 (1):58–67, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013456;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152606775321112.

Finch:2006:AIB

[3141] Caleb E. Finch. Aging, inflammation & the body electric. Dædalus, 135(1):68–76, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013457;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152606775321077.

Clark:2006:AGO

[3142] Kenneth Clark. The artist grows old. Dædalus, 135(1):77–90, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013458;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/001152606775321068.

Gokhale:2006:MSS

[3143] Jagadeesh Gokhale and Kent Smetters. Measuring social Security’s financial outlook within an aging society. Dædalus, 135(1): 91–104, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013459;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152606775320988.

Berkman:2006:HSS

[3144] Lisa F. Berkman and M. Maria Glymour. How society shapes aging: The centrality of variability. Dædalus, 135 (1):105–114, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20013460;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/001152606775321103.

Wright:2006:LS

[3145] Charles Wright. Last supper. Dædalus, 135(1):115, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20013461;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/001152606775321004.

Davis:2006:KBY

[3146] Ree Davis. I kneel before you. Dædalus, 135(1):116–119, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20013462;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/001152606775321031.

Bell:2006:SSP

[3147] Daniel Bell, Wolf Lepenies, and Howard Eiland. On society & sociology past & present. Dædalus, 135 (1):120–123, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20013463;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/001152606775321121.

Laber:2006:T

[3148] Jeri Laber. On torture. Dædalus, 135 (1):124–126, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013464;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152606775320997.

Nagel:2006:DF

[3149] Robert F. Nagel. On the decline of federalism. Dædalus, 135 (1):127–130, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20013465;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/001152606775321095.

Reagor:2006:CW

[3150] Simone Reagor, Howard Gardner, and Paula Marshall. On compromised work. Dædalus, 135(1):131–133, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013466.

Anonymous:2006:LED

[3151] Anonymous. Letters to the Editors of Dædalus. Dædalus, 135 (1):131–134, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/001152606775321059.

Bracewell:2006:PN

[3152] Ronald N. Bracewell. Poetry for nonpoets. Dædalus, 135(1):133–134, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20013467.

Anonymous:2006:BMa

[3153] Anonymous. Back matter. Dædalus, 135(1):??, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20013468.

Anonymous:2006:FMa

[3154] Anonymous. Front matter. Dædalus, 135(1):??, Winter 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20013449.

Bromwich:2006:RD

[3155] David Bromwich. A Republic divided. Dædalus, 135(2):5–10, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20028027;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/daed.2006.
                                        135.2.5.

Spacks:2006:RH

[3156] Patricia Meyer Spacks. Revolution in the humanities. Dædalus, 135 (2):11–14, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028028;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2006.135.
                                        2.11.

Marcus:2006:HCC

[3157] Steven Marcus. Humanities from classics to cultural studies: Notes toward the history of an idea. Dædalus, 135(2):15–21, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028029;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
2.15.

Delbanco:2006:ALV

[3158] Andrew Delbanco. American literature: a vanishing subject? Dædalus, 135(2):22–37, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028030;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
2.22.

Yu:2006:CLQ

[3159] Pauline Yu. Comparative literature in question. Dædalus, 135 (2):38–53, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028031;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
2.38.

Grafton:2006:HPF

[3160] Anthony Grafton. History’s postmodern fates. Dædalus, 135(2): 54–69, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028032;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
2.54.

Crow:2006:PAH

[3161] Thomas Crow. The practice of art history in America. Dædalus, 135 (2):70–90, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028033;

                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2006.135.
                                        2.70.

Early:2006:QBH

[3162] Gerald Early. The quest for a black humanism. Dædalus, 135 (2):91–104, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028034;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2006.135.
                                        2.91.

Balkin:2006:LHU

[3163] Jack M. Balkin and Sanford Levinson. Law & the humanities: an uneasy relationship. Dædalus, 135 (2):105–115, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028035;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2006.135.
                                        2.105.

Follesdal:2006:APT

[3164] Dagfinn Føllesdal and Michael Friedman. American philosophy in the Twentieth Century. Dædalus, 135 (2):116–126, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028036;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2006.135.
                                        2.116.

Boyers:2006:UDM

[3165] Peg Boyers. Unsent dedication for Martin Heidegger. Dædalus, 135 (2):127–129, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028037;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
2.127.

Braver:2006:DF

[3166] Adam Braver. A death in the family. Dædalus, 135(2):130–132, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028038;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2006.
135.2.130.

Quandt:2006:PPM

[3167] William B. Quandt. On the peace process in the Middle East. Dædalus, 135(2):133–135, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028039;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
2.133.

Schwarz:2006:JTB

[3168] Norbert Schwarz. On judgments of truth & beauty. Dædalus, 135 (2):136–138, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028040;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
2.136.

Waksman:2006:SP

[3169] Byron H. Waksman. On scientists as professionals. Dædalus, 135 (2):139–140, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028041;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2006.135.
                                        2.139.

Brown:2006:SSA

[3170] Edgar H. Brown, Jr., Jagadeesh Gokhale, and Kent Smetters. On social security & the aging of America. Dædalus, 135(2):140–142, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20028042;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/daed.2006.
                                        135.2.140.

Anonymous:2006:BMb

[3171] Anonymous. Back matter. Dædalus, 135(2):??, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028043.

Anonymous:2006:FMb

[3172] Anonymous. Front matter. Dædalus, 135(2):??, Spring 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028026.

Brown:2006:AUE

[3173] Marilyn A. Brown, Benjamin K. Sovacool, and Richard F. Hirsh. Assessing U.S. energy policy. Dædalus, 135(3):5–11, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028046;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2006.135.
                                        3.5.

Graham:2006:I

[3174] Jorie Graham. Incarnation: 9:30 am to 9:36 am. Dædalus, 135 (3):12–14, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028047;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
3.12.

Damasio:2006:MB

[3175] Antonio Damasio and Hanna Damasio. Minding the body. Dædalus, 135(3):15–22, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028048;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
3.15.

Edelman:2006:EM

[3176] Gerald M. Edelman. The embodiment of mind. Dædalus, 135(3): 23–32, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028049;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
3.23.

Ohman:2006:MSE

[3177] Arne Öhman. Making sense of emotion: Evolution, reason & the brain. Dædalus, 135(3):33–45, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028050;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2006.
135.3.33.

Johnson:2006:MID

[3178] Mark Johnson. Mind incarnate: From Dewey to Damasio. Dædalus, 135 (3):46–54, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028051;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2006.135.
                                        3.46.

Gilligan:2006:WML

[3179] Carol Gilligan. When the mind leaves the body … and returns. Dædalus, 135(3):55–66, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028052;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2006.135.
                                        3.55.

Connolly:2006:EE

[3180] William E. Connolly. Experience & experiment. Dædalus, 135(3): 67–75, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028053;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2006.135.
                                        3.67.

dAmboise:2006:MD

[3181] Jacques d’Amboise. The mind in dance. Dædalus, 135(3):76–77, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20028054;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/daed.2006.
                                        135.3.76.

Dolan:2006:BB

[3182] Ray Dolan. The body in the brain. Dædalus, 135(3):78–85, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028055;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2006.
135.3.78.

Fodor:2006:HMW

[3183] Jerry Fodor. How the mind works:

What we still don’t know. Dædalus, 135(3):86–94, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028056;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
3.86.

Gossy:2006:DW

[3184] Dorian Gossy. The door in the Woods. Dædalus, 135(3):95–103, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028057;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2006.
135.3.95.

Rosen:2006:LSC

[3185] Stanley Rosen. Leo Strauss in Chicago. Dædalus, 135(3):104– 113, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028058;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
3.104.

Handler:2006:WRH

[3186] Joel F. Handler. On welfare reform’s hollow victory. Dædalus, 135 (3):114–117, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028059;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
3.114.

Galston:2006:RPP

[3187] William A. Galston. On the reemergence of political pluralism. Dædalus, 135(3):118–122, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028060;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2006.
135.3.118.

Keddie:2006:HTC

[3188] Nikki R. Keddie and Anthony Grafton. On history in the Twentieth Century. Dædalus, 135(3):123–125, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028061;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2006.
135.3.123.

Golomb:2006:PME

[3189] Solomon W. Golomb and William B. Quandt. On peace in the Middle East. Dædalus, 135(3):126, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028062;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2006.
135.3.126a.

Dummett:2006:TEU

[3190] Michael Dummett, Dagfinn Føllesdal, and Michael Friedman. On teaching in European universities. Dædalus, 135 (3):126–127, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028063;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
3.126b.

Anonymous:2006:BMc

[3191] Anonymous. Back matter. Dædalus, 135(3):??, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028064.

Anonymous:2006:FMc

[3192] Anonymous. Front matter. Dædalus, 135(3):??, Summer 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028045.

Bilgrami:2006:NTD

[3193] Akeel Bilgrami. Notes toward the definition of ‘identity’. Dædalus, 135(4):5–14, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028067;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
4.5.

Appiah:2006:PI

[3194] Kwame Anthony Appiah. The politics of identity. Dædalus, 135 (4):15–22, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028068;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
4.15.

Hollinger:2006:IS

[3195] David A. Hollinger. From identity to solidarity. Dædalus, 135 (4):23–31, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028069;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
4.23.

Jung:2006:WLS

[3196] Courtney Jung. Why liberals should value ‘identity politics’. Dædalus, 135(4):32–39, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028070;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
4.32.

Shoemaker:2006:II

[3197] Sydney Shoemaker. Identity & identities. Dædalus, 135(4):40–48, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028071;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2006.
135.4.40.

Rovane:2006:WDI

[3198] Carol Rovane. Why do individuals matter? Dædalus, 135 (4):49–59, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028072;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
4.49.

Doniger:2006:MMM

[3199] Wendy Doniger. Many masks, many selves. Dædalus, 135(4):60–71, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028073;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2006.
135.4.60.

Feinberg:2006:OBO

[3200] Todd E. Feinberg. Our brains, our selves. Dædalus, 135(4):72–80, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028074;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2006.
135.4.72.

Hacking:2006:GBG

[3201] Ian Hacking. Genetics, biosocial groups & the future of identity. Dædalus, 135(4):81–95, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028075;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
4.81.

Altschuler:2006:CPD

[3202] Glenn C. Altschuler. The convictions of Peter Debye. Dædalus, 135(4):96–103, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028076;

                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2006.135.
                                        4.96.

diCamillo:2006:WAT

[3203] Kevin Carrizo di Camillo. Were all there for Geoffrey Hill. Dædalus, 135(4):104, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028077;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2006.135.
                                        4.104.

Dorfman:2006:S

[3204] Ariel Dorfman. Stew. Dædalus, 135(4):105–107, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028078;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2006.135.
                                        4.105.

Cook:2006:ICI

[3205] Michael Cook. On Islam & comparative intellectual history. Dædalus, 135(4):108–111, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028079;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2006.135.
                                        4.108.

Nocera:2006:FGE

[3206] Daniel G. Nocera. On the future of global energy. Dædalus, 135 (4):112–115, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028080;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2006.135.
4.112.

Anonymous:2006:BMd

[3207] Anonymous. Back matter. Dædalus, 135(4):??, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028081.

Anonymous:2006:FMd

[3208] Anonymous. Front matter. Dædalus, 135(4):??, Fall 2006. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028066.

McNeill:2007:VSH

[3209] William H. McNeill. Violence & submission in the human past. Dædalus, 136(1):5–12, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028084;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
1.5.

LeBlanc:2007:WWL

[3210] Steven A. LeBlanc. Why warfare?

Lessons from the past. Dædalus, 136(1):13–21, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028085;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
1.13.

Schell:2007:PPV

[3211] Jonathan Schell and Robert S. Boynton. People’s power vs. nuclear power: a conversation. Dædalus, 136(1):22–29, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028086;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
1.22.

Juergensmeyer:2007:GVT

[3212] Mark Juergensmeyer. Gandhi vs. terrorism. Dædalus, 136(1):30–39, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028087;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2007.
136.1.30.

Whitehead:2007:VCO

[3213] Neil L. Whitehead. Violence & the cultural order. Dædalus, 136 (1):40–50, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028088;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
1.40.

Todorov:2007:AGT

[3214] Tzvetan Todorov and Arthur Goldhammer. Avant-gardes & totalitarianism. Dædalus, 136(1): 51–66, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028089;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
1.51.

Michnik:2007:UMR

[3215] Adam Michnik and Elzbieta Matynia. The ultras of moral revolution. Dædalus, 136(1):67–83, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028090;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2007.
136.1.67.

Ness:2007:RFV

[3216] Cindy D. Ness. The rise in female violence. Dædalus, 136(1): 84–93, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028091;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
1.84.

Bloom:2007:FSB

[3217] Mia Bloom. Female suicide bombers:

a global trend. Dædalus, 136(1): 94–102, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028092;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
1.94.

Mnookin:2007:ECF

[3218] Robert H. Mnookin. Ethnic conflicts: Flemings & Walloons, Palestinians & Israelis. Dædalus, 136 (1):103–119, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028093;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
1.103.

Blight:2007:RMT

[3219] James G. Blight and Janet M. Lang. Robert McNamara: Then & now. Dædalus, 136(1):120–131, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20028094;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/daed.2007.
                                        136.1.120.

Jewitt:2007:PPP

[3220] David Jewitt and Jane X. Luu. Pluto, perception & planetary politics. Dædalus, 136(1):132–136, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20028095;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/daed.2007.
                                        136.1.132.

Gingerich:2007:PPP

[3221] Owen Gingerich. Planetary perils in Prague. Dædalus, 136(1): 137–140, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028096;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2007.136.
                                        1.137.

Kinsella:2007:S

[3222] John Kinsella. Into the Sun. Dædalus, 136(1):141–142, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028097;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2007.136.
                                        1.141.

Shinner:2007:LA

[3223] Peggy Shinner. Life is adequate.

Dædalus, 136(1):143–151, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028098;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2007.
136.1.143.

Rotberg:2007:INS

[3224] Robert I. Rotberg. On improving nation-state governance. Dædalus, 136 (1):152–155, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028099;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
1.152.

Fiske:2007:PB

[3225] Susan T. Fiske. On prejudice & the brain. Dædalus, 136(1): 156–159, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028100;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
1.156.

Acemoglu:2007:EOD

[3226] Daron Acemoglu and James A.

Robinson. On the economic origins of democracy. Dædalus, 136 (1):160–162, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028101;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
1.160.

Anonymous:2007:BMa

[3227] Anonymous. Back matter. Dædalus, 136(1):??, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028102.

Anonymous:2007:FMa

[3228] Anonymous. Front matter. Dædalus, 136(1):??, Winter 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028083.

Ehrlich:2007:GEB

[3229] Paul Ehrlich and Marcus W. Feldman. Genes, environments & behaviors. Dædalus, 136(2):5–12, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20028105;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/daed.2007.
                                        136.2.5.

Birkhead:2007:P

[3230] Tim Birkhead. Promiscuity. Dædalus, 136(2):13–22, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028106;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2007.136.
                                        2.13.

Roughgarden:2007:CDT

[3231] Joan Roughgarden. Challenging Darwin’s theory of sexual selection. Dædalus, 136(2):23–36, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20028107;   http://www.mitpressjournals.

org/doi/pdfplus/10.1162/daed.2007.
136.2.23.

Charlesworth:2007:WBE

[3232] Brian Charlesworth. Why bother? The evolutionary genetics of sex. Dædalus, 136(2):37–46, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028108;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
2.37.

Fausto-Sterling:2007:FD

[3233] Anne Fausto-Sterling. Frameworks of desire. Dædalus, 136(2):47–57, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028109;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2007.
136.2.47.

Benedict:2007:WLA

[3234] Elizabeth Benedict. What I learned about sex on the Internet. Dædalus, 136(2):58–65, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028110;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
2.58.

Doniger:2007:RKS

[3235] Wendy Doniger. Reading the “Kamasutra”: The strange & the familiar. Dædalus, 136(2):66–78, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028111;   http://www.mitpressjournals.

                                        org/doi/pdfplus/10.1162/daed.2007.
                                        136.2.66.

Corngold:2007:KS

[3236] Stanley Corngold. Kafka & sex. Dædalus, 136(2):79–87, Spring 2007. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20028112;
                                         http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2007.136.
                                        2.79.

Castle:2007:LPL

[3237] Terry Castle. The lesbianism of Philip Larkin. Dædalus, 136(2): 88–102, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028113;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2007.136.
                                        2.88.

Cohen:2007:SP

[3238] Lawrence Cohen. Song for Pushkin. Dædalus, 136(2):103–115, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20028114;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/daed.2007.
                                        136.2.103.

Marcus:2007:THH

[3239] Greil Marcus. A trip to Hibbing High School. Dædalus, 136 (2):116–124, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028115;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2007.136.
                                        2.116.

Simic:2007:DC

[3240] Charles Simic. Darkened chessboard. Dædalus, 136(2):125–126, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028116;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2007.
136.2.125.

Simic:2007:SH

[3241] Charles Simic. Secret history. Dædalus, 136(2):126, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028117.

Newland:2007:C

[3242] Peggy Newland. Clowns. Dædalus, 136(2):127–137, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028118;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
2.127.

Weber:2007:BA

[3243] Samuel Weber. On Benjamin’s ‘- abilities’. Dædalus, 136(2):138–141, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028119;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2007.
136.2.138.

Baker:2007:SAT

[3244] William F. Baker. On the state of American television. Dædalus, 136 (2):141–144, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028120;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
2.141.

Anonymous:2007:BMb

[3245] Anonymous. Back matter. Dædalus, 136(2):??, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028121.

Anonymous:2007:FMb

[3246] Anonymous. Front matter. Dædalus, 136(2):??, Spring 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028104.

Dunn:2007:CDE

[3247] John Dunn. Capitalist democracy: Elective affinity or beguiling illusion? Dædalus, 136(3): 5–13, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028124;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
3.5.

Appleby:2007:IUA

[3248] Joyce Appleby. The intellectual underpinnings of American democracy. Dædalus, 136(3):14–23, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028125;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2007.
136.3.14.

Bogle:2007:DCA

[3249] John C. Bogle. Democracy in Corporate America. Dædalus, 136(3): 24–35, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028126;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
3.24.

Blackburn:2007:EDM

[3250] Robin Blackburn. Economic democracy: Meaningful, desirable, feasible? Dædalus, 136(3):36–45, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028127;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2007.
136.3.36.

Friedman:2007:CEG

[3251] Benjamin M. Friedman. Capitalism, economic growth & democracy. Dædalus, 136(3):46–55, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028128;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2007.
136.3.46.

Loewenberg:2007:PL

[3252] Gerhard Loewenberg. Paradoxes of legislatures. Dædalus, 136(3): 56–66, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028129;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
3.56.

Epstein:2007:ZDD

[3253] Richard A. Epstein. Zoning: Deliberative democracy at zero prices. Dædalus, 136(3):67–76, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028130;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2007.136.
                                        3.67.

Muller:2007:DTC

[3254] Jerry Z. Muller. The democratic threat to capitalism. Dædalus, 136 (3):77–86, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028131;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2007.136.
                                        3.77.

Fogel:2007:CD

[3255] Robert W. Fogel. Capitalism & democracy in 2040. Dædalus, 136 (3):87–95, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028132;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2007.136.
                                        3.87.

Dugan:2007:SR

[3256] Lawrence Dugan. Silent Rome. Dædalus, 136(3):96, Summer 2007. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20028133;
                                         http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2007.136.
                                        3.96.

Gordimer:2007:TM

[3257] Nadine Gordimer. Tape measure. Dædalus, 136(3):97–99, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028134;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2007.
136.3.97.

Goldin-Meadow:2007:IL

[3258] Susan Goldin-Meadow. On inventing language. Dædalus, 136(3): 100–103, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028135;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
3.100.

Poole:2007:PPC

[3259] Keith T. Poole and Howard Rosenthal. On party polarization in Congress. Dædalus, 136(3):104– 107, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028136;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
3.104.

Navasky:2007:CLC

[3260] Victor S. Navasky. On civil liberties on campus. Dædalus, 136(3): 108–110, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028137;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
3.108.

Apter:2007:CG

[3261] David E. Apter. On Clifford Geertz. Dædalus, 136(3):111–113, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20028138;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/daed.2007.
                                        136.3.111.

Anonymous:2007:BMc

[3262] Anonymous. Back matter. Dædalus, 136(3):??, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028139.

Anonymous:2007:FMc

[3263] Anonymous. Front matter. Dædalus, 136(3):??, Summer 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028123.

Dionne:2007:WPI

[3264] E. J. Dionne, Jr. Why the public interest matters now. Dædalus, 136(4):5–9, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028142;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2007.136.
                                        4.5.

Galston:2007:ODR

[3265] William A. Galston. An old debate renewed: The politics of the public interest. Dædalus, 136 (4):10–19, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028143;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
4.10.

Wolfson:2007:PIL

[3266] Adam Wolfson. Public interest lost? Dædalus, 136(4):20–29, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028144;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2007.
136.4.20.

Glazer:2007:RPI

[3267] Nathan Glazer. Realizing the public interest: Reflections on an elusive goal. Dædalus, 136(4):30–36, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028145;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2007.
136.4.30.

Bhagwati:2007:EPP

[3268] Jagdish Bhagwati. Economic policy in the public interest. Dædalus, 136(4):37–44, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028146;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
4.37.

Hart:2007:COP

[3269] Gary Hart. The Commonwealth:

Our public legacy. Dædalus, 136 (4):45–50, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028147;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
4.45.

Whitman:2007:GPI

[3270] Christine Todd Whitman. Governing in the public interest: Then & now. Dædalus, 136(4):51–58, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028148;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2007.
136.4.51.

Bellah:2007:EPR

[3271] Robert N. Bellah. Ethical politics: Reality or illusion? Dædalus, 136(4):59–69, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028149;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
4.59.

Gutmann:2007:LDE

[3272] Amy Gutmann. The lure & dangers of extremist rhetoric. Dædalus, 136(4):70–78, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028150;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
4.70.

Taylor:2007:FEP

[3273] Lance Taylor. The flavors of economics & the public interest. Dædalus, 136(4):79–87, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028151;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
4.79.

McQuade:2007:SSW

[3274] Molly McQuade. Spring’s so sad, we want to know why. Dædalus, 136(4):88, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028152;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
4.88.

Gordon:2007:D

[3275] Mary Gordon. Dilly. Dædalus, 136(4):89–100, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028153;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
4.89.

West:2007:MMM

[3276] Cornel West and D. Graham Burnett. Metaphysics, money & the messiah: a conversation about Melville’s “The Confidence-Man”. Dædalus, 136(4):101–114, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028154;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2007.136.
4.101.

Bartov:2007:EGP

[3277] Omer Bartov. On Eastern Galicia’s past & present. Dædalus, 136 (4):115–118, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028155;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2007.136.
                                        4.115.

Ritvo:2007:AT

[3278] Harriet Ritvo. On the animal turn. Dædalus, 136(4):118–122, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20028156;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/daed.2007.
                                        136.4.118.

Anonymous:2007:BMd

[3279] Anonymous. Back matter. Dædalus, 136(4):??, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028157.

Anonymous:2007:FMd

[3280] Anonymous. Front matter. Dædalus, 136(4):??, Fall 2007. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028141.

Daston:2008:LCL

[3281] Lorraine Daston. Life, chance & life chances. Dædalus, 137 (1):5–14, Winter 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028160;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        1.5.

Kenny:2008:BIH

[3282] Anthony Kenny. The beginning of individual human life. Dædalus, 137 (1):15–22, Winter 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028161;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
1.15.

George:2008:EE

[3283] Robert P. George. Embryo ethics.

Dædalus, 137(1):23–35, Winter 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20028162;
 http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
1.23.

Rose:2008:VLS

[3284] Nikolas Rose. The value of life: Somatic ethics & the spirit of biocapital. Dædalus, 137(1):36–48, Winter 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028163;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2008.
137.1.36.

Broome:2008:WYL

[3285] John Broome. What is your life worth? Dædalus, 137(1): 49–56, Winter 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028164;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
1.49.

Lavi:2008:HDB

[3286] Shai Lavi. How dying became a ‘life crisis’. Dædalus, 137(1): 57–65, Winter 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028165;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        1.57.

McMahan:2008:EAN

[3287] Jeff McMahan. Eating animals the nice way. Dædalus, 137(1): 66–76, Winter 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028166;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        1.66.

Woolfson:2008:SL

[3288] Adrian Woolfson. Synthetic life. Dædalus, 137(1):77–83, Winter 2008. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/20028167;
                                         http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        1.77.

Richer:2008:A

[3289] Ted Richer. Anon. Dædalus, 137 (1):84–86, Winter 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028168;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        1.84.

Abani:2008:TLO

[3290] Chris Abani. Three letters, one song & a refrain. Dædalus, 137 (1):87–91, Winter 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028169;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
1.87.

Coley:2008:TGG

[3291] Phyllis D. Coley. On turning green into gold. Dædalus, 137 (1):92–95, Winter 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028170;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
1.92.

Harran:2008:JMR

[3292] Don Harrán. On a Jewish musical Renaissance. Dædalus, 137 (1):96–100, Winter 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028171;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
1.96.

Anonymous:2008:BMa

[3293] Anonymous. Back matter. Dædalus, 137(1):??, Winter 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028172.

Anonymous:2008:FMa

[3294] Anonymous. Front matter. Dædalus, 137(1):??, Winter 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028159.

McKibben:2008:CE

[3295] Bill McKibben. The challenge to environmentalism. Dædalus, 137 (2):5–7, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028175;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
2.5.

Marx:2008:INA

[3296] Leo Marx. The idea of nature in America. Dædalus, 137 (2):8–21, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028176;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
2.8.

Ritvo:2008:BJW

[3297] Harriet Ritvo. Beasts in the jungle (or wherever). Dædalus, 137 (2):22–30, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028177;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
2.22.

Parmesan:2008:WWT

[3298] Camille Parmesan. Where the wild things were. Dædalus, 137 (2):31–38, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028178;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
2.31.

Orians:2008:NHN

[3299] Gordon H. Orians. Nature & human nature. Dædalus, 137(2): 39–48, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028179;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
2.39.

Sunstein:2008:PN

[3300] Cass R. Sunstein. Precautions & nature. Dædalus, 137(2):49–58, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028180;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2008.
137.2.49.

Tetlock:2008:BT

[3301] Philip E. Tetlock and Michael Oppenheimer. The boundaries of the thinkable. Dædalus, 137(2): 59–70, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028181;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
2.59.

Schabas:2008:NDN

[3302] Margaret Schabas. Nature does nothing in vain. Dædalus, 137 (2):71–79, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028182;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
2.71.

Kevles:2008:CES

[3303] Daniel J. Kevles. The contested Earth: Science, equity & the environment. Dædalus, 137(2): 80–95, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028183;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        2.80.

Mason:2008:RE

[3304] John Hope Mason. Rousseau in England. Dædalus, 137(2): 96–101, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028184;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        2.96.

Wright:2008:BWB

[3305] C. D. Wright. Because we have been here before. Dædalus, 137 (2):102–103, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20028185;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        2.102.

Brown:2008:BS

[3306] Wesley Brown. Body and soul. Dædalus, 137(2):104–117, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20028186;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/daed.2008.
                                        137.2.104.

Ayres:2008:OE

[3307] Ian Ayres. On options & epidemics. Dædalus, 137(2):118–121, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20028187;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2008.
137.2.118.

Kraut:2008:PGW

[3308] Richard Kraut. On philosophy as a guide to well-being. Dædalus, 137 (2):122–125, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028188;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
2.122.

Anonymous:2008:BMb

[3309] Anonymous. Back matter. Dædalus, 137(2):??, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028189.

Anonymous:2008:FMb

[3310] Anonymous. Front matter. Dædalus, 137(2):??, Spring 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20028174.

Anonymous:2008:JAAa

[3311] Anonymous. Journal of the American Academy of Arts & Sciences. Dædalus, 137(3):3, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543792.

McMahon:2008:FTS

[3312] Darrin M. McMahon. Fear & trembling, strangers & strange lands. Dædalus, 137(3):5–17, Summer 2008. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/40543793;
 http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
3.5.

Jacob:2008:CLC

[3313] Margaret C. Jacob. The cosmopolitan as a lived category. Dædalus, 137 (3):18–25, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543794;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
3.18.

Cheah:2008:WWW

[3314] Pheng Cheah. What is a world? On world literature as world-making activity. Dædalus, 137(3):26–38, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/40543795;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2008.
137.3.26.

Smith:2008:PMC

[3315] Rogers M. Smith. Paths to a more cosmopolitan human condition. Dædalus, 137(3):39–49, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543796;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
3.39.

Long:2008:CCG

[3316] A. A. Long. The concept of the cosmopolitan in Greek & Roman thought. Dædalus, 137(3): 50–58, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543797;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
3.50.

Rosenblatt:2008:RA

[3317] Helena Rosenblatt. Rousseau, the anticosmopolitan? Dædalus, 137 (3):59–67, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543798;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
3.59.

Scheffler:2008:CJI

[3318] Samuel Scheffler. Cosmopolitanism, justice & institutions. Dædalus, 137 (3):68–77, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543799;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
3.68.

Nussbaum:2008:TGS

[3319] Martha C. Nussbaum. Toward a globally sensitive patriotism. Dædalus, 137 (3):78–93, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543800;
http://www.mitpressjournals.org/

                                        doi/pdfplus/10.1162/daed.2008.137.
                                        3.78.

Benhabib:2008:LHR

[3320] Seyla Benhabib. The legitimacy of human rights. Dædalus, 137(3): 94–104, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40543801;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        3.94.

Calhoun:2008:CMS

[3321] Craig Calhoun. Cosmopolitanism in the modern social imaginary. Dædalus, 137(3):105–114, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/40543802;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/daed.2008.
                                        137.3.105.

Benedict:2008:MOM

[3322] Elizabeth Benedict. Murder one: Mad Dog Taborsky & me. Dædalus, 137 (3):115–124, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40543803;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        3.115.

Pinsky:2008:P

[3323] Robert Pinsky. Paradise. Dædalus, 137(3):125, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40543804;
                                        http://www.mitpressjournals.org/

doi/pdfplus/10.1162/daed.2008.137.
3.125.

Ohlin:2008:T

[3324] Alix Ohlin. The teacher. Dædalus, 137 (3):126–134, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543805;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
3.126.

Finlayson-Pitts:2008:MNA

[3325] Barbara J. Finlayson-Pitts. On man, nature & air pollution. Dædalus, 137 (3):135–138, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543806;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
3.135.

Kirby:2008:CEA

[3326] William C. Kirby. On Chinese, European & American universities. Dædalus, 137(3):139–146, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/40543807;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2008.
137.3.139.

Stanish:2008:MPW

[3327] Charles S. Stanish. On museums in a postmodern world. Dædalus, 137 (3):147–149, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543808;
http://www.mitpressjournals.org/

                                        doi/pdfplus/10.1162/daed.2008.137.
                                        3.147.

Anonymous:2008:BMc

[3328] Anonymous. Back matter. Dædalus, 137(3):??, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40543809.

Anonymous:2008:FMc

[3329] Anonymous. Front matter. Dædalus, 137(3):??, Summer 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40543791.

Anonymous:2008:JAAb

[3330] Anonymous. Journal of the American Academy of Arts & Sciences. Dædalus, 137(4):3–4, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544053.

Greenhouse:2008:IWW

[3331] Linda Greenhouse. Independence: Why & from what? Dædalus, 137(4):5–7, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544054;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        4.5.

OConnor:2008:FIC

[3332] Sandra Day O’Connor. Fair & independent courts. Dædalus, 137 (4):8–10, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544055;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
4.8.

Jamieson:2008:WIP

[3333] Kathleen Hall Jamieson and Bruce W.

Hardy. Will ignorance & partisan election of judges undermine public trust in the judiciary? Dædalus, 137(4):11–15, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544056;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
4.11.

Burbank:2008:JIJ

[3334] Stephen B. Burbank. Judicial independence, judicial accountability & interbranch relations. Dædalus, 137(4):16–27, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544057;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
4.16.

Resnik:2008:IFJ

[3335] Judith Resnik. Interdependent Federal judiciaries: Puzzling about why & how to value the independence of which judges. Dædalus, 137(4):28–47, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544058;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
4.28.

Jackson:2008:PJI

[3336] Vicki C. Jackson. Packages of judicial independence: Implications for reform proposals on the selection & tenure of Article III judges. Dædalus, 137(4):48–63, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544059;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        4.48.

Dinh:2008:TJI

[3337] Viet D. Dinh. Threats to judicial independence, real & imagined. Dædalus, 137(4):64–73, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544060;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        4.64.

Schumer:2008:CCR

[3338] Charles E. Schumer. Congress & the Court: Restoring balance. Dædalus, 137(4):74–76, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544061;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        4.74.

Wilkinson:2008:CCJ

[3339] J. Harvie Wilkinson III. Congress & the Court: Judicial confirmation. Dædalus, 137(4):77–80, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544062;
                                        http://www.mitpressjournals.org/

doi/pdfplus/10.1162/daed.2008.137.
4.77.

Post:2008:CCS

[3340] Robert C. Post. Congress & the Court: The scope of national legislative power. Dædalus, 137 (4):81–85, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544063;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
4.81.

Geyh:2008:MJS

[3341] Charles Gardner Geyh. Methods of judicial selection & their impact on judicial independence. Dædalus, 137(4):86–101, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544064;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
4.86.

Brandenburg:2008:KCI

[3342] Bert Brandenburg and Roy A. Schotland. Keeping courts impartial amid changing judicial elections. Dædalus, 137(4):102–109, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544065;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2008.137.
4.102.

George:2008:WSC

[3343] Ronald M. George. Why state courts — and state-court elections — matter. Dædalus, 137 (4):110–121, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544066;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        4.110.

Marshall:2008:TRL

[3344] Margaret H. Marshall. Threats to the rule of law: State courts, public expectations & political attitudes. Dædalus, 137(4):122–128, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544067;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        4.122.

McGregor:2008:SCJ

[3345] Ruth V. McGregor. State courts & judicial outreach. Dædalus, 137 (4):129–138, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544068;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        4.129.

Breyer:2008:SAB

[3346] Stephen Breyer. Serving America’s best interests. Dædalus, 137 (4):139–143, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544069;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2008.137.
                                        4.139.

Anonymous:2008:BMd

[3347] Anonymous. Back matter. Dædalus, 137(4):??, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544070.

Anonymous:2008:FMd

[3348] Anonymous. Front matter. Dædalus, 137(4):??, Fall 2008. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544052.

Anonymous:2009:D

[3349] Anonymous. Dædalus. Dædalus, 138(1):3, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543866.

Spacks:2009:RH

[3350] Patricia Meyer Spacks and Leslie Berlowitz. Reflecting on the humanities. Dædalus, 138(1):5–7, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/40543867;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2009.
138.1.5.

Randel:2009:PGK

[3351] Don Michael Randel. The public good: Knowledge as the foundation for a democratic society. Dædalus, 138(1):8–12, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543868;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
1.8.

Franke:2009:PHC

[3352] Richard J. Franke. The power of the humanities & a challenge to humanists. Dædalus, 138(1): 13–23, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543869;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
1.13.

Ayers:2009:WHL

[3353] Edward L. Ayers. Where the humanities live. Dædalus, 138(1): 24–34, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543870;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
1.24.

Oakley:2009:HLA

[3354] Francis Oakley. The humanities in liberal arts colleges: Another instance of collegiate exceptionalism? Dædalus, 138(1):35–51, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543871;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
1.35.

Early:2009:HSC

[3355] Gerald Early. The humanities & social change. Dædalus, 138 (1):52–57, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543872;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
1.52.

Wood:2009:WL

[3356] Michael Wood. A world without literature? Dædalus, 138 (1):58–67, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543873;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
1.58.

Warren:2009:TDP

[3357] Rosanna Warren. The twelfth day [poem]. Dædalus, 138(1): 68–70, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543874;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
1.68.

Bynum:2009:PCO

[3358] Caroline W. Bynum. Perspectives, connections & objects, What’s happening in history now? Dædalus, 138(1):71–86, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543875;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
1.71.

Grafton:2009:ASR

[3359] Anthony Grafton. Apocalypse in the stacks? The research library in the Age of Google. Dædalus, 138 (1):87–98, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543876;
http://www.mitpressjournals.org/

                                        doi/pdfplus/10.1162/daed.2009.138.
                                        1.87.

ODonnell:2009:EHD

[3360] James J. O’Donnell. Engaging the humanities: The digital humanities. Dædalus, 138(1):99–104, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/40543877;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/daed.2009.
                                        138.1.99.

Shelemay:2009:PHE

[3361] Kay Kaufman Shelemay. Performing the humanities at the Ethiopian Millennium. Dædalus, 138(1): 105–109, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40543878;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2009.138.
                                        1.105.

Woodward:2009:FHP

[3362] Kathleen Woodward. The future of the humanities — in the present & in public. Dædalus, 138(1): 110–123, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40543879;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2009.138.
                                        1.110.

Zuckerman:2009:RTF

[3363] Harriet Zuckerman and Ronald G. Ehrenberg. Recent trends in funding for the academic humanities & their implications. Dædalus, 138 (1):124–146, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543880;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
1.124.

Anonymous:2009:BMa

[3364] Anonymous. Back matter. Dædalus, 138(1):??, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543881.

Anonymous:2009:FMa

[3365] Anonymous. Front matter. Dædalus, 138(1):??, Winter 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543865.

Song:2009:WDI

[3366] Sarah Song. What does it mean to be an American? Dædalus, 138 (2):1, 31–40, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543933;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
2.31.

Anonymous:2009:JAAa

[3367] Anonymous. Journal of the American Academy of Arts & Sciences. Dædalus, 138(2):3–4, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543929.

Anonymous:2009:F

[3368] Anonymous. Foreword. Dædalus, 138(2):5, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543930;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
2.5.

Greenberg:2009:TPT

[3369] David Greenberg. Torchlight parades for the television age: The Presidential debates as political ritual. Dædalus, 138(2):6–19, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543931;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
2.6.

Hsu:2009:HMG

[3370] Hsuan L. Hsu and Martha Lincoln. Health media & global inequalities. Dædalus, 138(2):20–30, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/40543932;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2009.
138.2.20.

Ratner-Rosenhagen:2009:AIR

[3371] Jennifer Ratner-Rosenhagen. Antiintellectualism as romantic discourse. Dædalus, 138(2):41–52, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/40543934;
 http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
2.41.

Mehrotra:2009:IFM

[3372] Ajay K. Mehrotra. The intellectual foundations of the modern American fiscal state. Dædalus, 138 (2):53–62, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543935;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
2.53.

Kaag:2009:PLE

[3373] John Jacob Kaag. Pragmatism & the lessons of experience. Dædalus, 138(2):63–72, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543936;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
2.63.

Klemek:2009:RFN

[3374] Christopher Klemek. The rise & fall of new left urbanism. Dædalus, 138(2):73–82, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543937;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
2.73.

Puskar:2009:RRE

[3375] Jason Puskar. Risking Ralph Ellison. Dædalus, 138(2):83–93, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/40543938;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2009.
138.2.83.

Colwell-Chanthaphonh:2009:RAA

[3376] Chip Colwell-Chanthaphonh. Reconciling American archaeology & native America. Dædalus, 138(2): 94–104, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543939;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
2.94.

Weiner:2009:LLC

[3377] Sharon K. Weiner. Looking out, looking in: competing organizational interests & the proliferation of Soviet WMD expertise. Dædalus, 138 (2):105–114, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543940;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
2.105.

MacDonald:2009:RAF

[3378] Paul K. MacDonald. Rebalancing American foreign policy. Dædalus, 138 (2):115–125, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543941;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
2.115.

Feimster:2009:GBB

[3379] Crystal N. Feimster. General Benjamin Butler & the threat of sexual violence during the American Civil War. Dædalus, 138(2): 126–134, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543942;
http://www.mitpressjournals.org/

doi/pdfplus/10.1162/daed.2009.138.
2.126.

Collins:2009:SF

[3380] Arda Collins. From speaking in the fall. Dædalus, 138(2):135, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/40543943;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2009.
138.2.135.

Dickman:2009:D

[3381] Matthew Dickman. Divinity. Dædalus, 138(2):136–137, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543944;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
2.136.

Martin:2009:EB

[3382] Dawn Lundy Martin. Excerpts from Discipline. Dædalus, 138(2): 138–139, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543945;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
2.138.

ORourke:2009:OC

[3383] Meghan O’Rourke. Ophelia to the Court. Dædalus, 138(2):140, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/40543946;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2009.
138.2.140.

Zapruder:2009:NL

[3384] Matthew Zapruder. The new lustration. Dædalus, 138(2):141–142, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/40543947;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/daed.2009.
                                        138.2.141.

Anonymous:2009:BMb

[3385] Anonymous. Back matter. Dædalus, 138(2):??, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40543948.

Anonymous:2009:FMb

[3386] Anonymous. Front matter. Dædalus, 138(2):??, Spring 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40543928.

Anonymous:2009:JAAb

[3387] Anonymous. Journal of the American Academy of Arts & Sciences. Dædalus, 138(3):3, Summer 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40543982.

Anonymous:2009:I

[3388] Anonymous. Introduction. Dædalus, 138(3):5–6, Summer 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40543983;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2009.138.
                                        3.5.

Rose:2009:CFH

[3389] Hilary Rose and Steven Rose. The changing face of human nature. Dædalus, 138(3):7–20, Summer 2009. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/40543984;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
3.7.

Gazzaniga:2009:HPA

[3390] Michael S. Gazzaniga. Humans:

The party animal. Dædalus, 138 (3):21–34, Summer 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543985;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
3.21.

Pippin:2009:NN

[3391] Robert B. Pippin. Natural & normative. Dædalus, 138(3):35–43, Summer 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/40543986;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2009.
138.3.35.

Hacking:2009:HAA

[3392] Ian Hacking. Humans, aliens & autism. Dædalus, 138(3):44–59, Summer 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/40543987;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2009.
138.3.44.

Darwin:2009:CMP

[3393] Charles Darwin. Comparison of the mental powers of man and the lower animals — continued. Dædalus, 138 (3):60–67, Summer 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543988;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
3.60.

Ritvo:2009:HH

[3394] Harriet Ritvo. Humans & humanists. Dædalus, 138(3):68–78, Summer 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/40543989;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2009.
138.3.68.

Harpham:2009:HDW

[3395] Geoffrey Galt Harpham. How do we know what we are? The science of language & human selfunderstanding. Dædalus, 138(3): 79–91, Summer 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543990;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
3.79.

Appiah:2009:EMP

[3396] Kwame Anthony Appiah. Experimental moral psychology. Dædalus, 138 (3):92–102, Summer 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543991;
http://www.mitpressjournals.org/

doi/pdfplus/10.1162/daed.2009.138.
3.92.

Dove:2009:CSC

[3397] Rita Dove. The countess shares confidences over Karneval chocolate. Dædalus, 138(3):103–104, Summer 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/40543992;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2009.
138.3.103.

Anonymous:2009:BMc

[3398] Anonymous. Back matter. Dædalus, 138(3):??, Summer 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543993.

Anonymous:2009:FMc

[3399] Anonymous. Front matter. Dædalus, 138(3):??, Summer 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543981.

Anonymous:2009:JAAc

[3400] Anonymous. Journal of the American Academy of Arts & Sciences. Dædalus, 138(4):3–4, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543996.

Miller:2009:NPN

[3401] Steven E. Miller and Scott D.

Sagan. Nuclear power without nuclear proliferation? Dædalus, 138(4):7–18, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543997;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
4.7.

Lester:2009:GNP

[3402] Richard K. Lester and Robert Rosner. The growth of nuclear power: Drivers & constraints. Dædalus, 138(4):19–30, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543998;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
4.19.

Socolow:2009:BRN

[3403] Robert H. Socolow and Alexander Glaser. Balancing risk: Nuclear energy & climate change. Dædalus, 138(4):31–44, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40543999;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
4.31.

Joskow:2009:EFN

[3404] Paul L. Joskow and John E. Parsons. The economic future of nuclear power. Dædalus, 138 (4):45–59, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544000;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
4.45.

Feiveson:2009:SVN

[3405] Harold A. Feiveson. A skeptic’s view of nuclear energy. Dædalus, 138(4):60–70, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544001;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
4.60.

Goldemberg:2009:NED

[3406] José Goldemberg. Nuclear energy in developing countries. Dædalus, 138(4):71–80, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544002;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
4.71.

Rowe:2009:NPC

[3407] John W. Rowe. Nuclear power in a carbon-constrained world. Dædalus, 138(4):81–90, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544003;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
4.81.

Lauvergeon:2009:NRO

[3408] Anne Lauvergeon. The nuclear Renaissance: an opportunity to enhance the culture of nonproliferation. Dædalus, 138(4):91–99, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544004;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
4.91.

Meserve:2009:GNS

[3409] Richard A. Meserve. The global nuclear safety regime. Dædalus, 138 (4):100–111, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544005;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2009.138.
                                        4.100.

Bunn:2009:RGR

[3410] Matthew Bunn. Reducing the greatest risks of nuclear theft & terrorism. Dædalus, 138(4):112–123, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/40544006;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/daed.2009.
                                        138.4.112.

Schelling:2009:WNW

[3411] Thomas C. Schelling. A world without nuclear weapons? Dædalus, 138(4):124–129, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544007;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2009.138.
                                        4.124.

Doty:2009:MDB

[3412] Paul Doty. The minimum deterrent & beyond. Dædalus, 138 (4):130–139, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544008;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2009.138.
                                        4.130.

Lodgaard:2009:TNW

[3413] Sverre Lodgaard. Toward a nuclearweapons–free world. Dædalus, 138 (4):140–152, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544009;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
4.140.

Nunn:2009:TST

[3414] Sam Nunn. Taking steps toward a world free of nuclear weapons. Dædalus, 138(4):153–156, Fall 2009. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/40544010;
 http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
4.153.

Sagan:2009:SRN

[3415] Scott D. Sagan. Shared responsibilities for nuclear disarmament. Dædalus, 138(4):157–168, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544011;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2009.138.
4.157.

Anonymous:2009:C

[3416] Anonymous. Contributors. Dædalus, 138(4):169–171, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/daed.2009.138.4.
169.

Anonymous:2009:BMd

[3417] Anonymous. Back matter. Dædalus, 138(4):??, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544012.

Anonymous:2009:FMd

[3418] Anonymous. Front matter. Dædalus, 138(4):??, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40543995.

Anonymous:2009:VI

[3419] Anonymous. Volume information. Dædalus, 138(4):??, Fall 2009. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40543994.

Goldschmidt:2010:MNF

[3420] Pierre Goldschmidt. Multilateral nuclear fuel supply guarantees & spent fuel management: What are the priorities? Dædalus, 139 (1):7–19, Winter 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544040;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2010.139.
                                        1.7.

Perkovich:2010:GIU

[3421] George Perkovich. Global implications of the U.S.–India deal. Dædalus, 139(1):20–31, Winter 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544041;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2010.139.
                                        1.20.

McCombie:2010:KRB

[3422] Charles McCombie and Thomas Isaacs. The key role of the back-end in the nuclear fuel cycle. Dædalus, 139 (1):32–43, Winter 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544042;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2010.139.
1.32.

Hecker:2010:LLN

[3423] Siegfried S. Hecker. Lessons learned from the North Korean nuclear crises. Dædalus, 139(1):44–56, Winter 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/40544043;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2010.
139.1.44.

Dhanapala:2010:MND

[3424] Jayantha Dhanapala. The management of NPT diplomacy. Dædalus, 139(1):57–67, Winter 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544044;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2010.139.
1.57.

Potter:2010:NSN

[3425] William C. Potter. The NPT & the sources of nuclear restraint. Dædalus, 139(1):68–81, Winter 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544045;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2010.139.
1.68.

Suzuki:2010:TRN

[3426] Atsuyuki Suzuki. Toward a robust nuclear management system. Dædalus, 139(1):82–92, Winter 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544046;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2010.139.
                                        1.82.

Shaker:2010:NPA

[3427] Mohamed I. Shaker. Nuclear power in the Arab world & the regionalization of the nuclear fuel cycle: an Egyptian perspective. Dædalus, 139 (1):93–104, Winter 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544047;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2010.139.
                                        1.93.

Maleki:2010:INF

[3428] Abbas Maleki. Iran’s nuclear file: Recommendations for the future. Dædalus, 139(1):105–116, Winter 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/40544048;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2010.139.
                                        1.105.

Diyakov:2010:NRP

[3429] Anatoly S. Diyakov. The nuclear “Renaissance” & preventing the spread of enrichment & reprocessing technologies: a Russian view. Dædalus, 139 (1):117–125, Winter 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544049;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2010.139.
1.117.

Miller:2010:ANF

[3430] Steven E. Miller and Scott D. Sagan.

Alternative nuclear futures. Dædalus, 139(1):126–137, Winter 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/40544050;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2010.139.
1.126.

Anonymous:2010:C

[3431] Anonymous. Contributors. Dædalus, 139(1):138–140, Winter 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/daed.2010.139.1.
138.

Ghiglione:2010:IFN

[3432] Loren Ghiglione. Introduction: the future of news. Dædalus, 139 (2):5–7, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20749818;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2010.139.
2.5.

Gans:2010:NNM

[3433] Herbert J. Gans. News & the news media in the digital age: implications for democracy. Dædalus, 139(2):8–17, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20749819;

                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2010.139.
                                        2.8.

Jamieson:2010:TLF

[3434] Kathleen Hall Jamieson and Jeffrey A. Gottfried. Are there lessons for the future of news from the 2008 Presidential campaign? Dædalus, 139 (2):18–25, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20749820;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2010.139.
                                        2.18.

Giles:2010:NEM

[3435] Robert H. Giles. New economic models for U.S. journalism. Dædalus, 139(2):26–38, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20749821;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2010.139.
                                        2.26.

Abramson:2010:SQJ

[3436] Jill Abramson. Sustaining quality journalism. Dædalus, 139(2): 39–44, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20749822;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2010.139.
                                        2.39.

Houston:2010:FIJ

[3437] Brant Houston. The future of investigative journalism. Dædalus, 139 (2):45–56, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20749823;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2010.139.
2.45.

Kennedy:2010:FSN

[3438] Donald Kennedy. The future of science news. Dædalus, 139(2): 57–65, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20749824;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2010.139.
2.57.

Zuckerman:2010:IRA

[3439] Ethan Zuckerman. International reporting in the age of participatory media. Dædalus, 139(2): 66–75, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20749825;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2010.139.
2.66.

Stephens:2010:CWJ

[3440] Mitchell Stephens. The case for wisdom journalism — and for journalists surrendering the pursuit of news. Dædalus, 139(2):76–88, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20749826;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/daed.2010.
139.2.76.

Singer:2010:JEA

[3441] Jane B. Singer. Journalism ethics amid structural change. Dædalus, 139(2):89–99, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20749827;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2010.139.
                                        2.89.

Schudson:2010:POD

[3442] Michael Schudson. Political observatories, databases & news in the emerging ecology of public information. Dædalus, 139(2):100–109, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20749828;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2010.139.
                                        2.100.

Fuller:2010:WHN

[3443] Jack Fuller. What is happening to news? Dædalus, 139(2): 110–118, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20749829;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2010.139.
                                        2.110.

Sagan:2010:IFN

[3444] Paul Sagan and Tom Leighton. The Internet & the future of news. Dædalus, 139(2):119–125, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20749830;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/daed.2010.139.
                                        2.119.

King:2010:CKI

[3445] Susan King. The Carnegie-Knight Initiative on the Future of Journalism Education: improving how journalists are educated & how their audiences are informed. Dædalus, 139 (2):126–137, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20749831;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2010.139.
2.126.

Ghiglione:2010:DSF

[3446] Loren Ghiglione. Does science fiction — yes, science fiction — suggest suggest futures for news? Dædalus, 139 (2):138–150, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20749832;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2010.139.
2.138.

Delanty:2010:DAL

[3447] Greg Delanty. In a diner above the Lamoille River [poem]. Dædalus, 139(2):151, Spring 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20749833;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/daed.2010.139.
2.151.

Loury:2010:CMI

[3448] Glenn C. Loury and Bruce Western.

The challenge of mass incarceration in America. Dædalus, 139(3):5–7, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20749837.

Loury:2010:ICM

[3449] Glenn C. Loury and Bruce Western. Introduction: The challenge of mass incarceration in America. Dædalus, 139(3):5–7, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_e_00018.

Western:2010:ISI

[3450] Bruce Western and Becky Pettit. Incarceration & social inequality. Dædalus, 139(3):8–19, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20749838;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00019.

Sampson:2010:PPL

[3451] Robert J. Sampson and Charles Loeffler. Punishment’s place: the local concentration of mass incarceration. Dædalus, 139(3):20–31, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20749839;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/DAED_a_
                                        00020.

Kruttschnitt:2010:PWI

[3452] Candace Kruttschnitt. The paradox of women’s imprisonment. Dædalus, 139(3):32–42, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20749840;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00021.

Fagan:2010:CJC

[3453] Jeffrey Fagan. The contradictions of juvenile crime & punishment. Dædalus, 139(3):43–61, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20749841;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00022.

Gottschalk:2010:CBR

[3454] Marie Gottschalk. Cell blocks & red ink: mass incarceration, the great recession & penal reform. Dædalus, 139 (3):62–73, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20749842;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00023.

Wacquant:2010:CRH

[3455] Loïc Wacquant. Class, race & hyperincarceration in revanchist America. Dædalus, 139(3):74–90, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20749843;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/DAED_a_
00024.

Simon:2010:CTA

[3456] Jonathan Simon. Clearing the “troubled assets” of America’s punishment bubble. Dædalus, 139(3): 91–101, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20749844;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00025.

Lacey:2010:AIC

[3457] Nicola Lacey. American imprisonment in comparative perspective. Dædalus, 139(3):102–114, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20749845;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/DAED_a_
                                        00026.

Kleiman:2010:TFP

[3458] Mark A. R. Kleiman. Toward fewer prisoners & less crime. Dædalus, 139 (3):115–123, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20749846;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00027.

Weisberg:2010:DPV

[3459] Robert Weisberg and Joan Petersilia. The dangers of Pyrrhic victories against mass incarceration. Dædalus, 139(3):124–133, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20749847;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/DAED_a_
                                        00028.

Loury:2010:CIS

[3460] Glenn C. Loury. Crime, inequality & social justice. Dædalus, 139 (3):134–140, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/20749848;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00029.

Knight:2010:WWV

[3461] Etheridge Knight. A Wasp woman visits a black junkie in prison [poem]. Dædalus, 139(3):141– 142, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/20749849;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00030.

Clifton:2010:CDT

[3462] Lucille Clifton. Cruelty. don’t talk to me about cruelty [poems]. Dædalus, 139(3):143–144, Summer 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/20749850;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/DAED_a_
00031.

Friedman:2010:I

[3463] Benjamin M. Friedman and Robert M.

Solow. Introduction. Dædalus, 139(4):5–8, Fall 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790420;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_e_00037.

Friedman:2010:OFS

[3464] Benjamin M. Friedman. Is our financial system serving us well? Dædalus, 139(4):9–21, Fall 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790421;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00038.

Solow:2010:BT

[3465] Robert M. Solow. The bigger they are …. Dædalus, 139 (4):22–30, Fall 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/25790422;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00039.

Zingales:2010:LLE

[3466] Luigi Zingales. Learning to live with not-so-efficient markets. Dædalus, 139(4):31–40, Fall 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/25790423;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00040.

Stein:2010:SSB

[3467] Jeremy C. Stein. Securitization, shadow banking & financial fragility. Dædalus, 139(4):41–51, Fall 2010. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/25790424;
                                         http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00041.

Bebchuk:2010:HFB

[3468] Lucian A. Bebchuk. How to fix bankers’ pay. Dædalus, 139 (4):52–60, Fall 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/25790425;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00042.

McCarty:2010:PFF

[3469] Nolan McCarty, Keith T. Poole, Thomas Romer, and Howard Rosenthal. Political fortunes: on finance & its regulation. Dædalus, 139(4):61–73, Fall 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790426;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00043.

Goodhart:2010:LFC

[3470] C. A. E. Goodhart. Lessons from the financial crisis for monetary policy. Dædalus, 139(4):74–82, Fall 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790427;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00044.

Hall:2010:FS

[3471] Robert E. Hall. Fiscal stimulus. Dædalus, 139(4):83–94, Fall 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790428;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00045.

Glaeser:2010:HPW

[3472] Edward L. Glaeser. Housing policy in the wake of the crash. Dædalus, 139(4):95–106, Fall 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790429;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00046.

Eichengreen:2010:IFR

[3473] Barry Eichengreen. International financial regulation after the crisis. Dædalus, 139(4):107–114, Fall 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/25790430;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00047.

Temin:2010:GRG

[3474] Peter Temin. The Great Recession & the Great Depression. Dædalus, 139(4):115–124, Fall 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/25790431;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00048.

Hamilton:2010:WAH

[3475] Alexander Hamilton. From the writings of Alexander Hamilton. Dædalus, 139(4):125–126, Fall 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/25790432;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00049.

Notley:2010:CEM

[3476] Alice Notley. The codex eats me [poem]. Dædalus, 139(4):127, Fall 2010. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/25790433;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/DAED_a_
                                        00050.

Early:2011:TWR

[3477] Gerald Early. The two worlds of race revisited: a meditation on race in the Age of Obama. Dædalus, 140(1):11–27, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/25790438;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00055.

Franklin:2011:TWR

[3478] John Hope Franklin. The two worlds of race: a historical view. Dædalus, 140(1):28–43, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790439;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00056.

Ferguson:2011:FER

[3479] Jeffrey B. Ferguson. Freedom, equality, race. Dædalus, 140(1): 44–52, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790440;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00057.

Geary:2011:RLM

[3480] Daniel Geary. Racial liberalism, the Moynihan Report & the “Dædalus” Project on “The Negro American”. Dædalus, 140 (1):53–66, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790441;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00058.

Martin:2011:PAA

[3481] Waldo E. Martin, Jr. Precious African American memories, postracial dreams & the American nation. Dædalus, 140(1):67–78, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/

                                        stable/25790442;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/DAED_a_
                                        00059.

Carpio:2011:RIB

[3482] Glenda R. Carpio. Race & inheritance in Barack Obama’s “Dreams from My Father”. Dædalus, 140 (1):79–89, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/25790443;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00060.

Gautier:2011:PRA

[3483] Amina Gautier. On post-racial America in the Age of Obama. Dædalus, 140(1):90–94, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/25790444;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00061.

Shelby:2011:JRC

[3484] Tommie Shelby. Justice & racial conciliation: Two visions. Dædalus, 140 (1):95–107, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/25790445;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00062.

Sundquist:2011:WDD

[3485] Eric J. Sundquist. “We dreamed a dream”: Ralph Ellison, Martin Luther King, Jr. & Barack Obama. Dædalus, 140(1):108–124, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/25790446;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00063.

Walker:2011:WLO

[3486] Clarence E. Walker. “We’re losing our country”: Barack Obama, race & the Tea Party. Dædalus, 140 (1):125–130, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790447;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00064.

See correction [3509].

Griffin:2011:LLM

[3487] Farah Jasmine Griffin. At last … ?: Michelle Obama, Beyoncé, race & history. Dædalus, 140(1): 131–141, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790448;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00065.

Sollors:2011:ONW

[3488] Werner Sollors. “Obligations to Negroes who would be kin if they were not Negro”. Dædalus, 140 (1):142–153, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790449;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00066.

Jocson:2011:PNR

[3489] Korina Jocson. Poetry in a new race era. Dædalus, 140(1): 154–162, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790450;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00067.

Hsu:2011:SJZ

[3490] Hua Hsu. Seeing Jay-Z in Taipei. Dædalus, 140(1):163–173, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/25790451;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/DAED_a_
00068.

Hollinger:2011:CPR

[3491] David A. Hollinger. The concept of post-racial: How its easy dismissal obscures important questions. Dædalus, 140(1):174–182, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790452;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00069.

McPherson:2011:PP

[3492] James Alan McPherson. Pursuit of the “Pneuma”. Dædalus, 140 (1):183–188, Winter 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/25790453;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00070.

Bobo:2011:SBJ

[3493] Lawrence D. Bobo. Somewhere between Jim Crow & post-racialism: Reflections on the racial divide in America today. Dædalus, 140 (2):11–36, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23047449;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00091.

Massey:2011:PFA

[3494] Douglas S. Massey. The past & future of American civil rights. Dædalus, 140(2):37–54, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23047450;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00076.

Wilson:2011:DSR

[3495] William Julius Wilson. “The Declining Significance of Race”: Revisited & revised. Dædalus, 140 (2):55–69, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23047451;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00077.

Heckman:2011:AFB

[3496] James J. Heckman. The American family in black & white: a postracial strategy for improving skills to promote equality. Dædalus, 140 (2):70–89, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23047452;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00078.

Nisbett:2011:AGP

[3497] Richard E. Nisbett. The achievement gap: Past, present & future. Dædalus, 140(2):90–100, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23047453;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00079.

Klarman:2011:SCB

[3498] Michael J. Klarman. Has the Supreme Court been more a friend or foe to African Americans? Dædalus, 140 (2):101–108, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23047454;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00080.

Sabbagh:2011:AAU

[3499] Daniel Sabbagh. Affirmative action: The U.S. experience in comparative perspective. Dædalus, 140 (2):109–120, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23047455;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00081.

Smith:2011:CHB

[3500] Rogers M. Smith, Desmond S. King, and Philip A. Klinkner. Challenging history: Barack Obama & American racial politics. Dædalus, 140 (2):121–135, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23047456;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00082.

Lee:2011:SRP

[3501] Taeku Lee. Somewhere over the rainbow? Post-racial & pan-racial politics in the Age of Obama. Dædalus, 140(2):136–150, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23047457;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00083.

Hochschild:2011:DAR

[3502] Jennifer L. Hochschild, Vesla M.

Weaver, and Traci Burch. Destabilizing the American racial order. Dædalus, 140(2):151–165, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23047458;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00084.

Richeson:2011:IMI

[3503] Jennifer A. Richeson and Maureen A. Craig. Intra-minority intergroup relations in the Twenty-First Century. Dædalus, 140(2): 166–175, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23047459;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00085.

Morgan:2011:HHG

[3504] Marcyliena Morgan and Dionne Bennett. Hip-hop & the global imprint of a black cultural form. Dædalus, 140(2):176–196, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23047460;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00086.

Cohen:2011:MMP

[3505] Cathy J. Cohen. Millennials & the myth of the post-racial society: Black youth, intra-generational divisions & the continuing racial divide in American politics. Dædalus, 140 (2):197–205, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23047461;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00087.

Young:2011:BMB

[3506] Alford A. Young, Jr. The black masculinities of Barack Obama: Some implications for African American men. Dædalus, 140(2):206–214, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/23047462;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/DAED_a_
                                        00088.

Waldinger:2011:INA

[3507] Roger Waldinger. Immigration: The new American dilemma. Dædalus, 140 (2):215–225, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23047463;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00089.

Biondi:2011:CBH

[3508] Martha Biondi. Controversial blackness: The historical development & future trajectory of African American studies. Dædalus, 140(2): 226–237, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23047464;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00090.

Anonymous:2011:CWL

[3509] Anonymous. Correction: “We’re losing our country”: Barack Obama, race & the Tea Party. Dædalus, 140(2):238, Spring 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23047447;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_x_00092.

See [3486].

Perry:2011:F

[3510] William J. Perry. Foreword. Dædalus, 140(3):5–9, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23047343;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_e_00093.

Kennedy:2011:I

[3511] David M. Kennedy. Introduction. Dædalus, 140(3):10–15, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/23047344;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/DAED_e_
00094.

Freedman:2011:CSA

[3512] Lawrence Freedman. The counterrevolution in strategic affairs. Dædalus, 140(3):16–32, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23047345;

                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00095.

Linn:2011:UAF

[3513] Brian McAllister Linn. The U.S. Armed Forces’ view of war. Dædalus, 140(3):33–44, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23047346;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00096.

Mahnken:2011:WGS

[3514] Thomas G. Mahnken. Weapons: The growth & spread of the precisionstrike regime. Dædalus, 140(3): 45–57, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23047347;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00097.

Goldich:2011:AMC

[3515] Robert L. Goldich. American military culture from colony to empire. Dædalus, 140(3):58–74, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/23047348;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/DAED_a_
                                        00098.

Korb:2011:MFT

[3516] Lawrence J. Korb and David R. Segal. Manning & financing the Twenty-First-Century all-volunteer force. Dædalus, 140(3):75–87, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23047349;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00099.

Avant:2011:MCA

[3517] Deborah D. Avant and Renée de Nevers. Military contractors & the American way of war. Dædalus, 140 (3):88–99, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23047350;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00100.

Winter:2011:FW

[3518] Jay M. Winter. Filming war.

Dædalus, 140(3):100–111, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/23047351;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/DAED_a_
00101.

Sheehan:2011:FCS

[3519] James J. Sheehan. The future of conscription: Some comparative reflections. Dædalus, 140(3): 112–121, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23047352;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00102.

Bacevich:2011:WA

[3520] Andrew J. Bacevich. Whose Army?

Dædalus, 140(3):122–134, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/23047353;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/DAED_a_
00103.

Dunlap:2011:MIC

[3521] Charles J. Dunlap, Jr. The Military– Industrial Complex. Dædalus, 140 (3):135–147, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23047354;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00104.

McSally:2011:DAM

[3522] Martha E. McSally. Defending America in mixed company: Gender in the U.S. Armed Forces. Dædalus, 140 (3):148–164, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23047355;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00105.

Fidell:2011:ML

[3523] Eugene R. Fidell. Military law. Dædalus, 140(3):165–178, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/23047356;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/DAED_a_
                                        00106.

Shay:2011:C

[3524] Jonathan Shay. Casualties. Dædalus, 140(3):179–188, Summer 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/23047357;
                                         http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00107.

Clark:2011:I

[3525] David D. Clark. Introduction. Dædalus, 140(4):5–16, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23046910;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00111.

Horrigan:2011:WCB

[3526] John B. Horrigan. What are the consequences of being disconnected in a broadband-connected world? Dædalus, 140(4):17–31, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23046911;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00112.

Nissenbaum:2011:CAP

[3527] Helen Nissenbaum. A contextual approach to privacy online. Dædalus, 140(4):32–48, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23046912;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00113.

Cheshire:2011:OTT

[3528] Coye Cheshire. Online trust, trustworthiness, or assurance? Dædalus, 140(4):49–58, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23046913;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00114.

Cerf:2011:SC

[3529] Vinton G. Cerf. Safety in cyberspace. Dædalus, 140(4):59–69, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/

                                        stable/23046914;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/DAED_a_
                                        00115.

Mulligan:2011:DC

[3530] Deirdre K. Mulligan and Fred B. Schneider. Doctrine for cybersecurity. Dædalus, 140(4):70–92, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/23046915;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/DAED_a_
                                        00116.

Camp:2011:RRS

[3531] L. Jean Camp. Reconceptualizing the role of security user. Dædalus, 140(4):93–107, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23046916;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00117.

Garrett:2011:RPF

[3532] R. Kelly Garrett and Paul Resnick. Resisting political fragmentation on the Internet. Dædalus, 140(4): 108–120, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23046917;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00118.

Schlozman:2011:WSC

[3533] Kay Lehman Schlozman, Sidney Verba, and Henry E. Brady. Who speaks? Citizen political Voice on the Internet commons. Dædalus, 140 (4):121–139, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23046918;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00119.

Sproull:2011:PBN

[3534] Lee Sproull. Prosocial behavior on the net. Dædalus, 140 (4):140–153, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23046919;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00120.

Benkler:2011:WPI

[3535] Yochai Benkler. WikiLeaks and the PROTECT–IP act: a new public-private threat to the Internet commons. Dædalus, 140 (4):154–164, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23046920;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00121.

Longley:2011:N

[3536] Michael Longley. Notebook. Dædalus, 140(4):165, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23046922.

Longley:2011:PB

[3537] Michael Longley. Puff-ball. Dædalus, 140(4):165, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23046921.

Longley:2011:PML

[3538] Michael Longley. Poems by Michael Longley. Dædalus, 140(4):165–167, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.

                                        mitpressjournals.org/doi/pdfplus/
                                        10.1162/DAED_a_00122.

Longley:2011:F

[3539] Michael Longley. Firewood. Dædalus, 140(4):166, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23046923.

Longley:2011:TO

[3540] Michael Longley. Tongue orchid. Dædalus, 140(4):167, Fall 2011. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23046924.

Donoghue:2012:I

[3541] Denis Donoghue. Introduction. Dædalus, 141(1):5–10, Winter 2012. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.jstor.org/stable/23240297;
                                         http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00124.

Chafe:2012:ANT

[3542] William H. Chafe. The American narrative: Is there one & what is it? Dædalus, 141(1): 11–17, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23240298;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00125.

Tribe:2012:ACN

[3543] Laurence H. Tribe. America’s constitutional narrative. Dædalus, 141 (1):18–42, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240299;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00126.

Brooks:2012:NCC

[3544] Peter Brooks. Narratives of the constitutional covenant. Dædalus, 141 (1):43–51, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240300;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00127.

Parini:2012:AM

[3545] Jay Parini. The American mythos.

Dædalus, 141(1):52–60, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/23240301;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00128.

Adorno:2012:WWA

[3546] Rolena Adorno. On Western waters:

Anglo–American nonfictional narrative in the Nineteenth Century. Dædalus, 141(1):61–75, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240302;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00129.

Hollinger:2012:APC

[3547] David A. Hollinger. The accommodation of Protestant Christianity with the enlightenment: an old drama still being enacted. Dædalus, 141 (1):76–88, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240303;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00130.

Kerber:2012:WDR

[3548] Linda K. Kerber. Why diamonds really are a Girl’s best friend: Another American narrative. Dædalus, 141 (1):89–100, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240304;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00131.

Lewis:2012:EEC

[3549] David Levering Lewis. Exceptionalism’s exceptions: The changing American narrative. Dædalus, 141 (1):101–117, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240305;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00132.

Doctorow:2012:NC

[3550] E. L. Doctorow. Narrative C. Dædalus, 141(1):118–125, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240306;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00133.

Jen:2012:S

[3551] Gish Jen. Spooked. Dædalus, 141 (1):126–129, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240307;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00134.

Wood:2012:OC

[3552] Michael Wood. The other case.

Dædalus, 141(1):130–138, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/23240308;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/DAED_a_
00135.

Ferris:2012:SLB

[3553] William Ferris. Southern literature: a blending of oral, visual & musical voices. Dædalus, 141(1): 139–153, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240309;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00136.

Greenspan:2012:DCB

[3554] Charlotte Greenspan. Death comes to the Broadway musical. Dædalus, 141 (1):154–159, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240310;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00137.

Greenlaw:2012:M

[3555] Lavinia Greenlaw. On the mountain. Dædalus, 141(1):160, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/23240311.

Greenlaw:2012:PLG

[3556] Lavinia Greenlaw. Poems by Lavinia Greenlaw. Dædalus, 141 (1):160–161, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00138.

Greenlaw:2012:O

[3557] Lavinia Greenlaw. Otolith. Dædalus, 141(1):161, Winter 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23240312.

Fri:2012:AEF

[3558] Robert W. Fri and Stephen Ansolabehere. The alternative energy future: Challenges for technological change. Dædalus, 141 (2):5–9, Spring 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23240275;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00142.

Greenstone:2012:PTM

[3559] Michael Greenstone and Adam Looney. Paying too much for energy? The true costs of our energy choices. Dædalus, 141(2):10–30, Spring 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23240276;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00143.

Graetz:2012:EPP

[3560] Michael J. Graetz. Energy policy: Past or prologue? Dædalus, 141 (2):31–44, Spring 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240277;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00144.

Aldy:2012:UMA

[3561] Joseph E. Aldy and Robert N. Stavins.

Using the market to address climate change: Insights from theory & experience. Dædalus, 141 (2):45–60, Spring 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240278;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00145.

Ansolabehere:2012:APE

[3562] Stephen Ansolabehere and David M.

Konisky. The American public’s energy choice. Dædalus, 141 (2):61–71, Spring 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240279;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00146.

Schrag:2012:SGG

[3563] Daniel P. Schrag. Is shale gas good for climate change? Dædalus, 141 (2):72–80, Spring 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240280;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00147.

Moniz:2012:SET

[3564] Ernest J. Moniz. Stimulating energy technology innovation. Dædalus, 141(2):81–93, Spring 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23240281;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00148.

Yanosek:2012:PFE

[3565] Kassia Yanosek. Policies for financing the energy transition. Dædalus, 141(2):94–104, Spring 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23240282;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00149.

El-Ashry:2012:NPP

[3566] Mohamed T. El-Ashry. National policies to promote renewable energy. Dædalus, 141(2):105–110, Spring 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/23240283;   http://www.mitpressjournals.
                                        org/doi/pdfplus/10.1162/DAED_a_
                                        00150.

Meinwald:2012:P

[3567] Jerrold Meinwald. Prelude. Dædalus, 141(3):5–7, Summer 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

                                        www.jstor.org/stable/23240261;
                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00156.

Bell:2012:SHW

[3568] James F. Bell III. The search for habitable worlds: Planetary exploration in the 21st Century. Dædalus, 141(3):8–22, Summer 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240262;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00157.

Tao:2012:PUC

[3569] Terence Tao. E pluribus unum: From complexity, universality. Dædalus, 141 (3):23–34, Summer 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240263;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00158.

McEuen:2012:SM

[3570] Paul L. McEuen. Small machines. Dædalus, 141(3):35–44, Summer 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/23240264;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/DAED_a_
00159.

Nocera:2012:CWP

[3571] Daniel G. Nocera. Can we progress from solipsistic science to frugal innovation? Dædalus, 141(3): 45–52, Summer 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240265;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00160.

Arkani-Hamed:2012:FFP

[3572] Nima Arkani-Hamed. The future of fundamental physics. Dædalus, 141 (3):53–66, Summer 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240266;

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00161.

Bassler:2012:MMM

[3573] Bonnie L. Bassler. Microbes as menaces, mates & marvels. Dædalus, 141 (3):67–76, Summer 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240267;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00162.

Shubin:2012:FE

[3574] Neil H. Shubin. Fossils everywhere. Dædalus, 141(3):77–88, Summer 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/23240268;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/DAED_a_
00163.

Somerville:2012:DPL

[3575] Chris Somerville. Deciphering the parts list for the mechanical plant. Dædalus, 141(3):89–97, Summer 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240269;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00164.

Petsko:2012:CEN

[3576] Gregory A. Petsko. The coming epidemic of neurologic disorders: What science is — and should be — doing about it. Dædalus, 141(3): 98–107, Summer 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/23240270;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00165.

Tilman:2012:BES

[3577] David Tilman. Biodiversity & environmental sustainability amid human domination of global ecosystems. Dædalus, 141(3):108–120, Summer 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/23240271;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/DAED_a_
00166.

Berenbaum:2012:P

[3578] May R. Berenbaum. Postlude.

Dædalus, 141(3):121–124, Summer 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL http://www.jstor.org/
stable/23240272;   http://www.mitpressjournals.
org/doi/pdfplus/10.1162/DAED_a_
00167.

Epstein:2012:IPO

[3579] Lee Epstein. On the importance of public opinion. Dædalus, 141(4):5–8, Fall 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/41721976;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00169.

Greiner:2012:LLE

[3580] D. James Greiner and Kevin M. Quinn.

Long live the exit poll. Dædalus, 141(4):9–22, Fall 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/41721977;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00170.

Stimson:2012:MMM

[3581] James A. Stimson. On the meaning & measurement of mood. Dædalus, 141(4):23–34, Fall 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/41721978;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00171.

Erikson:2012:POM

[3582] Robert S. Erikson. Public opinion at the macro level. Dædalus, 141(4):35–49, Fall 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/41721979;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00172.

Druckman:2012:POS

[3583] James N. Druckman and Thomas J. Leeper. Is public opinion stable? Resolving the micro/macro disconnect in studies of public opinion. Dædalus, 141(4):50–68, Fall 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/41721980;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00173.

Greenhouse:2012:POS

[3584] Linda Greenhouse. Public opinion & the Supreme Court: The puzzling case of abortion. Dædalus, 141(4):69–82, Fall 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/41721981;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00174.

Mutz:2012:GDC

[3585] Diana C. Mutz. The great divide: Campaign media in the American mind. Dædalus, 141 (4):83–97, Fall 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/41721982;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00175.

Segura:2012:LPO

[3586] Gary M. Segura. Latino public opinion & realigning the American electorate. Dædalus, 141 (4):98–113, Fall 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/41721983;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00176.

Gibson:2012:BFO

[3587] James L. Gibson. Being free in Obama’s America: Racial differences in perceptions of constraints on political action. Dædalus, 141 (4):114–129, Fall 2012. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://

www.jstor.org/stable/41721984;
http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00177.

Fri:2013:AEF

[3588] Robert W. Fri. The alternative energy future: The scope of the transition. Dædalus, 142 (1):5–7, Winter 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00181.

Harvey:2013:TT

[3589] Hal Harvey, Franklin M. Orr, Jr., and Clara Vondrich. A trillion tons. Dædalus, 142(1):8–25, Winter 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00182.

Krosnick:2013:DAP

[3590] Jon A. Krosnick and Bo MacInnis. Does the American public support legislation to reduce greenhouse gas emissions? Dædalus, 142 (1):26–39, Winter 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00183.

Oreskes:2013:CWC

[3591] Naomi Oreskes and Erik M. Conway. The collapse of Western civilization: a view from the future. Dædalus, 142(1):40–58, Winter 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00184.

Gallagher:2013:WHG

[3592] Kelly Sims Gallagher. Why & how governments support renewable energy. Dædalus, 142(1): 59–77, Winter 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00185.

Dietz:2013:RCB

[3593] Thomas Dietz, Paul C. Stern, and Elke U. Weber. Reducing carbon-based energy consumption through changes in household behavior. Dædalus, 142 (1):78–89, Winter 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00186.

Kasperson:2013:PAN

[3594] Roger E. Kasperson and Bonnie J.

Ram. The public acceptance of new energy technologies. Dædalus, 142 (1):90–96, Winter 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00187.

Keohane:2013:TPE

[3595] Robert O. Keohane and David G.

Victor. The transnational politics of energy. Dædalus, 142(1): 97–109, Winter 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00196.

Burtraw:2013:IBS

[3596] Dallas Burtraw. The institutional blind spot in environmental economics. Dædalus, 142(1): 110–118, Winter 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00188.

Carlson:2013:DDE

[3597] Ann E. Carlson and Robert W.

Fri. Designing a durable energy policy. Dædalus, 142(1): 119–128, Winter 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00189.

Dworkin:2013:RSS

[3598] Michael H. Dworkin, Roman V. Sidortsov, and Benjamin K. Sovacool. Rethinking the scale, structure & scope of U.S. energy institutions. Dædalus, 142(1):129–145, Winter 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00190.

Bierbaum:2013:ECS

[3599] Rosina M. Bierbaum and Pamela A. Matson. Energy in the context of sustainability. Dædalus, 142(1): 146–161, Winter 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00191.

Ansolabehere:2013:SSA

[3600] Stephen Ansolabehere and Robert W. Fri. Social sciences & the alternative energy future. Dædalus, 142 (1):162–171, Winter 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00192.

Berlowitz:2013:F

[3601] Leslie C. Berlowitz. Foreword. Dædalus, 142(2):5, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_e_00197.

Ornstein:2013:I

[3602] Norman J. Ornstein. Introduction. Dædalus, 142(2):6–8, Spring 2013. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00198.

Galston:2013:CGT

[3603] William A. Galston. The common good: Theoretical content, practical utility. Dædalus, 142 (2):9–14, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00199.

Mann:2013:FCG

[3604] Thomas E. Mann and Norman J. Ornstein. Finding the common good in an era of dysfunctional governance. Dædalus, 142(2):15–24, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00200.

Rosen:2013:CJB

[3605] Jeffrey Rosen. Can the judicial branch be a steward in a polarized democracy? Dædalus, 142 (2):25–35, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00201.

Stone:2013:SCC

[3606] Geoffrey R. Stone. The Supreme Court in the 21st Century. Dædalus, 142(2):36–48, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00202.

Hill:2013:SIW

[3607] Andrew A. Hill, Leonard Wong, and Stephen J. Gerras. “Self-Interest Well Understood”: The origins & lessons of public confidence in the military. Dædalus, 142 (2):49–64, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00203.

Jamieson:2013:CFC

[3608] Kathleen Hall Jamieson. The challenges facing civic education in the 21st Century. Dædalus, 142 (2):65–83, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00204.

Edwards:2013:WCG

[3609] Mickey Edwards. What is the common good? The case for transcending partisanship. Dædalus, 142 (2):84–94, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00205.

Leach:2013:CUR

[3610] Jim Leach. Citizens United: Robbing America of its democratic idealism. Dædalus, 142(2):95–101, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00206.

Gomory:2013:AC

[3611] Ralph Gomory and Richard Sylla. The American corporation. Dædalus, 142 (2):102–118, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00207.

Stern:2013:UCE

[3612] Andy Stern. Unions & civic engagement: How the assault on labor endangers civil society. Dædalus, 142 (2):119–138, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00208.

Hall:2013:PNS

[3613] Peter Dobkin Hall. Philanthropy, the nonprofit sector & the democratic dilemma. Dædalus, 142(2): 139–158, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00209.

Schudson:2013:RSJ

[3614] Michael Schudson. Reluctant stewards:

Journalism in a democratic society. Dædalus, 142(2):159–176, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00210.

Tannen:2013:ACA

[3615] Deborah Tannen. The argument culture: Agonism & the common good. Dædalus, 142(2): 177–184, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00211.

Gutmann:2013:VCC

[3616] Amy Gutmann and Dennis Thompson. Valuing compromise for the common good. Dædalus, 142(2): 185–198, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00212.

Gardner:2013:RCC

[3617] Howard Gardner. Reestablishing the commons for the common good. Dædalus, 142(2): 199–208, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00213.

Appiah:2013:DS

[3618] Kwame Anthony Appiah. The democratic spirit. Dædalus, 142 (2):209–221, Spring 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00214.

Massey:2013:AIP

[3619] Douglas S. Massey. America’s immigration policy fiasco: Learning from past mistakes. Dædalus, 142 (3):5–15, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00215.

Foner:2013:IPP

[3620] Nancy Foner. Immigration past & present. Dædalus, 142(3):16–25, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.

mitpressjournals.org/doi/pdfplus/
10.1162/DAED_a_00216.

Hirschman:2013:CIA

[3621] Charles Hirschman. The contributions of immigrants to American culture. Dædalus, 142(3): 26–47, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00217.

Tienda:2013:LAI

[3622] Marta Tienda and Susana M. Sánchez.

Latin American immigration to the United States. Dædalus, 142(3): 48–64, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00218.

Nee:2013:WAA

[3623] Victor Nee and Hilary Holbrow.

Why Asian Americans are becoming mainstream. Dædalus, 142(3): 65–75, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00219.

Singer:2013:CIG

[3624] Audrey Singer. Contemporary immigrant gateways in historical perspective. Dædalus, 142(3):76–91, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.

                                        mitpressjournals.org/doi/pdfplus/
                                        10.1162/DAED_a_00220.

Waters:2013:INY

[3625] Mary C. Waters and Philip Kasinitz. Immigrants in new York city: Reaping the benefits of continuous immigration. Dædalus, 142(3):92– 106, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00221.

Marrow:2013:AND

[3626] Helen B. Marrow. Assimilation in new destinations. Dædalus, 142(3): 107–122, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00222.

Bean:2013:ICL

[3627] Frank D. Bean, Jennifer Lee, and James D. Bachmeier. Immigration & the color line at the beginning of the 21st Century. Dædalus, 142 (3):123–140, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00223.

Rumbaut:2013:ILD

[3628] Rubén G. Rumbaut and Douglas S.

Massey. Immigration & language diversity in the United States. Dædalus, 142 (3):141–154, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00224.

Alba:2013:SDT

[3629] Richard Alba. Schools & the diversity transition. Dædalus, 142(3): 155–169, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00225.

Portes:2013:MED

[3630] Alejandro Portes and Adrienne Celaya.

Modernization for emigration: Determinants & consequences of the brain drain. Dædalus, 142(3): 170–184, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00226.

Jones-Correa:2013:ITP

[3631] Michael Jones-Correa and Els de Graauw.

The illegality trap: The politics of immigration & the lens of illegality. Dædalus, 142(3): 185–198, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00227.

Douglas:2013:CII

[3632] Karen Manges Douglas and Rogelio Sáenz. The criminalization of immigrants & the immigrationindustrial complex. Dædalus, 142 (3):199–227, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00228.

Rodriguez:2013:ICR

[3633] Cristina M. Rodríguez. Immigration, civil rights & the evolution of the people. Dædalus, 142(3): 228–241, Summer 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00229.

Early:2013:IOM

[3634] Gerald Early. Introduction: This is our music. Dædalus, 142 (4):5–10, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00230.

Burke:2013:S

[3635] Patrick Burke. The screamers. Dædalus, 142(4):11–23, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00231.

Yang:2013:YSW

[3636] Mina Yang. Yellow skin, white masks. Dædalus, 142(4):24–37, Fall 2013. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00232.

Robertson:2013:LN

[3637] David Robertson. Listening to the now. Dædalus, 142(4):38–44, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00241.

Hubbs:2013:HTC

[3638] Nadine Hubbs. Homophobia in Twentieth-Century music: The crucible of America’s sound. Dædalus, 142(4):45–50, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00237.

Hisama:2013:RCS

[3639] Ellie M. Hisama. The Ruth Crawford Seeger Sessions. Dædalus, 142(4):51–63, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00236.

Geary:2013:WWF

[3640] Daniel Geary. “The Way I Would Feel About San Quentin”: Johnny cash & the politics of country music. Dædalus, 142(4):64–72, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00234.

Greenspan:2013:HMM

[3641] Charlotte Greenspan. Hollywood as music museum & patron: Bringing various musical styles to a wide audience. Dædalus, 142 (4):73–81, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00235.

Tucker:2013:STT

[3642] Sherrie Tucker. Swing: From time to torque (dance floor democracy at the Hollywood canteen). Dædalus, 142(4):82–97, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00243.

Decker:2013:FMY

[3643] Todd Decker. Fancy meeting you here: Pioneers of the concept album. Dædalus, 142(4):98–108, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00233.

McWhorter:2013:LTN

[3644] John H. McWhorter. Long time, no song: Revisiting fats Waller’s lost Broadway musical. Dædalus, 142(4):109–119, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00238.

Ramsey:2013:NKB

[3645] Guthrie P. Ramsey, Jr. A new kind of blue: The power of suggestion & the pleasure of groove in Robert Glasper’s black radio. Dædalus, 142(4):120–125, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00240.

Radano:2013:SRF

[3646] Ronald Radano. The sound of racial feeling. Dædalus, 142(4): 126–134, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00239.

Teachout:2013:SSE

[3647] Terry Teachout. Satchmo’s shadow: an excerpt from Satchmo at the Waldorf. Dædalus, 142(4):135–139, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00242.

Duke:2013:EBP

[3648] Vernon Duke. Excerpts from Passport to Paris. Dædalus, 142 (4):140–145, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00247.

Kees:2013:GCB

[3649] Weldon Kees. A good chord on a bad piano. Dædalus, 142(4):146, Fall 2013. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00248.

Donoghue:2014:I

[3650] Denis Donoghue. Introduction. Dædalus, 143(1):5–18, Winter 2014. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_e_00249.

Miller:2014:WOH

[3651] J. Hillis Miller. What ought humanists to do? Dædalus, 143 (1):19–34, Winter 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00250.

Oakley:2014:PE

[3652] Francis Oakley. Politics & eternity. Dædalus, 143(1):35–45, Winter 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00251.

Bowlby:2014:HAB

[3653] Rachel Bowlby. “Half Art”: Baudelaire’s Le Peintre de la vie moderne. Dædalus, 143(1):46–53, Winter 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00252.

Beer:2014:EBW

[3654] Gillian Beer. Excerpts from The Waves, by Virginia Woolf. Dædalus, 143(1):54–63, Winter 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00253.

Spacks:2014:PM

[3655] Patricia Meyer Spacks. The power of Middlemarch. Dædalus, 143 (1):64–76, Winter 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00254.

Olney:2014:BNG

[3656] James Olney. Beckett’s “neither” & Giacometti’s Figurine entre deux boîtes qui sont des maisons. Dædalus, 143 (1):77–84, Winter 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00255.

Posnock:2014:IT

[3657] Ross Posnock. I’m not there. Dædalus, 143(1):85–95, Winter 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00256.

Cole:2014:LGM

[3658] Henri Cole. Louise Glück’s “Messengers”. Dædalus, 143(1):96–98, Winter 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.

mitpressjournals.org/doi/pdfplus/
10.1162/DAED_a_00257.

Putnam:2014:DLD

[3659] Michael C. J. Putnam. Dido’s long dying. Dædalus, 143(1): 99–106, Winter 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00258.

Holloway:2014:BAG

[3660] Karla F. C. Holloway. Beloved: America’s grammar book. Dædalus, 143 (1):107–114, Winter 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00259.

Sanders:2014:HBD

[3661] Scott Russell Sanders. Hooks baited with darkness. Dædalus, 143(1): 115–122, Winter 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00260.

Marcus:2014:RRF

[3662] Steven Marcus. On reading & rereading Freud’s introductory lectures on psychoanalysis. Dædalus, 143(1): 123–138, Winter 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00261.

Perry:2014:GPC

[3663] Elizabeth J. Perry. Growing pains: Challenges for a rising China. Dædalus, 143(2):5–13, Spring 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00268.

Naughton:2014:CEC

[3664] Barry Naughton. China’s economy: Complacency, crisis & the challenge of reform. Dædalus, 143 (2):14–25, Spring 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00269.

Davis:2014:DCR

[3665] Deborah S. Davis. Demographic challenges for a rising China. Dædalus, 143(2):26–38, Spring 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00270.

Whyte:2014:SIG

[3666] Martin King Whyte. Soaring income gaps: China in comparative perspective. Dædalus, 143 (2):39–52, Spring 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00271.

Hsiao:2014:CPH

[3667] William C. Hsiao. Correcting past health policy mistakes. Dædalus, 143(2):53–68, Spring 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00272.

Frazier:2014:SSS

[3668] Mark W. Frazier. State schemes or safety nets? China’s push for universal coverage. Dædalus, 143 (2):69–80, Spring 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00273.

Gallagher:2014:CWM

[3669] Mary E. Gallagher. China’s workers movement & the end of the rapid-growth era. Dædalus, 143 (2):81–95, Spring 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00274.

Liebman:2014:LRC

[3670] Benjamin L. Liebman. Legal reform: China’s law-stability paradox. Dædalus, 143(2):96–109, Spring 2014. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00275.

Yang:2014:IAP

[3671] Guobin Yang. Internet activism & the party-state in China. Dædalus, 143 (2):110–123, Spring 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00276.

Lee:2014:SSP

[3672] Ching Kwan Lee. State & social protest. Dædalus, 143(2): 124–134, Spring 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00277.

Weller:2014:PIR

[3673] Robert P. Weller. The politics of increasing religious diversity in China. Dædalus, 143(2): 135–144, Spring 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00278.

Kirby:2014:CCC

[3674] William C. Kirby. The Chinese century? The challenges of higher education. Dædalus, 143 (2):145–156, Spring 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00279.

Wasserstrom:2014:CG

[3675] Jeffrey Wasserstrom. China & globalization. Dædalus, 143(2): 157–169, Spring 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00280.

Fewsmith:2014:LGC

[3676] Joseph Fewsmith and Xiang Gao.

Local governance in China: Incentives & tensions. Dædalus, 143 (2):170–183, Spring 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00281.

Economy:2014:EGC

[3677] Elizabeth Economy. Environmental governance in China: State control to crisis management. Dædalus, 143 (2):184–197, Spring 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00282.

Greenhouse:2014:IIC

[3678] Linda Greenhouse. Introduction: The invention of courts. Dædalus, 143 (3):5–8, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00283.

Resnik:2014:RCD

[3679] Judith Resnik. Reinventing courts as democratic institutions. Dædalus, 143(3):9–27, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00284.

Lippman:2014:SCE

[3680] Jonathan Lippman. State courts: Enabling access. Dædalus, 143(3): 28–36, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00285.

Katzmann:2014:WLR

[3681] Robert A. Katzmann. When legal representation is deficient: The challenge of immigration cases for the courts. Dædalus, 143(3): 37–50, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00286.

Steiker:2014:GPP

[3682] Carol S. Steiker. Gideon’s problematic promises. Dædalus, 143 (3):51–61, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00287.

Simon:2014:ULA

[3683] Jonathan Simon. Uncommon law:

America’s excessive criminal law & our common-law origins. Dædalus, 143 (3):62–72, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00288.

Hensler:2014:JMA

[3684] Deborah R. Hensler. Justice for the masses? Aggregate litigation & its alternatives. Dædalus, 143 (3):73–82, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00289.

Hadfield:2014:IIA

[3685] Gillian K. Hadfield. Innovating to improve access: Changing the way courts regulate legal markets. Dædalus, 143 (3):83–95, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00290.

Graetz:2014:TCR

[3686] Michael J. Graetz. Trusting the courts: Redressing the state court funding crisis. Dædalus, 143(3): 96–104, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00291.

Schauer:2014:OID

[3687] Frederick Schauer. Our informationally disabled courts. Dædalus, 143 (3):105–114, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00292.

Galanter:2014:GCC

[3688] Marc Galanter and Angela M. Frozena. A grin without a cat: The continuing decline & displacement of trials in American courts. Dædalus, 143 (3):115–128, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00293.

Yeazell:2014:CIW

[3689] Stephen C. Yeazell. Courting ignorance: Why we know so little about our most important courts. Dædalus, 143 (3):129–139, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00294.

Silbey:2014:CAP

[3690] Susan S. Silbey. The courts in American public culture. Dædalus, 143 (3):140–156, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00295.

Greene:2014:ACC

[3691] Jamal Greene. (Anti)canonizing courts. Dædalus, 143(3):157–167, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.

mitpressjournals.org/doi/pdfplus/
10.1162/DAED_a_00296.

ORegan:2014:JMS

[3692] Kate O’Regan. Justice & memory: South Africa’s Constitutional Court. Dædalus, 143(3):168–178, Summer 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.

mitpressjournals.org/doi/pdfplus/
10.1162/DAED_a_00297.

Meinwald:2014:I

[3693] Jerrold Meinwald. Introduction. Dædalus, 143(4):5–8, Fall 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_e_00300.

Cummins:2014:PSL

[3694] Christopher C. Cummins. Phosphorus:

From the stars to land & sea. Dædalus, 143(4):9–20, Fall 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00301.

Thomas:2014:FUC

[3695] John Meurig Thomas. Foresight, unpredictability & chance in chemistry & cognate subjects. Dædalus, 143(4):21–30, Fall 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00302.

Wudl:2014:BFF

[3696] Fred Wudl. The bright future of fabulous materials based on carbon. Dædalus, 143(4):31–42, Fall 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00303.

Khosla:2014:CCH

[3697] Chaitan Khosla. The convergence of chemistry & human biology. Dædalus, 143(4):43–48, Fall 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00304.

Houk:2014:UCC

[3698] K. N. Houk and Peng Liu. Using computational chemistry to understand & discover chemical reactions. Dædalus, 143(4):49–66, Fall 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00305.

Ostriker:2014:AUR

[3699] Jeremiah P. Ostriker. From the atom to the universe: Recent astronomical discoveries. Dædalus, 143(4):67–70, Fall 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00306.

Frebel:2014:RCE

[3700] Anna Frebel. Reconstructing the cosmic evolution of the chemical elements. Dædalus, 143(4):71–80, Fall 2014. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00307.

Bakos:2014:E

[3701] Gáspár Áron Bakos. Exoplanets, 2003–2013. Dædalus, 143 (4):81–92, Fall 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00308.

Strauss:2014:MUS

[3702] Michael A. Strauss. Mapping the universe: Surveys of the sky as discovery engines in astronomy. Dædalus, 143(4):93–102, Fall 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00309.

Tremaine:2014:OCQ

[3703] Scott Tremaine. The odd couple:

Quasars & black holes. Dædalus, 143(4):103–113, Fall 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00310.

vanDokkum:2014:FEG

[3704] Pieter van Dokkum. The formation & evolution of galaxies. Dædalus, 143(4):114–124, Fall 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00311.

Spergel:2014:CT

[3705] David N. Spergel. Cosmology today. Dædalus, 143(4):125–133, Fall 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00312.

Gage:2015:NSN

[3706] Fred H. Gage. Neuroscience: The study of the nervous system & its functions. Dædalus, 144 (1):5–9, Winter 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_e_00313.

Wurtz:2015:BMA

[3707] Robert H. Wurtz. Brain mechanisms for active vision. Dædalus, 144(1):10–21, Winter 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00314.

Albright:2015:P

[3708] Thomas D. Albright. Perceiving. Dædalus, 144(1):22–41, Winter 2015. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00315.

Hudspeth:2015:EE

[3709] A. J. Hudspeth. The energetic ear.

Dædalus, 144(1):42–52, Winter 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00316.

Squire:2015:R

[3710] Larry R. Squire and John T. Wixted.

Remembering. Dædalus, 144(1): 53–66, Winter 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00317.

Watson:2015:SMB

[3711] Brendon O. Watson and György Buzsáki. Sleep, memory & brain rhythms. Dædalus, 144(1):67–82, Winter 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.

mitpressjournals.org/doi/pdfplus/
10.1162/DAED_a_00318.

Bizzi:2015:HSQ

[3712] Emilio Bizzi and Robert Ajemian.

A hard scientific quest: Understanding voluntary movements. Dædalus, 144(1):83–95, Winter 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00324.

LeDoux:2015:FWT

[3713] Joseph E. LeDoux. Feelings: What are they & how does the brain make them? Dædalus, 144(1): 96–111, Winter 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00319.

Miller:2015:WMC

[3714] Earl K. Miller and Timothy J. Buschman. Working memory capacity: Limits on the bandwidth of cognition. Dædalus, 144(1): 112–122, Winter 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00320.

Sejnowski:2015:C

[3715] Terrence J. Sejnowski. Consciousness. Dædalus, 144(1):123–132, Winter 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00321.

Rowe:2015:SAS

[3716] John W. Rowe. Successful aging of societies. Dædalus, 144 (2):5–12, Spring 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00325.

Olshansky:2015:DTA

[3717] S. Jay Olshansky. The demographic transformation of America. Dædalus, 144(2):13–19, Spring 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00326.

Hummer:2015:HOA

[3718] Robert A. Hummer and Mark D.

Hayward. Hispanic older adult health & longevity in the United States: Current patterns & concerns for the future. Dædalus, 144 (2):20–30, Spring 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00327.

Furstenberg:2015:FIR

[3719] Frank F. Furstenberg, Caroline Sten Hartnett, Martin Kohli, and Julie M. Zissimopoulos. The future of intergenerational relations in aging societies. Dædalus, 144(2): 31–40, Spring 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00328.

Berkman:2015:LFP

[3720] Lisa F. Berkman, Axel Boersch-Supan, and Mauricio Avendano. Laborforce participation, policies & practices in an aging America: Adaptation essential for a healthy & resilient population. Dædalus, 144 (2):41–54, Spring 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00329.

Carr:2015:PEA

[3721] Dawn C. Carr, Linda P. Fried, and John W. Rowe. Productivity & engagement in an aging America: The role of volunteerism. Dædalus, 144 (2):55–67, Spring 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00330.

Olshansky:2015:RSS

[3722] S. Jay Olshansky, Dana P. Goldman, and John W. Rowe. Resetting social security. Dædalus, 144 (2):68–79, Spring 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00331.

Bloom:2015:GPA

[3723] David E. Bloom, David Canning, and Alyssa Lubet. Global population aging: Facts, challenges, solutions & perspectives. Dædalus, 144 (2):80–92, Spring 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00332.

Zissimopoulos:2015:ISS

[3724] Julie M. Zissimopoulos, Dana P. Goldman, S. Jay Olshansky, John Rother, and John W. Rowe. Individual & social strategies to mitigate the risks & expand opportunities of an aging America. Dædalus, 144(2):93–102, Spring 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00333.

Michalak:2015:I

[3725] Anna M. Michalak and Christopher B. Field. Introduction. Dædalus, 144 (3):5–6, Summer 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_x_00336.

Field:2015:WCE

[3726] Christopher B. Field and Anna M.

Michalak. Water, climate, energy, food: Inseparable & indispensable. Dædalus, 144(3):7–17, Summer 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00337.

Witzel:2015:WM

[3727] Michael Witzel. Water in mythology.

Dædalus, 144(3):18–26, Summer 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00338.

Briscoe:2015:WSC

[3728] John Briscoe. Water security in a changing world. Dædalus, 144 (3):27–34, Summer 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00339.

Rissman:2015:PNP

[3729] Adena R. Rissman and Stephen R. Carpenter. Progress on nonpoint pollution: Barriers & opportunities. Dædalus, 144(3):35–47, Summer 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00340.

Schnoor:2015:WU

[3730] Jerald L. Schnoor. Water unsustainability. Dædalus, 144(3): 48–58, Summer 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00341.

Jacobs:2015:AWS

[3731] Katharine L. Jacobs and Lester Snow. Adaptation in the water sector: Science & institutions. Dædalus, 144 (3):59–71, Summer 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00342.

Luthy:2015:UWS

[3732] Richard G. Luthy and David L. Sedlak. Urban water-supply reinvention. Dædalus, 144(3):72–82, Summer 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.mitpressjournals.org/
                                        doi/pdfplus/10.1162/DAED_a_00343.

Anderson:2015:DMD

[3733] Terry L. Anderson. Dynamic markets for dynamic environments: The case for water marketing. Dædalus, 144(3):83–93, Summer 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00344.

Vorosmarty:2015:ITR

[3734] Charles J. Vörösmarty, Michel Meybeck, and Christopher L. Pastore. Impair-then-repair: a brief history & global-scale hypothesis regarding human–water interactions in the Anthropocene. Dædalus, 144(3): 94–109, Summer 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00345.

Tilman:2015:FHF

[3735] David Tilman. Food & health of a full Earth. Dædalus, 144(4):5–7, Fall 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_e_00349.

Tilman:2015:FAE

[3736] David Tilman and Michael Clark.

Food, agriculture & the environment: Can we feed the world & save the Earth? Dædalus, 144(4):8–23, Fall 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00350.

Bertini:2015:IW

[3737] Catherine Bertini. Invisible women.

Dædalus, 144(4):24–30, Fall 2015. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00351.

Jahn:2015:FHE

[3738] Jaquelyn L. Jahn, Meir J. Stampfer, and Walter C. Willett. Food, health & the environment: a global grand challenge & some solutions. Dædalus, 144(4):31–44, Fall 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00352.

Mueller:2015:CYG

[3739] Nathaniel D. Mueller and Seth Binder. Closing yield gaps: Consequences for the global food supply, environmental quality & food security. Dædalus, 144(4):45–56, Fall 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00353.

Balmford:2015:LFL

[3740] Andrew Balmford, Rhys Green, and Ben Phalan. Land for food & land for nature? Dædalus, 144 (4):57–75, Fall 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00354.

Robertson:2015:SA

[3741] G. Philip Robertson. A sustainable agriculture? Dædalus, 144 (4):76–89, Fall 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00355.

Henning:2015:EFF

[3742] Brian G. Henning. The ethics of food, fuel & feed. Dædalus, 144(4):90–98, Fall 2015. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00356.

Benkler:2016:I

[3743] Yochai Benkler and David D. Clark.

Introduction. Dædalus, 145(1):5–8, Winter 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.

mitpressjournals.org/doi/pdfplus/
10.1162/DAED_x_00360.

Clark:2016:CI

[3744] David D. Clark. The contingent Internet. Dædalus, 145(1): 9–17, Winter 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00361.

Benkler:2016:DFD

[3745] Yochai Benkler. Degrees of freedom, dimensions of power. Dædalus, 145 (1):18–32, Winter 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00362.

Kirstein:2016:END

[3746] Peter T. Kirstein. Edge networks & devices for the Internet of Things. Dædalus, 145 (1):33–42, Winter 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00363.

Estrin:2016:ROD

[3747] Deborah Estrin and Ari Juels. Reassembling our digital selves. Dædalus, 145(1):43–53, Winter 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00364.

Landau:2016:CPS

[3748] Susan Landau. Choices: Privacy & surveillance in a once & future Internet. Dædalus, 145(1): 54–64, Winter 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00365.

Tufekci:2016:PBC

[3749] Zeynep Tufekci. As the pirates become CEOs: The closing of the open Internet. Dædalus, 145(1): 65–78, Winter 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00366.

Palfrey:2016:DCL

[3750] John Palfrey. Design choices for libraries in the digital-plus era. Dædalus, 145(1):79–86, Winter 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00367.

Santirocco:2016:IRG

[3751] Matthew S. Santirocco. Introduction: Reassessing Greece & Rome. Dædalus, 145(2):5–19, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00371.

Holmes:2016:TCP

[3752] Brooke Holmes. Tragedy in the crosshairs of the present. Dædalus, 145(2):20–29, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00372.

Bartsch:2016:RLT

[3753] Shadi Bartsch. Roman literature:

Translation, metaphor & empire. Dædalus, 145(2):30–39, Spring 2016. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00373.

Hadas:2016:NBP

[3754] Rachel Hadas. The new “Brothers Poem” by Sappho. Dædalus, 145(2):40, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00384.

Greenwood:2016:RSC

[3755] Emily Greenwood. Reception studies: The cultural mobility of classics. Dædalus, 145(2):41–49, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00374.

Alexander:2016:THI

[3756] Caroline Alexander. On translating Homer’s Iliad. Dædalus, 145 (2):50–58, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00375.

Mitsis:2016:PCP

[3757] Phillip Mitsis. Philosophy & its classical past. Dædalus, 145(2): 59–67, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00376.

Putnam:2016:EC

[3758] Michael C. J. Putnam. Explicating Catullus. Dædalus, 145 (2):68, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00386.

Platt:2016:MCA

[3759] Verity Platt. The matter of classical art history. Dædalus, 145 (2):69–78, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00377.

Bagnall:2016:MAD

[3760] Roger S. Bagnall. Materializing ancient documents. Dædalus, 145 (2):79–87, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00378.

Chaniotis:2016:MCI

[3761] Angelos Chaniotis. Memory, commemoration & identity in an ancient city:

The case of Aphrodisias. Dædalus, 145(2):88–100, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00379.

Harper:2016:EFR

[3762] Kyle Harper. The environmental fall of the Roman Empire. Dædalus, 145 (2):101–111, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00380.

Wiener:2016:SSA

[3763] Malcolm H. Wiener. The scientific study of antiquity. Dædalus, 145(2):112, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00385.

Morris:2016:WAH

[3764] Ian Morris and Walter Scheidel. What is ancient history? Dædalus, 145 (2):113–121, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00381.

Struck:2016:CCP

[3765] Peter T. Struck. Classics: Curriculum & profession. Dædalus, 145 (2):122–126, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00382.

Crane:2016:GRS

[3766] Gregory Crane. Greco–Roman studies in a digital age. Dædalus, 145 (2):127–133, Spring 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00383.

Brown:2016:I

[3767] Archie Brown. Introduction. Dædalus, 145(3):5–7, Summer 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00392.

Keohane:2016:LED

[3768] Nannerl O. Keohane. Leadership, equality & democracy. Dædalus, 145 (3):8–20, Summer 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00393.

Haslam:2016:RPL

[3769] S. Alexander Haslam and Stephen D. Reicher. Rethinking the psychology of leadership: From personal identity to social identity. Dædalus, 145 (3):21–34, Summer 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00394.

Posner:2016:PLS

[3770] Eric A. Posner. Presidential leadership & the separation of powers. Dædalus, 145(3):35–43, Summer 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00395.

Swers:2016:WLL

[3771] Michele L. Swers. Women & legislative leadership in the U.S. Congress: Representing women’s interests in partisan times. Dædalus, 145(3): 44–56, Summer 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00396.

Elgie:2016:VPL

[3772] Robert Elgie. Varieties of presidentialism & of leadership outcomes. Dædalus, 145(3):57–68, Summer 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00397.

Huskey:2016:ALP

[3773] Eugene Huskey. Authoritarian leadership in the post-communist world. Dædalus, 145(3):69–82, Summer 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00398.

Kellerman:2016:LSP

[3774] Barbara Kellerman. Leadership — it’s a system, not a person! Dædalus, 145(3):83–94, Summer 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00399.

Stepan:2016:MCC

[3775] Alfred Stepan. Multiple but complementary, not conflictual, leaderships: The Tunisian democratic transition in comparative perspective. Dædalus, 145 (3):95–108, Summer 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00400.

Brown:2016:AFC

[3776] Archie Brown. Against the Führerprinzip: For collective leadership. Dædalus, 145 (3):109–123, Summer 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00401.

King:2016:FLP

[3777] Anthony King. In favor of “leader proofing”. Dædalus, 145(3):124– 137, Summer 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00402.

Sagan:2016:ETW

[3778] Scott D. Sagan. Ethics, technology & war. Dædalus, 145 (4):6–11, Fall 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_e_00407.

Walzer:2016:JUT

[3779] Michael Walzer. Just & unjust targeted killing & drone warfare. Dædalus, 145(4):12–24, Fall 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00408.

Horowitz:2016:EMR

[3780] Michael C. Horowitz. The ethics & morality of robotic warfare: Assessing the debate over autonomous weapons. Dædalus, 145(4):25–36, Fall 2016. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/pdfplus/10.1162/DAED_a_00409.

Fidler:2016:JUW

[3781] David P. Fidler. Just & unjust war, uses of force & coercion: an ethical inquiry with cyber illustrations. Dædalus, 145(4):37–49, Fall 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00410.

Kehler:2016:NWN

[3782] C. Robert Kehler. Nuclear weapons & nuclear use. Dædalus, 145 (4):50–61, Fall 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00411.

Lewis:2016:NNP

[3783] Jeffrey G. Lewis and Scott D. Sagan.

The nuclear necessity principle: Making U.S. targeting policy conform with ethics & the laws of war. Dædalus, 145(4):62–74, Fall 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00412.

Welsh:2016:RPA

[3784] Jennifer M. Welsh. The responsibility to protect after Libya & Syria. Dædalus, 145(4):75–87, Fall 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00413.

Axworthy:2016:NTP

[3785] Lloyd Axworthy and A. Walter Dorn. New technology for peace & protection: Expanding the r2p toolbox. Dædalus, 145(4):88–100, Fall 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00414.

Leaning:2016:PLR

[3786] Jennifer Leaning. The path to last resort: The role of early warning & early action. Dædalus, 145 (4):101–112, Fall 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00415.

Krause:2016:ACP

[3787] Keith Krause. From armed conflict to political violence: Mapping & explaining conflict trends. Dædalus, 145(4):113–126, Fall 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00416.

Valentino:2016:MCC

[3788] Benjamin Valentino. Moral character or character of war? American public opinion on the targeting of civilians in times of war. Dædalus, 145(4):127–138, Fall 2016. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00417.

Sagan:2017:CRW

[3789] Scott D. Sagan. The changing rules of war. Dædalus, 146 (1):6–10, Winter 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00418.

Savarese:2017:SEE

[3790] Laura Ford Savarese and John Fabian Witt. Strategy & entailments: The enduring role of law in the U.S. Armed Forces. Dædalus, 146 (1):11–23, Winter 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00419.

Sagan:2017:FBR

[3791] Scott D. Sagan. The face of battle without the rules of war: Lessons from Red Horse & the Battle of the Little Bighorn. Dædalus, 146 (1):24–43, Winter 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00420.

Felter:2017:LCC

[3792] Joseph H. Felter and Jacob N. Shapiro. Limiting civilian casualties as part of a winning strategy: The case of courageous restraint. Dædalus, 146 (1):44–58, Winter 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00421.

Weiner:2017:JWT

[3793] Allen S. Weiner. Just war theory & the conduct of asymmetric warfare. Dædalus, 146(1): 59–70, Winter 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00422.

Fazal:2017:RWA

[3794] Tanisha M. Fazal. Rebellion, war aims & the laws of war. Dædalus, 146(1):71–82, Winter 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00423.

Martins:2017:SHJ

[3795] Mark S. Martins and Jacob Bronsther. Stay the hand of justice? Evaluating claims that war crimes trials do more harm than good. Dædalus, 146(1):83–99, Winter 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        pdfplus/10.1162/DAED_a_00424.

Vinjamuri:2017:DPN

[3796] Leslie Vinjamuri. The distant promise of a negotiated justice. Dædalus, 146 (1):100–112, Winter 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00425.

Lazar:2017:ERC

[3797] Seth Lazar. Evaluating the revisionist critique of just war theory. Dædalus, 146(1):113–124, Winter 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00426.

Chayes:2017:WCN

[3798] Antonia Chayes and Janne E. Nolan.

What comes next. Dædalus, 146 (1):125–138, Winter 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00427.

Wise:2017:ECC

[3799] Paul H. Wise. The epidemiologic challenge to the conduct of just war: Confronting indirect civilian casualties of war. Dædalus, 146(1): 139–154, Winter 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
pdfplus/10.1162/DAED_a_00428.

Colton:2017:I

[3800] Timothy J. Colton. Introduction. Dædalus, 146(2):5–7, Spring 2017. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.mitpressjournals.org/
                                        doi/full/10.1162/DAED_a_00429.

Colton:2017:PP

[3801] Timothy J. Colton. Paradoxes of Putinism. Dædalus, 146(2):8–18, Spring 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.

                                        mitpressjournals.org/doi/full/10.
                                        1162/DAED_a_00430.

Bunce:2017:PCR

[3802] Valerie Bunce. The prospects for a color revolution in Russia. Dædalus, 146(2):19–29, Spring 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/DAED_a_00431.

Hale:2017:RPP

[3803] Henry E. Hale. Russian patronal politics beyond Putin. Dædalus, 146 (2):30–40, Spring 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/DAED_a_00432.

Hill:2017:NMP

[3804] Fiona Hill. The next Mr. Putin? The question of succession. Dædalus, 146(2):41–52, Spring 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/DAED_a_00433.

Taylor:2017:RSP

[3805] Brian D. Taylor. The Russian Siloviki & political change. Dædalus, 146 (2):53–63, Spring 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00434.

Popova:2017:PSR

[3806] Maria Popova. Putin-style “rule of law” & the prospects for change. Dædalus, 146(2):64–75, Spring 2017. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/full/10.1162/DAED_a_00435.

Chebankova:2017:III

[3807] Elena Chebankova. Ideas, ideology & intellectuals in search of Russia’s political future. Dædalus, 146 (2):76–88, Spring 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00436.

Laruelle:2017:NFC

[3808] Marlene Laruelle. Is nationalism a force for change in Russia? Dædalus, 146(2):89–100, Spring 2017. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/full/10.1162/DAED_a_00437.

Markus:2017:ASW

[3809] Stanislav Markus. The atlas that has not shrugged: Why Russia’s oligarchs are an unlikely force for change. Dædalus, 146(2):101–112, Spring 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/full/10.1162/DAED_a_00438.

Greene:2017:BBH

[3810] Samuel A. Greene. From boom to bust: Hardship, mobilization & Russia’s social contract. Dædalus, 146 (2):113–127, Spring 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/DAED_a_00439.

Darden:2017:RRE

[3811] Keith A. Darden. Russian revanche: External threats & regime reactions. Dædalus, 146(2):128–141, Spring 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.mitpressjournals.org/
                                        doi/full/10.1162/DAED_a_00440.

Breslauer:2017:IF

[3812] George W. Breslauer. Images of the future. Dædalus, 146(2): 142–150, Spring 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/DAED_a_00441.

Fishkin:2017:I

[3813] James S. Fishkin and Jane Mansbridge. Introduction. Dædalus, 146 (3):6–13, Summer 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/DAED_x_00442.

Offe:2017:RVI

[3814] Claus Offe. Referendum vs. institutionalized deliberation: What democratic theorists can learn from the 2016 Brexit decision. Dædalus, 146 (3):14–27, Summer 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00443.

Curato:2017:TKF

[3815] Nicole Curato, John S. Dryzek, Selen A. Ercan, Carolyn M. Hendriks, and Simon Niemeyer. Twelve key findings in deliberative democracy research. Dædalus, 146(3): 28–38, Summer 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00444.

Manin:2017:PDA

[3816] Bernard Manin. Political deliberation & the adversarial principle. Dædalus, 146(3):39–50, Summer 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/full/10.1162/DAED_a_00445.

Landemore:2017:DDO

[3817] Hélèns Landemore. Deliberative democracy as open, not (just) representative democracy. Dædalus, 146 (3):51–63, Summer 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00446.

Lupia:2017:IAR

[3818] Arthur Lupia and Anne Norton.

Inequality is always in the room: Language & power in deliberative democracy. Dædalus, 146(3): 64–76, Summer 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/DAED_a_00447.

Shapiro:2017:CRD

[3819] Ian Shapiro. Collusion in restraint of democracy: Against political deliberation. Dædalus, 146 (3):77–84, Summer 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/DAED_a_00448.

Lafont:2017:CDD

[3820] Cristina Lafont. Can democracy be deliberative & participatory? The democratic case for political uses of mini-publics. Dædalus, 146(3): 85–105, Summer 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/DAED_a_00449.

Bachtiger:2017:DCN

[3821] André Bächtiger and Simon Beste. Deliberative citizens, (non)deliberative politicians: A rejoinder. Dædalus, 146(3):106–118, Summer 2017. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.mitpressjournals.org/
                                        doi/full/10.1162/DAED_a_00450.

Siu:2017:DCI

[3822] Alice Siu. Deliberation & the challenge of inequality. Dædalus, 146 (3):119–128, Summer 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/DAED_a_00451.

Sunstein:2017:DDT

[3823] Cass R. Sunstein. Deliberative democracy in the trenches. Dædalus, 146 (3):129–139, Summer 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00452.

Fishkin:2017:ADD

[3824] James S. Fishkin, Roy William Mayega, Lynn Atuyambe, Nathan Tumuhamye, Julius Ssentongo, Alice Siu, and William Bazeyo. Applying deliberative democracy in Africa: Uganda’s first deliberative polls. Dædalus, 146 (3):140–154, Summer 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00453.

He:2017:ADC

[3825] Baogang He and Mark E. Warren. Authoritarian deliberation in China. Dædalus, 146(3):155–166, Summer 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/full/10.1162/DAED_a_00454.

Eikenberry:2017:I

[3826] Karl Eikenberry and Stephen D.

Krasner. Introduction. Dædalus, 146(4):6–17, Fall 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_e_00455.

Fearon:2017:CWC

[3827] James D. Fearon. Civil war & the current international system. Dædalus, 146(4):18–32, Fall 2017. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/full/10.1162/DAED_a_00456.

Jones:2017:CWP

[3828] Bruce D. Jones and Stephen John Stedman. Civil wars & the post-Cold War international order. Dædalus, 146(4):33–44, Fall 2017. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/full/10.1162/DAED_a_00457.

Patrick:2017:CWT

[3829] Stewart Patrick. Civil wars & transnational threats: Mapping the terrain, assessing the links. Dædalus, 146(4):45–58, Fall 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00458.

Crenshaw:2017:TJC

[3830] Martha Crenshaw. Transnational jihadism & civil wars. Dædalus, 146(4):59–70, Fall 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00459.

Wise:2017:CWG

[3831] Paul H. Wise and Michele Barry. Civil war & the global threat of pandemics. Dædalus, 146(4):71–84, Fall 2017. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/full/10.1162/DAED_a_00460.

Lischer:2017:GRC

[3832] Sarah Kenyon Lischer. The global refugee crisis: Regional destabilization & humanitarian protection. Dædalus, 146(4):85–97, Fall 2017. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.mitpressjournals.org/
doi/full/10.1162/DAED_a_00461.

Felbab-Brown:2017:OCI

[3833] Vanda Felbab-Brown. Organized crime, illicit economies, civil violence & international order: More complex than you think. Dædalus, 146(4):98–111, Fall 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00462.

Spruyt:2017:CWC

[3834] Hendrik Spruyt. Civil wars as challenges to the modern international system. Dædalus, 146 (4):112–125, Fall 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00463.

Biddle:2017:BSF

[3835] Stephen Biddle. Building security forces & stabilizing nations: The problem of agency. Dædalus, 146 (4):126–138, Fall 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00464.

Reno:2017:FSA

[3836] William Reno. Fictional states & atomized public spheres: A nonwestern approach to fragility. Dædalus, 146(4):139–151, Fall 2017. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.mitpressjournals.org/
                                        doi/full/10.1162/DAED_a_00465.

Matanock:2017:CPP

[3837] Aila M. Matanock and Miguel García-Sánchez. The Colombian paradox: Peace processes, elite divisions & popular plebiscites. Dædalus, 146 (4):152–166, Fall 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/DAED_a_00466.

Posen:2017:CWS

[3838] Barry R. Posen. Civil wars & the structure of world power. Dædalus, 146(4):167–179, Fall 2017. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/DAED_a_00467.

Anonymous:2017:CN

[3839] Anonymous. Correction notice. Dædalus, 146(4):168, Fall 2017. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        http://www.mitpressjournals.org/
                                        doi/full/10.1162/DAED_e_00468.

Eikenberry:2018:I

[3840] Karl Eikenberry and Stephen D. Krasner. Introduction. Dædalus, 147(1):8, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_e_00469.

Fukuyama:2018:LEC

[3841] Francis Fukuyama. The last English Civil War. Dædalus, 147 (1):15, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00470.

Fazal:2018:RRS

[3842] Tanisha M. Fazal. Religionist rebels & the sovereignty of the divine. Dædalus, 147(1):25, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00471.

Kalyvas:2018:JRC

[3843] Stathis N. Kalyvas. Jihadi rebels in civil war. Dædalus, 147(1):36, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi/full/10.
1162/DAED_a_00472.

Heydemann:2018:CWE

[3844] Steven Heydemann. Civil war, economic governance & state reconstruction in the Arab Middle East. Dædalus, 147(1):48, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00473.

Call:2018:PA

[3845] Charles T. Call and Susanna P. Campbell. Is prevention the answer? Dædalus, 147(1):64, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00474.

Ganguly:2018:ESL

[3846] Sumit Ganguly. Ending the Sri Lankan civil war. Dædalus, 147 (1):78, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00475.

Lockhart:2018:SSE

[3847] Clare Lockhart. Sovereignty strategies: Enhancing core governance functions as a postconflict & conflictprevention measure. Dædalus, 147 (1):90, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00476.

Risse:2018:LSD

[3848] Thomas Risse and Eric Stollenwerk. Limited statehood does not equal civil war. Dædalus, 147 (1):104, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00477.

Borzel:2018:BGE

[3849] Tanja A. Börzel and Sonja Grimm. Building good (enough) governance in postconflict societies & areas of limited statehood: The European Union & the Western Balkans. Dædalus, 147(1):116, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00478.

Mesfin:2018:PLF

[3850] Seyoum Mesfin and Abdeta Dribssa Beyene. The practicalities of living with failed states. Dædalus, 147(1):128, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00479.

Doucet:2018:SCE

[3851] Lyse Doucet. Syria & the CNN effect: What role does the media play in policy-making? Dædalus, 147(1):141, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00480.

Lindborg:2018:DAB

[3852] Nancy E. Lindborg and J. Joseph Hewitt. In defense of ambition: Building peaceful & inclusive societies in a world on fire. Dædalus, 147(1):158, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00481.

Gowan:2018:IRT

[3853] Richard Gowan and Stephen John Stedman. The international regime for treating civil war, 1988–2017. Dædalus, 147(1):171, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00482.

Guehenno:2018:UNC

[3854] Jean-Marie Guéhenno. The United Nations & civil wars. Dædalus, 147(1):185, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00483.

Krasner:2018:C

[3855] Stephen D. Krasner and Karl Eikenberry. Conclusion. Dædalus, 147 (1):197, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00484.

Deloria:2018:UFI

[3856] Philip J. Deloria, K. Tsianina Lomawaima, Bryan McKinley Jones Brayboy, Mark N. Trahant, Loren Ghiglione, Douglas Medin, and Ned Blackhawk. Unfolding futures: Indigenous ways of knowing for the Twenty-First Century. Dædalus, 147(2):6, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00485.

Stark:2018:NGF

[3857] Heidi Kiiwetinepinesiik Stark and Kekek Jason Stark. Nenabozho goes fishing: a sovereignty story. Dædalus, 147(2):17, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00486.

Ouden:2018:RAI

[3858] Amy E. Den Ouden. Recognition, antiracism & indigenous futures: a view from Connecticut. Dædalus, 147(2):27, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00487.

Worl:2018:ACO

[3859] Rosita Kaaháni Worl and Heather Kendall-Miller. Alaska’s conflicting objectives. Dædalus, 147 (2):39, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00488.

Goodyear-Kaopua:2018:MAI

[3860] Noelani Goodyear-Ka‘opua and Bryan Kamaoli Kuwada. Making ‘aha: Independent Hawaiian pasts, presents & futures. Dædalus, 147(2):49, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi/full/10.
1162/DAED_a_00489.

Garrison:2018:GAT

[3861] Nanibaa’ A. Garrison. Genetic ancestry testing with tribes: Ethics, identity & health implications. Dædalus, 147(2):60, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00490.

Eason:2018:RRI

[3862] Arianne E. Eason, Laura M. Brady, and Stephanie A. Fryberg. Reclaiming representations & interrupting the cycle of bias against Native Americans. Dædalus, 147 (2):70, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00491.

Brayboy:2018:WDM

[3863] Bryan McKinley Jones Brayboy and K. Tsianina Lomawaima. Why don’t more Indians do better in school? The battle between U.S. schooling & American Indian/Alaska Native education. Dædalus, 147 (2):82, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00492.

Bull:2018:RHE

[3864] Cheryl Crazy Bull and Justin Guillory. Revolution in higher education: Identity & cultural beliefs inspire tribal colleges & universities. Dædalus, 147(2):95, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00493.

Deloria:2018:NWI

[3865] Philip J. Deloria. The new world of the indigenous museum. Dædalus, 147(2):106, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00494.

Trahant:2018:SIH

[3866] Mark N. Trahant. The story of Indian health is complicated by history, shortages & bouts of excellence. Dædalus, 147(2):116, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00495.

Sandefur:2018:IL

[3867] Gary Sandefur and Philip J. Deloria. Indigenous leadership. Dædalus, 147(2):124, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00496.

Whyte:2018:CIR

[3868] Kyle Whyte. Critical investigations of resilience: a brief introduction to indigenous environmental studies & sciences. Dædalus, 147(2):136, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00497.

Bang:2018:IIP

[3869] Megan Bang, Ananda Marin, and Douglas Medin. If indigenous peoples stand with the sciences, will scientists stand with us? Dædalus, 147(2):148, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/DAED_a_00498.

McCarty:2018:HOL

[3870] Teresa L. McCarty, Sheilah E. Nicholas, Kari A. B. Chew, Natalie G. Diaz, Wesley Y. Leonard, and Louellyn White. Hear our languages, hear our voices: Storywork as theory and praxis in indigenouslanguage reclamation. Dædalus, 147(2):160, Spring 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/DAED_a_00499.

Rotberg:2018:AAP

[3871] Robert I. Rotberg. Accomplishing anticorruption: Propositions & methods. Dædalus, 147(3):5, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

                                        mitpressjournals.org/doi/full/10.
                                        1162/daed_a_00513.

Nash:2018:BT

[3872] Thomas Nash. Boss Tweed. Dædalus, 147(3):19, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_00515.

Mungiu-Pippidi:2018:SSC

[3873] Alina Mungiu-Pippidi. Seven steps to control of corruption: The road map. Dædalus, 147(3):20, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

                                        mitpressjournals.org/doi/full/10.
                                        1162/daed_a_00500.

Rothstein:2018:FSC

[3874] Bo Rothstein. Fighting systemic corruption: The indirect strategy. Dædalus, 147(3):35, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi/full/10.
1162/daed_a_00501.

Johnston:2018:RRR

[3875] Michael Johnston. Reforming reform: Revising the anticorruption playbook. Dædalus, 147 (3):50, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00502.

Taylor:2018:GAF

[3876] Matthew M. Taylor. Getting to accountability: a framework for planning & implementing anticorruption strategies. Dædalus, 147 (3):63, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00503.

Heywood:2018:CCT

[3877] Paul M. Heywood. Combating corruption in the twenty-first century: New approaches. Dædalus, 147(3):83, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00504.

Rose-Ackerman:2018:CP

[3878] Susan Rose-Ackerman. Corruption & purity. Dædalus, 147 (3):98, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_00505.

Teachout:2018:PMC

[3879] Zephyr Teachout. The problem of monopolies & corporate public corruption. Dædalus, 147(3): 111, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_00514.

Shelley:2018:CIT

[3880] Louise I. Shelley. Corruption & illicit trade. Dædalus, 147 (3):127, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_00506.

Wolf:2018:WNI

[3881] Mark L. Wolf. The world needs an International Anti-Corruption Court. Dædalus, 147(3):144, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

                                        mitpressjournals.org/doi/full/10.
                                        1162/daed_a_00507.

Moro:2018:PSC

[3882] Sérgio Fernando Moro. Preventing systemic corruption in Brazil. Dædalus, 147(3):157, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00508.

Bracking:2018:CSC

[3883] Sarah Bracking. Corruption & state capture: What can citizens do? Dædalus, 147(3): 169, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00509.

Suberu:2018:SAA

[3884] Rotimi T. Suberu. Strategies for advancing anticorruption reform in Nigeria. Dædalus, 147 (3):184, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00510.

Quah:2018:CCA

[3885] Jon S. T. Quah. Combating corruption in Asian countries: Learning from success & failure. Dædalus, 147(3):202, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00511.

Pei:2018:HFC

[3886] Minxin Pei. How not to fight corruption: Lessons from China. Dædalus, 147(3):216, Summer 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00512.

Diamond:2018:I

[3887] Shari Seidman Diamond and Richard O. Lempert. Introduction. Dædalus, 147(4):5, Fall 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_x_00516.

Jasanoff:2018:SCS

[3888] Sheila Jasanoff. Science, common sense & judicial power in U.S. courts. Dædalus, 147(4):15, Fall 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_00517.

Greenhouse:2018:SCS

[3889] Linda Greenhouse. The Supreme Court & science: a case in point. Dædalus, 147(4):28, Fall 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_00518.

Diamond:2018:WLC

[3890] Shari Seidman Diamond and Richard O. Lempert. When law calls, does science answer? A survey of distinguished scientists & engineers. Dædalus, 147(4):41, Fall 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_00519.

Lobel:2018:LNC

[3891] Jules Lobel and Huda Akil. Law & neuroscience: The case of solitary confinement. Dædalus, 147(4): 61, Fall 2018. CODEN DAEDAU.

ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi/full/10.
1162/daed_a_00520.

Eisenberg:2018:UFA

[3892] Rebecca S. Eisenberg and Robert Cook-Deegan. Universities: The fallen angels of Bayh–Dole? Dædalus, 147(4):76, Fall 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00521.

Rakoff:2018:IIE

[3893] Jed S. Rakoff and Elizabeth F. Loftus. The intractability of inaccurate eyewitness identification. Dædalus, 147(4):90, Fall 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00522.

Mnookin:2018:UFF

[3894] Jennifer L. Mnookin. The uncertain future of forensic science. Dædalus, 147(4):99, Fall 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00523.

Kadane:2018:CUR

[3895] Joseph B. Kadane and Jonathan J.

Koehler. Certainty & uncertainty in reporting fingerprint evidence. Dædalus, 147(4):119, Fall 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00524.

Gertner:2018:ATA

[3896] Nancy Gertner and Joseph Sanders. Alternatives to traditional adversary methods of presenting scientific expertise in the legal system. Dædalus, 147(4):135, Fall 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_00525.

Rubinfeld:2018:SES

[3897] Daniel L. Rubinfeld and Joe S. Cecil. Scientists as experts serving the court. Dædalus, 147(4): 152, Fall 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

                                        mitpressjournals.org/doi/full/10.
                                        1162/daed_a_00526.

Hans:2018:IJJ

[3898] Valerie P. Hans and Michael J. Saks. Improving judge & jury evaluation of scientific evidence. Dædalus, 147(4):164, Fall 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_00527.

Baltimore:2018:BSL

[3899] David Baltimore, David S. Tatel, and Anne-Marie Mazza. Bridging the science–law divide. Dædalus, 147(4):181, Fall 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_00528.

Levi:2018:I

[3900] John G. Levi and David M. Rubenstein. Introduction. Dædalus, 148(1):7, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_e_00529.

Frank:2018:HRI

[3901] Robert H. Frank. How rising income inequality threatens access to the legal system. Dædalus, 148 (1):10, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00530.

Caplan:2018:IJP

[3902] Lincoln Caplan. The invisible justice problem. Dædalus, 148 (1):19, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00531.

Levi:2018:RRL

[3903] David F. Levi, Dana Remus, and Abigail Frisch. Reclaiming the role of lawyers as community connectors. Dædalus, 148(1):30, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00532.

Hadfield:2018:MMM

[3904] Gillian K. Hadfield. More markets, more justice. Dædalus, 148 (1):37, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_00533.

Sandefur:2018:AW

[3905] Rebecca L. Sandefur. Access to what? Dædalus, 148(1):49, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

                                        mitpressjournals.org/doi/full/10.
                                        1162/daed_a_00534.

Brito:2018:RCC

[3906] Tonya L. Brito. The right to civil counsel. Dædalus, 148(1):56, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

                                        mitpressjournals.org/doi/full/10.
                                        1162/daed_a_00535.

Greiner:2018:NLE

[3907] D. James Greiner. The new legal empiricism & its application to access-to-justice inquiries. Dædalus, 148(1):64, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_00536.

Perlman:2018:PUN

[3908] Andrew M. Perlman. The Public’s unmet need for legal services & what law schools can do about it. Dædalus, 148(1):75, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_00537.

Ashar:2018:AP

[3909] Sameer Ashar and Annie Lai. Access to power. Dædalus, 148 (1):82, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00538.

King:2018:CCF

[3910] Shani M. King. The center on children & families. Dædalus, 148 (1):88, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00539.

Rostain:2018:TOA

[3911] Tanina Rostain. Techno-optimism & access to the legal system. Dædalus, 148(1):93, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00540.

Chambliss:2018:MLA

[3912] Elizabeth Chambliss. Marketing legal assistance. Dædalus, 148 (1):98, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00541.

Herrera:2018:CLP

[3913] Luz E. Herrera. Community law practice. Dædalus, 148(1):106, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi/full/10.
1162/daed_a_00542.

Sandman:2018:RLS

[3914] James J. Sandman. The role of the legal services corporation in improving access to justice. Dædalus, 148(1):113, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00543.

Hagan:2018:PDI

[3915] Margaret Hagan. Participatory design for innovation in access to justice. Dædalus, 148(1):120, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi/full/10.
1162/daed_a_00544.

Shanahan:2018:SCC

[3916] Colleen F. Shanahan and Anna E. Carpenter. Simplified courts can’t solve inequality. Dædalus, 148 (1):128, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00545.

Wallace:2018:CSL

[3917] Jo-Ann Wallace. Corporate support for legal services. Dædalus, 148(1):136, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00546.

Pleasence:2018:JCF

[3918] Pascoe Pleasence and Nigel J. Balmer.

Justice & the capability to function in society. Dædalus, 148 (1):140, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00547.

Frazier:2018:WBB

[3919] Kenneth C. Frazier. Why big business should support legal aid. Dædalus, 148(1):150, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00548.

Lash:2018:EBS

[3920] Karen A. Lash. Executive branch support for civil legal aid. Dædalus, 148(1):160, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00549.

Fisher:2018:WJS

[3921] Fern A. Fisher. Why judges support civil legal aid. Dædalus, 148 (1):171, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_00550.

Gordon:2018:LLP

[3922] Robert W. Gordon. Lawyers, the legal profession & access to justice in the United States: a brief history. Dædalus, 148(1):177, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_00551.

Hecht:2018:TZ

[3923] Nathan L. Hecht. The twilight zone. Dædalus, 148(1):190, Winter 2018. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

                                        mitpressjournals.org/doi/full/10.
                                        1162/daed_a_00552.

Early:2019:WJS

[3924] Gerald Early and Ingrid Monson. Why jazz still matters. Dædalus, 148(2):5, Spring 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01738.

Griffin:2019:FGL

[3925] Farah Jasmine Griffin. Following Geri’s lead. Dædalus, 148 (2):13, Spring 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01739.

Solis:2019:SAT

[3926] Gabriel Solis. Soul, Afrofuturism & the timeliness of contemporary jazz fusions. Dædalus, 148 (2):23, Spring 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01740.

Wells:2019:YCD

[3927] Christopher J. Wells. “You can’t dance to it”: Jazz music and its choreographies of listening. Dædalus, 148(2):36, Spring 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01741.

Klotz:2019:DBS

[3928] Kelsey A. K. Klotz. Dave Brubeck’s southern strategy. Dædalus, 148 (2):52, Spring 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01742.

Early:2019:KJM

[3929] Gerald Early. Keith Jarrett, miscegenation & the rise of the European sensibility in jazz in the 1970s. Dædalus, 148(2):67, Spring 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01743.

Tick:2019:EFC

[3930] Judith Tick. Ella Fitzgerald & “I can’t stop loving you”, Berlin 1968: Paying homage to & signifying on soul music. Dædalus, 148(2):83, Spring 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01744.

Gabbard:2019:LHI

[3931] Krin Gabbard. La La Land is a hit, but is it good for jazz? Dædalus, 148(2):92, Spring 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01745.

Monson:2019:YLA

[3932] Ingrid Monson. Yusef Lateef’s autophysiopsychic quest. Dædalus, 148(2):104, Spring 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01746.

Muller:2019:WJS

[3933] Carol A. Muller. Why jazz? South Africa 2019. Dædalus, 148 (2):115, Spring 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01747.

Lamont:2019:IGP

[3934] Michèle Lamont and Paul Pierson. Inequality generation & persistence as multidimensional processes: an interdisciplinary agenda. Dædalus, 148(3):5, Summer 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01748.

Grusky:2019:ROM

[3935] David B. Grusky, Peter A. Hall, and Hazel Rose Markus. The rise of opportunity markets: How did it happen & what can we do? Dædalus, 148(3):19, Summer 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01749.

LeGales:2019:SCG

[3936] Patrick Le Galès and Paul Pierson. “Superstar cities” & the generation of durable inequality. Dædalus, 148(3):46, Summer 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01750.

Bloemraad:2019:MSC

[3937] Irene Bloemraad, Will Kymlicka, Michèle Lamont, and Leanne S. Son Hing. Membership without social citizenship? Deservingness & redistribution as grounds for equality. Dædalus, 148(3):73, Summer 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01751.

Hing:2019:FRR

[3938] Leanne S. Son Hing, Anne E. Wilson, Peter Gourevitch, Jaslyn English, and Parco Sin. Failure to respond to rising income inequality: Processes that legitimize growing disparities. Dædalus, 148(3):105, Summer 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01752.

Jenson:2019:DCI

[3939] Jane Jenson, Francesca Polletta, and Paige Raibmon. The difficulties of combating inequality in time. Dædalus, 148(3):136, Summer 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01753.

Hochschild:2019:PIR

[3940] Jennifer L. Hochschild. Political inequality, “real” public preferences, historical comparisons & axes of disadvantage. Dædalus, 148 (3):164, Summer 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01754.

Newman:2019:NAI

[3941] Katherine S. Newman. New angles on inequality. Dædalus, 148 (3):173, Summer 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01755.

Rao:2019:PPO

[3942] Vijayendra Rao. Process-policy & outcome-policy: Rethinking how to address poverty & inequality. Dædalus, 148(3):181, Summer 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01756.

Baum:2019:ITS

[3943] Sandy Baum and Michael McPherson. Improving teaching: Strengthening the college learning experience. Dædalus, 148(4):5, Fall 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_e_01757.

Brighouse:2019:BBC

[3944] Harry Brighouse. Becoming a better college teacher (if you’re lucky). Dædalus, 148(4):14, Fall 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01758.

Coleman:2019:CAS

[3945] Mary Sue Coleman, Tobin L. Smith, and Emily R. Miller. Catalysts for achieving sustained improvement in the quality of undergraduate STEM education. Dædalus, 148(4):29, Fall 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01759.

Wieman:2019:EUT

[3946] Carl Edwin Wieman. Expertise in university teaching & the implications for teaching effectiveness, evaluation & training. Dædalus, 148(4):47, Fall 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01760.

Tatum:2019:TAC

[3947] Beverly Daniel Tatum. Together and alone? The challenge of talking about racism on campus. Dædalus, 148(4):79, Fall 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01761.

Hurtado:2019:NTC

[3948] Sylvia Hurtado. “Now is the time”:

Civic learning for a strong democracy. Dædalus, 148(4):94, Fall 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01762.

Greenstein:2019:FUE

[3949] Daniel I. Greenstein. The future of undergraduate education: Will differences across sectors exacerbate inequality? Dædalus, 148(4):108, Fall 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01763.

Hoskins:2019:CRU

[3950] Sally G. Hoskins. CREATE a revolution in undergraduates’ understanding of science: Teach through close analysis of scientific literature. Dædalus, 148(4):138, Fall 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01764.

Bailey:2019:FDB

[3951] Thomas R. Bailey and Clive R. Belfield. The false dichotomy between academic learning & occupational skills. Dædalus, 148(4): 164, Fall 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

                                        mitpressjournals.org/doi/full/10.
                                        1162/daed_a_01765.

Morton:2019:MEC

[3952] Jennifer M. Morton. Mitigating ethical costs in the classroom. Dædalus, 148(4):179, Fall 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01766.

Castleman:2019:FCC

[3953] Benjamin Castleman and Katharine Meyer. Financial constraints & collegiate student learning: A behavioral economics perspective. Dædalus, 148(4):195, Fall 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01767.

Lewis:2019:TCL

[3954] Earl Lewis. Toward a 2.0 compact for the liberal arts. Dædalus, 148(4):217, Fall 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01768.

Baum:2019:HFP

[3955] Sandy Baum and Michael McPherson. The human factor: The promise & limits of online education. Dædalus, 148(4):235, Fall 2019. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01769.

Keohane:2020:I

[3956] Nannerl O. Keohane and Frances Mc-Call Rosenbluth. Introduction. Dædalus, 149(1):6, Winter 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_e_01770.

Teele:2020:WV

[3957] Dawn Langan Teele. Women & the vote. Dædalus, 149(1):25, Winter 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

                                        mitpressjournals.org/doi/full/10.
                                        1162/daed_a_01771.

Sanbonmatsu:2020:WUU

[3958] Kira Sanbonmatsu. Women’s underrepresentation in the U.S. Congress. Dædalus, 149(1):40, Winter 2020. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

                                        https://www.mitpressjournals.org/
                                        doi/full/10.1162/daed_a_01772.

Dancygier:2020:APD

[3959] Rafaela Dancygier. Another progressive’s dilemma: Immigration, the radical right & threats to gender equality. Dædalus, 149 (1):56, Winter 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01773.

Chira:2020:DTG

[3960] Susan Chira. Donald Trump’s gift to feminism: The resistance. Dædalus, 149(1):72, Winter 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01774.

Iversen:2020:DGE

[3961] Torben Iversen, Frances McCall Rosenbluth, and Øyvind Skorge. The dilemma of gender equality:

How labor market regulation divides women by class. Dædalus, 149(1):86, Winter 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01775.

Michener:2020:WPG

[3962] Jamila Michener and Margaret Teresa Brower. What’s policy got to do with it? Race, gender & economic inequality in the United States. Dædalus, 149(1):100, Winter 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01776.

Lowes:2020:KSW

[3963] Sara Lowes. Kinship structure & women: Evidence from economics. Dædalus, 149(1):119, Winter 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi/full/10.
1162/daed_a_01777.

Jivani:2020:GLF

[3964] Anita I. Jivani. Gender lens to the future of work. Dædalus, 149 (1):134, Winter 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01778.

Htun:2020:FVA

[3965] Mala Htun and Francesca R. Jensenius. Fighting violence against women: Laws, norms & challenges ahead. Dædalus, 149(1):144, Winter 2020. CODEN DAEDAU.

ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

                                        mitpressjournals.org/doi/full/10.
                                        1162/daed_a_01779.

Goetz:2020:NCM

[3966] Anne Marie Goetz. The new competition in multilateral norm-setting: Transnational feminists & the illiberal backlash. Dædalus, 149 (1):160, Winter 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01780.

Folke:2020:SHW

[3967] Olle Folke, Johanna Rickne, Seiki Tanaka, and Yasuka Tateishi. Sexual harassment of women leaders. Dædalus, 149(1):180, Winter 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

                                        mitpressjournals.org/doi/full/10.
                                        1162/daed_a_01781.

Folbre:2020:CCP

[3968] Nancy Folbre. Cooperation & conflict in the patriarchal labyrinth. Dædalus, 149(1):198, Winter 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi/
                                        full/10.1162/daed_a_01782.

MacKinnon:2020:E

[3969] Catharine A. MacKinnon. Equality. Dædalus, 149(1):213, Winter 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

                                        mitpressjournals.org/doi/full/10.
                                        1162/daed_a_01783.

Spar:2020:GFM

[3970] Debora L. Spar. Good fellows:

Men’s role & reason in the fight for gender equality. Dædalus, 149 (1):222, Winter 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01784.

Keohane:2020:WPL

[3971] Nannerl O. Keohane. Women, power & leadership. Dædalus, 149 (1):236, Winter 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01785.

Legvold:2020:ISS

[3972] Robert Legvold and Christopher F.

Chyba. Introduction: The search for strategic stability in a new nuclear era. Dædalus, 149(2):6, Spring 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi//doi/full/
10.1162/daed_e_01786.

Miller:2020:NWT

[3973] Steven E. Miller. A nuclear world transformed: The rise of multilateral disorder. Dædalus, 149 (2):17, Spring 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01787.

Fink:2020:RNW

[3974] Anya Loukianova Fink and Olga Oliker. Russia’s nuclear weapons in a multipolar world: Guarantors of sovereignty, great power status & more. Dædalus, 149 (2):37, Spring 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01788.

Bin:2020:RNC

[3975] Li Bin. The revival of nuclear competition in an altered geopolitical context: a Chinese perspective. Dædalus, 149(2):56, Spring 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01789.

Roberts:2020:AND

[3976] Brad Roberts. On adapting nuclear deterrence to reduce nuclear risk. Dædalus, 149(2):69, Spring 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01790.

Brooks:2020:EAC

[3977] Linton F. Brooks. The end of arms control? Dædalus, 149 (2):84, Spring 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01791.

Wolfsthal:2020:WAC

[3978] Jon Brook Wolfsthal. Why arms control? Dædalus, 149(2):101, Spring 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi//doi/full/
10.1162/daed_a_01792.

Cameron:2020:WHC

[3979] James Cameron. What history can teach. Dædalus, 149(2):116, Spring 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi//doi/full/
10.1162/daed_a_01793.

Acton:2020:CWI

[3980] James M. Acton. Cyber warfare & inadvertent escalation. Dædalus, 149(2):133, Spring 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01794.

Chyba:2020:NTS

[3981] Christopher F. Chyba. New technologies & strategic stability. Dædalus, 149(2):150, Spring 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01795.

Muller:2020:NDN

[3982] Harald Müller and Carmen Wunderlich. Nuclear disarmament without the nuclear-weapon states: The nuclear weapon ban treaty. Dædalus, 149(2):171, Spring 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01796.

Timbie:2020:WF

[3983] James Timbie. A way forward. Dædalus, 149(2):190, Spring 2020. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

https://www.mitpressjournals.org/
doi//doi/full/10.1162/daed_a_01797.

Tannenwald:2020:LBA

[3984] Nina Tannenwald. Life beyond arms control: Moving toward a global regime of nuclear restraint & responsibility. Dædalus, 149 (2):205, Spring 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01798.

Chyba:2020:CSS

[3985] Christopher F. Chyba and Robert Legvold. Conclusion: Strategic stability & nuclear war. Dædalus, 149(2):222, Spring 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01799.

Audi:2020:RDI

[3986] Robert Audi. Religion & democracy: Interactions, tensions, possibilities. Dædalus, 149(3):5, Summer 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi//doi/full/
10.1162/daed_a_01800.

Greenawalt:2020:DRS

[3987] Kent Greenawalt. Democracy & religion: Some variations & hard questions. Dædalus, 149(3):25, Summer 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi//doi/full/
10.1162/daed_a_01801.

Freeman:2020:DRP

[3988] Samuel Freeman. Democracy, religion & public reason. Dædalus, 149(3):37, Summer 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01802.

Weithman:2020:LDT

[3989] Paul Weithman. Liberalism & deferential treatment. Dædalus, 149 (3):59, Summer 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01803.

Kaveny:2020:INR

[3990] Cathleen Kaveny. The ironies of the new religious liberty litigation. Dædalus, 149(3):72, Summer 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi//doi/full/
10.1162/daed_a_01804.

Campbell:2020:PPR

[3991] David E. Campbell. The perils of politicized religion. Dædalus, 149 (3):87, Summer 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01805.

Collins:2020:ORE

[3992] Stephanie Collins. Are organizations’ religious exemptions democratically defensible? Dædalus, 149 (3):105, Summer 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01806.

Loffler:2020:SRC

[3993] Winfried Löffler. Secular reasons for confessional religious education in public schools. Dædalus, 149 (3):119, Summer 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01807.

Zucca:2020:CTA

[3994] Lorenzo Zucca. Conscience, truth & action. Dædalus, 149(3):135, Summer 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi//doi/full/
10.1162/daed_a_01808.

Gunn:2020:DHR

[3995] T. Jeremy Gunn. Do human rights have a secular, individualistic & anti-Islamic bias? Dædalus, 149 (3):148, Summer 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01809.

Jacobs:2020:JPP

[3996] Jonathan A. Jacobs. Judaism, pluralism & public reason. Dædalus, 149(3):170, Summer 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01810.

Murphy:2020:RTJ

[3997] Colleen Murphy. Religion & transitional justice. Dædalus, 149 (3):185, Summer 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01811.

Hare:2020:PMT

[3998] John E. Hare. Patriotism & moral theology. Dædalus, 149 (3):201, Summer 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01812.

Rosenblum:2020:IPW

[3999] Nancy L. Rosenblum. Introduction: Paths to witnessing, ethics of speaking out. Dædalus, 149(4): 6, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi//doi/full/
10.1162/daed_e_01813.

Lifton:2020:BWP

[4000] Robert Jay Lifton. On becoming witnessing professionals. Dædalus, 149(4):25, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01814.

Oreskes:2020:WSR

[4001] Naomi Oreskes. What is the social responsibility of climate scientists? Dædalus, 149(4):33, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01815.

Socolow:2020:WMD

[4002] Robert H. Socolow. Witnessing for the middle to depolarize the climate change conversation. Dædalus, 149(4):46, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01816.

Thompson:2020:PEW

[4003] Dennis F. Thompson. The professional ethics of witnessing professionals. Dædalus, 149(4):67, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01817.

Gerrard:2020:ELF

[4004] Michael B. Gerrard. An environmental lawyer’s fraught quest for legal tools to hold back the seas. Dædalus, 149(4):79, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01818.

Mitchell:2020:RMC

[4005] Mark A. Mitchell. Racism as a motivator for climate justice. Dædalus, 149(4):96, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01819.

Kinney:2020:APC

[4006] Patrick L. Kinney. From air pollution to the climate crisis: Leaving the comfort zone. Dædalus, 149(4):108, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01820.

Henderson:2020:CBS

[4007] Rebecca Henderson. Climate in the boardroom: Struggling to reconcile business as usual & the end of the world as we know it. Dædalus, 149(4):118, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01821.

Titley:2020:TFC

[4008] David W. Titley. Task force climate change: a patron saint of lost causes, or just ahead of its time? Dædalus, 149(4):125, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01822.

Weber:2020:SBU

[4009] Elke U. Weber. Seeing is believing:

Understanding & aiding human responses to global climate change. Dædalus, 149(4):139, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi//
doi/full/10.1162/daed_a_01823.

Green:2020:LTM

[4010] Jessica F. Green. Less talk, more walk: Why climate change demands activism in the academy. Dædalus, 149(4):151, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi//
                                        doi/full/10.1162/daed_a_01824.

Rosenblum:2020:C

[4011] Nancy L. Rosenblum and Rafe Pomerance. A conversation. Dædalus, 149(4):163, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi//
                                        doi/full/10.1162/daed_a_01825.

Kormann:2020:CAD

[4012] Carolyn Kormann. The coral is not all dead yet. Dædalus, 149(4):180, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi//
                                        doi/full/10.1162/daed_a_01826.

Knowles:2020:SDA

[4013] Scott Gabriel Knowles. Slow disaster in the Anthropocene: a historian witnesses climate change on the Korean peninsula. Dædalus, 149(4):192, Fall 2020. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

                                        //www.mitpressjournals.org/doi//
                                        doi/full/10.1162/daed_a_01827.

Oposa:2020:LMT

[4014] Antonio Oposa Jr. Let me tell you a story. Dædalus, 149(4): 207, Fall 2020. CODEN DAEDAU.

ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi//doi/full/
10.1162/daed_a_01828.

Wood:2021:IW

[4015] Michael Wood. Introduction: In this world. Dædalus, 150(01):5, Winter 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi/full/10.
1162/daed_e_01829.

Moore:2021:WIL

[4016] Lorrie Moore. What is it like to write a novel? Dædalus, 150 (01):14, Winter 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01830.

Moretti:2021:TT

[4017] Franco Moretti. Two theories. Dædalus, 150(01):16, Winter 2021. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

https://www.mitpressjournals.org/
doi/full/10.1162/daed_a_01831.

Goldhill:2021:FTA

[4018] Simon Goldhill. Finding the time for ancient novels. Dædalus, 150 (01):26, Winter 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01832.

Armstrong:2021:SEF

[4019] Nancy Armstrong. Some endangered feeling. Dædalus, 150 (01):40, Winter 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01833.

Yeazell:2021:HJC

[4020] Ruth Bernard Yeazell. Henry James in — and out of — the classroom. Dædalus, 150(01):62, Winter 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi/full/10.
1162/daed_a_01834.

Cameron:2021:HCH

[4021] Sharon Cameron. The hole in the carpet: Henry James’s The Bostonians. Dædalus, 150(01):78, Winter 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi/full/10.
1162/daed_a_01835.

Brooks:2021:WST

[4022] Daphne A. Brooks. “A woman is a sometime thing”: (re)covering black womanhood in Porgy and Bess. Dædalus, 150(01):98, Winter 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi/full/10.
1162/daed_a_01836.

Chow:2021:W

[4023] Rey Chow and Austin Sarfan. We “other Victorians”? Novelistic remains, therapeutic devices, contemporary televisual dramas. Dædalus, 150(01):118, Winter 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01837.

Dimock:2021:SU

[4024] Wai Chee Dimock. The survival of the unfit. Dædalus, 150 (01):134, Winter 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01838.

Creswell:2021:PPG

[4025] Robyn Creswell. Poets in prose: Genre & history in the Arabic novel. Dædalus, 150(01):147, Winter 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https:

//www.mitpressjournals.org/doi/
full/10.1162/daed_a_01839.

Stewart:2021:ORR

[4026] Garrett Stewart. Organic reformations in Richard Powers’s The Overstory. Dædalus, 150(01):160, Winter 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi/full/10.
1162/daed_a_01840.

Hayot:2021:VGN

[4027] Eric Hayot. Video games & the novel. Dædalus, 150(01):178, Winter 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

mitpressjournals.org/doi/full/10.
1162/daed_a_01841.

Greenberg:2021:LTT

[4028] Jonathan Greenberg. Losing track of time. Dædalus, 150(01):188, Winter 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://www.

                                        mitpressjournals.org/doi/full/10.
                                        1162/daed_a_01842.

Massey:2021:BOW

[4029] Douglas S. Massey. The bipartisan origins of White nationalism. Dædalus, 150(2):5–22, Spring 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic).  URL https://direct.
                                        mit.edu/daed/article-pdf/150/2/
                                        5/1897830/daed_a_01843.pdf.

Hajnal:2021:IOW

[4030] Zoltan Hajnal. Immigration & the origins of White backlash. Dædalus, 150(2):23–39, Spring 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        150/2/23/1897807/daed_a_01844.pdf.

Hout:2021:IRP

[4031] Michael Hout and Christopher Maggio. Immigration, race & political polarization. Dædalus, 150(2):40–55, Spring 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/150/2/
                                        40/1897816/daed_a_01845.pdf.

Parker:2021:STM

[4032] Christopher Sebastian Parker. Status threat: Moving the right further to the right? Dædalus, 150(2):56– 75, Spring 2021. CODEN DAEDAU.

ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/150/2/
56/1901022/daed_a_01846.pdf.

Monk:2021:USC

[4033] Ellis P. Monk, Jr. The unceasing significance of colorism: Skin tone stratification in the United States. Dædalus, 150(2):76–90, Spring 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.
mit.edu/daed/article-pdf/150/2/
76/1897805/daed_a_01847.pdf.

Menjivar:2021:RI

[4034] Cecilia Menjívar. The racialization of “Illegality”. Dædalus, 150(2):91– 105, Spring 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/150/2/
91/1897814/daed_a_01848.pdf.

GarciaHernandez:2021:CM

[4035] César Cuauhtémoc García Hernández.

Criminalizing migration. Dædalus, 150 (2):106–119, Spring 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
150/2/106/1897803/daed_a_01849.
pdf.

Waters:2021:RLS

[4036] Mary C. Waters and Philip Kasinitz.

Race, legal status & social mobility. Dædalus, 150(2):120–134, Spring 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.

edu/daed/article-pdf/150/2/120/
1897828/daed_a_01850.pdf.

Gonzales:2021:LSD

[4037] Roberto G. Gonzales and Stephen P. Ruszczyk. The legal status divide among the children of immigrants. Dædalus, 150(2):135–149, Spring 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/150/2/135/
1897801/daed_a_01851.pdf.

Canizales:2021:LRT

[4038] Stephanie L. Canizales and Jody Agius Vallejo. Latinos & racism in the Trump Era. Dædalus, 150(2):150–164, Spring 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/150/2/150/
1897799/daed_a_01852.pdf.

Ceciliano-Navarro:2021:TMY

[4039] Yajaira Ceciliano-Navarro and Tanya Maria Golash-Boza. ”Trauma makes you grow up quicker”: The financial & emotional burdens of deportation & incarceration. Dædalus, 150(2):165–179, Spring 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/150/2/165/
1897812/daed_a_01853.pdf.

Lee:2021:AAA

[4040] Jennifer Lee. Asian Americans, Affirmative Action & the rise in anti-Asian hate. Dædalus, 150(2):180–198, Spring 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.

edu/daed/article-pdf/150/2/180/
1897837/daed_a_01854.pdf.

Alba:2021:SYA

[4041] Richard Alba. The surge of young Americans from minority-White mixed families & its significance for the future. Dædalus, 150(2):199–214, Spring 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/150/2/199/
1897823/daed_a_01855.pdf.

Tushnet:2021:IPF

[4042] Mark Tushnet. Introduction: The pasts & futures of the administrative state. Dædalus, 150(3):5–16, Summer 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/150/3/
5/1927453/daed_e_01856.pdf.

Strauss:2021:HAS

[4043] Peter L. Strauss. How the administrative state got to this challenging place. Dædalus, 150(3):17–32, Summer 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/150/3/
17/1927459/daed_a_01857.pdf.

Dudley:2021:MEA

[4044] Susan E. Dudley. Milestones in the evolution of the administrative state. Dædalus, 150(3):33–48, Summer 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.
mit.edu/daed/article-pdf/150/3/
33/1927451/daed_a_01858.pdf.

Farhang:2021:LCA

[4045] Sean Farhang. Legislative capacity & administrative power under divided polarization. Dædalus, 150(3):49–67, Summer 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/150/3/
                                        49/1927471/daed_a_01859.pdf.

Lewis:2021:FPR

[4046] David E. Lewis. Is the failed pandemic response a symptom of a diseased administrative state? Dædalus, 150(3):68–88, Summer 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        150/3/68/1927465/daed_a_01860.pdf.

Bell:2021:RBA

[4047] Bernard W. Bell. Replacing bureaucrats with automated sorcerers? Dædalus, 150(3):89–103, Summer 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/150/3/
                                        89/1927443/daed_a_01861.pdf.

Coglianese:2021:ALA

[4048] Cary Coglianese. Administrative law in the automated state. Dædalus, 150 (3):104–120, Summer 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        150/3/104/1927461/daed_a_01862.
                                        pdf.

Noveck:2021:IS

[4049] Beth Simone Noveck. The innovative state. Dædalus, 150(3):121–142, Summer 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/150/3/121/
1927463/daed_a_01863.pdf.

Nielson:2021:DD

[4050] Aaron L. Nielson. Deconstruction (not destruction). Dædalus, 150(3):143–154, Summer 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/150/3/
143/1927457/daed_a_01864.pdf.

Walker:2021:CBB

[4051] Christopher J. Walker. Constraining bureaucracy beyond judicial review. Dædalus, 150(3):155–171, Summer 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/150/3/155/
1927447/daed_a_01865.pdf.

White:2021:CPB

[4052] Avery White and Michael Neblo. Capturing the public: Beyond technocracy & populism in the U.S. administrative state. Dædalus, 150(3):172–187, Summer 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/150/3/172/
1927455/daed_a_01866.pdf.

Kessler:2021:UFA

[4053] Jeremy Kessler and Charles Sabel.

The uncertain future of administrative law. Dædalus, 150(3):188–207, Summer 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.

                                        edu/daed/article-pdf/150/3/188/
                                        1927441/daed_a_01867.pdf.

Sunstein:2021:SCB

[4054] Cass R. Sunstein. Some costs & benefits of cost–benefit analysis. Dædalus, 150(3):208–219, Summer 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/150/3/208/
                                        1927445/daed_a_01868.pdf.

Rao:2021:HFA

[4055] Neomi Rao. The hedgehog & the fox in administrative law. Dædalus, 150 (3):220–241, Summer 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        150/3/220/1927449/daed_a_01869.
                                        pdf.

Isaacman:2021:WSA

[4056] Allen Isaacman and Muchaparara Musemwa. Water security in Africa in the age of global climate change. Dædalus, 150(4):7– 26, Fall 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/150/4/
                                        7/1963555/daed_e_01870.pdf.

Musemwa:2021:USW

[4057] Muchaparara Musemwa. Urban struggles over water scarcity in Harare. Dædalus, 150(4):27–47, Fall 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic).  URL https://direct.
                                        mit.edu/daed/article-pdf/150/4/
                                        27/1963553/daed_a_01871.pdf.

Bender:2021:WBC

[4058] Matthew V. Bender. Water for Bongo: Creative adaptation, resilience & Dar es Salaam’s water supply. Dædalus, 150(4):48–63, Fall 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.
mit.edu/daed/article-pdf/150/4/
48/1963543/daed_a_01872.pdf.

Harris:2021:EEW

[4059] Leila M. Harris. Everyday experiences of water insecurity: Insights from underserved areas of Accra, Ghana. Dædalus, 150(4):64– 84, Fall 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/150/4/
64/1963566/daed_a_01873.pdf.

Livingston:2021:WSH

[4060] Julie Livingston. Water scarcity & health in urban Africa. Dædalus, 150(4):85–102, Fall 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
150/4/85/1963561/daed_a_01874.pdf.

Isaacman:2021:CBD

[4061] Allen Isaacman. Cahora Bassa Dam & the delusion of development. Dædalus, 150(4):103–123, Fall 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
150/4/103/1963569/daed_a_01875.
pdf.

Miescher:2021:GAD

[4062] Stephan F. Miescher. Ghana’s Akosombo Dam, Volta Lake fisheries & climate change. Dædalus, 150(4):124– 142, Fall 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/150/4/
124/1963559/daed_a_01876.pdf.

Derr:2021:DBT

[4063] Jennifer L. Derr. The dammed body: Thinking historically about water security & public health. Dædalus, 150(4):143–158, Fall 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
150/4/143/1963549/daed_a_01877.
pdf.

Verhoeven:2021:GER

[4064] Harry Verhoeven. The Grand Ethiopian Renaissance Dam: Africa’s water tower, environmental justice & infrastructural power. Dædalus, 150(4):159–180, Fall 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
150/4/159/1963563/daed_a_01878.
pdf.

Mwangi:2021:HVH

[4065] Oscar Gakuo Mwangi. Hydropolitics versus human security: Implications of South Africa’s appropriation of Lesotho’s highlands water. Dædalus, 150(4):181–193, Fall 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/

150/4/181/1963557/daed_a_01879.
pdf.

Hwang:2021:OYC

[4066] Jyhjong Hwang. An offer you can refuse: a host country’s strategic allocation of development financing. Dædalus, 150(4):194–219, Fall 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/150/4/194/
1964055/daed_a_01880.pdf.

Klug:2021:BPP

[4067] Heinz Klug. Between principles & power: Water law principles & the governance of water in post-apartheid South Africa. Dædalus, 150(4):220– 239, Fall 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/150/4/
220/1963541/daed_a_01881.pdf.

King:2021:ALR

[4068] Jackie King and Cate Brown. Africa’s living rivers: Managing for sustainability. Dædalus, 150(4):240–259, Fall 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/150/4/240/
1963545/daed_a_01882.pdf.

Verhoeven:2021:CWC

[4069] Harry Verhoeven. Climate & water in a changing Africa: Uncertainty, adaptation & the social construction of fragile environments. Dædalus, 150(4):260–277, Fall 2021. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        150/4/260/1963551/daed_a_01883.
                                        pdf.

Western:2022:VCP

[4070] Bruce Western and Sukyi McMahon. Violence, criminalization & punitive excess. Dædalus, 151(1):6–10, Winter 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/151/1/
                                        6/1978944/daed_a_01888.pdf.

Jackson:2022:SVA

[4071] Kellie Carter Jackson. The story of violence in America. Dædalus, 151(1):11– 21, Winter 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/151/1/
                                        11/1978940/daed_a_01884.pdf.

Butler:2022:PSV

[4072] Paul Butler. The problem of state violence. Dædalus, 151(1):22–37, Winter 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/151/1/
                                        22/1978950/daed_a_01885.pdf.

Webster:2022:PHA

[4073] Daniel W. Webster. Public health approaches to reducing community gun violence. Dædalus, 151(1):38– 48, Winter 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/151/1/
                                        38/1978981/daed_a_01886.pdf.

Hureau:2022:SGS

[4074] David M. Hureau. Seeing guns to see urban violence: Racial inequality & neighborhood context. Dædalus, 151(1):49–66, Winter 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.
mit.edu/daed/article-pdf/151/1/
49/1978954/daed_a_01887.pdf.

Keels:2022:DEP

[4075] Micere Keels. Developmental & ecological perspective on the intergenerational transmission of trauma & violence. Dædalus, 151(1):67–83, Winter 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/151/1/
67/1978936/daed_a_01889.pdf.

Richie:2022:EVC

[4076] Beth E. Richie. The effects of violence on communities: The violence matrix as a tool for advancing more just policies. Dædalus, 151(1):84– 96, Winter 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/151/1/
84/1978931/daed_a_01890.pdf.

Jones:2022:FAV

[4077] Barbara L. Jones. Faces of the aftermath of visible & invisible violence & loss: Radical resiliency of justice & healing. Dædalus, 151(1):97–106, Winter 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/151/1/
97/1978946/daed_a_01891.pdf.

Muhammad:2022:FLL

[4078] Khalil Gibran Muhammad. The foundational lawlessness of the law itself: Racial criminalization & the punitive roots of punishment in America. Dædalus, 151(1):107–120, Winter 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/151/1/107/
                                        1978948/daed_a_01892.pdf.

Chacon:2022:CLM

[4079] Jennifer M. Chacón. Criminal law & migration control: Recent history & future possibilities. Dædalus, 151 (1):121–134, Winter 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        151/1/121/1978942/daed_a_01893.
                                        pdf.

Cleve:2022:DPT

[4080] Nicole Gonzalez Van Cleve. Due process & the theater of racial degradation: The evolving notion of pretrial punishment in the criminal courts. Dædalus, 151(1):135–152, Winter 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        151/1/135/1978929/daed_a_01894.
                                        pdf.

Ward:2022:RRR

[4081] Geoff K. Ward. Recognition, repair & the reconstruction of “square one”. Dædalus, 151(1):153–169, Winter 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.

edu/daed/article-pdf/151/1/153/
1978933/daed_a_01895.pdf.

Simon:2022:KWW

[4082] Jonathan Simon. Knowing what we want: a decent society, a civilized system of justice & a condition of dignity. Dædalus, 151(1):170–180, Winter 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/151/1/170/
1978938/daed_a_01896.pdf.

Manyika:2022:GAR

[4083] James Manyika. Getting AI right:

Introductory notes on AI & society. Dædalus, 151(2):5–27, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.
mit.edu/daed/article-pdf/151/2/
5/2009531/daed_e_01897.pdf.

Shadbolt:2022:SBS

[4084] Nigel Shadbolt. ”From so simple a beginning”: Species of artificial intelligence. Dædalus, 151(2):28–42, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/151/2/
28/2009160/daed_a_01898.pdf.

Russell:2022:IWS

[4085] Stuart Russell. If we succeed. Dædalus, 151(2):43–57, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
151/2/43/2009166/daed_a_01899.pdf.

Dean:2022:GDD

[4086] Jeffrey Dean. A golden decade of deep learning: Computing systems & applications. Dædalus, 151(2):58– 74, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/151/2/
                                        58/2009186/daed_a_01900.pdf.

Scott:2022:DTI

[4087] Kevin Scott. I do not think it means what you think it means: Artificial intelligence, cognitive work & scale. Dædalus, 151(2):75–84, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/151/2/
                                        75/2009143/daed_a_01901.pdf.

Fei-Fei:2022:SCV

[4088] Li Fei-Fei and Ranjay Krishna. Searching for computer vision north stars. Dædalus, 151(2):85–99, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic).  URL https://direct.
                                        mit.edu/daed/article-pdf/151/2/
                                        85/2009150/daed_a_01902.pdf.

Rus:2022:MOF

[4089] Daniela Rus. The machines from our future. Dædalus, 151(2):100–113, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/151/2/100/
                                        2009193/daed_a_01903.pdf.

Gal:2022:MAS

[4090] Kobi Gal and Barbara J. Grosz. Multiagent systems: Technical & ethical challenges of functioning in a mixed group. Dædalus, 151(2):114–126, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/151/2/114/
2009176/daed_a_01904.pdf.

Manning:2022:HLU

[4091] Christopher D. Manning. Human language understanding & reasoning. Dædalus, 151(2):127–138, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/151/2/127/
2009183/daed_a_01905.pdf.

Choi:2022:CCC

[4092] Yejin Choi. The curious case of commonsense intelligence. Dædalus, 151 (2):139–155, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
151/2/139/2009208/daed_a_01906.
pdf.

Murati:2022:LCC

[4093] Ermira Murati. Language & coding creativity. Dædalus, 151(2):156–167, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/151/2/156/
2009202/daed_a_01907.pdf.

Rees:2022:NHW

[4094] Tobias Rees. Non-human words:

On GPT-3 as a philosophical laboratory. Dædalus, 151(2):168–182, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/151/2/168/
                                        2009141/daed_a_01908.pdf.

AguerayArcas:2022:DLL

[4095] Blaise Agüera y Arcas. Do large language models understand us? Dædalus, 151(2):183–197, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        151/2/183/2009147/daed_a_01909.
                                        pdf.

Elam:2022:STW

[4096] Michele Elam. Signs taken for wonders: AI, art & the matter of race. Dædalus, 151(2):198–217, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        151/2/198/2009152/daed_a_01910.
                                        pdf.

Gabriel:2022:TTJ

[4097] Iason Gabriel. Toward a theory of justice for artificial intelligence. Dædalus, 151(2):218–231, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        151/2/218/2009164/daed_a_01911.
                                        pdf.

Tasioulas:2022:AIH

[4098] John Tasioulas. Artificial intelligence, humanistic ethics. Dædalus, 151 (2):232–243, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/

151/2/232/2009174/daed_a_01912.
pdf.

Spence:2022:AAV

[4099] Michael Spence. Automation, augmentation, value creation & the distribution of income & wealth. Dædalus, 151(2):244–255, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
151/2/244/2009180/daed_a_01913.
pdf.

Tyson:2022:AAW

[4100] Laura D. Tyson and John Zysman. Automation, AI & work. Dædalus, 151 (2):256–271, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
151/2/256/2009172/daed_a_01914.
pdf.

Brynjolfsson:2022:TTP

[4101] Erik Brynjolfsson. The Turing trap:

The promise & peril of human-like artificial intelligence. Dædalus, 151 (2):272–287, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
151/2/272/2009138/daed_a_01915.
pdf.

Schmidt:2022:AGP

[4102] Eric Schmidt. AI, great power competition & national security. Dædalus, 151(2):288–298, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/

                                        151/2/288/2009162/daed_a_01916.
                                        pdf.

Carter:2022:MDA

[4103] Ash Carter. The moral dimension of AI-assisted decision-making: Some practical perspectives from the front lines. Dædalus, 151(2):299–308, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/151/2/299/
                                        2009191/daed_a_01917.pdf.

Dwork:2022:DAI

[4104] Cynthia Dwork and Martha Minow. Distrust of artificial intelligence: Sources & responses from computer science & law. Dædalus, 151(2):309–321, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/151/2/309/
                                        2009145/daed_a_01918.pdf.

Katyal:2022:DDE

[4105] Sonia K. Katyal. Democracy & distrust in an era of artificial intelligence. Dædalus, 151(2):322–334, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/151/2/322/
                                        2009189/daed_a_01919.pdf.

Cuellar:2022:AIR

[4106] Mariano-Florentino Cuéllar and Aziz Z. Huq. Artificially intelligent regulation. Dædalus, 151(2):335–347, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.

edu/daed/article-pdf/151/2/335/
2009200/daed_a_01920.pdf.

Coyle:2022:SD

[4107] Diane Coyle. Socializing data.

Dædalus, 151(2):348–359, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/151/2/348/
2009198/daed_a_01921.pdf.

Margetts:2022:RAG

[4108] Helen Margetts. Rethinking AI for good governance. Dædalus, 151 (2):360–371, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
151/2/360/2009178/daed_a_01922.
pdf.

Manyika:2022:ASI

[4109] James Manyika. Afterword: Some illustrations. Dædalus, 151(2):372–379, Spring 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/151/2/372/
2009170/daed_a_01923.pdf.

Berkowitz:2022:I

[4110] Carin Berkowitz, Norman Bradburn, and Robert B. Townsend. Introduction. Dædalus, 151(3):6–10, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.
mit.edu/daed/article-pdf/151/3/
6/2060688/daed_e_01924.pdf.

Townsend:2022:SHC

[4111] Robert B. Townsend and Norman Bradburn. The state of the humanities circa 2022. Dædalus, 151(3):11– 18, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/151/3/
11/2060696/daed_a_01925.pdf.

Liu:2022:WES

[4112] Alan Liu, Abigail Droge, Scott Kleinman, Lindsay Thomas, Dan C. Baciu, and Jeremy Douglass. What everyone says: Public perceptions of the humanities in the media. Dædalus, 151(3):19– 39, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/151/3/
19/2060702/daed_a_01926.pdf.

Butler:2022:PFH

[4113] Judith Butler. The public futures of the humanities. Dædalus, 151(3):40– 53, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/151/3/
40/2060686/daed_a_01927.pdf.

Guyer:2022:BSG

[4114] Sara Guyer. Beyond the survival of the global humanities. Dædalus, 151(3):54– 67, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/151/3/
54/2060674/daed_a_01928.pdf.

Berkowitz:2022:RPH

[4115] Carin Berkowitz and Matthew Gibson. Reframing the public humanities: The tensions, challenges & potentials of a more expansive endeavor. Dædalus, 151(3):68–81, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/151/3/
68/2060678/daed_a_01929.pdf.

Sanchez:2022:OHN

[4116] George J. Sánchez. Opening the humanities to new fields & new voices. Dædalus, 151(3):82–93, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.
mit.edu/daed/article-pdf/151/3/
82/2060670/daed_a_01930.pdf.

Meringolo:2022:CKP

[4117] Denise D. Meringolo, Lee Boot, Denise Griffin Johnson, and Maureen O’Neill. Creating knowledge with the public: Disrupting the expert/audience hierarchy. Dædalus, 151(3):94–107, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/151/3/
94/2060684/daed_a_01931.pdf.

Ruffins:2022:GMC

[4118] Fath Davis Ruffins. Grassroots museums & the changing landscape of the public humanities. Dædalus, 151 (3):108–123, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
151/3/108/2060680/daed_a_01932.
pdf.

Smulyan:2022:WPH

[4119] Susan Smulyan. Why public humanities? Dædalus, 151(3):124–137, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/151/3/124/
                                        2060690/daed_a_01933.pdf.

Balleisen:2022:CBE

[4120] Edward J. Balleisen and Rita Chin. The case for bringing experiential learning into the humanities. Dædalus, 151(3):138–152, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/151/3/138/
                                        2060692/daed_a_01934.pdf.

Hart:2022:CMA

[4121] Roderick P. Hart. Communication & media arts: Of the humanities & the future. Dædalus, 151(3):153–165, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/151/3/153/
                                        2060668/daed_a_01935.pdf.

Magness:2022:RSI

[4122] Jodi Magness and Margaret M. Mitchell. Religious studies & the imagined boundaries of the humanities. Dædalus, 151(3):166–179, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/151/3/166/
                                        2060672/daed_a_01936.pdf.

Appiah:2022:PHL

[4123] Kwame Anthony Appiah. Philosophy, the humanities & the life of freedom. Dædalus, 151(3):180–193, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/151/3/180/
2060676/daed_a_01937.pdf.

Wailoo:2022:PHT

[4124] Keith Wailoo. Patients are humans too: The emergence of medical humanities. Dædalus, 151(3):194–205, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/151/3/194/
2060694/daed_a_01938.pdf.

Pawelski:2022:PHF

[4125] James O. Pawelski. The positive humanities: A focus on human flourishing. Dædalus, 151(3):206–221, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/151/3/206/
2060682/daed_a_01939.pdf.

Chakrabarty:2022:PHS

[4126] Dipesh Chakrabarty. Planetary humanities: Straddling the decolonial/ postcolonial divide. Dædalus, 151 (3):222–233, Summer 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
151/3/222/2060700/daed_a_01940.
pdf.

Brady:2022:I

[4127] Henry E. Brady and Kay Lehman Schlozman. Introduction. Dædalus, 151(4):6–24, Fall 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        151/4/6/2058565/daed_e_01941.pdf.

Jasanoff:2022:DTT

[4128] Sheila Jasanoff. The discontents of truth & trust in 21st Century America. Dædalus, 151(4):25–42, Fall 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic).  URL https://direct.
                                        mit.edu/daed/article-pdf/151/4/
                                        25/2058542/daed_a_01942.pdf.

Brady:2022:FYD

[4129] Henry E. Brady and Thomas B. Kent. Fifty years of declining confidence & increasing polarization in trust in American institutions. Dædalus, 151(4):43– 66, Fall 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/151/4/
                                        43/2058597/daed_a_01943.pdf.

Blendon:2022:TMH

[4130] Robert J. Blendon and John M. Benson. Trust in medicine, the health system & public health. Dædalus, 151(4): 67–82, Fall 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/151/4/
                                        67/2058567/daed_a_01944.pdf.

Hatton:2022:ATS

[4131] C. Ross Hatton, Colleen L. Barry, Adam S. Levine, Emma E. McGinty, and Hahrie Han. American trust in science & institutions in the time of COVID-19. Dædalus, 151(4):83– 97, Fall 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/151/4/
83/2058577/daed_a_01945.pdf.

Oreskes:2022:AGA

[4132] Naomi Oreskes and Erik M. Conway.

From anti-government to anti-science: Why conservatives have turned against science. Dædalus, 151(4):98–123, Fall 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.
mit.edu/daed/article-pdf/151/4/
98/2058549/daed_a_01946.pdf.

Rainie:2022:NTF

[4133] Lee Rainie. Networked trust & the future of media. Dædalus, 151(4):124– 143, Fall 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/151/4/
124/2058584/daed_a_01947.pdf.

Schudson:2022:WDT

[4134] Michael Schudson. What does “Trust in the Media” mean? Dædalus, 151(4):144–160, Fall 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
151/4/144/2058562/daed_a_01948.
pdf.

Meares:2022:TMP

[4135] Tracey L. Meares. Trust & models of policing. Dædalus, 151(4):161–176, Fall 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/151/4/161/
                                        2058573/daed_a_01949.pdf.

Wu:2022:RPT

[4136] Cary Wu, Rima Wilkes, and David C. Wilson. Race & political trust: Justice as a unifying influence on political trust. Dædalus, 151(4):177–199, Fall 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/151/4/177/
                                        2058590/daed_a_01950.pdf.

Wuthnow:2022:RDT

[4137] Robert Wuthnow. Religion, democracy & the task of restoring trust. Dædalus, 151(4):200–214, Fall 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        151/4/200/2058571/daed_a_01951.
                                        pdf.

Levi:2022:TGO

[4138] Margaret Levi. Trustworthy government: The obligations of government & the responsibilities of the governed. Dædalus, 151(4):215–233, Fall 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/151/4/215/
                                        2058537/daed_a_01952.pdf.

Stewart:2022:TE

[4139] III Charles Stewart. Trust in elections. Dædalus, 151(4):234–253, Fall 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.

edu/daed/article-pdf/151/4/234/
2058544/daed_a_01953.pdf.

Margulies:2022:SST

[4140] Max Margulies and Jessica Blankshain.

Specific sources of trust in generals: Individual-level trust in the U.S. military. Dædalus, 151(4):254–275, Fall 2022. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/151/4/254/
2058552/daed_a_01954.pdf.

Levi:2023:MI

[4141] Margaret Levi and Zachary Ugolnik. Mobilizing in the interest of others. Dædalus, 152(1):7–18, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.
mit.edu/daed/article-pdf/152/1/
7/2072579/daed_a_01955.pdf.

Bowles:2023:FEC

[4142] Samuel Bowles and Wendy Carlin.

Foundations of an expanded community of fate. Dædalus, 152(1):19–24, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/152/1/
19/2072620/daed_a_01956.pdf.

Cuellar:2023:RPE

[4143] Mariano-Florentino Cuéllar. Reimagining political economy without “Yanking on a thread before it’s ready”. Dædalus, 152(1):25–30, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.

                                        mit.edu/daed/article-pdf/152/1/
                                        25/2072599/daed_a_01957.pdf.

Bednar:2023:GHS

[4144] Jenna Bednar. Governance for human social flourishing. Dædalus, 152(1):31– 45, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/152/1/
                                        31/2072607/daed_a_01958.pdf.

Singh:2023:ACH

[4145] Prerna Singh. All (cautiously) hail-and scale-community! Dædalus, 152(1):46– 51, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/152/1/
                                        46/2072595/daed_a_01959.pdf.

Kennedy:2023:PPH

[4146] III Joseph Kennedy. Power to pursue happiness. Dædalus, 152(1):52– 57, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/152/1/
                                        52/2072589/daed_a_01960.pdf.

Gopnik:2023:CPB

[4147] Alison Gopnik. Caregiving in philosophy, biology & political economy. Dædalus, 152(1):58–69, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic).  URL https://direct.
                                        mit.edu/daed/article-pdf/152/1/
                                        58/2072568/daed_a_01961.pdf.

Slaughter:2023:CR

[4148] Anne-Marie Slaughter. Care is a relationship. Dædalus, 152(1):70–76, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/152/1/
70/2072583/daed_a_01962.pdf.

Teles:2023:EP

[4149] Steven M. Teles. Egalitarian pluralism. Dædalus, 152(1):77–80, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.
mit.edu/daed/article-pdf/152/1/
77/2072613/daed_a_01963.pdf.

Iskander:2023:BIB

[4150] Natasha Iskander and Nichola Lowe.

Biophilic institutions: Building new solidarities between the economy & nature. Dædalus, 152(1):81–93, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.
mit.edu/daed/article-pdf/152/1/
81/2072562/daed_a_01964.pdf.

Beinhocker:2023:BM

[4151] Eric D. Beinhocker. Biophilic markets. Dædalus, 152(1):94–99, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.
mit.edu/daed/article-pdf/152/1/
94/2072560/daed_a_01965.pdf.

Livingston:2023:BMD

[4152] Julie Livingston. Biophilia & military degrowth. Dædalus, 152(1):100–104, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/152/1/100/
2072566/daed_a_01966.pdf.

Ahlquist:2023:MDJ

[4153] John S. Ahlquist. Making decent jobs. Dædalus, 152(1):105–118, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/152/1/105/
                                        2072577/daed_a_01967.pdf.

Naidu:2023:EJ

[4154] Suresh Naidu. Eudaimonic jobs. Dædalus, 152(1):119–124, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/152/1/119/
                                        2072626/daed_a_01968.pdf.

Miller:2023:MAS

[4155] Michelle Miller. Mutual aid as spiritual sustenance. Dædalus, 152(1):125–130, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/152/1/125/
                                        2072587/daed_a_01969.pdf.

Locke:2023:SCW

[4156] Richard M. Locke, Ben Armstrong, Samantha Schaab-Rozbicki, and Geordie Young. Supply chains & working conditions during the long pandemic: Lessons for a new moral political economy? Dædalus, 152(1):131–142, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/152/1/131/
                                        2072593/daed_a_01970.pdf.

Cohen:2023:DWD

[4157] Joshua Cohen. Doing well by doing right. Dædalus, 152(1):143–148, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/152/1/143/
2072597/daed_a_01971.pdf.

Charo:2023:UW

[4158] R. Alta Charo. Unchaining workers. Dædalus, 152(1):149–153, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/152/1/149/
2072605/daed_a_01972.pdf.

Chelwa:2023:IGS

[4159] Grieve Chelwa, Darrick Hamilton, and Avi Green. Identity group stratification, political economy & inclusive economic rights. Dædalus, 152 (1):154–167, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
152/1/154/2072611/daed_a_01973.
pdf.

Farrell:2023:RTV

[4160] Henry Farrell and Margaret Levi. Reducing the transactional value of identity & race. Dædalus, 152(1):168–173, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/152/1/168/
2072616/daed_a_01974.pdf.

Pastor:2023:NFW

[4161] Manuel Pastor. Neoliberal fragility:

Why it’s so hard for (some) economists to talk about racism. Dædalus, 152 (1):174–178, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        152/1/174/2072624/daed_a_01975.
                                        pdf.

Satz:2023:DNM

[4162] Debra Satz. Democracy & “Noxious” markets. Dædalus, 152(1):179–188, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/152/1/179/
                                        2072558/daed_a_01976.pdf.

Fleurbaey:2023:TPS

[4163] Marc Fleurbaey. Is there a proper scope for markets? Dædalus, 152 (1):189–193, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        152/1/189/2072575/daed_a_01977.
                                        pdf.

Thurston:2023:HSW

[4164] Chloe Thurston. How should we govern housing markets in a moral political economy? Dædalus, 152(1):194–197, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/152/1/194/
                                        2072585/daed_a_01978.pdf.

Henderson:2023:MF

[4165] Rebecca Henderson. Moral firms? Dædalus, 152(1):198–211, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/152/1/198/
                                        2072601/daed_a_01979.pdf.

Mayer:2023:MFC

[4166] Colin Mayer. Are moral firms committed firms? Dædalus, 152(1):212–216, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/152/1/212/
2072581/daed_a_01980.pdf.

OMara:2023:CFA

[4167] Margaret O’Mara. Can firms act morally? Dædalus, 152(1):217–224, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/152/1/217/
2072564/daed_a_01981.pdf.

Farrell:2023:MEH

[4168] Henry Farrell and Marion Fourcade.

The moral economy of high-tech modernism. Dædalus, 152(1):225–235, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/152/1/225/
2072591/daed_a_01982.pdf.

danahboyd:2023:SWA

[4169] danah boyd. The structuring work of algorithms. Dædalus, 152(1):236–240, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/152/1/236/
2072556/daed_a_01983.pdf.

Janeway:2023:HTM

[4170] William H. Janeway. High-tech modernism: Limits & extensions. Dædalus, 152(1):241–244, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
152/1/241/2072554/daed_a_01984.
pdf.

Carugati:2023:GAR

[4171] Federica Carugati and Nathan Schneider. Governance archaeology: Research as ancestry. Dædalus, 152 (1):245–257, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
152/1/245/2072572/daed_a_01985.
pdf.

Tsai:2023:TRT

[4172] Lily L. Tsai. Taking responsibility for tomorrow: Remaking collective governance as political ancestors. Dædalus, 152(1):258–264, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
152/1/258/2072570/daed_a_01986.
pdf.

Pendleton-Jullian:2023:SOE

[4173] Ann Pendleton-Jullian and John Seely Brown. In search of ontologies of entanglement. Dædalus, 152(1):265–271, Winter 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL https://direct.mit.
edu/daed/article-pdf/152/1/265/
2072609/daed_a_01987.pdf.

Welsh:2023:P

[4174] Jennifer M. Welsh, Paul H. Wise, and Jaime Sepúlveda. Preface. Dædalus, 152(2):6–12, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
152/2/6/2115306/daed_e_01988.pdf.

Miliband:2023:I

[4175] David Miliband and Ken Sofer. Introduction. Dædalus, 152(2):13–21, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/152/2/
13/2115286/daed_e_01989.pdf.

Shesterinina:2023:ICC

[4176] Anastasia Shesterinina. Identifying contemporary civil wars’ effects on humanitarian needs, responses & outcomes. Dædalus, 152(2):24–37, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/152/2/
24/2115297/daed_a_01990.pdf.

Freedman:2023:HCG

[4177] Lawrence Freedman. Humanitarian challenges of great power conflict: Signs from Ukraine. Dædalus, 152(2):40–51, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
152/2/40/2115284/daed_a_01991.pdf.

Barbar:2023:CEH

[4178] Ana Elisa Barbar. Challenges for ethical humanitarian health responses in contemporary conflict settings. Dædalus, 152(2):53–62, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.

mit.edu/daed/article-pdf/152/2/
53/2115333/daed_a_01992.pdf.

Stanski:2023:HHR

[4179] Keith Stanski. Humanitarian health responses in urban conflict zones. Dædalus, 152(2):70–82, Spring 2023. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

https://direct.mit.edu/daed/article-
pdf/152/2/70/2115279/daed_a_01993.
pdf.

Aguayo:2023:GEH

[4180] Sergio Aguayo and Sandra Sepúlveda. The great evasion: Human mobility & organized crime in Mexico & its borders. Dædalus, 152(2):86–99, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/152/2/
86/2115310/daed_a_01994.pdf.

Sjoberg:2023:TSH

[4181] Ann-Kristin Sjöberg and Mehmet Balci. In their shoes: Health care in armed conflict from the perspective of a non-state armed actor. Dædalus, 152(2):103–124, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
152/2/103/2115291/daed_a_01995.
pdf.

Fast:2023:GDR

[4182] Larissa Fast. Governing data: Relationships, trust & ethics in leveraging data & technology in service of humanitarian health delivery. Dædalus, 152(2):125–140, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
152/2/125/2115327/daed_a_01996.
pdf.

Bagshaw:2023:TCS

[4183] Simon Bagshaw and Emily K. M. Scott.

Talk is cheap: Security council resolution 2286 & the protection of health care in armed conflict. Dædalus, 152 (2):142–156, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
152/2/142/2115304/daed_a_01997.
pdf.

Toukan:2023:LRG

[4184] Dima M. Toukan. Localizing responses to gender-based violence: The case of women-led community-based organizations in Jordan. Dædalus, 152 (2):167–178, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
152/2/167/2115317/daed_a_01998.
pdf.

Murdie:2023:LND

[4185] Amanda Murdie and Morgan Barney.

Localizing the NGO delivery of health from the outside in. Dædalus, 152 (2):181–196, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
152/2/181/2115335/daed_a_01999.
pdf.

Wise:2023:C

[4186] Paul H. Wise, Jennifer M. Welsh, and Jaime Sepúlveda. Conclusion. Dædalus, 152(2):210–215, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/152/2/210/
                                        2115277/daed_a_02000.pdf.

Aguayo:2023:GEM

[4187] Sergio Aguayo. La gran evasión: Migración y crimen organizado en México y sus fronteras. Dædalus, 152(2):i– xiv, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/152/2/
                                        i/2115329/daed_a_02013.pdf.

Biedarieva:2023:MW

[4188] Svitlana Biedarieva. The morphology of war I. Dædalus, 152(2):23, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/152/2/
                                        23/2115281/daed_a_02001.pdf.

Neville:2023:CBI

[4189] Mark Neville. Children from Bogdanovka inside their burnt-out school bus, Kyiv region. Dædalus, 152(2): 39, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/152/2/
                                        39/2115288/daed_a_02002.pdf.

Murray:2023:HWP

[4190] Nina Murray. how to write a poem about Bucha. Dædalus, 152(2):52, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/152/2/
52/2115295/daed_a_02003.pdf.

Shah:2023:BOW

[4191] Bina Shah. A bird with one wing. Dædalus, 152(2):63–69, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic).  URL https://direct.
mit.edu/daed/article-pdf/152/2/
63/2115302/daed_a_02004.pdf.

Nguyen:2023:EC

[4192] Viet Thanh Nguyen. Excerpt from the committed. Dædalus, 152(2):83– 85, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/152/2/
83/2115293/daed_a_02010.pdf.

Fouad:2023:AD

[4193] Fouad M. Fouad and Norbert Hirschhorn.

Aleppo diary. Dædalus, 152(2):100– 102, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

mit.edu/daed/article-pdf/152/2/
100/2115308/daed_a_02005.pdf.

Ndoro:2023:S

[4194] Tariro Ndoro. Swept away. Dædalus, 152(2):141, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

direct.mit.edu/daed/article-pdf/
152/2/141/2115315/daed_a_02006.
pdf.

Ackerman:2023:E

[4195] Elliot Ackerman. An evacuation.

Dædalus, 152(2):157–166, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/152/2/157/
                                        2115319/daed_a_02007.pdf.

kokothett:2023:SGE

[4196] ko ko thett. A survival guide for exiles. Dædalus, 152(2):179–180, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL https://direct.mit.
                                        edu/daed/article-pdf/152/2/179/
                                        2115331/daed_a_02008.pdf.

Hussaini:2023:D

[4197] Hajar Hussaini. Disbound. Dædalus, 152(2):197–209, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        152/2/197/2115275/daed_a_02009.
                                        pdf.

Maidukov:2023:I

[4198] Sergiy Maidukov. Izium. Dædalus, 152 (2):1, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://direct.

                                        mit.edu/daed/article-pdf/152/2/
                                        1/2115312/daed_a_02011.pdf.

Maidukov:2023:J

[4199] Sergiy Maidukov. June 10. Dædalus, 152(2):1, Spring 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL https://

                                        direct.mit.edu/daed/article-pdf/
                                        152/2/1/2115299/daed_a_02012.pdf.

Wolfram:2023:LSJ

[4200] Walt Wolfram, Anne H. Charity Hudley, and Guadalupe Valdés. Language & social justice in the United States: an introduction. Dædalus, 152(3):5–17, Summer 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

URL /daed/article-pdf/152/3/5/
2155838/daed_e_02014.pdf.

Curzan:2023:LSL

[4201] Anne Curzan, Robin M. Queen, Kristin VanEyk, and Rachel Elizabeth Weissler. Language standardization & linguistic subordination. Dædalus, 152(3):18–35, Summer 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

URL /daed/article-pdf/152/3/18/
2155855/daed_a_02015.pdf.

Wolfram:2023:ALI

[4202] Walt Wolfram. Addressing linguistic inequality in higher education: a proactive model. Dædalus, 152(3):36–51, Summer 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

URL /daed/article-pdf/152/3/36/
2155851/daed_a_02016.pdf.

Valdes:2023:SJC

[4203] Guadalupe Valdés. Social justice challenges of teaching languages. Dædalus, 152(3):52–68, Summer 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

URL /daed/article-pdf/152/3/52/
2155848/daed_a_02017.pdf.

Leonard:2023:REL

[4204] Wesley Y. Leonard. Refusing endangered languages narratives. Dædalus, 152(3):69–83, Summer 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

                                        URL /daed/article-pdf/152/3/69/
                                        2155832/daed_a_02018.pdf.

Fine:2023:CLE

[4205] Julia C. Fine, Jessica Love-Nichols, and Bernard C. Perley. Climate & language: an entangled crisis. Dædalus, 152(3):84–98, Summer 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

                                        URL /daed/article-pdf/152/3/84/
                                        2155844/daed_a_02019.pdf.

Rosa:2023:RLB

[4206] Jonathan Rosa and Nelson Flores. Rethinking language barriers & social justice from a raciolinguistic perspective. Dædalus, 152(3):99–114, Summer 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

                                        URL /daed/article-pdf/152/3/99/
                                        2155840/daed_a_02020.pdf.

Clemons:2023:BWR

[4207] Aris Moreno Clemons and Jessica A. Grieser. Black womanhood: Raciolinguistic intersections of gender, sexuality & social status in the aftermaths of colonization. Dædalus, 152(3):115–129, Summer 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL /daed/article-pdf/
                                        152/3/115/2155846/daed_a_02021.
                                        pdf.

Yoo:2023:AAR

[4208] Joyhanna Yoo, Cheryl Lee, Andrew Cheng, and Anusha Ànand. Asian American racialization & model minority logics in linguistics. Dædalus, 152 (3):130–146, Summer 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL /daed/

article-pdf/152/3/130/2155842/daed_
a_02022.pdf.

Alim:2023:IWV

[4209] H. Samy Alim. Inventing “the White Voice”: Racial capitalism, raciolinguistics & culturally sustaining pedagogies. Dædalus, 152(3):147–166, Summer 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
152/3/147/2155826/daed_a_02023.
pdf.

Baugh:2023:LPA

[4210] John Baugh. Linguistic profiling across international geopolitical landscapes. Dædalus, 152(3):167–177, Summer 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
152/3/167/2155836/daed_a_02024.
pdf.

King:2023:LT

[4211] Sharese King and John R. Rickford.

Language on trial. Dædalus, 152 (3):178–193, Summer 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL /daed/

article-pdf/152/3/178/2155824/daed_
a_02025.pdf.

Mendoza-Denton:2023:CIC

[4212] Norma Mendoza-Denton. Currents of innuendo converge on an American path to political hate. Dædalus, 152(3):194–211, Summer 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/

                                        152/3/194/2155828/daed_a_02026.
                                        pdf.

Hudley:2023:LL

[4213] Anne H. Charity Hudley. Liberatory linguistics. Dædalus, 152(3):212–226, Summer 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL /daed/article-

                                        pdf/152/3/212/2155834/daed_a_02027.
                                        pdf.

Kleinman:2023:P

[4214] Arthur Kleinman. Preface. Dædalus, 152(4):6–7, Fall 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL /daed/

                                        article-pdf/152/4/6/2345394/daed_
                                        e_02028.pdf.

Becker:2023:IHM

[4215] Anne E. Becker, Giuseppe Raviola, and Arthur Kleinman. Introduction: How mental health matters. Dædalus, 152(4):8–23, Fall 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

                                        URL /daed/article-pdf/152/4/8/
                                        2345398/daed_e_02029.pdf.

Sampson:2023:MPP

[4216] Laura Sampson, Laura D. Kubzansky, and Karestan C. Koenen. The missing piece: a population health perspective to address the U.S. mental health crisis. Dædalus, 152(4):24–44, Fall 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

                                        URL /daed/article-pdf/152/4/24/
                                        2345400/daed_a_02030.pdf.

Swanson:2023:AGV

[4217] Jeffrey W. Swanson and Mark L.

Rosenberg. American gun violence & mental illness: Reducing risk, restoring health, respecting rights & reviving communities. Dædalus, 152(4):45–74, Fall 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

URL /daed/article-pdf/152/4/45/
2345412/daed_a_02031.pdf.

Hansen:2023:RPS

[4218] Helena Hansen, Kevin J. Gutierrez, and Saudi Garcia. Rethinking psychiatry: Solutions for a sociogenic crisis. Dædalus, 152(4):75–91, Fall 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

URL /daed/article-pdf/152/4/75/
2345388/daed_a_02032.pdf.

Metzl:2023:PPF

[4219] Jonathan M. Metzl. The protest psychosis & the future of equity & diversity efforts in American psychiatry. Dædalus, 152(4):92–110, Fall 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

URL /daed/article-pdf/152/4/92/
2345403/daed_a_02033.pdf.

Belkin:2023:DTL

[4220] Gary Belkin. Democracy therapy:

Lessons from ThriveNYC. Dædalus, 152(4):111–129, Fall 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL /daed/

article-pdf/152/4/111/2345406/daed_
a_02034.pdf.

Gone:2023:IHT

[4221] Joseph P. Gone. Indigenous historical trauma: Alter-native explanations for mental health inequities. Dædalus, 152(4):130–150, Fall 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
152/4/130/2345392/daed_a_02035.
pdf.

Jamison:2023:DME

[4222] Kay Redfield Jamison. Disorders of mood: The experience of those who have them. Dædalus, 152(4):151–165, Fall 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
152/4/151/2345384/daed_a_02036.
pdf.

Harrington:2023:MHS

[4223] Anne Harrington. Mental health’s stalled (biological) revolution: Its origins, aftermath & future opportunities. Dædalus, 152(4):166–185, Fall 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
152/4/166/2345396/daed_a_02037.
pdf.

Hyman:2023:BMD

[4224] Steven E. Hyman. The biology of mental disorders: Progress at last. Dædalus, 152(4):186–211, Fall 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
152/4/186/2345410/daed_a_02038.
pdf.

Horwitz:2023:TSD

[4225] Allan V. Horwitz and Jerome C. Wakefield. Two sides of depression: Medical & social. Dædalus, 152(4):212–227, Fall 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
152/4/212/2345408/daed_a_02039.
pdf.

Galatzer-Levy:2023:CMH

[4226] Isaac R. Galatzer-Levy, Gabriel J.

Aranovich, and Thomas R. Insel. Can mental health care become more human by becoming more digital? Dædalus, 152(4):228–244, Fall 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
152/4/228/2345382/daed_a_02040.
pdf.

Patel:2023:EEO

[4227] Vikram Patel and Atif Rahman. Empowering the (extra)Ordinary. Dædalus, 152(4):245–261, Fall 2023. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL /

daed/article-pdf/152/4/245/2345386/
daed_a_02041.pdf.

Kleinman:2023:GMH

[4228] Arthur Kleinman and Caleb Gardner.

Good mental health care: What it is, what it is not & what it could be. Dædalus, 152(4):262–279, Fall 2023. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
152/4/262/2345390/daed_a_02042.
pdf.

Baltimore:2024:PRI

[4229] David Baltimore, David S. Tatel, and Anne-Marie Mazza. Preface: Recognizing implicit bias in the scientific & legal communities. Dædalus, 153(1):6–7, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

                                        URL /daed/article-pdf/153/1/6/
                                        2345809/daed_e_02043.pdf.

Liu:2024:IIB

[4230] Goodwin Liu and Camara Phyllis Jones. Introduction: Implicit bias in the context of structural racism. Dædalus, 153(1):8–14, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

                                        URL /daed/article-pdf/153/1/8/
                                        2345792/daed_a_02044.pdf.

Holder:2024:SU

[4231] Jr. Eric H. Holder. Seeing the unseen. Dædalus, 153(1):15–17, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

                                        URL /daed/article-pdf/153/1/15/
                                        2345784/daed_a_02045.pdf.

Nunez-Smith:2024:CDV

[4232] Marcella Nunez-Smith. The case for data visibility. Dædalus, 153(1):18–20, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL /daed/article-pdf/
                                        153/1/18/2345788/daed_a_02046.pdf.

Morehouse:2024:SIR

[4233] Kirsten N. Morehouse and Mahzarin R. Banaji. The science of implicit race bias: Evidence from the implicit association test. Dædalus, 153(1):21–50, Winter 2024.

CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

URL /daed/article-pdf/153/1/21/
2345819/daed_a_02047.pdf.

Ratliff:2024:IAT

[4234] Kate A. Ratliff and Colin Tucker Smith. The implicit association test. Dædalus, 153(1):51–64, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

URL /daed/article-pdf/153/1/51/
2345814/daed_a_02048.pdf.

Meltzoff:2024:YCI

[4235] Andrew N. Meltzoff and Walter S.

Gilliam. Young children & implicit racial biases. Dædalus, 153(1):65–83, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
153/1/65/2345811/daed_a_02049.pdf.

Kubota:2024:UIR

[4236] Jennifer T. Kubota. Uncovering implicit racial bias in the brain: The past, present & future. Dædalus, 153(1):84–105, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic).

URL /daed/article-pdf/153/1/84/
2345801/daed_a_02050.pdf.

Galvan:2024:IBC

[4237] Manuel J. Galvan and B. Keith Payne. Implicit bias as a cognitive manifestation of systemic racism. Dædalus, 153(1):106–122, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
153/1/106/2345780/daed_a_02051.
pdf.

Hetey:2024:WCL

[4238] Rebecca C. Hetey, MarYam G. Hamedani, Hazel Rose Markus, and Jennifer L. Eberhardt. ”When the cruiser lights come on”: Using the science of bias & culture to combat racial disparities in policing. Dædalus, 153(1):123–150, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL /daed/article-pdf/
                                        153/1/123/2345806/daed_a_02052.
                                        pdf.

Glaser:2024:DEI

[4239] Jack Glaser. Disrupting the effects of implicit bias: The case of discretion & policing. Dædalus, 153(1):151–173, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL /daed/article-pdf/
                                        153/1/151/2345796/daed_a_02053.
                                        pdf.

Greenwald:2024:RIB

[4240] Anthony G. Greenwald and Thomas Newkirk. Roles for implicit bias science in antidiscrimination law. Dædalus, 153(1):174–192, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL /daed/article-pdf/
                                        153/1/174/2345776/daed_a_02054.
                                        pdf.

Kang:2024:LTM

[4241] Jerry Kang. Little things matter a lot: The significance of implicit bias, practically & legally. Dædalus, 153(1):193–212, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
153/1/193/2345782/daed_a_02055.
pdf.

Kalev:2024:RCS

[4242] Alexandra Kalev and Frank Dobbin.

Retooling career systems to fight workplace bias: Evidence from U.S. corporations. Dædalus, 153(1):213–230, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
153/1/213/2345786/daed_a_02056.
pdf.

Sigur:2024:IBV

[4243] Wanda A. Sigur and Nicholas M.

Donofrio. Implicit bias versus intentional belief: When morally elevated leadership drives transformational change. Dædalus, 153(1):231–249, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
153/1/231/2345790/daed_a_02057.
pdf.

Xiang:2024:MMW

[4244] Alice Xiang. Mirror, mirror, on the wall, who’s the fairest of them all? Dædalus, 153(1):250–267, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

tronic). URL /daed/article-pdf/
153/1/250/2345794/daed_a_02058.
pdf.

Walker:2024:DIB

[4245] Darren Walker. Deprogramming implicit bias: The case for public interest technology. Dædalus, 153(1):268–275, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL /daed/article-pdf/
                                        153/1/268/2345799/daed_a_02059.
                                        pdf.

Albright:2024:BIB

[4246] Thomas D. Albright, William A. Darity, Diana Dunn, Rayid Ghani, Deena Hayes-Greene, Tanya Katerí Hernández, and Sheryl Heron. Beyond implicit bias. Dædalus, 153(1):276–283, Winter 2024. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (elec-

                                        tronic). URL /daed/article-pdf/
                                        153/1/276/2345804/daed_a_02060.
                                        pdf.

Holton:1973:TOS

[4247] Gerald James Holton. Thematic origins of scientific thought: Kepler to Einstein. Harvard University Press, Cambridge, MA, USA, 1973. ISBN 0-674-87745-4. 495 pp. LCCN Q125 .H722.